A Complete Bibliography of Publications in Ecological Modelling (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

29 August 2024 Version 1.07

Title word cross-reference

1 [BGMP06, BP04, TSJ08, UZ01]. $17.99 [Jør09b]. 2 [MMB07, Mur01, NSO^{+}08, UZ01, dlFN08]. 3 [AYK07, BW01b, DWD07, GBS00, GNA^{+}06, HY07a, HY07b, JCB^{+}02, KMN^{+}07, MBGP08, NBT^{+}09, SLGS00, TSF^{+}05, WMM^{+}07, XRM08]. $64.95 [Log02]. $69.95 [Jor05i]. $73 [Jør04d]. $79.95 [Nie06]. ^{1} [WHP01]. 137 [DBBS03]. _{10} [DVJ^{+}08, ESZ^{+}00]. _{2} [AW06b, ARL^{+}06, AS09, BW01a, BMT07, CGY08, CGdPRY09, GAA^{+}05, GZY^{+}06, GTJ^{+}00, Gut07, HB06, HHB^{+}08, KBE^{+}06, LIS07, MOP^{+}05, MOP^{+}06, MMB07, MBF09, MWD05, NJF^{+}08, RFA^{+}01, RNKG03, RGF00, RMDC04, dSLSD02, Tan02, VA00, YY06, ZXC03, ZGF^{+}05, vODFS04]. _{2.5} [DVJ^{+}08]. _{3} [WHP03, vODFS04]. _{3}^{−} [VL08]. _{4} [WHP03]. _{F} [WFB^{+}08]. a [ÇDKK05, WR01b]. : [TFM01]. F [CM05a, Nad07]. k [WW09, WDBK08]. Q_{10} [Xia00c]. R [WGS^{+}02].

-D [UZ01, AYK07, BGMP06, BW01b, GNA^{+}06, JCB^{+}02, KMN^{+}07, UZ01].

-fixing [HB06]. -fold [WDBK08]. -layer [WFB^{+}08]. -means [WW09]. -PG [XRM08].

0 [Jor05a, Jor05b, Jor05c, Jor05e, Jor05g, Jor05i, Jor05j, Jor05k, Jor05l, Nie06]. 0-08-044015-0 [Xu03]. 0-08-044028-2 [Xu03]. 0-12-088459-3 [Nie06]. 0-19-926124-5 [Jor05e]. 0-387-00835-7 [Jor05i]. 0-444-50240-8 [Hey01a]. 0-470-86671-3 [Jor05k]. 0-470-87093-1 [Jor05a]. 0-471-08498-0 [Jor05g]. 0-471-49618-9 [Jor05b]. 0-521-82230-0 [Jor05l]. 0-85199-563-2 [Jor05c]. 0-87893-096-5 [Log02].

1 [BPE^{+}07, FSE^{+}04a, Jor05a, Jor05h, Zha06]. 1-84564-059-4 [Zha06]. 1-85312-954-2 [Jor05h]. 1-day [BCCR^{+}02]. 10 [Pet04a]. 119 [CPH00]. 120 [BGF00]. 122 [MLGV00a]. 123 [JT01, RLHD01]. 137 [BG01a]. 137/2 [BG01a]. 140 [Str01a]. 140/3 [Str01a]. 144 [Jag09, Met02]. 153 [For03, HDBM02a]. 157 [LOL03a]. 158 [Ano03-27]. 170 [Ano04v, CSU04]. 171 [Swa06]. 188 [LH06]. 189 [Suh06]. 191 [CG07]. 193 [Tia06a]. 195 [CM07a]. 1950s [Kan04a]. 1970s [HSJ04]. 1973 [Zuc04]. 1980s [HSJ04, ZC07]. 1990s [HSJ04, ZC07]. 1998 [Nie00a]. 1d [PWC^{+}09, BP04].

2 [FSE^{+}04b, Jor05c, Jor05f, Jor05h, KCZ^{+}03, PBE^{+}07]. 2002 [Jør04g, ZCY09]. 2003 [Ano06-47]. 2004 [Ano06-45, Ano06-46, LMPR06]. 2006 [LBL^{+}08, Ano08a]. 2008 [Ano09-33]. 2009 [Ano09c, Ano09d, Ano09e, Ano09f, Ano09j, Ano09k, Ano09l]. 201 [KL09, San09]. 212 [Dun08]. 215 [BLB09, Ned09]. 21st [LMPR06]. 220 [Fat09b, MG09]. 23^{o} [PG08]. 25 [Bar04]. 29.95 [Nie08a]. 292 [Ned09]. 2nd [Log02].

3 [CWL05, FSE^{+}04c, Jor05f, Jor05j, Jor05k, Nie06, WCS00]. 3-540-00153-0 [Jor05j]. 3-540-43373-2 [Jor05f]. 3-box [IMK^{+}07]. 3-dimensional [RWM^{+}07]. 3-PG [CWL05, WCS00]. 316 [KL09]. 3rd [Xu03].

4 [Zha06]. 4-nonylphenol [MLTN06]. 4th [Ano07a].

5 [Jor05e]. 5-model [ZMB^{+}08].

609 [Hey01a]. 64 [JAN^{+}03].

7 [Jor05i].

89 [RLHD01].

9 [Jor05b]. 921 [LLCL04].

=pkgDivGame [AS00a].

Îles [GCG^{+}07]. Îles-de-la-Madeleine [GCG^{+}07].

aapa [LRJ^{+}09]. Aare [MSP^{+}08]. Aat [Jor05l]. abalone [BDLL06]. abandoned [BCMR03, RSB09]. abdominal [LGL02]. Abies [BAP^{+}06, Oga09b, SK04, WBB05]. abilities [AP06]. Ability [SBJ^{+}02, FH08, HLH^{+}06, PR01, Zha03b]. Abiotic [LA07, Lud04, Mon09b]. Above [Urs09, DAC^{+}09, WW08, XSY^{+}09, ZBSA07, vNM02]. above- [vNM02]. above-ground [DAC^{+}09, ZBSA07]. abscission [MBK^{+}03]. absence [CWS09, Den08, Ned09, VM09, VSGW09]. absences [CL08a]. absorption [Che06, DdSG01]. abstraction [BWJZ01]. abundance [BA08, BP08, CSHY08, EG09, EPQdCA06, HS01, HMGK05, KID^{+}07, LGS03, LGD01, OW05, PEPB09, PBS05, PE06, SHN09, Sin07, WZKL09, WR01c, YHC04]. abundances [YPR^{+}05]. Acacia [KBM^{+}03]. Academic [Jør04c, Jør04d, Jør04f, Nie06]. Acari [CM07a, CM06a]. acarine [BG03]. accelerate [RB02]. accelerated [RP09]. accelerometer [GPC^{+}09]. acceptable [AHKB01]. access [MWWZ01]. accessibility [Mat03b]. acclimation [Gut07]. account [BKMB08, CQ07, MSB07]. Accounting [BF07, MF06, BM00, BKC^{+}07, CCY06, CC06, GW04, TB06b, VFTB06]. accumulation [DBBS03, KK07, LB01, Roe00]. accuracies [Kir06a]. accuracy [Cou03, ECZ^{+}06, EN08, GG00, MXC^{+}04, SP02, TSBH09, WO01b]. accurate [OJD04]. Acer [Lar02]. Achieving [AB06, BBGM06, CKBH00]. acid [BPC04, GVC09, TBP^{+}07, ZAM^{+}07]. acidic [Ito07]. Acidification [NKK^{+}07, HRJ^{+}00, OHH07]. Acidithiobacillus [AMSM06]. acinaciformis [SMTR07]. Acipenser [WX09]. Acknowledging [LRJ^{+}09]. Acknowlodgement [Ano05-55]. acquisition [MLPK01]. Acropora [Lir03]. Acroptilon [GPB01]. across [BW01a, HSC^{+}04, HTA^{+}08, KRZ07, KGA06, LGR^{+}09, Mal03, PCC^{+}07, TTJ07, YSB08, ZMB^{+}08, ZRCA08]. action [GPP02, Ort04]. active [BRK07, CUF^{+}09, LR07b]. activities [Hua03, MB05a, TN06]. activity [Ale07, JLF08, MH02, MBGP08, MNZC08]. actual [VM09]. acutus [RMD04]. acyclic [SAB^{+}06]. Ad [Ano01a]. Adams [BSB^{+}09]. Adaptation [RKH^{+}07, Hul01, Hul02, Hul04, JF04, PDBJ09]. Adaptations [KFJ^{+}09, KS04b]. Adapting [GFG09]. Adaptive [LM07b, VB06, CC00, JWLA00, KBF^{+}08, LKR06, LEH06, Mac00, MB07, Pra00, Pra05a, Rai01, Rec03, Tsc04, UDB07, YK00]. Adding [WC07, Bor06]. additions [CC05]. Additive [Ric05, BW02, FAA^{+}02, GEH02, Hee02, LEH06, MFB^{+}06, YM02, vNM02]. adehabitat [Cal06]. adequate [SBDD04a]. Adirondack [MMR06]. adjacent [JM04, LXP^{+}08, MRG09]. adjoint [ZWXF05]. adjusted [MDVG09]. Adour [CGC01]. Adour-Garonne [CGC01]. Adriatic [KL09, Leg03b, MNZ06, San09, JAK^{+}06, KL07, Leg08, Mal01]. Adsorption [WTMG09]. adults [JJWF07, Shi04a]. advanced [KWS^{+}07]. Advances [Ano06-46, Foh05, Sca01, Kri04b, MEO06]. advantage [MB07]. Advantages [Uus07]. advective [KMN^{+}07]. Advisory [Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano07b, Ano07c, Ano07d]. Aegean [MSM^{+}07]. AEI [ZC00]. aerosol [ESZ^{+}00]. aerosols [Alo04]. aeruginosa [BP02, JKWJ03, KCJ^{+}07]. aestivum [MWD05]. affect [MTKM^{+}06, SHZL09, TCGL03]. affected [ABP05, BGF00, BRGS09, DBD^{+}08, MH02, Met03, PLTT05, SYC04, ZPK^{+}07]. affecting [DDLD07, Ond07, PPP05, SF07]. affects [EWH^{+}02, JKJ06, LYC08, VF07]. affine [SMTR07]. affinis [DSD^{+}09]. afforestation [MGCK^{+}03, PPR03]. Africa [ESZ^{+}00, HRH^{+}05, LS09a, RvGC^{+}08, SA07b]. African [CG06c, GGV^{+}09, PMLM08]. after [AHKB01, Bir06, EPS04, GPDF09, Kai00, MWK07, MMF^{+}09b, MMF^{+}09a, MBD^{+}00, MY02b, OMHR06, Pen09, PBH^{+}07, RF09, SSH^{+}07, TMM05]. aftermath [Kri04b]. again [SLPP05]. against [BBD^{+}04, CCG07, CBSLS07, IWM^{+}06, LGR^{+}09, Sel00, TMHJ06]. Age [AALM05, JM01, Leb05, Aub04a, Aub04b, BLHB06, BMD09, CC00, CDAGK06, CK07b, DFF07, EB06, FLS06, HA08a, Jen00, KT03, MBM06, MM00, SMB^{+}06, SMR08, SB00, SW03, SZ03, Tar08, TW00, Xia02]. age- [Xia02]. age-dependent [SB00]. age-specific [CK07b, HA08a]. Age-structured [AALM05, Aub04b, CDAGK06, KT03, Tar08]. ageing [BH00, JJWF07]. Agelaius [ÖTÖR06]. Agent [BPW^{+}03, MBLA03, ADS^{+}07, BL04, DH01, GNA^{+}06, GBB^{+}06, JWLA00, KWW^{+}09, LPFL09, LPM^{+}09, MDGV09, RvGC^{+}08, THJ^{+}03, YMD08]. Agent-based [MBLA03, GBB^{+}06, LPM^{+}09, THJ^{+}03, YMD08]. agents [GLS02, LM07a, Rec03]. aggradation [UJF06]. aggregated [TH06b, WR03, ZH06, vHPS02]. aggregates [GPL05]. aggregating [Jua09]. Aggregation [ACVP00, KPAK02, PF01, PBM^{+}05, AAU02, ABS05, AA05, BFU^{+}09, EB07b, JNA^{+}09, JO09, Ken02, MAdlPR02, MG01, Sil04, SF04]. aggregative [MAL06]. aggressive [CDAGK06]. aggressiveness [CMSB07]. Ago [KCY^{+}08]. Agricultural [MG02, AFT09, BDR01, BvdW04, BRGS09, BH06, DFGC04, FRZ00, GGV^{+}06, Gri08, JM04, KvKV05, LLA^{+}09a, mLMfT05, MHvIR00, MMHE06, MBPS04, PEM06, PBMRE08, RS04, Sav00, SB03, SFM08, WHX^{+}03, ZSKV05, vdPLG^{+}00]. agriculture [BBC03, CJY^{+}09, LGS03, SCP05, ZLZM02, Jør04d]. agro [PPP05, SS00]. agro-ecosystem [PPP05]. agro-ecosystems [SS00]. agroclimatic [Kir06b, Kir06b]. agroecosystem [Sep00, dG02a, dG02b]. agroecosystems [CD05, EBH^{+}01, FCC^{+}00, Kir06a]. agroecozones [LS02b]. agroforestry [MGCK^{+}03]. ahead [BCCR^{+}02]. AHP [YZC^{+}07]. aided [Bye00, ESG06]. Aiming [GJY07]. Aims [Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-37, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27]. Air [Cor05, GPSM08, CZG05, DGOZ04, GPV08, GFGZ08, KMRV07, MLA^{+}02, RGO^{+}06, SGH^{+}08, TYT^{+}09, WHP01]. air-mediated [KMRV07]. air-to-soil [GFGZ08]. airborne [DVJ^{+}08, YSG^{+}06]. al [AB06, BBGM06, Dun08, JT01, RLHD01]. Alan [Jor05e]. Alaska [CNG06, Pot04]. alba [BAP^{+}06]. albedo [Wan05]. Alberta [VSHC08, YTH03]. ALBIOC [Roe01]. Alborán [Per07]. alder [LPD08, MB05a]. Alfred [Jør04d]. algae [HMBG03, MWM02, RT08, SCB^{+}09, SGLRE05]. algal [BR01, CM06b, HJ02, KFS06, LHDJ03, LLL^{+}07, ML05, OAL^{+}07, WR01c, YSM00]. algocoenoses [Lev00]. algorithm [ACJT08, BHFMG05, BPJM00, CA04, CC05, GL01b, HTS^{+}07, HSMN08, IC02, KLPP07, LC07a, MB00, OHM^{+}06, PCWP06, Sel00, SBS^{+}06, SBC07]. algorithms [BR01, CPP00, DGD03, DGD06, DLC07, Haw00, KCJ^{+}07, Lat06, LHZ^{+}06, Par00, PFR09, PF00b, TMP06, WR01b]. alien [KBM^{+}03]. Alignment [ATA03]. Alison [Jor05l]. All-scale [BL02]. Allee [ADdSC06, Bra01b, CL08b, EWH^{+}02, JBR07, KDS03, LL02b, POF08]. Allee-like [CL08b]. allelochemicals [ALJL03, Mar06]. allelopathic [DH00]. Allelopathy [FCP^{+}07, ALJL03, Bla07b, GPB01, MB06b, MVV07, SGLRE05]. alliances [MF02a]. Allocation [BMR07, BPJM00, IWW08, Kai00, Kin05, RVWH06, YM05, ZPD06, DP03, DP07]. allometric [BPW00, Hen07]. allometrically [Kin05]. Allowing [YMD08]. Alluvial [ZBW05, DS02]. ALMaSS [THJ^{+}03]. ALMIS [Esc05]. Alnus [LPD08]. ALOMYSYS [CCGMJ07]. along [AEP^{+}04, ASP^{+}07, Dam03, Dam08, HHL08, KR04, LKLK07, MBM00, MSA^{+}03, OM02, PM06a, SK04, Sin07, WHHH07, ZvBS05]. Alopecurus [CS01]. Alpine [VZG05, YY06, AEK^{+}07, FE04, JU01, RPRV09, SMG07, SMSR00, TTA^{+}01, WTST08, WTS^{+}06, XZS^{+}07, YYZ06]. Alps [BLDCM06, Hör03, RBEZ08]. Altar [SF07]. alter [Fie04]. alteration [OFK08]. Alternative [FWHL06, MMLR07, vNS03, BG01a, BG01b, DFF07, ECL^{+}02, FK03, HNS08, LGS^{+}00, MWM02, MFJM04, MG02, OW02, PF00b, PFFM07]. alternatives [LHK00]. Amazon [DSA08, EMdC^{+}01, PdAD^{+}04, Tol06]. Amazonia [ACE07, PR01]. Amazonian [SFCP02, ZXC03]. ambient [dSLSD02]. Ambio [Zuc04]. America [KW02b, LN08, LGR^{+}09, MKB00, Wel04, YJJG09]. American [HL04, MWP07, RMD04, ZC07]. americana [BB08b]. Americans [BS08b]. americanus [FPS03, ZC07]. AmeriFlux [KBE^{+}06]. Amistad [DCG01]. ammonia [RMSS02]. ammonium [KM00]. AMOEBA [WDW00]. among [BJK09, FYN^{+}07, GL09, Gan06, GBSP08, MNOS08, MOJ01, Nal00, OzDBS07, SGLRE05, Tsc02, Uch00, Xia02, Zha03b]. amphibian [SGF09, WSP08]. amphipod [SKJvdHG06]. Amphipoda [MMM02]. Amsterdam [Hey01a, Hey01b, Nie00a]. Amudarya [RSM05]. Anabaena [MSL01]. anaerobic [NWP06]. ANAFORE [DVdB^{+}08]. analogs [Kan04b]. analyse [CDM05, IVC^{+}08, Mun09]. analysed [Cal05]. Analyses [ZYL06, CBD^{+}09, CMCD04, FJ05, GGB^{+}06, Gre06, Håk00, HMF00, LBL^{+}08, NCM07, San07]. Analysing [BBGH07, BGWC07, DFM07, JWLA00, MM00]. Analysis [Ano09-33, Bar00, BLDCM06, CMB05, FCH04, Il’08, Jen02, Jør04h, Jor05a, Jor05l, KWW^{+}03, KWD^{+}04, Log02, MR06, NDM00, SS06a, SLS03, Sch03, SBC07, TYT^{+}09, Tis06, TMM05, VREA06, WFB^{+}08, ZSPV08, ÅSCP09, AAU02, AF09, AWL04, ASG^{+}05, ACE07, APG^{+}03, ABV^{+}06, Ano04-96, AN00, AB05a, AB05b, ASS^{+}06, BFU^{+}09, Ban05, BCM^{+}08, BM00, BMG08, Ber02, BH01, BSM08, BL02, BBGH07, BWPC06, BSR04, BFOS08, BHAV03, BGL01, Bru05, BB03b, CDM05, Cal06, CSR08, CGLS07, CW01, CGC01, CG09, CH06, CLTH08, CSdZ09, CMPO05a, CS08, CGG^{+}05, DC08a, DUASCM07, DFF07, DLPN06, DBBS03, DDH^{+}09, DDF^{+}05, EG09, FMM^{+}07, Fat04c, FSUH07, FWF07, FWHL06, FC05b, FM07, Fis09, FPS03, GYY00, GRW04, GSB^{+}06b, GKT07, GWHS03, GDL06b, GH07, GLS07, HB04, HL04, Hen07]. analysis [HSRD09, Her04c, HB00a, HB00b, HUB02, Hin09, ICGÁ05, IVP08a, IVP08b, JKS^{+}05, JJWS08, JNA^{+}09, JLBS09, Kaz07, KSG05, Kri03, KIL^{+}03, Kri04a, KGJ08, KVL^{+}09, KKH^{+}08, LD06a, LBBC08, Las06b, Lat06, LCF09, LKY^{+}06, LOL02, LOL03a, LOL03b, Len07, Li00a, LCY09, LYCU09, LW09, LCH^{+}00, MB08, MC08, MKS^{+}09, MEJ06, MRK^{+}07, MDB^{+}02, MRE^{+}06, Mon05, MPS02, MMPN07, MS03, MCK07, ML05, NAC04, OI07, Odu04, Oga09b, Oli03a, OBR01, PLL04, PRL07, PE08, PT02, PLA06, PBSOMG^{+}05, PPR03, PPE^{+}07, dSPdBB08a, PBCZ01, PPS08, PWC^{+}09, PRRB09, PSVH09, PHBF07, PTGI09, Pra00, PAF06, PBML08, PMJ08, PKG00, RST05, RBSG06, Ray08, RKH05, RGO^{+}06, RFdV07, SA07a, SPTP01, SM04a, SF09, SHB06, SFM08, SGP^{+}06, SGP^{+}07, SFC05, SRB06, SCH05, SKP09, SK04, SL06, SLE07]. analysis [SV03b, SYCU09, Suh05, Suh06, Svi04, TPJ08, TTA^{+}03, TKSP09, TBPF08, TKK07, VSG^{+}08, VLA^{+}06, WKZP04, WIH^{+}09, WBP^{+}07, Whi05, WGT05, WT04b, WH09, XRM08, Xia02, Xia07, XG09, YFH03, Yea04, ZSD^{+}08, ZKH09, ZGH05, ZZZ06, ZZLC06, ZYY09a, ZPD^{+}08, ZHW05, ZLW09, vNM02, vdBDR02, GBB^{+}09]. analytic [LSY^{+}09, Sti06]. Analytical [TSF07, ZPO09, Kir06a, Kir06b, MTVC05, MHN00, ZXC03]. analytically [KA01]. analyze [DC08b, ECHN08, MBM06, vNLS02]. analyzed [LAPAMD07]. Analyzing [CK07a, GL09, GL08, dFOV07, HDH00, MGSdG07, MWMN06, SHG^{+}08, SRN05, STH06, TN06]. anchovy [KL07, KL09, MSM^{+}07, San09, SFM04]. and/or [BA05]. Andaman [RPC^{+}05]. Anderson [Jør04e]. Andes [KBK07]. Andrea [Jor05a]. anemochorous [SW01, WWRZ04]. Angra [MTVC05]. angustifolia [VCM07]. Aniansson [Jør04f]. Animal [VM06, Ben04, BNTK04, CM05b, dSDSM08, HAA08, HBDA09, HGR08, JBB^{+}05, MAL07, PLP^{+}04, Sti08, iWLSN00]. animals [Cal06, Eza05, HGD05, Mat03b, Nal01, OSHO09, PKS08, SHG^{+}08, THJ^{+}03]. anion [Met03]. anisopliae [SGHM01]. Anisotropy [WNW09, VMG05]. ANN [DGGD04, VLD02]. ANNA [WRCB01]. Anne [Jor05e]. Anne-Marie [Jor05e]. annealing [RMA08]. Annnouncement [Ano09j]. Announcement [Ano09c, Ano09d, Ano09e, Ano09f, Ano09a, Ano09b, Ano09k, Ano09l, Ano09-31, Ano09-32, Ano01m]. announcement/call [Ano01m]. ANNs [RMDC04]. Annual [CDM^{+}00, MGH^{+}05, MRS05, MB05b, WV05, WVP08, WKL^{+}03, WCH08, WH07, ZSPV08]. annuals [SSNHP08]. anomalies [Las06b]. anopheline [SBC07]. Anostraca [MVZM05]. ANOVA [GGB^{+}06]. ANOVA-based [GGB^{+}06]. anoxic [CM09]. Answer [Bio03, GSB06a]. ant [BLDCM06, LH05, LH06, SBB06, WH07]. ant-dispersed [LH05, LH06, WH07]. antagonistic [KVPA07]. antagonists [GTRF01]. Antarctic [CDA08, RC08]. Antarctica [MKS^{+}02, MMRLP06, OW05, OW08]. anthropogenic [BdB05, BFS05, HOK^{+}09, LRT^{+}08, MH02, MFG^{+}08, TN06]. Anticosti [PLP^{+}04]. antifrost [CT07a]. antillarum [MHZ^{+}06]. Antofagasta [PG08]. ants [Rod07]. Apalachicola [WHH^{+}08]. Aphid [PEM06, BDBS08, FPDP06, GTRF01, MB08, MKvdW^{+}09, PEPB09]. aphids [MKM^{+}07, MKM^{+}08]. Aphis [dSMZ09]. Apis [AGM^{+}08, WS02]. Apollo [BF07]. Appalachian [RG05]. Apparent [CGR03, DKRŠS00, XZS^{+}07]. appearance [XWB04]. apple [GL08]. Applicability [Deb02, BPC04, GVC09]. applicable [WHP01]. Application [AGZ05, AQS^{+}07, ATDK08, BSB00, BP04, BG08, CMM02, CMM^{+}07, CMCD04, DWH06, DSD^{+}09, GGB^{+}06, GLS07, HBO07, Ito07, JDTD06, Jor05c, KLL^{+}07, KLPP07, KS07, LLA^{+}09a, MNZ06, MP07, MKM^{+}08, MBGP08, MCE^{+}07, NMJ07, ORS^{+}09, OEK^{+}06, RŽČ^{+}04, SPK^{+}09, SL01, SHG^{+}08, TPCS06, YKO05, ZCB08, Zue07, dCBL09, AOY02, BWAM09, BW01b, BMR06, BL01, Buz08, CT07a, CMB^{+}02, CTH^{+}00, DLC07, DBBS03, Fle01, GDG^{+}00, GBEB06, GORV06, GACO04, GM05, GTJ^{+}00, HTS^{+}07, HFSH06, JG08, Jør02b, KA01, LTP06, LPC05, LL03b, MVPB02, MKM^{+}07, MR00, MH05, MG09, MHP^{+}06, NMC^{+}06, NS08, NF04, NSEDP06, OW05, OZL07, PLL04, fPzWhSY07, Pet04b, PCM^{+}03, RR05a, RS04, Sal06, SHB04, SM04b, SDS^{+}07, SA07b, SFC05, SK04, TP06, TNO^{+}09a, TYLL07, WFB^{+}08, WCH08, WBR08, WT04b]. application [WLZ^{+}09, YLSH03, ZH06, ZJC^{+}07, ZCG^{+}08, ZLL05, ZAM^{+}03b, RLLB09]. Applications [ARF08, Dže01, DŽD06, GICB09, PCCL03, Rec01a, ATAK00, BB08b, BKPS08, BBKN03, BW04, CM06a, CM07a, Fat04a, Gau06, Li00a, Mon09b, ÖTÖR06, ÖTÖ06, PG04, PPS08, RMD04, VLD02, WAB^{+}07, WA02, WD02, Xia03, Xia05b]. Applied [Jør04h, CT07a, Fat04c, FBDM09, GdBD08, Jen00, JLBS09, KK04, LKY^{+}06, NWP06, Wil08, dSMZ09]. Applying [BBC03, CSSCC07, CCM02, Com04, CP09, DRDD01, Saa00, WGT05]. Approach [Jør09b, MSC09, ATA03, AM02, AMB07, ASN02, BHC04a, BH00, BB08b, BSG07, BRC04, BPW00, CLM^{+}09, CBMP07, CD07, CDMM08, CFCP04, CMA^{+}06, CNG06, DDLB02, DD00, DM03, DLK01, DFM07, EG09, ES06, ES01, Eza05, FLS09, Gan06, Gas05, GMPC08, GdBD08, Gri08, HB01, HAG^{+}00, IOIA08, JKWJ03, JNM^{+}06, Jor00a, JcLM09, KMT09, KPH02, KZH07, KE07a, KFH07, KSvOO09, LBBN06, LLLP00, LSY^{+}09, LCSZ09, LNW00, LS02b, LCM^{+}09a, LPCC05, LJ09, Lud09, MD06a, MBMP06, Man00, MGCK^{+}03, MCKNM09, MKS^{+}09, MEJ06, MBW09b, MMLR07, Mon09a, Mon09b, MWP07, NK06, NW07, NK04, OBE^{+}07, OIR^{+}08, OJ02, OS02, ÖÖ04, PK02, PBA06, PSVH09, Pra05b, RDS07, RM09, Ric05, RV05, RLLB09, RMWW07, Sak09, SIK07, SAL07, SKP09, Sta07, Sti06, SJ03, SZL^{+}04, SKM^{+}06, SS01, SGF09]. approach [TB05, TM05a, TH06b, TP08, THA02, TS09, Tsc04, VGK^{+}04, VFTB06, VCM07, VMR09, WLG07, WZKL09, WZ01, WL06, WHW^{+}05, WDW00, WKZ03, WR01a, WLL^{+}08, WD02, Xia00b, YLSH03, YY06, YZS^{+}04, ZBD^{+}09, Zav08, ZKH09, ZXC03, ZVK05, ZWXF05, ZSPV08, ZLW09, vW07, Jør04e]. approaches [AWL04, Aus07, BBD^{+}04, CM06b, FK03, GRBT08, JBT^{+}05, JLAM09, Liu01, MSB07, NU00, RSC09, SV03b, TD06, TK07, TKSP09, URB06, Weg00, vdBDR02]. appropriate [DBB^{+}08, Mit09, PLL04]. approximation [GCM^{+}05, MKRH03, PGW00, WV05, WVP08]. April [Ano00-33, Ano00-43, Ano01-36, Ano02-48, Ano02-33, Ano03-44, Ano03-55, Ano04-54, Ano04-64, Ano05-29, Ano05-36, Ano06-38, Ano07s, Ano08r, Ano08j, Ano09-30, Ano09n, Ano09-33]. Aquaculture [MMPT06, BFHR05, Dow05, GCG^{+}07, JKPL09, LBS08b]. aquaculture-derived [JKPL09]. aquacultures [JGL07]. aquatic [AW00, AMB07, Aok08, ARF08, ATDK06, GSG^{+}04, GW08, HMBG03, JLH01, JVL02, Jør07b, KE07a, KBvVK08, MRRJ06, MBDB09, MPS02, MMPN07, NG09, OWWS01, OPL^{+}09, PCCL03, PCW08, PEAS01, RBSJ01, Rec03, RKS^{+}07, Str01a, Str01b, TN09, VCD05, ZK08, ZA09]. AQUATOX [PCW08]. aqueous [KK07]. aqueous-phase [KK07]. aquifer [IF07, SHZ05]. aquifers [IVP08b, LLH^{+}06]. Arabian [GGS08]. arabicus [SF07]. arable [HKB02, KH02]. araneus [WG07b]. arborescent [CDM^{+}00]. Arcachon [BSB^{+}09]. Archipelago [MFG^{+}06, MFG^{+}08]. architectural [DWD07, EVF^{+}07, PS05, Pet04a, RKH05]. Architecture [DP03, DP07, GBS00, GLS02, Han02, PBB04, Vil01, WMM^{+}07]. Arctic [BRK07, FKIL08, TFF07]. Arctocephalus [BDP^{+}02]. Arctophila [RAH07]. Area [LE04, ZSZ06, CGS08, CV07, CCG07, CMB^{+}02, CS07, CFS09, CSF^{+}04, CDA08, DBD^{+}08, DD02, Eti04, GBN^{+}06, Gri08, ICSC05, JM04, KFB09, Kri03, LSY^{+}09, LLCL04, Mat03a, MFB^{+}06, OAAF07, OIP^{+}08, PG08, TTA^{+}01, TAL^{+}05, WLJ00, YTS03, ZC00]. area-wide [Gri08]. areal [Szi00]. areas [Ayd08, CM06a, CM07a, DDFP07, HGB04, HVVK09, KMRV07, Las06a, MEKL08, MEJ06, MSD06, PFR09, PBPZ04, PM06a, POF08, PCP07, VTB^{+}08, WSY^{+}07]. arena [WC07]. arenaria [RC06]. arenas [WC07]. Argentina [CGH^{+}05, AGM^{+}08, CGG^{+}05, GCM^{+}05, SCP05]. Argopecten [WTMY07]. argus [JLAM09]. arid [ES06, HRH^{+}05, KKH^{+}08, KKH^{+}09, PSBJ07, RHH05, SvL04, TM05b, TCGL03, THA02, Urs09, WGB^{+}08, WSWL08, ZPW05]. ARIMA [BMBOCR03]. arion [GS02]. Arizona [FCC^{+}04]. Arm [RBW09, PM06a]. armigera [SvdWB^{+}06]. Arnulf [Jør04h]. aromatic [WLB^{+}05]. arrotAge [LBS^{+}06]. Art [Cam04, MLHT09]. arthropod [LWJ06]. arthropods [TS09]. Artificial [MH05, NK06, SBMJ09, AMSM06, ABM^{+}06, BPBF^{+}00, BGF00, BDP^{+}02, CKP^{+}01, DLG06, FH08, GDL03, HO01, HFSH06, JJK^{+}01, KPKP06, KZH07, KST^{+}07, KKCC06, KHLS07, LKLK07, LMM^{+}07, LMG08, MZWM05, MSTK08, NT07, NSEDP06, OWWS01, OAL^{+}07, OJ02, OJD04, OHM^{+}06, ÖTÖR06, ÖTÖ06, PKC^{+}01, PCCL03, PRL07, RMA08, RMDC04, SRW05, TB06a, TPJ08, VLD02, WR01c, DGGD04]. artificially [TMLV07]. artisanal [ASHHRRPE04, TN08]. as-built [KM04]. ascendency [JU09, LS04b]. Ascoglossa [CTH^{+}00]. ascospores [dJBPG02]. Asellus [AGD06]. Ash [BMF^{+}06]. Asia [CZL05, SOK03, TSBH09, WSF^{+}00, WSF^{+}02]. Asian [SOK06]. Asiatic [AFLB09]. aspect [BHW^{+}08]. aspects [KIL^{+}03, LG00, RFA^{+}01]. aspen [SRB06, SVB09, WGVB01]. aspera [CMM02, LvNMvdB04]. assemblage [DRDD01, Pen09, Sin07]. assemblages [BGL01, CM06a, CM07a, FK07, HRMC01, ICGÁ05, KLPP07, LMM^{+}07, MFG^{+}06, PGA06, VPR^{+}09, VML^{+}06]. assembling [RL09]. assembly [GDP09, LK07, Rey03]. assess [CBMP07, CB07a, DMH^{+}03, FRZ00, GL06, Jan01, KBK07, KBF^{+}08, LS02a, LPFL09, MHP^{+}06, PK08, Sav00, Sno08]. Assessing [BCM^{+}08, BB04, BRP^{+}06, CL08a, CWCH09, CWL05, CNG06, Cou03, CG06c, DDLB02, Dor07, EGE^{+}08, FJ05, FBDM09, GSBN03, GdBD08, HHM01, IVP08a, IVP08b, Jor05a, KTL^{+}05, LKP03, LTLH08, LCSZ09, MM06a, MG09, Met01, Met02, MLF^{+}06, MC01, PT02, PBA06, PCB07, RBEZ08, TEMJ06, TNO^{+}09a, WX09, XHH^{+}04, AFLB09, BS04, CWF03, Dow05, IG02, JLAM09, ORS^{+}09, PM09, RLR09, SSNB07, SRK06, TB06b, WLL^{+}08]. Assessment [BCLR04, EDKF06, FK03, HKB02, IGP^{+}03, KPKP06, KP09, KH02, LWBW05, SRR06, SGHG04, URB06, WYT09, WG00, AF09, ADSO08, AQS^{+}07, Aus07, BdB05, BMS^{+}08, BKC^{+}07, BFS03, BFS05, CCC04, CUS01, CFPV08, CdQO06, CJY^{+}09, CGHW05, CBD^{+}09, CBSLS07, FPDP06, zGsLcXwZ09, HXP^{+}09, HHKH09, JW09, KFR06a, mLMfT05, LZZ^{+}07, LSY^{+}09, LLCL04, MW03, MCJ^{+}04, MSB08, MHSP^{+}06, NI08, NS07, OW02, Ort08, OACB09, PKS^{+}07, PWSS07, RCL06, Sch00, SHK^{+}07, SBHH07, SYCU09, TOS09, TSJA02, VCM07, WMH08, Wim04, WGT05, WT04b, Xia00a, Xia02, XTDL01, XLZ^{+}04, XZZ^{+}05, ZGL08]. assessments [HRMC01, WSCR03, WEW01, vWSH08]. assimilating [MCJB08]. assimilation [CLTH08, FL09, JF00, KK04, MBK^{+}03, PCS03, ZWXF05]. assimilative [KMN^{+}07]. assisted [ZNC^{+}06]. associated [BBGH07, HY07a, KWD^{+}04, OBE^{+}07, SCAP05, SKvdW09]. Associates [Log02]. association [CM05b]. assumed [WL04]. assumption [Gra05, ZAM^{+}03a]. assumptions [GG00, Wim04]. aster [ZHMN02, TR03]. asymmetric [BBP08, OT03]. asymptotic [SMR08]. Asynchronous [CLHB^{+}08, MY05]. at-sea [RHB06, YHC04]. Ateweberhan [GB08]. Athens [Ano09-33]. atherinids [BMBOCR03]. Atlantic [KR04, BC03, BC01, EFJ^{+}08, GAB^{+}09, HRC^{+}07, HDB^{+}06, HD01, RHE06, SMB^{+}06, WHHH07]. atlas [SBS^{+}06]. atmosphere [Alo04, BPC04, CCJ07, OIR^{+}08, RMSS02]. Atmospheric [VLD02, BB07a, BMT07, CCC00, HPA00, LH09, PG04, PWC^{+}09, RGF00, ZAM^{+}07]. Atoll [Bar04, HE09]. attain [LE04]. attenuation [BCD^{+}05, Lar02]. attraction [DH01]. attractors [FWHL06, vNS03]. Attribute [RIGJM06, BMS^{+}08, BKC^{+}07, Pra00, Pra08]. attributes [CP01, GE05, PGFM04]. attribution [Loe04, Swa06]. attrition [SKvdW09]. August [Ano00-41, Ano00-29, Ano01-27, Ano01-40, Ano02-43, Ano02-32, Ano03-46, Ano03-56, Ano04-65, Ano04-52, Ano05-35, Ano05-40, Ano06y, Ano06-33, Ano06-45, Ano07m, Ano07r, Ano08p, Ano08a, Ano09s, LMPR06, LBL^{+}08]. aurata [HGLLV03, LS08b]. aurita [LS09a]. Australia [HRMC01, WRCB01, AN06, AFLB09, CMPO05b, KKH^{+}09, LR03, MSL01, MB06a, PLC08, RRB^{+}01, RH04, RBE^{+}08, SBG06, SBC07, ZSKV05, ZBL03]. Australian [Wil08, Gri04, LELR02, LC07b, Mur01, TCD02]. Austria [Jør04h]. Author [Ano01l, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano02l, Ano02k, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano03l, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03k, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano04j, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano05g, Ano05h, Ano05i, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g]. Authors [Ano04-44, GB08]. Autocorrelation [FRM09, ASI^{+}08, BKO08, BDF^{+}06, CK07a, EN06, OdKV03, SMET06, WLZ^{+}09, WWC^{+}07, dFOV07]. autologistic [Dor07]. Automata [dSPdBB08a, WBvDHZ09, AS00a, BDR06, CMM02, CM06b, zGsLcXwZ09, KST^{+}07, KPS03, LZ07, LLS^{+}08, MKDV08, SM09, SAL07, SFCP02, WLZ^{+}09, YDS08]. Automated [ATD^{+}06, DMF07, KNCP04, ATDK06, ATDK08, KLL^{+}07, RMWW07]. Automatic [ŽJDK06, LIS07]. automaton [AP06, BH02, CHW07, MT02, SKT00]. autonomous [SB06c]. autoregression [Gan06]. Autoregressive [GB02, FBC02, JKJ^{+}08, LLL^{+}07]. autotroph [BBM04, Sar04]. availability [CFCP04, PFBBJ08, SM04a, TPJ08, WHP03]. available [JAB^{+}06, Jør08c, PDS07]. Aveiro [LCM^{+}09b, ROQ^{+}09, LS06]. average [GB02, LS04b]. averaging [PMJ08]. avian [CM07b, SO06]. avoid [LQL05]. avoidance [JBB^{+}05, LHB08]. Awassa [FM07]. Axios [NKK^{+}09]. Ayres [Jør04a]. ayu [NSO^{+}08]. azoricus [MCD^{+}08]. Azov [TSJA02].

B [Bro09, Jor05c, Ned09, Nie06, MSL01]. B-spline [MSL01]. Back [Ano06v, IGP^{+}03, Jor05b, Jor05k, KE07b, Ric00, Xia06]. back-propagation [IGP^{+}03]. backward [VPB08]. bacteria [AN00, SKK^{+}03, ZvBS05]. bacterial [KVPA07, KVPA08, MY05, SU08]. bacterioplankton [BH03]. Bactrocera [YZS^{+}04]. baculoviruses [SvdWB^{+}06]. bad [Nie08b]. badger [SRS^{+}03]. Bagmati [KLL^{+}07]. Bagré [VOM06]. Bai [Jor06b]. Bai-Lian [Jor06b]. bairdi [RG05]. Baja [ASHHRRPE04]. BAL [LE04]. Balaenoptera [MDB^{+}06]. Balance [ZRR^{+}08, BBGH07, BCLR04, Cai05, ESWG02, FMdV01, FM07, HBM04, Håk09, HG07, KJB07, LBS08b, OSHO09, OW02, PTGI09, RSF^{+}01, Roe01, SV03b, TFTO07, WM06]. balanced [LBBN06, LCG^{+}09, Ort08, OACB09, PK08, PM06a]. balances [ABC04, Weg00]. Balancing [CRM09, ZPP06, AB03, BFU^{+}09]. Balkan [GPSM08]. ballistics [RG06]. balsam [BP08]. Baltic [CT01, AFTB07, HE05, KGNK03, KNB08b, MEKL08, San07]. Bampfylde [AB06]. banana [TMD04, TRDM06, VTL^{+}09]. banded [SM07a]. bandicoot [TIJ^{+}01]. bang [Par02]. Bangkok [FMdV01]. Bank [ALAS09, MRS05, SDDC07, CGH^{+}03]. Baranov [Xia05a]. bare [AYKY05]. Barendregt [Jor05l]. bark [Bye00, JPD^{+}06, SBR^{+}07]. Barley [FPDP06, ARL^{+}06, BMD09]. barnacle [BBM06]. barnacles [DAdSC08]. barred [TIJ^{+}01]. Barrier [Cha02, Gri04, KE07b, MBW09a]. barriers [HC06]. Basal [LE04, CS07, CSF^{+}04, MFB^{+}06]. basal-area [CSF^{+}04]. Basal-Area-in-Larger [LE04]. base [BPC04, KE07b, TBP^{+}07]. Based [SBC^{+}09, AF09, AGD06, ASJD01, AVP08, AOY02, AQS^{+}07, ACJT08, ASZRRR08, BTPL06, BC01, BPP09, BBP^{+}07, BMG08, BPW^{+}03, BKGC05, Ber02, BB08b, Bia03, Bir06, BSG07, BHAV03, BMR^{+}05, BMR06, BBB03, CAB08, CLHB^{+}08, CHL08, CSKP08, Cha08, CSCB04, CCC00, CCY06, CC06, CJ07, CCJ07, CBP^{+}06, CSdZ09, CEK08, CGR03, CWS09, CMPO05a, CMPO05b, CGH^{+}05, DDLB02, DVdB^{+}08, DGD06, DGRU06, DLK01, DFM07, DJ05, DSD^{+}09, DMRP07, EB07b, FSBD01, FPDP06, FL09, Fie04, FMRS09, FAB^{+}07, FBDM09, FH08, Gas05, GFGZ08, GDP09, GP07, GLS02, GGB^{+}06, GRBT08, GM05, zGsLcXwZ09, GS02, GBB^{+}06, GP06, GAPE06, Håk09, HKB02, HD00a, HB02a, HMGK05, HXP^{+}09, HHK07, HB09a, HR03, HHP06, HB05, Hos06, IO02, JT01, JAN^{+}03, JNM^{+}06, JU09, KBW00, KKW08, KM06, KMT09]. based [KST^{+}07, KTKR08, KDK06, KBM^{+}03, KWW^{+}09, KH02, LPCZ07, LBL^{+}08, LS09a, LBBN06, LFJ^{+}06, LZZ^{+}07, Lin01, LPM^{+}09, LB01, LLS^{+}08, LYCU09, MB08, MU02, MM06b, MSHL00, MRT05, MS07, MCKNM09, MEJ06, MBLA03, MKM^{+}07, MKvdW^{+}09, MXC^{+}04, MM06c, MBM06, MR00, MRK^{+}07, MBW09b, Met03, MBS^{+}09, MFB^{+}06, MY02b, MKiIK07, MMJN03, Mun09, MM00, NW07, NHP^{+}06, NWH^{+}06, OG08, OBE^{+}07, OW05, OSHO09, ÖÖ04, PBB04, PK02, PEM06, PE08, Pet02, PVS^{+}09, PBY^{+}03, PdAD^{+}04, PF01, PCS01, PBG09, PBK03, PHWH^{+}09, PS01, RR06, Rai01, RLHD01, RG05, RH09, RvGC^{+}08, RVWH06, Rem04, RLR09, RB02, RHM^{+}05, RGO^{+}06, RLLB09, RCL06, Saa00, SR08, SPS03, Sak09, SAH03, SPB^{+}06, SIK07, SR02, SS06a, SE04, SOA03, SV03a, SBL03, SOK03, SRS^{+}03, SLGS00, SNRW06, SV03b, Sti08]. based [SJ03, SSM^{+}09, SYCU09, SSdlMP^{+}08, SSS00, SHS08, iTI02, TTA^{+}01, TPJ08, TTA^{+}03, TF05, TSF07, TYK03, TYS^{+}09, THJ^{+}03, dJTMBGMP^{+}09, TTJ^{+}09, TKTB07, Uch00, VREA06, VADV06, VH07, VJ01, VCRD06, WTST08, WKZP04, WG07b, WR01b, WW08, WH07, WTMY07, WBS^{+}02, WBR^{+}06, Xia00b, XCW07, XLD01, XBM^{+}07, XBM^{+}08a, XBM^{+}08b, XTDL01, XJM07, YTS03, YDS08, YMD08, YZS^{+}04, ZPGG03, ZYL06, ZYY09b, ZNC^{+}06, ZAM^{+}03a, ZAM^{+}03b, vNLS02, AHP01, FYN^{+}07, GGB^{+}06, LBS^{+}06]. basic [APJ03, JAK^{+}06, KYH00, SM09]. Basin [Ano06-47, BCLR04, CGC01, CUF^{+}09, DGGD04, DVGD07, Foh05, HKHB06, MRiI^{+}07, OPL^{+}09, OFK08, SR02, VZG05, WAB^{+}07, DSA08, FPK^{+}07, LCY09]. basin-scale [MRiI^{+}07]. basins [CMSC01, HKB02, KH02, MBDB09, MBT03, PKS^{+}07, Van04]. basis [BT01, BAP^{+}06, HMBG03, JAK^{+}06, MAB01, PM06b, RIE01]. Basque [GBB^{+}09]. bass [CCLS06, HL04]. Bastianoni [Bro09]. bat [MVGH00]. Bathymodiolus [MCD^{+}08]. Bay [ASHHRRPE04, BB01, KCY^{+}08, LR07b, TM05a, TTHH07, WHH^{+}08, FSJ04a, KSH^{+}03, KGNK03, MSTK08, Mur01, BSB^{+}09, DUASCM07, EFS^{+}03, GBB^{+}09, JLBS09, KHJ^{+}08, Mal01, MMF^{+}09b, SPB^{+}06, TL03]. Bayes [WNW09]. Bayesian [Sno08, ASG^{+}05, ASS^{+}06, AQS^{+}07, BHSR01, BSR04, BRP^{+}06, BHAV03, CB07a, DM03, EG09, FPDP06, FLS09, HL04, MLH05, MDPC06, PWC^{+}09, PWH07, Pra00, Pra05a, PMJ08, QSB03, RR06, RM09, RPPB04, SJG^{+}08, TH08, TMP06, Uus07, WLG07, Wan07, War08, WLB^{+}05, WSP08, ZA09]. bays [Kan04a, SH04]. be [BB07b, Che06, EKBF04, MCSG06, OSS02, RBSG06, SLE07, Szi00, vLLv^{+}07]. BEAM4 [dCPC01]. bear [JDD^{+}03, KA00]. bearded [HMGK05]. bearing [DSD^{+}09]. bears [FPS03]. beaver [PBC09]. Beck [BEM00]. bed [LGC07]. Beddington [NG07]. Beeby [Jor05e]. beech [DDG^{+}05, DDF^{+}05, JDPI07, JAB^{+}06, Kno03, KS08, KBF^{+}08, OHH07, VSG^{+}08, WFB^{+}08]. Beer [TYLL07]. beet [SDDC07]. beetle [BPBL00, CRL09, Ken02, Mat03a, SBR^{+}07]. beetles [Bye00, JR05, JPD^{+}06]. before [EPS04, MWK07, MMF^{+}09b, MMF^{+}09a, OSS02, SSH^{+}07]. began [Kan04a]. begets [QXW02]. beginnings [Arm04]. Behavior [Van08b, AER^{+}07, BG01a, BG01b, BKB00, EB07b, FBC02, FAB^{+}07, FCH04, FCC^{+}04, GNA^{+}06, Hua03, JBB^{+}05, KSH^{+}03, KKCC06, KW02a, LHB08, LC04, Liu01, RJGO00, TM05a, WMH08, WL04, YK00]. behavioral [PAdlPS00, RL05b]. behaviors [MAL06]. behaviour [AMB07, CDAGK06, DLK01, EJG05, For02, For03, GPC^{+}09, MDGV09, MB05a, MHP^{+}06, Mon09b, NJB^{+}09, OB04, RM02, SJH05, TKTB07, UDB07, Xia04, YDS08]. behaviour-based [DLK01]. behavioural [PPP05]. behaviours [DMBO00]. behind [VCC03]. Beijing [ZYY09b]. being [SLPP05]. Beisner [Nie06]. BelEUROS [DVJ^{+}08]. BelEUROS-model [DVJ^{+}08]. Belgium [AGD06, DGGD04, DVGD07, DVJ^{+}08, MLR08, RVRL05]. belief [RM09]. below [HTK07, Urs09]. below-ground [HTK07]. belowground [LL02a, RL05a, vNM02]. Bendoricchio [Xu03, Ano05q]. beneficial [ADS^{+}07]. benefit [DFM07, LTLH08, PCP07, RDS07]. benefits [TB06b, WAB^{+}07]. Bengal [SPB^{+}06]. Benguela [HB00a, HB00b, HSJ04, SFM04, TSJ^{+}09]. Bénoué [MSHP04]. benthic [ASAC02, BPAB^{+}06, BCL^{+}09, BHSR01, CKP^{+}01, DGD03, Gra05, HHD01, JAK^{+}06, JGL07, JKPL09, LTP06, MM06b, NWP06, OW02, OACB09, PKC^{+}01, PLS^{+}06, PSP^{+}07, SSKN08, SHK^{+}07, TN06, VTB^{+}08]. benzene [KPKP06]. Berchtesgaden [WTS^{+}06]. Bergh [Jor05l]. Berlin [Jor05f, Jor05j]. berry [OI07]. best [Kir06a, LQL05]. beta [dSDSM08, LT06]. between [AF05, AYK07, AB03, ABR05, Aus02, BLDN00, BP03, CBS09, CMCD04, CNG06, CM08, DSD^{+}08, DdSG04, DLG06, ECHN08, EWH^{+}02, FCP^{+}07, GPSM08, GYY00, GYW^{+}04, GML05, HCJ^{+}09, HLK06, Hui07, IVC^{+}08, ITDD09, IP00, JPB03, KFS06, KBF^{+}08, KRvL^{+}02, KM00, LGC07, LC07a, MHMHKA04, MBLA03, MML00, MCGO05, Met03, MBKD02, NiTT01, NU00, OCK01, OAAF07, Oku09, OIR^{+}08, Ond07, PEAS01, PHP04, RJR04, Ric02, RA06, SBB06, Sav00, SDA^{+}03, SMSV09, Shi04b, SLE07, SZL^{+}04, SPJB06, SWBH08, TR03, TVS00, VF07, vN02]. Beven [Jor05k]. Beverton [dlS07]. Beyers [Bro04a]. Beyond [Jør04c]. BGC [CMB^{+}02, CMM^{+}07, WBR08, WIH^{+}09]. bias [EB06, JG08, PS09, PMLM08, RB02]. biased [GHP08]. Biennal [Ano08a]. biennial [LBL^{+}08]. Bifurcation [TKK07, VPB08]. bifurcations [GLS07]. Big [MCB06, MD04, VCM07]. big-leaf [VCM07]. bigeye [NCM07]. Bight [GRW04, Wil08]. binary [FM08, Jua09, Mal03, Wil03]. binomial [YSB08]. bio [Dow05, JCB^{+}02, MSS02]. bio-economic [MSS02]. bio-physical [Dow05, JCB^{+}02]. bioaccumulation [KMB08, NG09, SD08, WLB^{+}05, ZK08]. bioassay [SKJvdHG06]. bioavailability [GJ00]. Biochemical [ABV^{+}06, BP04, HK02, YK00]. BIOCLIM [BHP05]. bioclimatic [FK03]. biocontrol [CTH^{+}00, JBR07, SB06c]. biodegradation [CM08, MY05]. Biodiversity [ABR05, RPC^{+}05, CRM09, FWSB05, ICGÁ05, IP00, MGS^{+}09, PMLM08, Ric02, RCL06, SBS^{+}06, WK04, vN02]. bioeconomic [AS03, KS08, MG05, PMD^{+}09, dCPC01]. bioenergetic [BDI04, LS08b]. bioenergetics [CNG06, MWWZ01, MRK^{+}07, MRiI^{+}07, SJH05]. bioenergetics-based [MRK^{+}07]. biogenic [vBBE^{+}08]. Biogeochemical [FEP^{+}04, FPSJ04, FSJ04a, ORF01, OBR01, Bog04, ECBD09, GRPF07, HOK^{+}09, Ito07, KC01, LIS07, LFJ^{+}06, LS09b, LSS^{+}00, LLLT08, MMR06, MR06, PHBF07, RBSG06, RAI04, VJ01, XS08, ZA09, ZAM^{+}07]. Biogeochemistry [Lim04, Ano04w, FYN^{+}07, RNKG03, ZLZM02]. Biogeoeconomics [Bog04]. biogeographic [Gut07]. biogeographical [Pen00]. biogeomorphic [MLM06]. bioindicators [Håk09]. Biological [Jor06b, Sem08, BvdW04, Bor06, CJS^{+}02, CIM07, CP09, CMSC01, DCP^{+}07, FL09, GRHS00, KWW^{+}09, LKR03, LKH^{+}08, LD06b, LC01, MDGV09, MG09, MBGP08, NT07, RDS07, SS07, SSNB07, SGLH04, SvdWB^{+}06, TN09, Tia06a, Tia06b, TNO^{+}09a, Wil07, XV03, Xia05b, YSM00, dCPC01, dSMZ09]. Biologically [PC07, LR07b]. Biologically-inspired [PC07]. biology [HN09, Xia00c, Xia03]. Biomass [ZAM^{+}03b, ASAC02, BKGC05, BLC^{+}07, BB08b, BH03, BHP08, CGGE08, DUH03, EPS04, EFS^{+}03, Fle01, FMP^{+}00, GSM08, HB02b, HB03a, HS01, HS06, IWW08, KA01, KTKR08, KK07, LB01, MMM02, PGG06, Roe00, SSK^{+}07, SM04b, SP01, STK00, SDS01, SKM^{+}06, TYS^{+}09, Urs09, WMSW09, WCS00, WHP03, XSY^{+}09, ZCG^{+}08, ZBSA07, ZAM^{+}03a, dlS07]. biomass-based [KTKR08]. biomasses [HB03b, PCM^{+}03]. Biome [RJS^{+}06, CMM^{+}07, WIH^{+}09]. BIOME-BGC [CMM^{+}07, WIH^{+}09]. biomes [SOA03, WSZS08]. biomimetic [RL09]. biomonitoring [KKA^{+}08]. Biondini [Zha03a]. Biophysical [MCJ^{+}04, Mat06, RSC09, TCGL03, ZNC^{+}06, ZGL08]. bioreactors [HG07]. bioregions [GW08]. biosecurity [AFLB09]. Biosphere [DCG01, KK08, SF07, BDE08, Kir01]. biota [CBSLS07]. Biotic [Lud04, MGH^{+}05, BNM09, IGP^{+}03, LA07, WLLY04]. biotime [Kir06b]. biotope [MNEB01]. biphenyls [RBW09]. bird [BDF^{+}06, GCLG03, LKLK07, NBP05, OW06, SBDD04b, SBDD04a, SCAP05, TDL^{+}07]. birds [MW03]. birnavirus [LYC08]. Birth [SÅ06]. Biscay [GBB^{+}09]. Bison [HGD05]. bistability [SAI09]. bistable [GJ07]. bits [dMPVNO07]. bivalve [Dow05, SM04a, dSSGR00]. bivalves [ZDR03]. Black [BEM00, CHHP02, FAB^{+}07, LKY^{+}06, AMSW07, FPS03, HBUS02, LPD08, MB05a, New06, OJ02, PT02, SPM^{+}08, LMGM^{+}09, OMRD01]. black-capped [PT02]. Black-legged [FAB^{+}07]. Black-tailed [LKY^{+}06, HBUS02]. blackbird [ÖTÖR06]. Blackgrass [CS01, CCGMJ07]. blackwater [HBRW07]. Blackwell [Jør09b]. blade [DCP^{+}07]. blade-strike [DCP^{+}07]. bleaching [RP09, YSB08]. Bled [Ano06-48, ATD^{+}06, DDŽ06, DŽD06]. blocking [EFJ^{+}08]. bloom [CM04, HB06, JKWJ03, OAL^{+}07, Ond07, PCpDC09, RH04, VBFM^{+}08]. blooms [CM06b, JPB03, JCX^{+}08, LHDJ03, LLL^{+}07, ML05, Oke04a, PWZ^{+}09, RT08]. blowdown [RF09]. Blue [GS02, BR01, HMBG03, KTKR08]. blue-green [BR01]. blue-greens [HMBG03]. bluefin [LG04, Wil08]. boar [HAS07]. Board [Ano04t, Ano04u, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano09h]. boating [AF09]. bobwhite [LGD01]. bodied [KD05]. bodies [BB08b, Mon02]. body [OSHO09, OzDBS07, RBSJ01, Yos06b]. bog [LAPAMD07, Zav08]. Bolivian [KBK07]. bollworm [GP06]. Boltzmann [SNRW06]. Boltzmann-based [SNRW06]. Book [Bro09, Jor01a, Jør02a, Jor05b, Jor05c, Nie08a, SRB06, Zha06]. Boophilus [CTG03, CTG04, TCGL03]. boost [vNS03]. boosting [MFB^{+}06]. bootstrap [LdVA06]. border [Hee02]. Boreal [NSEDP06, ZPD^{+}08, DFF07, GGPBEK07, GRBT08, GAA^{+}05, GZY^{+}06, HFSH06, KW00, KCZ^{+}03, KLL01, MHM^{+}03, NW06, Nun03, PLH09, PJAZ02, PGFM04, SPM^{+}08, SJG^{+}08, TBP^{+}07, VSHC08, WGVB01, Wan05, WSS^{+}06, YP02, YL07, ZPD06]. Borer [BMF^{+}06, TMS^{+}07]. Borneo [PBY^{+}03]. Bornhöved [RHM^{+}05]. Botanic [KWW^{+}03]. Botswana [MMWMH07, ZPW05]. Bottlenecks [ABB06]. Bottom [Her04a, BCL^{+}09, BSG07, HY07b, MSTK08]. Bottom-up [Her04a]. bottomland [NH07, ZBW05]. Bound [Jor05a, Jor05c, Jor05d, Jor05f, Jor05g, Jor05j]. boundaries [MWP07, TTJ07]. boundary [FGB08, GVDF07, Met03, VM01, vdHGF09]. bounds [CG07, CG06a]. bout [WC07]. Bovine [CG06c, SRS^{+}03]. box [BA05, GCG^{+}07, IMK^{+}07, OJ02, LdVA06]. Brachyramphus [YHC04]. Bradhyrizobium [PMH00]. branch [vNM02]. branches [LSM^{+}04]. branching [Lir03, SB00, SLE07]. Brandenburg [HVVK09]. Brassica [BHM^{+}06, CBP07, CML09, HC09, PvdBW^{+}02, SF07]. brassicae [PvdBW^{+}02]. Brazil [GRW04, ADSO08, AP00, CdQO06, FMC^{+}08, Phi04, RdQFO07, SSK^{+}07, dCSB08, dSA01, dBWG04, dBWG05]. Brazilian [ACE07, PEE09, PR01]. bream [LvNMvdB04]. breeding [BDP^{+}02, BKMB08, HD01, LKLK07, Oli03a, ÖTÖR06, POF08, WZC^{+}05, YHC04]. Brennan [Jor05e]. Breton [GB01]. breviscapa [dCSB08]. Bridging [KBF^{+}08, LS09b, Ric02, RA06]. brief [Jør09a, SH04]. Brit [Jor05a]. British [APHV08, Jør04c, Jør04d, Jør04a, KRK05, WKZ03, YHC04]. Britt [Jør04f]. Brittany [GGV^{+}06]. Broa [AP00, RdQFO07]. broad [CKPP03, ZBZ^{+}09]. broad-leaved [CKPP03, ZBZ^{+}09]. broadleaved [BW01a]. bromoides [DKRŠS00]. Bromus [WCH08]. brook [WPD04, ZAM^{+}07]. brown [BEM00, BRP^{+}06, CDM^{+}00, JDD^{+}03, KA00]. Brows [BEM00]. browsing [RPRV09, SL01, WBB05]. Brunham [Jør04e]. Brusc [BMT07]. bryozoan [LK03]. Bt [GP06, GAPE06, LT04]. Bt-corn [LT04]. buck [Par02]. bud [Kai00, PAF06]. budburst [CB07b]. Budget [WLL^{+}08, BFOS08, CCC00, HvG07, HC09, MCM^{+}09, MKS^{+}02, QXW02, RR01, RR05a, SW03, Tan02, TYT^{+}09]. budgetary [MMPN07]. budgets [GCM^{+}05, KCNN06, PGW00, SPM^{+}08]. budworm [BP08, FBC02, SGLH04]. buffalo [CG06c]. buffer [DCG01]. Buffon [Bir06, MSHP04]. Bug [MHP^{+}06]. build [WBS^{+}02]. building [AHKB01, BLDCM06, OSS02, ÖTÖ06, PGW00, WCH08, WGV^{+}08, Jor05h]. built [KM04]. Bulgaria [GPSM08, RNC08]. bullhead [CSKP08]. bunch [HB09a]. bunchgrasses [TDHO07]. buoyancy [SJH05]. buoyant [NP06]. Burkina [KRvL^{+}02, VOM06]. Burlington [Ano08a, LBL^{+}08]. Burman [BH01]. burned [Las06a, SLPP05]. burning [RCZ^{+}06]. Burrinjuck [WRCB01]. bushcricket [Har08]. bustard [SSdlMP^{+}08]. butterfly [AKB07, GS02].

C [Jør04d, Jor05l, Swa06, Ulg04, ZGF^{+}05, BHI^{+}06, FCC^{+}00, IWM^{+}06, KNB08b, LBS^{+}06, MCD^{+}08, SYC04, TFM01, VVF06, WHP03]. C-CLASS [ZGF^{+}05]. C-Fix [VVF06]. C-flux [MCD^{+}08]. C. [ZSPV08]. CA [WKL^{+}03]. cabbage [ZBL07]. CABI [Jor05c]. Cabras [FU05]. cachinnans [BL01]. CAEDYM [TSJ08]. Caiman [dCS05]. Calabria [Las06a]. Calanoida [DSD^{+}09]. calculate [LH01]. calculated [SMR08]. calculating [Xia06]. calculation [Deb02, TF09, WCC02, ZLL05]. Calculations [JLDM05, JU09, KCY^{+}08]. Caledonia [PE02]. calibrated [AAO06, RSF^{+}01, SDL08]. calibrating [ZA09]. Calibration [BRGS09, CSR08, CMB^{+}02, HTA^{+}08, LLW^{+}06, MLF^{+}06, RMWW07, ZAM^{+}03b, AB05b, AQS^{+}07, BMG08, CMPO05a, FPCA00, GFG09, Jon07, Jor01c, JRBS02, LIS07, LHC07, LK02, Log08, MRK^{+}07, MMJN03, Mur01, PWSY07, RSW07, SKvdW09, SJG^{+}08, VA00, WIH^{+}09, XHC^{+}08, ZRR^{+}08]. calibrations [BVC^{+}01]. California [ASAC02, ASHHRRPE04, DUASCM07, MZASLMLC04, FSM^{+}01, GB02, GKG05, KTB07, KSG05, MWWZ01, Pyk04, RLSZK^{+}08, SF04, WFHP07]. call [Ano01m, Oku09]. caller [Par02]. calling [Par02]. Callitris [RBE^{+}08]. caloric [LC08]. Calvi [EFS^{+}03]. Camargue [MBLA03]. Cambridge [Jor05l]. Cameroon [BLBT01, MSHP04]. Campeche [ALAS09]. Campololongo [Jor05a]. Can [BP07, BB07b, Den09, LQL05, MCSG06, RBSG06, Str09, Szi00, Jag01, Jag09, KK00b, LL00, MP08, MLGG09, PBC09, POF08, vLLv^{+}07]. Canada [Ano06-45, LMPR06, SSH^{+}07, FBDM09, GRBT08, GB01, HRC^{+}07, KRK05, LLW^{+}06, LPC05, MMF^{+}09b, NW06, PJAZ02, YHC04, ZMB^{+}08, ZAM^{+}07]. Canadian [LAD06, PLTT05, SPM^{+}08, WGVB01, WLL^{+}08, ZGF^{+}05, ZPD^{+}08]. Canal [AAA00]. canaliculus [RR05a]. canid [PBK03]. canids [RMS08]. Canis [FCKH06]. Cannibalism [KT03, PSC04]. canonical [FGFB05, Kai00, RKH05]. canopies [Che06, JNM^{+}06, MBM00, Ski04, SLT^{+}09]. Canopy [WBN^{+}03, AJJ^{+}05, BSJ^{+}02, DBD^{+}08, ESWG02, FCC^{+}04, GQB05, GBS00, HTS00, Jon07, KH07, MBK^{+}03, MLA^{+}02, PvdBFJ02, SI06, SKvdW09, TH06a, TYT^{+}09, WCC02, WYT09, WLJ00]. Cantabrian [SO04]. Canyon [FCC^{+}04, YBM^{+}05, KPKP06, RSB02]. capabilities [CW04]. capability [LL07b]. capable [LMH05]. Capacity [BK07, DGM08, DMH^{+}03, GCG^{+}07, HHK07, Hui06, MLF^{+}06, RR05b, SM04a, SG05, XS08, YTLF08, dlS07]. Cape [GB01, LXP^{+}08]. Capercaillie [KK00a]. capped [PT02]. capture [SRL^{+}00]. captured [WBvDHZ09]. captures [BJFM06]. capybara [FC05a]. carabid [JR05]. Carbon [BRGS09, DP07, LPPS05, PG08, dCSB08, dSLSD02, SHSP04, WLL^{+}08, YL07, Ale07, BB08b, BFOS08, BCG08, Cai05, CCC00, CB07b, Chu08, DPL^{+}04, DDG^{+}05, DVdB^{+}08, DP03, DDF^{+}05, EM08, FCC^{+}00, GSBM08, GARB09, GS08, HHB^{+}08, HB00a, HFSH06, IO02, Ito05, JL07, JGL07, KA01, KW00, KS04a, KvKV05, KFR07, KKH^{+}08, KKH^{+}09, KGNK03, KDW^{+}09b, LRJ^{+}09, LBL^{+}08, LBBC08, LPC05, LLLT08, LAZ^{+}08, LH09, MH02, MBK^{+}03, MCBA07, MCM^{+}09, MGCK^{+}03, MMRLP06, MH05, MLF^{+}06, MPS02, MMPN07, MMN08, NW06, Oga09b, OBE^{+}07, OIR^{+}08, OHM^{+}06, PPR03, PJAZ02, PLD^{+}02, POH02, Pot04, QXW02, RGF00, RBC09, Roe01, SC04, SOHY07, SW03, SPM^{+}08, SJG^{+}08, SHS08, Svi02, SSH08, TYT^{+}09, TR05, TDLP03, WGVB01, WCHM02, WMM^{+}07, YWHW02, YYZ06, YY06, Zav08, ZGF^{+}05, ZJN^{+}06, ZPD^{+}08, vN02, vdPvOV00]. carbon- [ZGF^{+}05]. carbon-14 [KGNK03]. carbon-budget [CCC00]. carbon-dynamics [KDW^{+}09b]. Caribbean [MHZ^{+}06, PCC^{+}07]. caribou [FCH04, JG08, TFF07, WH04]. Carlo [Ann01, GTJ^{+}00, KSvOO09, QSB03, WSCR03, WT04b]. Carlos [AP00, RdQFO07]. carnivore [PSBJ07]. carnivorous [RRB^{+}01]. Carolina [RBW09, LXP^{+}08, SRW05, WDW00]. carp [BW04]. carpet [SM07a]. carpio [BW04]. Carpobrotus [SMTR07]. Carposina [KL03]. Carposinidae [KL03]. carps [LCLC07]. CARPSIM [BW04]. Carrying [BK07, Hui06, DGM08, DMH^{+}03, GCG^{+}07, HHK07, RR05b, SM04a, XS08, YTLF08, dlS07]. Carson [CWH^{+}00, CW01]. CART [GSC09]. Cartagena [TL03]. cascade [Her04b]. Case [Jor05j, KSÇ^{+}00, OW06, PSVH09, Rem04, AW06b, AN06, AP02, BRV09, BBC03, BMR^{+}05, BCG08, CBP07, CC06, CL08b, CGHW05, Don06, DDFP07, DAR^{+}07, FJ05, FEP^{+}04, FBDM09, GHP^{+}09, GGPBEK07, GGV^{+}06, GWK^{+}06, GGB^{+}06, GP02, zGsLcXwZ09, HvI01, HAS07, HB09b, IOIA08, JAN^{+}03, JJ00, KWD^{+}04, KHJ^{+}08, LMMK09, LWLZ06, LZZ^{+}07, LCY09, LS09b, LTLH08, LMG08, MI01, MEJ06, MHP^{+}06, MY02b, MAK^{+}04, Oga09b, OPL^{+}09, PBA05, PBPZ04, PS09, PHH00, Rai08, RHM^{+}05, RNC08, RPVR03, RVRL05, RCL06, SSK^{+}07, SFM08, SPK^{+}09, SGY01, SCP05, SBMJ09, SS00, SvL04, TMJ04, Tol06, WR08, WGS^{+}02, WM06, WYMS07, XLZ^{+}04, YZC^{+}07, ZLO02, ZJTB03, ZBL07, ZZ08, ZYY09a, ZYY09b, ZCY09, ŽJDK06]. Case-based [Rem04]. case-study [BCG08, DAR^{+}07, PS09]. cases [Zha07]. CASTEAUR [DBBS03]. Caswell [Log02]. cat [PAdlPS00, PFBBJ08]. catabolic [AVTP05, VATP05]. Catalan [CPTD08]. Catalina [KSG05]. Catastrophe [WRP07]. Catastrophic [TB05, MHZ^{+}06]. Catastrophic-like [TB05]. Catch [MTKM^{+}06, Xia05a, Xia06, PPE^{+}07, SSS^{+}09, WTMY07]. catches [LG04]. catching [PLLdCB06]. Catchment [JWW^{+}07, AFT09, BPW^{+}03, BBGH07, Bor07, ESWG02, HKPH08, HSV07, HZF07, KKL^{+}06, NKK^{+}07, NKK^{+}09, RVRL05, WHX^{+}03, ZSKV05, ZBL03]. Catchment-scale [JWW^{+}07]. catchments [BDR01, HHF05, MCJ^{+}04, TMLV07, ZGL08]. CATCHSCAPE [BPW^{+}03]. Categorical [Met03]. categorization [PSVH09]. cattle [CTG03, CTG04, Eza05, GZ05, GPC^{+}09, TCGL03]. cattle-fever [CTG03, CTG04, TCGL03]. Caulerpa [CTH^{+}00]. causal [LiWZ00]. Causasilty [Gan06]. causation [IG02]. cause [Yos08]. caused [APCdlR07, LLCL04, PMC03, SOK06]. causes [CAB08, MM00]. cautionary [CGR03]. Cave [BLDN00]. cavity [LGL02]. CBM [KDW^{+}09b, WLL^{+}08]. CBM-CFS3 [KDW^{+}09b, WLL^{+}08]. CCAMLR [CDA08]. CCM [SAI09]. CD [Hey01a, Jor05i, CWCH09, VCD05]. CD-ROM [Hey01a, Jor05i]. CDE [Jor05g]. CDE-ROM [Jor05g]. CDPD [Ben06]. cedar [Ito07]. cell [BM07, XHH^{+}04]. cell-level [XHH^{+}04]. Cellular [dSPdBB08a, WBvDHZ09, AS00a, AP06, BH02, BDR06, CMM02, CM06b, CHW07, zGsLcXwZ09, KPS03, LZ07, LLS^{+}08, MKDV08, MT02, SM09, SKT00, SFCP02, WLZ^{+}09, YDS08]. cement [AW06a]. census [CGR03, Oli03a]. central [BJJ06, GRBT08, Hee02, Hör03, Kri03, LvGC^{+}04, OW02, PJAZ02, RLSZK^{+}08, SHCS04, WG07a, WKZ03, KW02b, LLCL04, NAC04, OIP^{+}08, TC09, DDJ^{+}01]. Central-European [TC09]. centrality [JBP07]. Century [LMPR06, YLJ^{+}01, MBD^{+}00]. CERES [ETH^{+}04, SC01]. certain [OEK^{+}06]. Cervus [DDJ^{+}01, PSVH09]. CFS3 [KDW^{+}09b, WLL^{+}08]. Chagas [CAG03]. chain [Bal00, BAP^{+}03, GLS07, HK02, KBvVK08, LK02, MdRdA06, SHZ05, lXzL02a]. chains [DD00, GWHS03, Her04a, HH04, KPAK02, LBBN06, LL00, VP01]. chalcographus [JPD^{+}06]. Chalk [JWW^{+}07, BHW^{+}08]. Challenges [SMSV09, SBC^{+}09, AMRS08, DGOZ04, GH09, Uus07]. chamber [BSJ^{+}02]. chambers [HHD01]. chamois [FE04]. chamomile [dCBL09]. chance [BP07]. Chandoli [IKS09]. Change [AEP^{+}04, BK07, Jør04h, Jor05f, AFLB09, BLDN00, BMBOCR03, BBT06, BFOS08, CV07, CG00, CT07b, CN07b, EBH^{+}01, EMdC^{+}01, GYY00, GGPBEK07, Gar04, HY07a, HPA00, HFSH06, HVVK09, HHKH09, HCL00, KW02b, KFH07, KR04, KBM^{+}03, KDW^{+}09b, LHC09, Loe04, MSHL00, MVPB02, MSM^{+}08, Mey04, MB00, MDJ09, MGSdG07, MSB08, NG09, NI08, Oga09b, OHH01, PEM06, PPR03, PSCS^{+}01, PHD04, PDBJ09, PKS^{+}07, Pot04, Pra09, RNKG03, RIE01, RWW07, RGL^{+}07, RGF00, RBEZ08, RB09, SDA^{+}03, SNF01, SMR08, SS08b, Sti08, Swa06, TEMJ06, TFF07, WGB^{+}08, WHP03, YLJY03, ZH06, ZWCL05, Nie06]. changed [GTJ^{+}00, Saa00]. Changes [CN07a, FCC^{+}04, HSJ04, SBC^{+}09, YLJ^{+}01, BBGH07, BP08, CCC00, CPT09, DS01, DLL^{+}09, EFJ^{+}08, FHE06, FBC02, GRW04, GGS08, JMN02, JRT02, KPK^{+}07, Kir06b, Kri03, LZZA09, Lud09, MDH^{+}09, MPC06, Met01, Met02, OIP^{+}08, OL09, SB06b, SVS04, SDL08, SSdlMP^{+}08, TC09, TDdSLS^{+}08, TSJ^{+}09, VZG05, VSFM03, WZ01, WFM01, WSY^{+}07, ZM08, dBWG04]. Changing [BSG07, FPS03, GRBT08, LNW00, PAF06, RP09, SK04, Sin07, SBC^{+}09]. Changjie [Jør06a]. Changsha [LZZ^{+}07]. Channa [JLAM09]. Channel [SM07a, DS02, LGC07, MLM06, SFU01, UZ01, AMSH08, VBFM^{+}08]. Chaos [CS03, MdRdA06, Den08, Gut07, MRRJ06, Ned09]. Chaotic [VP01, MRRJ07, Wil01]. Chapala [BMBOCR03]. Chapalichthys [BMBOCR03]. Chapman [Jor06b]. Chara [CMM02, LvNMvdB04]. characterisations [MFB^{+}05]. characteristic [Håk00, MHN00, PPS08]. Characteristica [MC04]. characteristics [CG06b, FYN^{+}07, Jag01, Jag09, KKA^{+}08, Kri03, LR09, MF02b, NG09, SB06b, SAS06]. Characterization [KMR^{+}07, RPC^{+}05, SL04, ZSD^{+}08, APG^{+}03, AGM^{+}08, FCM05, Fis09, LL07b, UZSM05]. characterize [YDS08]. Characterizing [FHE06, PDL^{+}08, RMBM06, VMG05, KYL07]. Charisma [vNSvdBC03]. charr [MY02b]. Chatelier [Kir06b]. Cheltenham [Jør04a]. Chem [HBO07]. Chemical [CJ07, LJW03, Per07, RL09, HBO07, MWH00, WLB^{+}04, vWSH08]. chemicals [BB07a, BRW^{+}05, MHSP^{+}06, vWSH08]. chemistry [CCC00, HBO07]. chemostat [VPV09]. chemostats [BB08a]. Chernigov [vdPLG^{+}00]. Chesapeake [JLBS09]. Chi [LTLH08]. Chi-Jia-Wan [LTLH08]. chicks [MPRJ04]. Chief [Fat09b, Fat09a]. Chile [ECHN08, MWM02, MTD^{+}09, NAC04, OW02, Ort08, PG08]. chin [MI01]. China [JCX^{+}08, VPR^{+}09, YLLW02, ZYY09b, CCY06, CQ07, CL08b, CZG05, zGsLcXwZ09, HHL^{+}02, HJZ06, LG00, mLMfT05, LWLZ06, LZZ^{+}07, LWC^{+}07, LSY^{+}09, LL08, LM07b, MLL^{+}05, SGY01, SWBJ01, TYZ^{+}05, WZJ08, XSY^{+}09, XHH^{+}04, XLZ^{+}04, YZC^{+}07, YLJ^{+}01, YWL^{+}05, YFLW06, YTLF08, ZH06, Zha07, ZCG^{+}08, ZBZ^{+}09, ZNC^{+}06, ZCY09, ZSZ06]. chinensis [JZY07, WLZ^{+}09]. Chinese [CC06, CJY^{+}09, SYCU09, WX09, WLZ^{+}09, ZYY09a]. Chiricahua [BM08]. chiricahuensis [BM08]. Chironomidae [KKCC06]. chironomids [KKCC06]. Chironomus [CFCP04]. Chirostoma [BMBOCR03]. Chlorella [FCP^{+}07]. chlorophyll [AP02, BW01b, EPS04, GL01a, HMBG03, PEE09, WR01b, ÇDKK05]. chlorophyll- [WR01b]. Chlorophyta [KM00]. choice [LMGM^{+}09, WL04]. choices [Eti04, VM06]. cholinesterase [PGM08]. cholinesterase-inhibiting [PGM08]. Choosing [Tol06, TBPF08]. chorus [Par02, Har08]. chronic [SM06]. chronological [IC02]. Chrysomelidae [SBL03]. chubs [DBBS03]. Ciénaga [LT01]. Cilento [RMBM06]. circadian [ML04]. Circulation [XIX^{+}08, CD05]. cities [MCB06, SYCU09, ZYY09a]. citri [AFLB09]. Citrus [dSPdBB08a, AFLB09, YSG^{+}06]. City [LZZ^{+}07, VLD02]. clade [Yos06b]. Cladocera [ZPGG03]. clam [BSB^{+}09, LYC08, MG05, RC06]. clams [HC03a]. CLASS [ZGF^{+}05, GC02, HM01, hMzL08, Met03, WGVB01]. classes [TBPF08]. classic [LUKD06]. Classical [LUKD06, LJ09, Lud09, PG08]. Classification [GG04b, GW08, LKLK07, BBB03, CWS09, DDJ^{+}01, DDL^{+}06, ECZ^{+}06, FM08, GICB09, KCD^{+}04, Met03, MF02a, SFC05, SAS06, TMP06]. classified [PBHGF07]. classifiers [PVS^{+}09, SAS06]. clear [ORS^{+}09, PG04, WTS^{+}06]. clear-cuttings [ORS^{+}09]. clearing [AN06]. Cliffs [SRB06]. Climate [BK07, Loe04, PDBH02, RWW07, Swa06, AFLB09, BLDN00, BHW^{+}08, BBT06, BFOS08, CCC00, CT07b, CN07b, EM08, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, FCF^{+}04, GGPBEK07, GRBT08, GTJ^{+}00, HRH^{+}05, HPA00, HFSH06, HVVK09, KLM^{+}02, Kir06b, KFH07, KBM^{+}03, LZZA09, LLF02, LGD01, MBMP06, MVPB02, MXC^{+}04, MSM^{+}08, MRiI^{+}07, MSA^{+}03, Mey04, MDJ09, MGSdG07, NHP^{+}06, NI08, OLK^{+}04, OHH01, PLTT05, PLA06, PSCS^{+}01, PDBJ09, Pot04, Pra09, RGF00, RP09, Sem08, SBDD04b, SNF01, TC09, TEMJ06, TFF07, TPJ08, TMS^{+}07, WHP03, XHC^{+}08, ZM08]. climates [BEF03, HO01, Saa00, TM05b, vLLv^{+}07]. climatic [AKMB01, Aru05, BHP05, BKPS08, BAP^{+}06, GYY00, KBGJS06, KBK07, KBB00, MTD^{+}09, MSHL00, Mii00, MC01, PT08, RNKG03, SB06b, VSFM03]. climato [MSA^{+}03]. climato-topographic [MSA^{+}03]. Clonal [SMTR07, IWW08, LUKD06, OCK01, TDHO07, WvS06]. clones [BPBF^{+}00]. closed [KL01, MSD06, MHN00]. closure [CUS01, Mit09]. closures [ZR04]. cloud [WHZ06]. clouds [DWD07]. clover [BPBF^{+}00, LCF09, LWL^{+}02, ZJG^{+}06]. Clumped [PJGW06, BSTS^{+}02]. Clupea [MRK^{+}07]. cluster [BHAV03, Gre06, MML02, SNRW06]. clustered [YSB08]. clustering [DDT07, LS02b, RMBM06, WW09]. clusters [LS09a]. CN [DP07]. CO [RNKG03, DMBO00, BM00, IVC^{+}08, JWLA00, NWH^{+}06, TDL^{+}07, ARL^{+}06, AS09, BW01a, BMT07, CGY08, CGdPRY09, GAA^{+}05, GZY^{+}06, GTJ^{+}00, Gut07, HHB^{+}08, KBE^{+}06, MOP^{+}05, MOP^{+}06, MBF09, MWD05, RFA^{+}01, RGF00, dSLSD02, Tan02, VA00, YY06, ZXC03, ZGF^{+}05, vODFS04]. Co-evolution [DMBO00, JWLA00]. co-existence [IVC^{+}08, NWH^{+}06]. co-production [BM00]. co-viability [TDL^{+}07]. CO2FIX [MGCK^{+}03]. coal [AM02]. coarse [MJR06, TK01, ZKH09]. coarse-grained [MJR06]. coast [OACB09, PCM^{+}03, RHE06, GGV^{+}09, KR04]. Coastal [Ano06-47, ZCKR07, ATAK00, ABP05, AACIS^{+}08, BGF00, BHSR01, BMT07, CPH00, CHL08, CM04, CM06b, CSM^{+}06, CUF^{+}09, Dow05, DLL^{+}09, DMH^{+}03, Esp03, FU05, GRW04, HGB04, HPF08, JRT02, LHDJ03, LXP^{+}08, LT01, LCM^{+}09b, MEKL08, MNZ06, MMPT06, MG02, ML05, NKK^{+}09, PBA05, PBPZ04, PG08, RŽČ^{+}04, RMWW07, SNF01, SMTR07, TDdSLS^{+}08, TSZdRR03, TPCS06, VTB^{+}08, VCAS01, VFTB06, VCRD06, WHHH07, XLZ^{+}04, XIX^{+}08, YHC04, ZBD^{+}09, ZRASC04, Zue07]. coastline [BDP^{+}02]. coccolithophore [BSM08]. cod [AFTB07, PNN^{+}08, SMB^{+}06, TNK04]. cod-dominated [PNN^{+}08]. code [Sal06, SA07a, dBWG04]. codling [TTJ07]. coefficient [Xia00c]. Coefficients [HMBG03, Håk00, HvI01, Her08, LSAGF05]. coevolving [Pat07]. coexist [BIS09]. Coexistence [PDS07, AN00, AB06, BBGM06, Bor06, CMR09, CBP07, EGE^{+}08, Gri04, PJGW06, RD07, RR05b, Sil07, WZW05, WW08, Zha03b]. coexisting [CSdZ09, ZLX09]. coffee [AZM^{+}06, CR06]. cognitive [ÖÖ04]. COHERENS [MNZ06]. coherent [RBB05]. cohort [BKMB08, TMD04, YZS^{+}04]. cohort-based [YZS^{+}04]. cohort-level [BKMB08]. Coimplication [Pat07]. Col [JPD^{+}06]. cold [MHSP^{+}06, FSE^{+}04b]. Cold-water [FSE^{+}04b]. Coleoptera [JJWF07, SBL03]. Colinus [LGD01]. collapse [BKMB08, SSH^{+}07]. collapsed [ASZRRR08]. collected [CPB^{+}08, PSP^{+}07, SPB^{+}06]. collective [RHM^{+}05, VS08a]. Collembola [CR03]. Collembolan [LAPAMD07]. Cololabis [IMK^{+}07]. Colombian [Tol06]. colonial [OW06, ZPP06]. colonies [NBP05, SKK^{+}03]. colonization [AP06, LS01, SFCP02]. colony [FAB^{+}07, HD01]. color [Bro09, GLD07]. Colorado [Ken02, YBM^{+}05]. colour [Jor05e]. Columbia [KRK05, YHC04, APHV08, HP09, SR02, WKZ03]. column [BHP08, HHD01]. Combination [CV07, SMSR00, BPC07, OHM^{+}06, RKH05]. combine [Sta07]. Combined [GGB^{+}06, MBPS04, DWH06, DDH^{+}09, EPM^{+}04, KA00]. Combining [CWS09, RGL^{+}07, YZC^{+}07, HSRD09, KFR06b, Plu00, RR06, YYZ06, ZWCL05]. combustion [AM02]. Comment [AB06, BBGM06, BFS05, Dun08, GB08, Zha03a, Bio03, Ric02]. Comments [Bar04, Nad07, Ned09, Paw00, Pet04a, San09, Sno08, Swa06, Cha07, KL09, MMV08]. commercial [CNG06, IMS07, LvNMvdB04, NAC04, VCRD06]. Common [Jør09b, WG07b]. Common-Sense [Jør09b]. commonly [KM04, Oku09, War08]. communal [HRH^{+}05]. communities [AW06a, Aok08, BCL^{+}09, BB03a, BKS05a, BCPM09, BPE^{+}07, BHAV03, CP09, DDL^{+}06, DJ05, EIRT00, GSG^{+}04, JKJ06, LKLK07, LAPAMD07, LC01, LTP06, LCSZ09, MMPT06, MR00, MFG^{+}08, MWMN06, OW02, PBE^{+}07, PKC^{+}01, PLS^{+}06, PP04, RVWH06, Reu05, RA06, SHK^{+}07, SRK06, SF04, TPC^{+}07, TN06, WYMS07, YPR^{+}05, Yos08, vNLS02]. Community [OAL^{+}07, PSP^{+}07, Reu05, AS00c, AS01, Bar04, CPP00, CKP^{+}01, DAC^{+}09, EBM06, Fat07a, FK07, FWS^{+}05, GL09, GDP09, GL01b, KPK^{+}07, LS04a, LA07, LK07, LELR02, LS08a, LMPT05, Lud09, MM06b, MS07, MT02, MMLR07, Nal00, Nal01, Oke04a, PBC09, PRRB09, PG02, Rey03, SB06a, SPB^{+}06, SHK^{+}07, Tan02, TT05, TDL^{+}07, Tsc02, VHG05, WLLY04, WT04a, WK04, WSZS08, ZRCA08]. community-level [Fat07a]. Comparative [Fis09, LEH06, NAC04, Ort08, PK08, Ray08, VLA^{+}06, WBP^{+}07, FWHL06, KTKR08, Kri04a, PE08, RT08, SYCU09]. compare [HC06, WWC^{+}07]. compared [KLM^{+}02, KW00]. Comparing [BKB00, CSM^{+}06, MF02b, OW05, PE06, PFFM07, RPVR03, SOHY07, SSH^{+}07, SF09, YSB08, DFGC04, GSC09, HHP06, LS08b, SPM^{+}08]. Comparison [ETH^{+}04, FRB^{+}05, FYN^{+}07, GPK00, HHD01, HZF07, LJ09, MHM^{+}03, PGA06, PMJ08, SPJB06, TR03, TVK^{+}08, vWBV02, AB03, BBGH07, BKC^{+}07, BF07, CRT07, DDT07, FM08, GDL03, GRBT08, GL01b, Her04c, HUB02, JBT^{+}05, JBP07, KCY^{+}08, LMM^{+}07, Met03, MBS^{+}09, MMF^{+}09b, MMF^{+}09a, MFB^{+}06, MAB01, OJD04, RTB04, SR08, SPS03, SC01, SV03a, SAS06, SY07, SM06, TSBH09, Tud01, WFHP07, War08, XSY^{+}09, ZMB^{+}08, vdBDR02]. Comparisons [CWBR01, OHM^{+}06, TYZ^{+}05]. compartment [BWPC06, GSB^{+}06b, GBG02, HGB04, SGP^{+}06, SGP^{+}07, SPK^{+}09, WBP^{+}07, YWHW02, Zav08]. compartmental [SKP09]. compatible [KGBC03]. compensated [ŠH09b]. compensation [JNSS02]. Compensatory [MJ06, BDLL06]. Competing [SBDD04b, BSR06, DH00, EYB^{+}02, KC03, Uch00]. Competition [CMR09, WR03, AB06, BTPL06, BBGM05, BBGM06, BH00, Bio01, Bio03, CAB08, CNG06, Dam03, EWH^{+}02, FG08, FCP^{+}07, Fie04, GLM02, HS06, Il’08, KAN^{+}09, KM00, LE04, Lid01, MT02, MB02a, MWWM07, PCpDC09, PTS^{+}04, RL05a, RMA08, SRA05, SAR^{+}09, TYS^{+}09, Tsc02, WT04a, ZdlP05, Zha03a, ZLH06, vdHGF09]. competitive [AP06, CMM02, Gra07, HCJ^{+}09, TB08, VPV09, Wit02, Zha03b, vLDHP08]. competitor [GLM02, Gri04]. competitors [LQL05, LLZ^{+}08, Nal00, WZW05, Zha03b]. complementary [HPD09]. Complex [FAA^{+}02, MSHP04, AKMB01, Aum07, BBP08, Bru05, CCM02, DJ05, FLS06, HG07, HDBM02a, HDBM02b, HvI06, Hua03, JWLA00, KC05, KIL^{+}03, LLS^{+}08, MDJ09, PE02, Pet07, Rai01, RSB02, SBB06, SCPC^{+}07, SMSV09, Ski04, TB06a, TA05, VS05, WCC02, WTS^{+}06, Weg00, WLNW08, Wir00, WM02, WD02, ZYY06]. Complexity [RB09, AO00, HL03, Hui07, JF00, LH07, Lin06, MBW09b, Pet04a, PdMVN09, Ric00, RA06, SJLL08, YBM^{+}05, Jor05b]. component [BGL01, ÇDKK05, HK02, Las06b, MB05a, PBB04, PLL04, PBCZ01, PCS01, RBSG06, SHSP04, VA00, WBS^{+}02, WMM^{+}07]. component-based [PBB04, WBS^{+}02]. components [BCD^{+}05, KFR06b, Mon09b, VML^{+}06]. composition [DS01, EBH^{+}01, EFJ^{+}08, FC06, IGP^{+}03, LA07, PE02, dSLSD02, SVB09]. compositions [Pen09, TT05]. compound [KDW^{+}09a]. compounds [BPC07]. comprehensive [CCBB05, DLPN06, SAI09]. compressed [HJZ06]. comprise [Hua03]. compromising [KvKV05]. Computation [GLS07, SB00, JKWJ03, KW00, Rec03, WR01a, WDR06]. Computational [DGOZ04, Li02, PE08, PARH07, SKP09]. computations [ZEGS08]. compute [Lar02]. Computer [Esc05, dSMZ09, MU02]. computing [DBBS03, SBS^{+}06]. concentrating [Eza05]. concentration [CPH00, MAK^{+}04, RV05]. concentrations [AW06b, BJK09, BCCR^{+}02, BHP08, DVJ^{+}08, HFSH06, KPKP06, KS04a, PvdBW^{+}02, PPE^{+}07, STH06, TDdSLS^{+}08, VCD05, ŽŽV^{+}06, ZAM^{+}05]. concept [AFLB09, BKB00, KPS03, RKH^{+}07, VCC03, WXYMZZF03, ŽJDK06]. Concepts [Rai01, Com04, EBR02, GORV06, GKT07, MHKW00, Odu02, Paw00, RHM^{+}05, RHH05, SS06a, YCVA01]. conceptual [AF09, EM08, GSBN03, Hua03, Kin04, Rec03, SLS03]. Conceptualization [JMVvDV02]. concerned [Yun08]. concerning [LS09a]. concerns [CRM09]. conchilega [WGV^{+}08]. conclusions [YLL^{+}05]. condensation [MHSP^{+}06]. condition [IVP08a, IVP08b, KDW^{+}09a, hMzL08, PHH00, SMB^{+}06, SGY01, Wil01, ZAM^{+}05]. conditional [WAB^{+}07]. Conditioning [PCB07]. Conditions [RD07, AMSW07, AN00, BPA08, BNM09, CUS01, CG00, DKRŠS00, DBD^{+}08, ETH^{+}04, FDCH08, GTJ^{+}00, KBGJS06, KVPA07, KBK07, KNZ04, MVZM05, NSO^{+}08, RWM^{+}07, Sil00, SRK06, SSNHP08, TR05, VMR09, YL07, ZAM^{+}03a, ZAM^{+}03b]. conductance [BMD09, WYT09]. Conference [Ano01m, Ano03z, Ano06-48, Ano07a, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09i, DDŽ06, KL08, Leg01, Ano06-45, LMPR06, Ano06-47, Ano09-31, Ano09-32]. Confidence [MLGC03]. configuration [DFM07, HLS06]. configurations [BMB07, GCLG03]. confinement [FG07, Nal00, Nal01]. conflicting [GBSP08]. conflicts [PWH07]. conifer [COB^{+}06, HNF09, RSB09, SRB06, YABM07]. coniferous [CKPP03, GPK00, GAA^{+}05, KW00, MKB00, STH06]. connectance [SI07]. connected [SM09]. connecting [MVZM05]. connection [CYHK04]. Connectivity [Fig09, XIX^{+}08, BBP08, GdBD08, HNF09, PKS08]. consecutive [KCBS00]. Consequences [BPE^{+}07, KDS03, KP00, OzDBS07, PBE^{+}07, ABS05, BFU^{+}09, CM07b, CS01, CK07b, GORJ03, HGD05, HP06, LKR03, LKR06, WAB^{+}07]. consequent [TYT^{+}09]. Conservation [AKB07, DLP09b, BIS09, GS02, HVVK09, JDD^{+}03, LTM^{+}04, Oke04b, RMD04, SC01, Tis06, TNK04, WH04, WTL00, Jør04b]. conservative [Jen05]. conserve [Pra07]. conserving [KE07a]. Considerable [YLL^{+}05]. Consideration [Fuk09, Oga09a]. considerations [KM04, RLFB04, SCR03]. considered [VTB^{+}08]. Considering [VJ06]. consistency [Hen07, NU00]. consistent [EN08, VL08]. conspecific [DH01]. Constance [CWCH01]. constant [LJW03, Mul07, ZKH09]. constants [WLB^{+}05]. constellation [PHP04]. constituents [Alo04]. Constrain [Rd06]. Constrained [WBvDHZ09, ZHW05, BDR06, LHZ^{+}06, SGHG04]. constraint [LS02a]. Constraints [BA05, Wel04, GPP02, LOM06, MCQA08, SBB09, VM09]. construct [ZRR^{+}08]. constructed [AM02, KM04, MLC05, SL06, TAP07, WTMG09]. Constructing [ATDK06, US08]. Construction [CCGMJ07, Log02, BA05, Ber02, FSUH07, HLHZ06, HLxY04, Van08a, ŽJDK06]. consumer [BB01, Mul07, RMF09, SF09, Van08b]. Consumption [Ben06, AO08, BB01, CPH00, Che05, HvG07, Kai00, MLL^{+}05, RLFB04]. contact [APHV08, SM09]. contagion [RG06, SSR07]. contagious [MCB06]. Contaminant [VPSG05, CM08, Mon09a]. contaminants [Mon09b]. contaminated [RIGJM06]. contamination [IF07, Mon02, PWSS07]. contemporary [BWJZ01, Kri04b]. content [BKC^{+}07, ETH^{+}04, JZY07, RSW07, TTPS09]. Contents [Ano01s, Ano01t, Ano01u, Ano01q, Ano01p, Ano01v, Ano01r, Ano01w, Ano01n, Ano01o, Ano02m, Ano02q, Ano02o, Ano02n, Ano02r, Ano02s, Ano02t, Ano02u, Ano02v, Ano02w, Ano02p, Ano02x, Ano03m, Ano03n, Ano03w, Ano03o, Ano03p, Ano03q, Ano03x, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04-92, Ano04-93, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05-66, Ano05-67, Ano05-68, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06v, PCM^{+}03]. context [Bia03, BMR^{+}05, CP04, GL08, IA08, KVL^{+}09]. contiguous [FSE^{+}04a, FSE^{+}04b, FSE^{+}04c]. continental [AS00b, GBB^{+}09, OW08]. continued [Ano05j, Ano06h, Ano06i]. continuing [Ano04w, Hal04]. Continuous [MSB07, GG08a, Mal03, MBD^{+}00, vLLv^{+}07]. continuously [SK04]. Continuum [OM02, ASP^{+}07, MBM00, NF04, TYT^{+}09]. contraception [Zha00]. contract [Don06]. Contradiction [HN09]. contrasted [SM07a]. Contrasting [HCL00, PE07, HLS06, MP07, VCD05]. Contribution [GRW04, VML^{+}06, BT01, CUS01, CT01, GDL03, MRG09, MB06a, Sem08, SHS08, Til04, TRDM06, VS05, WG00, MJ02]. contributions [Ano04-46, BU04, Lim04, OJ02]. Control [Fat04a, GWHS03, HCL06, TKSP09, AW00, AVTP05, AS03, BB08a, BvdW04, BWPC06, BL01, BW04, Bul08, CGY08, CPV07, CFPV08, CN09, CG06c, Fat04b, GSB^{+}06b, HA08b, HRC^{+}07, HLK06, HC06, HP06, Hua08, KWW^{+}09, LKH^{+}08, LTLH08, Loe00, MDGV09, MMRLP06, MAdlPR02, MG09, PCpDC09, PBC02, PGS03, PR07, PPP05, PFFM07, RDS07, SGHM01, SBG06, SGP^{+}06, SGP^{+}07, Sep00, SD08, SC02, Str01a, Str01b, SvdWB^{+}06, TPJ08, TNO^{+}09a, VATP05, WBP^{+}07, XV03, ZSKV05, ZHMN02, dSMZ09, vLLv^{+}07]. control-like [VATP05]. controllable [WKL^{+}03]. controlled [MBP^{+}08]. Controlling [IF07, BA08, KMB08, Nak08]. controls [HY07b, JCX^{+}08, KW00, LPC05, PSPD09]. controversial [SMSV09]. convection [WBH01]. convective [GVDF07]. conventional [CMB^{+}02, FCC^{+}00, GL01b, IVC^{+}08, ITDD09]. convergence [BCS09, CKBH00, HBCL07]. conversion [MBD^{+}00]. Convexity [Log08]. Cook [DD03]. Cool [FSE^{+}04a]. Cool-water [FSE^{+}04a]. Cooperation [HZHL05, IA08, Zha03b]. coordinate [DLPN06]. coordinates [BL02]. Copepod [Roe00, RWW07]. Copepoda [DSD^{+}09, TK08, TSZdRR03]. copepods [DSD^{+}09]. copiotrophic [ZvBS05]. copper [CMG00]. Coral [LS04a, AGNLGSG04, Bar04, CHL08, DDFP07, Dun08, Fig09, GB08, HE09, Lir03, MVMA08, MMV08, MFG^{+}06, MHZ^{+}06, RM09, RP09, SAL07, YSB08, YMD08, AVP08]. corals [BIS09, KVV01]. core [Mog02, YTCV01]. cormorant [Jen01]. Corn [TMS^{+}07, LT04]. Corophium [SKJvdHG06]. correlated [dSDSM08, TN01, WT04b]. correlates [SRR06]. Correlation [LJ07, Bla07a, Hui07, LMMK09, PWSS07]. Correlations [Ond07, Epp00, WLNW08]. correlative [RPVR03]. corridor [Jor00a]. corridors [GGS08, Hoy07, JBT^{+}05, LN08]. Corrigendum [BGF00, CPH00]. Corumbataí [PdC06]. cosine [MB00]. Cosmopolites [VTL^{+}09]. COSMOS [VTL^{+}09]. Cost [FBC08, DFM07, DAR^{+}07, GdBD08, LN08, LL03a, RKH^{+}07]. cost-surface [GdBD08]. Costa [DCG01]. Costanza [Jor05i]. costs [GHP08, LJW03, RP07, WAB^{+}07]. cotton [GP06, GAPE06, SvdWB^{+}06, WLG07]. Cottus [CMA^{+}06, RG05]. cougars [LN08]. Could [Che06, BIS09]. count [BW02, MDB^{+}06, SHN09]. counterpart [Pil04]. counting [OEK^{+}06]. countries [Com04, SOK06, Jor05c]. country [GPSM08]. country-to-country [GPSM08]. County [KTB07, PV04]. Coupled [Roe01, CJS^{+}02, ECBD09, FGB08, GPDF09, GPP02, HBO07, HHB^{+}08, JCB^{+}02, KE07b, MBDB09, MRK^{+}07, MRiI^{+}07, SSKN08, SKK^{+}03, Tan02, TR05, UZ01, VP01, XRM08, ZGF^{+}05, dCBL09]. Coupling [Lar02, RMSS02, SP01, SAR^{+}09, XGL^{+}09, AE02, CHSB00, JZY07, TSJ^{+}09, WHH^{+}08]. CoupModel [NJF^{+}08]. course [GL08, Man00, SL04, vWSH08]. covariance [KW00, MCJB08, SJ06, WKZP04]. Cove [MKS^{+}02]. Cover [Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-37, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano06v, Dam08, HB02b, LLLT08, MBM00, PSCMMNS06, RBC09, RBEZ08, SB06b, SRB06, SS08b, TTA^{+}01, TEMJ06, TYZC05, VZG05, WHZ06, YP02].

Cover-Contents [Ano06v]. Cover/Aims [Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-37, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27]. coverage [ZSZ06]. cowardice [PM06b]. cows [Shi04b]. Cox [LdVA06]. coyote [CEK08, LG03]. coyotes [RMS08]. CQESTR [LGR^{+}09]. crab [PBH^{+}07]. Cracidae [PSCS^{+}01]. cranes [DGM08]. crassifolia [ZNC^{+}06]. Crassostrea [RR01]. crayfish [SSS^{+}09]. CRC [Jor06b]. created [WM00, YLJY03]. Creating [Kin04]. Creation [MAHvD08]. Creek [RBW09, Buz08, TYLL07]. crews [Don06]. crickets [WC03]. crisp [Met03, RR07]. cristatus [TK08]. criteria [ALP03, CFS09, FM08, FLS09, KFR06a, LKP03]. criterion [PBG09, URB06]. Critical [Sch00, Aus07, GPK00, GH07, SvL04, WK04]. criticality [RADD^{+}01, SWBJ01]. Croatia [Ano03z, APJ03, JAN^{+}03, JAB^{+}06, Leg01]. Croatian [BPAB^{+}06]. crocea [HC03a]. crocodile [RMD04]. Crocodylus [RMD04]. Crop [Jor05c, WM06, BvdW04, Che06, CWCH09, FPCA00, FDP^{+}03, GARB09, GGH08, HTA^{+}08, HHC02, Jon07, LBS^{+}06, LFJ^{+}06, LOM06, ME07, MLL^{+}05, MDH^{+}08, PPE^{+}07, PvdBFJ02, PS01, Saa00, SBJ^{+}02, SHP^{+}07, SKvdW09, TS03, TYZ^{+}05, TMD04, WSF^{+}02, XHC^{+}08, XGL^{+}09, ZLZM02, dBWG04, dBWG05, vN02, vdBDR02, MLL^{+}05]. crop-pasture [XGL^{+}09]. Crop-Soil [Jor05c]. cropping [CCGMJ07, HvI01, LR03, LWC^{+}07, SDDC07, TRDM06]. crops [BMS^{+}08, IF07, IVC^{+}08, MTKM^{+}06, PGM08, RS04, TTA^{+}03, WSY^{+}07, vODFS04]. CropSyst [CABD09]. Cross [San07, SG09, CML09, DDH01, HC09, MZWM05, OAAF07, WDBK08, XHC^{+}08]. cross-diffusion [DDH01]. Cross-disciplinary [SG09]. Cross-ecosystem [San07]. cross-pollination [CML09, HC09]. cross-sectional [OAAF07]. cross-validation [MZWM05, WDBK08]. crowding [PHWH^{+}09, Van08b]. crown [GB02, MSL06]. crucial [RF09]. Crustacea [MVZM05, ZPGG03]. cryptic [PBH^{+}07]. Cryptolestes [JJWF07]. Cryptomeria [TYS^{+}09]. cryptophytes [HMBG03]. Cs [DBBS03]. Cu [HvG07, VCD05]. cucumber [DD03]. Cuddington [Nie06]. cues [YK00]. culling [BP07, SC02]. cultivars [FPCA00]. cultivation [CWCH09, PMH00, WBvDHZ09, vN02]. cultural [SDA^{+}03]. culture [BLVC03, CRC08, DMH^{+}03, MLGV00a, MLGV00b]. cumminsii [OI07]. cumulative [MKM^{+}07, MKvdW^{+}09]. cuniculus [FSBD01, SBG06]. cups [SMSV09]. Current [GH09, BHP05, CWCH01, KHT06, MOP^{+}06, Saa00, SRK06, WFHP07]. currents [BVD05]. curve [CG06b, DD02, MKvdW^{+}09]. curves [CG06b]. curzoniae [LZ07]. cuttings [ORS^{+}09]. CVaR [WAB^{+}07]. cyanobacteria [HB06, KNB08b, Ond07, PGG06, KM00]. cyanobacterial [GH07, SAI09]. cyanobacterium [MSL01]. cyber [Fat04c]. cyber-ecosystems [Fat04c]. Cybernetic [Jør06a]. cybernetics [CT07a]. Cycle [SHSP04, URB06, CSG02, Chu08, CCGMJ07, EM08, GS08, HB06, IO02, Ito05, KFS06, LBBC08, Li02, LSS^{+}00, OMRD01, Oli03a, Oli03b, SCPC^{+}07, SBHH07, Svi02, SSH08, TR05, ZPGG03, Zav08, IO02]. cycles [ARF08, Bog04, DDG^{+}05, DDF^{+}05, HM01, IG02, ML04, MML00, MML02]. Cyclic [FH07a, ZLH06]. Cycling [KMT09, ZAM^{+}03b, BP04, BHI^{+}06, CMPO05a, CMPO05b, CMSC01, DR08, KC01, KE07b, KCZ^{+}03, LBL^{+}08, MWWM07, MMN08, Pot04, RŽČ^{+}04, dCSB08, STK00, Van08b, VJ01, WMM^{+}07, YABM07, ZAM^{+}03a, ZAM^{+}07]. Cyclopoida [TSZdRR03]. Cydia [TTJ07]. cypress [HUB02, RBE^{+}08, Gof04]. Cyprinus [BW04]. cyst [MH07]. Czech [NKK^{+}07].

D [Jør04h, UZ01, AYK07, BGMP06, BW01b, BP04, DWD07, GBS00, GNA^{+}06, HY07a, HY07b, JCB^{+}02, KMN^{+}07, MMB07, MBGP08, Mur01, NSO^{+}08, NBT^{+}09, SLGS00, TSF^{+}05, TSJ08, UZ01, WMM^{+}07, ZC00, dlFN08]. D-AEI [ZC00]. D-NEMURO [HY07a]. Dôme [GPDF09]. Dacus [YZS^{+}04]. Daechung [OAL^{+}07]. daily [AFT09, BRK07, Lar02, LG04, PvdBW^{+}02, RLFB04, SA07b, SL04, SSKN08, Tud01, WHP01, ZPW05]. Daisy [BJ02, MMJN03, MTKM^{+}06]. Dakota [HBUS02]. Dall [WMH08]. Daly [BPP09]. Dam [PHW08]. damage [BS04, BAP^{+}06, GPK00, RPRV09, SKP^{+}07]. dams [CDV05]. damselflies [VS08b]. Danger [KTL^{+}05]. Daniel [KMiIW07]. Danjiangkou [LSY^{+}09]. DANUBIA [KFN^{+}08]. Daphnia [BPC07, GR09, GvNK09, KS04b, MB07, PHWH^{+}09, RV05, Van08b]. Daqing [ZH06]. Darwin [JF04]. Data [FL09, GvNK09, MB02b, AN06, Asp02, ATD^{+}06, ATDK08, ABM^{+}06, BW02, BPC04, BHFMG05, BBD^{+}04, BLB08, BLB09, BL02, BCG08, BHV06, Cal05, CCG07, CM03, CCJ07, CLTH08, CMB^{+}02, COB^{+}06, CPB^{+}08, Coh09, DSA08, Dam08, DKL^{+}09, DAR^{+}07, DC08b, DDH^{+}09, Dun08, DT03, EFK^{+}03, EDKF06, ESZ^{+}00, FRB^{+}05, FC06, GFG09, GRC^{+}07, GICB09, GB08, GPC^{+}09, HMPF05, HE05, Håk09, HDH00, HHP06, HB09b, IWM^{+}06, Jen02, JDTD06, JO09, KM06, KKL^{+}06, KMN^{+}07, KK04, LHC07, Las06a, LL07b, LBS^{+}06, LD06b, LLF02, Loe04, MVMA08, MMV08, MP00, MCM^{+}09, MEJ06, MKvdW^{+}09, MXC^{+}04, MCSG06, MJR06, Met03, MLL^{+}05, MCJB08, MDB^{+}06, MMJN03, OJD04, Oli03a, OW06, OSS02, OLB04, OZL07, PLTT05, PDL06, Par00, PCS03, PT08, PBCZ01, PMLM08, Plu00, PCL^{+}05, RH09, RGO^{+}06]. data [RPPB04, RLR05, RSW07, RVRL05, RJS^{+}06, SSS^{+}09, SPB^{+}06, SI06, SA07b, SMG07, SBL03, SPM06, SHN09, SGH^{+}08, SHG^{+}08, SRN05, Sto06, SZL^{+}04, SWBH08, Swa06, SBC07, TPJ08, TYZC05, TD06, VM09, VSGW09, VBD06, VCP09, WLB^{+}04, WO01a, WCH08, WHP01, XSY^{+}09, XBM^{+}08b, YSM00, YTCV01, YKO05, ZLO02, ZWXF05]. Data-directed [GvNK09]. data-driven [TD06]. data-mining [SZL^{+}04]. Database [CWA^{+}09, DWD07, KVL^{+}09]. Database-driven [CWA^{+}09]. dataset [OHM^{+}06]. datasets [Sto06]. dates [GL08]. David [Jør04e]. Dawyck [KWW^{+}03]. day [BCCR^{+}02]. DayCent [HBO07, SHP^{+}07]. DayCent-Chem [HBO07]. DBEM [Mun09]. DBH [LvGC^{+}04]. DC [Jør04h]. Dealing [RLLB09]. DeAngelis [NG07]. Death [GKG05, dSPdBB08a, iT03]. debris [TK01, ZKH09]. DEBtox [BPC07, LPCC05]. Decadal [CPT09, BW01a, JRT02]. decades [HSJ04]. decay [HHD01, KYL07, ZKH09]. decaying [GVDF07]. December [Ano00-37, Ano01-30, Ano02-30, Ano02-42, Ano03-40, Ano04-68, Ano04-61, Ano04-69, Ano05-39, Ano07y, Ano09y]. deciduous [Bir01, GZY^{+}06, OI07, TMM05, WBR08]. Decision [CT07a, Jør08a, KTL^{+}05, LMPR06, RIGJM06, AMRS08, BDD^{+}01, CM04, CKA^{+}02, DDT07, Dei04, DGD03, DGD06, EMdC^{+}01, FG08, HK01, KFN^{+}08, Mat06, MG01, New09, PSC^{+}01, Pet04b, PR07, WZKL09, Wil09, ZVK05]. decision-making [AMRS08, EMdC^{+}01, HK01, Mat06, Pet04b]. decision-support [New09]. decisions [DCG01, GRR04, MBLA03, MG02, RP07, Sti06]. declarative [Vil01]. decline [BRP^{+}06, DFGC04, LMGM^{+}09, MM00]. decomposer [RMR08]. Decomposition [MMJN03, AMB07, CBMP07, CMPO05a, CWWS01, Hua03, LPPS05, MBGP08, PLTT05, PT08, dCSB08, SU08, TTJ^{+}09, XRM08, YWHW02, ZMB^{+}08]. Dedication [Leg03a]. deduction [BdB05, BFS05]. deep [CM09, ET04, GL01a, HBM04, LCLC07, NBT^{+}09, RFdV07, TSJ08]. deep-reservoir [RFdV07]. Deer [PSVH09, CK07b, DDJ^{+}01, FHT^{+}09, GW04, HGL^{+}06, Jen00, JM01, PGLS03, PLP^{+}04, RPRV09, SLL^{+}06, VBRS07, Wal01, XLD01]. DeerKBS [XLD01]. defence [LLA^{+}09b]. defenders [PM06b]. defenses [RJFAA07, RJDF08]. deficient [BB07b]. deficits [EPM^{+}04]. define [HBM04]. defined [CGdPRY09, FBC08]. Defining [ABP05, CDMM08, Fig09, TK07, vdHGF09]. Definition [TF09, GACO04, TPC^{+}07, ZDR03]. defoliation [EB07a, Kai00]. deforestation [ASG^{+}05, MWK07, PR01, SSR07]. degradation [MFB^{+}05]. del [CGH^{+}05, FC05b, FCM05, SBVB05, SCP05]. Delay [SB03, DR08, JJWF07, LWL^{+}02, MSB07, SBL03, ZZZ06]. delayed [TIJ^{+}01]. delays [SB06c]. delineate [LS02b]. delineating [PKG00, ZC00]. Delineation [ASP^{+}07]. delphinid [FBF^{+}06]. Delphinidae [FBF^{+}06]. Delta [MMWMH07, KMPB03, RSM05, JRT02, YLJY03]. DEM-based [AHP01, JAN^{+}03]. demand [BHSR01, CBMP07, Med06, SRB06, SOK06, SB02, SM03]. demersal [Leg08]. Democracy [Jør04f]. Demographic [KD05, LK03, RWW07, VH07, WEW01, ALO^{+}01, CG06a, CG07, EN08, FPS03, GHP08, HGD05, JLAM09, RHHM08, Xia02]. demographics [GP06, GAPE06]. demographies [DFGC04]. demography [CMCD04, Fig09, Liu01]. demonstrated [MCD^{+}08]. demonstrating [Gof04]. Demonstration [EB07a, SGLH04]. dendritic [dlFN08]. Dendroctonus [CRL09]. Deng [Ned09]. denitrification [IF07, LKKL09, VPV09, YHTH08]. denitrifying [VPV09]. Denmark [GZJ06, Kri03, MNEB01, TSJ08, ZM08]. densa [MTD^{+}09]. dense [Che06]. densiflora [KK07, LvGC^{+}04]. densities [BMB07, Gra05, MLGG09, SKK^{+}03]. Density [HG07, MJH02, TNK04, AS03, CL06, EVF^{+}07, FRM09, FAB^{+}07, GBS00, GAB^{+}09, HPHP04, HS06, Her08, HvG07, HTMO06, KBW00, KJB07, LL06, MMM02, MKvdW^{+}09, Mat03a, MSTK08, MJ06, New09, RLR09, Ros09, SRL^{+}00, SJM03, SHS08, TB08, TYZC05, TNO^{+}09b, WPD04, WC03, WPMT07, WMH08, Whi00, Wit02, lXzL02a, YFH03, ZBW05]. Density-dependence [TNK04]. Density-dependent [HG07, AS03, Her08, YFH03, ZBW05]. departure [SRK06]. depend [FWSB05, Her04a]. Dependence [Mii00, AY07, CBSLS07, CL06, GTRF01, HS01, Her04a, LL06, MF02a, MFA07, RPRV09, Ros09, TNK04, TVK^{+}08, WC03, lXzL02a]. dependency [RJR04, Xia00c]. dependent [AS03, AZM^{+}06, DH06, EBM06, HG07, Her08, HMFM00, LKP03, LWJ06, LSAGF05, MCD^{+}08, MKvdW^{+}09, MBM06, MAdlPR02, MJH02, Mon02, NBP05, Nie09b, PF01, SB00, TSZdRR03, Wit02, Xia02, YFH03, ZBW05]. depending [PBC02]. depends [BFE07, MLHT09]. depensation [SvL04]. deposition [BPC04, GVC09, HRJ^{+}00, HNF09, Ito07, JL07, ZAM^{+}07]. depth [ASN02, AHP01, BA08, CM05a, HBM04, JGL07, KHT06, LL07a, Nad07]. derasa [HC03a]. Derivation [MK00, CG06b, LR09]. derive [Bra01a]. derived [Asp02, BFU^{+}09, CWBR01, DGRU06, JKPL09, Las06a, RJS^{+}06, Szi00]. Deroceras [SB03]. describe [CSKP08, SBJ^{+}02]. described [FYN^{+}07, JMN02, Ken02]. describing [Fle01, GBB^{+}06, Sai07, SJG^{+}08, TVK^{+}08, VA00]. Description [BPW^{+}03, CCG07, DMRP07, Jør07b, MRK^{+}07, PTGI09, RE03, AA05, AB05a, DP03, DDF^{+}05, HG07, Haw00, IO02, IWM^{+}06, KPS03, Li00a, LGR^{+}09, SSKN08, SGLH04, WXYMZZF03, WMM^{+}07]. descriptions [KA01]. descriptors [Mit09]. desert [MY02a, SFC05]. Desertification [IMS07, IVP08a, MWP07]. Desierto [SF07]. Design [DLP^{+}09a, Gar04, Jor05j, NH07, PCWP06, SDS^{+}07, SL06, YDS08, ABC03, BWJZ01, DGGD04, ECZ^{+}06, HSMN08, HSC^{+}04, JG08, Jor00a, KM04, LNW00, Mur01, Pet04a, PBD00, RMF09, SM04b, TRDM06, Vil05, YCVA01]. designation [VGK^{+}04]. designed [SFCP02, XBM^{+}08b]. designing [BDD^{+}01]. designs [PLLdCB06, RHB06, Xia00b]. desman [BRV09]. Destabilising [Cha00]. Destabilizing [RG04]. Destruction [KTL^{+}05, Bey03, CL08b, Jør07b, LQL05, Lin05, LLZ^{+}08, NiTT01, NiT04]. destructive [LGL02]. destructor [WS02]. detail [Bor06, FPSJ04]. detailed [Lar02, MSL06, PHBF07, WM00]. detect [HMG06, TA05]. detectable [BF07]. detected [ASI^{+}08]. Detecting [KJ08, WLG07, Ski03]. detection [CTG03, IC02, KKCC06, Loe04, QKR03, Swa06, TCGL03, YLJY03]. detention [KFB09]. determinants [ACE07]. determination [BdB05, BFS05, FMP^{+}00, HPD09, HD01]. Determine [TOS09, Ito07, OEK^{+}06, PWSS07, Rod07, SJ05, SZL^{+}04, SWBH08, TYS^{+}09]. Determining [BBB03, CPP00, Don06, KID^{+}07, KZH07, Sti06]. deterministic [KPKP06, KFR06b, LHC07, MSS02, WRCB01]. Detrital [RMR08]. detritivorous [RS01]. detritus [AMB07, ABS05, HFL07]. develop [GPC^{+}09, MU02, Mon09a]. developed [BHV06, MHvIR00, PF00a, PBG09]. Developing [AMSW07, Jor05c, LMG08, ZBSA07, Aum07, Com04, KFR06a, MRRJ07, PBB04]. Development [Buz08, DVGD07, FPK^{+}07, GBEB06, GW04, GZJ06, LKH^{+}08, LHT^{+}08, MH04, New09, PHWH^{+}09, RAM^{+}03, SMG07, TP06, WSF^{+}00, ZCZ04, Ano03-27, Aok08, AS01, BJJ03, BPE^{+}07, DVdB^{+}08, FL07, GGP03, GL04, HFV03, HDH00, HTS00, HHC02, HTK07, HC05a, HKL07, JBMBP02, JBMBP03, Jør07b, Jør08c, KWS^{+}07, Kir01, KVL^{+}09, LBBC08, LS02a, LLS^{+}08, LYCU09, LP03a, LP03b, NW06, Oga09b, OLK^{+}04, OZL07, PWSY07, PS05, PHBF07, RSB09, SDS^{+}07, Sch00, STH06, SGF08, TIJ^{+}01, TMS^{+}07, WAB^{+}07, WB00, WH07, XZD^{+}06, ZYY09a, dSSGR00]. developments [Bor07, LG00]. Deviation [SI06]. Dexin [Jør06a]. DGVM [SIK07]. Diadegma [TNO^{+}09b]. Diadema [MHZ^{+}06]. diagenesis [KCY^{+}08, SHB04, SHB06]. diagenetic [dG04a]. diagnosis [ADSO08]. Diagonal [Bir06]. diagram [Vil05]. diagrams [Abe04]. dialect [Str09]. dialogue [SMSV09]. diameter [Bra01a, GC02, SDL08, TBPF08]. diamondback [TNO^{+}09b]. Diaphorina [AFLB09]. Diatom [BLVC03, CSR08, SMSR00, TPC^{+}07]. diatoms [HMBG03]. Dicyphus [MDGV09]. die-off [MHZ^{+}06]. diel [Hel08, PRRB09, RJGO00]. difference [BPJM00]. differences [CGR03, MRiI^{+}07]. different [AMSW07, AYKY05, BG01a, BG01b, BSB00, BF07, CKPP03, ETH^{+}04, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, FPCA00, FSM^{+}01, HTS00, KBGJS06, KFB09, KKH^{+}09, MPRJ04, MR06, MR08, MWMN06, MMJN03, Nal01, PLA06, PDS07, TB08, TVK^{+}08, WZJ08, Weg00, WBR07, vWBV02]. differential [DD02, GM05, JZC^{+}07, Nal00]. differently [RGF00]. differing [RAI04, WBTC00, WBN^{+}03]. diffuse [GBEB06, GBG^{+}03, LH01, WTS^{+}06]. diffusers [Mal01]. diffusion [BJJ06, DDH01, ZPO09]. diffusive [LSAGF05]. diffusivities [GVDF07]. digestibility [GL04]. digestion [LJW03, WJL05]. Digital [PARH07, SGY01, Hör03, LJR06]. Dilemma [iTI02]. dilution [HH04]. dimension [JNM^{+}06, TBPF08]. dimension-based [JNM^{+}06]. dimensional [Bio01, Bio03, BDI04, DCI01, FK05, HJZ06, MDH^{+}08, MS00, ORS^{+}09, RŽČ^{+}04, RH04, ROQ^{+}09, RAI04, RWM^{+}07, UZSM05, Zha03a, ZZZ06, ZCB08, ZWXF05]. dimensionality [GS06]. dimensions [Mon05, iWLSN00]. DIN [MRG09]. dinamica [SFCP02]. Dinaric [JPD^{+}06]. dinoflagellate [YK00]. Dioscoreophyllum [OI07]. dioxide [Ale07, KW00, LRJ^{+}09, NK06, OIR^{+}08]. dioxide/temperature [Ale07]. dioxins [CHSB00]. diploid [HPD09]. Diptera [KKCC06]. Dipterocarp [HD00b]. Direct [BPBL00, AMK00, WTS^{+}06]. directed [GvNK09]. directional [BLB08, BLB09]. Directionality [WWRZ04]. directions [WIMK07]. Directive [TSJ08]. DirectScience [Mau07]. disappearance [MTD^{+}09]. discharges [PLB^{+}06]. Disciplinary [Jor05h, SG09]. discontinuous [CKBH00]. discovered [KCJ^{+}07]. discovery [ATDK08, BR01, BHV06, ITDD09, TDLP03]. Discrete [YMT03, BPW00, Den08, Den09, KPS03, MSB07, NS08, Ned09, OVK^{+}06, SDD^{+}04, TK07, TNO^{+}09b]. discretization [KFB09]. discriminate [MCGO05]. discrimination [VPR^{+}09]. Discussion [JT01, RLHD01]. disease [BRW^{+}05, CAG03, DDS^{+}04, FSBD01, FPDP06, Gri08, Mog02, MB09, SDD^{+}04, TM01]. diseases [CMSB07, MCB06, OVK^{+}06, PvdBFJ02, SC02]. Dispersal [BKO08, CDV05, JR05, KW02a, LS01, Mal03, BTPL06, BF07, Bye00, CBFLS09, CS01, DAR^{+}07, EWH^{+}02, GLD07, GG04a, GHP08, HGL^{+}06, HPHP04, JBT^{+}05, JBR07, KMRV07, LN08, LZ07, MDGV09, MT02, MRS05, PBS02, PvdBFJ02, PLLdCB06, PJGW06, SPD^{+}07, SCBG09, SJ05, SJ06, SBJ^{+}02, TMJ04, TTJ07, VM09, VM06, VF07, WWRZ04, WCH08, WLNW08, YMT03, ZPO09]. dispersed [DDLD07, LH05, LH06, WH07]. dispersion [AAA00, CCG07, Ken02, MD07, MTVC05, NK06, SW01, SLT^{+}09]. displacement [Ben04, iWLSN00]. display [Vil05]. disrupting [BRW^{+}05]. dissolution [KNZ04]. Dissolved [WHX^{+}03, AÖÇ05, AG03, BP04, CMG00, DFHP04, FMM^{+}07, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, GML09, HPF08, JLH01, KMM^{+}00, LS06, LCR06, MRG09, MFB^{+}05, SR02, VCD05]. dissonance [Gia04]. Distance [Mun09, Eti04, FK03, FBC08, HC09, MLR08, NBP05, OMA01, SHM07, WRP07]. Distance-based [Mun09]. distance-dependent [NBP05]. distance-independent [MLR08]. distances [BF07, FAB^{+}07, HPHP04, LA04, PBS02, Shi04b, VF07]. distinct [MBP^{+}08]. Distributed [BKS^{+}05b, Fat04b, BFD02, BWPC06, ESG06, Fat04a, GSB^{+}06b, HZF07, JJWF07, Loe00, MP04b, SGP^{+}06, SGP^{+}07, SBL03, Svi02, TMLV07, WBP^{+}07]. distributed-delay [JJWF07]. Distribution [Ano06-46, Lin01, NJB^{+}09, AN06, ALAS09, APJ03, AG03, Asp02, AFLB09, Aus02, Aus07, BPAB^{+}06, BCL^{+}09, BLHB06, BDR01, BSTS^{+}02, BFE07, BSJ^{+}02, CCBB05, CGS08, CGC01, CM05a, CL08a, CSHY08, CWCH09, CZG05, CWL05, DSA08, DGRU06, ECZ^{+}06, EBR02, EFHL05, EN06, FH08, GG02, GGP06, Gri04, GZ00, GKG05, GSC09, HO01, Hör03, HD01, HLxY04, JJ00, JG08, JNM^{+}06, JR05, KFR07, Lar02, LJR06, LR07b, LRT^{+}08, LS06, Mac00, MBDB09, MLHT09, MKRH03, McK01, MCA^{+}06, MSA^{+}03, MF02a, MDB^{+}06, MDVG09, MCK07, Nad07, OW05, OSS02, OPL^{+}09, PF00b, PDV^{+}07, PVS^{+}09, PF01, RM02, SSK^{+}07, SF07, SH09a, SMG07, SBDD04b, SLL^{+}06, SWSF06, SP02, SBHH07, SZL^{+}04, SSdlMP^{+}08, SRR06, SHS08, dJTMBGMP^{+}09, TMS^{+}07, VM09, VTB^{+}08, VSGW09, WCC02, WHHH07, WTMG09]. distribution [Wel04, WW03, Wil08, YHC04, YWL^{+}05, ZLH06, ZJC^{+}07, ZNC^{+}06, dFOV07]. distribution-based [PF01]. distributional [Gra05, GW08]. distributions [ALP03, Aru05, ABM^{+}06, BKO08, BHP05, BKB00, CK07a, CZL05, CUSZ09, GEH02, HS01, HSRD09, HB09b, ICGÁ05, KMN^{+}07, LEH06, LCM^{+}09b, MBM00, MRT05, MMR02, MJR06, Mey04, MEO06, Mon05, PSCS^{+}01, PAS06, PMLM08, RPVR03, SFU01, SBDD04a, TMP06, TVS00, VM09, WÖW05, YPR^{+}05, ZLO02]. Disturbance [MKB00, APEA09, BSR06, CD07, CCC00, GdBD08, KH07, LKR03, LS04a, LK07, Lid01, MCKNM09, MT02, MMR06, MWH00, PBM^{+}05, PG02, RAH07, SGH04, WM08, WYO05, Yos06a]. Disturbances [WKZ03, DMBM05, MDB^{+}02, SGLH04, lXzL02a]. disturbed [CM06a, CM07a]. diurnal [KMM^{+}00, Rob05, WCC02]. diverse [AWL04, CP09, TT05]. diversification [PPP05]. diversion [HNS08, LHB08]. Diversity [LC01, MFG^{+}06, AEP^{+}04, AB06, BBGM05, BBGM06, BPE^{+}07, BCH09, CG09, CN07a, GG08b, HLK06, IP00, KvKV05, LL06, LT06, Man00, ML06, MPP05, MFG^{+}08, MLGG09, PBE^{+}07, RVWH06, Ric00, Ric02, Ric05, RTB04, RMR08, YH04, Yos03, Yos06a, YLJ^{+}01, YLL^{+}05, ZZL^{+}09]. diving [vNL^{+}08]. division [Hel08, IWW08]. DNA [FMP^{+}00]. DNDC [LIS07, MLF^{+}06, NJF^{+}08]. Dnieper [MHP^{+}06]. Do [HE09, KPT^{+}09, LJR06, LG04, PE02, Hua03, KW02a, TPJ08, VIK^{+}08]. DOC [MY05]. documentation [BHL01]. Does [Bor06, FWSB05, PPE^{+}07, SLPP05, WZW05, XCW07, PDBJ09, SHZL09]. dogs [HBUS02]. DOM [MMWMH07]. Domain [DLP^{+}09a, ATD^{+}06, ATDK06, ATDK08, CBD^{+}09, DT03, FF01, GGV^{+}06]. domain-specific [FF01, GGV^{+}06]. domestic [PAdlPS00, SRW05]. dominance [BP02, Pet02, TDdSLS^{+}08]. dominant [MAA^{+}09, SL04]. dominated [BBC03, LH01, LCLC07, PNU03, PNN^{+}08, RBC09]. Dominica [PGCK04]. Doppler [Xia03]. Dordrecht [Jør04c, Jør04d]. dormancy [MVZM05]. dose [BPBF^{+}00, BJFM06]. Dosidicus [RLSZK^{+}08]. Douglas [MBK^{+}03, MLR08, TP06, vWBV02]. Douglas-fir [MBK^{+}03, MLR08, TP06, vWBV02]. Douro [RT08]. DOVE [PARH07]. doves [MGH^{+}05]. down [Ano04-97, Bar00, HY07b, Her04a, KGSB01, MMRLP06, Ort04, Ulg04, vLLv^{+}07]. Downscaling [CMSB07, MBW09b]. downstream [EPM^{+}04]. downy [FK05, RCGB08]. Dr [Cha09]. DRAIN [TMLV07]. Drainage [Ayd08, CGC01, EJG05, KJB07, OPL^{+}09]. drained [TMLV07]. draining [WHX^{+}03]. drawing [GPP02]. Dreissena [vNL^{+}08]. dreissenid [ZCB08]. drift [BDI04, HHK07, RB02]. drift-based [RB02]. drift-feeding [BDI04, HHK07]. drilling [CGH^{+}03]. drive [PBC09]. driven [BVD05, Ben06, CWA^{+}09, CWCH01, EYB^{+}02, MP07, MGSdG07, OZL07, SE04, SSCS06, TD06, WSF^{+}02]. droplet [JDPI07]. Drought [MD06b, BEM00, MVPB02, MBKD02, Zav04]. drought-induced [MVPB02]. dry [IO02, SSNHP08, TYT^{+}09]. dry-matter [IO02]. dryades [MKB00]. drying [MBP^{+}08]. dryland [MM06a]. DSS [CHL08]. dual [Ano03-27, JBMBP02, JBMBP03, LPU^{+}07, Ula09]. Dubrovnik [Ano03z]. duck [MBLA03]. ducks [vNL^{+}08]. duckweed [PWSY07, fPzWhSY07]. due [IMS07, JRBS02, WFM01, Wel04, WGB^{+}08]. duftschmidi [SBL03]. dune [FWS^{+}05]. Dunne [MMV08]. duorarum [ASZRRR08]. duration [BKPS08, HB01, TYT^{+}09, TTA^{+}01]. durations [SKCM07]. durian [OAAF07]. during [DWR09, DP03, EFJ^{+}08, ETH^{+}04, HBRW07, Kan04a, LE04, Lin05, MWK07, MP07, NJF^{+}08, RNC08, VBFM^{+}08, WLG07, YHC04]. Durio [OAAF07]. dust [AW06a]. Dutch [CM04, CM06b]. dwarf [Cai05, FPDP06, RG06]. dwelling [GJ00, JZC^{+}07, KBW00, ŽJDK06]. DyLEM [BP04]. DyLEM-1D [BP04]. Dynamic [DADGA06, GORJ03, GB03, HSMN08, Her04b, KBE^{+}06, OZL07, RJDF08, SSR07, SIK07, SKP09, SSPL^{+}08, TB06b, WH09, Zav08, AMSW07, AKB07, BGF00, BPC04, BK05b, BLDCM06, BH03, CCC04, CSSCC07, CLM^{+}09, CHL08, CRL09, EMdC^{+}01, GDG^{+}00, GQB05, GGB^{+}06, GZ05, GZJ06, HB03a, HGB04, HCJ^{+}09, HvG07, HJ02, HHKH09, HG01, IMS07, JMVvDV02, KCY^{+}08, KCNN06, LKR06, LS08a, LL02b, MD06a, MH03, MEKL08, MLC05, Mat03a, MPC06, MG01, MSW^{+}09, MJ06, Mur01, NG09, Ort08, PGCK04, Pen00, PP04, PR01, PV04, RST05, RR01, RR05a, RPPB04, SB01, SS06b, SDL08, Sti06, SSM^{+}09, Svi02, USK^{+}06, WZ01, Wil03, WR01c, ZJTB03, ZJM04, dSA01, vODFS04, Jør04b]. Dynamical [BDLL06, HL03, OW02, SJLL08, AER^{+}07, HB02b, Hui07, JDPI07, VS08b]. dynamically [PCS01, ZS09]. Dynamics [BBM04, Eza05, HVB06, Jør06a, KVPA07, KVPA08, LCF09, LT04, MAdlPR02, NMC^{+}06, NG07, SO02, VPV09, VATP05, Wan05, AE02, AR02, AVTP05, APEA09, AS00a, ALJL03, ABR05, AB05b, ASP^{+}07, AS01, BBM06, BSB^{+}09, Bal00, BDBS08, BSR06, BBT00, BDE08, BMG08, Ben06, BLC^{+}07, BH00, BB08b, BHC04b, BPC07, Bir01, BPM01, BR01, BRC04, BDJ^{+}09, BW04, BNTK04, BDG01, BPW00, BCDJ00, CC00, CBMP07, CAB08, CAG03, CRC08, Cha02, CDAGK06, CDM^{+}00, CFCP04, CSKP08, CS03, Cha00, CMA^{+}06, CJS^{+}02, CKN^{+}01, CBP^{+}06, CPJ06, CPP00, CKP^{+}01, CEP06, CCLS06, CKPP03, CSdZ09, Coh09, CHW07, CPTD08, CTH^{+}00, CTG04, CN09, CA04, DFHP04, DPL^{+}04, Dei04, DPT09, DAdSC08, DN06, DP07, DFCM01, DJ05, DT03, EFS^{+}03, EDKF06, EBM06, EWH^{+}02, FPDP06, FF01, FCF^{+}04]. dynamics [FC05a, FC05b, Fle01, FBC02, FEP^{+}04, FLS06, FCC^{+}00, FPK^{+}07, GIZ^{+}03, GSG^{+}04, GGS08, GBG02, Gil08, GML09, GP02, GTRF01, GvNK09, GSM08, HK02, HGL^{+}06, HSC^{+}04, HTS00, Her08, HBDA09, HVVK09, HLxY04, HZHL05, HPF08, IWM^{+}06, JM01, JJK^{+}01, JKWJ03, JKJ^{+}08, JLAM09, JLBS09, KYH00, KNB08a, KRK05, KA01, KPH02, KCD^{+}04, KLM^{+}02, KCJ^{+}07, KFN^{+}08, KE07b, KFS06, KNM09, KGZR05, KBM^{+}03, KWW^{+}09, KKZK03, KKH^{+}08, KKH^{+}09, KAN^{+}09, KS07, KDW^{+}09b, LKR03, LGC07, LMH05, LR07a, LS04a, LS09a, LKH^{+}08, LHC09, LG03, Li00a, LS09b, LAB^{+}05, Lir03, LLLT08, LUKD06, LPCC05, MB08, Mac00, MDH^{+}09, MP04a, MPM02, MBMP06, MSM^{+}07, MCBA07, Mar06, MCKNM09, MKvdW^{+}09, MT02, Mat01, MML02, MSHP04, MM06c, McG05, McK00, MAL06, MRK^{+}07, MB02a, MZWM05, MBD^{+}00, MMHE06, MMB07, Moi04, MLGV00a]. dynamics [MLGV00b, MWMN06, MLGG09, MWH00, MH07, Mul07, MTKM^{+}06, MCK07, NDD^{+}07, NW06, NS08, NW07, NK04, NKC04, NHLPA06, Nie09b, NSEDP06, NV03, OWWS01, OVK^{+}06, OT03, OWHS09, PKC^{+}01, PNU03, PBMRE08, Pen00, PJAZ02, PLD^{+}02, PGFM04, PE07, Pet02, PGLS03, PF01, PF03, PBA06, PCM^{+}03, PHW08, PAdlPS00, PCL^{+}05, PBS05, PKS^{+}07, Pot04, PHWH^{+}09, RSF^{+}01, RBSG06, RJGO00, RHE06, RAH07, RSA04, RL05b, RM09, RHH05, RGL^{+}07, RV05, Rob03, RLR05, RE03, RBE^{+}08, RMS08, SR08, SSS^{+}09, SPD^{+}07, Sak09, SB06a, SBG06, SB03, SC07, SU08, SMB^{+}06, SCPC^{+}07, SAR^{+}09, Sem08, SDDC07, SSCS06, SAB03, SWK^{+}05, SRS^{+}03, SRRS04, SMTR07, SFCP02, SSKN08, SPC05, SG05, SHZL09, SSS00, iTI02, TB08, TK08, TEMJ06, TTHH07, TH06b, TK01, TRDM06, TSZdRR03, TS09, TMM05, TA05]. dynamics [Ula09, VS08a, Van08b, VP01, VCP09, VMR09, WTST08, Wal01, WVP08, WM00, WGVB01, WCHM02, WHX^{+}03, WG07b, Wan07, WRB08, WRC^{+}08, WBvDHZ09, WW08, WG04, Wit02, WSZS08, WLZ^{+}09, XV03, XG09, Yea04, YZS^{+}04, Yos03, YTCV01, YCVA01, ZdlP05, ZvBS05, ZJN^{+}06, ZLH06, ZZ08, ZZL^{+}09, ZLX09, ZBZ^{+}09, dGdMPC05, vNLS02, vdPvOV00]. DYRESM [TSJ08].

E. [Ulg04]. early [AMSW07, BC01, KCY^{+}08, Phi04, ZTF08, Ano04w]. earlywood [Mii00]. earth [Che05, Phi02, MRT05, TDLP03, WLSE06]. Earthquake [LLCL04]. earthworm [HvG07]. earthworms [VMR09]. EASS [CCJ07]. east [Ito05, LGS^{+}00, MB02b, RNKG03, WG07a, CZL05, EFJ^{+}08, GGV^{+}09, JCX^{+}08, OW05, OW08, PBY^{+}03, PMLM08, SMW02, SOK03]. east-central [WG07a]. eastern [BCCR^{+}02, BC01, BLBT01, DLK01, EMdC^{+}01, FBF^{+}06, Mii00, OG08, OB04, PdAD^{+}04, RBE^{+}08, SBG06, SOK06, TIJ^{+}01, USK^{+}06, VREA06, ZBL03, FPK^{+}07]. easy [NT07]. easy-to-use [NT07]. ECEM [KL08]. Echinogammarus [MMM02]. ECLPSS [WBS^{+}02]. Eco [ABC03, HKPH08, LWLZ06, WZN05, BKPS08, CB01, CSB03, HHF05, Jør07b, JU09, LZZ^{+}07, LSY^{+}09, MB09, PKC06, PKS^{+}07, SHZL09, WHW^{+}05, XV03, YZC^{+}07, Zha06]. eco-environment [LZZ^{+}07, YZC^{+}07]. Eco-environmental [LWLZ06, LSY^{+}09]. eco-epidemic [SHZL09, XV03]. eco-epidemiological [CB01, CSB03, MB09, PKC06]. eco-exergy [Jør07b, JU09, Zha06]. eco-hydro-climatic [BKPS08]. Eco-hydrological [HKPH08, HHF05, PKS^{+}07, WHW^{+}05]. Eco-Industrial [WZN05]. Eco-sustainable [ABC03]. ECOBAS [BHL01]. Ecodynamics [Pat07]. EcoFL [RCZ^{+}06]. ecohydrological [PDV^{+}07]. ecohydrology [MPP05]. Ecol [Ano04v, BGF00, BG01a, BLB09, CSU04, CPH00, CM07a, CG07, Dun08, Fat09b, Jag09, KL09, LH06, MG09, MLGV00a, RLHD01, San09, Str01a, Suh06, Tia06a]. ecologic [AM02]. ecologic-economic [AM02]. Ecological [Ano03z, Ano03-27, Ano06-45, Ano06-48, Ano07a, Ano08a, Ano09i, Ano09-33, Bar04, CYHK04, Cha09, CC06, CBD^{+}09, DDŽ06, Esp03, FSUH07, For03, HDBM02a, Hey01a, Hey01b, Jør00b, Jør05m, Jor05l, JFGN06, Jør07c, Jør07a, Jør09a, Kan04a, KL08, Kri04b, LMPR06, LBL^{+}08, Leg01, LOL03a, LCY09, MWK07, MBT03, Met02, ÖÖ04, Rey02, SCR03, Swa06, Tie05, Vil06, Xu03, ZYY09a, AN06, AMM04, APG^{+}03, AO08, Ale08, ABS05, ABB06, AO00, Ano04-28, ABR05, ASS^{+}06, Arm04, AMRS08, ACVP00, Aus02, Aus07, BFU^{+}09, BCM^{+}08, BS08b, BCPM09, BJJ03, BLB08, BLB09, BMS^{+}08, BL02, BP03, BSG07, BPE^{+}07, BM08, BMR^{+}05, BMR06, Bru05, CCC04, CSSCC07, CLM^{+}09, CGLS07, CJY^{+}09, CGHW05, CFS09, CWF03, CWWS01, CV01, CP01, DDT07, DC08a, Den08, Dže01, Epp00, Fat04b]. ecological [FC04, FH07a, FK07, FAA^{+}02, FJ05, Fle01, FRZ00, GHP^{+}09, GPDF09, GL09, GKT07, Gau06, GJY07, GL01b, GBA08, zGsLcXwZ09, GICB09, GB03, GJ07, HJZ06, HMFM00, IP00, JW09, JLH01, JKJ06, JOBL08, Jør02b, JVL02, JF06, Jør08c, JLF08, Kaz07, Kir06a, KS08, KGBC03, KCBS00, LPCZ07, LMGM^{+}09, LS02a, Lek07, LKP03, Li00c, LG00, LiWZ00, Lin05, LP03a, LP03b, Lud09, MRRJ07, MM06b, MBDB09, MJ02, MLGC03, MW03, MPC06, MLPK01, McG05, Med06, MSTK08, MGSdG07, MBPS04, Ned09, NT07, Odu02, OM02, OL09, ÖTÖR06, ÖTÖ06, PBE^{+}07, PC07, PCW08, PM09, PSCMMNS06, PPS08, PF03, PCB07, Plu00, PSVH09, PR07, PWSS07, RH09, Rec01a, RIE01, Ric00, Ric02, RA06, SPB^{+}06, SG09, SPK^{+}09, SAB^{+}06, SBB09, SPM06, SGY01, Shi04a, SHN09, SBC^{+}09, Sta07]. ecological [SYC04, Sto06, SPJB06, SF04, TYK03, TPC^{+}07, TK07, TKSP09, TSKP09, TSF^{+}05, TKK07, TA05, TF09, VS05, VBD06, Van04, Van08a, VJ06, Vil01, VF07, WPBM06, WSCR03, WO01a, WXYMZZF03, WEW01, WO01b, WR01a, Whi05, WM02, WD02, WFT05, YLL^{+}05, ZCB08, ZLL05, ZGL08, vNS03, Ano04-45, GBB^{+}09, Jør00b, Bro04a, Jør06a, Jør09b, LMPR06, Nie06]. ecological-niche [BCM^{+}08, GBB^{+}09]. ecologically [KKA^{+}08, RADD^{+}01]. ecologico [CT07a]. ecologico-cybernetics [CT07a]. Ecology [Bro09, Zuc04, BOB07, Cha08, GKT07, GG04b, GZ00, HA08b, JLBS09, KWD^{+}04, Kri04a, LMG08, Liu01, MS07, Nal00, Nal01, PPP05, Ric00, RvdKK^{+}02, Sal02, Suh05, Suh06, Svi00b, Svi08, TLW01, YM02, dlPP05, vPJK04, Ano04-95, Jør04a, Jor05e, Nie00a]. ecology-economical [HA08b]. EcoNet [Kaz07]. Economic [Jor05l, AM02, ABR05, BJJ03, BMS^{+}08, BCS09, CJY^{+}09, CV01, FA03, GGPBEK07, GB03, KvKV05, LHC09, MB06a, MLPK01, MSS02, MBPS04, PGCK04, Pil04, SOK06, SCR03, VFTB06, WFM01, Whi05, XZD^{+}06, ZS09]. economical [CFS09, HA08b]. Economics [MSC09, Ano04-95, Kan04a, Suh05, Suh06, Zuc04, Jør04c, Jør04d]. Economy [Jør04c, Jør04d]. Ecopath [CHSB00, CW04, FM07, HB00b, KNCP04]. ecophysiological [BB08b, KBF^{+}08, vW07]. Ecosim [CW04, WC07]. ecosphere [Pat07]. Ecosystem [AVP08, BK07, CDA08, HY07a, LBS08b, MZASLMLC04, PARH07, PWZ^{+}09, SBC^{+}09, Svi04, TM05a, TR05, YY06, ARL^{+}06, AYK07, AB03, Ano04v, AMSH08, AS00b, ASAC02, ASHHRRPE04, ASZRMH^{+}04, ASZRRR08, BS07, BBM04, BGF00, BP05, BKGC05, Bla07b, BW01b, BVC^{+}01, BMT07, BL04, BCD^{+}05, BMR^{+}05, BPW00, Buz08, CSU03, CSU04, CLBH^{+}09, CCC00, CCJ07, CLTH08, CKN^{+}01, CKL^{+}06, CCM02, CP04, CSM^{+}06, CPTD08, DPT09, DM03, Dow05, DLL^{+}09, DG04b, DFCM01, ECBD09, EGE^{+}08, ECL^{+}02, FL09, FMRS09, FYN^{+}07, FSJ03, FPSJ04, FSJ04a, FSJ04b, GE05, GRW04, GMPC08, GP09, GML06, GvNK09, GRPF07, HHL08, HBO07, HY07b, HB00b, HSJ04, HE09, HCL00, JCB^{+}02, Jør02b, Jør07b, JGL07, KNCP04, KP09, KKW^{+}07, KE07b, KGSB01, Kri04a, KKH^{+}08, KGNK03, KMPB03, KK04, KNB08b]. ecosystem [LAD06, LC07a, Len07, LKP03, LSHG08, LTP06, LS09b, LNW00, LPC05, LCLC07, LLLT08, LT01, LL03b, MDH^{+}09, MRRJ06, MHMHKA04, Man00, MNPJ03, MS07, MXC^{+}04, MK04, MRiI^{+}07, MH05, MBW09b, MB05a, MMF^{+}09b, MMF^{+}09a, MBF09, MD04, MCJB08, MDJ09, MHKW00, MG02, Nie00b, NM09, Nie08c, NI08, OMRD01, OACB09, OB04, PK08, PK01, PNN^{+}08, PR01, PPP05, Pra00, Pra05b, Pra07, QXW02, RBSJ01, Ray08, Rey02, Rey03, RBR05, RMSS02, RBC09, RLSZK^{+}08, RMWW07, RTB04, RAM^{+}03, RMDC04, SPS03, SBA06, SO04, San07, SC04, SIS^{+}07, SSH^{+}07, SAK00, SPK^{+}09, SFM04, SWK^{+}05, SSKN08, SW03, SHP^{+}07, SYCU09, SSS00, SGF08, TN09, iT03, TEMJ06, TDLP03, TSJ^{+}09, TSJ08, Ula09, VCC03, VL08, VHG05, VCAS01, VVF06, WC07, WV05, WM00, WIH^{+}09, WDW00, WKZ03, WIMK07, WGB^{+}08]. ecosystem [WSZS08, XBM^{+}07, XBM^{+}08a, XBM^{+}08b, lXzL02c, XLZ^{+}04, XZZ^{+}05, XZS^{+}07, XGL^{+}09, YHTH08, YYZ06, YYR^{+}07, ZR04, ZYL06, ZYY06, ZC07, ZWXF05, ZCQ^{+}09, Zue07, dlFN08, vW07]. ecosystem-based [AVP08, ASZRRR08, FMRS09, SPS03]. ecosystem-chemical [HBO07]. ecosystem-level [Bla07b, HHL08, MDJ09]. ecosystemic [MK00]. Ecosystems [Ano07l, CWA^{+}09, JPS00, AF05, APHV08, AEP^{+}04, AGNLGSG04, ATDK06, BBC03, BWPC06, BBKN03, BDJ^{+}09, CG09, DFHP04, DPL^{+}04, DLR03, DVPS08, DGD06, DRDD01, DDFP07, DS01, DBBS03, Esp03, FRB^{+}05, Fat04c, Fat07a, FWHL06, FG07, FSJ04a, Gas05, GMPC08, GSB^{+}06b, GSM08, HvI06, HUB02, HTK07, Hua08, IO02, JLH01, KMT09, KB04, KE07a, KCZ^{+}03, KVL^{+}09, KKH^{+}09, LBL^{+}08, LBBC08, mLMfT05, LSS^{+}00, LBBR01, LP03a, LP03b, Lud04, MD06a, MDH^{+}09, ML04, NiT04, NU00, NV04, Ort08, PCW08, PK02, PEAS01, PBV05, Pot04, Pra05a, RS04, RIGJM06, RKS^{+}07, SLPP05, SF09, SS07, SGP^{+}06, SGP^{+}07, SLS03, Sch03, SE04, SLGS00, SS00, SHS00, Str01a, Str01b, TP06, TS09, TKK07, UJF06, VJ01, WMSW09, WBP^{+}07, WBS^{+}02, YFLW06, YTLF08, Yun08, ZBD^{+}09, Zav08]. ecosystems [ZBL07, ZPD06, LBL^{+}08]. ecotone [Pet02]. ECOTOX [Hey01a]. ecotoxicological [CHSB00, CCBB05]. Ecotoxicology [LPCC05, BDBS08, Hey01a, Hey01b]. EcoTroph [GP09]. ecotypes [AGM^{+}08]. Ecuador [Odu04]. Ed [Jør04c, Jor05e, Log02, Nie00a]. eddies [BVD05]. eddy [FRB^{+}05, GVDF07, KW00, MCJB08, WKZP04]. Edge [FWF07, ZC00, BBGH08, FAA^{+}02, VH09]. edges [OMHR06]. Edited [Hey01a, Hey01b, Cha09]. Edition [Jør04e, Xu03, Hey01b]. Editor [Fat09b, Mai08, Fat09a]. Editor-in-Chief [Fat09b, Fat09a]. Editorial [Ano03y, Ano03z, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-37, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Fat09b, Fat09a, ZEGS08, Jør05m, JFGN06, Ano04t, Ano04u, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano09h]. Eds [Jør04b, Jør04d, Jør04a, Jør04f, Jør04h, Jor05b, Jor05c, Jor05d, Jor05g, Jor05i, Jor05j, Jor05l, Jør06a, Jor06b, Nie06, Xu03]. educational [GBSP08]. Edward [Jør04a]. EDYS [CCM02]. eel [LR07a]. Effect [CDAGK06, CBSLS07, CM09, DR08, FGB08, HPA00, NHLPA06, OT03, PCHG03, TM01, Wil01, Yos06a, dSSGR00, AGNLGSG04, BJK09, BDR01, BJH01, BDK01, Bra01b, CGY08, CBP07, CBFLS09, CFCP04, Cha00, CL08b, CWCH09, ČKBB06, DH00, DDH^{+}09, EWH^{+}02, FAA^{+}02, FPSJ04, GSBM08, GBS00, GG05, HvG07, IA08, Jen01, JBT^{+}05, JNM^{+}06, JBR07, KFH07, LHB08, Lin06, LPM^{+}07, MM06c, MK00, MMA02, MBKD02, MGSdG07, NKK^{+}09, OCK01, OHH01, OHH07, Par02, PT02, POF08, PHP04, PBA06, PBV05, PR01, PGG06, RMD04, RV05, Rob03, RLDL09, SH09a, SBB09, SRL^{+}00, SMW02, Sil07, Sti08, SJM03, TC09, THA02, Van08b, VPG05, WLG07, WLNW08, WK04, Xia03, XHH^{+}04, ZHL07, ZLX09]. Effective [BCCR^{+}02, SAB^{+}06, CPV07, GHP^{+}09, GSZ08]. effectiveness [BSTS^{+}02, GJY07]. Effects [APEA09, AMB07, BLHB06, CSCB04, DUH03, ECZ^{+}06, FSJ04b, GLD07, GGPBEK07, GGS08, KLM^{+}02, KMN^{+}07, KKH^{+}09, KM00, LvNMvdB04, LH04, Lir03, MAL06, Mey04, MY05, OIP^{+}08, PJAZ02, PSCS^{+}01, PRRB09, SO04, SSH^{+}07, SU08, SMB^{+}06, SRB06, SNF01, SMR08, SRG^{+}09, SP02, TSBH09, VTB^{+}08, WSZS08, WLJ00, vLDHP08, AAU02, ADdSC06, AA05, ATAK00, BS08a, BKS05a, BEM00, BPC07, BBGH07, BWPC06, BP08, BM08, BCLR04, BMD09, BRW^{+}05, CL08a, CF09, CCGMJ07, CML09, CGH^{+}03, DKRŠS00, DdSG01, Dow05, DS01, DH01, EB06, EVF^{+}07, FCF^{+}04, FWF07, FWHL06, FDP^{+}03, FJ05, FSM^{+}01, GL09, GRW04, Gas05, GSB^{+}06b, GBA08, Gra07, GAA^{+}05, GZY^{+}06, GZJ06, Han09, HB06, HS06, Her04a, JCE06, JNA^{+}09, KDS03, KS08, KBvVK08, KLL01, KGSB01, LLLP00, LC07b, Lin03, LK03, LBS08b, MBM00]. effects [MSHL00, MMM02, MPRJ04, MCKNM09, MOP^{+}05, MT02, MBM06, MFJM04, MLTN06, MSTK08, MLGV00a, MLGV00b, MWMN06, MB02b, MG02, NiT04, NV04, Oke04a, ORS^{+}09, OFK08, PCW08, PARH07, PA09b, PP04, PG02, Pot04, QXW02, RJDF08, Rob05, RCZ^{+}06, SPTP01, SGP^{+}06, SGP^{+}07, SL01, SDDC07, SHW04, SOK03, SOK06, SKT00, SM06, SGHG04, SF04, TH06a, TFF07, TN06, VJ06, VH09, WPD04, WNW09, WC03, WRCB01, Wal01, WZ01, WFM01, WBB05, WÖW05, WBP^{+}07, WL04, WG04, WBN^{+}03, WBS^{+}02, YZS^{+}04, YLL^{+}05, ZSKV05, ZWCL05, ZBWR06, ZC00, vNL^{+}08, vODFS04, vdPvOV00, Mon09a]. efficacy [PPP05]. efficiency [DGSBG09, LJW03, MY05, MLL^{+}05, MB07, Oga09b, Par02, SBB09, TYZ^{+}05, VCC03, VPV09, XZS^{+}07, ZDR03, vODFS04]. Efficient [HvI06, PGW00, WLSE06]. Efficiently [KYL07]. effort [Jen02, MCA^{+}06, OG08, ZR04]. efforts [Eza05, MDB^{+}06]. EFIMOD [CKL^{+}06, KCZ^{+}03]. Egeria [MTD^{+}09]. Egg [JZC^{+}07, DSD^{+}09, SBL03, ZPP06]. EHIM [XZZ^{+}05]. eigenvector [Mun09]. EKF [PCS03]. Ekko [Jør04d]. ELAM [GNA^{+}06]. elaphus [DDJ^{+}01]. elasticity [ÅSCP09, HMF00, Mul07]. Elbe [CSR08, SCB^{+}09]. electric [AHKB01]. electrical [Kan04b]. Electricity [AO08]. electromagnetic [BDJ^{+}09]. Electronic [Hey01b]. element [ARF08, DAdSC08]. elementary [Wil03]. Elements [PGS03, KCZ^{+}03]. elevated [Gut07]. elevation [BSM08, Hör03, LJR06, MOP^{+}05, MOP^{+}06]. elevational [HHL08]. ELFSim [LPM^{+}07]. Elgar [Jør04a]. Eliciting [YEMZ03]. elk [FJ05, WG04]. ellerii [LK03]. elliptica [MKS^{+}02]. ELMO [BW01b]. elongata [Har08]. elongatus [SKCM07]. Elsevier [Hey01a, Hey01b, Nie00a, Nie06, Xu03]. Elster [WR08]. elucidation [JJK^{+}01]. elvers [GM05]. Elysia [CTH^{+}00]. embayment [BGF00, Sin07, WYMS07]. embedded [Nie00b]. embedding [KŞE04]. embodied [Rec03]. embryo [WT04b]. Emerald [BMF^{+}06]. emerge [YMD08]. Emergence [DH06, JT09, ACVP00, BMR06, JR05, LL07a, LPFL09, MDPC06, Mur06, SDDC07]. Emergent [BMR^{+}05, Esc05, ARF08, CCMT09, HB05, RHM^{+}05, Reu05, VPG05]. emerges [MB05a]. Emerging [VS08a, JPS00]. Emergy [CdQO06, CR06, JOB04, LR03, LYCU09, Odu04, WZN05, ADSO08, BM00, BU04, CC06, HB04, LD06a, SL06, SYCU09, TB06b, ZYY09a, ZYY09b, Her04c, Ano04-96, YLSH03]. Emergy-based [LYCU09]. emigration [FHT^{+}09, HP06, JM04]. Emiliania [BSM08]. emission [AAKO^{+}08, CGY08, CGdPRY09, GBG^{+}03, LIS07, MSHL00, OEK^{+}06, Ski03, XJM07]. emissions [AW06a, BH06, GSZ08, LL02a, NK06, NKK^{+}09, NJF^{+}08, PBA05, PGS03, RMDC04, SDB03, SHS08, vWSH08]. emoryi [dJTMBGMP^{+}09]. emphasis [FK03, LBL^{+}08, NSEDP06, YMD08]. emphasising [TFTO07]. emphasizing [RJGO00]. Empirical [HB02b, WWL^{+}05, BB04, BKC^{+}07, BBKN03, CMD06, DAR^{+}07, GFG09, GRBT08, HMPF05, HE05, LCM^{+}09a, MH02, RR06, RBR06, SFU01, SSNB07, VCD05, BLC^{+}07]. empirically [Bra01a, Lin01, OBE^{+}07]. empiricism [RBB05]. EmSim [Mau07]. Emternalities [Pil04]. emulation [PWC^{+}09]. encaustus [BMBOCR03]. encephalitis [USK^{+}06]. Enchytraeid [Nie09b]. enclosed [ATAK00, GSM08, HE09, KSH^{+}03, MSTK08, Mur01]. encounter [FBF^{+}06, LBS08a, RJFAA07]. encounters [CUSZ09]. encrasicolus [KL07, KL09, MSM^{+}07, San09]. End [Jør04g, LS09b, GLS02]. End-to-End [LS09b]. end-user [GLS02]. endangered [GS02, PGLS03, PWH07, RAH07, RBR06, SGF09]. ended [TF05, TSF07]. endemics [HB09b]. ENDLESS [ABC03]. endocrine [BRW^{+}05]. endogenous [GBA08]. enemies [ZS09]. enemy [GAPE06]. Energetic [HKL07, BCS09]. energetics [HC05a, Pet04b, SPK^{+}09]. Energy [BOJ04, BU04, Her04c, Jør04c, Jør04d, MC04, SBB09, VCAS01, YABM07, ABC03, ABC04, BPP09, Bog04, Bro04b, Dei04, ECP08, FRB^{+}05, FH07a, FGB08, GAA^{+}05, GZY^{+}06, HHK07, HvG07, Hua03, HBCL07, Hua08, KB04, KCNN06, KWD^{+}04, LC08, LBBR01, MK04, Med06, MB05b, MKS^{+}02, MZASLMLC04, Odu02, OSHO09, Pet04a, RR01, RR05a, SC04, STK00, Suh05, Suh06, Til04, VCC03, VA00, WKZP04, WLJ00, ZGF^{+}05, ZZZ06, Zuc04]. energy-matter [STK00]. ENFA [HSRD09]. engineering [Ano04-45, Arm04, GKT07]. engineers [CH04]. England [Jor05a, SMW02]. English [AMSH08, SM07a, VBFM^{+}08]. Engraulis [KL07, KL09, MSM^{+}07, San09]. enhance [GLS02, HLS06]. Enhanced [Ano09i, GORJ03, LKKL09]. enhancement [IiTS^{+}04, vBBE^{+}08]. Enhancing [KM04]. Eniwetok [Bar04]. Enlightenment [MC04]. enriched [dG04a]. enrichment [EPS04, JKPL09, vdPvOV00]. ensemble [MCJB08, PVS^{+}09]. ensembles [KDW^{+}09a]. ENSO [Gan06]. entities [Jua09, SS07]. entomopathogenic [SGHM01]. Entrée [GCG^{+}07]. Entropic [CG09, TP08]. Entropy [Aok08, SS00, Lud04, MRT05, PAS06, Ric02, SSdlMP^{+}08, ZYL06, LP03a]. entropy-based [MRT05]. ENTS [WLSE06]. Envelope [PDBH02, PMLM08]. Environ [BO07, GSB^{+}06b, SGP^{+}06, SGP^{+}07, SKP09, WBP^{+}07]. Environment [ÅSCP09, Jør04c, Jør04d, Jør04h, ALJL03, Bia03, Bio01, Bio03, DMBO00, Del04, DMBM05, Eza05, GBA08, Hui06, Jen05, Kin04, LKR06, LZZ^{+}07, MPM02, Mar06, MEJ06, MMPT06, ME07, Mon09b, MPS02, OzDBS07, Pat07, QMH00, RSB02, SdH02, TMJ04, WBS^{+}02, WBR07, YZC^{+}07, Zha03a, dlS07, Ano04-94]. Environment-specific [ÅSCP09]. Environmental [Ano08a, DŽD06, GLD07, Jor05f, LBL^{+}08, RVWH06, RR05a, VTB^{+}08, AMD^{+}03, AHP01, ABC03, ABC04, AMRS08, BVD05, BCL^{+}09, BRV09, BS08a, BM00, BNM09, BSB00, BP02, BCLR04, BCS09, BCH09, CUS01, CSCB04, CM09, DKRŠS00, Dam03, Dam08, Dow05, Dun08, ES01, ESG06, Fat04a, FDCH08, Fis09, GPP02, GB08, Han09, Hee02, HLK06, HD01, HLxY04, IKS09, Kar04, KW00, KC05, KFR06b, Kri04a, KGNK03, LG04, LHK00, LWLZ06, LSY^{+}09, MVZM05, MI01, MVMA08, MMV08, MP00, MEJ06, MFG^{+}06, MHKW00, NG09, NCM07, NS07, Ond07, OL09, PCpDC09, PK08, PBB04, PCW08, PGCK04, PT08, PG04, Pet07, PSVH09, PHH00, PKS^{+}07, PLC08, PKG00, QKR03, RIE01, RBR05, RWW07, RJS^{+}06, Sav00, SGHM01, Sch00, SVS04, SMSV09, SOK03, SOK06, Sil00, SL06, SRN05, Sti08, Sto06]. environmental [SZL^{+}04, SRR06, SBC07, TB06b, TKHS^{+}07, dJTMBGMP^{+}09, TOS09, Uus07, VGK^{+}04, VML^{+}06, WA02, lXzL02b, YM05, YSB08, YTT^{+}09, ZSD^{+}08, ZEGS08, vWSH08, Jor05b, Jor05d]. environmentally [DS02]. environments [AYKY05, BC03, FCF^{+}04, GLP05, IWW08, Jen02, LW09, MBP^{+}08, SGHG04, SSNHP08, VCD05, WJM^{+}03]. environs [BBP^{+}07]. Enzo [Jør04g]. epibenthic [DJ05]. EPIC [BRGS09, IWM^{+}06, TS03, dBWG04, dBWG05]. epidemic [AZM^{+}06, CG06c, LG03, MOJ01, SKT00, SHZL09, XV03]. epidemics [PvdBW^{+}02, ZHMN02]. Epidemiological [HXP^{+}09, CB01, CSB03, GGV^{+}09, GH00, MB09, PKC06, SM09]. epidemiology [DDS^{+}04, MSB07, SDD^{+}04, VTL^{+}09]. Epinephelus [ALAS09]. epiphytes [dSA01]. episode [BKPS08, EFJ^{+}08]. episodic [BGMP06, Lid01]. epistemic [EBR02]. equal [Wil05]. equation [ASS^{+}06, GSB05, ITDD09, JZC^{+}07, OSHO09, TDLP03, WFB^{+}08, WGS^{+}02, Xia05a, Xia05b, XBM^{+}07, dlS07]. equations [ALM00, BdB05, BPC04, BFS05, GFG09, GM05, ORF01, TLW01, Tud01, Xia05a, Xia06, Xia07, ZBSA07, vODFS04]. Equilibrium [WIH^{+}09, CG09, CKBH00, GLS07, HBO07, HZHL05, Hui06, KTB07, LC01, hMzL08, MKRH03, RHH05, Wel04, WÖW05, YWHW02, YFH03]. equilibriums [MCBA07]. equine [USK^{+}06]. Equivalence [BBP^{+}07, RF04, YTT^{+}09]. equivalent [PNU03]. Eragrostis [MCSG06]. Ergodicity [KK08]. Erie [MWH00, ZCB08]. Erik [Cha09]. Erinaceus [DAR^{+}07]. Eritrea [TN08]. Erken [EPTB07]. erosion [GSBM08, JL07, LTLH08, RSC09, RB02, ZSZ06]. erosivity [WGS^{+}02]. Erratum [Ano03-27, Ano04v, BG01a, BLB09, CSU04, CM07a, CG07, Fat09b, For03, HDBM02a, Jag09, LH06, LOL03a, MG09, Met02, MLGV00a, Str01a, Suh06, Tia06a]. error [BS08a, EN08, JG08, LBBC08, LH04, PGW00, PT02, PBML08, TLW01, TW02]. error-in-variable [TLW01, TW02]. errors [HHC02, KC05, SCAP05, SWBH08]. ERSEM [KE07b]. eruptive [BPM01]. Erzgebirge [BAP^{+}03]. Escape [BM07, TW00, dJBPG02]. espresso [SMSV09]. ESS [XCW07]. Essence [Jør04g]. essential [SGH^{+}08, VGK^{+}04]. essentials [HZ01]. establish [Aru05]. established [KGBC03]. establishing [CBD^{+}09]. establishment [GWK^{+}06, Gra04, LMMK09, SMET06]. Esteros [CGH^{+}05, FC05b, FCM05, SBVB05, SCP05]. Estimate [EJG05, AFLB09, BG08, CM05a, Che06, DRDD01, FMP^{+}00, Håk09, KKA^{+}08, KK04, LBBC08, NWP06, Nad07, Saa00, SSKN08, ZCB08]. Estimated [MP08, Hee02, MLGC03, WHP01]. Estimates [HHL08, TTJ^{+}09, WCS00, BS08a, CSF^{+}04, CK07b, Gra04, HTS00, MBF09, SMR08, TYZ^{+}05, TMS^{+}07, WLG07, XBM^{+}08a, YBM^{+}05]. Estimating [BDK01, BCG08, DGM08, GWK^{+}06, GC02, JAN^{+}03, KPH02, Las06a, LA04, RHB04, SW01, TR05, VHG05, WW09, vBBE^{+}08, AYKY05, BPC04, BH06, DCP^{+}07, EG09, GW04, LLLT08, OBE^{+}07, RLFB04, WO01b, Xia00b, XSY^{+}09, ZXC03, ZWXF05]. Estimation [BJK09, CPB^{+}08, Li02, MLH05, MXC^{+}04, MSL06, TM05b, VCD05, WTS^{+}06, YSM00, YSG^{+}06, ABV^{+}06, BCPM09, CA04, GP07, HTS^{+}07, HGR08, JJWS08, Jor01c, KC01, KRN04, NBP05, PT02, PS09, RBR06, SSK^{+}07, SAH03, SJ06, SBL03, She06, SOK03, SDB03, TS03, TW02, VCRD06, WNW09, Wan07, WLB^{+}05, WFT05, YHTH08]. Estimator [PBHGF07, AHP01]. estimators [EB06, PMC08, WPMT07, WTMY07]. estuaries [BHSR01]. estuarine [BdB05, BFS03, BFS05, DR08, FWSB05, RS01, RWW07, RLLB09, She06, TSZdRR03, VHG05]. Estuary [BB01, BWPC06, GSB^{+}06b, MMM02, SGP^{+}06, SGP^{+}07, ZCZ04, CBD^{+}09, DSD^{+}09, JMN02, LXJ^{+}03, MRG09, MBCS07, PM06a, RH04, SSPL^{+}08, SSKN08, TSF^{+}05, WBP^{+}07, dSA01, DLL^{+}09, LXP^{+}08, MHP^{+}06]. Ethiopia [FM07]. ethological [GGP03]. Eucalypt [PWH07]. Eucalyptus [CMPO05b, MLF^{+}06, MBS^{+}09, MB02b]. Euler [KE07b]. Eulerian [ATA03, CUF^{+}09, GNA^{+}06]. Eulogy [BHJ04]. euphotic [RFdV07]. euphratica [RSM05]. EUR [Nie06]. Eurasian [SSNB07]. euro [Jør04c, Jør04d]. europaeus [DAR^{+}07]. Europe [SM07a, AAKO^{+}08, DFGC04, EFJ^{+}08, HTA^{+}08, LGS03, SSPL^{+}08, SGH^{+}08]. European [Ano03z, Ano06-47, Ano06-48, DDŽ06, KL08, Leg01, APHV08, BCG08, FSBD01, GBB^{+}09, HB01, KS08, KBF^{+}08, LR07a, MSA^{+}03, MSP^{+}08, OHH07, RBEZ08, SBG06, TC09, THJ^{+}03, TMS^{+}07, WFB^{+}08, WSZS08]. Eurytemora [DSD^{+}09]. eutrophic [BLVC03, BP04, CBS09, MI01, SP01, SSS00]. Eutrophication [AB05a, AB05b, AQS^{+}07, MH04, ATAK00, BSR04, CM03, DWR09, DCI01, Håk00, Håk09, Jan01, JRBS02, JRT02, KSÇ^{+}00, KHLS07, LCM^{+}09a, MNPJ03, NV04, PM06a, RdQFO07, RAM^{+}03, She06, WRCB01, XTDL01, ZJM04]. eutrophied [Sri04]. evaluate [CMA^{+}06, ECP08, GBS00, PPGP08, RV05, RdQFO07, THA02]. evaluated [LBS08b, MWWZ01]. Evaluating [AM02, ALP03, AS09, BTdFC05, BVNS02, CEK08, DAR^{+}07, GFGZ08, HLH^{+}06, JBT^{+}05, KC05, MDB^{+}02, PF00a, Pra09, RHB06, RHHM08, RMD04, SK01, SB07, YLSH03, DKL^{+}09, GCLG03, HHP06, Las06b, LCG^{+}09, LPM^{+}07, MP04a, MSD06, Oke04b, Pra05b, SRS^{+}03, Tud01, Xia00b, XZD^{+}06, XWB04, ZC07]. Evaluation [ABM^{+}06, BBD^{+}04, DC08a, DCP^{+}07, EVF^{+}07, HHB^{+}08, IKS09, KRN04, MAA^{+}09, NSO^{+}08, PDBH02, PMD^{+}09, RMWW07, SAL07, TN09, TKHS^{+}07, TNO^{+}09b, VBRS07, ZYY09b, Bar00, BvdW04, CCG07, CJ07, CCGMJ07, CABD09, CR06, EFHL05, ES01, Gar04, GW04, HHF05, KBE^{+}06, LR03, LWLZ06, LYCU09, McK01, Mon09b, PLL04, PF00b, Pra07, Pra08, PBD00, RSC09, RHB05, SV03a, SBVB05, SL06, SHK^{+}07, VPSG05, WZN05, WBR08, WSF^{+}00, YZC^{+}07, ZYY06, dSSGR00]. Evapoclimatonomy [Bey03]. Evaporation [Ayd08, AYKY05, HTS^{+}07]. evapotranspiration [OIP^{+}08]. evasion [zLGH^{+}05]. Event [EYB^{+}02, IC02, MTD^{+}09, NJF^{+}08, RNC08, RH04]. Event-driven [EYB^{+}02]. events [DDLD07, GWK^{+}06, Ond07, RF09, ZTF08]. Eventual [EM08]. Everglades [BCDJ00, DRDD01, GDG^{+}00, HUB02, RSF^{+}01, RMD04, TDdSLS^{+}08]. evergreen [ZBZ^{+}09]. Evidence [RCL06, Leg08, MMRLP06, SM04a, TTHH07]. Evolution [GHP08, Jør07d, BTPL06, BRK07, DMBO00, GPDF09, HZHL05, IA08, JWLA00, LJ07, SVS04, SSH08, Wal04, Yos06b]. Evolutionary [Wit02, Yos03, Yos08, ZZL^{+}09, AFTB07, BR01, BDJ^{+}09, JKWJ03, KCJ^{+}07, McG05, Rec03, Sel00, TKK07, WR01a, WDR06, Yos06a]. evolutions [CMCD04]. evolving [SVS04]. Exact [Mar06]. Exactly [Jor06b]. Examination [JF06, PWH07, ZJM04, Mat03a]. examine [BKS05a, Paw00].

Examining [AP06, BLC^{+}07, ES06, MFJM04, CB07c, MMR06]. example [BS08b, BBT06, DRDD01, FJ05, FLS06, GL09, HTS^{+}07, HL04, JLAM09, OMRD01, PT02, RNKG03, SBL03, TCD02, TKTB07, VPR^{+}09, WKL^{+}03]. examples [Håk00, KA01]. exceptional [IC02]. excess [SOK06]. exchange [ARL^{+}06, BW01a, DLG06, FRB^{+}05, FPCA00, FGB08, GPSM08, GS08, GAA^{+}05, GZY^{+}06, Kir06a, KBE^{+}06, LRJ^{+}09, MBF09, MWD05, OIR^{+}08, OT03, PHP04, PTGI09, RMSS02, Sin07, SPM^{+}08, Tan02, WBTC00, Zav04, ZPD^{+}08, vWBV02]. exchanges [VSFM03, YABM07, ZGF^{+}05]. exclusion [PBV05]. excretion [GZ05, TFM01]. exemplified [DS01, Håk09]. exercise [EIRT00, Sal02, Wir00]. Exergetic [CJY^{+}09]. Exergy [Che05, CCY06, FC04, LTP06, Lud09, SS01, SSP03, UJF06, VFTB06, CJ07, CQ07, Deb02, FMP^{+}00, Jor01c, JVL02, JMN02, JOB04, JLDM05, Jør07b, Jør07d, JU09, LP03b, LJ09, PKC^{+}01, PLS^{+}06, RBSJ01, SPJB06, LP03b, Zha06]. Exergy-based [CCY06]. Exergy-economic [VFTB06]. exhaust [NK06]. exist [XCW07]. Existence [HM01, IVC^{+}08, NWH^{+}06]. existing [SBDD04a]. exogenous [LLL^{+}07]. exotic [AP06, SMET06, VML^{+}06]. exotics [VIK^{+}08]. expand [MLPK01]. Expansion [KGSB01, LvNMvdB04, LPP^{+}07, SHS08, iT03]. expected [Ben04, CWF03, iWLSN00]. experience [PDBJ09]. Experiences [MHP^{+}06]. experiment [AFT09, BOB07, DDH^{+}09, ESZ^{+}00, GvNK09, JCB^{+}02, PLTT05, PSPD09, SB00, Tud01]. Experimental [Lug04, SS08a, WCHM02, WHZ06, BFHR05, GVC09, HC03b, JZC^{+}07, MMJN03, Par00, TSBH09, Vil05, VCM07, WCH08, WM06, Xia00b]. Experiments [FCP^{+}07, BJK09, CS03, Esc05, KH02, LH04, LGR^{+}09, LS08b, MMV^{+}09, Nal00, Nal01, SKCM07]. Expert [Ano02y, ATDK08, Mac00, Met01, Met02, VBD06, YEMZ03, YSM^{+}06]. explain [LL02a, LA07, POF08, Str09]. explained [NJB^{+}09, Shi04b, vLLv^{+}07]. explaining [AOY02]. Explanation [Jør02b, JVL02, BR01, YJJG09]. Explanations [Odu02]. explanatory [LKP03]. explicit [ASG^{+}05, ASJD01, ACJT08, BS08b, BHW^{+}08, Ber02, BH00, BM07, CEK08, DC08b, DFCM01, ECHN08, FAB^{+}07, GH09, Gri04, Her04a, HD01, JBT^{+}05, Kin05, KJ08, LC07a, LC04, MP04a, MCKNM09, MSD06, NH07, PEM06, PGFM04, PE07, PBY^{+}03, PdAD^{+}04, PHD04, PWSS07, RHB06, RL05a, RHB04, SIK07, SBM04, SV02, SRG^{+}09, SVB09, SRR06, SSNHP08, SHM07, TFF07, VTL^{+}09, WBS^{+}02, WD02, XHH^{+}05, YP02, ZHMN02, vNSvdBC03]. exploitation [Aub04b, BFD02, CPT09, FH07b, OCK01]. exploited [AS00b, ASAC02, ASZRMH^{+}04, DDFP07, PNN^{+}08, SPS03, SSH^{+}07, Wit02]. Exploration [OCK01, ZRASC04, CGH^{+}03, MHvIR00, RH09]. Exploratory [SBA06, Jan01]. explore [KBM^{+}03, KWW^{+}09, OL09, SGF08]. Exploring [ACE07, AMSH08, ASS^{+}06, ASZRRR08, GDP09, LC04, MBKD02, RSF^{+}01, TN08, WDR06, HDBM02a, HDBM02b, MHMHKA04]. exponential [Ken02]. export [CR06]. exposed [CBS09, KK07]. exposure [BJK09, CPV07, CFPV08, CGHW05, CBSLS07, KNB08a, SR02, VBD06, vWSH08]. expressed [FSJ03]. expression [WLLY04]. Extended [Sak07, KDK06, MEJ06, WBvDHZ09, PCS03]. extension [MWD05, PvdBFJ02, SSM^{+}09, XBM^{+}07]. extensive [Gri08]. extent [BDF^{+}06, LM07b, MM06a, MBKD02]. external [BHP08, lXzL02a]. externalities [Pil04, SCP05]. Extinction [LLZ^{+}08, WJM^{+}03, Bra01b, GG00, HMG06, IiTS^{+}04, Jag01, Jag09, JKJ06, KP00, LQL05, Mat01, NiTT01, YJJG09]. extinctions [ABB06]. extirpations [PSCMMNS06]. extract [DWD07]. extraction [HB02a]. extrapolation [KE07a]. extremal [Gau06]. extreme [Leg08]. extrinsic [DMBM05].

F [Bro09, JT01, Jor05g, RLHD01]. F.A.L.C.A.D.E. [ABC04]. face [SGF09]. facilitate [CSSCC07, WZW05]. facilitates [Yos06b]. facilitating [ZDR03]. facilitation [MML00, MML02, ZdlP05]. facilitative [vLDHP08]. facilities [HB02a]. Factor [GPV08, GBB^{+}09, BCM^{+}08, EKBF04, Jua09, Vil05, WGS^{+}02, YH04, vBBE^{+}08]. Factors [DDLD07, Nak08, VPG07, APJ03, BP02, BM08, BAP^{+}06, DSD^{+}08, FMP^{+}00, IGP^{+}03, KMB08, LA07, LG04, MDH^{+}09, Mii00, MVV07, OAL^{+}07, Ond07, PvdBW^{+}02, Phi02, PPP05, RG04, SF07, SGHM01, SZL^{+}04, SBC07, TCGL03, dJTMBGMP^{+}09, VML^{+}06, WFM01]. facultative [SMM^{+}02]. Fagaceae [dJTMBGMP^{+}09]. Fagus [Kno03, TBPF08]. FAHP [LSY^{+}09]. failures [JG05, Xia04]. fall [Jør04g]. Fallopia [SWCO07]. fallow [MBD^{+}00, vN02]. fallows [SL04]. falls [Rob03]. family [ADSO08, EB06, FBF^{+}06, KA01, MML02]. far [VSGW09]. Farfantepenaeus [ASZRRR08]. farm [Kri03]. farming [AKB07, CMDP^{+}00, MBLA03, MG05, MTKM^{+}06, RMGR09]. farmland [TDL^{+}07]. farms [ADSO08, CdQO06, Mur06]. Faroe [ZR04]. Faso [KRvL^{+}02, VOM06]. fast [HD00a, RG04]. Fate [SBHH07, AF05, BG08, CHSB00, CWH^{+}00, HSV07, KGNK03, MH02, MBCS07, PCW08, RSF^{+}01, dG04a]. Fath [Bro09]. fault [AF09]. fauna [CWF03, FDCH08, HR03]. favourability [BRV09]. Fear [LXP^{+}08]. feasibility [IVC^{+}08]. feasible [HC09]. features [DJ05, MLM06, VM06]. February [Ano00-34, Ano00-44, Ano01y, Ano01-38, Ano02-29, Ano02-41, Ano03-45, Ano04-66, Ano04-47, Ano05-43, Ano05-28, Ano06-36, Ano06-28, Ano06-37, Ano07-27, Ano08y, Ano08n, Ano08-27, Ano09x, Ano09z]. fecundity [Han09]. federal [Don06]. feed [Cor05, MD06b, MLGV00a, MLGV00b]. feed-forward [Cor05, MD06b]. Feedback [LCH^{+}00, Ale07, BB08a, EM08, MKvdW^{+}09, PSPD09]. feeders [KS04b, SM04a]. feedforward [PSPD09]. Feeding [KS04b, BDI04, DGSBG09, GH00, HHK07, JVL02, LS08b, LCLC07, MLGV00a, MLGV00b, SHG^{+}08, SO02, Tud01, WC07, ZDR03, ZJG^{+}06]. FEFLOW [ZWCL05]. female [VS08b]. female-limited [VS08b]. FEMME [SdH02]. fenced [JZY07]. feral [GDL06a, GL06, PDL^{+}08]. ferns [ZLO02]. Ferrara [Ano06-47]. ferrugineus [JJWF07]. fertility [LK07, MOP^{+}06, SC02]. fertilizer [CC05]. fertilizing [IF07]. fetal [PGLS03]. Fever [GGV^{+}09, CTG03, CTG04, TCGL03]. few [Bro09]. fewer [GORJ03]. Field [dBWG05, Ber02, BH06, BDK01, CR03, CML09, DPT09, ECP08, EFK^{+}03, EDKF06, HC03b, JcLM09, KBW00, LL02a, LKKL09, LiWZ00, LFJ^{+}06, MS07, NKK^{+}07, PDL^{+}08, RPPB04, SPB^{+}06, SKJvdHG06, TNO^{+}09b, VCP09, WC03, WM06, XWB04, ZCKR07]. field-grown [XWB04]. field-scale [LL02a, ZCKR07]. fields [FAA^{+}02, KKW08, PBMRE08, SC01, TFTO07, VTL^{+}09]. Fifth [KL08]. figures [Bro09]. filial [PSC04]. filled [JJWF07]. filling [OHM^{+}06, XBM^{+}07, XBM^{+}08a, XBM^{+}08b]. Filter [PCS03, KS04b, LCLC07, LLL^{+}07, MCJB08, NMJ07]. filter-feeding [LCLC07]. Finches [JF04]. find [HSMN08, HKPH08]. Finding [Jor05b]. finds [RLR09]. fine [LMH05, MB05a]. fine-scale [LMH05]. finer [MJR06]. finishing [GZ05]. finite [BPJM00, DAdSC08]. Finland [GBG^{+}03, KK00a, LSM^{+}04, LR07b, Mii00]. Finnish [RCL06]. FINS [SR02]. fir [BP08, BAP^{+}06, JAB^{+}06, MBK^{+}03, MAA^{+}09, MLR08, MAK^{+}04, TP06, vWBV02]. Fire [Wim04, BLDN00, BLHB06, BS08b, BCRSTVG07, BH02, BP08, CL06, DFF07, Don06, FCF^{+}04, FSM^{+}01, FCC^{+}04, GdBD08, HGT^{+}00, HSC^{+}04, Ito05, KCD^{+}04, KLM^{+}02, KVL^{+}09, LS01, Li00c, Li02, LC07b, MOLN06, MY02a, MNZC08, PCHG03, PBV05, PCB07, PFFM07, SPTP01, SLPP05, SHCS04, SWBJ01, SSM^{+}09, TAL^{+}05, VCM07, WGB^{+}08, YHG04, YDS08]. fire-exclusion [PBV05]. fire-managed [LS01]. firefly [KMR^{+}07]. fires [LR09, PMC03, RM02, SWSF06]. First [CMA^{+}06, Hey01b, BW01b, CCBB05, FC05a, FG07, RNC08, TPC^{+}07, WV05, WVP08, Jor05e]. first-order [WVP08]. Fish [KLPP07, Pen09, SR02, ATA03, BGL01, CdQO06, CMA^{+}06, CDV05, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, Fig09, FLS06, Fuk09, GDG^{+}00, GL09, GG02, GGP03, GGB^{+}06, GNA^{+}06, GW08, HB02a, Han09, HL04, HB05, HB09b, HBSM04, IGP^{+}03, JPB03, JMVvDV02, KPK^{+}07, KC03, KSH^{+}03, KD05, KGSB01, LS09a, LHB08, LPM^{+}07, MPM02, MPRJ04, MR00, MAdlPR02, MRiI^{+}07, MBW09a, MFG^{+}06, MFG^{+}08, MSD06, PGA06, PCC^{+}07, PGG06, RLHD01, RD07, RLBL01, RLFB04, RHB04, Sin07, SZL^{+}04, SRR06, TB06a, TSJ^{+}09, Tud01, VGK^{+}04, VPG05, VPG07, WL04, Xia00a, Xia00b, dG02a, dG02b, vLLv^{+}07, vNLS02, MP04a, PMD^{+}09, JT01]. Fisher [FC04, MPC06]. fisheries [AVP08, AMSH08, ASHHRRPE04, CGS08, CP04, DUASCM07, FMRS09, GL09, JCX^{+}08, LPM^{+}09, MP04a, McK00, Oke04b, PMD^{+}09, PCP07, SO04, TN08, dCPC01]. fishers [MSD06, Xia04]. fishery [ASZRRR08, Aub04b, Dew01, DLK01, DLC07, Kar04, LvNMvdB04, Leg08, MSS02, MAdlPR02, NAC04, PBH^{+}07, TMJ^{+}07, VREA06, ZRASC04]. fishes [DMBO00, DG04b, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, KP09, PSC04, RS01, SJH05]. fishing [AFTB07, AGNLGSG04, BDLL06, CSM^{+}06, CPT09, DLG06, FWHL06, GL09, GRW04, Gas05, GP09, MDH^{+}09, MHMHKA04, MMPT06, MAdlPR02, MCA^{+}06, OG08, PP04, SSH^{+}07, SCR03, TN08, Xia06, ZR04]. fishpond [dG02a, dG02b]. fit [MDH^{+}09, WSCR03]. fitness [McK01, WC03, YTT^{+}09]. fitophenology [ČKBB06]. FITPOP [McK00]. Fitting [Coh09, EFK^{+}03, PF00b, RPPB04, WNW09]. five [MF02b]. fix [SWBH08, VVF06]. fixation [BB07b, KVPA08]. fixing [HB06, KVPA07, KVPA08]. fjordic [PTGI09]. Flames [LC07b]. Flanders [AGD06, DGGD04, DVGD07]. flat [SSKN08]. flava [TMJ04]. fleet [PMD^{+}09]. fleshy [TMJ04]. fleshy-fruited [TMJ04]. flexible [KS07, POH02, SDS^{+}07, SdH02, ZHW05]. flies [WC03]. flight [SPS03, YTS03]. Floater [POF08]. floating [GCM^{+}05]. flock [SDD^{+}04]. flood [KFB09, RNC08]. flooded [CMB05, XJM07, dG02a]. flooding [GPDF09, GIKS08, HBRW07, PBS05]. floodplain [HR03, KJB07, MMWMH07, PLC08, AA05]. floor [Oke04a]. Florida [HUB02, Oke04a, WHH^{+}08, CP01, HGL^{+}06, KB04, MBKD02, PGLS03, RSF^{+}01, YMD08]. Flow [RS04, AB03, AS01, BB08b, BWPC06, CBP07, CBFLS09, CMCD04, GMPC08, GSB^{+}06b, GACO04, HB00a, JWW^{+}07, JSM03, KB04, KKA^{+}08, KS07, LGC07, Lat06, LLA^{+}09a, LBBN06, LXP^{+}08, MMRLP06, NP06, NSEDP06, SGP^{+}06, SGP^{+}07, SBB09, SOK03, SNRW06, Suh05, Suh06, WC07, WTMG09, WBP^{+}07, WG04, YKO05]. flow-balanced [LBBN06]. flow-based [BB08b]. flowering [GL08]. flowing [Rey03]. Flows [ASAC02, CM09, CSM^{+}06, Dei04, DBB^{+}08, LD06a, PLC08, RLBL01, US08]. fluctuating [Jen02, Jen05, LKR06, WJM^{+}03]. fluctuations [AÖÇ05, CSCB04, CR03, HPH00, KS04a, MY05, PCpDC09, WG07a, ZSPV08]. fluvial [DBBS03]. fluviatilis [BRW^{+}05, Han09]. flux [BGMP06, BO07, COB^{+}06, FRB^{+}05, KW00, LAXP07, MCD^{+}08, MH05, MCJB08, MZASLMLC04, NJF^{+}08, OHM^{+}06, SOHY07, WYT09, YY06]. fluxes [AFT09, BC03, BDR01, BPC04, DRDD01, ECP08, FGB08, JMVvDV02, JRT02, LBBR01, MMWMH07, PG08, RFA^{+}01, SC04, SOHY07, SYC04, SJG^{+}08, VCAS01, VA00, WV05, WKZP04, WLJ00, XBM^{+}07, ZDR03, ZXC03]. fly [YZS^{+}04]. flying [Bye00]. focus [HVVK09]. Fogliano [HMFM00]. fold [WDBK08]. Foliage [TYS^{+}09, Fle01, KBW00, Lar02, ŽJDK06]. foliage-dwelling [KBW00, ŽJDK06]. foliar [MBK^{+}03, STH06]. following [Bey03, PPR03, WYO05]. fontinalis [WPD04]. Food [AA05, CFCP04, CPTD08, GRR04, JKJ06, OB04, SB02, APHV08, ATD^{+}06, BMT07, CPT09, CN09, DRDD01, FH07a, Fat07b, GPDF09, GE05, GLS07, HFL07, HK02, Her04a, HH04, HvG07, JSM03, JBP07, JO09, KP09, KSvOO09, KPAK02, KBvVK08, KFJ^{+}09, KKZK03, LBS08a, LCP06, LS09b, LCLC07, MdRdA06, MPRJ04, MMRLP06, MCE^{+}07, NKC04, Nie08c, OL09, PM06a, PG08, PARH07, PA09b, PBM^{+}05, QHM05, Rec03, RLFB04, RV05, RLLB09, Roe00, RMWW07, RKS^{+}07, San07, SM04a, SOK03, SOK06, UDB07, Van08b, VP01, WPD04, WVP08, WLB^{+}05, WL04, Wil03, Wil05, lXzL02a, lXzL02b, ZK08, dG02b]. food-quality [Roe00]. Food-web [CPTD08, PBM^{+}05, WLB^{+}05]. Foodweb [RBW09]. footprint [CC06, Med06, NBT^{+}09, ZLL05, ZCY09]. footprints [KWD^{+}04]. forage [TPJ08, vLDHP08]. foraging [BG01a, BG01b, BKMB08, DH01, FAB^{+}07, GRR04, GR09, LC04, LJW03, NBP05, PBC09, Sti08, WC07, WMH08]. foraminifera [VTB^{+}08]. forced [DVPS08, MBDB09]. forcing [ASP^{+}07, MDH^{+}09, MCBA07, TSJ^{+}09, VM01]. forcings [Ano04v, CSU03, CSU04, PLA06]. forcings/temperatures [PLA06]. FORE [SS08b]. FORE-SCE [SS08b]. forecast [KFH07, WTMY07]. Forecasting [GNA^{+}06, MSL01, zGsLcXwZ09, GZJ06, HTMO06, MD06b, MDPC06]. Forest [Ale08, DFF07, ESWG02, New06, SAB03, WCHM02, WLL^{+}08, ZAM^{+}03b, AHKB01, APJ03, ACJT08, BW01a, BBGM05, BBGM06, BLHB06, BWJZ01, BK05a, BPC04, BAP^{+}03, BCRSTVG07, BH00, BJJ03, BLJS05, BDF^{+}06, Bir01, Bla07b, BS04, BDR06, BTdFC05, BRC04, CD07, CAB08, CKN^{+}01, CKA^{+}02, CKL^{+}06, CB07b, CMM^{+}07, CKPP03, CWBR01, CMPO05a, CMPO05b, DDG^{+}05, DBD^{+}08, DVdB^{+}08, DKL^{+}09, DFCM01, DMBM05, DDF^{+}05, DMRP07, EBH^{+}01, ECHN08, FRB^{+}05, FCF^{+}04, FGB08, FPK^{+}07, Gar04, GS09, GAB^{+}09, HHL^{+}02, HSC^{+}04, HvI06, Her03, HRJ^{+}00, HNF09, HD00b, Ito05, Ito07, JDPI07, JAB^{+}06, JNM^{+}06, KLM^{+}02, KW00, KS04a, KRN04, KCZ^{+}03, KLL01, KvKV05, KBE^{+}06, LLW^{+}06, LIS07, LAD06, LBL^{+}08, LBBC08, LE04, LPP^{+}07, LAPAMD07, LKP03, LMMK09, LH04, LR09, Lin01, LLF02, LL06, LZB^{+}06, LPPS05, LNW00, LHZ^{+}06, LLLT08]. forest [LH09, LL00, MGV00, MH02, MKB00, MRG09, MBM00, MBP^{+}08, MCM^{+}09, MGCK^{+}03, MKDV08, MMR06, MFJM04, MDB^{+}02, MCGO05, MAA^{+}09, MB05a, Mla04, MF02b, MM00, NMC^{+}06, NKK^{+}07, NJF^{+}08, OI07, Oga09b, OIP^{+}08, OIR^{+}08, PLD^{+}02, PGFM04, PE02, PE07, PSPD09, PdAD^{+}04, PTS^{+}04, PF01, PF03, PLP^{+}04, PMC03, PB02, RF09, RM02, RJR04, RPRV09, RHB05, RE03, RCL06, SPD^{+}07, SSK^{+}07, SM04b, SI06, SE04, SMG07, SL01, SWBJ01, SW03, SWSF06, Sti06, SW01, SLT^{+}09, SGH04, SPM^{+}08, SJG^{+}08, SZ03, TC09, TAL^{+}05, TYS^{+}09, TP06, TCD02, VSG^{+}08, VJ01, VSFM03, WWRZ04, WGVB01, WKZP04, Wan05, WM08, WHW^{+}05, WBB05, WRC^{+}08, WCS00, WW08, Wim04, WGT05, WBS^{+}02, WSZS08, WWC^{+}07, XHH^{+}04, XHH^{+}05, YSM^{+}06, YDS08, YP02, YPR^{+}05, YABM07, YL07, ZJN^{+}06, ZMB^{+}08, ZCG^{+}08, ZBZ^{+}09, ZBWR06]. forest [ZB04, ZPD06, ZAM^{+}03a, ZAM^{+}05, vWBV02, CMB^{+}02, FPK^{+}07, LIS07, Lug04, MLF^{+}06, SZ03, WHZ06]. FOREST-BGC [CMB^{+}02]. Forest-DNDC [LIS07, MLF^{+}06]. forest-fire [TAL^{+}05]. forest-landscape [LZB^{+}06]. forest-line [WW08]. forest-shrubland [PE07]. forested [EB07a, KKL^{+}06, KMRV07]. forestry [BFOS08, KDW^{+}09b, LA04]. forests [ASJD01, Bye00, CKPP03, FC06, GGPBEK07, GCLG03, GAA^{+}05, GZY^{+}06, Hör03, HBRW07, IOIA08, JPD^{+}06, KKH^{+}08, KS07, LA07, LvGC^{+}04, LAZ^{+}08, MHM^{+}03, NW06, NH07, PJAZ02, PDV^{+}07, PBY^{+}03, RFA^{+}01, RNKG03, RGL^{+}07, RSB09, Rob03, RCZ^{+}06, SBR^{+}07, SHCS04, SSNB07, SRG^{+}09, SPM^{+}08, TBP^{+}07, TK01, VSHC08, WRB08, WKZ03, ZdlP05, ZPD^{+}08]. Foreword [CF08, Des03]. form [Kin05, SH09a]. formalised [DP07]. format [FGFB05]. formation [FWS^{+}05, HB06, Hos06, HBSM04, Kir06b, Kno03, MdRdA06, MP00, SKK^{+}03, VM01]. Formica [BLDCM06]. FORMIX3 [HD00b]. Formosa [GE05]. forms [Cai05, KVV01, YMD08]. formula [MHN00]. Formulating [RTB04, ZPD06]. Formulation [PA09a, CM06a, CM07a, CCJ07, Her03, LUKD06, LWL^{+}02, SBM04]. formulations [MLGV00b, TR03]. ForNBM [ZAM^{+}03a, ZAM^{+}03b, ZAM^{+}07]. forsteri [BDP^{+}02]. forward [Cor05, MD06b]. found [LJ09]. foundation [Hul04]. foundational [COC04]. Foundations [Met01, Met02]. founder [IA08]. four [Bey03, CSF^{+}04, Hul02, Hul04, MBS^{+}09, MF02a, War08, WR03, ZYY09a]. Fourier [HD00a]. Fourth [DŽD06, Ano06-48, DDŽ06]. Fox [Jen05, KSG05]. foxes [Sel00, SHW04]. Fractal [GPL05, Li00a, UZSM05, APG^{+}03, JNM^{+}06, MCGO05, SHZ05, SLE07, VMG05, vNM02]. fractals [MPP05]. fractions [JLH01]. fragmentation [ECHN08, GAB^{+}09, JCE06, LK03, MY02b, PBA06]. fragmented [BWAM09, CN07a, KJ08, LL07b, RBC09, ZC00, ZPO09]. frailty [WZC^{+}05]. Framework [TSJ08, BMR07, BLJS05, BG08, BGWC07, DN06, DLPN06, DBB^{+}08, EPQdCA06, Her04a, HBDA09, Kin04, KS07, LBBC08, MHvIR00, MJ02, MNPJ03, MSM^{+}08, PBG09, Rai01, Rec03, RS04, RPPB04, SLS03, Sch03, SCH05, Vil01, ZA09, ZCY09]. frameworks [RBB05]. France [BSB^{+}09, BG08, MBLA03, RLBL01, TPCS06, BMT07, CGC01, DSD^{+}09, FEP^{+}04, LPFL09, RNKG03, RLLB09, VSG^{+}08, VML^{+}06]. Francesca [Jor05a]. free [AGZ05, KVPA07, KVPA08, hMzL08]. free-living [KVPA07, KVPA08]. free-path [AGZ05]. freeze [NJF^{+}08]. freeze/thaw [NJF^{+}08]. French [MFG^{+}06, MFG^{+}08, BLDCM06, PCM^{+}03, TPC^{+}07]. frequencies [GC02]. Frequency [CSHY08, BLHB06, EBM06, Li02, Lir03, NHLPA06, PCS03, TKSP09, WC03, YHG04]. Fresh [MHZ^{+}06, Mon09b]. freshwater [ASP^{+}07, CBSLS07, DRDD01, EIRT00, FDCH08, GW08, GRPF07, GB03, HB09b, JKJ^{+}08, MFB^{+}05, MWMN06, RD07, RC08, RIGJM06, SM03, WR01b, WR01c]. frog [BJJ06, BM08]. Front [Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-37, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Sil07]. frontier [SFCP02]. fruit [OAAF07, SW01, WWRZ04, YZS^{+}04]. fruited [TMJ04]. fruiting [KWW^{+}03]. fruits [SWB06]. Fryxell [MMRLP06]. Fuel [KVL^{+}09, SHCS04, HSC^{+}04, HNF09, KRZ07, LL07b, RCZ^{+}06]. fuels [BCRP09, ESZ^{+}00, FCC^{+}04]. fuelwood [CRM09]. fugacity [PBC02]. fugitive [AW06a]. full [Mat06]. fulva [RAH07]. function [AY07, AO00, APJ03, BvdW04, BPE^{+}07, CD05, FBC08, JAN^{+}03, JOBL08, KCJ^{+}07, KP00, LL07c, Mat03b, MMF^{+}09b, MMF^{+}09a, MAB01, PBE^{+}07, PBC09, RKH05, RBR05, SM06, TR03, TF05, TSF07, vLDHP08]. Functional [RST05, RKH05, vNM02, BCL^{+}09, BHC04b, CBD^{+}09, CWWS01, DH06, DSD^{+}08, Esc05, EGE^{+}08, FF06, FSM^{+}01, GR09, GdBD08, Her04a, HM01, JVL02, KJ08, mLMfT05, LL08, hMzL08, ML06, MR08, NM09, Oku09, PS05, SB02, SLZ09, TDHO07, VSFM03, WH09, Zav04, ZRCA08]. functional-structural [PS05, TDHO07]. functioning [GPDF09, GMPC08, GP09, KSvOO09, KGSB01, LZZA09, Lev00, RMSS02, TMHJ06, WSZS08, Zav08, vW07]. functions [BVNS02, FRZ00, GP02, JG08, KGBC03, Lar02, LLLP00, McK01, Mul07, OW05, Sta07, SÅ06, TN09, WFM01, WDBK08, WB00, ZHW05, BLC^{+}07]. fundamental [BBGH08, Den08, Ned09, Xia07]. Fundamentals [Xu03]. funding [Tis06]. Fundulus [MWWZ01]. fungal [KWW^{+}03, LDM00]. fungi [CGD04]. fungus [SGHM01]. Fure [GZJ06]. Further [MEO06]. fuscipes [OMA01, OMA01]. Future [AMRS08, Jør04f, BDE08, BHP05, BBT06, CWL05, GRR04, HvI01, HO01, KK00a, Med06, Nie09a, OZL07, SVB09, TMS^{+}07, WIMK07, Wil01, WSY^{+}07, Jør04h]. future-oriented [HvI01]. fuzziness [Fuk09, Met03]. Fuzzy [AGD06, AÖÇ05, BBT00, ES01, ML04, NCM07, PEE09, Sil00, VADV06, WLLY04, ABC04, BDR06, CM03, CM06b, CH06, GML09, GIKS08, HKB02, HTS^{+}07, KBW00, LS08a, LSHG08, LSY^{+}09, LS02b, Mac00, MM06a, MM06b, Met01, Met02, Met03, MSP^{+}08, MDVG09, MAG01, ÖÖ04, dSPdBB08a, dSPdBB08b, Pra05b, RSM05, Sti06, SGHG04, WXYMZZF03]. fuzzy-constrained [BDR06]. FVS [LLW^{+}06, WBR08]. FVS-BGC [WBR08].

G [Jor05d, Jor05h, Jør06a, Xu03]. G. [MC04]. GA [Ano09-33]. Gadus [SMB^{+}06]. Gaertn. [LPD08]. gaeumannii [MBK^{+}03]. gain [Esc05, RHP07, TYT^{+}09]. Galápagos [Oke04b]. Galemys [BRV09]. Galicia [DADGA06]. gall [KH00]. gall-midge [KH00]. game [MAL06, MG01, iTI02, ZZL^{+}09]. Games [CT07b]. Gammaridae [MMM02]. Gammarus [AGD06, GBB^{+}09]. GAMs [WA02]. GaoShan [TYLL07]. Gap [Jør08a, BMB07, DFCM01, FPK^{+}07, KBF^{+}08, MKB00, MBP^{+}08, OHM^{+}06, Pet02, PMJ08, Ric02, RA06, RHB05, SL01, XBM^{+}07]. gap-filling [OHM^{+}06]. gaps [SI06, XBM^{+}08a, XBM^{+}08b]. Garden [KWW^{+}03, SBB06]. Gardner [MWM02]. Gargano [TAL^{+}05]. Garonne [CGC01, RLBL01]. GARP [SF07, SBS^{+}06]. gas [Alo04, BCST05, BB07b, FPCA00, FGB08, MWD05, OT03, SR02, VSFM03, WBTC00, Zav04, vWBV02]. gas-exchange [OT03]. Gaussian [CK07a, CG06b, PWC^{+}09, SBJ^{+}02]. Gaztelugatxe [BBM06]. geese [HRC^{+}07]. gene [CBP07, CBFLS09, CMCD04, KS07, LLA^{+}09a, SSPR03, TTA^{+}03]. General [Kir06a, Aok08, BH03, BVC^{+}01, CCM02, DD00, FPK^{+}07, HB03a, HBM04, Håk09, KTB07, KC05, KJ08, MJ02, OSHO09, PG02, SS06a, Xia00a, YP02, ZJC^{+}07, Ano04-28]. generalised [GH07, Hee02]. generalist [RBR05]. generalists [WÖW05]. generality [SB07]. Generalizability [ÖTÖR06, VSHC08]. generalization [MZWM05]. Generalized [BW02, GG08b, GEH02, CM05a, LEH06, LOL02, LOL03a, LOL03b, Mat03a, MF02a, MFB^{+}06, Nad07, WHZ06, Xia05a, YM02]. generalizing [Xia00c]. generate [SA07b]. generated [GGH08, RWM^{+}07, SRN05]. generates [EB07a]. Generating [TN01, EKBF04, KK07]. generation [Hel08, Leb05, SIS^{+}07, TM05b]. generation-tracing [SIS^{+}07]. generational [Gra04]. Generic [RvGC^{+}08, BMR07, BFOS08, CWS09, HAS07, MP04a, PK01, PK02, PLD^{+}02, RS04, WR01c, ZC00]. GeneSys [CMCD04]. Genetic [DGD06, HGD05, ML05, CA04, CC05, DGD03, DLC07, Gre06, HLHZ06, JOB04, KBF^{+}08, LL07c, McK01, OHM^{+}06, PCWP06, PFR09, RB02, RS04, SBS^{+}06, SBC07, TMP06, VF07, Whi00, WR01b]. genetically [BMS^{+}08, CBP07, IVC^{+}08, ITDD09, LA04, RS04, SvdWB^{+}06, TTA^{+}03]. genetics [McK00, PTS^{+}04]. genome [Deb02, LJ09]. genotype [CMCD04]. genotyped [Shi04a]. genotypes [KS07]. geo [ZGL08]. geo-information [ZGL08]. geochemical [CYHK04, KKW08]. geogas [ZTXD03]. geographic [BRV09, JJ00, Mey04, OEK^{+}06, PSCS^{+}01, PAS06, dJTMBGMP^{+}09]. Geographical [ADSO08, Jor05g, DSA08, DSD^{+}08, KMR^{+}07, MLHT09, DDS^{+}04]. geologic [Loe04, Swa06]. geological [Sem08]. geometric [CM06a, CM07a, MBGP08, MSL06, YTT^{+}09]. geometrical [Ric00]. geometry [KP00, Li00a]. geomorphological [BAP^{+}06]. Georg [Jor05g]. George [MKS^{+}02]. Georges [CGH^{+}03]. Georgia [Ano09-33, FCC^{+}00, OB04, Wim04, ZCZ04]. Geospatial [RPC^{+}05]. Geostatistical [MDB^{+}06, WLB^{+}04, BKC^{+}07, LMPT05, XHH^{+}05]. Geostatistics [SM07b, LOM06, SSK^{+}07]. Germany [WR08, BAP^{+}03, BRGS09, FRB^{+}05, HZF07, KJB07, MAA^{+}09, SS00, WTS^{+}06]. germination [HFV03]. GETLAUS01 [GTRF01]. giant [ALO^{+}01, Bas04, HC03a, LAB^{+}05, NW07]. Gibbs [WPBM06]. Gibraltar [Per07]. Gigartina [MWM02]. gigas [RR01, RLSZK^{+}08]. Gilbert [Jor05l]. gilthead [HGLLV03]. Gironde [RLLB09]. girth [MLR08]. GIS [GB08, Met02, AN06, ESG06, KA00, LWLZ06, LZZ^{+}07, MVMA08, MBT03, MEJ06, MMR06, OW05, PS01, RSC09, RCL06, SJ03, TS03, TTA^{+}01, dJTMBGMP^{+}09, VGK^{+}04, VCRD06, XTDL01, YSM^{+}06, YDS08, YZC^{+}07, ZNC^{+}06]. GIS-aided [ESG06]. GIS-assisted [ZNC^{+}06]. GIS-based [MEJ06, OW05, PS01, SJ03, TTA^{+}01, dJTMBGMP^{+}09, VCRD06, XTDL01, YDS08]. GIS-integrated [AN06]. GIS. [Met01]. Giuseppe [Ano05q]. glacial [GTJ^{+}00, LMM^{+}07]. glass [KWD^{+}04]. glaucophylla [RBE^{+}08]. glide [LHT^{+}08]. Global [Ale07, CS08, KBB00, KBE^{+}06, Phi02, SPTP01, SIK07, SBC^{+}09, SHSP04, TS03, WSY^{+}07, AMK00, CKBH00, DJ05, FJ05, GGB^{+}06, GL08, HY07a, HVB06, HCL00, KC01, LSS^{+}00, Mey04, MSB08, MHSP^{+}06, NHLPA06, Pen00, PSCS^{+}01, PWC^{+}09, SOHY07, Sem08, SNF01, SHP^{+}07, Svi00a, Svi02, SSP03, SZ03, Svi04, TSBH09, TTJ^{+}09, WHP03, Jor05f]. Global-scale [WSY^{+}07, HVB06]. globalization [AOY02]. globally [WHP01]. GLOBE [RLR05]. globosa [CM04]. globulus [CMPO05b, MLF^{+}06, MBS^{+}09, MB02b]. Glossina [OMA01]. GLUE [PHBF07]. Glumsø [ATDK08, VKB06]. glutionsa [LPD08]. GMM [WMH08]. Goal [WB00, KGBC03]. goals [BJJ03]. Gobi [RR05b]. gobio [CMA^{+}06]. golden [iTI02, AÖÇ05]. golf [Man00]. Gompertz [Kar04]. good [Oli03b, TPC^{+}07]. goodeid [BMBOCR03]. goodness [WSCR03]. goodness-of-fit [WSCR03]. goods [SYC04]. gooseneck [BBM06]. gopher [BWAM09]. Gopherus [BWAM09]. gorgonians [BIS09]. Goro [VCRD06, MNZ06]. goshawk [RJR04]. gossypii [dSMZ09]. government [Pet04b]. GPP [VSG^{+}08]. GPS [GPC^{+}09, SWBH08]. GRAAL [DP03, DP07]. GRAAL-CN [DP07]. gracefully [TW00]. gradations [AOY02]. gradient [AEP^{+}04, Dam03, Dam08, HHL08, LKLK07, MSA^{+}03, MFB^{+}06, PM06a, RKH^{+}07, SK04, VM07, ZRCA08]. gradients [EKBF04, LA07, OM02, Sin07, WHHH07]. grained [MJR06]. grains [CdQO06]. graminoid [HUB02]. Gran [SF07]. granaries [JJWF07]. Grand [YBM^{+}05, FEP^{+}04, FCC^{+}04, RSB02]. Grande [GCG^{+}07, HTS^{+}07]. Grant [Jør09b]. grape [MDPC06]. grapevine [RCGB08]. graphical [Van04, Vil05]. graphs [DDT07]. GRASP [LOL03a, LOL02, LOL03b]. grass [AR02, IMS07, LMH05, LCF09, MC01, RAH07, RSA04, SRA05, SAR^{+}09, SLGS00, SZ03, dSA01]. grasshoppers [Fie04, WJL05]. grassland [BHW^{+}08, DAC^{+}09, FRB^{+}05, MBM00, MT02, MBD^{+}00, MY02a, Pet02, RMDC04, SLS03, Sch03, SOA03, STK00, TMHJ06, WW08, WHP03, WBTC00, XSY^{+}09, ZGF^{+}05, ZPK^{+}07, LCF09]. grassland/forest [MBM00]. grassland/oak [MY02a]. grasslands [AEK^{+}07, GPB01, HG01, LS01, RGF00, RMSS02, SLS03, Sch03, WBN^{+}03]. gravity [NDD^{+}07]. gray [MJH02]. grayling [MSP^{+}08]. grazing [AEP^{+}04, BLDN00, BKS05a, BKGC05, CTG03, GL01a, MAHvD08, PCHG03, PFFM07, RHH05, RKH^{+}07, Shi04b, SB02, SvL04, SHM07, TDL^{+}07, WZJ08, WRB08, ZPK^{+}07]. Great [Gri04, BB01, Cha02, EBH^{+}01, RE03, Wil08]. greater [VF07]. Greece [AG03, GPSM08, NKK^{+}09]. GREek [FPK^{+}07]. Green [NS08, BR01, Cha02, CB07a, HMBG03, Sno08]. greenhouse [CGY08]. greenness [PLH09]. greens [HMBG03]. greenshell [RR05a]. Grewia [TMJ04]. grey [DFGC04]. Grid [RGO^{+}06, WW08, Bir06, HXP^{+}09, SBS^{+}06]. Grid-based [RGO^{+}06, WW08, Bir06]. grids [BOB07, BM07]. gross [Lar02, VHG05]. Ground [Ayd08, WL06, DAC^{+}09, HTK07, MCM^{+}09, SRW05, Urs09, WX09, XSY^{+}09, ZBSA07]. Ground-level [WL06]. groundfish [SSH^{+}07]. groundfishery [CNG06]. groundwater [AHKB01, BKS^{+}05b, IF07, LLH^{+}06, PBD00, ZWCL05, Jor05g]. group [DH01, FBF^{+}06, Mog02, Sil04]. grouper [ALAS09, AS00b]. grouping [MAL07]. groups [APJ03, BHC04b, DSD^{+}08, FSM^{+}01, GICB09, HY07b, JVL02, MR08, PF03, SLZ09, ZRCA08]. grouse [MML00, MML02, PCHG03]. growing [ETH^{+}04, GZ05, HB01, HB06, ZvBS05]. growing-finishing [GZ05]. grown [BGWC07, CS07, LM07b, XWB04, ZBSA07]. Growth [AR02, HH04, KVV01, RAH07, RSA04, SRA05, ÅSCP09, ARL^{+}06, AO08, Ale08, ABR05, ARF08, BC01, BSJ^{+}02, Bra01a, BMD09, Cai05, CSR08, CGD04, CM07b, CSF^{+}04, CGR03, CWBR01, CTH^{+}00, CMPO05b, DADGA06, DMRP07, Esc05, GGH08, GRBT08, GIKS08, GH07, HCJ^{+}09, HC03a, HTA^{+}08, HHC02, HL04, HS03, HS06, HGLLV03, HRJ^{+}00, HD00b, IMK^{+}07, Jon07, KBGJS06, Kin05, KCZ^{+}03, KFS06, KRvL^{+}02, LKR06, LPD08, LH05, LH06, LE04, LvGC^{+}04, LS08b, LdVA06, LWL^{+}02, MWWZ01, MGL04, MBK^{+}03, MPRJ04, MHM^{+}03, MOP^{+}05, MOP^{+}06, MKRH03, MKM^{+}08, MKvdW^{+}09, MMR02, MAA^{+}09, MWWM07, MBS^{+}09, MY05, MC01, MLL^{+}05, MMA02, MDH^{+}08, MLGV00a, MKiIK07, NP06, OCK01, OAAF07, Oga09a, OSHO09, Par00, PPE^{+}07, PLD^{+}02, PBY^{+}03, PB02, PV04, RFA^{+}01, RPRV09, RR05a, RC06, RE03, RBR06, RWM^{+}07, Sak07]. growth [SIS^{+}07, SM07a, SM04b, SBM04, SOK06, Sil07, SMTR07, SO06, TC09, TR03, TF05, TSF07, TDHO07, TSZdRR03, TBPF08, TFM01, WPD04, WHH^{+}08, WBR08, Weg00, WM06, WS02, WSF^{+}02, WvS06, WWC^{+}07, WBR^{+}06, YM05, YMD08, Zei04, ZLZM02, ZBL07, ZCG^{+}08, ZBWR06, ZB04, dlS07, vODFS04, vdBDR02, DP03, DP07, JPS00]. growth-limited [MMR02]. growths [CMM02]. Grübler [Jør04h]. Guangzhou [zGsLcXwZ09]. Guide [Jor05a]. guided [KKW08]. guidelines [SBVB05]. guild [DS01]. guilds [IGP^{+}03]. Guinotte [Mai08]. Gulf [LR07b, MZASLMLC04, NKK^{+}09, ATAK00, ASAC02, ASZRMH^{+}04, ASZRRR08, BVD05, CS08, DUASCM07, Håk09, LCG^{+}09, MHMHKA04, MAG01, OL09, RLSZK^{+}08, SSH^{+}07, WHHH07, XIX^{+}08, ZC07]. gull [BL01]. gullies [GSC09]. Gulls [LKY^{+}06]. gunnii [TIJ^{+}01]. Gwydir [PLC08]. gypsum [KNZ04]. gypsy [Gra04, LRT^{+}08, Wil01].

H [Jor05h, Jor05j, KBE^{+}06, MMB07, MWD05]. H. [Ano04-46, BU04, Cam04, Com04, Kan04b, Lim04, Lug04, MC04, MD04, SH04, Ulg04]. HAART [Wal04]. Habitat [CGHW05, DDJ^{+}01, KDK06, RCL06, SGHM01, BLHB06, BCM^{+}08, BDI04, BBGH08, CDM05, Cal06, CL08b, CCLS06, DGM08, DFM07, DMF07, DC08b, DD03, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, FE04, FKIL08, FH08, Fuk09, GBB^{+}09, GBN^{+}06, GGP03, GGP06, Gra07, GZ00, HP09, HDB^{+}06, HHM01, HG02, HLH^{+}06, HRMC01, HBUS02, HLS06, HGR08, IKS09, JCE06, KMR^{+}07, KA00, LTM^{+}04, LQL05, Lin05, LAB^{+}05, LLZ^{+}08, MCQA08, MNOS08, MBW09a, MP08, MSB08, MY02b, MSP^{+}08, MG02, NiTT01, NiT04, OW06, OW08, OLB04, Par02, PF00a, PHP04, PBA06, PLP^{+}04, PSVH09, PWSS07, RB06, RPRV09, Rd06, Rem04, RKS^{+}07, RSM05, SM07b, SAL07, SCAP05, SV03a, SCH05, Sin07, SHG^{+}08, SJ03, SB07, SHZL09, VGK^{+}04, VADV06, VSHC08, VCRD06, VBRS07, WÖW05, WL04, WGV^{+}08, YEMZ03, ZPO09, ZGL08]. habitat-contamination [PWSS07]. habitat-islands [PHP04]. habitat-suitability [HHM01]. habitats [ALO^{+}01, BPAB^{+}06, BvdW04, BHC04b, BF07, JAK^{+}06, OMA01, SHG^{+}08, SGHG04, TDdSLS^{+}08]. Habitone [SRB06]. haemorrhagic [FSBD01]. Hägerhäll [Jør04f]. Haibei [YY06]. Hakea [LKH^{+}08]. Hal [Log02]. halepensis [RNKG03]. half [YLJ^{+}01]. halibut [CH06]. Halimeda [YMD08]. Hall [Jor06b]. Hall/CRC [Jor06b]. Hampshire [BB01]. Handbook [Jør04a, Cha09, Hey01b].

Handling [VM06]. Hangzhou [HJ02]. Hanoi [ABV^{+}06]. hantavirus [WLG07]. harassment [VS08b]. Harbour [XLZ^{+}04]. Hard [Jor05a, Jor05c, Jor05f, Jor05j, BCL^{+}09, Jor05d, Jor05g, LYC08, SAS06]. Hardbound [Bro09, Jør06a, Jor06b, Xu03, Zha06, Jør04b, Jør04c, Jør04d, Jør04a, Jør04e, Jor05h, Jor05i, Jor05l, Jør04h]. hardwood [DFCM01, FBDM09, GB02, NH07, SHCS04, ZBW05]. Hare [LGS03]. harengus [MRK^{+}07]. Harlem [BS08b]. Harmonic [KGJ08]. harmonious [Gia04]. harpacticoid [RWW07]. Hartwell [RBW09]. Harvest [Jen05, CB07a, Cha00, CK07b, Gan06, HAA08, HCL06, Jen02, MI01, Ort08, Sno08, TMD04, ZPP06]. harvested [FLS06, TK01, TSJA02]. harvesting [ACJT08, ASZRMH^{+}04, ASZRRR08, CKBH00, GLHV08, GG05, Jen00, KC03, LM07a, NG07, PJAZ02, RRB^{+}01, SvL04, ZRASC04, ZAM^{+}05]. hatchery [SCH05]. hatching [JPB03, JZC^{+}07, MVZM05, SBL03]. Hawaii [TM05a]. Hawaiian [Sno08, CB07a]. Hayden [ZAM^{+}07]. hazard [OG08, RSC09]. haze [MOLN06]. headwater [KRK05]. health [CYHK04, LYCU09, LL03b, Rey03, SYCU09, XLZ^{+}04, XZZ^{+}05]. healthy [BRW^{+}05]. heartwood [Kno03]. heat [BGMP06, SS08a, SJG^{+}08, ZXC03]. heavily [NKK^{+}07, SSH^{+}07]. heavy [BA05, Del04, LH01, ZTXD03]. hebes [TSZdRR03]. hedgehogs [DAR^{+}07]. Heidelberg [Jør04b, Jør04e, Jor05d, Jor05f, Jor05i, Jor05j]. height [KBK07, MBM00, PS09, PDS07, XCW07]. held [Ano06-46]. Helicoverpa [SvdWB^{+}06]. help [LH09, LBBR01, Sel00]. helps [PFBBJ08]. hemispherical [JAB^{+}06]. Heracleum [NW07]. herb [LH05, LH06, WH07]. herbaceous [Pet02, SGHG04, WBTC00]. herbage [DAC^{+}09, TMHJ06]. herbicide [CBP07]. herbivore [BBM04, BJH01, HG01, OzDBS07, OS02, Sar04, SHM07]. herbivores [OzDBS07, vLDHP08]. herbivorous [BB01, HB03b]. herbivory [JNSS02]. herbs [JAB^{+}06, SRL^{+}00]. Herd [HGD05, CM05b, DDS^{+}04, Eza05, HA08a, SDD^{+}04]. Herding [Jua09]. hermatypic [KVV01]. herpetological [DRDD01]. Herring [DLC07, BL01, Mac00, MRK^{+}07, RWM^{+}07, VS08a]. hesperus [MDGV09]. Heterogeneity [Bla07a, ACE07, BLHB06, CMA^{+}06, FCF^{+}04, Fie04, FMC^{+}08, GLD07, Gra07, HLxY04, JBT^{+}05, LMH05, LRJ^{+}09, LS01, LLZ^{+}08, MRT05, MB06b, PG02, STK00, SRRS04, SHM07, TM01, VMG05, YM05]. heterogeneous [ADS^{+}07, BH02, Bio01, Bio03, GG04a, GLHV08, GLP05, Gil08, HGT^{+}00, IWW08, LDM00, Mat01, MDB^{+}06, RKH^{+}07, SHZ05, SMET06, WW03, WÖW05, Zha03a]. Heterotrophic [TVK^{+}08, HHB^{+}08, TFM01]. heuristic [CM03, LCM^{+}09a, LK02, McK00]. hexagonal [BOB07]. HFCW [WTMG09]. Hg [LYC08]. Hg-stressed [LYC08]. hidden [FCH04, FCKH06, TA05, VCC03]. Hierarchical [JLAM09, ASG^{+}05, AO00, BLJS05, BHSR01, BRC04, CD07, CB07c, DDS^{+}04, GS09, HL04, Kir01, KNM09, KFJ^{+}09, MCKNM09, Ric05, SV03b, WIH^{+}09, WD02, YHG04, ZA09]. hierarchies [LK07, ŽJDK06]. hierarchy [Bog04, BOJ04, HBCL07, LSY^{+}09, Nie00b, Sti06, TM01]. high [Asp02, CHW07, DSA08, EPS04, GAA^{+}05, GPC^{+}09, HDB^{+}06, KWS^{+}07, LJR06, MLGG09, PL04, PCS03, SWBH08, TFF07]. high-level [CHW07]. high-nutrient [EPS04]. high-value [DSA08]. higher [MRiI^{+}07, PK02]. Highlighting [Sal06]. highly [HKPH08, RBC09, ZPO09]. hill [WZKL09]. hill-pasture [WZKL09]. hillslope [KKL^{+}06]. hilly [LZZ^{+}07]. hindcasting [HHD01]. hippopotamus [LC04]. HIRVAC [FGB08]. hispid [WLG07]. hispida [FKIL08]. hispidus [WLG07]. historic [RWM^{+}07]. historical [APHV08, Håk09, KPH02, KRZ07, SRK06]. histories [KGA06]. history [BC01, BCH09, CM07b, Fie04, FPS03, Jag01, Jag09, Jør09a, KGZR05, LK03, NG09, RV05, RG06, SR02, SCPC^{+}07, SW03, SO06, Wal04]. HIV’s [Wal04]. HMM [LBS^{+}06]. HMM-based [LBS^{+}06]. hoarding [GRR04]. hog [Sav00]. Hogweed [NW07]. Holdridge [YLJ^{+}01]. holistic [ML06]. hollow [OW06]. hollows [Nun03]. Holocene [LLF02, NHP^{+}06]. holometabolous [ZZ08]. Holothuria [DD03]. Holt [dlS07]. Homarus [GBB^{+}09, ZC07]. Home [WG07b, HGR08, KM06, MP04b, MP08, TKTB07, WPMT07]. honeybee [SC07]. honeybees [AGM^{+}08, WS02]. honeydew [JDPI07]. Hong [XLZ^{+}04]. Hooghly [MRG09]. Hopkins [CRL09]. HoPoMo [SC07]. hopper [HC06]. Hordeum [BMD09]. horizontal [DW01, LMH05, LMPT05, MLC05, NSO^{+}08, WX09, WTMG09]. Horn [AÖÇ05]. Hornem [PCM^{+}03]. host [BSTS^{+}02, CRL09, GLL^{+}07, MB02a, MMT^{+}07, NHLPA06, RD07, SB06c, SRRS04, TNO^{+}09b]. hosts [HNS08]. hotspot [PMLM08]. Hourly [BRK07, AMK00, BCCR^{+}02, BW01a]. House [KTB07]. household [ALO^{+}01, EMdC^{+}01]. households [LAB^{+}05]. Howard [Bro04a, Phi04, Til04, Zuc04]. HSB [FCC^{+}00]. HSB-C [FCC^{+}00]. HSPF [RNC08]. Hubei [mLMfT05]. Huds [CS01]. Hui [GSB06a]. human [ASJD01, Ale07, BCRSTVG07, BBC03, CYHK04, CM06a, CM07a, CBSLS07, CP01, DN06, EBH^{+}01, Fis09, GPP02, HNS08, Liu01, Mat06, MK04, NJB^{+}09, RSB02, RBC09, SF07, TCGL03, YWL^{+}05, YTLF08]. human-caused [APCdlR07]. human-disturbed [CM06a, CM07a]. human-dominated [RBC09]. humans [KLM^{+}02]. humic [MFB^{+}05]. Hummingbird [BG01a, BG01b]. hummocks [Nun03]. humus [WFB^{+}08]. Hunan [YZC^{+}07]. hundred [Jør07a]. hunting [BLBT01, LMG08, MBLA03]. Huron [TTHH07]. Hurricane [KW02b]. husbandry [MTKM^{+}06]. huxleyi [BSM08]. hyacinth [MI01]. Hybrid [KFR06b, LK02, GRBT08, PLD^{+}02, RE03, WPBM06, WL06]. hybridising [WR01b]. hydraulic [BPA08, BDI04, COB^{+}06, ESWG02, MVPB02, NSO^{+}08, WX09]. hydraulic-bioenergetic [BDI04]. hydro [AHKB01, BKPS08, ZGL08]. hydro-ecological [ZGL08]. hydro-electric [AHKB01]. hydrocarbon [WLB^{+}05]. Hydrochaeris [FC05a]. hydrochemical [PDL06]. Hydrodynamic [FU05, UZ01, CG00, MBDB09, MNZ06, NP06, TSF^{+}05, WHH^{+}08]. hydrodynamic-ecological [TSF^{+}05]. Hydrodynamics [GACO04, AAA00, SM04a]. hydrogen [VPV09]. hydrogeochemical [LHC07]. hydrogeological [CYHK04]. hydroinformatics [MWMN06]. hydrologic [AMSW07, ASP^{+}07, HZF07, KKA^{+}08, Pyk04, VZG05, XS08]. hydrological [AAO06, BSB00, BBGH07, Bor07, FCM05, GQB05, HKPH08, HHF05, KMR^{+}07, KKL^{+}06, PKS^{+}07, SCP05, TH06a, WHW^{+}05]. hydrologically [SSCS06]. hydrology [FWSB05, PCS01, TB06a]. hydrometeorological [FCM05]. hydropower [VZG05]. hydrosystem [TPC^{+}07]. Hydrothermal [HFV03]. hyper [BP04]. hyper-eutrophic [BP04]. hypercube [XHH^{+}05]. hyperdynamism [WM08]. hyperspectral [Asp02, YSG^{+}06]. hypertrophic [RvGC^{+}08]. hyphae [DdSG01, DdSG04]. hyphae-induced [DdSG04]. hypotheses [PGLS03]. hypothesis [BSR06, HZHL05, LS09a, MJ02, MML00, MML02, SLPP05]. hypothetical [BM08]. hypoxic [KSH^{+}03, SSKN08]. Hysteresis [ZJBI03, SAK00].

Ibera [SBVB05, CGH^{+}05, SCP05, CGG^{+}05, FC05b, FCM05, dCS05, GCM^{+}05]. Iberian [BRV09, GBN^{+}06]. ibex [LMG08]. ICBM [KA01]. ice [RC08, TMM05]. Idaho [PCHG03]. identical [WZW05]. identifiability [ABV^{+}06, DWR09, OBR01]. Identification [BPBF^{+}00, BHFMG05, LR07a, MLC05, DC08b, OAL^{+}07, PBCZ01, dJTMBGMP^{+}09, ZLW09]. identify [DBB^{+}08, DD03, Eza05]. Identifying [GH07, JOBL08, KA00, TTB06, TKTB07, LCP06, MAB01, PSVH09]. IEMSS [Ano08a, LBL^{+}08]. Ierland [Jør04d]. if [KPKP06]. IFC [Ano04x, Ano04y, Ano04z]. IGBEM [FSJ04a]. ignition [APCdlR07]. II [ALJL03, AB05b, CCG07, CSB03, CMPO05b, DDG^{+}05, DDS^{+}04, FPSJ04, GARB09, GAPE06, HKB02, Kir06b, SDS^{+}07, Sch03, SSM^{+}09, SPM^{+}08, YTCV01, ZAM^{+}03b, dG02a, dBWG05, vdBDR02]. II.Ecosystemsandmetabolism [Ano04-27]. iii [Ano02-52, Ano03-59, Ano03-64, Ano03-60, Ano03-61, Ano03-62, Ano03-63, Ano03-65, Ano04-76, Ano04-77, Ano04-70, Ano04-71, Ano04-72, Ano04-73, Ano04-74, Ano04-75, Ano05-46, Ano05-47, Ano05-48, Ano05-49, Ano05-50, Ano05-51, Ano05-52, Ano05-53, Ano05-54, Ano06-39, Ano06-40, Ano06-44, Ano06-41, Ano06-42, Ano06-43, Ano04-28, CV01]. Ikonos [KDK06]. ilex [VSFM03]. Illinois [MNOS08, WG07a]. Illuminating [OJ02]. illustrated [Bro09, LK03]. illustration [LPM^{+}07]. illustrations [Jør04b, Jør04e]. image [DGRU06, KDK06]. imagery [CGG^{+}05, RLR05, SAS06]. images [YSG^{+}06]. imaging [ASI^{+}08]. imitation [Str09]. Immigration [DDLD07, EGE^{+}08, GDL06a, HLK06]. immunization [Sel00]. Impact [AW06a, GGH08, GJ07, HVVK09, JGL07, LHC07, Leg08, zLGH^{+}05, MNPJ03, NHP^{+}06, PDBH02, PS09, PLB^{+}06, WRC^{+}08, ZM08, BHC04a, BHP05, BBT06, CBMP07, CUS01, DGSBG09, DMBM05, EBH^{+}01, ES01, FRZ00, GP09, GIKS08, GAB^{+}09, HPH00, HHC02, Ito07, KW02b, KFR07, KNB08b, MP04a, MG09, MSB08, PK08, PCB07, PGM08, PAdlPS00, RS01, RRB^{+}01, RMS08, SDB03, SHS00, SBHH07, SHM07, TB08, TEMJ06, TNO^{+}09a, VZG05, WR08, Wal04, WTMG09, CMDP^{+}00]. impacted [BRW^{+}05, NKK^{+}07]. Impacts [KH07, CSM^{+}06, DDLB02, HFSH06, JRT02, KP09, KJB07, KBM^{+}03, KWW^{+}09, LKH^{+}08, MHMHKA04, MBK^{+}03, MRiI^{+}07, MMF^{+}09b, MMF^{+}09a, MDJ09, PFR09, PSBJ07, PKS^{+}07, Pra09, SB06b, SBVB05, SW03, TS03, XHC^{+}08, ZCB08]. implement [WHW^{+}05]. Implementation [GHP^{+}09, KKCC06, PKC^{+}01, VVF06, AYK07, CKL^{+}06, LPC05, Pen09, YDS08]. implemented [PC07, SHK^{+}07]. implementing [KDW^{+}09b]. Implications [KKL^{+}06, LCG^{+}09, MSS02, MJ06, PGLS03, PSPD09, TSKP09, AER^{+}07, ALO^{+}01, Ash06, BP08, FA03, GGPBEK07, Han09, ICGÁ05, Kri04a, KRvL^{+}02, MY02a, PV04, RBSJ01, Roe00, VS08b, WHP03]. implicit [BHW^{+}08, PE07, PBA06]. implies [LS04b]. Importance [BNM09, MCQA08, PMH00, RLSZK^{+}08, VH09, WL04, ZK08, BDF^{+}06, BA08, BHP08, CN07b, Esp03, EGE^{+}08, Hal04, Hos06, JBP07, KZH07, LL07a, LBL^{+}08, LOM06, MWWZ01, MWH00, MKiIK07, MVV07, OJD04, SO06]. important [FSJ03, JOBL08, KD05, MXC^{+}04, WS02]. importing [HGD05]. impounded [TMJ^{+}07, VOM06]. improve [Eza05, GGM06, LJR06, NMJ07, VS05]. Improved [JRBS02, BH02, CLTH08, PSVH09, PHBF07, PWH07]. Improving [CB07b, DGRU06, HDB^{+}06, SSK^{+}07, SBM04, Sel00, Sto06, WNW09, Ash06, HKPH08, PE08]. impulse [HA08b, Tud01]. impulse-input [Tud01]. impulsive [NG07, WH09]. in-stream [CBFLS09]. inadequacy [URB06]. inadequate [SHN09]. inadmissible [KE07a]. inbreeding [GHP08]. incidence [PKC06, TM01]. incidental [SHW04]. include [BJH01]. included [SLE07]. includes [DVdB^{+}08, Jor05g, Jor05i]. Including [HFL07, CSKP08, MDH^{+}08, RMR08, SIS^{+}07, SSKN08]. inclusion [PT08, XCW07]. incoming [NP06]. incomplete [WSP08]. Incorporating [GQB05, KTL^{+}05, LRT^{+}08, MDGV09, MGV00, MC08, MFA07, PKS08, PS05, RMF09, WLZ^{+}09, ZS09, CFS09, KKH^{+}08, NCM07, PGFM04, YZS^{+}04]. incorporation [AMRS08, KBB00]. increase [BP07, Bor06, BMT07, CT01, Eti04, RNKG03, RGF00]. increased [RWW07]. increasing [SOK06]. increment [Bra01a, LPD08, MLR08]. incubation [MSB07]. independent [LA07, LLF02, MZWM05, MLR08, PF01, dlS07, vWSH08]. independent-validation [MZWM05]. Index [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano01d, Ano01e, Ano01g, Ano01-45, Ano01-46, Ano01-48, Ano02k, Ano02a, Ano02h, Ano02i, Ano02j, Ano02-63, Ano02-53, Ano02-60, Ano02-61, Ano02-62, Ano03b, Ano03c, Ano03d, Ano03e, Ano03k, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03-69, Ano03-70, Ano03-71, Ano03-77, Ano03-72, Ano03-73, Ano03-74, Ano03-75, Ano03-76, Ano04j, Ano04a, Ano04b, Ano04c, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04-91, Ano04-81, Ano04-82, Ano04-83, Ano04-84, Ano04-86, Ano04-87, Ano04-88, Ano04-89, Ano04-90, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05-57, Ano05-58, Ano05-59, Ano05-60, Ano05-61]. Index [Ano05-62, Ano01l, Ano01b, Ano01c, Ano01f, Ano01h, Ano01i, Ano01j, Ano01k, Ano01-54, Ano01-44, Ano01-47, Ano01-49, Ano01-50, Ano01-51, Ano01-52, Ano01-53, Ano02l, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02-64, Ano02-54, Ano02-55, Ano02-56, Ano02-57, Ano02-58, Ano02-59, Ano03l, Ano03a, Ano03-78, Ano03-67, Ano03-68, Ano04d, Ano04-85, Ano05g, Ano05h, Ano05i, Ano05-63, Ano05-64, Ano05-65, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06-49, Ano06-50, Ano06-51, Ano06-52, Ano06-53, Ano06-54, Ano06-55, KDW^{+}09a, MRT05, MSA^{+}03, NMJ07, RSM05, SMSR00, Szi00, XZZ^{+}05]. India [IKS09, KP09, MRG09, Rai08, RPC^{+}05, RJS^{+}06, SAS06]. Indian [NCM07]. Indicating [MHKW00]. indication [SMSR00]. indicative [Szi00]. indicator [AO08, Bar00, BO07, DFF07, EN08, LTP06, LGS03, RCL06, SS00, WG00]. indicators [CSM^{+}06, GFGZ08, HDH00, JNA^{+}09, KSvOO09, LHK00, LP03a, LP03b, MK00, PWZ^{+}09, Sch00, TKHS^{+}07]. indices [FC04, GSZ08, HTK07, IGP^{+}03, IP00, JBP07, KRN04, KSvOO09, LP03a, LP03b, MP00, Mii00, PHH00, PKG00, RMA08, Ric02, Sch00, Sil00, TKHS^{+}07, YLL^{+}05]. indigenous [AF09, JLAM09]. Indirect [AP02, BWPC06, GSB^{+}06b, KCBS00, NiTT01, NiT04, SGP^{+}06, SGP^{+}07, WBP^{+}07, BPBL00, Kri04a, PP04, SGHG04]. Individual [BMR06, CBP^{+}06, CSF^{+}04, FWS^{+}05, HD00a, HB09a, Jag01, Jag09, MLR08, RLHD01, ZPGG03, AR02, ASJD01, ACJT08, BTPL06, BC01, BMG08, Ber02, Bia03, BSG07, BMR^{+}05, BGWC07, CAB08, CLHB^{+}08, CSKP08, Cha08, CSdZ09, CEK08, DLK01, DJ05, DSD^{+}09, EB07b, FSBD01, Fie04, GDP09, GP07, GLS02, GGB^{+}06, GM05, GS02, GBB^{+}06, HMGK05, HHP06, HB05, Hos06, JT01, KMT09, KTKR08, LS09a, MU02, MW03, MM06c, MBM06, MR00, MAL06, MY02b, NW07, NWH^{+}06, OG08, PK02, PEM06, PE08, Pet02, PBY^{+}03, PdAD^{+}04, PF01, PBG09, PBK03, PAdlPS00, PHWH^{+}09, Rai01, RG05, RHM^{+}05, RV05, SR08, SIK07, SR02, SRS^{+}03, Sti08, TYS^{+}09, TTB06, Uch00, VH07, Van08b, WTST08, WG07b, WH07, Xia00b, XCW07, YTS03, YTH03, ZBWR06, vNLS02]. Individual-based [BMR06, HD00a, HB09a, RLHD01, ZPGG03, ASJD01, ACJT08, BTPL06, BC01, BMG08, Ber02, Bia03, BMR^{+}05, CAB08, CLHB^{+}08, CSKP08, Cha08, CSdZ09, CEK08, DJ05, DSD^{+}09, EB07b, FSBD01, Fie04, GDP09, GP07, GLS02, GGB^{+}06, GM05, GBB^{+}06, HHP06, HB05, Hos06, JT01, KTKR08, LS09a, MU02, MR00, NW07, NWH^{+}06, OG08, PK02, PEM06, PE08, Pet02, PBY^{+}03, PdAD^{+}04, PF01, PBG09, PBK03, Rai01, RG05, RHM^{+}05, SR08, SIK07, SR02, SRS^{+}03, Sti08, Uch00, WTST08, WG07b, WH07, Xia00b, XCW07, YTS03, vNLS02]. individual-level [VH07]. individual-module [AR02]. individual-oriented [MW03]. individuals [BJK09, GRC^{+}07, HB09a, PHP04, Shi04b, Tsc02, Uch00]. Indonesia [OIP^{+}08, AVP08]. Indonesian [PBY^{+}03]. induce [LG04]. induced [DdSG04, EBM06, HLxY04, KVV01, LZZA09, LLA^{+}09b, MVPB02]. inducible [RJGO00, RJFAA07, RJDF08]. inducing [YH04]. Induction [Whi00]. Inductive [ALBA06, WR01a]. Industrial [Jør04a, WZN05, KWD^{+}04, Ski03, YLSH03]. industrialization [DUH03]. industry [RMGR09, Sav00]. IndVal [Pen09]. inequality [WSWL08]. infected [CSG02]. infection [CP02, PCL^{+}05]. infections [RCGB08]. Infectious [RMGR09, Mur06, OVK^{+}06]. Inference [Jør04e, LdVA06, VF07, MLH05, QSB03, SH09a, TH08, US08]. inferior [GLM02]. inferred [YTCV01]. infestans [CAG03]. infestation [CS01, SBR^{+}07]. infestations [BDR06, CTG03]. inflated [BW02, SHN09]. inflation [FRM09]. inflow [LH01]. Influence [AAKO^{+}08, Ano06-47, CZL05, Gau06, JBR07, Kar04, PMC08, SG05, lXzL02a, YBM^{+}05, AP02, BdB05, BTdFC05, BFS05, BCH09, CTG04, CT07b, Dei04, GARB09, Gra07, GG00, HvG07, HNF09, IGP^{+}03, Jag01, Jag09, MDH^{+}09, Med06, MB02a, PHWH^{+}09, RR05a, WBH01, WHP03, lXzL02b]. influenced [FCF^{+}04]. Influences [JCB^{+}02, BPBF^{+}00, MP07, MFG^{+}08, Phi04, YSB08, ZC00]. influencing [Roe00, SGHM01, SBC07, dJTMBGMP^{+}09, TOS09, VPG07]. influential [VPG05]. Info [Jør08a]. Info-Gap [Jør08a]. informatics [PC07, Ano07a, Cha09]. Information [ADSO08, FC04, HTK07, Jør04e, Jor05g, Kir01, MPC06, RSW07, BKC^{+}07, DLG06, For02, For03, HTA^{+}08, JOB04, KMR^{+}07, LS04b, LT06, LJ09, Lud09, MCE^{+}07, PSCMMNS06, PBG09, PLP^{+}04, RS04, SS01, Svi04, TKHS^{+}07, UJF06, WSP08, ZYL06, ZGL08, ZCQ^{+}09]. Information-hierarchical [Kir01]. informational [DN06]. informed [SG09]. Informing [GKT07]. ingestive [ZDR03]. inhabitation [GG02]. inhibiting [PGM08]. Inhomogeneous [Kar03, KNB08a]. initial [KGZR05, Mal01, Wil01]. Initialisation [BJ02]. initiation [RMGR09]. initiative [JFG^{+}08]. injuries [WSF^{+}02]. Inland [YHTH08, KMPB03, LR07a, RBE^{+}08, TN09]. inlets [CUS01]. Inner [JZY07, ZPK^{+}07, XSY^{+}09]. innovative [WL06]. inoculum [PvdBW^{+}02]. inorganic [LR07b, LCR06, MRG09, MPS02, MMPN07]. input [DGD03, GDL06b, KZH07, LH04, MB02b, SGH^{+}08, Tud01, ZvBS05, ZCY09, dCPC01]. inputs [CMDP^{+}00, PLB^{+}06, RMGR09, TPCS06]. Insect [PPP05, ADS^{+}07, BDR06, BJFM06, JJWS08, JZC^{+}07, LT04, PCCL03, SLT^{+}09, WW09, dSMZ09]. insecticide [HC06]. insects [Han02, JDPI07, OWWS01, Sai07, ZZ08]. insight [WKL^{+}03]. Insights [Hoy07, LAXP07, MHZ^{+}06, MCK07]. insolation [YBM^{+}05]. inspired [PC07]. instability [Ale07]. Instantaneous [vLDHP08, Xia06]. Institute [Jør04h]. Institutionalized [SPS03]. Instream [BNM09, GACO04]. Instructions [Ano04-44]. instrument [SPS03]. insular [GLP05, Yos08]. intake [HHK07, WPD04, vLDHP08]. Integral [HS03]. Integrated [HOK^{+}09, KMPB03, PKS^{+}07, ZGL08, vPJK04, BB07a, BG08, BFS03, BFHR05, CdQO06, CCC00, CM06b, GBEB06, GML06, Jan01, KKA^{+}08, LZZ^{+}07, Man00, New09, NJF^{+}08, RSC09, SIS^{+}07, SAL07, TOS09, VMR09, WKZP04, WA02, YDS08, ZRR^{+}08, ZBL03, AN06]. Integrating [BPC07, BMD09, CKA^{+}02, HKHB06, KCNN06, KS08, Liu01, LLLT08, LCM^{+}09a, MOLN06, PGCK04, Sti06, TD06, Vil01, BSR06, NG09, YEMZ03]. Integration [CM03, PHBF07, WMM^{+}07, Zav04, ALM00, ACVP00, MCM^{+}09, Mat06, SM04a, TS03, TLW01]. Integrative [GGP06]. integrity [BNM09, IGP^{+}03, MHKW00]. intelligence [FH08]. intelligence-based [FH08]. intensification [LGS03]. intensities [WZJ08, WBR07]. intensity [AZM^{+}06, CG06b, Lir03, WG00, Yos06a, ZPK^{+}07]. intensive [LWC^{+}07, MLGV00a, MLGV00b]. Inter [Tud01, Hel08, HGR08, WRP07, WVP08]. inter-annual [WVP08]. Inter-comparison [Tud01]. inter-generation [Hel08]. inter-reserve [WRP07]. inter/intra [HGR08]. inter/intra-specific [HGR08]. interacting [NMJ07, Sak09, WBB05, WTL00]. Interaction [Cal05, CM08, EDKF06, Nal01, AF05, ASJD01, AS00b, CL06, DdSG04, DDH01, EWH^{+}02, GDL06b, GVC09, Her08, KPT^{+}09, KFS06, MS00, Pet04b, PGA06, QHM05, Reu05, SNRW06, VJ06, Yos03, ZB04]. Interactions [Ano06-47, BLDN00, HLK06, AGZ05, BG03, CT07b, CN09, DR08, DVPS08, DS01, DJ05, EBM06, GDP09, GML05, GS09, GP06, GAPE06, HRH^{+}05, Her08, HC03b, HG01, JLH01, KIL^{+}03, LAB^{+}05, LLA^{+}09b, MU02, MBLA03, MVV07, NF04, Oku09, OzDBS07, RMF09, Rec03, RSB02, SBB06, SKP^{+}07, SS06a, SLGS00, VLA^{+}06, VIK^{+}08, Wit02, ZYY06, vLDHP08]. interactive [CKA^{+}02, JBB^{+}05]. interannual [IMK^{+}07, Las06b]. interception [AJJ^{+}05, WCC02, XBM^{+}07]. intercepts [RHHM08]. Intercomparison [GAA^{+}05, GZY^{+}06]. Interdecadal [AYK07]. interests [Ano04w, GBSP08]. interface [Aus02]. interference [Hul01]. interior [Pot04, WKZ03]. intermediate [BSR06]. internal [BHP08, HBM04, MMB07, YK00]. International [Ano07a, Ano08a, DŽD06, Jør04h, Jør09a, LMPR06, LBL^{+}08, Ano06-45, Ano09i]. Internet [SBS^{+}06]. interpatch [Eti04, KP00]. interplanetary [KL01]. interplay [DPT09, MBKD02]. interpolating [KHT06]. interpolation [AKMB01, KKW08]. interpret [LOM06, YMT03]. interpretation [FSJ04b, LiWZ00, TSKP09, Van04, WO01a, Zav04, Log02]. interpreted [MJ02, Nie00b]. Interpreting [Lud09]. intersite [PLTT05]. interspecific [KAN^{+}09]. intertidal [DAdSC08]. Interval [GRHS00, GLS07, MBKD02]. intervals [BP08, KM06]. intervention [RIGJM06]. intestinalis [LGL02]. Intra [LL06, Yos06b, ACE07]. Intra-clade [Yos06b]. intra-regional [ACE07]. Intra-specific [LL06, HGR08]. intracolonial [SC07]. Intraguild [RLDL09, Bor06]. Intraspecific [Fie04, BPE^{+}07, PBE^{+}07, lXzL02a]. Intrinsic [Zei04, PS09, dlS07]. introduced [GLP05]. Introducing [HBDA09, MOP^{+}05, SKP^{+}07, BMR^{+}05]. Introduction [LMPP00, LSAGF05, CP04, SGH04, WM02, ZEGS08, Jor05h]. introductions [JBR07]. introgression [TTA^{+}03]. intrusion [MMB07]. inundation [PLC08]. invaded [MCSG06]. Invading [GLP05, KK00b, SMET06]. Invariance [Den08, Ned09]. Invariants [TKSP09]. Invasion [dCBL09, ADdSC06, DDLD07, GL06, GPB01, JLAM09, MMF^{+}09b, MMF^{+}09a, NW07, WW09, Jor06b]. invasions [LD06b, MWH00]. Invasive [CH04, MMF^{+}09b, MMF^{+}09a, VM09, GWK^{+}06, GPB01, KP09, KBM^{+}03, LKH^{+}08, RDS07, SCPC^{+}07, Sil07, SMTR07, Wel04]. invasives [SF07]. inventory [DKL^{+}09, SGH^{+}08, WGT05]. Inverse [GTJ^{+}00, CA04, DM03, JNA^{+}09, PEAS01, VHG05, WNW09, WV05, WVP08, vBBE^{+}08]. inversion [LAZ^{+}08]. invertebrate [AMB07, HHK07, MLGG09]. invertebrates [CGC01, JVL02]. inverted [WMSW09]. invested [LL07c]. investigate [HB05, MDH^{+}09, WS02]. Investigating [LT06, MM06c, SRS^{+}03, TBP^{+}07, GRW04, OB04]. Investigation [KNZ04, MRRJ07, Yun08, AGNLGSG04, CMD06, GHP^{+}09, GGB^{+}06, LKKL09, MNZ06, ML06, NW06, SHZ05]. Investigations [CN07b, Kri04a, Ayd08, SV03b]. investments [PDS07]. invitation [JFG^{+}08]. involving [KIL^{+}03, Nal01]. Iowa [HPA00, LFJ^{+}06]. IPCC [KDW^{+}09b]. IPNV [RMGR09]. Ips [JPD^{+}06]. Irish [KHJ^{+}08, RMGR09]. iron [EPS04]. irradiance [WHP01]. irregular [KM06, Mun09]. irreplaceable [FL07]. irreversible [DPT09]. irrigated [LWC^{+}07, PGM08, SB01]. irrigation [BDD^{+}01, PSPD09]. ISBN [Hey01a, Jor05a, Jor05b, Jor05c, Jor05e, Jor05f, Jor05g, Jor05h, Jor05i, Jor05j, Jor05k, Jor05l, Log02, Nie06, Xu03, Zha06]. iSDM [VM09]. ISEM [Ano06-45, Ano08a, Ano09i, LMPR06, LBL^{+}08, Ano03z, Ano09c, Ano09d, Ano09e, Ano09f, Ano09j, Ano09k, Ano09l, Ano09-31, Ano09-32, Leg01]. ISIS [MP04a, PMD^{+}09]. ISIS-Fish [MP04a, PMD^{+}09]. Iskar [RNC08]. Island [GB01, KSG05, LBS08b, MKS^{+}02, PLP^{+}04, ZVK05, CWCH01, KSG05, Leg03b]. Islands [RPC^{+}05, ICSC05, PHP04, PFBBJ08, Ray08, DD03, ZR04]. isolation [LA04]. isoprene [MSHL00]. isoproturon [BBD^{+}04]. isotope [dSLSD02, YYZ06]. isotopic [YY06]. Israel [BHI^{+}06, HPH00]. Israeli [GGS08]. Issue [Ano06-44]. Issues [HDH00, JFG^{+}08, ÖTÖ06, SS06a]. Italian [MNZ06]. Italy [Ano06-47, ABP05, FMM^{+}07, HMFM00, LBS08b, LMPT05, RMBM06, VTB^{+}08, BD05, CFPV08, FU05, JT09, OPL^{+}09, PBA05, TAL^{+}05, VCRD06]. iterative [KC01]. IV [Ano04-45, PBC02]. Ivlev [WH09]. Ivlev-type [WH09]. IX [Ano04-46]. Izsák [Ric02].

J [Bro04a, Bro09, Jor05k, Jor05l, Jør06a, Ula05]. jack [New09]. James [Jør04b, Ula05]. January [Ano00y, Ano00-36, Ano01x, Ano01-37, Ano02-27, Ano02-39, Ano02-44, Ano03-38, Ano03-51, Ano04-51, Ano04-56, Ano05-32, Ano05-44, Ano06x, Ano06-32, Ano06-31, Ano07n, Ano07v, Ano08m, Ano08t, Ano09m, Ano09-29]. Japan [HOK^{+}09, KCY^{+}08, SWSF06, YHTH08]. Japanese [FH08, Fuk09, Ito07, IOIA08, MLTN06, OYiI09, SWCO07]. japonica [SWCO07, TYS^{+}09]. japonicum [PMH00]. jellyfish [JCX^{+}08]. Jeroen [Jor05l]. Jia [LTLH08]. Jiangsu [CL08b]. John [Jor05a, Jor05b, Jor05g, Jor05k, Mai08]. Johnson [Mar06]. join [MG01]. Joint [WZC^{+}05]. Jørgensen [Bro09, Hey01a, Hey01b, Xu03, Zha06, Cha09].

Jou [LLCL04]. Jou-Jou [LLCL04]. Journal [Ano06-42, Ano06-43, Jør09a]. July [Ano00-42, Ano00-31, Ano01-29, Ano02-38, Ano02-31, Ano03-41, Ano03-52, Ano04-55, Ano04-58, Ano05-37, Ano06z, Ano06-34, Ano07o, Ano07u, Ano08w, Ano09r]. jumbo [RLSZK^{+}08]. June [Ano00-30, Ano00-32, Ano01-41, Ano02-40, Ano02z, Ano03-39, Ano03-50, Ano04-67, Ano04-49, Ano05-38, Ano05-30, Ano06-30, Ano07w, Ano08z, Ano09p, Ano09q]. Junin [PCL^{+}05]. juniper [APEA09]. jury [JG05]. Jutland [Kri03]. juvenile [AGZ05, BDI04, GR09, HDB^{+}06, SR02]. juveniles [Aub04a].

K-select [CDM05]. Kachchh [BVD05]. Kalimantan [PBY^{+}03]. Kalman [PCS03, MCJB08]. Kaneohe [TM05a]. Kaplan [DCP^{+}07]. kappa [FM08]. Karenia [VBFM^{+}08]. Karoo [HRH^{+}05, RHH05]. kauri [CS07]. Kay [Ula05]. KBS [LCM^{+}09a]. Keith [Jor05k]. Kelkit [YKO05]. kelp [MMLR07, Ort08]. kelp-urchin [MMLR07]. Kenneth [Jør04e]. Kenting [CHL08]. Keoladeo [Rai08]. Kern [KTB07]. Kernel [KM06, KDK06, LLS^{+}08, SCBG09]. Kernel-based [KM06, KDK06, LLS^{+}08]. kernels [PLLdCB06]. kernoviae [HXP^{+}09]. Kerr [PSVH09]. kestrel [dFOV07]. Key [HGL^{+}06, Håk09, WZKL09, PGLS03]. keystone [JcLM09, LCP06, VJ06]. Khieo [PSVH09]. kill [FCKH06]. kill-sites [FCKH06]. killifish [MWWZ01]. kimi [CR03]. Kimio [Jør04c]. Kin [BTPL06, MML00, MML02]. kin-facilitation [MML00]. kinetic [BA05, WLB^{+}05]. kinetics [GMPC08, HG07]. King [MKS^{+}02, OG08]. Kingdom [SWCO07]. Kinneret [BHI^{+}06, HPH00, GIZ^{+}03, GHP^{+}09, GFG09]. kite [MBKD02]. Kittiwake [FAB^{+}07, BKMB08, WZC^{+}05]. Kivu [VIK^{+}08]. Klanjscek [San09]. Kluwer [Jør04c, Jør04d, Jør04f]. knotweed [SWCO07]. know [VSGW09]. Knowledge [BR01, BHV06, Tis06, AGD06, ATD^{+}06, ATDK06, ATDK08, CM03, DT03, KBF^{+}08, LCM^{+}09a, ÖÖ04, TD06, XLD01, YEMZ03]. knowledge-based [AGD06, XLD01]. knowledge-driven [TD06]. KNP [Rai08]. kob [MSHP04, MSHP04]. Kobus [MSHP04]. Kohonen [CGC01, GS06, WW09]. Kong [XLZ^{+}04]. koraiensis [AJJ^{+}05]. Korea [JJK^{+}01, KCJ^{+}07, OAL^{+}07, LvGC^{+}04, PL02, PSP^{+}07]. Krakatau [Bey03]. kriging [HSRD09]. Krkonose [WW08]. Krummholz [Cai05, WW08]. Kullback [LT06]. Kuparuk [WVP08]. Kuroshio [KMN^{+}07]. Kushiro [Nak08]. Kuwayama [AFLB09]. Kyoto [Jør04c].

L [Hey01a, Hey01b, Jor05d, Jor05g, Nie00a, RKH05]. L-systems [RKH05]. L. [BKS05a, BMD09, BW04, CML09, CMB05, DDJ^{+}01, DADGA06, GBN^{+}06, KL07, KL09, Kno03, LPD08, LGL02, MSP^{+}08, MWD05, NV03, San09, TBPF08, VCM07, WH07]. labile [DVdB^{+}08]. labor [IWW08]. laboratory [SBL03, SKCM07, VPSG05, dSSGR00]. labour [MMPT06]. lacunarity [FWF07]. lacustrine [BW01b, CGG^{+}05]. Ladoga [RAM^{+}03]. ladybeetles [BvdW04]. Laemophloeidae [JJWF07]. Lagoon [Ano04v, BMT07, CSU04, CCMT09, CT01, CU06, FMM^{+}07, dSA01, AACIS^{+}08, CUS01, CCBB05, CMDP^{+}00, FU05, GBEB06, GML06, GML09, HMFM00, LS09b, LT01, LS06, MM06b, MNZ06, NDM00, PSC^{+}01, PCM^{+}03, PLB^{+}06, SCR03, SPC05, TSZdRR03, TPCS06, VCAS01, VML^{+}06, VCRD06, ZRASC04, ZSPV08, BD05, CSU03, CPB^{+}08, GML05]. Lagoons [Ano06-47, HPF08, VLA^{+}06]. Lagrange [KE07b]. Lagrangian [ATA03, CUF^{+}09, GNA^{+}06, KE07a, SCB^{+}09]. Lagrangian-agent [GNA^{+}06]. Lagune [GCG^{+}07]. LAI [JNM^{+}06, Jon07]. Lake [CWCH01, DWR09, HJ02, MH04, MBMP06, MMF^{+}09b, MMF^{+}09a, RBW09, SR08, TTHH07, AER^{+}07, AQS^{+}07, ATD^{+}06, BGMP06, BCD^{+}05, BHP08, CBS09, CGH^{+}05, FMC^{+}08, GB03, Håk00, HvRIZ08, Jan01, JU01, KFH07, LZZA09, LAXP07, LP03a, LP03b, Lud04, MLH05, MP07, MFB^{+}05, MR00, MR06, NP06, PDL06, PLL04, RvGC^{+}08, RC08, RKS^{+}07, SA07a, dCSB08, SHB04, SD08, TMJ^{+}07, TSJ08, XTDL01, XZZ^{+}05, ZCB08, vNLS02, AG03, AB05a, AB05b, ATDK08, BMBOCR03, BLC^{+}07, BO07, BHI^{+}06, CMM02, CJS^{+}02, CM03, CG00, DWH06, EPTB07, FM07, FMC^{+}08, GIZ^{+}03, GHP^{+}09, GFG09, GZJ06, HPH00, HJZ06, JCB^{+}02, KPK^{+}07, KFJ^{+}09, LvNMvdB04, LAXP07, LMPT05, MR00, MMRLP06, MWH00, NDD^{+}07, ORF01, OBR01, OB04, RAM^{+}03, Sin07, VIK^{+}08, VKB06, ZJBI03, ZJTB03, ZRR^{+}08, ZCB08]. lake-model [AER^{+}07]. lakes [AER^{+}07, BR01, BH03, BB07b, CF09, EJG05, ET04, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, FHE06, FLS09, HB01, HB02b, HB03a, HB03b, HBM04, HS06, Lin01, LH01, MH03, MR06, MR08, MDJ09, MS00, RvGC^{+}08, SMSR00, Sri04, TN09, TB05, TB06a, WWL^{+}05, WR01b, WR01c, vNL^{+}08, vPJK04, RE03]. Lambertian [Che06]. lamproides [RRB^{+}01]. Land [WGVB01, ZGF^{+}05, ACE07, AS03, BBGH07, CV07, CCM02, DS01, DCG01, EBH^{+}01, ECP08, HKB02, HvI01, HHF05, HBCL07, HHKH09, IO02, KWS^{+}07, KW02b, KJB07, KvKV05, KH02, KDW^{+}09b, LHC09, LLLT08, MHvIR00, MLPK01, MSM^{+}08, MCJ^{+}04, MBPS04, MG02, OIP^{+}08, OdKV03, OFK08, PSCMMNS06, PHD04, Pra08, RRB^{+}01, RGL^{+}07, RBC09, Roe01, RVRL05, RBEZ08, SB06b, SFM08, SV02, SV03a, SS08b, TEMJ06, TYZC05, THA02, VZG05, WFM01, WG00, WBvDHZ09, WBTC00, WBN^{+}03, XGL^{+}09, ZSKV05, ZH06, vdBDR02]. land-change [PHD04]. land-cover [PSCMMNS06, RBC09, SS08b]. land-surface [Roe01]. land-use [ACE07, CV07, DCG01, HHF05, KW02b, KJB07, KDW^{+}09b, MSM^{+}08, MG02, OIP^{+}08, RGL^{+}07, RBEZ08, THA02, ZSKV05, ZH06]. landcover [EMdC^{+}01]. LANDFIRE [KRZ07]. LANDIS [HSC^{+}04, MFJM04, Mla04, SM04b, SDS^{+}07, SGLH04, SGH04, SSM^{+}09, YHG04]. LANDIS-II [SDS^{+}07, SSM^{+}09]. LandNEP [ECL^{+}02]. lands [SB01, Urs09]. Landsat [Met03, XSY^{+}09]. Landsberg [XBM^{+}07]. Landscape [BBC03, CBP07, Mat06, MS03, RPC^{+}05, VF07, Wal01, YLJY03, ADS^{+}07, Ash06, AKB07, Bar00, BLJS05, BS04, BB03b, CN07a, CKA^{+}02, CZG05, CWWS01, CP01, DFF07, ECL^{+}02, FF01, FH07b, FBC08, FPS03, FRZ00, GDG^{+}00, Gar04, GGV^{+}06, GH09, GSBN03, GdBD08, Gra07, Gra04, HHL^{+}02, HPHP04, HLR06, HLS06, HB09b, HKL07, JBT^{+}05, KKW08, KPH02, KCD^{+}04, KJB07, Kri03, LTM^{+}04, LL07b, LLLP00, LZB^{+}06, Mat03a, MFJM04, MMHE06, Mla04, MSW^{+}09, MNEB01, PEM06, PLH09, PGFM04, PE02, PG02, Ric00, RBEZ08, SM04b, SDS^{+}07, SDA^{+}03, SS06a, SBM04, SFC05, SBDD04b, SV02, SMSV09, SFCP02, SV03b, SMET06, SVB09, SF04, TTB06, VM06, WZ01, WFM01, WG00, WW03, WK04, XHH^{+}04, XHH^{+}05, ZZLC06, vN02, Jor05i]. landscape-guided [KKW08]. landscape-scale [MFJM04]. landscape-wide [Gra04]. landscapes [AKB07, BWAM09, BvdW04, DFGC04, DLP^{+}09a, DLP09b, DMF07, EB07a, GG04a, GFAD06, Gau08, Gil08, HGT^{+}00, HDBM02a, HDBM02b, HSC^{+}04, KRZ07, KW02a, KJ08, LLA^{+}09a, LAB^{+}05, Mal03, MP08, MJ06, PL04, RBC09, RR07, SDA^{+}03, SS06b, SRK06, SGH04, THJ^{+}03, WÖW05, Wim04, ZC00, ZPO09]. landslide [LLCL04]. landsliding [BTdFC05]. Landuse [PDBJ09, Bar00, SW03]. Language [DLP^{+}09a, MC04, Bro04b, FF01, GGV^{+}06]. Lanice [WGV^{+}08]. Lanier [BO07, OB04]. Lansink [Jør04d]. LAPS [MLA^{+}02]. larch [BLDCM06, Ito05, PAF06]. larch-tree [BLDCM06]. Large [CWA^{+}09, ZEGS08, AKMB01, CGS08, CCM02, DKL^{+}09, DGOZ04, Fat04c, FMC^{+}08, HKB02, HvRIZ08, KPK^{+}07, KLPP07, KH02, KRvL^{+}02, LCG^{+}09, OMHR06, OSS02, OB04, OL09, PBS02, PCWP06, PE08, SFM08, SRK06, Sto06, WAB^{+}07, WGT05, XBM^{+}08b, YHC04, ZK08, ZCB08, ZW06, GS02]. large-scale [CGS08, CCM02, DKL^{+}09, Fat04c, LCG^{+}09, OSS02, OL09, PBS02, SFM08, WGT05, YHC04]. largemouth [HL04]. Larger [LE04, KD05, Yos06b]. larger-bodied [KD05]. Larus [BL01]. larvae [JPB03, JZC^{+}07]. Larval [RHE06, LPM^{+}07, Xia03, XIX^{+}08, dSSGR00]. Late [YJJG09]. Latency [VPB08]. latent [Wan07]. lateral [GVDF07, TFTO07]. Laternula [MKS^{+}02]. latest [GTRF01]. latewood [Mii00]. Latin [XHH^{+}05]. latipes [FH08, Fuk09, MLTN06]. latitude [PSPD09]. lattice [LPP^{+}07, LBS08a, SI06, SNRW06, iT03]. lattices [dCBL09]. law [Aok08, LL03a, RM02, SI06, SWSF06, TSF07, TYLL07, Wil08]. Lawrence [SSH^{+}07]. laws [Til04]. Laxenburg [Jør04h]. layer [BRK07, CSdZ09, FGB08, GVDF07, HC03b, OMRD01, OFK08, Tan02, WFB^{+}08]. layers [BB07a, YL07]. lazy [Cor05]. LCA [BH06]. Leach [MMM02]. leaching [BBD^{+}04, BDK01, GORV06, HKB02, KH02, LFJ^{+}06, SFM08]. leading [BP02]. leads [BKO08]. Leaf [TTJ^{+}09, AMB07, BSJ^{+}02, BMD09, DBD^{+}08, FPCA00, FDP^{+}03, JNM^{+}06, KKH^{+}09, MSHL00, NS08, PvdBW^{+}02, PDS07, SL04, Tan02, VSFM03, VCM07, WLJ00, XWB04, YLLW02, Zav04, ZdlP05, ZMB^{+}08]. leaf-litter [ZMB^{+}08]. LEAFC3 [BMD09, MWD05]. LEAFC3-N [BMD09, MWD05]. Learning [DT03, DŽD06, MLM06, BNM09, CPP00, Cor05, DKRŠS00, Dže01, GBN^{+}06, HTS^{+}07, IVC^{+}08, JDTD06, KA00, KDK06, KFR06b, MSD06, PC07, PDL^{+}08, Rec01a, SPM06, VKB06, Xia04]. learning-extended [KDK06]. leasing [MBLA03]. Least [US08, DAR^{+}07, KC05, LN08]. least-cost [DAR^{+}07, LN08]. Least-inference [US08]. leaved [CKPP03, ZBZ^{+}09]. leaves [Che06, SRL^{+}00]. left [Sai07]. legacy [MMR06, MD04]. Legendre [Nie00a]. legged [BL01, FAB^{+}07]. legislation [BDK01]. Legović [San09]. Lehmann [MCSG06]. lehmanniana [MCSG06]. Leibler [LT06]. Leibniz [MC04]. length [AGZ05, BP03, MMR02]. leopard [BM08, FMRS09]. Lepidoptera [GS02, KL03]. lepidus [ZJG^{+}06]. leptokurtic [ZJC^{+}07]. Leptonycteris [MVGH00]. Leslie [Jør04a, Jor05d, Jor05g, CFCP04, KH00, LPCC05, SKJvdHG06]. Leslie-matrix [SKJvdHG06]. lesser [dFOV07]. Lessons [TFF07, VG08, OB04]. LESV [PTGI09]. lethal [BPC07]. Letter [Fat09b, Fat09a, Mai08]. lettuce [ZHMN02]. leucomaenis [MY02b]. leucoryx [GGS08]. leucospilota [DD03]. Level [PBC02, RPC^{+}05, ALO^{+}01, Bla07b, BKMB08, Bru05, CKA^{+}02, CHW07, DFF07, DP07, DFCM01, DDS^{+}04, EMdC^{+}01, Fat07a, FWS^{+}05, FC05b, Gas05, HPH00, Han09, HHL08, HB05, KKW^{+}07, KE07a, MRK^{+}07, MRiI^{+}07, MLTN06, MDJ09, OMRD01, PLH09, PdMVN09, PAdlPS00, SDD^{+}04, SM06, TCD02, VH07, WL06, WSS^{+}06, XHH^{+}04, YYR^{+}07, ZWCL05, Zue07, dMPVNO07]. levels [ACVP00, CS03, CKL^{+}06, IF07, KST^{+}07, Mon02, New06, PK02, ROQ^{+}09, WLG07, ZM08]. Leymus [JZY07]. Li [Jor06b]. Lian [Jor06b]. library [ATD^{+}06, ATDK06, RvGC^{+}08]. LIDAR [MEJ06]. Lie [ZVLP05]. Life [FPS03, GS08, URB06, BC01, BCH09, CM07b, CCGMJ07, Fie04, Gra04, HB06, Hua03, Jag01, Jag09, KL01, KGZR05, KGA06, NG09, Oli03a, Oli03b, Pat07, PV04, RV05, RG06, SCPC^{+}07, SO06, SBHH07, TMS^{+}07, Wal04, YLJ^{+}01, ZZZ06, ZTF08]. Life-cycle [GS08, CCGMJ07, Oli03a, Oli03b]. life-history [BC01, BCH09, RV05, SCPC^{+}07, SO06]. Lifemapper [SBS^{+}06]. light [Ale08, Che06, DMBM05, HS03, HS06, JAB^{+}06, Lar02, MBM00, PvdBW^{+}02, PDS07, Rob05, SRL^{+}00, SRG^{+}09, TR03, XBM^{+}07, vODFS04]. light-use [vODFS04]. lightly [PNN^{+}08]. lightning [APCdlR07, PMC03]. lightning-caused [PMC03]. lightning/human [APCdlR07]. lightning/human-caused [APCdlR07]. Ligula [LGL02]. like [CL08b, TB05, VATP05, ZvBS05]. likelihood [GP02, TTA^{+}03, War08]. limit [CSG02, HM01]. limitation [BSM08, KH07, Nie09b, RHHM08, Van08b, VVF06]. limitations [CW04, DKL^{+}09, HHF05, KPH02, MJR06, dBWG04, dBWG05]. Limited [PFBBJ08, CSR08, CRC08, DPL^{+}04, DVPS08, MMR02, Mul07, VS08b]. limiters [Aru05]. limits [EPS04]. limnetic [CM05a, Nad07]. limnological [RvGC^{+}08]. Lindeman [BBT06]. Lindenmayer [PS05]. Lindl [Oga09b]. line [EG09, WTST08, WW08]. line-transect [EG09]. Linear [DRDD01, BCCR^{+}02, BPBF^{+}00, ÇDKK05, CIM07, Gau08, GEH02, JKJ^{+}08, KC05, LKR03, LMM^{+}07, LLS^{+}08, MHvIR00, MF02a, Nie08c, RBSG06, SAS06, SMTR07, Vil05, WHZ06, XWB04, YKO05]. linearization [BB08a]. lines [KBK07]. Lingholm [Jør04b]. link [IP00, MKvdW^{+}09, VJ06]. linkage [WPMT07]. linkages [CBS09, DMF07]. linked [HXP^{+}09, JL07, PCS01]. Linking [COB^{+}06, JOBL08, Kin05, LTM^{+}04, LOM06, MDJ09, Nie08c, RBC09, RKS^{+}07, SBB09, TEMJ06, VH09, WMH08, ZGL08, COC04, EPTB07, KS07, MS07, ZLZM02, PBE^{+}07, dG02a]. links [DMF07, JCX^{+}08, PBM^{+}05]. Lion [FG06]. lions [CNG06]. Liptovska [Ond07]. Litter [KKH^{+}08, PLTT05, HvG07, KKH^{+}09, TTJ^{+}09, YTCV01, ZJN^{+}06, ZMB^{+}08]. litter-removed [ZJN^{+}06]. litterbag [PLTT05]. litterbags [ZMB^{+}08]. litterfall [LSM^{+}04, MRG09]. litters [KKH^{+}09]. Little [MMRLP06, SSdlMP^{+}08]. littoral [CF09, PP04]. Liu [Mar06]. live [Svi08]. lived [RLDL09, vdPLG^{+}00]. Liverpool [KHJ^{+}08]. livestock [GGV^{+}09, HRH^{+}05, IMS07, MTKM^{+}06, RR05b]. Living [BBGH08, ALJL03, Aok08, KVPA07, KVPA08, TP08]. lizard [TKTB07]. LM2 [ZRR^{+}08]. LM2-Toxic [ZRR^{+}08]. load [CG00, HOK^{+}09, Hui06, LH01, LGL02, SHCS04]. loading [BKS05a, BHSR01, GBEB06, HBM04, JGL07, LZZA09, RCZ^{+}06, TSJ08, vPJK04]. loadings [ZAM^{+}05]. loads [BHP08, PBA05, ZSKV05]. lobster [GBB^{+}09, XIX^{+}08, ZC07]. Local [Jør04f, Oku09, SCPC^{+}07, SLL^{+}06, AE02, BDG01, CKL^{+}06, DJ05, GGB^{+}06, Gra07, LG04, LCH^{+}00, MC08, MB02a, MB02b, NW07, NHLPA06, PLH09, Phi02, SM04a, SJ06, WWC^{+}07, YTS03]. localized [BPBL00]. locally [SM09]. location [APCdlR07, DSD^{+}08, HBUS02, RJR04]. locations [HD01, Pyk04, RJR04]. locus [HLHZ06, HPD09]. locust [HC06]. Locusta [SGHM01]. locusts [SGHM01]. lodgepole [BPBL00, TK01, WKZ03]. Loehle [Swa06]. Loess [ZSZ06]. log [YPR^{+}05]. log-sech [YPR^{+}05]. LogCauchy [YPR^{+}05]. logic [AÖÇ05, CM03, CH06, GIKS08, LSHG08, MM06b, Met01, Met02, NCM07, Paw00, Pra05b, SGHG04]. logistic [Asp02, BS08a, CG06b, GSB05, MKRH03, MKvdW^{+}09, PF00b, PF00a, Sak07, SMET06, TH08, TF05, TSF07, YSM^{+}06, vHPS02]. logistic-based [TF05, TSF07]. logit [CK07a]. lognormal [YPR^{+}05]. Loire [BG08]. Long [LH09, TC09, WFM01, CWCH09, CKPP03, GvNK09, HRH^{+}05, IWM^{+}06, KPK^{+}07, KFH07, LGR^{+}09, Loe04, LLL^{+}07, LGD01, MGH^{+}05, MBD^{+}00, MVGH00, NW06, OHH07, PDL06, PGFM04, RFA^{+}01, RSB09, RHP07, RLDL09, SSH08, Swa06, TMD04, WKZ03, YTCV01, YCVA01, vdPLG^{+}00]. long-lived [RLDL09, vdPLG^{+}00]. long-nosed [MVGH00]. Long-term [LH09, TC09, WFM01, CWCH09, CKPP03, GvNK09, HRH^{+}05, IWM^{+}06, KPK^{+}07, KFH07, LGR^{+}09, Loe04, LGD01, MGH^{+}05, MBD^{+}00, NW06, OHH07, PDL06, PGFM04, RFA^{+}01, RSB09, RHP07, SSH08, Swa06, TMD04, WKZ03, YTCV01, YCVA01]. longitudinal [GVDF07]. longleaf [CL06, DW01]. loosestrife [Wel04]. Loss [Aub04a, ECHN08, JCE06, Mit09, RHP07, SHZL09, WGS^{+}02]. losses [LPU^{+}07, WSF^{+}00]. Lost [Nie06]. Lotka [MG09, FGFB05, HZ01, LSAGF05, Svi00a, TNO^{+}09a]. Lour [WLZ^{+}09]. lovegrass [MCSG06]. Lovett [Mar06]. low [EPS04, ESZ^{+}00, MLGG09, SM06]. low-chlorophyll [EPS04]. low-level [SM06]. low-smoke [ESZ^{+}00]. lower [AYK07, HP09, JJ00, KKW^{+}07, KMN^{+}07, MRK^{+}07, MRiI^{+}07, ROQ^{+}09, WFHP07, YYR^{+}07, Zue07, ZBW05]. lowland [HKPH08, HZF07, HBRW07, HD00b, KJB07, KLPP07, TMLV07, TFTO07]. Ltd [Jor05a]. luga [MWM02]. Lugano [Kri04b]. lugubris [BLDCM06]. Luhmann [GPP02]. Lumbricus [PBMRE08]. lumped [BG03]. lupus [FCKH06]. Luquillo [HHL08, Lug04, WCHM02, WHZ06]. lusori [LYC08]. Lutra [OPL^{+}09]. Luxembourg [FDCH08]. Lyapunov [Tar08]. Lycaenidae [GS02]. lysimeters [BBD^{+}04]. Lysina [NKK^{+}07]. Lysiphlebus [dSMZ09]. Lythrum [Wel04].

M [Jor05d, Jor05g, Jor05l, Jør09b, KMiIW07]. maccoyii [Wil08]. Maccullochella [TNK04]. MacGillivray [LK03]. Machine [DŽD06, SPM06, DKRŠS00, Dže01, GBN^{+}06, IVC^{+}08, JDTD06, KA00, KDK06, KFR06b, PC07, PDL^{+}08, Rec01a]. machines [GKG05, MAB01]. mackerel [BC01]. macquariensis [TNK04]. macro [ESZ^{+}00, HHKH09, JVL02, MCGO05, PKS^{+}07]. macro- [MCGO05]. macro-invertebrates [JVL02]. macro-scale [ESZ^{+}00, HHKH09, PKS^{+}07]. macroalgae [BHC04b, RM09, TSF^{+}05, YMD08]. macrocopa [ZPGG03]. macroeconomics [Abe04]. macrofaunal [BHAV03]. Macroinvertebrate [LMM^{+}07, CKP^{+}01, CWF03, FDCH08, GSG^{+}04, HRMC01, PKC^{+}01, PLS^{+}06, SHK^{+}07, SM03, TN06, VADV06]. macroinvertebrates [DGGD04, DVGD07, DGD03, Gra05, PSP^{+}07, WYO05]. Macrophyte [MVV07, BB03a, MS00, PNU03, SMSR00, VML^{+}06, WvS06]. macrophytes [HB02b, HS03, HS06, WWL^{+}05, vNSvdBC03]. macroporous [LPU^{+}07]. macroscale [BKS^{+}05b]. macroscopic [HG07]. mactroides [dSSGR00]. maculata [GR09]. Maculinea [GS02]. Madagascar [ASG^{+}05]. made [CT01, TMJ^{+}07]. Madeleine [GCG^{+}07]. magna [BPC07, GR09, PHWH^{+}09]. magnesium [MAK^{+}04]. magnetometer [GPC^{+}09]. magnitude [iWLSN00]. Mahalanobis [FK03]. main [APJ03, CAG03, DDG^{+}05, LLW^{+}06, TC09, dCPC01]. Maina [GB08, Dun08]. Mainau [CWCH01]. Maine [LCG^{+}09, OL09, XIX^{+}08, ZC07]. mainly [SM09]. maintain [LL06, MLGG09, RVWH06]. maintenance [HLHZ06, VPG07]. maize [BHC04a, BDD^{+}01, DP03, ITDD09, NV03]. major [EKBF04, WSY^{+}07, YFLW06]. Making [SPS03, AMRS08, CKA^{+}02, Dei04, EMdC^{+}01, HK01, Mat06, Pet04b, PR07]. malaianus [HCJ^{+}09]. malaria [CSCB04, LPFL09, WBR07]. male [HPD09, Sai07, TTJ07, WC03, YTS03]. Mali [KMPB03]. malignant [CSCB04]. mammal [VPR^{+}09]. mammalian [PSBJ07]. mammals [MB05b, Reu05]. man [COC04, CT01, FCF^{+}04, TMJ^{+}07]. man-influenced [FCF^{+}04]. man-made [CT01, TMJ^{+}07]. Managed [GCLG03, LS01, RGF00, RCZ^{+}06, SMG07, TBP^{+}07, TH06b, ZBZ^{+}09]. Management [Ano09i, NKK^{+}09, OACB09, PR07, SBC^{+}09, AVP08, ASHHRRPE04, BBM06, BSB^{+}09, BWJZ01, BPW^{+}03, BRGS09, BB07a, Bor07, BL04, BIS09, BCLR04, CLM^{+}09, CGY08, CPH00, CHL08, CCM02, CKPP03, CB07c, CK07b, CEK08, CHHP02, DVGD07, DS02, Dew01, DUASCM07, DFF07, ECL^{+}02, Eza05, FC05a, Foh05, FMRS09, GHP^{+}09, GBSP08, GGPBEK07, GDL06a, Gri08, HMBG03, HAA08, HvI06, Her03, JWLA00, Jan01, JDD^{+}03, JW09, KLL^{+}07, KNM09, KBM^{+}03, LCF09, LKP03, LWC^{+}07, LMG08, LPM^{+}07, LH09, LCH^{+}00, MP04a, MWM02, MNPJ03, MGCK^{+}03, MCKNM09, MSHP04, MFJM04, MG05, MCJ^{+}04, New09, OW02, PK08, PLL04, PdAD^{+}04, PCB07, PWH07, PB02, Pra00, Pra05a, PFFM07, RSC09, RF09, Roe00, RP07, SPS03, SC07, Sep00, SSPL^{+}08, Sti06, SVB09, TMJ^{+}07, TDL^{+}07, TB06b, TNK04, TCD02, VREA06]. management [VB06, WRP07, WH04, Wil07, Wil09, XLD01, ZRASC04, Zha00, ZS09, Jor05l]. management-responsive [KNM09]. Managing [KvKV05, LD06b, PSC^{+}01, Man00, Pra09]. mangrove [BH00, MH02, MRG09, PBH^{+}07, RS01, Ray08, VCAS01]. Mangueira [FMC^{+}08]. Manila [BSB^{+}09]. manipulations [SRG^{+}09]. mantegazzianum [NW07]. many [BH01, ZW06]. many-parametered [BH01]. map [EPQdCA06, GL01b, KLPP07, LAPAMD07, LRT^{+}08, PLP^{+}04, RJS^{+}06, dCBL09]. maple [Lar02]. Mapping [APCdlR07, EBR02, FC06, YSM^{+}06, DMF07, JAK^{+}06, KDK06, Las06a, LKY^{+}06, Met03, ÖÖ04, PCC^{+}07, Rem04, RJS^{+}06, SHK^{+}07, WLB^{+}04]. maps [Asp02, CGC01, DGRU06, MCK07, Mun09, PLS^{+}06, RB06, RLR09, SBDD04a, WW09]. maquis [PE02, VCM07]. Mara [Ond07]. Marbled [RHB06, YHC04]. March [Ano00-35, Ano00-45, Ano01-31, Ano01-34, Ano02-47, Ano02-34, Ano03-42, Ano03-53, Ano04-48, Ano05-33, Ano06-29, Ano07t, Ano08u, Ano09-27, Ano09-28]. Marco [Jor05a]. margin [OIP^{+}08, TDdSLS^{+}08]. marginal [KvKV05, vdPvOV00]. margins [KBW00, PDL^{+}08]. mariculture [Odu04]. Marie [Jor05e]. marina [AACIS^{+}08]. Marine [Ano07l, CWA^{+}09, Jør04b, LBS08b, OACB09, XLZ^{+}04, AF09, APHV08, ATAK00, ASP^{+}07, BVD05, BPAB^{+}06, BBM06, BMT07, CGS08, CIM07, Dow05, DDFP07, DJ05, FL09, FWHL06, FYN^{+}07, FPSJ04, FSJ04a, FSJ04b, Gas05, GMPC08, GP09, JAK^{+}06, JGL07, KKW^{+}07, KK04, LTP06, LS09b, MD06a, MDH^{+}09, MRiI^{+}07, MGS^{+}09, MLGG09, PCP07, RBSG06, SPB^{+}06, SIS^{+}07, SHG^{+}08, SKJvdHG06, SY07, SGLRE05, SHS00, TVS00, Tia06a, Tia06b, VTB^{+}08, VL08, WRP07, WIMK07, YYR^{+}07, ZR04]. marinus [MMM02]. maritima [dSA01]. Mark [Jor05b]. marked [GRC^{+}07]. market [CP09]. markets [MMPT06, SYC04]. Markov [Bal00, BAP^{+}03, CT07a, FCH04, FCKH06, JLBS09, KLL01, LBBN06, LL00, LK02, PM09, SHZ05, TA05, Wil09]. Markov-chain [LK02]. Markovian [GS06, Wil07]. marmoratus [YHC04]. Marques [Bro09]. MARS [GSC09]. marsh [MWWZ01, SFU01]. Marsh. [Lar02]. marshes [CCMT09, HUB02, SNF01, WHHH07]. marsupial [Whi00]. Maryland [PV04]. Mass [KE07a, Ort08, PM06a, ZRR^{+}08, BCLR04, DBD^{+}08, DAC^{+}09, FM07, GPL05, HBM04, Håk09, HG07, HC09, LBS08b, MRT05, OSHO09, OW02, OACB09, PK08, RSF^{+}01, RP09, SS08a, TMHJ06, WFB^{+}08, vLDHP08]. mass-balance [BCLR04, FM07, HBM04, Håk09, HG07, LBS08b, OW02]. mass-balanced [PK08]. Masssachusetts [Log02]. masting [ASI^{+}08]. mate [MB02a]. material [ZDR03]. materials [Suh05, Suh06]. Mathematical [ALJL03, AN00, BIS09, DMH^{+}03, MD07, MPS02, Nie08a, Oga09b, TL03, YTT^{+}09, Zha00, dCS05, BK05b, BFHR05, CRC08, CSG02, CP02, DUH03, FCP^{+}07, GG04b, MRS05, NWP06, NS08, PKC06, PGS03, RLFB04, Sch03, SC02, Sri04, SKK^{+}03, Svi08, TR03, TYK03, UZ01, Yun08, Zha07, ZTXD03, dlPP05]. mathematically [SdH02]. mathematics [GJY07, LL00, Nie08b]. Matla [MRG09]. matrices [BL02, DLPN06, GS06, Log08, RSA04]. matrix [BPC07, CR03, CL06, HPHP04, HC05b, Jen00, JM01, KCNN06, Leb05, MSM^{+}07, MJH02, PBHGF07, PMC08, RBR06, SR08, SCPC^{+}07, SKJvdHG06, dSLSD02, TSZdRR03, TBPF08, VH09, Yea04, ZBW05, Log02]. matsutake [YSM^{+}06]. matter [BA05, BHSR01, BJ02, CBMP07, CKN^{+}01, HPH00, HGB04, HMPF05, HE05, IO02, KRK05, KW02a, KGJ08, LGR^{+}09, MH03, MBD^{+}00, MBGP08, NMC^{+}06, PPE^{+}07, PKS^{+}07, SU08, STK00, XRM08, ZTXD03]. Matthews [Jor05c]. Matthias [Jør04b]. maturation [AFTB07, CC00]. mature [LAZ^{+}08, MAA^{+}09]. maturity [FHE06, PEAS01]. maximal [Hui06]. maximise [RHP07]. maximizing [KGBC03, New06, Tsc02]. Maximum [COC04, HC09, PAS06, SSdlMP^{+}08, CMD06, GL01a, Hal04, HA08b, JVL02, LL02b, MMA02, SMR08, War08, XZS^{+}07]. May [Ano00-27, Ano00-38, Ano01-28, Ano02-46, Ano02-28, Ano03-43, Ano03-54, Ano04-59, Ano05-41, Ano07q, Ano07p, Ano08q, Ano08x, Ano09o, EKBF04, RR07, vNS03]. Mayfield [EB06]. mays [NV03]. McClanahan [GB08]. McNary [PHW08]. Meadow [YY06, PCM^{+}03, WZJ08, XZS^{+}07, YYZ06]. Mean [AGZ05, Bra01b, AMK00, Ber02, BP08, EN08, WV05, WX09, YTT^{+}09]. mean-field [Ber02]. means [AFLB09, BL02, CSM^{+}06, HB00b, JJK^{+}01, JKWJ03, KC05, PFR09, SHG^{+}08, WRCB01, WW09]. measure [MHSP^{+}06, WDW00]. measured [ATD^{+}06, JAB^{+}06, VBD06]. Measurement [ZYY06, BS08a]. measurements [ECP08, KW00, LAZ^{+}08, MSL06, TMHJ06, WYT09, XBM^{+}08b]. measures [HKPH08, IP00, LT06, MP04a, PK08, PLH09, Ric02, RA06]. Measuring [PdMVN09, dMPVNO07]. meat [BLBT01, SWB06]. mechanism [Ale07, BDLL06, Roe00, WKL^{+}03]. Mechanisms [PSPD09, HC05a, HKL07, KFS06, PRRB09, Str01a, Str01b, WGB^{+}08, Zav04]. Mechanistic [SL02, CGW^{+}06, HHD01, MBMP06, McG05, MP04b, RPVR03, RCGB08, TKTB07, ZdlP05]. Mecopoda [Har08]. medaka [FH08, Fuk09, MLTN06]. media [CRC08, COB^{+}06]. mediated [HvG07, KMRV07, LLA^{+}09b, PSC04]. Mediterranean [CM07a, CPT09, ECBD09, EFS^{+}03, FPK^{+}07, MSM^{+}07, MDB^{+}06, AEP^{+}04, ATAK00, BC03, BMT07, CM06a, CMDP^{+}00, CMB^{+}02, CMM^{+}07, CSM^{+}06, CPTD08, CTH^{+}00, CMSC01, GBEB06, GTJ^{+}00, IVP08b, KNM09, MVPB02, MCM^{+}09, MCKNM09, MLGG09, PP04, PBV05, PCM^{+}03, PLB^{+}06, RKH^{+}07, SPTP01, SLPP05, SM04a, SMTR07, TS09, VML^{+}06, VCRD06, VSFM03, Zav04, ZdlP05]. Mediterranean-type [TS09]. medium [CML09, Il’08, JU01, XBM^{+}08a, ZW06]. medium-sized [XBM^{+}08a]. medium-term [CML09]. Medvednica [JAN^{+}03]. Meeting [Ano08a, LBL^{+}08]. megafaunal [YJJG09]. megaherbivore [LC04]. Mejillones [Ort08]. Mekong [WAB^{+}07]. Melampyrum [WH07]. melanostictus [OYiI09]. melding [PSCMMNS06]. Melicertus [OG08]. mellifera [AGM^{+}08, WS02]. memoriam [KMiIW07]. memory [DH01, GBA08, Hel08, LLL^{+}07]. Menai [KGJ08]. Mendota [MR00]. mercury [CWH^{+}00, CW01, Lin01, RŽČ^{+}04, SD08]. Meretrix [LYC08]. mesh [OFK08]. mesh-typed [OFK08]. Meso [MMV08, SV03b, HHKH09, MCGO05, PKS^{+}07]. meso- [HHKH09, PKS^{+}07]. meso-pore [MCGO05]. Meso-scale [MMV08, SV03b]. mesocarnivores [MNOS08]. mesocosm [DDH^{+}09, SSS00]. mesofauna [KIL^{+}03]. mesopredator [RLDL09]. mesoscale [HHF05, SB06b, WFM01]. mesotrophic [AQS^{+}07]. meta [HL04, Hen07, ZKH09]. meta-analysis [HL04, Hen07, ZKH09]. metabolic [ZYY09a]. metabolism [LHC09, MB05b, YHTH08, ZYY09b]. metacommunities [MLGG09]. metacommunity [BCH09]. metal [LH01, VTB^{+}08]. metalimnetic [JU01]. metals [BA05, Del04, SBHH07, ZTXD03]. Metamodelling [GORV06]. metaphysiological [OS02, RJGO00]. Metapopulation [HLxY04, MMHE06, Bra01b, BL01, Eti04, EN06, GS06, HGL^{+}06, HHP06, HL03, HZHL05, Hui07, KP00, zLGH^{+}05, Moi04, MJ06, MCK07, Mun09, RL05b, SRRS04, TS09, ZLH06]. metapopulations [BBP08, CL08b, EPQdCA06, Fig09, Gra07, ZLX09]. Metarhizium [SGHM01]. meteorology [AW06b, SE04]. methane [XJM07]. method [AE02, AMD^{+}03, Ash06, Bra01a, CDM05, CT07a, CWS09, DAdSC08, DWD07, EFHL05, GNA^{+}06, JNA^{+}09, KM06, KC01, KA00, LE04, LCP06, MK00, MGSdG07, MM00, NDM00, Pen09, RBR06, Ski03, TYK03, Vil05, WCH08, XTDL01, ZLL05, vBBE^{+}08]. Methodological [ÖTÖ06, Mon09a, SS06a]. methodologies [SF09]. Methodology [SSPL^{+}08, Aum07, BFS03, BB03b, CCC04, CSSCC07, Mon02, SV02, SV03a, SRK06, XZZ^{+}05]. Methods [CZG05, KHT06, Ann01, Bul08, CW04, GDL03, GFG09, GL01b, GSC09, HBM04, JDTD06, KRN04, KPAK02, LN08, LLA^{+}09a, LMM^{+}07, LMPT05, MFB^{+}06, MCE^{+}07, OJD04, OHM^{+}06, OLB04, PE08, PT08, PKG00, QSB03, QKR03, SBC07, US08, WCC02, dlPP05]. methylmercury [KMB08]. metric [CABD09]. metrics [BBP08, CSSCC07, IGP^{+}03, Ric00]. Metropolitan [LHC09]. Mexican [MVGH00, PSCS^{+}01, ZRASC04]. Mexico [HTS^{+}07, AS00b, ASAC02, ASHHRRPE04, ASZRMH^{+}04, ASZRRR08, DUASCM07, VCAS01, ALAS09, BMBOCR03, MZASLMLC04, dJTMBGMP^{+}09]. MHYDAS [TMLV07]. MHYDAS-DRAIN [TMLV07]. mice [KTB07]. Michigan [ZRR^{+}08, CJS^{+}02, JCB^{+}02, KFJ^{+}09]. Micro [OLK^{+}04, CBMP07, GvNK09]. Micro-climate [OLK^{+}04]. micro-ecosystem [GvNK09]. micro-organisms [CBMP07]. microalgae [JF00]. microalgal [GSM08]. microarthropods [DDL^{+}06]. microbes [HB09a]. microbial [BMT07, GPDF09, HE09, KA01, MBGP08, PG08, SKM^{+}06, VPV09, VP01, WR03]. microbubbles [ZTXD03]. Microclimate [CTG04, BHW^{+}08]. microcolonial [CGD04]. microcosms [BKS05a, Nie08c, SKK^{+}03, Bro04a]. Microcystis [BP02, JKWJ03, KCJ^{+}07, KM00, RST05, RvGC^{+}08, RH04]. microecosystem [Hoy07]. microhabitat [RLBL01]. microorganism [FL07]. micropollutants [CCBB05]. Micropterus [HL04]. microtopography [Nun03, VMG05]. Microula [LM07b]. mid [PSPD09]. mid-latitude [PSPD09]. Middle [HTS^{+}07]. midge [CFCP04, KH00]. midwestern [LN08]. might [CLBH^{+}09]. migration [AGZ05, DVGD07, GGS08, IA08, KP00, KBB00, Mon09a, MVGH00, OG08, PRRB09, RST05, RJGO00, SCBG09, SJLL08, ZLX09]. migrations [CDM^{+}00, HBSM04, MAdlPR02, OYiI09]. migratoria [SGHM01]. migratory [GCLG03, MBW09a, MMT^{+}07, MSD06, SGHM01]. Mikhailovich [Ano08-27]. mikimotoi [VBFM^{+}08]. Milan [Cor05, Jør01b]. mildew [FK05, RCGB08]. Mill [BAP^{+}06]. mimicry [KAN^{+}09]. mineralisation [CMPO05b]. mineralization [SC01, ZMB^{+}08, ZCKR07]. minimal [DVPS08, MDJ09]. Minimise [RHP07]. minimization [WFT05]. Minimum [MMR02, JU01, LL03a]. mining [BHV06, CM03, LBS^{+}06, Sto06, SZL^{+}04, SBC07]. Minjiang [LWLZ06]. Minnesota [BJJ06, MFJM04, WGT05]. minor [Alo04]. Miq. [HCJ^{+}09]. Mire [Nak08, LRJ^{+}09]. misrepresent [MP08]. missing [IC02, SYC04, Xia05a]. missions [KL01]. Mississippi [ZBW05, JJ00, JRT02]. mist [KK07]. mistletoe [RG06]. Mitch [KW02b]. mite [Ski04, WS02]. mitigation [BFOS08]. mix [Don06]. mixed [ATA03, CSdZ09, DMRP07, HN09, HNF09, KS08, KLL01, LH05, LH06, LvGC^{+}04, MAA^{+}09, MC01, PB02, SL02, SLS03, Sch03, SSM^{+}09, WFB^{+}08, WRB08, WHZ06, ZKH09]. mixed-conifer [HNF09]. mixed-severity [SSM^{+}09]. mixed-species [DMRP07]. mixedwood [LAD06, SPM^{+}08, VSHC08]. mixing [BGMP06, CF09, MSTK08]. mixotrophy [SHS00]. mixture [AG01, Jua09, SAS06]. mixtures [LCF09, LWL^{+}02]. mobilities [KKH^{+}09, Nal01]. mobility [Nal00]. mode [AVTP05, VATP05]. Model [Ano04v, BGF00, BG01a, BvdW04, CSU04, CM07a, CCJ07, CG06a, CG07, DDF^{+}05, Dun08, Fat09b, IWM^{+}06, Jør04f, JKPL09, KL09, KBE^{+}06, KNB08b, LH06, LMMK09, LGR^{+}09, Mat06, MLGV00a, PWSY07, fPzWhSY07, RLHD01, RPRV09, RP09, RF04, SPS03, San09, SIK07, SWK^{+}05, SKvdW09, Str01a, TTA^{+}03, WLL^{+}08, WMM^{+}07, ZRR^{+}08, ZYY09a, ZAM^{+}03b, AF05, AR02, AF09, ARL^{+}06, ASJD01, AMSW07, AYK07, AOY02, Ale07, ADdSC06, AS00a, AER^{+}07, AP00, AA05, Ano03-27, AMK00, AJJ^{+}05, AAO06, AMSH08, AB05a, AB05b, AP06, AS00c, AS01, ATDK08, AFLB09, AYKY05, Ayd08, BTPL06, BKO08, BS07, BBM06, BSB^{+}09, BDLL06, BS08a, BT01, BSR06, BKS05a, BC01, BS08b, BMG08, BPW^{+}03, BCPM09, BPC04, BGMP06, BD05, BK05b, BH02, BPBL00, BRGS09, BKPS08, Bio01, Bio03, BH06, Bir01, BMS^{+}08, BDR06]. model [BW01b, BP04, BDI04, BJJ06, Bor06, BWPC06, BO07, BHSR01, BFOS08, BSG07, BH03, BVC^{+}01, BHAV03, BDP^{+}02, BMD09, BBKN03, BJ02, BG03, BFHR05, CSR08, CC00, CAB08, CPH00, CLBH^{+}09, CHSB00, CCBB05, CW01, CAG03, CCG07, CRC08, CM05a, CSKP08, Cha08, CB01, CP02, CSB03, CSCB04, CL08a, CCC00, CJS^{+}02, CKN^{+}01, CGD04, CKL^{+}06, CB07b, CS07, CCM02, CR03, CBP^{+}06, CEP06, COB^{+}06, CRL09, CMG00, CB07c, CHW07, CMCD04, CEK08, CWL05, CWS09, CMPO05a, CMPO05b, CDA08, CWWS01, CTG04, CS08, CGH^{+}05, CN07b, CN09, CC05, CL06, DMBO00, DWH06, DDLB02, DLR03, DD02, DVPS08, DVdB^{+}08, DGGD04, DDL^{+}06, DVJ^{+}08, Dew01, DADGA06, DWR09, DAdSC08, DWD07, DdSG01, DdSG04, Dow05, DDFP07, DCI01, DS01, DLK01, DAR^{+}07, DP03, DP07, DC08b]. model [DFCM01, DH00, DDH01, DUH03, DBBS03, DSD^{+}09, DDS^{+}04, DMRP07, EBH^{+}01, EB07a, ECBD09, EB07b, EM08, EFS^{+}03, Esc05, EMdC^{+}01, EVF^{+}07, EN06, FSBD01, FL09, FG06, FGEG06, FG08, FWS^{+}05, FCP^{+}07, FM07, Fie04, FK05, FGB08, Fle01, FBC02, FAB^{+}07, FLS09, FRZ00, FHT^{+}09, FYN^{+}07, FPSJ04, FSJ04a, FL07, GDG^{+}00, GBEB06, GPDF09, GL09, GQB05, Gar04, GL06, Gas05, GSB^{+}06b, GFGZ08, GBS00, GG02, GGP03, GSG^{+}04, GGP06, GDP09, GGV^{+}09, GBG02, GGB^{+}06, GML06, GS08, GSBN03, zGsLcXwZ09, GG08a, GORJ03, GdBD08, GLM02, GTRF01, GAA^{+}05, GZY^{+}06, GCG^{+}07, Gra04, Gri04, GG00, GS02, Gri08, GBG^{+}03, GL04, GW04, GVC09, GZ05, GPC^{+}09, GZJ06, GSC09, GH07, HB03a, HGB04, Håk09, HB02a, HRH^{+}05, HGD05, HFL07, HLHZ06, HCJ^{+}09, HMGK05, HTS^{+}07, Har08, HBO07, HY07b]. model [HNS08, HP09, HC03a, HTA^{+}08, HB06, HHB^{+}08, HGLLV03, HHD01, HR03, HB00a, HB05, HAS07, HRJ^{+}00, HC06, HJ02, HZF07, HGR08, HLR09, HC09, HTMO06, HJZ06, HZ01, HBSM04, HD01, HVB06, HMFM00, HMF00, HCL00, HG01, IO02, Ito07, IMK^{+}07, JBB^{+}05, JDPI07, JWLA00, Jan01, Jen00, JM01, Jen05, JJWF07, JW09, JMVvDV02, JZC^{+}07, JG08, JBMBP02, JBMBP03, Jon07, JVL02, JRBS02, Jor05g, Jør08c, Kai00, KC08, KBW00, Kar04, KW00, Ken02, KKA^{+}08, KC01, KL03, KM04, Kin05, KBB00, KKW^{+}07, KC05, KDW^{+}09a, KE07b, KKL^{+}06, KCZ^{+}03, KMN^{+}07, KFR06a, KNM09, KBM^{+}03, KWW^{+}09, KH00, KGSB01, KH02, KKH^{+}08, KGNK03, KK04, KDW^{+}09b, LLW^{+}06, LIS07, LS04a, Lar02, LAD06, LHC07, LPU^{+}07, LS04b, LH07, LKH^{+}08, LvGC^{+}04, LPP^{+}07, LBS08a, LHT^{+}08, LC07a, LG03]. model [LELR02, LC04, LH04, LFJ^{+}06, LWC^{+}07, LS08a, LS08b, Lid01, LC07b, Lin06, Lin01, LH01, LRT^{+}08, Lir03, LLF02, LZB^{+}06, LPPS05, LPM^{+}07, LPM^{+}09, LDM00, LB01, LPC05, LZ07, LLLT08, LAZ^{+}08, LUKD06, LdVA06, LBS08b, LWL^{+}02, LL03b, LLL^{+}07, MB08, MGV00, MH02, MVZM05, MKB00, MH03, MH04, MBMP06, MRRJ07, MHMHKA04, MSM^{+}07, MM06b, MBDB09, MTD^{+}09, MNZ06, ML06, MSHL00, MS07, MVPB02, Mar06, MCD^{+}08, MHM^{+}03, MOP^{+}05, MKDV08, MKRH03, MT02, MW03, MXC^{+}04, Mat03a, MSHP04, MM06c, MBM06, MMPT06, MR00, MLPK01, MSS02, McG05, McK00, MSM^{+}08, MCA^{+}06, MRK^{+}07, MRiI^{+}07, MG05, MDB^{+}02, MG09, MWWM07, MLF^{+}06, MBS^{+}09, MR06, MBD^{+}00, MJH02, MMB07, MG01, MRS05, MP04b, MLL^{+}05, MCJB08, MDH^{+}08, MHP^{+}06, Mon09b, MLR08, MDJ09, MWMN06, MTVC05, MVGH00, MY02b]. model [MSL06, MSD06, MSW^{+}09, MSP^{+}08, MS00, MKiIK07, MB06b, MB09, MMJN03, MWD05, MTKM^{+}06, MHZ^{+}06, MNEB01, MOJ01, Mur01, Mur06, MM00, NWP06, Nad07, NMC^{+}06, NDD^{+}07, NiT04, NW06, NK04, NF04, New06, New09, NG09, NHP^{+}06, NWH^{+}06, NJF^{+}08, NH07, OG08, OCK01, Oga09a, OLK^{+}04, Oke04b, OYiI09, OIP^{+}08, ORS^{+}09, Oli03a, OVK^{+}06, ORF01, OBR01, OSHO09, Ond07, OMHR06, OL09, OFK08, PdC06, PLTT05, PBC02, PNU03, PK01, PEM06, PE08, PGCK04, PA09a, PM09, PSC04, PPE^{+}07, PNN^{+}08, dSPdBB08a, dSPdBB08b, PBMRE08, Pen00, PJAZ02, PLD^{+}02, PS05, Pet02, Pet04b, PWC^{+}09, PBY^{+}03, PdAD^{+}04, PF01, PCS01, PP04, PEPB09, PBK03, PMLM08, PG02, PHBF07, PR01, PTGI09, PHWH^{+}09, PBML08, PS01, PMJ08, QHM05, RST05, RHB06, RSF^{+}01, RBSG06, RŽČ^{+}04, RJGO00, RJFAA07, RG05]. model [RNKG03, RBSJ01, RL05a, RL05b, RC08, RVWH06, RR01, RR05a, RKH05, RV05, RHB05, RPPB04, RLR05, RE03, RSW07, Roe01, RVRL05, RMWW07, RWM^{+}07, RCGB08, RHB04, RPC^{+}05, RMR08, RBEZ08, Saa00, SDD^{+}04, SSS^{+}09, Sai07, Sak07, SAH03, Sal06, SA07a, SCBG09, SPTP01, SFU01, SH09a, SC04, SM04a, Sar04, SIS^{+}07, SB01, SCB^{+}09, SHB04, SHB06, SR02, SM04b, SDS^{+}07, SB06c, SC07, SGP^{+}06, SGP^{+}07, SBM04, SA07b, SE04, SMG07, SL01, SOA03, SCPC^{+}07, SRA05, SAR^{+}09, SAI09, SSCS06, SGY01, She06, SOK03, SRS^{+}03, Shi04b, Sil04, SLGS00, Sin07, SMTR07, SKT00, SB00, SB02, SKJvdHG06, SC02, SNRW06, SWCO07, SK01, SFCP02, SS08b, SSKN08, SW03, SM03, SBJ^{+}02, SKCM07, SHP^{+}07, SDL08, dSLSD02, SvL04, Sti08, SVB09, SLT^{+}09, SKK^{+}03, SJLL08, SRR06]. model [SJG^{+}08, SHS08, Svi02, SZ03, SSH08, SHM07, TN09, Tan02, TYT^{+}09, TLW01, TW02, TH08, TSBH09, Tar08, TT05, TDdSLS^{+}08, TMP06, TK08, TEMJ06, TFF07, TBP^{+}07, TVS00, TSF07, TR05, TDL^{+}07, TMLV07, TMD04, TYS^{+}09, TDLP03, TSKP09, TDHO07, TP06, TNO^{+}09a, THJ^{+}03, TSZdRR03, TBPF08, TFM01, TMHJ06, TSF^{+}05, TMM05, TMS^{+}07, TSJ08, TN08, TFTO07, TL03, Tud01, TN06, TYLL07, TTJ^{+}09, Uch00, USK^{+}06, VS08b, VH07, VBD06, VB06, VSG^{+}08, VJ01, VVF06, VCD05, VCP09, Vil05, VTL^{+}09, VCRD06, VPSG05, WLB^{+}04, WNW09, WC03, WTST08, WRCB01, WV05, WVP08, WM00, WKL^{+}03, WKZP04, WG07b, WCH08, WBR08, WIH^{+}09, War08, WT04a, WLB^{+}05, WYO05, WM06, WIMK07, WRC^{+}08, Whi00, WBP^{+}07, WCS00, WBvDHZ09, WSWL08, WW08, Wil03, WS02, Wil08, WSF^{+}00, WSF^{+}02]. model [WR01c, WSP08, WHP01, Wir00, WBTC00, WJL05, WTMY07, WA02, WBS^{+}02, WT04b, WLZ^{+}09, WWC^{+}07, WBR07, XRM08, XS08, XV03, XCW07, XBM^{+}07, XBM^{+}08a, XBM^{+}08b, XHC^{+}08, XHH^{+}04, XHH^{+}05, XZD^{+}06, XJM07, XGL^{+}09, YMT03, YTS03, YK00, YWHW02, YHG04, YDS08, YSG^{+}06, YFH03, YP02, YHC04, YY06, YMD08, YYR^{+}07, Yun08, YL07, ZPGG03, ZPP06, ZH06, ZdlP05, ZKH09, ZLZM02, ZW02, Zha03a, ZJM04, ZZZ06, ZJC^{+}07, Zha07, ZMB^{+}08, ZCB08, ZCG^{+}08, ZWXF05, ZBW05, ZC00, ZCZ04, ZTXD03, ZB04, ZAM^{+}03a, ZAM^{+}07, ZCQ^{+}09, Zue07, dG02a, dG02b, dSA01, dBWG04, dBWG05, dCPC01, dJBPG02, dlFN08, vHPS02, vNLS02, vNSvdBC03, vN02, AB05a, BBC03, BPE^{+}07, NW06, ZPD^{+}08, Jor05h, Jør04e]. Model-based [TTA^{+}03, BHAV03, DDLB02, HR03, MKiIK07, MM00, NHP^{+}06, SOK03]. Model-Building [Jor05h]. Model-Making [SPS03]. Modeled [JL07, BMB07, BMBOCR03, Esp03, MBF09]. Modeling [Ano07l, BW01a, BPA08, BB01, BMF^{+}06, BHC04b, CD05, CBS09, Chu08, CM07b, CV01, CP01, CU06, Dam03, DGSBG09, DLG06, DLL^{+}09, DDH^{+}09, ECL^{+}02, FC05a, FPCA00, GSBM08, GARB09, GBEB06, GPB01, GRPF07, HB09b, HNF09, JM04, JDD^{+}03, JZY07, Jør04b, Jør09b, JRT02, KNB08a, KCY^{+}08, KKW08, KFS06, KR04, KMRV07, LPD08, LLLP00, LAXP07, LFJ^{+}06, LAB^{+}05, LZ07, MGCK^{+}03, MT02, Mat03a, MBCS07, MBD^{+}00, MF02a, MY02a, MB05b, Mul07, MNZC08, OHH07, PBS02, PCW08, PWSY07, fPzWhSY07, PHH00, Rai08, RNC08, RSB02, RMS08, SIS^{+}07, SC01, Sil07, ŠH09b, SGLH04, SGH04, SJM03, TB06a, TMJ04, TYZC05, Van08b, VIK^{+}08, WHHH07, WHH^{+}08, WZJ08, WMSW09, WBB05, WGB^{+}08, WM02, XZS^{+}07, YTH03, YYZ06, ZBD^{+}09, ZvBS05, ZC07, ZMB^{+}08, ZBWR06, ZSPV08, ZSZ06, ZAM^{+}05, ZAM^{+}07, ZCKR07, dBWG04, dBWG05]. Modeling [dGdMPC05, AM02, AY07, Alo04, AÖÇ05, ABV^{+}06, ASN02, Ano04-28, ASS^{+}06, BK07, BKGC05, Bia03, BKS^{+}05b, BLDCM06, BPW00, CMSB07, CJS^{+}07, CFCP04, CMA^{+}06, CZL05, CKN^{+}01, DPL^{+}04, Den08, DGRU06, DBB^{+}08, EDKF06, ESWG02, FK03, FU05, GHP^{+}09, GGM06, GL01a, GPP02, GTJ^{+}00, HG07, HB09a, HvG07, HSV07, JJ00, JCB^{+}02, JLF08, JPD^{+}06, KLL^{+}07, KSÇ^{+}00, Kaz07, KMB08, KFH07, KFR06b, LLH^{+}06, LG00, LXJ^{+}03, LPCC05, MWWZ01, MM06a, MCKNM09, MB00, MMLR07, NP06, NKK^{+}07, NT07, PL02, PGS03, PBC09, PSCMMNS06, PPS08, PAS06, PCL^{+}05, PV04, PFFM07, RBW09, RJR04, ROQ^{+}09, Roe00, RAM^{+}03, SF07, Sca01, SPK^{+}09, SU08, SLL^{+}06, SBMJ09, SRG^{+}09, SJ03, SKM^{+}06, SSNHP08, TB05, Tia06a, Tia06b, TD06, dJTMBGMP^{+}09, Tsc04, UZ01, VM09, VSGW09, VG08, WZ01, WKZ03, WIMK07]. modeling [WL04, Wim04, WD02, XHH^{+}05, YSM00, YTLF08, Zei04, ZGH05, ZBL07, ZK08, ZZ08, ZPD06, ZLW09, vW07, Ned09, Jor05i]. Modell [BLB09, CG07, Jag09, Suh06, Tia06a]. Modelled [vWSH08, DDF^{+}05, Esc05, HvRIZ08]. modellers [Den09]. Modelling [AN06, AW00, ALAS09, Ann01, Ano03z, Ano03-27, Ano04v, Ano06-45, Ano06-48, Ano08a, Ano09i, AHKB01, AB06, ADS^{+}07, AACIS^{+}08, BVD05, BC03, BPAB^{+}06, BBGM05, BBGM06, BCL^{+}09, BGF00, BDR01, BHM^{+}06, BEM00, BLB08, BLB09, BJH01, BMT07, BAP^{+}06, BCD^{+}05, BJFM06, BBGH08, BHP08, CSU03, CSU04, CFPV08, CHSB00, CCBB05, Cha09, CMDP^{+}00, CBFLS09, CM06b, CT01, CT07b, CGH^{+}03, ČKBB06, DKRŠS00, Dam08, DDG^{+}05, DBD^{+}08, DVJ^{+}08, DDF^{+}05, DAC^{+}09, DDŽ06, EIRT00, ESZ^{+}00, EPM^{+}04, FCF^{+}04, FF06, FMRS09, For03, FMC^{+}08, GP09, Gil08, GML05, GIKS08, GS09, GBG^{+}03, GSM08, GB08, GJ00, HPH00, HB03b, HDBM02a, HTS00, Hey01a, HBDA09, Hör03, HBRW07, HPF08, IGP^{+}03, Ito05, ITDD09, JAK^{+}06, Jen01, JKWJ03, JU01, JF00, Jør00b, JF04, Jor05b, Jør05m, Jor05d, Jor05f]. Modelling [Jor05g, JFGN06, Jør07a, Jør09a, JGL07, KBGJS06, KRK05, KMM^{+}00, KBK07, KFN^{+}08, KE07b, KW02b, KL08, KJB07, Kri04b, KK07, KFR07, LN08, LPCZ07, LIS07, LL02a, LMPR06, LAD06, LBL^{+}08, Leg01, LOL03a, LR07b, LS08a, LSHG08, LC07b, LSS^{+}00, LLA^{+}09b, LBBR01, LCM^{+}09b, LMPT05, LLL^{+}07, MI01, MVMA08, MRG09, MBK^{+}03, MMM02, MPRJ04, MCM^{+}09, MOP^{+}06, MNOS08, MFG^{+}06, Met02, MR08, ME07, MMA02, MAG01, MMN08, MB06b, Nal00, Ned09, Nun03, OWWS01, OMA01, OIP^{+}08, OW06, OPL^{+}09, PBA05, PvdBW^{+}02, PL04, PBV05, PCM^{+}03, PLB^{+}06, PBS05, PB02, PLC08, PGG06, RFA^{+}01, RRB^{+}01, RM09, RR05b, Reu05, RLBL01, RG06, RBC09, RBE^{+}08, RMGR09, SSS^{+}09, SPD^{+}07, SB06a, Sav00, SAL07, SBR^{+}07, SS08a, SMM^{+}02, SDDC07, SCP05, SGLRE05, SLZ09, SMET06, SJH05, SvdWB^{+}06, SSS00]. Modelling [Swa06, TH06b, TRDM06, TAP07, TTPS09, TSF^{+}05, TTJ07, VBFM^{+}08, VPR^{+}09, VSFM03, VKB06, VMR09, WPD04, WR08, WZKL09, WGVB01, WTL00, WvS06, iWLSN00, Xia00c, Xia04, YM05, YZS^{+}04, YTCV01, YCVA01, ZSKV05, ZR04, ZGF^{+}05, ZVK05, ZJG^{+}06, ZPO09, ZTF08, dG04a, vNL^{+}08, AE02, AAA00, AWL04, AMB07, ALJL03, ABP05, ARF08, Ash06, ATD^{+}06, ATDK06, ACVP00, Aus02, BBT00, BW02, BDE08, BHL01, BBT06, BH00, BB08b, BBD^{+}04, Bey03, BM07, BP02, Bor07, BSG07, BG08, BIS09, BGWC07, BW04, BB03b, CCC04, CLM^{+}09, CBMP07, CD07, CYHK04, CGLS07, Cha02, CB07a, CDM^{+}00, CMM02, CM03, CDV05, CGHW05, CZG05, CKPP03, CMCD04, CCGMJ07, CMB05, CNG06, DDJ^{+}01, DLP^{+}09a, dSDSM08, DD00, DLPN06, DMH^{+}03, Dun08, DDS^{+}04, Dže01, ES06, EPTB07, Eza05, FPDP06, FEP^{+}04, FG07, Fuk09]. modelling [GORV06, GDL06a, GKT07, GGV^{+}06, GH09, GRC^{+}07, GLS02, GRBT08, GACO04, GvNK09, Gra05, GW08, GB03, HK01, HMBG03, HKB02, HXP^{+}09, HKHB06, HHK07, HOK^{+}09, HKPH08, HLR06, HO01, HG02, HRMC01, HHKH09, IKS09, JWW^{+}07, JKJ^{+}08, JBT^{+}05, JLH01, JNM^{+}06, KID^{+}07, KST^{+}07, KP00, Kir06a, Kir06b, KS08, KE07a, KHJ^{+}08, KVL^{+}09, KS07, KMPB03, LMH05, LRJ^{+}09, LR07a, LMGM^{+}09, LLA^{+}09a, LS09a, LEH06, LHDJ03, Lev00, Li00c, Lin06, LFB07, LT01, LCR06, MD06a, MMV08, Mal01, Man00, MLHT09, MBT03, McK01, Met03, MD07, MZWM05, MF02b, MDB^{+}06, Mon02, MWP07, MPS02, MS03, MBPS04, NK06, NW07, NV04, NI08, OBE^{+}07, OIR^{+}08, OB04, OS02, PK02, PLH09, PLA06, PBSOMG^{+}05, Pen00, PEE09, PG04, Per07, PE02, PDV^{+}07, PF03, PBA06, PBG09, PBK03]. modelling [PKS^{+}07, PBD00, RDS07, Rai01, RHE06, Rec01a, RKH05, RS04, RGL^{+}07, RPPB04, RPVR03, RH04, RKH^{+}07, RCL06, SDD^{+}04, Sak09, Sal06, SM07a, SCAP05, SAS06, SGH^{+}08, SHG^{+}08, Sno08, SK01, SdH02, SBC^{+}09, Sta07, SYC04, SBHH07, SSdlMP^{+}08, TM05a, TTA^{+}01, TCD02, TKTB07, Uus07, VGK^{+}04, Vil01, VCM07, VM06, WFB^{+}08, WPBM06, WWL^{+}05, WHW^{+}05, WRB08, Weg00, WO01b, WR01a, WDR06, WGV^{+}08, WLSE06, WZC^{+}05, WA02, WSZS08, WSY^{+}07, YEMZ03, YWL^{+}05, Zav08, ZDR03, ZJTB03, ZNC^{+}06, ZGL08, ZEGS08, ZBL03, dCS05, dFOV07, dlPP05, vPJK04, vWBV02, vdPLG^{+}00, Jør00b, Jor05k, Nie08a, Xu03, Jor05j]. modelling/reconstructing [LRJ^{+}09]. Models [Ano06-46, BPP09, BB07a, HC03b, Jor05c, Jor06b, LMPR06, Log02, AALM05, AN06, AGD06, ASG^{+}05, APHV08, AS03, AG01, AP02, ALP03, AQS^{+}07, AGNLGSG04, ACJT08, Aru05, ARF08, ALBA06, Aum07, ABM^{+}06, Aus07, BCCR^{+}02, Bal00, BRV09, BGF00, BA05, BLVC03, BAP^{+}03, BHW^{+}08, Ber02, BH01, BRK07, BB04, BPC07, Bla07b, BS04, BSB00, BKB00, BBGH07, BSR04, BCLR04, BMR^{+}05, BMR06, BKMB08, BEF03, BBB03, Bru05, BNTK04, BHV06, Buz08, CRT07, ÇDKK05, CK07a, CLHB^{+}08, CV07, CS03, CH06, CSHY08, CLTH08, CKA^{+}02, CWA^{+}09, CIM07, CWF03, Coh09, CSF^{+}04, CSM^{+}06, CABD09, CWBR01, Cou03, CP09, CGW^{+}06, CHHP02, DSA08, DVGD07, DCP^{+}07, DGD03, DGD06, DKL^{+}09, DGOZ04, DGM08, DJ05, DT03, ECHN08, ECZ^{+}06, ETH^{+}04, ECP08, EFHL05, EPTB07, EFK^{+}03]. models [Epp00, EN08, FRB^{+}05, FF01, FBF^{+}06, FPCA00, FA03, FRM09, FBDM09, FCH04, FCKH06, FCC^{+}00, FH08, FSJ03, FPSJ04, FSJ04a, FSJ04b, GGM06, GPK00, GBN^{+}06, GFAD06, Gau08, GG04b, GDL03, GWK^{+}06, GDL06b, GP07, GB02, GP02, GM05, GS06, GH00, GBB^{+}06, GB08, GZ00, GEH02, GLS07, Håk00, HB02b, HBM04, HMPF05, HE05, HD00a, HFL07, HAA08, HA08a, HA08b, HG07, HDB^{+}06, Hee02, HMG06, HHP06, Hin09, HHM01, HLH^{+}06, HHF05, Hör03, Hos06, HC05b, Il’08, IVC^{+}08, JT01, JKS^{+}05, JJWS08, JG08, JLBS09, Jør02b, JMN02, Jor05i, Jør08c, KPKP06, Kar03, KWS^{+}07, KA01, KNCP04, KCD^{+}04, KTB07, KW05, Kir06a, KTKR08, KCNN06, KFB09, KPS03, KLL01, KSRWG07, KFR06b, LBL^{+}08, LBBC08, LTM^{+}04, LJR06, LEH06, Leb05, Lek07, LCP06, LS09b]. models [LCG^{+}09, LW09, LOM06, Loe00, LJW03, LL00, LCH^{+}00, LPCC05, MDGV09, MDH^{+}09, MSL01, MVMA08, MGL04, MHvIR00, MP00, MEKL08, MU02, MLC05, MC08, MKM^{+}07, MKvdW^{+}09, Mat06, MOLN06, MMR06, MAL06, MBW09b, MBS^{+}09, MF02a, MFA07, Mit09, Mla04, Mog02, Moi04, MFB^{+}06, Mon09a, MDVG09, Mur01, NS08, NKC04, Nie09a, Nie08c, OMRD01, Oli03b, OW05, OW08, OW02, Ort08, OACB09, OSS02, OLB04, OdKV03, ÖÖ04, ÖTÖR06, PK08, PBB04, PRL07, PGW00, PBPZ04, PM06a, PF00b, PF00a, PKS08, PEAS01, PE07, PVS^{+}09, Pet07, PTS^{+}04, PGA06, PF01, PMC08, PPGP08, PBM^{+}05, PCB07, PBG09, Plu00, PHD04, PB02, PE06, RR06, RLHD01, RvGC^{+}08, RM02, Rd06, RHM^{+}05, RHHM08, RLFB04, RBB05, RdQFO07, RPVR03, RBC09, RLLB09, RBR06, Ros09, RAM^{+}03, SR08, SBB06, SBA06]. models [SAK00, SI06, SS06a, Sch00, SBDD04b, SV02, SBL03, Shi04a, SSNB07, SHN09, Ski04, SY07, SP02, Sto06, SDS01, Str01a, Str01b, SB07, SÅ06, Svi00a, Svi08, SGF08, TLW01, TYZ^{+}05, TR03, TYK03, TN01, TIJ^{+}01, TNK04, TNO^{+}09b, TSJ^{+}09, TA05, TVK^{+}08, TF09, TKTB07, UZ01, VL08, VADV06, VS05, Van04, VMG05, VSHC08, WWRZ04, WFHP07, WSCR03, WC07, Wan07, WG07a, WHH^{+}08, WYT09, WMH08, WR01b, Wil05, WZC^{+}05, WHZ06, WBR^{+}06, Xia00a, Xia02, XSY^{+}09, XWB04, Yea04, YM02, YKO05, Zha00, ZA09, vNS03, vODFS04, vdBDR02, vdHGF09, Jor05a]. MODERATO [BDD^{+}01]. modern [DFGC04]. Modification [GVC09, TYLL07, XBM^{+}08a]. modifications [XBM^{+}07, ZSKV05]. modified [BMS^{+}08, CBP07, HFL07, IVC^{+}08, ITDD09, LFJ^{+}06, LA04, RS04, SvdWB^{+}06, TTA^{+}03, ZLL05, dBWG04, dBWG05]. modify [SLPP05]. modifying [HMFM00, KGSB01]. modular [LK03]. modulated [YABM07]. modulation [WJL05]. module [AR02, CCG07, HSC^{+}04, SM04b, SGLH04]. modules [OCK01]. Mogan [ZJTB03, ZJBI03]. Moina [ZPGG03]. moisture [HVVK09, Kir06a, PSPD09, TTPS09, YL07, ZCQ^{+}09]. mold [SWB06]. molecular [Shi04a]. mollusc [CTH^{+}00]. mollusk [dSSGR00]. momentum [KGA06]. Mondego [MMM02, PM06a, SSPL^{+}08]. money [Bog04]. Mongolia [XSY^{+}09, JZY07, RR05b, ZPK^{+}07]. Monitoring [BA08, BCMR03, HSV07, JPD^{+}06, KHJ^{+}08, WSP08, CSSCC07, CPB^{+}08, Han09, KKCC06, PCWP06, TYK03, Whi05, Jor05f]. monoculture [BGWC07]. Monographs [Bar04]. monomictic [RvGC^{+}08]. monsoon [AP02, ZBZ^{+}09]. Mont [PE02]. montane [Hör03]. Monte [Ann01, GTJ^{+}00, KSvOO09, QSB03, WSCR03, WT04b]. Montgomery [PV04]. monthly [AMK00, SA07b, YKO05, ZAM^{+}03a, ZAM^{+}03b]. morhua [SMB^{+}06]. Morin [FEP^{+}04]. morio [ALAS09]. Moroccan [RHE06]. MORPH [Sti08]. morphological [DWD07, GL04]. morphology [SRL^{+}00, WR08]. Morphometric [AGM^{+}08, LGL02]. Mortality [CN07b, FSJ03, BB04, Haw00, JNSS02, KSH^{+}03, MVPB02, MBM06, MWWM07, Nie09b, POF08, RP09, SBR^{+}07, WRC^{+}08, WGT05, WBR^{+}06, Xia06, YTH03]. mosaic [AKB07, DFCM01, FCF^{+}04, GGV^{+}06, PE07, SS06b]. mosaics [vN02]. mosquito [HTMO06, PBS05, SSCS06]. mosquitoes [CMSB07, HNS08, SBC07]. mosquitofish [BMG08]. most [PPE^{+}07, TOS09]. moth [Gra04, LRT^{+}08, MDPC06, NS08, PAF06, TNO^{+}09b]. moths [TTJ07, Wil01, YTS03]. mottled [RG05]. mound [BLDCM06]. mound-building [BLDCM06]. mountain [BPBL00, CV07, CRL09, JPD^{+}06, PL04, RR05b, RGL^{+}07, WBB05, WBTC00, WBN^{+}03, HBO07, LLCL04]. mountainous [LWLZ06, RBEZ08]. Mountains [HHL08, ZNC^{+}06, CZG05, OHH07, WW08]. Movement [RLHD01, WÖW05, BWAM09, BJJ06, DMBO00, FHT^{+}09, GM05, GNA^{+}06, GPC^{+}09, GH07, HB02a, HPHP04, Hos06, LC07a, MCQA08, NHLPA06, NJB^{+}09, SR02, Sil04, SKT00, Ski04, SJM03, TFTO07, WLNW08, iWLSN00, Xia00b, ZJG^{+}06, ZJC^{+}07, JT01]. movements [Ben04, CP01, FCH04, FHT^{+}09, PBH^{+}07]. moving [GB02, Han02, Mon02]. MTB [JAN^{+}03]. MTB/64 [JAN^{+}03]. much [JF00]. Mucropetraliella [LK03]. Muller [Bro09, Jor05j]. Müllerian [KAN^{+}09]. Mulligan [Jor05b]. Multi [BL04, CABD09, DH01, Jor05h, Mon09b, RIGJM06, RSW07, Tan02, VSG^{+}08, ASI^{+}08, ADS^{+}07, BMS^{+}08, BLDCM06, BB03b, CLM^{+}09, CP09, DDL^{+}06, DMH^{+}03, GLS02, Gra04, Gri04, KDW^{+}09a, KFR06a, KNM09, LKP03, LPFL09, LPM^{+}09, MSM^{+}08, MCA^{+}06, MH05, MFG^{+}06, MHP^{+}06, OMRD01, OIR^{+}08, OdKV03, OFK08, ÖÖ04, PMD^{+}09, SA07a, SJ03, SGF08, VA00, Vil01, WH09, Xia07, vW07]. Multi-agent [BL04, DH01, ADS^{+}07, LPFL09]. multi-agents [GLS02]. Multi-Attribute [RIGJM06, BMS^{+}08]. multi-competitor [Gri04]. multi-component [VA00]. multi-criteria [KFR06a, LKP03]. Multi-Disciplinary [Jor05h]. multi-ecosystem [MH05]. multi-fleet [PMD^{+}09]. multi-generational [Gra04]. Multi-layer [Tan02, OMRD01, OFK08]. multi-level [OMRD01]. Multi-metric [CABD09]. Multi-model [Mon09b, MSM^{+}08, MHP^{+}06]. multi-mutualist [CP09]. Multi-objective [RSW07, DDL^{+}06]. multi-paradigm [Vil01]. multi-predator [WH09]. multi-scale [BLDCM06, BB03b, CLM^{+}09, KNM09, MFG^{+}06, OdKV03, SJ03, Vil01, vW07]. multi-species [DMH^{+}03, LPM^{+}09, OIR^{+}08, PMD^{+}09, SGF08, Xia07]. multi-spectral [ASI^{+}08]. multi-step [ÖÖ04]. multi-target [KDW^{+}09a]. Multi-year [VSG^{+}08, SA07a]. Multiagent [BLBT01]. multidimensional [GSBN03]. Multifractal [CEMS05, DW01]. multilayer [TYT^{+}09, WL06]. Multilevel [BKMB08, VM07, HH04, IA08, LMMK09, Reu05]. Multimodel [BSM08, Jør04e]. Multiple [MBF09, Pra00, ZLW09, ASJD01, ÇDKK05, CCJ07, FA03, KvKV05, LWBW05, MMT^{+}07, OW05, Pra08, RPPB04, VREA06, WLG07, WSF^{+}00, WSF^{+}02, ZBSA07]. Multiple-pattern [ZLW09]. multiple-pest [WSF^{+}00]. multiple-species [FA03]. multiple-stemmed [ZBSA07]. multiple-year [CCJ07]. multiplicative [WNW09]. Multiscale [DDS^{+}04, SDD^{+}04, HB09b, SV03a]. Multispecies [DUASCM07, GL09, TN08]. multispectral [SAS06]. multitrophic [dG02a, dG02b]. Multivariate [HR03, Kri03, LKY^{+}06, BCM^{+}08, ES06, GS09, Hin09, KŞE04, LEH06, LMM^{+}07, LMMK09, LMPT05, PDL06, ZH06]. Multiyear [HHB^{+}08, KBE^{+}06]. Murray [MSL01, OAAF07]. Murrelets [RHB06, YHC04]. mushrooms [YSM^{+}06]. Mussel [MWMN06, GCG^{+}07, MMF^{+}09b, MMF^{+}09a, RD07, RR05a, SNRW06]. mussels [KTKR08, vNL^{+}08]. mustelid [JCE06]. mutual [LS04b]. Mutualism [Zha03b, DH06, Fat07a, McG05, NF04]. mutualist [CP09]. mutualistic [FGEG06]. mutually [FL07]. Mya [RC06]. mycorrhizae [NF04]. mycorrhizal [Ano03-27, CP09, JBMBP02, JBMBP03]. MyLake [SA07a]. myosuroides [CS01]. myths [LCH^{+}00].

N [Hey01a, Hey01b, KNB08b, NJF^{+}08, TFM01, BMD09, BHI^{+}06, DBB^{+}08, HvI01, HB06, LIS07, MSM^{+}07, MWD05, NJF^{+}08, PWSY07, fPzWhSY07, RMDC04, SC01, SYC04, TFM01, ZMB^{+}08, ZCKR07]. N-flux [NJF^{+}08]. N-relationships [HvI01]. Nadarajah [Cha07]. nakai [AJJ^{+}05]. Nakdong [JJK^{+}01, JKWJ03, KCJ^{+}07, PL02]. Nakicenovic [Jør04h]. Namaqualand [HRH^{+}05]. Namibia [HB00a]. napus [BHM^{+}06, CBP07, CML09, HC09, PvdBW^{+}02]. Narew [KLPP07, Pen09]. narrow [HB09b]. Narva [LR07b]. natality [GGS08]. National [BLDN00, BCDJ00, IKS09, MSHP04, PS01, Rai08, RMBM06, RSB09, VB06, ZCY09, BCS09, MLPK01, PSP^{+}07, WTS^{+}06]. native [MCSG06, MBD^{+}00, ZLO02, BS08b]. native-range [MCSG06]. Natke [Jor05h]. Natural [Jor05j, MFG^{+}08, Str01a, Str01b, ABR05, BCST05, GBSP08, GG08a, GPP02, GAPE06, HOK^{+}09, KNB08a, Li00c, MNEB01, OMA01, RLBL01, Roe01, Sal02, SF07, WB00, Wil09, ZS09, ZBWR06]. natural/human [GPP02]. naturalised [ZVK05]. Nature [JAN^{+}03, BCRSTVG07, COC04, Cam04, HVVK09, LS09a, LL08, NM09, Ula09]. NDVI [Bir01, Las06b]. near [CMB05, IiTS^{+}04, JRT02, KŞE04, LGC07, LC01, LCM^{+}09b, Ort08, PL04]. near-bed [LGC07]. near-surface [KŞE04, PL04]. nearby [AW06a, HUB02]. nearly [MHvIR00]. NEATS [MSC09]. Nebojsa [Jør04h]. Nebraska [FHT^{+}09]. necessary [Fuk09, Mit09]. Necrosis [RMGR09, Mur06]. NEE [DDF^{+}05]. need [Pet07]. needed [VM09]. needle [MBK^{+}03]. needles [MAK^{+}04]. Needs [Liu01, LBBC08]. Negative [Hui07]. Negev [GGS08]. negotiating [GBSP08]. negotiations [CT07b]. neighborhood [ZBWR06]. neighbors [VPG05]. neighbour [BL02, DLPN06]. neighbourhood [BH00, GLM02]. neighbouring [NBP05]. neighbours [Bir06]. nematode [MGS^{+}09]. nematodes [MH07, SB06c, TRDM06]. NEMURO [AYK07, BK07, FYN^{+}07, HY07a, KKW^{+}07, KK04, RMWW07, RWM^{+}07, WIMK07, YYR^{+}07]. NEMURO-based [FYN^{+}07]. NEMURO.FISH [IMK^{+}07, MKiIK07]. Neocalanus [TK08]. Neotropical [SAR^{+}09, LH05, LH06, PGA06]. Nepal [CRM09, KLL^{+}07]. Nest [OW08, LKY^{+}06, ÖTÖR06, PFBBJ08, RJR04, SHW04, VH09]. nest-site [LKY^{+}06]. Nested [ICGÁ05, MSS02]. nesting [OW06]. nests [RJR04]. Net [RGF00, AWL04, Ben04, CWBR01, ECL^{+}02, HHK07, MBK^{+}03, MXC^{+}04, MBF09, New06, TFF07, TDLP03, VHG05, VVF06, WBN^{+}03, iWLSN00, XBM^{+}07, XBM^{+}08a, XBM^{+}08b, YHTH08, ZPD06]. Netherlands [NMC^{+}06, vPJK04]. Network [Ano09-33, DGGD04, Fat04c, Fat07a, GSB^{+}06b, HB00b, HUB02, JU09, LS04b, Lat06, MSC09, RH09, AAU02, AFT09, BFU^{+}09, BBP08, BSR04, BRP^{+}06, CHSB00, CDM^{+}00, CH06, CKP^{+}01, CBD^{+}09, CPB^{+}08, DC08a, FSUH07, Fis09, GKT07, GDL03, GDL06b, HXP^{+}09, HB00a, IGP^{+}03, IOIA08, JJK^{+}01, JNA^{+}09, JLBS09, Jør07c, Kaz07, KST^{+}07, KSvOO09, KHLS07, LHDJ03, LKLK07, LBBN06, LLLP00, LCY09, LCG^{+}09, MH05, MD06b, NK06, NU00, NT07, OHM^{+}06, Ort08, ÖTÖR06, PCWP06, PLA06, Ray08, RM09, SFU01, Sca01, SF09, SM09, SG09, SMG07, SBMJ09, TSKP09, UZ01, VLD02, WRCB01, WO01b, WBP^{+}07, WR01c, WFT05, YSG^{+}06, ZBL07, ZZ08, ZYY09a]. Network-based [RH09]. Networks [JKJ^{+}08, AMSM06, AMD^{+}03, AMM04, AB03, ABS05, ABB06, AKMB01, BFU^{+}09, BCCR^{+}02, BPBF^{+}00, BGF00, BP03, BDP^{+}02, BGL01, CBFLS09, CB07c, Cor05, CGGE08, DGD06, DLG06, Fat04b, FC06, Gau08, GS06, HO01, HFSH06, JOBL08, JF06, KPT^{+}09, KPKP06, KSÇ^{+}00, KZH07, KST^{+}07, KKCC06, LJ07, LS02a, LMM^{+}07, Len07, LM07a, MZWM05, MRE^{+}06, MAB01, NSEDP06, OWWS01, OAL^{+}07, OJ02, OJD04, ÖTÖ06, PKC^{+}01, PCCL03, PRL07, PBSOMG^{+}05, PWH07, RMA08, RMDC04, SRW05, SGP^{+}06, SGP^{+}07, SAB^{+}06, SBB09, SWCO07, TB06a, TPJ08, US08, Uus07, VJ06, WRP07]. Netwrk [HB00b]. Neural [CH06, DGGD04, JKJ^{+}08, LHDJ03, PLA06, ZBL07, ZZ08, AMSM06, AMD^{+}03, AMM04, AFT09, AKMB01, BCCR^{+}02, BPBF^{+}00, BGF00, BDP^{+}02, BGL01, CKP^{+}01, Cor05, DGD06, DLG06, FC06, GDL03, GDL06b, HO01, HFSH06, IOIA08, JJK^{+}01, KPKP06, KSÇ^{+}00, KZH07, KST^{+}07, KKCC06, KHLS07, LKLK07, LMM^{+}07, MH05, MZWM05, MD06b, NK06, NT07, NSEDP06, OWWS01, OAL^{+}07, OJ02, OJD04, OHM^{+}06, ÖTÖR06, ÖTÖ06, PKC^{+}01, PCCL03, PRL07, PBSOMG^{+}05, RMA08, RMDC04, SRW05, Sca01, SMG07, SBMJ09, TB06a, TPJ08, VLD02, WRCB01, WO01b, WR01c, WFT05, YSG^{+}06]. Neuroet [NT07]. neurofuzzy [MSL01]. Neuse [BWPC06, GSB^{+}06b, SGP^{+}06, SGP^{+}07, WBP^{+}07]. Neutral [GFAD06, Gau08, LLA^{+}09a, BCPM09, CN07a, KW05]. neutrally [SVS04]. Nevada [CWH^{+}00, HNF09]. newly [TMJ^{+}07, YLJY03]. NH4 [VL08]. Nhue [ABV^{+}06]. Ni [VCD05]. Nicaragua [CR06]. niche [BCM^{+}08, BBGH08, HFL07, HLHZ06, HLxY04, LS08a, PSCMMNS06, PPS08, PSVH09, Sto06, SSdlMP^{+}08, Van08a, GBB^{+}09]. niche-based [SSdlMP^{+}08]. Nicobar [RPC^{+}05]. Nielsen [Bro09, Hey01a, Hey01b]. Niger [KMPB03, SL04]. Nigeria [OI07]. nigra [AMSW07, OMHR06]. nilotica [KBM^{+}03]. nitrate [AFT09, BSM08, EPS04, GORV06, LFJ^{+}06, MBT03, NWP06, PPE^{+}07, SFM08]. nitrate-nitrogen [AFT09]. nitrification [IOIA08, PA09a]. Nitrogen [DP07, LLH^{+}06, LPC05, SPC05, AFT09, BC03, BDR01, BDK01, BWPC06, BMD09, BB07b, CMDP^{+}00, CMPO05a, CMPO05b, DPL^{+}04, GCM^{+}05, GSB^{+}06b, GZ05, HKB02, HOK^{+}09, HFSH06, HNF09, IF07, JMVvDV02, Jon07, KA01, KVPA07, KVPA08, KFN^{+}08, KH02, LBL^{+}08, LR07b, LWC^{+}07, LSS^{+}00, LAZ^{+}08, LCR06, MRG09, MLGV00a, MWD05, MTKM^{+}06, NK06, NMJ07, NK04, PJAZ02, PLD^{+}02, PCM^{+}03, PLB^{+}06, PHBF07, PBML08, SC01, SGP^{+}06, SGP^{+}07, SMM^{+}02, SOK03, SOK06, TR05, TP06, TPCS06, WR08, WGVB01, WM06, WBP^{+}07, WMM^{+}07, YYZ06, YABM07, ZGF^{+}05, ZJN^{+}06]. nitrogen-coupled [ZGF^{+}05]. nitrogen-fixing [KVPA07, KVPA08]. nitrogen-sensitive [MWD05]. nitrous [LL02a]. nivalis [MVGH00]. nivea [OW08]. NK [SGF08]. NLG [Hey01a]. no [BJK09, BKMB08, FCC^{+}00, NJF^{+}08, VL08]. no-till [FCC^{+}00]. nocturnal [PG04]. node [Wil05]. nodosa [ZSPV08]. noise [GLD07, GJ07, KVV01, Par02, VTB^{+}08, lXzL02b, ZB04]. noise-induced [KVV01]. noises [Kar04]. noisy [MPM02]. noltii [PCM^{+}03]. nominally [Xia05a]. Non [BVC^{+}01, CG09, Hul01, JKJ^{+}08, KYH00, LKR03, AF09, BPBF^{+}00, BPP09, BvdW04, BGL01, CIM07, CBSLS07, FDCH08, Hee02, JLAM09, KC03, KLL01, KWW^{+}09, LLS^{+}08, LUKD06, LGL02, OWHS09, PEPB09, PGM08, RBSG06, RHH05, RA06, SB06c, SMTR07, VJ06, Wan07, Wel04, Wil03, Wil07, XWB04, YSM^{+}06, vNM02]. non-additive [vNM02]. non-autonomous [SB06c]. non-Bayesian [Wan07]. non-binary [Wil03]. non-classic [LUKD06]. non-crop [BvdW04]. non-destructive [LGL02]. Non-equilibrium [CG09, RHH05, Wel04]. non-human [CBSLS07]. non-indigenous [AF09, JLAM09]. Non-interference [Hul01]. Non-linear [JKJ^{+}08, LKR03, BPBF^{+}00, CIM07, LLS^{+}08, RBSG06, SMTR07, XWB04]. non-Markov [KLL01]. non-Markovian [Wil07]. non-parametric [Hee02]. non-point [OWHS09]. non-polluted [FDCH08]. non-probabilistic [RA06]. non-renewable [BPP09]. non-sedentary [PEPB09]. non-selective [KC03]. Non-spatial [BVC^{+}01]. non-supervised [BGL01]. non-target [KWW^{+}09, PGM08]. non-trophic [VJ06]. Non-unique [KYH00]. non-wood [YSM^{+}06]. nonconservative [KYL07]. Nonlinear [FMM^{+}07, FDP^{+}03, KŞE04, ALM00, BLDCM06, CM04, GVC09, LAZ^{+}08, SDS01, TB06a, VHG05, Wan07, Yun08]. Nonlinearities [Svi08]. Nonlinearity [ZB04, PS09, QXW02]. nonparametric [GSC09, LC07a]. Nonpoint [RBB05, ATAK00]. nonylphenol [MLTN06]. Nordhaus [Jør04h]. normal [HK01]. Nors [Bro09]. north [DADGA06, MB02b, OW02, SBG06, Ano07l, BC01, EFJ^{+}08, ECBD09, FYN^{+}07, FPK^{+}07, HY07a, HL04, KKW^{+}07, LN08, LWC^{+}07, LGR^{+}09, LXP^{+}08, LMG08, MKB00, MRiI^{+}07, Mey04, MLL^{+}05, OYiI09, SRW05, SMW02, SS00, WDW00, Wel04, YLLW02, YJJG09, ZVK05]. north-central [OW02]. north-east [MB02b, SMW02]. North-eastern [FPK^{+}07]. north-western [DADGA06]. northeast [KJB07, ZH06, WZJ08]. northeastern [CGH^{+}03, HHL^{+}02, MFJM04, XHH^{+}04, AN06, RWM^{+}07]. Northern [BBM06, LGD01, AYK07, AS00b, BPW^{+}03, BRC04, DFCM01, FBDM09, HB00a, HB00b, HSJ04, JLAM09, LRJ^{+}09, Mal01, MC01, Ort08, RJR04, RCZ^{+}06, SBC07, ZGF^{+}05, APHV08, CFPV08, GB01, HZF07, MCJ^{+}04, MZASLMLC04, RSM05, SSNB07, YL07]. Northwest [CWL05, CWS09, RRB^{+}01, Don06, HD01, YABM07]. Northwestern [EFS^{+}03, LMGM^{+}09, MKB00, TK08, ZNC^{+}06, MDB^{+}06]. Norway [BB04, KS08, MAA^{+}09, Mii00, MAK^{+}04, OHH07, SBR^{+}07, WFB^{+}08]. nosed [MVGH00]. Note [Ano01-43, Ano03-66, Ano04-78, Ano04-79, Ano04-80, Ano05-56, CGR03, Ros09]. Nothofagus [JDPI07]. Notonecta [GR09]. Nova [GB01]. novacekii [KM00]. novel [AMD^{+}03, MMB07, PA09a, Sta07, VFTB06]. Novelties [JT09, CCMT09]. November [Ano00z, Ano01-33, Ano01-35, Ano02-50, Ano02-35, Ano02-49, Ano03-47, Ano03-58, Ano04-63, Ano04-50, Ano04-62, Ano05-34, Ano05-31, Ano06w, Ano06-47, Ano07z, Ano08s, Ano08o]. NPP [AOY02, MXC^{+}04, TYS^{+}09, WZJ08]. NPT [TSKP09]. Nuclear [FMP^{+}00, MTVC05]. null [SLPP05]. Number [Jør04c, Jør04d, BP03, Rod07, Sai07, SM09, VPG05]. numbers [Vil06]. Numerical [Alo04, ALM00, BP02, CJS^{+}07, JJWS08, LKKL09, Mal01, MPM02, MSTK08, PBD00, AALM05, ADdSC06, BHI^{+}06, CM06b, DGOZ04, DCI01, KFR06b, MD07, NSO^{+}08, OLK^{+}04, PHW08, SR02, Tud01, VPSG05, WLSE06, Nie00a].

Nutrient [CRC08, FLS09, ZAM^{+}03b, AN00, ATAK00, ASP^{+}07, BBM04, BKS05a, BPC04, BP02, BHP08, CG00, CC05, DR08, EKBF04, EPS04, FMdV01, JWW^{+}07, JRT02, KE07b, KPAK02, LZZA09, LXJ^{+}03, LXP^{+}08, MRK^{+}07, MRiI^{+}07, Mit09, MMN08, MB06b, Mul07, MB07, NDD^{+}07, NKK^{+}09, NV03, OWHS09, PBA05, RL05a, RNC08, RMWW07, RWM^{+}07, Sar04, SU08, ŠH09b, STH06, TTPS09, VJ01, YH04, ZSKV05, ZvBS05, ZAM^{+}03a, ZAM^{+}05, dBWG04, dBWG05, vPJK04, vdPvOV00]. Nutrient-limited [CRC08]. nutrient-phytoplankton [MRiI^{+}07]. nutrient-plankton [MB06b]. nutrient-rich [BP02]. Nutrients [CMSC01, GL01a, BW01b, BP04, BG08, DDH^{+}09, GML05, GBG^{+}03, RFdV07]. nutrition [NMJ07]. nutritional [Ano03-27, JBMBP02, JBMBP03, MCD^{+}08]. NW [BC03, CSM^{+}06, CPTD08, CPT09]. NY [MMR06, BS08b].

O [KBE^{+}06, LIS07, MMB07, MWD05, NJF^{+}08, RMDC04, WFB^{+}08, WM06, vODFS04]. Oak [GKG05, MY02a, NS08]. oasis [SS08a]. obesus [NCM07]. Obituary [Jør08b]. Object [BWJZ01, BDD^{+}01, BB03b, GL04, KVL^{+}09, MKDV08]. object-orientated [KVL^{+}09]. Object-oriented [BWJZ01, BDD^{+}01, GL04, MKDV08]. objective [CPV07, DDL^{+}06, RSW07]. objectives [KvKV05, VREA06]. Observability [Wil09]. observation [BOB07, HHC02, Jør02b, MDB^{+}06]. observations [ESWG02, LFJ^{+}06, LLLT08, MJ02, SSS00]. Observed [CM05b, PPGP08, JVL02, TSBH09, TAL^{+}05]. observers [GRHS00]. Observing [HLR06]. obtain [WKL^{+}03]. obtained [HY07b]. occupancy [APEA09, Moi04, New06, SHN09]. occurred [VBFM^{+}08]. occurrence [APCdlR07, BDF^{+}06, CSCB04, HSRD09, JAN^{+}03, MJR06, ÖTÖR06, PCpDC09, PSBJ07, RLR09, SLPP05]. occurrence-only [HSRD09]. occurrences [PLP^{+}04]. ocean [EPS04, DLK01, FBF^{+}06, FMRS09, NCM07]. oceanic [MMR02, WBH01]. oceanica [EFS^{+}03]. ocher [NWP06]. Ochotona [LZ07]. October [Ano00-39, Ano01z, Ano01-32, Ano01-42, Ano02-51, Ano02-36, Ano03-49, Ano04-60, Ano05-42, Ano06-35, Ano06-48, Ano07x, Ano08k, Ano09v, Ano09w, DDŽ06, DŽD06, JT09]. Octopus [AS00b]. Odum [Ano04-46, BPP09, Bro04a, BU04, Cam04, Com04, Kan04b, Lim04, Lug04, MC04, MD04, Phi04, SH04, Til04, Ulg04, Zuc04]. off [LS09a, LBS08b, MHZ^{+}06, NAC04, PG08]. offer [CBMP07]. offs [ABR05, RVWH06, Sav00, Sil07, vN02]. offsets [FRM09]. offspring [Shi04a]. offstream [BNM09]. OGCM [KMN^{+}07]. OH [KK07]. oil [BCST05, Per07]. oilseed [BHM^{+}06, CBP07, CML09, DSD^{+}08, GL06, PvdBW^{+}02, PDL^{+}08]. Oithona [TSZdRR03]. Okavango [MMWMH07]. old [MS07]. olds [BIS09]. oligotrophic [BC03, BLVC03, ZvBS05]. olive [MSB08, ZBSA07]. Oman [AW06b]. Omnivory [KKZK03]. One [RAI04, ZW06, FGEG06, KVPA08, RJDF08, TSJ^{+}09, VATP05, WH09, YWHW02]. One- [RAI04]. one-compartment [YWHW02]. one-predator [KVPA08, VATP05]. one-prey [RJDF08, WH09]. one-way [TSJ^{+}09]. Oneida [MMF^{+}09b, MMF^{+}09a, SR08]. ongoing [CG06c]. onion [BBT06]. only [HSRD09, MCSG06, OW06, OLB04, VSGW09, ZLO02]. Onondaga [LAXP07]. onset [KRvL^{+}02]. Ontario [Sin07, LLW^{+}06, PMC03]. ontogenesis [KPS03]. Ontogenetic [MGL04]. ontogenic [RMR08]. Onychiruidae [CR03]. OO [RvGC^{+}08]. open [APEA09, LGC07, MSTK08, TF05, TSF07, ZBSA07]. open-channel [LGC07]. open-ended [TF05, TSF07]. open-grown [ZBSA07]. operating [PPS08]. Optimal [ASHHRRPE04, BFD02, CGY08, For02, For03, GLHV08, HRC^{+}07, IWW08, Loe00, MCA^{+}06, MDGK01, New06, RHP07, SD08, She06, SvL04, Wil07, ALP03, Bra01a, Bul08, Don06, Eti04, FAB^{+}07, GRR04, HAA08, HS01, HG02, LJW03, MHvIR00, MEKL08, MP04b, PF03, Sti06, TN08, WTL00, ZS09]. optimally [CGdPRY09]. Optimisation [LHZ^{+}06, MSP^{+}08, RMA08, TCD02, WBR^{+}06, DGD06, LK02, MDVG09]. Optimising [LL07c]. Optimization [DGGD04, HA08a, MCJB08, PFR09, RBSJ01, SV02, SV03a, ZCQ^{+}09, CD07, CFS09, DGM08, KNCP04, LS04b, YTT^{+}09]. optimize [GICB09]. optimized [HTS^{+}07, iT03]. Optimizing [BL01, CC05, HBUS02, HLS06, LBS08a]. Optimum [DS02, Sep00]. options [BFOS08, HHF05, LPM^{+}07, PMD^{+}09, THA02]. oral [Sel00]. Orbetello [FMM^{+}07, GML05, GML06, GML09]. Orbigny [RLSZK^{+}08]. orchard [TTJ07]. orchid [RLR09]. Order [MRRJ06, BCRSTVG07, CS03, Eza05, HBCL07, Lin05, WV05, WVP08, Wir00, WT04b, ZW02]. ordered [BHFMG05, MRRJ07]. Ordinal [Ban05]. Ordinary [Ayd08]. ordination [GL01b, ZHW05]. Ore [OHH07]. Oregon [CWBR01, Gar04, Pet04b]. organ [ARF08, DDG^{+}05, DP07, Oga09a]. organ-specific [ARF08]. organic [BHSR01, BJ02, CBMP07, CKN^{+}01, JKPL09, KRK05, LGR^{+}09, MH02, MBD^{+}00, MBGP08, MHN00, MHSP^{+}06, MTKM^{+}06, NMC^{+}06, PPE^{+}07, PKS^{+}07, SU08, XRM08, YL07, dG04a, vWSH08]. organically [dG04a]. organisational [BMR06]. organising [TYK03, WW09]. organism [ZJC^{+}07]. Organismal [MB07, HB05]. Organisms [PARH07, CBMP07, FMP^{+}00, GJ00, JLDM05, LK03, Mon02, MAB01, NG09, PTGI09, YMT03]. organization [Fis09, Kir01, LCSZ09, RL09, SAK00, WK04]. organizational [Dei04, ZW02]. Organized [Ano08a, AEK^{+}07, Ano06-45, LMPR06, LBL^{+}08, RADD^{+}01, SWBJ01]. organizing [CGC01, GL01b, GS06, KLPP07, LKY^{+}06, LAPAMD07, PLS^{+}06, SHK^{+}07]. organs [DWD07]. oribatid [CM06a, CM07a]. Oribatida [CM07a, CM06a]. orientated [KVL^{+}09]. oriented [BWJZ01, BDD^{+}01, CV07, GL04, HvI01, LR07a, LNW00, MKDV08, MW03, PBB04, PBG09, SGF09, WSWL08, YTS03]. orientor [JMN02]. origin [KYL07]. Origins [TPCS06, Str09]. orographic [WHZ06]. Orthogonal [BLC^{+}07, PT08, Bir06]. Oryctolagus [FSBD01, SBG06]. Oryx [GGS08, GGS08]. Oryza [CMB05]. Oryzias [FH08, Fuk09, MLTN06]. oscillating [GGM08]. oscillation [SKK^{+}03]. oscillations [CAB08, CSG02]. other [Cam04, MVV07, PPP05]. otter [OPL^{+}09]. Oude [Jør04d]. Oulema [SBL03]. Outbreak [BFE07, Zha07]. outbreaks [BP08, Wil01]. outcomes [PE07]. outcrossing [ITDD09]. outer [Hee02]. outflow [HBM04]. outlook [RHM^{+}05]. output [BGWC07, DC08a, IVC^{+}08, LH07, MC08, SMB^{+}06, WLB^{+}04, ZCY09]. outputs [WWC^{+}07]. over-compensatory [MJ06]. Overcoming [URB06]. Overcrowding [ZHL07]. overexploitation [IVP08b]. overfished [KL07, KL09, San09]. overflows [EPM^{+}04]. overgrazing [IMS07]. overland [CBFLS09]. overlap [JcLM09, MNOS08]. overlying [KFS06]. overmature [ZPD06]. overstory [SRG^{+}09]. Overview [Jør08c, SS06a, Ano04-46, Jør05m, KHJ^{+}08, KVL^{+}09, LMPP00, PKG00]. overwintering [ADS^{+}07]. Oviposition [KL03]. owing [Aub04a]. oxbow [dCSB08]. Oxford [Jor05e]. oxide [LL02a]. oxidizing [VPV09]. oxygen [AÖÇ05, AG03, BGMP06, BP04, BHSR01, CPH00, CMDP^{+}00, DFHP04, EPM^{+}04, FMM^{+}07, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, GML09, HHD01, HPF08, JU01, KMM^{+}00, LS06, MLH05, NDM00, PNU03, PSC04, SM03, TL03, WHX^{+}03]. oxygen-mediated [PSC04]. oyster [RR01, WHH^{+}08]. oysters [CMDP^{+}00, TVS00]. ozone [AG01, AAKO^{+}08, BCCR^{+}02, BPBF^{+}00, CPV07, PBSOMG^{+}05, WLB^{+}04, WL06, WBS^{+}02, ŽŽV^{+}06].

P [Jør04e, Jør06a, Nie00a, BHI^{+}06, KFS06, KNB08b, MDH^{+}08, TFM01]. P-cycle [KFS06]. P-dynamics [KFS06]. P-response [MDH^{+}08]. Pacific [Ort08, OACB09, AYK07, Ano07l, CH06, CWL05, CWS09, Don06, DLK01, DLC07, FBF^{+}06, FLS06, FYN^{+}07, HY07a, IMK^{+}07, KKW^{+}07, MRK^{+}07, MRiI^{+}07, MKiIK07, OYiI09, RR01, RWM^{+}07, SCH05, TK08, VREA06, YABM07]. Pacina [JT09]. package [Cal06]. paddy [XJM07]. Page [Ano02-52, Ano03-59, Ano03-64, Ano03-60, Ano03-61, Ano03-62, Ano03-63, Ano03-65, Ano03-80, Ano04-77, Ano04-70, Ano04-71, Ano04-72, Ano04-73, Ano04-74, Ano04-75, Ano05-46, Ano05-47, Ano05-48, Ano05-49, Ano05-50, Ano05-51, Ano05-52, Ano05-53, Ano05-54, Ano06-39, Ano06-40, Ano06-44, Ano06-41, Ano06-42, Ano06-43, Ano03-79]. pages [Jør04b, Jør04c, Jør04d, Jør04g, Jør04e, Jør04h, Jør06a, Nie08a, Xu03, Ano00-33, Ano00-43, Ano00-41, Ano00-29, Ano00-37, Ano00-34, Ano00-44, Ano00y, Ano00-36, Ano00-30, Ano00-32, Ano00-42, Ano00-31, Ano00-35, Ano00-45, Ano00-27, Ano00-38, Ano00z, Ano00-39, Ano00-40, Ano00-28, Ano01-36, Ano01-27, Ano01-40, Ano01-30, Ano01y, Ano01-38, Ano01x, Ano01-37, Ano01-41, Ano01-29, Ano01-31, Ano01-34, Ano01-28, Ano01-33, Ano01-35, Ano01z, Ano01-32, Ano01-42, Ano01-39, Ano02-48, Ano02-33, Ano02-43, Ano02-32, Ano02-30, Ano02-42, Ano02-29, Ano02-41, Ano02-27, Ano02-39, Ano02-44, Ano02-40, Ano02z, Ano02-38, Ano02-31, Ano02-47, Ano02-34, Ano02-46, Ano02-28, Ano02-50, Ano02-35, Ano02-49, Ano02-51, Ano02-36, Ano02-45, Ano02-37, Ano03-44, Ano03-55, Ano03-46, Ano03-56, Ano03-40, Ano03-45, Ano03-38, Ano03-51, Ano03-39, Ano03-50, Ano03-41, Ano03-52, Ano03-42]. Pages [Ano03-53, Ano03-43, Ano03-54, Ano03-47, Ano03-58, Ano03-49, Ano03-48, Ano03-57, Ano04-54, Ano04-64, Ano04-65, Ano04-52, Ano04-68, Ano04-61, Ano04-69, Ano04-66, Ano04-47, Ano04-51, Ano04-56, Ano04-67, Ano04-49, Ano04-55, Ano04-58, Ano04-48, Ano04-59, Ano04-63, Ano04-50, Ano04-62, Ano04-60, Ano04-57, Ano04-53, Ano05-29, Ano05-36, Ano05-35, Ano05-40, Ano05-39, Ano05-43, Ano05-28, Ano05-32, Ano05-44, Ano05-38, Ano05-30, Ano05-37, Ano05-33, Ano05-41, Ano05-34, Ano05-31, Ano05-42, Ano05-45, Ano06-38, Ano06y, Ano06-33, Ano06-36, Ano06-28, Ano06-37, Ano06x, Ano06-32, Ano06-31, Ano06-30, Ano06z, Ano06-34, Ano06-29, Ano06w, Ano06-35, Ano06-27, Ano07s, Ano07m, Ano07r, Ano07y, Ano07-27, Ano07n, Ano07v, Ano07w, Ano07o, Ano07u, Ano07t, Ano07q, Ano07p, Ano07z, Ano07x, Ano08r, Ano08j, Ano08p, Ano08y, Ano08n, Ano08m, Ano08t, Ano08z]. Pages [Ano08w, Ano08u, Ano08q, Ano08x, Ano08s, Ano08o, Ano08k, Ano08v, Ano08l, Ano09-30, Ano09n, Ano09s, Ano09y, Ano09x, Ano09z, Ano09m, Ano09-29, Ano09p, Ano09q, Ano09r, Ano09-27, Ano09-28, Ano09o, Ano09v, Ano09w, Ano09t, Ano09u, Jør04f]. Pagodroma [OW08]. PAH [TVS00]. pair [SYCU09]. palaeoclimates [GTJ^{+}00]. palaeoecological [JDTD06]. palaeological [KLM^{+}02]. pallasi [MRK^{+}07]. palm [CA04, Mat06]. palmata [Lir03]. palustris [CL06]. Panama [DCG01]. Pancreatic [RMGR09, Mur06]. panda [ALO^{+}01, LAB^{+}05]. panther [CP01]. Paper [Jor05b, Jor05k, Zuc04]. Paperback [Nie06, Xu03]. Papers [Ano06-47, Ano06-48, Ano08a, Ano09-33, Ano01m, GH09, LBL^{+}08]. Paphia [SM07a]. Papp [Ric02]. paradigm [KFR06b, MNZC08, Van08a, Vil01]. Paradigms [Nie06]. Paradoxes [GSB05, JG05]. paralic [FG07]. parallel [PE08]. Parallelization [CWWS01]. Parameter [HMF00, Jor01c, KC01, SBL03, WSZS08, WFT05, YYR^{+}07, BS08a, BCPM09, BEF03, BG03, CLTH08, CIM07, CSF^{+}04, CA04, JJWS08, KNCP04, KZH07, Lin06, LAZ^{+}08, MGSdG07, dSPdBB08a, SAH03, SDL08, TW02, ZLW09]. parametered [BH01]. Parameterisation [SKCM07, DP03]. parameterised [BMD09, MWD05]. parameterization [EVF^{+}07, FPK^{+}07, GSZ08, SWK^{+}05, VL08, WBR08]. parameterizations [FL09, Tia06a, Tia06b]. Parameterizing [HHP06, ZPD06]. Parameters [HK01, Hey01b, ABV^{+}06, Ann01, AAO06, BCL^{+}09, BT01, BMG08, BHP05, GP07, GH07, HMFM00, JAK^{+}06, KSRWG07, KK04, Las06a, MGV00, MLGC03, MKM^{+}08, MCJB08, MLGV00a, MAG01, PPE^{+}07, RSC09, RJS^{+}06, She06, SM06, TM05b, VPSG05, WS02, WT04b, YMT03, YSM00, ZWXF05, ZNC^{+}06, ZPW05, ZCQ^{+}09, dCPC01]. parametric [Hee02, RM02]. parametrisation [RVRL05]. parametrization [BPBL00]. Paraná [AA05]. Parasite [MMT^{+}07, SRRS04, WS02]. parasitic [TRDM06]. parasitism [DH06, NF04]. parasitoid [GLL^{+}07, MB02a, NHLPA06, SB06c, TNO^{+}09b, WC03, dSMZ09]. parcel [EMdC^{+}01]. Paria [MHMHKA04]. Park [BLDN00, IKS09, JAN^{+}03, MSHP04, MMR06, Rai08, RMBM06, WZN05, WTS^{+}06, BCDJ00, RSB09, VB06]. Paronychiurus [CR03]. Part [AB05b, CV01, DDG^{+}05, DDF^{+}05, HKB02, Kir06b, XIX^{+}08, AB05a, BPE^{+}07, CMPO05a, CMPO05b, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, Kir06a, KH02, PBE^{+}07, ZAM^{+}03a, ZAM^{+}03b]. Partial [Oli03a, Oli03b, BMT07, GM05, JZC^{+}07, McK01]. participatory [VG08]. particle [CUSZ09, DWH06, Mon05, NMJ07, SR02, TSKP09]. particle-based [SR02]. particle-size [Mon05]. particular [CYHK04, LBL^{+}08]. particulate [BA05, CMG00, HPH00, HGB04, HMPF05, HE05, JLH01, KGJ08, Lin01, MH03]. partitioned [MAB01, OSS02]. partitioning [AO00, CCBB05, PDS07, RD07, Ric05, Uch00, YM05]. partly [MH02]. partners [Sai07]. Partnership [Cam04]. partridge [DFGC04]. parvipinnis [MWWZ01]. passage [LHB08, MBW09a]. passive [Kan04b]. past [BDE08, CWL05, HO01, KTB07]. pastoral [RHH05]. pasture [Gil08, WZKL09, XGL^{+}09, ZVK05, ZBL03, dJBPG02]. pasture-woodland [Gil08]. pastures [GBG02]. Patch [HPHP04, iTI02, APEA09, AMB07, BSR06, BG01a, BG01b, DFCM01, Eti04, KPH02, KP00, LH04, Li00a, Moi04, MSW^{+}09, PHP04, SWBH08, VH09]. patch-dynamic [MSW^{+}09]. patch-matrix [HPHP04]. patch-occupancy [APEA09]. patches [FAA^{+}02, MWWM07, OCK01]. Patchiness [SAI09, Bla07a, HvRIZ08]. patchy [CUSZ09, GGV^{+}06, GFAD06, MP08]. path [AGZ05, LN08, Len07]. pathogen [MBK^{+}03]. pathway [MCD^{+}08]. Pathways [RFdV07, BP03, FH07a]. Patos [dSA01]. Patten [Bro09]. Pattern [Hos06, SS06b, AF05, APEA09, AO08, Bar00, BHFMG05, CV07, CPP00, DS02, FWS^{+}05, GML09, HPHP04, HSRD09, KW02a, KGSB01, LR07a, LYCU09, MdRdA06, ML04, MBM00, OHH01, PE02, PBG09, Rey02, SS06a, SV03a, Svi00a, Svi04, SGF09, SF04, TTA^{+}01, TKTB07, VM01, VCP09, WO01a, Wal01, WG00, WSWL08, ZZLC06, Zha07, ZLW09]. pattern-based [TKTB07]. pattern-oriented [CV07, LR07a, PBG09, SGF09, WSWL08]. Patterning [CKP^{+}01, KPK^{+}07, PLS^{+}06, OAL^{+}07, PKC^{+}01, PCCL03]. Patterns [AN02, KSRWG07, APG^{+}03, AEP^{+}04, AEK^{+}07, ASS^{+}06, ASP^{+}07, BMBOCR03, BBP08, BSTS^{+}02, BKC^{+}07, BPE^{+}07, CLM^{+}09, CSHY08, CML09, CWBR01, DW01, EB07a, GYW^{+}04, Gau06, HGT^{+}00, Hel08, HLK06, JR05, KYH00, KWW^{+}03, Li00a, zLGH^{+}05, LZB^{+}06, LOM06, LGD01, MWK07, MMRLP06, MFJM04, MAHvD08, NP06, NJB^{+}09, PSP^{+}07, PRL07, PA09b, PG02, PMC03, PR01, RB02, RHB05, RBC09, SV02, SV03a, SK04, TSBH09, TTB06, Urs09, Wel04, WDR06, WBR^{+}06, YTS03, YHG04, Yos06a, YJJG09, ZLH06, ZRCA08]. PATUMOD [GBG02]. Patuxent [BBC03]. Paulo [dCSB08]. Paz [ASHHRRPE04, DUASCM07]. PBL [MTVC05]. PCA [Las06b]. PCBs [RBW09]. PCNM [DLPN06]. pCO2 [BSM08]. peach [GBS00]. peak [BKPS08]. peaked [AY07]. peaks [AG01]. peanut [FPCA00, FDP^{+}03]. pear [GL08]. Pearl [DLL^{+}09]. Peary [TFF07]. peat [LAPAMD07, WSS^{+}06, YTCV01]. peat-core [YTCV01]. peatland [GCM^{+}05, YTCV01, YCVA01]. peatlands [Nun03, WSS^{+}06]. pectinatus [CMM02]. Pedodiversity [ICSC05, SI07]. Pedodiversity-area [ICSC05]. pedogenetic [APJ03]. pedological [ICGÁ05]. peduncle [OAAF07]. peeper [Par02]. Peipsi [KPK^{+}07]. pelagic [BS07, CGGE08, DMBO00, DG04b, HE09, LS09a, MSS02, MMRLP06, PG08, RLSZK^{+}08, San07, SSKN08]. pelagics [LCG^{+}09]. Pelicans [CB01, CSB03]. pellets [NWP06]. penaeid [Gri04]. penalized [WA02]. pendulina [RAH07]. Peng [Jor05d]. Peninsula [CDA08, VCAS01, GBN^{+}06, Ort08]. People [Mat06, ÖÖ04]. Perameles [TIJ^{+}01]. Perca [BRW^{+}05, Han09]. perceptron [WL06]. perceptual [PKS08]. perch [BRW^{+}05, Han09, SR08, TTHH07, TSJA02]. perennial [GL04, MMHE06]. perfoliatus [BKS05a, WvS06]. perform [PWSS07, Rd06]. Performance [CCC04, WPMT07, BHC04a, BB04, CL08a, DCP^{+}07, DLG06, EVF^{+}07, FM08, FPSJ04, FSJ04b, GIKS08, JNA^{+}09, KM04, LEH06, MLHT09, MPRJ04, PE08, PF00a, Sch00, TPJ08, TAP07]. period [DWR09, Eza05, GH00, LR07b]. Periodic [MSM^{+}07, MCBA07, Cha00, Il’08, PBC02, VM01]. periodically [WJM^{+}03]. Periodicities [ZPW05]. periods [SSH^{+}07]. periphyton [AS00c, AS01, BCDJ00, FEP^{+}04, LGC07]. permafrost [ASN02]. permanence [LSAGF05]. permutation [HC05b]. Perna [RR05a]. PERPEST [VBD06]. Persistence [JSM03, BHM^{+}06, BBP08, CRT07, Eti04, HL03, Hui07, JBR07, MJ06, Uch00]. persistent [GH00, MHN00, iWLSN00, dG04a, vWSH08]. persisting [BP07, Lin05]. perspective [BVD05, CZL05, DUASCM07, HK02, Li02, OZL07, Pen00, VG08, Bro09]. Perspectives [Plu00]. Perturbation [Bul08, iT03, BS08a, LCG^{+}09]. perturbations [DDLB02]. Perugia [VLD02]. Peruvian [WTMY07]. pest [BB07a, BW04, GP06, GAPE06, LPP^{+}07, PRL07, SB03, TMS^{+}07, WSF^{+}00, WSF^{+}02, ZS09]. Pesticide [SRW05, HSV07, JW09, MBCS07, RS04, VBD06]. pesticide-resistance [RS04]. Pesticides [SBVB05, BDBS08, BH06, GJ00, PGM08]. PestLCI [BH06]. petrel [OW05, OW06]. petrels [OW08]. petroleum [BCST05, BCRP09, CGH^{+}03]. Petrovskii [Jor06b]. PG [CWL05, WCS00, XRM08]. pH [SD08]. PHABSIM [NSO^{+}08]. Phaeocryptopus [MBK^{+}03]. Phaeocystis [CM04]. phase [DP03, IG02, KK07]. phenology [CB07b, Gra04, KWS^{+}07, KBE^{+}06, LWJ06, PWZ^{+}09]. Phenomena [Svi08, ASI^{+}08, HG07, Hua08, PBH^{+}07, SAK00]. phenomenological [SCBG09]. phenotypic [PARH07]. pheromone [BJFM06, SLT^{+}09, YTS03]. philippinarum [BSB^{+}09, PSC^{+}01, VCRD06]. Philippine [DS01]. phillipinarum [SCR03]. Phillyrea [VCM07]. Philosophical [Ano04-46]. philosophy [Sal02]. Phoca [FKIL08]. phoeniceus [ÖTÖR06]. phosphate [JF00]. phosphorus [AP02, BO07, BCDJ00, DWH06, DdSG01, GCM^{+}05, Håk09, HFSH06, LPU^{+}07, LAXP07, MBMP06, MEKL08, MDH^{+}08, MLGV00b, NSEDP06, RSF^{+}01, SHB04, SHB06, TSJ08, WM00]. phosphorus/eutrophication [Håk09]. photoacclimation [SY07]. photographs [JAB^{+}06]. photos [Jor05e]. photosynthesis [BSJ^{+}02, BMD09, DKRŠS00, FDP^{+}03, GQB05, GBS00, Lar02, Wir00, WBN^{+}03, YLLW02]. photosynthesis-stomatal [BMD09]. photosynthetic [Lar02]. Photothermal [BHC04a]. Phu [PSVH09]. Phu-Khieo [PSVH09]. physalus [MDB^{+}06]. Physical [HB02a, Ano04v, BP04, CSU03, CSU04, CJS^{+}02, Dow05, GPDF09, JAK^{+}06, JCB^{+}02, KKH^{+}08, MSP^{+}08, OMRD01, Roe01, ZTXD03]. Physical-based [HB02a]. physical-ecosystem [OMRD01]. physical-microbial [GPDF09]. Physicochemical [SS07]. physics [Ric00]. physiological [BP02, CTG04, FPSJ04, MHM^{+}03]. Physiologically [GP06, GAPE06]. physiology [GQB05, MBK^{+}03, WBN^{+}03, ZdlP05]. Phytophthora [HXP^{+}09]. Phytoplankton [EPTB07, HvRIZ08, MD06a, NWH^{+}06, AN00, ASP^{+}07, BLC^{+}07, BBKN03, BHP08, CRC08, CP02, CSdZ09, DR08, EB07b, EIRT00, ET04, FWSB05, FMC^{+}08, HB03a, HY07b, Hel08, JJK^{+}01, JKJ^{+}08, LR07b, LELR02, LCM^{+}09b, MP07, MRK^{+}07, MRiI^{+}07, MR08, Mit09, MMA02, MB06b, MVV07, NP06, OMRD01, PCpDC09, PRRB09, PWZ^{+}09, RBSJ01, RIE01, Roe00, RMWW07, RWM^{+}07, RAM^{+}03, Sca01, SY07, SDS01, TR03, YH04, vLLv^{+}07]. phytoplanktonic [CPB^{+}08]. phytoremediation [MI01]. Picea [WBB05, ZNC^{+}06]. PICES [BK07]. Picoides [RCL06]. picture [Bro04b]. Piecewise [Nie08c, CM04]. Piediluco [CG00]. Piedmont [FCC^{+}00, Wim04]. pig [CdQO06, HMGK05]. pika [LZ07, RR05b]. pike [TSJA02]. pike-perch [TSJA02]. pilchards [MOJ01]. pilchardus [RHE06]. pin [Dam08]. pin-point [Dam08]. Pinacate [SF07]. Pinder [Jor05g]. pine [APEA09, BPBL00, CRL09, CWL05, CL06, DADGA06, DW01, Fle01, LA07, LSM^{+}04, Mii00, New09, RBE^{+}08, TK01, WKZP04, WKZ03, ZJN^{+}06]. pink [ASZRRR08, GP06]. Pinus [AJJ^{+}05, CL06, DADGA06, GBN^{+}06, KK07, LvGC^{+}04, OMHR06, RNKG03, WRB08]. Piran [Mal01]. PISCATOR [vNLS02]. piscivore [MR00]. pitfall [Den08, Ned09]. Pityogenes [JPD^{+}06]. PIXGRO [ARL^{+}06]. placed [ZMB^{+}08]. placement [CFS09, LTLH08]. Placing [CP04]. Plackett [BH01]. Plain [LWC^{+}07, MLL^{+}05, NSEDP06, ZCKR07, YLLW02]. Plains [BS08b, EBH^{+}01]. plan [BJJ03, FC05a]. Planktic [Rey03]. planktivorous [PGG06]. Plankton [ZRCA08, AP00, CN07b, CN09, DDH^{+}09, FEP^{+}04, JPB03, MPM02, MMR02, MBW09b, MB06b, NDD^{+}07, Oke04a, PCpDC09, SPB^{+}06, SNRW06, SPC05, TSJ^{+}09, VM01, ZCB08]. planktonic [BP02, BP04, IC02, KMN^{+}07, LMPT05, PA09b, SHS00]. Planning [MRE^{+}06, BVD05, BK05a, BLJS05, DLP09b, LTM^{+}04, LNW00, LHZ^{+}06, MKDV08, Pet04a, PdAD^{+}04, Sch00, ZBL03]. Plant [BEF03, Gut07, LC08, Pet02, TH06a, WZN05, APG^{+}03, AEP^{+}04, AY07, AHKB01, Aru05, ABM^{+}06, Bio01, Bio03, BSJ^{+}02, BFHR05, Che06, CS01, Dam03, Dam08, DWD07, DP03, EGE^{+}08, FWS^{+}05, FGB08, FSM^{+}01, GDL06a, GH00, Han02, Haw00, HG01, IWW08, JAN^{+}03, KPT^{+}09, KPS03, KJ08, KBM^{+}03, KS07, LJR06, LA07, LK07, LCSZ09, LUKD06, MMHE06, MRS05, MTVC05, MMJN03, NF04, Oga09a, OzDBS07, PLH09, PvdBFJ02, PCS01, RDS07, RVWH06, RKH05, RMSS02, SFU01, STK00, SLGS00, Ski04, SRK06, SLZ09, SMET06, SGHG04, SF04, Tan02, TYT^{+}09, TT05, TRDM06, TM01, WZKL09, WGVB01, WKL^{+}03, WT04a, Weg00, Wel04, WBTC00, XCW07, YM05, YM02, YABM07, Zha03a, vW07]. plant-parasitic [TRDM06]. plantation [CS07, LA04, SPD^{+}07, TYT^{+}09, ZBL03]. plantations [CMPO05a, CMPO05b, DADGA06, MLF^{+}06, MBS^{+}09]. planthopper [Gri08]. plants [AW06a, ALJL03, ARF08, BGWC07, DdSG01, DdSG04, DP07, GLM02, MVPB02, NJB^{+}09, PDS07, SRA05, Ski03, SKvdW09, Tsc02]. plastic [KWD^{+}04]. plasticity [PARH07, PAdlPS00]. Plateau [LM07b, ZSZ06, LZ07, CZL05, XZS^{+}07, YLLW02]. play [VIK^{+}08]. plebejus [OG08]. Pleistocene [YJJG09]. plerocercoid [LGL02]. plot [DH01, IO02]. plot-scale [IO02]. plots [LE04, WM06]. plotting [EFHL05]. plume [SBJ^{+}02]. Plutella [MG09, TNO^{+}09a]. PM [DVJ^{+}08, ESZ^{+}00]. PM10 [CFPV08]. PnET [CB07b, NJF^{+}08]. PnET-N-DNDC [NJF^{+}08]. Point [WLG07, APCdlR07, Dam08, DWD07, GS09, HSRD09, KE07a, hMzL08, OWHS09, SK04, ZTXD03]. points [SLE07, TH08]. Poland [KLPP07]. policies [CPV07, CFPV08, CDMM08, JWLA00, MG02, Pra08, Sav00]. policy [CKBH00, JFGN06, LHK00, PMD^{+}09, Sal06]. policy-relevant [Sal06]. pollen [KMRV07, LLF02]. Pollicipes [BBM06]. pollination [CML09, DH06, HC09]. pollinator [KPT^{+}09]. Pollutant [MD07, AAA00, DCI01]. pollutants [MHP^{+}06, MHN00, PGS03, dG04a]. polluted [FDCH08, LW09, NDM00]. pollution [DGOZ04, DUH03, GBEB06, GPSM08, JW09, MBT03, MEJ06, Mon02, MAG01, OWHS09, RNC08, RGO^{+}06, SGH^{+}08, VTB^{+}08, VLD02, ZVLP05, ZM08]. polychaete [WGV^{+}08]. polychlorinated [RBW09]. polyculture [FL07]. polycyclic [WLB^{+}05]. polygon [LC07a]. polygonal [Gau08]. polymictic [BHP08]. polymorpha [vNL^{+}08]. Polymorphism [HLHZ06, GGP03, VS08b]. Polynesia [MFG^{+}06, MFG^{+}08]. polyphemus [BWAM09]. pomonella [TTJ07]. pond [MMN08, PWSY07, fPzWhSY07]. ponderosa [CWL05, LA07]. ponderosae [CRL09]. ponds [BK05b, FLS09, JMVvDV02, KMM^{+}00, SMM^{+}02]. pondweed [JNSS02]. ponticum [SMET06]. pool [PPE^{+}07, Pyk04]. pools [ABS05, BJ02, CM09]. poorest [LLZ^{+}08]. poorly [SPM06]. POP [TMD04]. Population [CCLS06, CA04, FLS06, GGM08, Han09, Hin09, KGA06, KAN^{+}09, LKR06, LH05, LH06, Log02, MKM^{+}07, MKvdW^{+}09, MY02b, OVK^{+}06, SBG06, TK08, TSZdRR03, AALM05, ALM00, BDBS08, BDLL06, BS08a, BBT00, BFD02, BMG08, BEM00, BWAM09, Ben06, BPC07, BPM01, BSTS^{+}02, BM08, BW04, BG03, CRT07, CBMP07, CAB08, CAG03, Cha02, CDAGK06, CDM^{+}00, CFCP04, CSKP08, CS03, Cha00, CMA^{+}06, CBP^{+}06, CM07b, Coh09, CEK08, CTH^{+}00, CTG04, CKBH00, CN07b, CL06, dSDSM08, DAdSC08, DN06, DDH^{+}09, DT03, ECBD09, EFK^{+}03, EDKF06, EWH^{+}02, EVF^{+}07, FPDP06, FC05a, FBC02, GLD07, GLHV08, GG02, GGP06, GGS08, GP07, GGB^{+}06, GP02, GBA08, GTRF01, GW04, HGD05, HLHZ06, HRC^{+}07, Her08, HB05, HAS07, HBDA09, HZ01, Hui06, HMF00, HC05b, IG02, IiTS^{+}04, JKS^{+}05, Jen00, JDD^{+}03, JJWS08]. population [KYH00, KCJ^{+}07, KTB07, KCNN06, KSG05, KGZR05, KBM^{+}03, KWW^{+}09, KH00, LKR03, LR07a, LvNMvdB04, LTM^{+}04, LKH^{+}08, LG03, Lir03, LW09, LUKD06, LPCC05, MB08, MU02, MSM^{+}07, MGH^{+}05, MBLA03, MKRH03, MKM^{+}08, MML00, MML02, MSHP04, MM06c, MBM06, McK01, McK00, MSM^{+}08, MRK^{+}07, MLTN06, MJH02, MRS05, Mon09a, MH07, Nie09b, OWWS01, OI07, PCpDC09, PEM06, PT02, PBMRE08, PSCMMNS06, PTS^{+}04, PBA06, PAdlPS00, PCL^{+}05, PHWH^{+}09, RJGO00, RMD04, RGL^{+}07, RP09, RV05, RC06, RPPB04, RP07, Ros09, SR08, SSS^{+}09, SB06a, SC07, SMB^{+}06, SCPC^{+}07, SRA05, SRL^{+}00, SVS04, SBL03, SSCS06, SK04, SKT00, SMR08, SKJvdHG06, SSPR03, SM06, SO06, SKCM07, SKK^{+}03, SHS08, SGF09, Tar08, TEMJ06, TYZC05, TMD04, TRDM06, TN01, TIJ^{+}01, TNK04, TNO^{+}09b, TBPF08, TSJA02, Uch00, VH07]. population [Van08b, VCP09, VPG05, VMR09, WKL^{+}03, WG07b, Wan07, WCH08, WHH^{+}08, WMH08, WW03, WS02, WG04, WH07, XCW07, Xia07, XG09, YFH03, Yea04, YZS^{+}04, YWL^{+}05, lXzL02b]. population-level [MLTN06, SM06]. populations [AS00c, AKB07, BRW^{+}05, BNTK04, CR03, CDV05, CHHP02, FHE06, GLL^{+}07, GDL06a, GL06, GCLG03, HAA08, HK02, HCL06, Hin09, HP06, JBB^{+}05, JM01, Jen02, JJ00, KNB08a, KVPA07, KVPA08, KPS03, KD05, KS07, LPP^{+}07, LC07b, LPM^{+}07, LDM00, Mat01, MMA02, MSB07, Mul07, PCHG03, POF08, PGLS03, PBK03, PDL^{+}08, SPS03, Sak07, Sak09, SBB06, SJ06, SRS^{+}03, SC02, Sti08, TB08, VPV09, Wal01, WG07a, WL04, Wil07, WSP08, Wit02, XIX^{+}08, ZCB08]. Populus [RSM05]. pore [KFR07, MCGO05, MBGP08]. porosity [LPU^{+}07]. porous [COB^{+}06]. Portugal [GE05, MMM02, PM06a, ROQ^{+}09, BRV09, LCM^{+}09b, RT08, SSPL^{+}08]. Posidonia [EFS^{+}03]. position [Par02]. positional [JBP07]. positions [SAB^{+}06, SBB09, YYZ06]. Positive [Fat07a]. possibilities [MJR06]. possible [Aus07, BF07, MHMHKA04, TMJ^{+}07]. post [ES06, GGS08, HK01, Oli03a, RJFAA07, WM08]. post-breeding [Oli03a]. post-disturbance [WM08]. post-encounter [RJFAA07]. post-normal [HK01]. post-reintroduction [GGS08]. post-wildfire [ES06]. Potamogeton [BKS05a, CMM02, HCJ^{+}09, WvS06]. potato [HTA^{+}08, Ken02, MH07, SBJ^{+}02]. Potential [Bra01a, BBKN03, FRZ00, LSM^{+}04, SHW04, TFF07, Ale07, AFLB09, BM08, CB07a, CTH^{+}00, CGH^{+}03, FCC^{+}04, GBN^{+}06, GSC09, HHC02, IOIA08, Kir06b, KBF^{+}08, LN08, MDB^{+}02, RPVR03, Rod07, Roe01, SF07, dCSB08, SL04, Sno08, XIX^{+}08]. Potentials [DKL^{+}09, BDJ^{+}09, RHB04]. Potter [MKS^{+}02]. Pounds [Jor05a, Jør04c, Jør04d, Jør04a]. Power [RM02, WZN05, AHKB01, COC04, CMD06, Hal04, Hen07, Kin04, MMV^{+}09, MTVC05, SI06, SWSF06, TSF07]. Power-law [RM02, SWSF06, TSF07]. pp [Bro09, Jør04a, Jor05a, Jor05b, Jor05c, Jor05e, Jor05f, Jor05g, Jor05h, Jor05i, Jor05j, Jor05k, Jor05l, Jor06b, Jør09b, Log02, Zha06]. pp. [Jor05d].

Practical [ABV^{+}06, Jør04e, CKL^{+}06, HB01, WLL^{+}08]. Practice [Jor05a, Jør09b]. practices [AKB07, BDR01, GH09, SCH05]. practitioners [VG08]. Prague [MEJ06]. prairie [HBUS02, MC01, MG02]. pratense [WH07]. pratensis [MMHE06]. prawn [OG08]. Pre [APHV08, RJFAA07, AN06, CGdPRY09, SMSR00, ZDR03]. pre-alpine [SMSR00]. pre-clearing [AN06]. pre-defined [CGdPRY09]. Pre-encounter [RJFAA07]. Pre-European [APHV08]. pre-ingestive [ZDR03]. precautionary [HK01, MHSP^{+}06]. precipitation [GTJ^{+}00, HNF09, WHP01]. Predation [LYC08, Bor06, CMR09, FF06, FMRS09, FSJ03, JM01, Jen01, LWJ06, PFBBJ08, RLDL09, SSH^{+}07, SHW04, Yos06b, ZPP06]. Predator [DS01, dSPdBB08b, AVTP05, AGZ05, BPM01, BG03, CSG02, CEP06, CKBH00, CN09, DH06, DDH01, FG06, FGEG06, GG05, HM01, JCE06, KT03, KC08, KVPA08, zLGH^{+}05, hMzL08, MAL06, MB09, NG07, NKC04, PGA06, RJGO00, RJFAA07, SG05, SO02, SJLL08, Tsc04, VATP05, WH09]. predator-dependent [DH06]. predator-free [hMzL08]. predators [BFE07, GORJ03, KDS03, Oku09, PEPB09, RJDF08, VH09]. predatory [DMBO00, HB03b, KH00, Ski04]. predict [AMSW07, AMM04, BB03a, BKGC05, BBP08, BHSR01, BH03, CPH00, CWS09, DJ05, GZ05, HB01, HB02b, HB03a, HGB04, HLH^{+}06, HvG07, HBDA09, JJWF07, LL00, LGL02, MH03, MSHL00, MVPB02, MOLN06, MJR06, MBS^{+}09, Mon09b, MLM06, RBSG06, SPD^{+}07, SBDD04a, SBJ^{+}02, Sti08, SVB09, TT05, TSJ^{+}09, VM09, VSFM03]. Predictability [MGS^{+}09]. predictable [SPM06]. predicted [BPE^{+}07, EFHL05, FM08, MCSG06, MC01, PPR03, PS09, PPGP08, VBD06]. Predicting [BHP05, BDI04, BRW^{+}05, BCS09, CM04, DSA08, DJ05, FKIL08, GBB^{+}09, GBN^{+}06, GR09, HBRW07, JBB^{+}05, JKPL09, KK00a, Kno03, MBW09a, MFB^{+}06, MAK^{+}04, ÖTÖR06, PP04, PMLM08, Pot04, RDS07, RJR04, RSB09, TPC^{+}07, WRCB01, WBR08, WYMS07, WW03, WR01b, WHZ06, XWB04, ZLO02, ZKH09, vW07, AW06b, ABM^{+}06, BDF^{+}06, BHAV03, CRT07, CR03, CWF03, EDKF06, FDCH08, FH08, GPK00, GKG05, KP00, LHB08, Mac00, MSA^{+}03, MF02b, MEO06, PCCL03, PLD^{+}02, RT08, RPVR03, SGHG04, TVS00, Wil01, ZCG^{+}08]. Prediction [AMSM06, AG01, FCKH06, FH08, IC02, JJK^{+}01, Leg03b, MLL^{+}05, SOK06, ŽŽV^{+}06, AGD06, AMD^{+}03, Aus02, BCCR^{+}02, BNM09, BR01, BSR04, ÇDKK05, Cor05, DGGD04, DGD03, DLC07, FBDM09, GB01, HK01, HSRD09, HR03, IOIA08, JKJ^{+}08, KPKP06, KŞE04, KHLS07, LHC07, LOL02, LOL03a, LOL03b, MLH05, MHM^{+}03, MNEB01, ML05, OLK^{+}04, OW05, PKC^{+}01, SRW05, SH09a, SFC05, SL01, SCAP05, SWBH08, WGS^{+}02, WL06, WYO05, Jor05d]. predictions [Ash06, BHP05, CKP^{+}01, CWBR01, DLR03, DAR^{+}07, EBR02, KM04, LM07b, MB02b, NMJ07, PHBF07, RR06, Rem04, VBRS07, WC07, Wel04, WRC^{+}08, WR01c, vHPS02]. Predictive [Ano06-46, GZ00, HRMC01, KCJ^{+}07, PCC^{+}07, ALP03, CL08a, DGD06, HMBG03, HN09, KC05, Lin01, MFA07, MBF09, OW08, PF00a, RPVR03, SBDD04b, WGV^{+}08]. predicts [VH07]. predisposition [RPRV09]. predominant [OAL^{+}07]. Preface [Ari00, Bre05, GH09, HCLR05, JT09, KMIW07, Li00b, Rec01b, Sca06, Sek08, ZSTM06]. preference [DD03, FH08, Fuk09, Mat03b]. preferences [SO02, WGV^{+}08]. Preferential [VL08, WTMG09]. Prelim [Ano02-52, Ano03-64, Ano03-60, Ano03-61, Ano03-62, Ano03-63, Ano03-65, Ano04-76, Ano04-77, Ano04-70, Ano04-71, Ano04-72, Ano04-73, Ano04-74, Ano04-75, Ano05-46, Ano05-47, Ano05-48, Ano05-49, Ano05-50, Ano05-51, Ano05-52, Ano05-53, Ano05-54, Ano06-39, Ano06-40, Ano06-44, Ano06-41, Ano06-42, Ano06-43, Ano03-59]. Preliminary [CDA08, CS03, LWL^{+}02]. preprocessing [KST^{+}07]. prescribed [PBV05, RCZ^{+}06, Ski03]. presence [CWS09, KC03, MFB^{+}06, OW06, OLB04, PCpDC09, PDL^{+}08, VSGW09, ZLO02]. presence-only [OW06, OLB04, VSGW09, ZLO02]. presences [HLH^{+}06]. Present [Med06, APHV08, HO01, NHP^{+}06, TMS^{+}07]. preservation [ABR05, MAHvD08, Tis06]. Press [Jør04g, Jor05e, Jor05h, Jor05l, Jør06a, Nie06, Nie08a, Zha06]. pressure [AFTB07, BMT07, FRZ00]. pressures [Gra07]. prevailing [DBD^{+}08, KGSB01]. Prevalence [MDVG09, FM08, PCL^{+}05, WLG07]. Prevalence-adjusted [MDVG09]. prevent [CSG02]. Prey [AVTP05, AGZ05, BvdW04, BFE07, BG03, CSG02, CEP06, CKBH00, CN09, DMBO00, DS01, DDH01, FG06, FGEG06, FAB^{+}07, GORJ03, GG05, HM01, KT03, KC08, KVPA08, KDS03, zLGH^{+}05, hMzL08, MPRJ04, MAL06, MF06, MB09, NG07, NKC04, Oku09, dSPdBB08b, PGA06, RJGO00, RJFAA07, RJDF08, RLDL09, SG05, SJLL08, Tsc04, VATP05, WH09]. PRI [Bra01a]. Price [Jør09b, Nie06, Zha06]. primary [AWL04, BGF00, CMDP^{+}00, CWBR01, DVJ^{+}08, GYY00, GYW^{+}04, KST^{+}07, MDH^{+}09, MXC^{+}04, NMC^{+}06, PK01, PBPZ04, PLB^{+}06, RGF00, Rob05, RCGB08, Sca01, SOA03, SMM^{+}02, SWK^{+}05, SSNHP08, WBH01, ZBD^{+}09, ZPD06]. primer [HDBM02a, HDBM02b]. primordial [ZTF08]. Princeton [Nie08a]. Principal [PLL04, BL02, BGL01, ÇDKK05, DLPN06, Las06b, PBCZ01, RBSG06]. Principle [BDE08, Den08, Ned09, Aok08, BPP09, COC04, CMD06, HK01, HA08b, JVL02, Kir06b, Leg08, Jør06a]. principles [Com04, Gau06, Hul02, Hul04, Kir06a, Lev00, NM09, Pet04b]. Prioksko [KK08]. priori [RDS07]. priorities [HMF00]. prioritize [KRZ07]. Prisoner [iTI02]. pro [BPJM00]. Probabilistic [DPL^{+}04, FHT^{+}09, BBD^{+}04, HFV03, HBDA09, LBBN06, RGL^{+}07, RA06, TVS00, WSP08, YTT^{+}09]. probabilistically [Tol06]. probabilities [SMET06]. probability [BS04, BRP^{+}06, Coh09, JBT^{+}05, KYL07, SLPP05, SHZ05, TCGL03, WHZ06]. probability/Markov [SHZ05]. PROBE [EPTB07]. problem [BM00, CPV07, DM03, KE07a, LK02, Log08, MVZM05, WLJ00]. Problematic [GP02]. Problems [dlPP05, ADdSC06, Bul08, HB04, Haw00, Kir01, MKDV08, Sep00, SV03b, Yun08, ZVLP05]. procedure [GICB09, HTA^{+}08, LIS07, VPSG05]. procedures [AB03]. Proceedings [Ano06-45, LMPR06]. Process [Bru05, HHK07, PBCZ01, AOY02, AQS^{+}07, AP06, ALBA06, CV07, CWL05, CWS09, CMPO05a, CMPO05b, DVdB^{+}08, DMRP07, GRBT08, Håk09, Ito07, JLBS09, KFR06a, KBM^{+}03, KWW^{+}09, KH02, LD06a, LBL^{+}08, LSY^{+}09, LTP06, LB01, Loe00, MGV00, MSHL00, MS07, MXC^{+}04, MLF^{+}06, MBS^{+}09, PGW00, PJAZ02, PCS01, Plu00, RR06, RVWH06, RAM^{+}03, Sal06, SS06a, SE04, SLGS00, Sta07, Sti06, SSM^{+}09, TYS^{+}09, TF09, WR01b, WDR06, XS08, XBM^{+}08b, XJM07]. Process-based [HHK07, AOY02, AQS^{+}07, CWS09, CMPO05a, CMPO05b, DVdB^{+}08, DMRP07, GRBT08, Håk09, KBM^{+}03, KWW^{+}09, KH02, LBL^{+}08, LB01, MSHL00, MS07, MXC^{+}04, MBS^{+}09, PCS01, RR06, RVWH06, SE04, SLGS00, SSM^{+}09, TYS^{+}09, WR01b, XBM^{+}08b, XJM07]. process-oriented [CV07]. Processes [RST05, ALO^{+}01, BB08b, BLB08, BLB09, CJS^{+}07, CMM^{+}07, CM08, DDG^{+}05, DLP^{+}09a, DPT09, ESG06, FWS^{+}05, FAA^{+}02, GS09, HE09, KPT^{+}09, KTB07, KMN^{+}07, LL02a, MWK07, MLGG09, Nie08c, Reu05, San07, SS06b, SRN05, SF04, TKK07, TTB06, VJ01, WPBM06, Wil09, ZH06, YTCV01]. processing [CR06, KKL^{+}06, PE08, ZAM^{+}03a]. producer [PK01]. product [YSM^{+}06]. production [AZM^{+}06, BGF00, BM00, BH03, BBKN03, CdQO06, CRC08, CMDP^{+}00, CPB^{+}08, CCGMJ07, CWBR01, CR06, EFS^{+}03, GL09, GYY00, GYW^{+}04, HB02b, HB03a, HB03b, IO02, JDPI07, Jen05, KST^{+}07, LD06a, LP03a, Lud04, MD06a, MB06a, Mul07, MB07, MG02, New06, PBPZ04, PEE09, PCS01, PLB^{+}06, PS01, RGF00, Rob05, RHB04, RCZ^{+}06, Sca01, SP01, SLS03, Sch03, SOA03, SOK03, SHP^{+}07, SSNHP08, TYZ^{+}05, TDLP03, VHG05, WBH01, WSF^{+}00, WSF^{+}02, ZBD^{+}09, ZCG^{+}08, ZBL03, dBWG04, dBWG05]. PROductive [LCF09]. Productivity [KB04, AWL04, AY07, Bar04, CWL05, CWS09, ECL^{+}02, Eza05, HvRIZ08, MDH^{+}09, MGH^{+}05, MXC^{+}04, MLF^{+}06, MB02b, MM00, RNKG03, Roe00, SWK^{+}05, SSNB07, TS03, VVF06, Wal01, WKZ03, XBM^{+}07, XBM^{+}08a, XBM^{+}08b, lXzL02c, ZVK05, ZPD06, ZAM^{+}07, vN02]. products [AM02, ABC04, LH09, QMH00, TFM01, vdPLG^{+}00]. profile [CM05a, Nad07]. profiles [GB02, HFV03, TYS^{+}09, WLJ00]. prognoses [ZJBI03]. prognosis [GG00]. prognostic [KWS^{+}07]. program [BK07, TW02]. Programming [CT07a, DRDD01, MHvIR00, McK01, ML05, Sti06, Whi00]. PROGRASS [LCF09]. progress [Haw00, JLF08, PvdBW^{+}02, WDW00]. project [KRZ07, SS08b, WDW00, ZRR^{+}08]. Projecting [SGF09, Wal04, MKM^{+}08, TDdSLS^{+}08]. projection [Log08, MJH02, RSA04, SR08]. projections [LE04]. projects [MGCK^{+}03, Pra07]. Promise [HB04]. promotes [Zha03b]. Proof [vWSH08]. propagation [GdBD08, IGP^{+}03, LBBC08, MWP07, PVS^{+}09, SPTP01, SKT00]. Proper [KNB08b]. properties [BFU^{+}09, BPA08, BCRSTVG07, BBGH07, BMR^{+}05, BMR06, BHV06, ESWG02, Esc05, GRPF07, HB05, KP09, LCLC07, MRRJ07, PM09, RHM^{+}05, Reu05, Roe01, SI07, SM09, Svi02, TH06a, TAL^{+}05, VPG05, WHW^{+}05, YP02, ZPK^{+}07, vNM02]. property [THA02]. prophecies [Cam04]. Prophet [Cam04]. proposal [GSBN03, Ort04, SBVB05]. proposed [YCVA01]. Proposing [NM09, PBG09]. propositional [Paw00]. prospecting [RST05]. Prospects [HHF05]. prosperous [Ano04-97, Ort04, Ulg04]. PROTBAS [MP07]. PROTECH [ET04, EPTB07, LELR02, RIE01]. protected [CGS08, CFS09, DDFP07, OACB09, PFR09, PCP07]. Protection [CBSLS07, SBVB05]. protocol [GBB^{+}06, TSBH09]. prototype [KRZ07]. Provence [RNKG03]. provide [PR01]. provides [MCK07]. Province [CL08b, mLMfT05, YZC^{+}07, MGH^{+}05, ZCG^{+}08]. pruned [Cor05]. pruning [PBSOMG^{+}05]. Pseudo [dlFN08, CL08a, VSGW09]. pseudo-absence [VSGW09]. pseudo-absences [CL08a]. Pseudocalanus [SKCM07]. Pseudokirchneriella [FCP^{+}07]. psyllid [AFLB09]. pubescens [WRB08]. Publisher [Ano03-66, Ano04-78, Ano04-79, Ano04-80, Ano05-55]. Publishers [Ano01-43, Ano05-56, Jør04c, Jør04d, Jør04f]. Publishing [Jor05c, Jør09b]. Puerto [HHL08, WCHM02, WHZ06]. Pula [Leg01]. pulse [WYO05]. pulse-disturbance [WYO05]. Pulsed [YH04]. pump [CMSC01]. pure [LMH05]. purification [TN09]. Purple [Wel04]. purpuratus [WTMY07]. pursuit [zLGH^{+}05]. pus [WM06]. Pushchino [KL08]. Puy [GPDF09]. puzzle [YJJG09]. PVA [HMG06]. pyramids [FK07, WMSW09]. pyrenaicus [BRV09]. pyrene [DDH^{+}09]. Pyrenopeziza [PvdBW^{+}02].

q [dMPVNO07]. q-bits [dMPVNO07]. Q2 [VKB06, ŽŽV^{+}06]. Qalabotjha [ESZ^{+}00]. Qilian [CZG05, ZNC^{+}06]. Qinghai [LM07b, ZNC^{+}06]. Qinghai-Tibet [LM07b]. qua [MC04]. quadrants [JAN^{+}03]. quadrat [Nal01]. quadrat-sampling [Nal01]. quadratic [Ric05]. quadricauda [KM00]. quaking [SRB06]. Qual2e [PdC06]. QUAL2Kw [KLL^{+}07]. qualification [SLS03]. qualify [PDBJ09]. Qualitative [DLR03, SBB06, SBA06, SSPR03, BMS^{+}08, FBC02, LCH^{+}00, MMLR07, RLLB09, SB06a, TN06, Van04]. quality [AMB07, BD05, BW01b, BDI04, BU04, CLM^{+}09, Cal05, CJS^{+}07, CJ07, CBS09, CWF03, Cor05, DWH06, GPV08, GACO04, HPH00, HKPH08, HBRW07, HHKH09, KLL^{+}07, KKCC06, KFB09, KFH07, Leg03b, LXP^{+}08, Lin06, LFB07, LT01, Man00, MPRJ04, MS00, MMJN03, MB02b, MG02, NP06, PLTT05, PLL04, PL02, PNU03, PCWP06, PCS03, PA09b, PBCZ01, PTGI09, PV04, PKG00, RSW07, Roe00, RCL06, Sal06, Sav00, SBVB05, SBMJ09, SHK^{+}07, SJM03, TFM01, VJ06, WG00, WYMS07, YSM00, YZC^{+}07, ZRR^{+}08, ZCZ04, ZLW09, vLDHP08]. quality-macrophyte [MS00]. Quantification [ATAK00, Sch03, TPCS06, ZSD^{+}08]. quantify [JcLM09, LPU^{+}07, LHK00]. Quantifying [DFCM01, JBP07, KW05, LS02a, MLGG09, BU04, HDBM02a, HDBM02b, HTK07, OJD04]. Quantitative [APG^{+}03, GPP02, SY07, ALBA06, PGFM04, RLLB09, SCH05]. quantum [DPT09, XZS^{+}07]. quarantine [VPB08]. quarries [BCMR03]. quasi [BPP09]. quasi-sustainability [BPP09]. Quebec [Ano06-45, LMPR06, FBDM09, GCG^{+}07, PLP^{+}04]. Queensland [HRMC01, Mey04, YZS^{+}04]. Quercus [ASI^{+}08, LvGC^{+}04, dJTMBGMP^{+}09, VSFM03, WRB08]. question [Oku09]. QuickBird [LL07b]. Quinte [MMF^{+}09b]. Quo [BP05]. quota [Dew01, LPM^{+}09].

R [Bro09, Jør04e, Jor05i, FE04, Cal06]. R. [Jor05c]. Rab [Leg03b]. rabbit [FSBD01, SBG06]. rabbits [FSBD01]. rabies [BSTS^{+}02, Sel00]. raccoon [CGHW05]. radial [LPD08, MAB01]. radiation [AMK00, BCD^{+}05, CBSLS07, DBD^{+}08, Dun08, GGH08, MSL06, ORS^{+}09, PG04, SS01, SSP03, WCC02]. radicals [KK07]. Radio [WPMT07, CDM05]. Radio-tracking [WPMT07, CDM05]. radiocarbon [BCG08]. radiometer [KWS^{+}07]. radionuclide [RIGJM06]. radionuclides [BA05, DBBS03, vdPLG^{+}00]. radiotracking [DAR^{+}07]. radish [FK05]. Railsback [JT01]. rain [BBGM05, BBGM06, GAB^{+}09, HD00b, OIP^{+}08, OIR^{+}08]. Rainfall [HTMO06, AJJ^{+}05, DVPS08, HHC02, LC07b, WGS^{+}02, ZPW05, Jor05k]. Rainfall-Runoff [Jor05k]. Rainfed [Ayd08, TFTO07]. Rainfed-areas [Ayd08]. rainforest [AS00a, AB06, OHH01]. rainforests [PMC08, TH06b]. ramorum [HXP^{+}09]. Ramsar [MTD^{+}09]. Rana [BM08]. Ranchi [MMN08]. Random [PDV^{+}07, Ben04, dSDSM08, TN01, Vil06, iWLSN00, YMT03, ZJC^{+}07]. randomization [OJ02]. Range [LPP^{+}07, BPA08, DAC^{+}09, GBS00, HB09b, HGR08, ICGÁ05, KM06, KPH02, LAZ^{+}08, MCSG06, MP04b, MHN00, PKS08, PMD^{+}09, SJ05, SBC07, WG07b, WPMT07, Wel04]. rangeland [ES06, JWLA00]. rangelands [HRH^{+}05, MDGK01, PCS01, PSBJ07, RKH^{+}07, THA02]. ranges [MP08, NBP05, QMH00, TKTB07]. Rangifer [FCH04]. rank [BMD09]. Ranking [BT01]. Rao [Ric05]. rape [BHM^{+}06, CBP07, CMCD04, CML09, DSD^{+}08, GL06, PvdBW^{+}02, PDL^{+}08]. rapid [Per07, SOK06]. rare [RRB^{+}01]. Rarotonga [DD03]. raster [DFM07]. raster-based [DFM07]. rata [BPJM00]. rate [ÅSCP09, COB^{+}06, FBF^{+}06, GPC^{+}09, HLK06, KYL07, LKR06, LBS08a, MKM^{+}08, NDM00, Par00, PMC08, Sil07, SO06, SHM07, SWBH08, iT03, TR03, TTPS09, TBPF08, WHH^{+}08, WLB^{+}05, WTMY07, Xia06, YWHW02, dlS07, vLDHP08]. rates [BDBS08, BMBOCR03, CBMP07, CML09, CGR03, JJWF07, KRvL^{+}02, KM00, LKH^{+}08, LHB08, MLH05, MSHL00, MMA02, MMPN07, PBHGF07, PMC08, SJ06, SMR08, Ski03, TN01, WC07, Xia00b, YTCV01, ZKH09]. ratio [DH06, KNB08b, LL07c, PGLS03]. ratio-dependent [DH06]. ration [RLFB04, Tud01, WPD04]. Rationale [Mon09b]. ratios [Met03, TVS00, TFM01, YYZ06]. rats [WLG07]. Ratto [Jor05a]. re [FL07, LPFL09]. re-emergence [LPFL09]. re-supply [FL07]. reach [HHK07]. reaches [LWLZ06]. Reactive [DWR09, Lin01]. Reactive-transport [DWR09]. Reactivity [NKC04]. reactor [LJW03]. reading [Vil06]. real [AER^{+}07, AN02, ATDK08, CPB^{+}08, ML05]. real-time [CPB^{+}08, ML05]. real-world [ATDK08]. realised [HPHP04]. Realising [MC04]. realism [WC07]. realistic [RF09]. realities [BCS09]. really [KW02a, LL00, SS07]. reappraisal [SWB06]. rearing [HP09, PSC^{+}01, SCR03]. reasoning [SB06a]. receiver [PPS08]. recently [VOM06]. recharge [BKS^{+}05b]. reciprocal [Hul01]. recirculation [LKKL09]. reclaimed [Ray08]. reclassification [KDK06]. reclycing [Sar04]. recognise [MM06b]. recognition [GML09, ML04, WO01a]. Recolonization [SPD^{+}07]. reconciliation [UJF06]. reconstruct [GTJ^{+}00]. Reconstructing [APHV08, KTB07, LRJ^{+}09]. Reconstruction [Li00c, MCE^{+}07, Par00]. record [KLM^{+}02, LLF02]. records [HSRD09]. recover [ASZRRR08]. recoveries [Xia00b]. Recovery [LMGM^{+}09, PBH^{+}07, CB07a, LAXP07, LTP06, LLCL04, NKK^{+}07, PMC08, PBM^{+}05, RMS08, SBA06, Sno08, VB06, WYO05]. recreational [AF09]. recruit [Xia02]. Recruitment [DLC07, RC06, SKvdW09, AMSW07, CH06, JPB03, KH07, NCM07, OMHR06, RHE06, SPD^{+}07]. recruits [Rod07]. Rectangular [BOB07]. recurrence [ZSD^{+}08]. Recurrent [JKJ^{+}08, CKP^{+}01, JJK^{+}01, WRCB01]. recursive [MD06b]. recycle [YLSH03]. recycled [AM02]. recycling [AN00, ABC03, BBM04, KPAK02, KWD^{+}04, dSLSD02]. red [ALAS09, AS00b, DDJ^{+}01, Kno03, LZZ^{+}07, MWM02, MML00, MML02, ÖTÖR06, RMS08, SHW04]. red-winged [ÖTÖR06]. redistribution [Han02, HLR06, JJWS08]. Redthenbacher [SBL03]. reduce [LH09]. Reduced [vLLv^{+}07, Han09, NDD^{+}07]. reduced-gravity [NDD^{+}07]. Reducing [GS06, LH07]. Reduction [CGdPRY09, AS09, CGY08, Eti04, MPRJ04]. reductions [CG00, LXP^{+}08, SHCS04]. redundant [LSHG08]. reed [AMB07]. reef [AGNLGSG04, Bar04, CHL08, DDFP07, Fig09, KD05, LS04a, LPM^{+}07, MFG^{+}06, MFG^{+}08, Oke04b, RM09, SAL07, Cha02, Gri04]. reefs [Dun08, GB08, MVMA08, MHZ^{+}06, YMD08]. Reeve [BSB^{+}09]. reference [CBD^{+}09, Håk09, KW02b, McK00, TPC^{+}07]. referenced [LWC^{+}07]. references [Nie06]. refinements [BCM^{+}08]. refinery [AW06b]. Reflections [Zuc04]. reforestation [LTLH08]. refuge [GP06, MLPK01]. refuges [GORJ03, GAPE06, WMSW09]. regarding [Ned09]. regeneration [WBB05]. Regime [Sri04, AER^{+}07, AHKB01, CMSC01, EPS04, Ito05, MLH05, MPC06, NCM07, SFM04, TL03, ZSD^{+}08, ZBD^{+}09]. regimes [DFF07, ECL^{+}02, Fat07b, FSM^{+}01, HS03, Li00c, PJAZ02, Pyk04, SAK00, SSM^{+}09, YHG04, Zav08]. Region [LHC09, AKMB01, GPSM08, GTJ^{+}00, JJ00, KMN^{+}07, Las06a, LWLZ06, LZZ^{+}07, LXP^{+}08, MCM^{+}09, OLK^{+}04, PGA06, PTGI09, RPRV09, RE03, SM07a, TYT^{+}09, ZH06, ZXC03, dGdMPC05, vdPLG^{+}00]. Regional [LOM06, TMS^{+}07, ACE07, BRGS09, CKL^{+}06, DDS^{+}04, GSBM08, GARB09, GYY00, GYW^{+}04, GdBD08, HY07b, MXC^{+}04, MOLN06, MB02b, OIP^{+}08, Saa00, SW03, SHS08, TCD02, WHW^{+}05, XHC^{+}08]. regionalisation [HKB02]. Regionalised [Sep00]. regions [ASN02, JM04, MLGC03, OEK^{+}06, WHHH07]. Regression [BPC04, ASG^{+}05, Asp02, ÇDKK05, CM04, CG06b, DGRU06, Dor07, DD03, HSRD09, KDW^{+}09a, LEH06, LOL02, LOL03a, LOL03b, LAZ^{+}08, MBS^{+}09, NDM00, Ond07, PF00b, PF00a, SRN05, SMET06, STH06, WA02, YSM^{+}06, YSB08, ZH06, vHPS02]. regression-kriging [HSRD09]. regressions [Hen07]. regressive [APG^{+}03]. regularities [Gut07]. Regularization [Rd06]. regularizing [Par00]. regulate [HE09]. regulated [AVTP05, GACO04, HKPH08, KCJ^{+}07, RLBL01, VATP05]. regulating [MWH00]. Regulation [Ano03-27, JBMBP02, JBMBP03, BS08a, GAB^{+}09, HZHL05, KCBS00, MJ06, OSHO09, RL05b, SJH05, WG07b]. reindeer [MDGK01]. reintroduced [GG04a]. reintroduction [FJ05, GGS08]. reintroductions [VB06]. related [CJS^{+}07, Håk00, Mar06, MOP^{+}06, VM06, WLLY04]. Relating [RJGO00, SM06, PLP^{+}04, TFM01, YLL^{+}05]. relation [CPT09, LL07a, LGD01, NiTT01, OMA01, PvdBW^{+}02, SRL^{+}00]. Relations [DSD^{+}08, KGZR05, Fat07a, SF09]. Relationship [KRvL^{+}02, OAAF07, Pet04a, SDA^{+}03, Xia02, AF05, BB03b, GYY00, HZ01, JPB03, LGC07, Mat03a, NCM07, PEAS01, SI07, SB07, SZL^{+}04, TTPS09, lXzL02c]. Relationships [GYW^{+}04, BP03, DS01, Hee02, HvI01, HGR08, ICSC05, Kri04a, LC01, MP08, Odu02, PLA06, SWBH08, Vil05, VOM06, Whi00]. Relative [RLR09, SO06, Bra01a, BHP08, Gra05, KZH07, MDH^{+}09, MWH00, MVV07]. relatives [MML00, TTA^{+}03]. Release [SJ05, NWP06, RLDL09, SU08, TFM01]. relevance [FK07, RHH05]. Relevant [BB03a, KKA^{+}08, Sal06]. reliability [Jor00a, JSM03]. reliability-theory [Jor00a]. remediation [PBD00]. Remote [BKGC05, CCJ07, CGG^{+}05, TYZ^{+}05, BCMR03, CMB^{+}02, GB08, HDB^{+}06, LWLZ06, LBBR01, MVMA08, Met01, Met02, PWZ^{+}09, RSC09, SAL07, SOA03, Szi00]. remotely [Asp02, DSA08, Dun08, FC06, LLLT08, MMV08, MLL^{+}05, Plu00, SAS06, TPJ08]. Removal [CSG02, AMSM06, Del04, KKH^{+}08, PWSY07, fPzWhSY07, PBM^{+}05, PGG06, SMM^{+}02]. removals [BL01]. removed [ZJN^{+}06]. renewable [BPP09]. Rényi [Mon05]. repens [GPB01]. Reply [BFS05, KL09, MMV08, Mai08, RLHD01, Hui07]. representation [Bia03, Bla07b, LS09b, NBT^{+}09, SHZ05]. representations [Mal03]. Representing [Rob05]. repression [AVTP05, VATP05]. reproducing [LMH05]. Reproduction [MSB08, DSD^{+}09, LL06, SM09, ZPGG03, vLLv^{+}07]. reproductive [BGWC07, LH05, LH06, SMB^{+}06, TPJ08]. Republic [NKK^{+}07]. required [GPK00, LL06]. requirements [HLS06, MHKW00]. rescue [KDS03, LL07c]. Research [LBBC08, AMRS08, BMR^{+}05, CKL^{+}06, HMF00, NS07, WHP03, XCW07]. reseeding [ES06]. reservation [ESG06]. Reserve [OACB09, BBM06, HSMN08, LCH^{+}00, MRE^{+}06, RMF09, SSR07, SHW04, SG05, Tol06, WRP07, DCG01, KK08, SF07]. reserves [LL08]. Reservoir [HAG^{+}00, OAL^{+}07, RdQFO07, WRCB01, AP00, BP02, BP04, CLM^{+}09, GPDF09, GE05, KP09, KFS06, KHLS07, LELR02, LCLC07, Ond07, PK08, PCL^{+}05, PGG06, RFdV07, SJM03, VOM06, dlFN08, LSY^{+}09]. reservoirs [AP02, ÇDKK05, FLS09, KST^{+}07, OB04, RAI04, RMOA06]. residence [CU06, RMOA06]. resident [MY02b, SSPR03]. residential [OLK^{+}04]. residual [BVD05, WFT05, ZGH05]. residues [MMJN03]. Resilience [FH07b, VCM07, GBA08, KGBC03, MHZ^{+}06, OW02, lXzL02c]. resistance [GP06, LT04, RS04]. resistant [BPBF^{+}00]. resolution [Asp02, BDF^{+}06, FSJ04b, HDB^{+}06, KWS^{+}07, LJR06, LE04, MB02b, SDS^{+}07, SV03b]. resolutions [MJR06]. resolved [BS07]. Resolving [LAZ^{+}08]. Resource [Nie09b, Tsc02, Uch00, AS01, ARF08, BB08a, BVNS02, BPJM00, CGY08, CCY06, DUH03, Esc05, GBSP08, JG08, Kai00, KRvL^{+}02, LM07a, MG02, OCK01, OW05, Pet04b, RMF09, RHHM08, SPS03, Sak09, SF09, SC07, Van08b, WDBK08, WG04, YM05]. resource-based [Sak09]. Resources [Jør04h, BFD02, BPP09, Bla07a, FL07, LL07c, MCJ^{+}04, MP04b, PHH00, RD07, TOS09, WAB^{+}07, Wil09]. respect [ALAS09, MHSP^{+}06, SFU01, TBPF08]. respiration [CPB^{+}08, GQB05, HHL08, HHB^{+}08, JZY07, MLH05, QXW02, TH06a, TVK^{+}08, VHG05]. Response [Cha07, CG00, GRBT08, LCG^{+}09, AMSW07, ASZRMH^{+}04, BPBF^{+}00, BCL^{+}09, BDLL06, BSM08, BMT07, BJFM06, CCLS06, CG06b, CP01, DMBO00, FF06, GR09, Gra07, HFV03, KPKP06, KKCC06, LCF09, LK07, LAXP07, hMzL08, MOP^{+}06, MSM^{+}08, MDH^{+}08, NI08, OT03, OL09, PKC^{+}01, PEM06, Per07, RJGO00, RP09, SH09a, Sin07, SB02, Sta07, TKSP09, VH07, WBR08, WO01b, WSF^{+}02, lXzL02b, YK00, YM05, ZvBS05, ZHW05, ZAM^{+}05, ZAM^{+}07]. Responses [JAB^{+}06, Ano04v, BMB07, CSU03, CSU04, DH06, EFHL05, HM01, HCL00, KJ08, LKR03, LXP^{+}08, MRiI^{+}07, OM02, Oku09, RNKG03, RIE01, RWW07, RGL^{+}07, RWM^{+}07, SBG06, SKJvdHG06, SM06, TB06a, WZJ08]. responsive [KNM09]. resting [HB06]. restocking [LGS^{+}00]. Restoration [LGS^{+}00, Bey03, DVGD07, GDG^{+}00, GZJ06, Lin03, MD04, RMD04, SHB04, WDW00]. restoring [Xia05a]. restricted [PTGI09]. restrictions [ZR04]. result [KPT^{+}09]. resulting [FHE06, SDL08, VCP09]. Results [Ano06-46, NWH^{+}06, vW07, BA08, KIL^{+}03, LUKD06, Mon09b, NKK^{+}07, ORF01, SKJvdHG06, YJJG09, ZPO09]. resuspension [CBS09, CGH^{+}05]. retention [GBG^{+}03, LXJ^{+}03, WR08]. Rethinking [PPS08]. reticulatum [SB03]. retrieval [JNM^{+}06]. return [BP08]. reuse [YLSH03]. reveal [BKMB08, SM07b]. revealed [LMGM^{+}09, Zha07]. Revealing [Gre06]. reveals [Hen07]. Review [Bro04a, Bro09, GDL03, Jor05b, Jor05c, Nie08a, Zha06, BK05a, BCPM09, BL04, Kri04a, Lim04, Odu04, PB02, SH04, War08, YCVA01]. reviewers [Ano02y]. Reviews [Jor01a, Jør02a]. revise [TDLP03]. revised [WGS^{+}02]. Revision [ZDR03, ALBA06]. revisited [LC01, OM02, RADD^{+}01]. Revisiting [Del04, MHZ^{+}06]. Reynolds [SPK^{+}09]. Rhine [VZG05]. rhino [CHHP02]. rhizomatous [WvS06]. rhizome [SWCO07]. rhizosphere [SKM^{+}06]. Rhododendron [SMET06]. rhomboïdes [SM07a]. rhythmic [MB00]. rhythms [MB05b, WPD04]. Ria [GE05, LS06, ROQ^{+}09]. Rica [DCG01]. rice [CMB05, CABD09, DS01, MZWM05, SCP05, TFTO07, WSF^{+}00, WSF^{+}02, XJM07, dG02a, dG02b]. rice-fish [dG02b]. RICEPEST [WSF^{+}02]. rich [BP02, EGE^{+}08, ZBWR06]. Richards [LdVA06]. richness [FC06, KH07, MFG^{+}06, PCCL03, PLH09, PCC^{+}07, SLZ09]. Rico [HHL08, WCHM02, WHZ06]. Riederalp [Ano06-46]. Riga [Håk09]. rights [MBLA03]. rigorous [MVZM05]. Rinconada [OACB09]. ring [BB04, Mii00]. ringed [FKIL08]. Rio [HTS^{+}07]. riparian [HTS^{+}07, HKHB06, SL01, WM00]. riparius [CFCP04]. Risk [TSJA02, AF09, AQS^{+}07, AFLB09, CB01, CSB03, CGHW05, FPDP06, GWK^{+}06, GG00, GP06, HXP^{+}09, HSC^{+}04, IVP08a, JBB^{+}05, Jag01, Jag09, JKJ06, KP00, LPFL09, LRT^{+}08, MW03, NS07, PWSS07, RDS07, SSR07, Sav00, SHCS04, WLB^{+}04, WEW01, WW09, WAB^{+}07, WJM^{+}03, WT04b]. risks [EYB^{+}02, VBD06]. Rissa [BKMB08]. River [AA05, ABV^{+}06, Ano06-47, BWPC06, CWH^{+}00, CGHW05, DLL^{+}09, GSB^{+}06b, HP09, JJK^{+}01, JRT02, KLL^{+}07, KCJ^{+}07, KLPP07, LWLZ06, LCY09, LXP^{+}08, PL02, RT08, RH04, SGP^{+}06, SGP^{+}07, WBP^{+}07, YBM^{+}05, YLJY03, ZCZ04, ABP05, CBFLS09, CDM^{+}00, DGGD04, DVGD07, FEP^{+}04, Foh05, GACO04, HKB02, HKHB06, HBRW07, KH02, KMPB03, Lin06, LFB07, MI01, MNPJ03, MBT03, MSP^{+}08, NV04, OFK08, PBA05, PNU03, PCWP06, PA09a, PBCZ01, PKS^{+}07, RLBL01, RSB02, RSW07, RMOA06, RFdV07, SBMJ09, TTPS09, WR08, WYO05, WAB^{+}07, vdPvOV00, CSR08, CW01, EPM^{+}04, HTMO06, JKWJ03, JJ00, MSL01, NKK^{+}09, PBA05, PdC06, Pen09, RNC08, SCB^{+}09, SR02, WVP08, WX09]. riverine [JRT02]. Rivero [CMG00]. rivers [BRP^{+}06, CWF03, DS02, HMBG03, HMPF05, Rey03, SP01]. RIVPACS [CWF03]. roach [LGL02]. road [GSZ08, GDL06a, JBB^{+}05]. roadmap [HK01]. roads [BTdFC05, JBB^{+}05]. Robert [Bro04a, Jør04a]. robust [EFHL05, KSvOO09, MRE^{+}06]. robustness [LHK00, PBPZ04]. rock [CGD04]. rocky [Oke04b, PP04, HBO07]. Rodent [BPM01, PCL^{+}05]. roe [VBRS07, RPRV09]. Roe-deer [RPRV09]. roja [MWM02]. Role [ABM^{+}06, DD02, MB09, PKC06, PCpDC09, AW06b, ASP^{+}07, AB06, ARF08, AKB07, BB08a, BBGM05, BBGM06, BCRSTVG07, BDJ^{+}09, BHI^{+}06, CGLS07, CMSC01, DG04b, GPB01, GHP08, Håk00, HGL^{+}06, Hel08, HvRIZ08, IA08, Kan04b, KGSB01, Leb05, LCG^{+}09, MKS^{+}02, NSEDP06, PLH09, RF09, RBR05, RKS^{+}07, RMGR09, SB06c, SRRS04, VM01, WSWL08, ZRCA08, dFOV07]. Roles [STK00, SBDD04b, VIK^{+}08]. roller [NS08]. ROM [Hey01a, Jor05g, Jor05i]. Romania [GPSM08]. ROMUL [CKN^{+}01, NMC^{+}06]. Rondônia [SSK^{+}07]. Root [SLE07, CWCH09, ŠH09b, WMM^{+}07, ZvBS05, ZJG^{+}06]. root-feeding [ZJG^{+}06]. root-zone [ZJG^{+}06]. rooting [TT05]. roots [MB05a, ZvBS05]. Roseate [MPRJ04]. Ross [HTMO06]. Rostherne [KGSB01]. rot [SWB06]. rotation [GARB09]. rotational [CTG03]. rotifer [KNB08a]. rough [LGC07]. round [ZAM^{+}03a, ZAM^{+}03b]. Routes [Nie06]. routine [SAH03]. row [LFJ^{+}06]. row-crop [LFJ^{+}06]. RS [dSA01]. Ruditapes [BSB^{+}09]. rule [BBB03, HKB02, KBW00, KCBS00, LiWZ00, RvGC^{+}08, iTI02, VADV06]. rule-based [BBB03, KBW00, RvGC^{+}08, VADV06]. rules [JT01, Jør02b, KCJ^{+}07, LC01, LNW00, RLHD01]. rumen [WR03]. ruminants [SB02]. run [Jon07, Kir06b]. run-time [Jon07]. running [BB03a, CCC04, CSSCC07, PCCL03, VADV06]. runoff [LH01, OFK08, SB06b, WGS^{+}02, Jor05k]. Rupicapra [FE04]. Ruppia [dSA01]. rupture [RB09]. rural [BDR01, CRM09, MCJ^{+}04, Sch00]. Russia [BCG08, Kir06b, KL08, KLL01]. Russian [MB08, ZBSA07]. Russian-olive [ZBSA07]. rust [AZM^{+}06]. Ruth [Jør04b]. Rutilus [LGL02]. ryegrass [GL04, LWL^{+}02].

S [Bro09, Hey01a, Hey01b, JT01, PG08, RLHD01, Xu03, Zha06, MMB07]. S. [Bro09, Cha07]. Sacca [MNZ06, VCRD06]. saccharum [Lar02]. sachalinensis [SK04]. saddlepoint [MKRH03]. SAFE [BPC04, NKK^{+}07, TBP^{+}07]. SAFE AIR [CCG07]. sage [PCHG03]. Saginaw [TTHH07]. Sahel [KRvL^{+}02]. Sahelian [BKGC05, SL04, TMHJ06, VA00]. saira [IMK^{+}07]. salicaria [Wel04]. salinity [CT01, KBGJS06, MM06a, MMM02, Met01, Met02, RMD04, TDdSLS^{+}08, UZ01, WHHH07]. salinity-stressed [KBGJS06]. salinization [SB01]. Salix [AMSW07]. SALMO [RvGC^{+}08, WRCB01, BRP^{+}06]. SALMO-OO [RvGC^{+}08]. salmoides [HL04]. salmon [AGZ05, HDB^{+}06, Mur06, RPPB04, RMGR09, SCH05, SM06]. salmonid [BDI04, SR02, WT04b]. salmonids [HHK07, MY02b]. salmons [FLS06]. Salt [CCMT09, LD06a, Met03, SFU01, SNF01, WHHH07]. salt-affected [Met03]. Salt-marshes [CCMT09]. Saltelli [Jor05a]. Salton [CB01, CSB03]. Salvelinus [MY02b, WPD04]. Sambar [PSVH09]. sample [ECZ^{+}06, GPC^{+}09, SP02, WPMT07]. samplers [WPBM06]. samples [PBHGF07]. sampling [CG06b, For02, For03, GTJ^{+}00, HG02, KM06, KRN04, Mun09, Nal01, PLLdCB06, RHB06, WLG07, XHH^{+}05]. Sanctuary [PSVH09]. sand [FWS^{+}05]. sandhill [DGM08]. sandy [NMC^{+}06]. Santa [KSG05]. Santojanni [KL09]. São [RdQFO07, AP00, dCSB08]. sap [COB^{+}06]. Sapling [BMB07]. saplings [FRM09]. Sardina [RHE06]. sardine [OYiI09, RHE06, SFM04]. Sardinella [LS09a]. Sardinia [FU05, LBS08b]. Sardinops [OYiI09]. SARS [Zha07]. sasakii [KL03]. Saskatchewan [LPC05]. SATCHMO [MWWM07]. satellite [KDK06, Las06a, LL07b, MCM^{+}09, PCL^{+}05, RLR05, RJS^{+}06, YSM00]. Satilla [ZCZ04]. sativa [CMB05]. saturated [KNZ04]. saturation [EM08]. Sau [HAG^{+}00]. saury [IMK^{+}07, MKiIK07]. Savanna [BH02, AR02, DW01, FCF^{+}04, LC07b, MWWM07, PSBJ07, RSA04, SRA05, SAR^{+}09, TMJ04, TEMJ06]. savannah [VA00, CGHW05]. savannas [MSW^{+}09, WSWL08]. scabies [LG03]. scalar [KYL07]. scale [Ash06, BNM09, BL02, BLDCM06, BSG07, BB03b, CLM^{+}09, CGS08, CCM02, CZG05, DDG^{+}05, DVdB^{+}08, DUASCM07, DKL^{+}09, DMBM05, ESZ^{+}00, Fat04c, FEP^{+}04, GSBM08, GARB09, HDBM02a, HDBM02b, HHL^{+}02, HSV07, HHKH09, HVB06, IO02, JWW^{+}07, JDPI07, KNM09, LMH05, LL02a, LL03a, LCG^{+}09, LZB^{+}06, MMV08, MMR02, MRiI^{+}07, MFJM04, MFG^{+}06, MDB^{+}02, MB02b, Nal01, OW08, OSS02, OPL^{+}09, OL09, OdKV03, PBS02, PSP^{+}07, PBH^{+}07, PKS^{+}07, PAF06, RvdKK^{+}02, Rob05, SPTP01, SOHY07, SM07b, SFM08, SMSV09, SV03b, SJ03, SSP03, TPC^{+}07, TYS^{+}09, TM01, TTB06, Vil01, WLNW08, WGT05, WLJ00, WSY^{+}07, YHC04, Yos06a, ZEGS08, ZCKR07, vW07]. scale-space [HDBM02a, HDBM02b]. scales [BW01a, CUF^{+}09, DBB^{+}08, FDP^{+}03, GBS00, GdBD08, KRvL^{+}02, LWBW05, MBKD02, OW05, PLA06, RG06, SF09, YLL^{+}05]. Scaling [BCRSTVG07, BSJ^{+}02, FBC02, MPP05, PTS^{+}04, ZdlP05, vN02, BMR07, CEMS05, GPL05, LC01, MCGO05, TAL^{+}05, Wir00, ZPW05, vNM02]. Scaling-up [FBC02]. scallop [CGH^{+}03, SIS^{+}07, WTMY07]. scatter [BJK09]. scavengers [WG04]. scavenging [DGSBG09]. SCE [SS08b]. scenario [BDE08, BBGH07, DDH^{+}09, PRRB09]. Scenarios [SDA^{+}03, YFLW06, AAKO^{+}08, ASHHRRPE04, BS08b, CKPP03, CN07b, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, GYY00, HHKH09, KPKP06, Med06, MBPS04, OW02, SVB09, TN08]. scene [GEH02]. Scenedesmus [KM00]. scentless [dCBL09]. schedules [BDD^{+}01]. Scheme [WGVB01, ZGF^{+}05, AOY02, CCJ07, GSZ08, MLA^{+}02, VA00, WLSE06]. schemes [PBD00]. Schismus [SF07]. school [ATA03, VS08a, Wil08]. schools [HBSM04, MOJ01, VPG05, VPG07]. Science [Hey01a, Jør06a, HK01, MRT05]. ScienceDirect [Mau07]. Sciences [Jor05j, Kri04a]. Scientific [Jor05a, LL08, Her03]. sclerophylls [Zav04]. Sclerotinia [dJBPG02]. sclerotiorum [dJBPG02]. Scolytidae [JPD^{+}06]. Scomber [BC01]. scombrus [BC01]. Scope [Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-37, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27]. scores [ÇDKK05]. Scotia [GB01]. Scots [DADGA06, Fle01, LSM^{+}04, Mii00]. Scottish [KIL^{+}03, Mur06]. scramble [EWH^{+}02]. scrapie [DDS^{+}04, SDD^{+}04]. scrofa [HAS07]. scrub [WCS00]. sculpin [RG05]. SD [Ano01a, BLDN00]. SE [Ort08, OACB09]. Sea [CT01, KL09, KGNK03, Leg03b, MSM^{+}07, San09, YHTH08, CB07a, CNG06, DD03, MM06c, MBM06, MSTK08, Oke04a, RHB06, Sno08, VFTB06, YHC04, Zue07, ATAK00, CB01, CSB03, CPTD08, CPT09, CTH^{+}00, EKBF04, ECBD09, HE05, JAK^{+}06, JCX^{+}08, KL07, KHJ^{+}08, LMGM^{+}09, Leg08, MNZ06, MDB^{+}06, MLGG09, OMRD01, Per07, SO04, San07, TSJA02]. sea-coastal [VFTB06]. Seabird [HD01, ZPP06]. seabirds [PFBBJ08]. seabream [HGLLV03]. seagrass [BHC04b, EFS^{+}03, MB06a, PCM^{+}03]. seagrasses [MMB07]. seals [BDP^{+}02, FMRS09, FKIL08]. search [SHM07]. searching [For02, For03]. seas [RŽČ^{+}04]. seascapes [PCC^{+}07]. seashore [RAH07]. season [ETH^{+}04, HB01, KRvL^{+}02, MKiIK07, YHC04]. Seasonal [HY07b, HS06, JNSS02, WKZP04, ASN02, BVD05, BPBF^{+}00, Bir01, CPH00, CR03, CWBR01, EFS^{+}03, FYN^{+}07, JLBS09, MMWMH07, OMRD01, OzDBS07, PCM^{+}03, Rob05, SSKN08, SDL08, SSdlMP^{+}08, WCHM02, WG07a, WDR06, WHP03]. seasonal/daily [SSKN08]. seasonality [GTJ^{+}00]. seasonally [HvRIZ08]. seawater [Leg03b]. sech [YPR^{+}05]. Second [Ano09-31, Ano09-32, Jør04e, Leg01, Wir00, WT04b, Jor05e, Nie00a, Ano06-46]. Second-order [WT04b]. Secondary [ABB06, BKC^{+}07, DVJ^{+}08, HC03b, MB06a, NMC^{+}06, QMH00, Roe00]. sectional [OAAF07]. Sector [WLL^{+}08, BFOS08]. sectors [MM06b]. security [zGsLcXwZ09]. sedentary [PEPB09]. Sediment [CGH^{+}05, AFT09, CJS^{+}07, CBS09, HHD01, KCY^{+}08, KFS06, MWH00, NWP06]. sedimentation [HGB04, SP01]. sedimentological [JLH01]. sediments [CPJ06, DWR09, JCB^{+}02, LAXP07, SP01, SU08, TVS00, TTPS09, dG04a]. seed [CS01, CMCD04, CCGMJ07, DH06, EGE^{+}08, GDL06a, LM07b, MRS05, SPD^{+}07, SJ05, SDDC07, TMJ04, WCH08]. seedbank [DSD^{+}08]. seedbeds [LL07a]. seedling [CCGMJ07, CL06, LL07a, LMMK09, ML06, SPD^{+}07]. seedlings [KK07]. seeds [KMRV07, SWB06]. Seedscape [EBH^{+}01]. seeking [lXzL02c]. segment [GACO04]. segmentation [BB03b]. segmentation/object [BB03b]. segregated [Mat01]. SEIB [SIK07]. Seine [CMG00, DSD^{+}09, EPM^{+}04]. select [CDM05, DGD03]. Selected [Ano06-47, Ano06-48, Ano09-33, LBL^{+}08, CSSCC07, GH09, KS08, MJ02, Wit02, Ano08a]. Selecting [BMG08, CPV07, Gra05, VSGW09, ALP03, RIGJM06]. Selection [Jør04e, Pra07, ACJT08, BA05, BCM^{+}08, BVNS02, CDM05, CRL09, CG06a, CG07, GG08a, GICB09, HGR08, IA08, JG08, LJ07, LKY^{+}06, MCQA08, NV04, OW05, OW06, OW08, PMLM08, PMJ08, SSR07, Sal06, SM07b, SK01, SKvdW09, SWBH08, VM07, War08, WDBK08, WA02, ZDR03, ZPO09]. Selective [BDBS08, JF04, KC03, PBC09]. selectivity [GR09, MF06]. selenium [VH07]. Self [AEK^{+}07, LiWZ00, RADD^{+}01, SWBJ01, SHK^{+}07, CGC01, DDH01, GL01b, GS06, HZHL05, Kar03, KLPP07, LKY^{+}06, LAPAMD07, LCSZ09, PLS^{+}06, Ric00, RMBM06, RL09, SAK00, TYK03, TYS^{+}09, WW09, WSWL08, WK04]. self-assembling [RL09]. self-organising [TYK03, WW09]. self-organization [LCSZ09, RL09, SAK00, WK04]. Self-organized [AEK^{+}07, RADD^{+}01, SWBJ01]. Self-organizing [SHK^{+}07, GL01b, GS06, KLPP07, LKY^{+}06, LAPAMD07, PLS^{+}06]. self-regulation [HZHL05]. self-similar [Ric00, RMBM06]. Self-thinning [LiWZ00, Kar03, TYS^{+}09, WSWL08]. Semi [SSNB07, AJJ^{+}05, ATAK00, GSM08, HRH^{+}05, KE07a, MY02a, MSTK08, MTVC05, MNEB01, Mur01, OI07, OPL^{+}09, PSBJ07, RHH05, RLLB09, RBR06, SvL04, TCGL03, THA02, WGB^{+}08, ZPW05]. semi-analytical [MTVC05]. semi-aquatic [OPL^{+}09]. semi-arid [HRH^{+}05, PSBJ07, RHH05, SvL04, TCGL03, THA02, WGB^{+}08, ZPW05]. semi-deciduous [OI07]. semi-desert [MY02a]. Semi-empirical [SSNB07, RBR06]. semi-enclosed [ATAK00, GSM08, MSTK08, Mur01]. semi-Lagrangian [KE07a]. semi-natural [MNEB01]. semi-quantitative [RLLB09]. semi-theoretical [AJJ^{+}05]. semiarid [ZGF^{+}05, dBWG04, dBWG05]. semiclausum [TNO^{+}09b]. Sense [Jør09b]. sensed [Asp02, DSA08, Dun08, FC06, LLLT08, MMV08, MLL^{+}05, Plu00, SAS06, TPJ08]. sensible [GP02]. sensing [BKGC05, BCMR03, CCJ07, CMB^{+}02, CGG^{+}05, GB08, HDB^{+}06, LWLZ06, LBBR01, MVMA08, Met01, Met02, PWZ^{+}09, RSC09, SAL07, SOA03, Szi00, TYZ^{+}05]. sensing-based [BKGC05, CCJ07, SOA03]. sensitive [BPBF^{+}00, CG06b, MWD05]. Sensitivity [CMCD04, DBBS03, HRJ^{+}00, HLR09, JG08, Jon07, LD06a, PRL07, PPR03, RVRL05, SHB06, WRB08, XRM08, YFH03, dCPC01, ÅSCP09, AB05a, BMG08, BH01, BO07, Bru05, CGLS07, CIM07, CS08, DFF07, FJ05, FH08, GQB05, GDL06b, GGB^{+}06, GH07, Håk00, HAS07, JO09, LBL^{+}08, LH07, Lin06, MC08, MRK^{+}07, OBR01, PBSOMG^{+}05, PWC^{+}09, PHBF07, PTGI09, QXW02, RPRV09, RGF00, SA07a, SPTP01, TBPF08, WLL^{+}08, Wil01, Yea04, YYR^{+}07, ZPD^{+}08, vNS03, Jor05a]. Sep [GPDF09]. separability [Las06a]. separating [HBM04]. separation [GPV08]. September [Ano00-40, Ano00-28, Ano01-39, Ano02-45, Ano02-37, Ano03z, Ano03-48, Ano03-57, Ano04-57, Ano04-53, Ano05-45, Ano06-27, Ano06-48, Ano08v, Ano08l, Ano08-27, Ano09t, Ano09u, DDŽ06, DŽD06, KL08, Leg01]. sequences [BCRSTVG07, LBS^{+}06, SVS04, TAL^{+}05]. Sequential [PLLdCB06]. sequestration [BRGS09, GSBM08, GARB09, JL07, LPC05, MGCK^{+}03, MLF^{+}06, OBE^{+}07]. serendipity [OI07]. Sergei [Jor06b]. sericea [LKH^{+}08]. Series [Jør04c, Jør04d, AMD^{+}03, Ban05, BB04, BWPC06, CM06a, CM07a, CG06a, CG07, FMM^{+}07, FBC02, HTMO06, IC02, KRZ07, KŞE04, LHC07, Las06b, NDM00, PAF06, RPPB04, SGP^{+}07, SDS01, TNO^{+}09b, WR01a, WBP^{+}07, ZSD^{+}08]. serrata [ASI^{+}08]. service [DGSBG09, PBB04]. service-oriented [PBB04]. services [NM09, PR01, Pra07, SYC04]. sessile [MBDB09]. seston [PA09b]. set [Hul04, LS08a, SYCU09]. Setchell [MWM02]. Seto [YHTH08]. sets [BHFMG05]. Setting [TSJ08, GEH02, HMF00]. settings [CP01]. settlement [POF08, WTMY07]. seven [BWPC06, GSB^{+}06b, SGP^{+}06, SGP^{+}07, WBP^{+}07]. seven-compartment [BWPC06, GSB^{+}06b, SGP^{+}06, SGP^{+}07, WBP^{+}07]. several [Hin09, Ond07, ZW06]. severe [TMM05]. severity [SSM^{+}09]. sewage [Mal01]. sewer [EPM^{+}04]. Sex [Jen00, GHP08, HPD09, LL07c, PGLS03, SÅ06, YTS03]. sex-biased [GHP08]. sex-pheromone-oriented [YTS03]. sex-specific [GHP08].

Sexual [GG08a, Mog02, VS08b]. shading [EVF^{+}07, Oke04a, SKvdW09, WBB05]. shallow [BTdFC05, BCLR04, BBKN03, BHP08, CBS09, CGH^{+}05, DFHP04, EJG05, ET04, FMC^{+}08, GSM08, GB03, KPK^{+}07, LKKL09, MP07, MS00, PDL06, PBPZ04, PCC^{+}07, SSPL^{+}08, SSS00, TB05, TB06a, TSF^{+}05, ZBD^{+}09]. shallow-water [GSM08, PCC^{+}07]. shallow-zone [LKKL09]. Shallowness [EKBF04]. Shan [LMG08]. Shannon [Ric02]. Shao [Jor05d, Jor05g]. shape [BS08a, OM02]. Shared [WC07]. sharing [MML00, McK01]. Shearwaters [YFH03, HMF00]. sheep [DH01, PCHG03, SDD^{+}04, TPJ08, WMH08]. shelf [AS00b, GBB^{+}09, SO04, Oke04a]. shell [RC06, SM07a]. shellfish [WDW00]. Shelterbelt [ZBSA07]. Shelterbelt-grown [ZBSA07]. Shelters [MCB06]. shift [IG02, NCM07, ZJBI03]. shifting [WBvDHZ09, vN02]. shifts [AER^{+}07, SFM04, Sri04, TB05, ZSD^{+}08, ZBD^{+}09]. shoal [MOJ01]. shoals [Mac00]. shoot [SRA05, SAR^{+}09, SRL^{+}00]. shore [MNZ06]. shorebird [WYMS07]. Short [CUS01, DFHP04, KW00, YFH03, CKP^{+}01, RFA^{+}01, RHP07, Yea04]. Short-tailed [YFH03]. Short-term [CUS01, KW00, CKP^{+}01, RFA^{+}01, RHP07, Yea04]. shortage [SMW02]. Should [OSS02, SLE07, VSGW09]. shoulders [Bas04]. shrew [VF07, WG07b, WG07a]. shrimp [ASZRRR08, Gri04, MLGV00a, MLGV00b, Odu04]. shrub [GB01, SL04, TMJ04, TEMJ06, WW08]. shrubland [PE07, Pet02]. shrublands [SPTP01]. Siberian [Ito05]. Sichuan [VPR^{+}09]. Sicilian [SL06]. Sicily [VTB^{+}08]. Sidney [BO07]. Siena [JT09]. Sierra [HNF09]. Sigmodon [WLG07]. significance [BBB03, KZH07, PG08, vWSH08]. significant [dJTMBGMP^{+}09]. sikkimensis [LM07b]. silica [CSR08]. silica-limited [CSR08]. SILVA [MAA^{+}09]. Silver [MAA^{+}09, MAK^{+}04, BAP^{+}06, JAB^{+}06, KB04]. Sim [IO02]. Sim-CYCLE [IO02]. SIMBA [TMD04]. SIMBA-POP [TMD04]. similar [Ric00, RMBM06]. similarities [BNM09]. similarity [RLR09, SAB03]. similarity-based [RLR09]. SIMPLB [Ito07]. Simple [CN09, Svi02, vODFS04, AW00, AYKY05, Bir01, CSR08, CHW07, Dew01, DMRP07, FL09, GORJ03, Gut07, HFL07, HG07, HZF07, Hua03, HCL00, IMK^{+}07, IOIA08, JSM03, KBW00, KK00b, KBvVK08, Lid01, MDGV09, MKRH03, MHSP^{+}06, Mur01, NK04, Ond07, SFU01, SC04, SCB^{+}09, SLT^{+}09, SKM^{+}06, TA05, VL08, VCD05, XBM^{+}07, YY06, vdBDR02]. simpler [Pet07]. simplest [KC08]. Simplicity [Jor05b]. simplification [CGW^{+}06]. Simplified [PG04, GVC09, Ito07, Oli03a, WHW^{+}05]. Simpson [GG08b]. Simpson-diversity [GG08b]. SimSphere [PWC^{+}09]. simulate [BGMP06, CMM^{+}07, LDM00, NK04, RGL^{+}07, RMDC04, SCB^{+}09, SFCP02, VJ01, VTL^{+}09, WSF^{+}00, YHG04, YP02]. Simulated [ASZRMH^{+}04, BM08, GGS08, JCE06, MMRLP06, RWM^{+}07, SBG06, BSTS^{+}02, BBGH07, Bye00, CN07a, CC05, FJ05, GGH08, KRZ07, LBS08a, LK07, LXP^{+}08, MWK07, MMPN07, MMN08, MCK07, MG02, Nal01, OJD04, PGLS03, RMD04, RMA08, TSBH09, VF07, WM08, WLJ00]. simulates [FHT^{+}09]. Simulating [ALO^{+}01, BWAM09, BJJ06, Cai05, CPJ06, CTG03, DW01, DMBM05, DCG01, EBH^{+}01, FSM^{+}01, GIZ^{+}03, GG04a, HGT^{+}00, HSC^{+}04, IWM^{+}06, LGR^{+}09, Lin03, LLS^{+}08, Oke04a, PBH^{+}07, PGM08, PSBJ07, Pyk04, RF09, RHB05, RCZ^{+}06, SVS04, SFM04, SWK^{+}05, SSM^{+}09, SPM^{+}08, TCGL03, WG04, WBR07, ZJN^{+}06, ZBZ^{+}09, ZPD^{+}08, ARL^{+}06, ASJD01, BPC04, BH02, BRC04, CMG00, CMCD04, ETH^{+}04, GTRF01, LG03, LZB^{+}06, LPM^{+}09, MSM^{+}08, MMV^{+}09, PG02, RNKG03, RIE01, RCGB08, SR08, SMSV09, SLGS00, WZ01, WCH08, Weg00, WSF^{+}02, vWBV02, CMPO05b]. Simulation [AG03, BCDJ00, CWH^{+}00, CKPP03, CTH^{+}00, EFJ^{+}08, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, FE04, HD00b, Jør04c, Jor05c, Jor05i, KSH^{+}03, KFB09, LB01, MWM02, MGH^{+}05, MLTN06, MLA^{+}02, MMWMH07, MLGV00a, MLGV00b, NJF^{+}08, OFK08, PBC09, PvdBFJ02, Rec03, Rob03, SB06b, SMW02, SHP^{+}07, WH04, WSS^{+}06, Wil08, YLLW02, ZWCL05, vdPvOV00, AFT09, ACJT08, Aum07, BKO08, BHC04a, BKS05a, BG01a, BG01b, BPW^{+}03, Ber02, BOB07, BKC^{+}07, BW04, Bro04b, BHI^{+}06, CD07, Cha02, CCJ07, CKA^{+}02, CDV05, CML09, Cou03, CP09, DMBO00, DH01, ET04, ESG06, GYY00, GBSP08, GGP03, GSG^{+}04, GCLG03, GPP02, HRH^{+}05, HGD05, HAG^{+}00, HNS08, HHL^{+}02, HPHP04, HPD09, HHC02, HZF07, IO02, IVC^{+}08, JMVvDV02, Jon07, KMR^{+}07, Kaz07, KPH02, KSRWG07, KM00, LS04a, LTM^{+}04, LLA^{+}09a, LD06b, LELR02, LC07b, LPFL09]. simulation [Lir03, LOM06, LGS^{+}00, MB08, MP04a, MU02, Mat06, MMPT06, MLPK01, McK00, MH05, MD07, MWWM07, MDH^{+}08, MTVC05, MVGH00, MMJN03, MHZ^{+}06, MNZC08, NSO^{+}08, NWH^{+}06, NH07, OYiI09, OIR^{+}08, OW02, OL09, PK01, PK02, PPE^{+}07, PJAZ02, PGFM04, POH02, PdAD^{+}04, PTS^{+}04, PCS01, PP04, PS01, RG05, RvGC^{+}08, RGF00, Roe01, SA07a, SKP^{+}07, SM04b, SDS^{+}07, SFM08, SL02, SHCS04, SRS^{+}03, SB02, SWCO07, dSLSD02, SLT^{+}09, SF04, TC09, TDdSLS^{+}08, TFF07, TR03, TMD04, TYS^{+}09, TK07, VS08a, VS05, VPSG05, WC03, WSCR03, Wan05, WAB^{+}07, WM06, WSWL08, WH07, XHH^{+}04, XHH^{+}05, Yos06a, ZPP06, ZLZM02, ZHMN02, dSA01, dSSGR00, dSMZ09, vNSvdBC03, KH02]. simulation-through-optimization [CD07]. Simulations [BP08, FGB08, LM07a, MTKM^{+}06, OMRD01, AVP08, AN02, BRGS09, Bir06, BDK01, BVC^{+}01, BKC^{+}07, BLBT01, BL04, CUS01, CWCH01, GPV08, HHP06, KCD^{+}04, KLM^{+}02, KJ08, LLF02, LL06, LWL^{+}02, MKB00, MBLA03, MAL07, Ort08, RLLB09, RAI04, SA07a, TMM05]. Simulator [LCF09, FPK^{+}07, LLW^{+}06, MAA^{+}09, RSB02, SR02, VBRS07, AS00a]. simulators [LMMK09, Zav04]. Simultaneous [HGR08, TLW01]. Sinauer [Log02]. sinensis [WX09]. Single [FA03, ADdSC06, BKC^{+}07, CS01, FBDM09, HPD09, IC02, KDW^{+}09a, LW09, MB05a, Sak07, URB06]. single- [KDW^{+}09a]. single-criterion [URB06]. Single-species [FA03, ADdSC06, LW09, Sak07]. singular [Yun08]. sink [AHP01, Wal01]. sinks [Fig09]. site [BNM09, LKY^{+}06, MTD^{+}09, MOP^{+}06, New06, PFBBJ08, RGF00, Saa00, SSR07, SRB06, WKZ03, WYMS07, CGHW05]. site-specific [RGF00]. sites [FDCH08, FCKH06, KKA^{+}08, KBE^{+}06, Tol06, WBTC00, YL07]. Sitona [ZJG^{+}06]. Sitting [Bas04]. situ [BCMR03, SC01, LKKL09]. situation [WSF^{+}00, WSF^{+}02, YWHW02]. situation-driven [WSF^{+}02]. situations [AZM^{+}06]. six [HB09b, ZGH05]. Size [EN06, HS01, MCD^{+}08, AFTB07, ALM00, BS07, BLHB06, CM05b, CGGE08, Deb02, DH01, FBF^{+}06, FAB^{+}07, GLHV08, GR09, GPL05, Hen07, ICGÁ05, JRBS02, MEKL08, MBP^{+}08, MRT05, MKM^{+}07, Mon05, Nie09b, OzDBS07, PBHGF07, PBV05, RM02, RFdV07, SSS^{+}09, SMB^{+}06, SHW04, SP02, THA02, VH09, VPG05, WPMT07, WSWL08, Wil08, YFH03, Yos06b]. size-classified [PBHGF07]. Size-dependent [MCD^{+}08, Nie09b]. size-distribution [RM02]. size-heterogeneous [GLHV08]. size-resolved [BS07]. size-structured [ALM00, SSS^{+}09]. size/age [SMB^{+}06]. sized [JU01, XBM^{+}08a]. sizes [GW04, MKM^{+}08, MCGO05, RBSJ01]. skewed [Aub04a]. skies [WTS^{+}06]. skottsbergii [MWM02]. sky [PG04]. Slope [BHW^{+}08]. Slovakia [Ond07]. Slovenia [Ano06-48, BBT06, DDJ^{+}01, DDŽ06, DŽD06, JDD^{+}03, JPD^{+}06, LD06a, Med06]. slow [RG04]. slug [CBP^{+}06, SB03, SB06c]. small [ADSO08, AFT09, BPW^{+}03, CGS08, CdQO06, DDLB02, DUASCM07, DG04b, DMBM05, GB01, HGD05, LS09a, LE04, LCG^{+}09, MBT03, MBCS07, MDB^{+}02, NSEDP06, ORS^{+}09, PBHGF07, PBH^{+}07, PSBJ07, Reu05, RFdV07, TMLV07, VPR^{+}09, WPMT07, WLNW08, ZW06]. small- [CGS08]. small-scale [DUASCM07, DMBM05, MDB^{+}02, WLNW08]. smallmouth [CCLS06]. Smart [PV04]. smoke [ESZ^{+}00]. snail [MBKD02, RRB^{+}01]. snakehead [JLAM09]. snappers [Esp03]. snow [HLR06, LHT^{+}08, OW05, OW06, OW08, TTA^{+}01]. snow-glide [LHT^{+}08]. SOC [WZJ08]. social [CFS09, GPP02, JLBS09, Jua09, LM07a, PGCK04, PM06b, PBK03]. socially [CEK08]. societal [CQ07, GPP02, LMGM^{+}09]. Society [Ano06-45, Ano07a, Ano08a, Ano09i, LMPR06, LBL^{+}08, Ano04-94, CC06, Kin04]. socio [LPCZ07, LMGM^{+}09, LHC09, MLPK01, WFM01]. socio-ecological [LPCZ07, LMGM^{+}09]. socio-economic [LHC09, WFM01]. socio-economic-ecological [MLPK01]. socioeconomic [ALO^{+}01, SDA^{+}03, WZ01]. socioeconomics [Liu01]. soft [RC06]. soft-shell [RC06]. Softbound [Jor05e, Jør09b]. Software [Ano08a, LBL^{+}08, ABC03, ABC04, Cal06, Kaz07, LBS^{+}06, Moi04, RNC08, WBS^{+}02, ZWCL05]. Soil [FCC^{+}00, Jor05c, MBP^{+}08, RSC09, XS08, APJ03, BPA08, BPC04, Bio01, Bio03, BBGH07, BKC^{+}07, BCG08, BJ02, BHV06, CBMP07, CEMS05, CKN^{+}01, DVPS08, DSD^{+}08, DDL^{+}06, ETH^{+}04, FCF^{+}04, FGB08, GSBM08, GARB09, GFGZ08, GRPF07, GVC09, GPL05, GJ00, HBO07, HTS00, HHB^{+}08, HRJ^{+}00, HVVK09, IMS07, Ito07, IOIA08, IWM^{+}06, JZY07, JZC^{+}07, KA01, KFN^{+}08, KIL^{+}03, KBE^{+}06, KRvL^{+}02, LIS07, LZZ^{+}07, LGR^{+}09, LWBW05, MCBA07, MBD^{+}00, MBGP08, Mon05, MMJN03, NMC^{+}06, NKK^{+}07, Nie08c, ORS^{+}09, OHH07, PPR03, PPE^{+}07, POH02, PWC^{+}09, PHH00, PKS^{+}07, Pot04, PBML08, QXW02, RMSS02, RLR05, RVRL05, SI07, SFC05, SAR^{+}09, TH06a, TYT^{+}09, TBP^{+}07, TVK^{+}08, VMG05, VPSG05, WCHM02, WGS^{+}02, WHHH07, WM06, XRM08, YWHW02, YABM07, ZLZM02, Zha03a, ZPK^{+}07, ZSZ06, ZAM^{+}07, ZCQ^{+}09, dBWG04, dBWG05, vdPLG^{+}00, Kir06a]. soil-dwelling [GJ00, JZC^{+}07]. Soil-moisture [Kir06a]. soils [AYKY05, BRK07, BRGS09, GCM^{+}05, LPU^{+}07, LPPS05, MPP05, MCGO05, Met03, NMC^{+}06, NK04, OBE^{+}07, TKHS^{+}07, YL07, ZCKR07]. solar [AMK00, BCST05, BCRP09, Dun08, GGH08, LD06a, SS01, SSP03, WCC02, WHP01, YBM^{+}05]. solitary [JCE06, MAL06]. solute [SHZ05]. solution [AALM05, LK02, Mar06, XS08]. solutions [Jor05g, LL08, MHvIR00, MCBA07, TSF07]. Solvable [Jor06b]. solve [ADdSC06, DAdSC08, KE07a]. solved [HD00a, KA01]. Solving [ZVLP05, CPV07]. SOM [GWK^{+}06]. Some [BCM^{+}08, KIL^{+}03, LG00, RLFB04, AWL04, Aus07, GL01b, Tud01, WXYMZZF03, Yun08]. songbird [VSHC08]. Sonoran [SWK^{+}05]. Sons [Jor05a, Jor05g, Jor05k]. sooty [JDPI07]. sooty-beech [JDPI07]. sordidus [VTL^{+}09]. Sorex [WG07b]. sorption [CM08]. SORTIE [BMB07, MDB^{+}02, TMM05]. sound [DS02]. source [OWHS09, RBB05, Wal01, dBWG04]. sources [ATAK00, CCG07, Fig09, HOK^{+}09, MBF09, OEK^{+}06, ZTXD03]. South [HRH^{+}05, KCJ^{+}07, MSL01, RBW09, CMPO05b, GE05, KRK05, OB04, PM06a, RNKG03, RBE^{+}08, SL04, SSPL^{+}08, ZBZ^{+}09, ZBL03, CdQO06, CPTD08, DDJ^{+}01, ESZ^{+}00, FMC^{+}08, GRW04, HUB02, HBUS02, MB06a, MAA^{+}09, RvGC^{+}08, RR05b, TCGL03]. South-central [DDJ^{+}01]. south-east [RNKG03]. south-eastern [OB04, RBE^{+}08, ZBL03]. south-west [SL04, MAA^{+}09]. south-western [CMPO05b, KRK05]. Southampton [Jor05h]. southeastern [KFJ^{+}09, LAD06, LR07b, PCHG03, SS08b]. Southern [Ano06-47, MBLA03, RMBM06, VTB^{+}08, VML^{+}06, ASZRRR08, BLDCM06, Cha02, CZG05, ECHN08, FBDM09, LSM^{+}04, LPFL09, MWM02, MNOS08, MLR08, RG05, SSH^{+}07, SAS06, SFM04, SF04, TM05a, TAL^{+}05, Wil08, FMRS09, FSM^{+}01, MTD^{+}09]. Southhampton [Jør04g]. Southwest [MWP07]. southwestern [ASZRMH^{+}04, HBUS02, LR03, MCSG06, CWBR01]. sown [WSY^{+}07]. soybean [HPA00, PMH00]. sp [AMSM06, HCJ^{+}09]. space [AE02, BRV09, BNTK04, Cal06, CB07a, DDLD07, Epp00, GG02, GRC^{+}07, HDBM02a, HDBM02b, Hin09, HGR08, KFR07, Lid01, LCR06, Mat03b, MBGP08, PF01, RF09, RPPB04, RB09, Sno08, Wan07, WG07a]. space-dependent [PF01]. space-independent [PF01]. spaces [Pet07]. spacing [BH00]. SPACSYS [WMM^{+}07]. Spain [BBM06, DADGA06, HAG^{+}00, BCCR^{+}02, BRV09, SLPP05]. SPAnDX [KBM^{+}03]. Spanish [SDA^{+}03]. Sparse [STH06, LD06b, MDB^{+}06]. Sparus [HGLLV03, LS08b]. Spatial [ACE07, ASI^{+}08, APJ03, AS00c, Aus02, BK05a, BMBOCR03, BJJ03, CGC01, CFS09, DLPN06, Epp00, FBF^{+}06, GLL^{+}07, GLM02, GB01, HSRD09, HAS07, IWW08, IiTS^{+}04, JKS^{+}05, JJ00, Ken02, LXJ^{+}03, MBM00, MFB^{+}05, MAL07, OACB09, OdKV03, PDBH02, PE02, PMC03, PBML08, RB02, RA06, RMS08, SK04, Shi04a, SF04, WCHM02, WGS^{+}02, WM08, ZGH05, ZPK^{+}07, ZBL03, vdPLG^{+}00, AMM04, APG^{+}03, AEP^{+}04, AEK^{+}07, APEA09, ALAS09, AN02, Aru05, Asp02, BKO08, BPAB^{+}06, BMB07, BCL^{+}09, BDR01, BLJS05, BDF^{+}06, Bio01, Bio03, BLB08, BLB09, BSTS^{+}02, BFE07, BSJ^{+}02, BL02, BLDCM06, BVC^{+}01, BKC^{+}07, BCH09, CLM^{+}09, CEMS05, CK07a, CGS08, CMA^{+}06, CSHY08, CKA^{+}02, CBP^{+}06, CZG05, Cou03, DGM08, DLK01, DMF07, DH01, DBB^{+}08, EN06, FL09, Fie04, FMC^{+}08, FH08, FSJ04b, GLD07, GL06]. spatial [Gau06, GCLG03, GS09, Gre06, GRPF07, HP09, Hör03, HC05b, IKS09, KCD^{+}04, KW02a, KW05, KFB09, KKL^{+}06, LRJ^{+}09, LS01, LHC09, LOL02, LOL03a, LOL03b, LHT^{+}08, LR07b, LG03, zLGH^{+}05, LWBW05, LOM06, LS06, LMPT05, MKB00, MdRdA06, MWK07, MM06a, MBDB09, MJR06, MCA^{+}06, Met01, Met02, MWWM07, MF02a, MY02a, MFA07, MDB^{+}06, Mon02, MBKD02, MB06b, MHN00, NP06, Nal01, NHLPA06, OCK01, OWHS09, dSPdBB08a, PS01, PFFM07, PWSS07, QMH00, RAH07, RMF09, RJR04, SI07, SSK^{+}07, SFU01, SDS^{+}07, SI06, SAR^{+}09, SHCS04, SLL^{+}06, SRS^{+}03, Sil07, SMET06, SB07, SHZL09, SJLL08, SHM07, TTA^{+}01, TK01, TP06, TMS^{+}07, TM01, TKTB07, VTB^{+}08, VPR^{+}09, VML^{+}06, WZW05, WCH08, WW03, Whi00, Wim04, WLJ00, WHZ06, WLZ^{+}09, WWC^{+}07, XCW07, XHH^{+}05, YLL^{+}05, ZLO02, ZR04, Zha03a]. spatial [ZZL^{+}09, ZNC^{+}06, dFOV07, vNSvdBC03, Jor05l]. Spatially [BHW^{+}08, BM07, ECHN08, PGFM04, SRG^{+}09, SSNHP08, ZHMN02, ASG^{+}05, ASJD01, ACJT08, BFD02, BS08b, Ber02, BH00, BH02, CEK08, DC08b, DFCM01, ESG06, EN08, FAB^{+}07, GLD07, GH09, Gri04, HLHZ06, HZF07, HD01, JBT^{+}05, KW02a, KJ08, KD05, LC07a, LC04, LWC^{+}07, LPM^{+}07, LDM00, LHZ^{+}06, Loe00, MP04a, MPM02, MSM^{+}07, MCKNM09, Mat01, MP04b, MSD06, NH07, OG08, OSS02, PEM06, PE07, PBY^{+}03, PdAD^{+}04, PHD04, PWSS07, RHB06, RL05a, SIK07, SBM04, SV02, SVB09, SRR06, Svi02, SHM07, TFF07, TMLV07, VTL^{+}09, WT04a, WBS^{+}02, WD02, XHH^{+}05, YP02, vHPS02]. spatially-explicit [DFCM01, KJ08]. spatially-structured [GLD07, LDM00]. Spatio [AKMB01, EBM06, GML09, BKGC05, CBMP07, KS07, LAB^{+}05, LL06, LZB^{+}06, MEJ06, MDB^{+}02, PL04, SRN05, SZL^{+}04, VMR09, ZCQ^{+}09]. Spatio-temporal [AKMB01, EBM06, GML09, BKGC05, CBMP07, KS07, LAB^{+}05, LL06, LZB^{+}06, MEJ06, MDB^{+}02, PL04, SRN05, SZL^{+}04, VMR09, ZCQ^{+}09]. Spatiotemporal [Her08, ZLH06, HB05, SRRS04, ZZL^{+}09]. spawner [NCM07]. spawner-recruitment [NCM07]. spawning [CDM^{+}00, MSP^{+}08, MKiIK07, NSO^{+}08, SMB^{+}06, VS08a, WX09, WTMY07, dlS07]. Special [Ano06-44, FK03, KW02b, McK00]. specie [ALP03, Aru05, BHP05, HSRD09, RPVR03, Tis06, dJTMBGMP^{+}09]. Species [Ano06-46, Aus07, FPK^{+}07, KBW00, PDBH02, SCAP05, AR02, AF09, ADdSC06, AP06, Asp02, Aus02, ABM^{+}06, BKO08, BSR06, BHP05, BP07, BBGH08, BCH09, CK07a, CM06a, CM07a, CN07a, CMM02, CPJ06, CSdZ09, CWS09, CG06b, DD02, DMH^{+}03, DMBM05, DH00, DMRP07, ECZ^{+}06, EBR02, EFHL05, EGE^{+}08, FGEG06, FG08, FA03, FC06, FLS06, GG04a, GLP05, GWK^{+}06, GB02, GIKS08, GEH02, Han09, HHL08, Har08, Hee02, HPD09, HB06, HLK06, HHM01, HLH^{+}06, HLS06, JAN^{+}03, JDD^{+}03, JLAM09, JG08, JKJ06, JOBL08, JcLM09, KC03, Kar04, KH07, KK00b, LLW^{+}06, LJ07, LJR06, LKH^{+}08, LEH06, LCP06, Lin05, LRT^{+}08, LL06, LPM^{+}09, LW09, LSAGF05, MBDB09, MLHT09, MPRJ04, MCSG06, McK01, MJR06, MBW09a, MFG^{+}06, MMF^{+}09b, MMF^{+}09a, MEO06, MFB^{+}06, MDJ09, MWH00, MDVG09, MCK07, NiTT01, NAC04, OM02]. species [OIR^{+}08, OHH07, PCCL03, PRL07, PLH09, PF00b, PMD^{+}09, Pet02, PAS06, PF03, PCC^{+}07, PWH07, PJGW06, PE06, RDS07, RD07, RSA04, Rec03, RVWH06, Rem04, RBR05, RBR06, RCL06, RHP07, Sak07, SCBG09, SH09a, SF09, SCPC^{+}07, SAR^{+}09, SL04, Sil07, SGH^{+}08, SM06, SLZ09, SvL04, SP02, SBS^{+}06, SB07, SGF08, SGF09, TC09, TDdSLS^{+}08, TMP06, TMM05, VM09, VJ06, VCP09, VML^{+}06, WZKL09, WBR08, WW09, Wel04, WR03, WTL00, WBTC00, Xia07, lXzL02a, YSB08, YPR^{+}05, Yos03, ZLO02, ZBWR06, ZBSA07, vdHGF09, PDBH02]. species- [HHL08, MDJ09]. species-distribution [JG08]. species-rich [EGE^{+}08, ZBWR06]. Species-specific [SCAP05]. Specific [DLP^{+}09a, ÅSCP09, ARF08, CK07b, FF01, GGV^{+}06, GHP08, HA08a, HGR08, LL06, LP03b, RGF00, SCAP05, SHSP04, TR03, WT04a, WSF^{+}00]. spectra [CGGE08, HS01, MFG^{+}06]. Spectral [MCK07, ASI^{+}08, GLD07, Las06a, SAS06]. speeds [GPK00]. spent [BKMB08]. spiders [KBW00, ŽJDK06]. Spiekeroog [KE07b]. SPIEL [SHB04, SHB06]. spill [Per07]. spin [TR05]. spin-up [TR05]. Spirogyra [HCJ^{+}09]. splash [PvdBFJ02]. spline [MSL01]. splines [LEH06, WA02]. SPM [HMPF05, HE05, KHJ^{+}08]. spoils [SWB06]. SPOMSIM [Moi04]. spore [SBJ^{+}02]. spores [KMRV07]. spot [PvdBW^{+}02, Las06a, Las06b, YSM00]. SPOT-Vegetation [Las06a]. SPOT/VEGETATION [Las06b]. spotted [MY02b]. spp [AACIS^{+}08, BMBOCR03, CTG03, CTG04, JDPI07, MSL01, TCGL03, dGdMPC05]. sprawl [Gan06]. Spread [WH07, AF09, BMF^{+}06, DDS^{+}04, GL06, GdBD08, LH05, LH06, LKH^{+}08, Mal01, PvdBFJ02, RMGR09, SSPR03, WKL^{+}03, YMD08, dCBL09]. spreading [MSB07, MNZC08]. spreadsheet [VJ01]. spreadsheet-based [VJ01]. spreadsheets [MBT03]. spring [ARL^{+}06, EVF^{+}07, Par02]. Springer [Jør04b, Jør04e, Jor05d, Jor05f, Jor05i, Jor05j]. Springer-Verlag [Jør04b, Jør04e, Jor05f]. Springs [KB04]. Spruce [JPD^{+}06, BB04, BP08, FBC02, KS04a, KS08, MAA^{+}09, Mii00, MAK^{+}04, New06, NJF^{+}08, OHH07, SBR^{+}07, SGLH04, SPM^{+}08, WFB^{+}08, YTH03, ZNC^{+}06]. SPUR [PCS01]. square [KC05]. squared [EN08]. squid [RLSZK^{+}08]. squirrels [WPMT07]. St [SSH^{+}07]. Stability [AER^{+}07, hMzL08, ZZZ06, Aub04a, BFD02, CG09, Esp03, GBA08, GORJ03, HE09, HLR09, IVP08a, IVP08b, LC01, MC01, MLM06, MWP07, OS02, PRL07, PM09, PEAS01, PBM^{+}05, SLS03, Sch03, Tar08, vdHGF09]. stabilization [ABR05, BK05b, KMM^{+}00]. stabilize [KK00b]. stable [Aub04a, KVV01, YYZ06, YY06, ZHL07]. Stage [HZHL05, CAG03, DAdSC08, GL06, Gra04, GG05, RPPB04, SKCM07, SÅ06, TSZdRR03, TMS^{+}07, dCBL09]. stage-dependent [TSZdRR03]. Stage-equilibrium [HZHL05]. stage-structured [CAG03, DAdSC08, GL06, RPPB04, dCBL09]. stages [BC01, Bey03, GL08, HB06, KCBS00, Leb05]. stakeholders [GBSP08]. stale [LL07a]. stand [BLDCM06, BAP^{+}06, CKL^{+}06, CS07, CKPP03, DDG^{+}05, DVdB^{+}08, Kar03, LZB^{+}06, New09, Oga09b, PGFM04, Rob03, RE03, SAB03, SRG^{+}09, SPM^{+}08, SJG^{+}08, TYS^{+}09, WFB^{+}08, WRB08, ZdlP05]. stand-scale [DVdB^{+}08, TYS^{+}09]. stand-types [New09]. standard [GBB^{+}06, PKC06, Tia06a, Tia06b]. standardized [CSM^{+}06]. standards [KDW^{+}09b]. standpoint [BCPM09]. stands [BS04, DMRP07, FBDM09, GS09, KRN04, KS08, Lar02, MAA^{+}09, MM00, New06, SMG07, ZBW05, ZPD06]. State [BNTK04, GRC^{+}07, HGD05, WG07a, AB03, CSSCC07, CB07a, CLTH08, ESG06, GSB^{+}06b, Hin09, KCY^{+}08, KB04, KGZR05, LP03a, LP03b, MLHT09, MR06, MR08, Pet04b, RPPB04, SGP^{+}06, Sno08, TR05, Wan07]. state-of-the-art [MLHT09]. state-parameter [CLTH08]. State-space [BNTK04, GRC^{+}07, WG07a, CB07a, Hin09, RPPB04, Sno08, Wan07]. States [KR04, GLS07, KVV01, MRRJ07, KRZ07, MCSG06, OB04, SHCS04, SS08b]. static [HNS08, MD06a, Pen00]. station [WLB^{+}04]. stationary [TA05]. stations [DGSBG09, FYN^{+}07, VZG05]. statistical [AN06, ACE07, Aus02, ABM^{+}06, CCG07, CWS09, GL01b, JKWJ03, LEH06, LR09, MHM^{+}03, MOP^{+}05, MBS^{+}09, OLB04, PDL06, QKR03, RSC09, Ric00, Shi04b, TM05a, ZW02]. statistics [BCM^{+}08, HR03, RBEZ08, SA07b]. status [BdB05, BPC04, BAP^{+}06, BFS03, BFS05, TBP^{+}07, Tis06, TPC^{+}07, XZD^{+}06]. stay [TW00]. StDM [SSPL^{+}08, CSSCC07]. Steady [AB03, GSB^{+}06b, SGP^{+}06, KCY^{+}08, KB04, TR05]. Steady-state [GSB^{+}06b, SGP^{+}06, KB04]. Stefano [Jor05a]. Stella [Jor05i, CV01]. Steller [CNG06]. stem [KK07, PvdBFJ02, SDL08]. stemmed [ZBSA07]. stems [HHL08]. step [BHFMG05, CCBB05, CMA^{+}06, FC05a, FG07, ÖÖ04, TPC^{+}07]. Stephens [Jor05c, JJWF07]. steppe [JZY07, RR05b, WZJ08]. stepwise [PMJ08]. sterile [TTJ07]. sterilis [WCH08]. stick [Bir06]. still [JG05, NW07, XCW07]. Stipa [DKRŠS00]. Stochastic [Cha02, GDL06a, KID^{+}07, LCR06, NS07, Pra08, WPBM06, WYO05, lXzL02c, ÅSCP09, AS00a, ASN02, BKPS08, BW04, CCC04, CSSCC07, CAG03, CGR03, DVPS08, FK05, GGB^{+}06, KPT^{+}09, LMMK09, LG03, Lid01, LDM00, LW09, MC08, MKS^{+}09, MKRH03, MSS02, MG05, Moi04, MFB^{+}06, OG08, RC06, Sai07, Sar04, SCPC^{+}07, SAR^{+}09, SFCP02, SDS01, TH08, TN01, TIJ^{+}01, VCP09, WG07a, lXzL02a, YP02, Yos06a, YKO05, SSPL^{+}08]. stochastic-dynamic [CCC04, CSSCC07]. stochastic/deterministic [MSS02]. stochasticity [GHP08, JBR07]. stock [Aub04b, CB07a, LAZ^{+}08, MAdlPR02, PMC08, Sno08, WTMY07, Xia00a, Xia02, dlS07]. stocked [Jen01]. stocking [MR00]. stocks [CGH^{+}03, FLS06, RGF00, SRR06, vN02]. stoichiometric [TFM01]. stoichiometry [MB07]. Stolby [ESG06]. stomatal [BMD09, GQB05, OT03, WYT09]. stops [HPD09]. Storage [OSHO09, BBP^{+}07, DVdB^{+}08, LC08, RBC09, YL07]. storage-based [BBP^{+}07]. stored [JOB04, LJ09]. storm [Lir03, TMM05]. stormwater [CB07c, TB06b]. Strait [KGJ08, Per07]. Straskraba [Jør01b]. strategic [LMG08, PdAD^{+}04]. strategies [AS09, AMSH08, ABC03, ASZRMH^{+}04, ASZRRR08, BG01a, BG01b, CEK08, GBSP08, GRW04, LS01, LBS08a, LGS^{+}00, MVZM05, MWM02, MFJM04, MLGV00b, MDGK01, Oke04b, Ort08, PCB07, PWH07, PPP05, PFFM07, RSC09, RIGJM06, RHP07, TMJ^{+}07, ZRASC04, vWBV02]. strategy [ADS^{+}07, CG06c, GGP03, HG02, LH05, LH06, LC08, OACB09, RR06, SRN05, VS05]. stratification [MSTK08, WLJ00]. stratified [CPH00, HvRIZ08, KFS06, NP06]. stray [VSGW09]. Stream [HBCL07, LTLH08, BNM09, CGC01, CBFLS09, HBO07, HRMC01, JT01, KKA^{+}08, KHT06, LGC07, LMM^{+}07, LCM^{+}09a, MY02b, NV04, RLHD01, RG05, SBA06, Sin07, TYLL07, ZAM^{+}05, YKO05]. stream-resident [MY02b]. streams [HFSH06, KRK05, NSEDP06, PKC^{+}01, WHX^{+}03]. street [KPKP06]. strength [LS04b, OzDBS07]. strengths [Nal01, QHM05]. stress [Dun08, FDP^{+}03, GIKS08, GZY^{+}06, GB08, HPA00, MVMA08, MMV08, PSPD09, TKHS^{+}07, WBR08]. stress-using [GB08]. stressed [KBGJS06, LYC08]. stresses [TB08]. stressors [MWMN06, TOS09]. strike [DCP^{+}07]. strip [EFHL05]. strongly [MAB01, NDM00, RBSG06, SJ05]. Strontium [Lim04]. Structural [Fat07b, JMN02, LZZA09, Len07, LFB07, TIJ^{+}01, ASS^{+}06, Esc05, HFL07, KRN04, KvKV05, PS05, RKH05, RHB05, SSS00, TDHO07, WFB^{+}08]. structurally [GZJ06, ZJTB03, ZJM04]. Structure [BP03, LUKD06, WSF^{+}02, AMM04, ASAC02, AS00c, Aub04a, Bar04, BCPM09, BPE^{+}07, BCH09, DFF07, FLS06, GYY00, GLM02, GS09, Gre06, GG05, HAG^{+}00, HLR09, Hua03, JKJ06, JOBL08, JVL02, KB04, KP09, KW05, KBE^{+}06, LSHG08, LCLC07, LMPT05, Mac00, MFG^{+}08, MMF^{+}09b, MMF^{+}09a, MZASLMLC04, MSL06, Mun09, NAC04, Ort08, PBE^{+}07, PPE^{+}07, PBC09, PBK03, PSBJ07, QHM05, RJR04, Rob03, San07, SAK00, SDA^{+}03, SL02, SMB^{+}06, SBB09, SW03, SM03, SRG^{+}09, SHZL09, SZ03, Tsc02, UDB07, VLA^{+}06, VPG07, VF07, WZW05, WK04, WBN^{+}03, WSZS08, XCW07, lXzL02b, YP02, ZLX09, ZPD^{+}08, ZAM^{+}03a, CMPO05a]. structured [AALM05, ALM00, Ano03-27, Aub04b, CAG03, CDAGK06, CEK08, DAdSC08, EN08, EN06, GLD07, GL06, HLHZ06, HCL06, HMG06, Jen00, JM01, JBMBP02, JBMBP03, KT03, KW02a, KD05, LPM^{+}07, LDM00, MPM02, MSM^{+}07, RPPB04, SSS^{+}09, SÅ06, Tar08, WT04a, dCBL09]. structures [BKB00, Cal05, LHB08, MAL07, MBGP08, MMLR07, VS08a, WFM01]. Structuring [CGGE08, Håk00]. student [RLR05]. studied [EM08]. studies [BMR^{+}05, CDM05, GEH02, HvI01, ICGÁ05, KSÇ^{+}00, KWD^{+}04, MCQA08, MD04, MSTK08, PT08, RHM^{+}05, SBA06, SH04, WIMK07, XHC^{+}08, Jor05j]. Study [HJ02, NDD^{+}07, WVP08, AW06b, AAA00, AN06, AP02, BGL01, BCG08, CSG02, CSCB04, CJS^{+}02, CC06, CL08b, CGHW05, CML09, DWR09, Don06, DDFP07, DS01, DAR^{+}07, DSD^{+}09, FEP^{+}04, FBDM09, GHP^{+}09, GPDF09, GGPBEK07, GGV^{+}06, GDL03, GWK^{+}06, GGB^{+}06, GP02, zGsLcXwZ09, HPA00, HHL^{+}02, HPHP04, HPD09, HvI01, HB09b, IF07, IOIA08, JAN^{+}03, JJ00, KMB08, KTKR08, Kri03, KHJ^{+}08, KM00, KNB08b, LS09a, LD06b, LWLZ06, LZZ^{+}07, LCY09, LS09b, LTLH08, LMG08, MHvIR00, MPM02, MSM^{+}07, MEJ06, MHP^{+}06, MY02b, MKiIK07, MAK^{+}04, NP06, NHP^{+}06, Oga09b, OIP^{+}08, OW06, PKC06, PBA05, PDL06, PL02, PBPZ04, PARH07, PS09, PSVH09, PHW08, PHH00, Ray08, RNC08, RT08, RPVR03, Roe00, RVRL05, RPC^{+}05, SSK^{+}07, SPB^{+}06, SM07a, SM09, SFM08, SPK^{+}09, SU08, SGY01, SHCS04, SBMJ09, Sri04, SS00, SYCU09]. study [TMJ04, TYS^{+}09, Tol06, VS08a, WR08, WGS^{+}02, WRB08, WM06, WYMS07, Wil05, WH07, XLZ^{+}04, YWHW02, YZC^{+}07, Yos06a, YYR^{+}07, YL07, ZLO02, ZJTB03, ZBL07, ZZ08, ZYY09a, ZYY09b, ZCY09]. Studying [BCPM09, LBS^{+}06, PLP^{+}04, Ale07, EB07b, JNM^{+}06, LCH^{+}00]. sturgeon [HP09, SJM03, WX09]. Sub [CDA08]. Sub-area [CDA08]. subalpine [Hör03, RSB09]. subarctic [FYN^{+}07, TK08]. subcapitata [FCP^{+}07]. subcritical [dlS07]. subdomains [LC07a]. Subject [Ano01-54, Ano01-44, Ano01-45, Ano01-46, Ano01-47, Ano01-48, Ano01-49, Ano01-50, Ano01-51, Ano01-52, Ano01-53, Ano02-64, Ano02-63, Ano02-53, Ano02-54, Ano02-55, Ano02-56, Ano02-57, Ano02-58, Ano02-59, Ano02-60, Ano02-61, Ano02-62, Ano03-78, Ano03-67, Ano03-68, Ano03-69, Ano03-70, Ano03-71, Ano03-77, Ano03-72, Ano03-73, Ano03-74, Ano03-75, Ano03-76, Ano04-91, Ano04-81, Ano04-82, Ano04-83, Ano04-84, Ano04-85, Ano04-86, Ano04-87, Ano04-88, Ano04-89, Ano04-90, Ano05-63, Ano05-64, Ano05-65, Ano05-57, Ano05-58, Ano05-59, Ano05-60, Ano05-61, Ano05-62, Ano06-49, Ano06-50, Ano06-51, Ano06-52, Ano06-53, Ano06-54, Ano06-55, BGMP06]. Subjective [RB06]. Sublethal [KBvVK08, BPC07]. submerged [vNSvdBC03]. submersed [GML05, HS03, HS06, WWL^{+}05]. submodel [NJF^{+}08, PCS01, SC01]. subornata [CTH^{+}00]. subpopulations [EN06, KD05]. subroutine [CB07b]. subset [ICGÁ05]. subsidy [LPM^{+}07]. substances [JLH01, MFB^{+}05]. substantial [CKN^{+}01]. substrate [ALAS09, TFM01]. substrates [WR03]. subsurface [GBEB06, MLC05]. subtidal [Oke04b]. subtropical [FMC^{+}08, GCM^{+}05, TB06b, ZCG^{+}08]. success [BKMB08, DDLD07, KW02a, LS01, ÖTÖR06]. successes [Xia04]. successful [SB00, VG08]. Succession [KLL01, APG^{+}03, BJH01, BP02, BP04, DKL^{+}09, FWS^{+}05, GAB^{+}09, GB01, HHL^{+}02, JRBS02, KCD^{+}04, KCBS00, LAD06, MKB00, MP07, MS07, MDB^{+}02, NMC^{+}06, Nak08, RF09, Roe00, RAM^{+}03, SAK00, SBM04, SL01, SGH04]. successional [AEP^{+}04, DLP09b]. successions [BAP^{+}03, KK08, LL00, Lud09]. Succisa [MMHE06]. succulent [RVWH06, RHH05, HRH^{+}05]. Sudden [GKG05, dSPdBB08a, MTD^{+}09]. Suez [AAA00]. sufficiently [MCSG06]. sugar [Lar02, LB01]. sugarcane [LB01]. suitability [BDP^{+}02, BBGH08, DDJ^{+}01, DGM08, GBN^{+}06, HHM01, HG02, HLH^{+}06, HC06, HLS06, LTM^{+}04, MBW09a, MSB08, OLB04, RB06, RPRV09, RSM05, SCAP05, SHG^{+}08, SJ03, VADV06, VCRD06, WX09]. suitable [GBB^{+}09, IKS09, NSO^{+}08, PSVH09, SA07a]. Sulawesi [OIP^{+}08]. sulfur [CPJ06]. sulphur [AMSM06]. sum [MOP^{+}06]. Summary [RHM^{+}05]. summer [AP02, VBFM^{+}08]. summering [dFOV07]. Sundarban [Ray08]. Sunderland [Log02]. super [PE08]. super-individual [PE08]. superiority [iTI02]. Supervised [DDT07, BGL01]. supplementation [FHE06]. supply [CCLS06, FL07, KM00, MPRJ04, MY05, SOK03, SB02, YH04]. Support [GKG05, KTL^{+}05, MAB01, RIGJM06, FG08, HN09, KL01, KFN^{+}08, New09, PSC^{+}01]. supporting [AVP08]. suppression [BFE07]. Surface [WGVB01, YWL^{+}05, YTLF08, ZGF^{+}05, BCCR^{+}02, BB07a, ECP08, FRB^{+}05, GBEB06, GdBD08, HBM04, Han02, HSV07, KWS^{+}07, KŞE04, Lin01, MLA^{+}02, MSTK08, PL04, PBSOMG^{+}05, Phi02, RKH^{+}07, Roe01, VMG05, WCHM02, Wan05]. surfaces [Che06, CGD04, Tan02]. surplus [Jen05]. surprises [QXW02]. surrogate [Sto06]. surrounding [LAPAMD07]. survey [ECZ^{+}06, WLG07, WSP08]. surveying [BCMR03]. survivability [Sil07]. Survival [LW09, AMSW07, AHKB01, DD02, GLM02, HP06, JJWF07, JZC^{+}07, MB08, SAH03, TN01, WZC^{+}05, WGT05, WT04b, ZJG^{+}06, ZZ08, ZBWR06]. survival-analysis-based [MB08]. survive [PFBBJ08]. Sus [HAS07]. susceptibility [Dun08, GB08, LYC08, MVMA08, MMV08]. susceptible [TOS09]. Suspended [HMPF05, HE05, AFT09, BA05, DWH06, GCG^{+}07, HPH00, HGB04, JCB^{+}02, KGJ08, MH03, SP01]. suspension [SM04a, ZDR03]. Sustainability [Ano09i, DDFP07, SBC^{+}09, BPP09, BCS09, DFF07, Fis09, GB03, LPCZ07, MMPT06, PMH00, Paw00, PMD^{+}09, PCP07, PHH00, Pra05b, RHH05, STK00, SS00, WDW00, WKZ03, XZD^{+}06, Zha06, Jør04f]. Sustainable [Jør04d, ABC03, Aub04b, Bor07, CHL08, Foh05, GBSP08, HDH00, HvI06, Kir01, LL02b, Sch00, TRDM06, TCD02, ZLZM02]. SVAT [MLL^{+}05, OIP^{+}08, PWC^{+}09, VA00]. SVAT-crop [MLL^{+}05]. Svedin [Jør04f]. Sven [Cha09]. Svirezhev [Ano08-27, Jør08b]. SW [BRGS09]. SW-Germany [BRGS09]. swamp [SSCS06]. swamps [Gof04, Gof04]. swan [JNSS02, RH04]. Swannack [Jør09b]. SWAP [ETH^{+}04]. sward [LMH05]. swards [SL02]. SWAT [GBG^{+}03, PHBF07, RVRL05, SA07b]. Sweden [EPTB07, KGNK03, YL07]. Swedish [Jør04f, Lin01, SJG^{+}08]. swimbladders [SJH05]. swimming [UZSM05]. Swiss [BRP^{+}06, RSB09]. switching [AVTP05, MB09, TH08, VATP05]. Switzerland [Ano06-46, MSP^{+}08]. sylvatica [Kno03, TBPF08]. sylvestris [DADGA06, GBN^{+}06, WRB08]. symbiosis [Ano03-27, JBMBP02, JBMBP03, PMH00]. symbolic [Dže01]. symbols [Ano04-28]. Symmetric [CAB08]. symmetries [ZVLP05]. Symmetry [RL05a, MKvdW^{+}09, RB09]. sympatric [MNOS08, RMS08]. symptoms [PM06a]. synchronizing [Har08]. synchronous [CLHB^{+}08]. Synchrony [Mat01, PAF06, GLL^{+}07, zLGH^{+}05]. synergetic [KFR06b]. synoptic [HGR08]. Synthesis [LHC09, WIMK07, BSR04, Hua03, Mau07, Ric00, ZYY09b]. synthetic [YZC^{+}07]. System [ADSO08, Jør04h, KTL^{+}05, Kir06a, Kir06b, RIGJM06, WFHP07, AW00, AP00, ABR05, AB05b, ADS^{+}07, BFU^{+}09, BBM06, BSB^{+}09, BH02, BPBL00, BPM01, BS04, BDG01, CYHK04, CMR09, CHL08, CSG02, CP02, CJY^{+}09, COB^{+}06, CPB^{+}08, CGG^{+}05, DR08, Dew01, GGM08, GORJ03, Gri08, GG05, HCJ^{+}09, HB00a, HLR09, HKL07, IMS07, IVP08a, IVP08b, JWLA00, JL07, JNSS02, KT03, KVPA08, KMR^{+}07, KL01, KFN^{+}08, KE07b, KLPP07, LKKL09, LSHG08, LL02b, LFB07, LSAGF05, Mac00, Met01, Met02, MB05a, MLGV00a, MLGV00b, MMT^{+}07, MPS02, MMPN07, MMN08, NG07, OZL07, PKC06, PBC02, PLL04, PCWP06, PLA06, PSC^{+}01, PARH07, Pen09, PEE09, PBM^{+}05, Rai08, RJDF08, RS01, Sar04, SOK03, SLE07, SvdWB^{+}06, SHM07, TKSP09, TAP07, Tsc04, Van08b, VP01, VATP05, WKL^{+}03, Weg00, WLSE06]. system [WFT05, WH09, XLD01, YLSH03, YSM^{+}06, Yos03, ZZZ06, ZBL07]. System-analytical [Kir06a, Kir06b]. system-dynamic [HCJ^{+}09]. systematic [CGW^{+}06]. systematically [RB02]. systemic [Fis09]. Systems [Abe04, Bro09, CQ07, MC04, MSC09, RdQFO07, Tie05, AMD^{+}03, AMM04, AMB07, Ale08, Ano04-28, Aok08, AN00, Aum07, BPP09, BSB00, BMR06, BU04, Bro04b, CdQO06, Chu08, CCGMJ07, CSM^{+}06, CTG03, CKBH00, CV01, Fat04a, FAA^{+}02, FJ05, FH07b, GKT07, GRHS00, GPP02, GJ07, HvI01, HM01, Hua03, Kan04b, KK00b, LPCZ07, LR03, LMM^{+}07, LWC^{+}07, LCY09, LWJ06, hMzL08, MLGC03, MPC06, MD07, MGSdG07, MSW^{+}09, MTKM^{+}06, Nie00b, Odu02, OS02, PBA05, PGCK04, PS05, Pet04b, Phi02, PdMVN09, PR07, PKG00, RDS07, Rai01, RJFAA07, RKH05, RHH05, RG04, RL09, SMSR00, SMSV09, SDDC07, SKP09, SMTR07, SvL04, TLW01, TP08, TB06b, TRDM06, TKSP09, VFTB06, WB00, WM02, WD02, XJM07, ZK08, ZYY09a, dG02a, dMPVNO07, vdBDR02, Jor05g].

T [Ano04-46, Bro04a, BU04, Cam04, Com04, Jor05j, Kan04b, Lim04, Lug04, MC04, MD04, SH04, Til04, Ulg04, Zuc04]. T. [HC03a]. tabaci [BBT06]. Table [Ano04-92, Ano04-93, Ano05-66, Ano05-67, Ano05-68, WSS^{+}06]. tag [Xia00b]. Taihu [HJZ06, CM03]. tail [PLLdCB06]. tailed [CK07b, HBUS02, Jen00, LKY^{+}06, SLL^{+}06, XLD01, YFH03]. Taipei [LHC09]. Taiwan [CHL08, LLCL04, LTLH08, TYLL07]. taking [MSB07]. tamarisk [WLZ^{+}09]. Tamarix [WLZ^{+}09]. Tamaulipan [MGH^{+}05]. tandem [MB00]. Tanganyika [BLC^{+}07, NDD^{+}07]. tapering [SLE07]. Tapes [PSC^{+}01, SCR03, VCRD06]. tarandus [FCH04]. Tarantola [Jor05a]. target [CGdPRY09, Eza05, KDW^{+}09a, KWW^{+}09, LNW00, PGM08, TSJ08]. target-oriented [LNW00]. targets [HC06]. Tasmania [MB02b, RRB^{+}01]. Tasmaphena [RRB^{+}01]. taxa [ICGÁ05]. taxa-range [ICGÁ05]. taxifolia [CTH^{+}00]. taxonomic [AAU02, CSSCC07]. Teaching [Nie09a]. teak [TYT^{+}09]. technical [HvI01]. technique [BH01, DFM07, JJWS08, KST^{+}07, KK04, LGL02, OEK^{+}06, VHG05]. Techniques [Ber02, CM03, GAA^{+}05, GZY^{+}06, LdVA06, MLHT09, MF02b, PDL06, PBSOMG^{+}05, PHD04, RPVR03, ZGH05, ZGL08]. Technological [Jør04h, SPJB06]. temperate [BW01a, BB08b, BEF03, GAA^{+}05, GZY^{+}06, IOIA08, PSPD09, RFA^{+}01, RMDC04, SSPL^{+}08, THJ^{+}03, TSF^{+}05, VSG^{+}08, YABM07, ZBWR06]. Temperature [LWJ06, QXW02, AY07, Ale07, AG03, Ash06, BMD09, CM05a, CSKP08, CZG05, GFGZ08, GAA^{+}05, GG00, HvG07, KŞE04, KBE^{+}06, LCM^{+}09b, MOP^{+}05, MOP^{+}06, MLA^{+}02, MMA02, Nad07, ORS^{+}09, OFK08, PL04, PBC02, PHW08, RV05, RWM^{+}07, TR03, TYLL07, TVK^{+}08, VVF06, VSFM03, WL04, WHP01, Xia00c, ZK08, dSSGR00]. temperature-dependency [Xia00c]. Temperature-dependent [LWJ06]. temperatures [CMB05, ME07, PLA06, WSS^{+}06]. Temporal [BWPC06, JKJ^{+}08, KS04a, LS06, OWHS09, TK01, AKMB01, AAO06, BKGC05, CBMP07, CKA^{+}02, CPP00, CZG05, DC08b, DBB^{+}08, EBM06, GML09, JO09, KS07, Las06b, LAB^{+}05, LL06, LZB^{+}06, LLZ^{+}08, MEJ06, MM06c, MFB^{+}05, MDB^{+}02, Met01, Met02, MBKD02, NP06, PDL06, PL04, PKC^{+}01, dSPdBB08a, PDS07, RMBM06, RJS^{+}06, SCB^{+}09, SDS^{+}07, SMB^{+}06, SRN05, SB07, SZL^{+}04, VMR09, WGS^{+}02, WHZ06, YHG04, Yos06a, ZCQ^{+}09, ZBL03, GAPE06]. Temporal-spatial [OWHS09]. temporally [Ann01, OG08, RHB04, SSNHP08]. temporary [CUS01]. tendency [Jua09]. term [CUS01, CWCH09, CKP^{+}01, CKPP03, CML09, GvNK09, HRH^{+}05, IWM^{+}06, KPK^{+}07, KW00, KFH07, LGR^{+}09, LH09, Loe04, LGD01, MGH^{+}05, MBD^{+}00, NW06, OHH07, PDL06, PGFM04, RFA^{+}01, RSB09, RHP07, SSH08, Swa06, TC09, TMD04, WFM01, WKZ03, XBM^{+}07, Yea04, YTCV01, YCVA01]. termite [LBS08a]. terms [FM08, MJ02, Mit09, Xia05a]. Tern [MPRJ04]. terraces [DS01]. terrain [AKMB01, WCC02, WTS^{+}06]. Terrasnyi [KK08]. terrestrial [AWL04, AOY02, CCC00, GS08, HTK07, HCL00, KC01, KVL^{+}09, LBL^{+}08, LSS^{+}00, PK02, Roe01, RKS^{+}07, SOHY07, SYC04, TR05, TKHS^{+}07, TP06, WIH^{+}09, WLSE06, YFLW06, YTLF08]. terrestris [PBMRE08]. territorial [MML02]. territoriality [PBK03]. territories [SHSP04]. territory [MML00]. Tesca [LT01]. tessellation [LLA^{+}09a]. Test [AYKY05, Weg00, YJJG09, BS08b, CCJ07, DMRP07, HAS07, MKvdW^{+}09, SLPP05, VA00, dBWG05]. testaceipes [dSMZ09]. tested [MWD05]. Testing [FGEG06, MSA^{+}03, MZWM05, MGSdG07, PLTT05, PBPZ04, PEPB09, RvGC^{+}08, SSPL^{+}08, SDS01, TMHJ06, FBDM09, IWM^{+}06, ÖTÖ06, PA09a, ZCG^{+}08]. tests [RF04, VPSG05]. Tetrao [KK00a]. Tetraopes [Mat03a]. tetraophthalmus [Mat03a]. Tetrax [SSdlMP^{+}08]. Texas [HGD05, Kan04a, LGS^{+}00, SH04, TCGL03]. text [Svi04]. Tha [MI01]. Tha-chin [MI01]. Thailand [BPW^{+}03, MI01, MCJ^{+}04, PSVH09, TMJ^{+}07]. Thalassiosira [BLVC03]. Thau [PCM^{+}03, TPCS06, VML^{+}06]. thaw [ASN02, NJF^{+}08]. Their [KPKP06, NKK^{+}09, AALM05, AF05, BVD05, BCL^{+}09, BT01, BPC04, BBGH08, CP04, CBD^{+}09, Del04, HvRIZ08, Hul04, OMA01, PR01, SM04a, SHG^{+}08, SV03b, VPV09, Xia04, vNM02]. thematic [JFG^{+}08]. theorem [SPK^{+}09]. theorems [Lev00]. Theoretic [Jør04e]. Theoretical [Dew01, dlPP05, AJJ^{+}05, BB04, Gas05, GMPC08, GLM02, JG05, Kri04a, LBBC08, MJ02, MHKW00, Ric00, RBB05]. theoretically [Kir06a]. theories [MNPJ03]. Theory [GORV06, HC05a, Jør08a, Jør09b, Nie06, Suh05, Suh06, Tie05, ZJC^{+}07, Ale08, AGZ05, Aus02, ABM^{+}06, Aus07, BPP09, BMR06, DPT09, GJY07, GPP02, Hul04, IO02, Jor00a, Jør02b, Jør07c, KYL07, LiWZ00, MGL04, MP00, MCE^{+}07, OM02, Rey02, SG09, SAB03, SSPR03, VS05, Wir00, WD02, Xia00a, Xia03, ZZLC06]. Thermaikos [MAG01, NKK^{+}09]. thermal [FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, GIZ^{+}03, GFGZ08, HAG^{+}00, Pot04, TH06a]. thermal/dissolved [FSE^{+}04a, FSE^{+}04b, FSE^{+}04c]. thermally [CPH00, MBDB09]. thermocline [CM05a, Nad07]. Thermodynamic [Tie05, Ale08, LCSZ09, LP03a, LP03b, LJ09, Lud09, MRRJ07, MJ02, RC08, ZW02]. thermodynamical [NU00]. Thermodynamics [DPT09, Nie00b, Svi00b, CG09, TK07]. these [FSJ03]. theta [BS08a]. theta-logistic [BS08a]. think [RR07]. Thinking [MD04, Kan04b]. thinning [Kar03, LiWZ00, TYS^{+}09, WSWL08]. Third [Ano03z, Ano08a, LBL^{+}08, Ano09-33]. Thomas [Phi04]. thornshrublands [TCGL03]. Thou [Cam04]. thousand [Bro04b]. threat [NW07]. threatened [BP07, JDD^{+}03, RHP07]. Three [MJ02, MBP^{+}08, RH04, ROQ^{+}09, SWSF06, WDBK08, AR02, BHC04b, Bio01, Bio03, BDI04, BBGH07, CMR09, CS03, DCI01, FG08, GFG09, HSJ04, HFSH06, HJZ06, JBT^{+}05, KBE^{+}06, LR03, MD06a, MR08, MS00, ORS^{+}09, PM06a, PdMVN09, RŽČ^{+}04, RG06, RAI04, SI06, SAR^{+}09, SM06, UZSM05, VPV09, WZKL09, WM06, WR03, lXzL02a, Zha03a, ZWXF05, dMPVNO07, vdBDR02, vdHGF09]. Three-dimensional [RH04, ROQ^{+}09, Bio01, Bio03, BDI04, DCI01, HJZ06, MS00, ORS^{+}09, RŽČ^{+}04, RAI04, UZSM05, Zha03a, ZWXF05]. three-species [lXzL02a]. three-trophic [PdMVN09, dMPVNO07]. Threshold [Aub04b, BBB03, CBP07, FM08, IiTS^{+}04, Roe00, ZS09]. thresholds [HMG06, QKR03, RHHM08, SJ05]. thrips [BBT06, BBT06]. throughfall [BPC04, KS04a]. Throughflow [MKS^{+}09, BBP^{+}07]. throughflow- [BBP^{+}07]. Thun [MSP^{+}08]. Thunnus [LG04, NCM07, Wil08]. Thyle [RVRL05]. Thymallus [MSP^{+}08]. thynnus [LG04]. Tibet [LM07b]. Tibetan [CZL05, XZS^{+}07, YLLW02]. tick [CTG03, CTG04, EPQdCA06]. ticks [TCGL03]. tidal [Buz08, CUF^{+}09, MBCS07, SFU01, SSKN08, WHHH07]. tide [BVD05, KE07a]. tide-driven [BVD05]. Tiefan [Jør06a]. Tien [LMG08]. tiered [CBSLS07]. Tiezzi [Bro09, Jør04g]. tiger [ASJD01, IKS09]. tiger/human [ASJD01]. TIGMOD [ASJD01]. till [FCC^{+}00]. tillage [FCC^{+}00, GSBM08, GARB09, LL07a, SC01]. tillering [LMH05, MZWM05]. timber [DSA08, Gan06, GGPBEK07, RRB^{+}01, RCZ^{+}06]. Time [BDBS08, BWPC06, Den08, Ned09, SGP^{+}07, TAL^{+}05, WBP^{+}07, AMD^{+}03, BW01a, Ban05, Bra01b, BKMB08, CPB^{+}08, CBSLS07, CMCD04, CG06a, CG07, CU06, CUF^{+}09, DFHP04, DDLD07, Epp00, FMM^{+}07, FBC02, GLHV08, GG08a, GL08, HTMO06, JZC^{+}07, Jon07, KRZ07, KŞE04, LHC07, Leb05, LSAGF05, LCR06, MCBA07, Mon02, MSB07, ML05, NS08, NDM00, OVK^{+}06, OW02, RPPB04, Rob05, RMOA06, SB06c, SHZ05, SB00, SDS01, TNO^{+}09b, TW00, Van08a, WR01a, Xia02, Yos06a, ZSD^{+}08, vWSH08, Jør04g, PAF06, Jør04g]. time-course [GL08]. time-dependent [LSAGF05, Xia02]. time-scale [Rob05]. Time-scaling [TAL^{+}05]. time-series [AMD^{+}03, FBC02, HTMO06]. Time-varying [BDBS08, MCBA07]. timed [SAL07]. timely [LH09]. timescale [Sem08]. timing [CGR03, LAXP07]. tiny [HB09a]. tip [ZvBS05, Den09]. tissue [DVdB^{+}08]. Title [Ano02-52, Ano03-59, Ano03-64, Ano03-60, Ano03-61, Ano03-62, Ano03-63, Ano03-65, Ano03-79, Ano03-80, Ano04-77, Ano04-70, Ano04-71, Ano04-72, Ano04-73, Ano04-74, Ano04-75, Ano05-46, Ano05-47, Ano05-48, Ano05-49, Ano05-50, Ano05-51, Ano05-52, Ano05-53, Ano05-54, Ano06-39, Ano06-40, Ano06-44, Ano06-41, Ano06-42, Ano06-43]. Tivela [dSSGR00]. TM [Met03]. tobit [ES06]. Todd [Jør09b]. tolerance [Zav04]. tolerant [CBP07]. Tolo [XLZ^{+}04]. tomato [KBGJS06]. tool [BDD^{+}01, Cal06, CKN^{+}01, DMF07, EDKF06, FG08, FDCH08, GHP^{+}09, GBEB06, GBSP08, Håk09, HR03, HTK07, JDD^{+}03, LMG08, MP04a, MU02, MWMN06, NI08, PLL04, PCS03, PDV^{+}07, POH02, RST05, SSPL^{+}08, vNM02]. Toolbox [MCJ^{+}04]. Tools [LMPR06, BMR06, DKRŠS00, DVGD07, JLBS09, War08]. toothed [VF07]. top [Bar00, HY07b, Her04a, KGSB01, MMRLP06, vLLv^{+}07]. top-down [Her04a, KGSB01, MMRLP06]. Topics [Nie08a]. topographic [BHW^{+}08, MSA^{+}03, WHHH07, YBM^{+}05]. topography [Rob03, SBDD04b]. Topological [QHM05, VJ06, FBC08, SM09]. topology [JO09, SBB09, Sil07]. toposequence [TFTO07]. Torr. [dJTMBGMP^{+}09]. tortoise [BWAM09]. total [AP02, RSF^{+}01, WTS^{+}06]. touchstone [RvGC^{+}08]. tourism [LPCZ07, PFR09, PGCK04]. tourism-based [LPCZ07]. tournefortii [SF07]. tower [SOHY07, XBM^{+}07, XBM^{+}08a, XBM^{+}08b]. tower-based [XBM^{+}07, XBM^{+}08a, XBM^{+}08b]. town [SL06]. Toxic [ZRR^{+}08, BPC07, KKCC06, KBvVK08, ZTXD03]. toxicant [KNB08a]. toxicity [KC03, SM06, VH07]. toxin [CRC08]. tracers [KE07a]. tracing [SIS^{+}07]. track [Dei04]. Tracking [NG09, PSCMMNS06, CGdPRY09, CDM05, MNZC08, SHG^{+}08, TSKP09, WPMT07]. Trade [ABR05, RVWH06, Sav00, Sil07, vN02]. Trade-offs [ABR05, RVWH06, Sav00, Sil07, vN02]. tradeoffs [CM07b, GL09, KvKV05]. trading [LPM^{+}09]. Traditional [SHN09, DFGC04, SS08a, SAS06]. trained [WL06]. training [AMD^{+}03, DS02, ÖTÖ06, WFT05]. trait [Hua08, LK07, ML06, MBW09b, WT04a]. trait-based [MBW09b]. traits [BCH09, CM07b, DWD07, Fie04, SCAP05, vW07]. trajectories [SAL07, UZSM05]. transect [EG09]. transfer [DDS^{+}04, GFGZ08, LH09, MMPN07, PWC^{+}09, SBB09, SS08a, vdPLG^{+}00]. Transferability [BRV09, AAO06, MR06]. transfers [DBBS03]. transform [BHFMG05]. transformation [LLH^{+}06, LdVA06, PWSY07, fPzWhSY07, QMH00, SMM^{+}02, TTPS09, VCC03]. transformations [JMVvDV02]. transformity [BCST05, BCRP09, BOJ04, BU04]. transforms [HD00a]. transgenic [GL06, KS07]. Transient [KGZR05, Yea04, DWR09, FMRS09, NKC04, XG09]. transients [Wil01]. transition [EB06, GS06, JLBS09, PBHGF07, SHZ05]. transitions [KVV01, YP02]. translate [VBD06]. Translating [SKJvdHG06]. translocation [ARF08, LL07c, Oga09a, RHP07]. transmission [HTMO06, LDM00, Mog02, MMT^{+}07, MOJ01, SDD^{+}04, USK^{+}06]. Transpiration [NV03, COB^{+}06, KBGJS06, Szi00]. transport [BC01, CWH^{+}00, CW01, CMG00, CM08, CUF^{+}09, DWR09, DCI01, EFJ^{+}08, GPSM08, GSZ08, HHK07, HHB^{+}08, JWW^{+}07, JF00, KGNK03, LLH^{+}06, OCK01, RBB05, SPK^{+}09, VPSG05, WLB^{+}04, WTMG09, XIX^{+}08, ZTXD03]. trap [BJFM06, LG04]. trappability [OMA01]. Trasimeno [LMPT05]. travel [FAB^{+}07, SHZ05]. traveling [Sil07]. travelling [PAF06]. treated [Che06]. treatment [DGOZ04, SL06, TAP07, WBR07, YLSH03]. treatments [KRZ07]. tree [AF09, BB04, BKB00, BLDCM06, Bra01a, Cai05, COB^{+}06, CSF^{+}04, CWS09, DGD03, ECZ^{+}06, Esc05, FBDM09, GBS00, GRBT08, GC02, GL08, GICB09, GB01, Kai00, Kar03, KBK07, Kin05, KH07, LPD08, LC07b, ML06, MBM00, MOP^{+}05, MOP^{+}06, Mii00, MFB^{+}06, MSB08, MSL06, NDM00, NI08, OHH07, PS09, PS05, PBY^{+}03, PTS^{+}04, PMLM08, RFA^{+}01, RGL^{+}07, Rob03, SKP^{+}07, SBM04, SFC05, SCPC^{+}07, SBR^{+}07, SLGS00, SGH^{+}08, TC09, TMM05, WTST08, WZKL09, WBR08, WBB05, WRC^{+}08, WGT05, WSZS08, WBR^{+}06, YTH03, ZdlP05, ZVK05, ZBSA07]. tree-based [FBDM09, MFB^{+}06]. tree-ring [BB04, Mii00]. TREED [FLS09]. Treegrass [SLGS00]. treeline [Cai05, GFGZ08, HLR06]. TreeMig [LZB^{+}06]. Trees [LE04, BH00, BAP^{+}06, CM04, DSA08, DDT07, DDJ^{+}01, DVdB^{+}08, DGD06, DD03, GPK00, Kno03, KDW^{+}09a, LA04, LAZ^{+}08, MAA^{+}09, MF02a, SW01, STH06, TYS^{+}09, WWRZ04, WNW09, ZBWR06, vNM02]. trends [Loe04, MGH^{+}05, SGF09, Swa06, VSFM03, XZD^{+}06]. tri [BDG01, HK02]. tri-trophic [BDG01, HK02]. Triangle [AVP08, XZD^{+}06]. Triatoma [CAG03]. Tridacna [HC03a]. tridactyla [BKMB08]. tridactylus [RCL06]. Trieste [CS08]. trigger [MTD^{+}09, SAK00]. Trinidad [MHMHKA04]. trip [RSB02]. TRIPLEX [PLD^{+}02, ZCG^{+}08]. Triticum [MWD05]. tritium [MTVC05]. tritrophic [Cha08, LLA^{+}09b, MdRdA06]. Trophic [AGNLGSG04, Bar04, DG04b, FM07, GE05, GMPC08, JCX^{+}08, JcLM09, MSC09, PNN^{+}08, VOM06, AW00, AYK07, BdB05, BFD02, BFS03, BFS05, BDG01, CSSCC07, CHSB00, CMR09, CS03, CSM^{+}06, CDA08, CGGE08, CMSC01, DUASCM07, DD00, DS01, Gas05, GWHS03, GDP09, HK02, Her04b, JLBS09, KB04, KKW^{+}07, KMN^{+}07, KK00b, LBS08b, MHMHKA04, MRK^{+}07, MRiI^{+}07, MR06, MR08, MZASLMLC04, NAC04, Oke04b, OW02, Ort08, OACB09, PK08, PK02, PdMVN09, ROQ^{+}09, SAK00, SAB^{+}06, SBB09, SM03, TN08, US08, VJ06, VLA^{+}06, VIK^{+}08, WFHP07, WC07, YYZ06, YYR^{+}07, ZRCA08, Zue07, dMPVNO07]. trophic-level [Gas05]. Tropho [GGP03]. Tropho-ethological [GGP03]. trophodynamic [CS08, HMFM00]. Tropical [ASG^{+}05, AR02, BB08b, BHC04b, CA04, DMBO00, DG04b, Eza05, FBF^{+}06, FRM09, FC06, Her03, KP09, KH07, LC07b, LAZ^{+}08, MBP^{+}08, OIP^{+}08, OIR^{+}08, OHH01, PK08, PLL04, PTS^{+}04, PMC08, RSA04, Rob03, dCSB08, SC04, SRA05, TYT^{+}09, TH06b, VLA^{+}06, WRC^{+}08, WSF^{+}00, WSF^{+}02, dG02b, dGdMPC05]. tropics [AN06]. trout [BEM00, BRP^{+}06, CDM^{+}00, TNK04, VH07, WPD04]. truly [VCD05]. trutta [BRP^{+}06]. tryoni [YZS^{+}04]. Tschirhart [CEP06]. tsetse [OMA01]. Tuamotu [MFG^{+}06, MFG^{+}08]. tube [WGV^{+}08]. tube-building [WGV^{+}08]. tuberculosis [CG06c, SRS^{+}03]. tuna [DLK01, LG04, NCM07, VREA06, Wil08, YMD08]. Tung [Nie08a]. tunnel [CCG07]. tunnels [LBS08a]. turbidimetry [CGH^{+}05]. turbines [DCP^{+}07]. turbulence [GVDF07, LGC07]. turbulent [CF09]. turkey [LGS^{+}00, ZJTB03]. Turkish [TB05]. Turnover [LK07, BCG08, HHB^{+}08, KFR07, POH02, VCP09]. turtle [Cha02, CB07a, MBM06, Sno08]. turtles [MM06c]. Twelvemile [RBW09]. twin [LL03a]. twin-scale [LL03a]. Two [GDL06b, Jør07a, LBS08a, Mog02, OLB04, QKR03, TSJ^{+}09, CPV07, CRC08, CMM02, CSdZ09, CWBR01, DH00, FGEG06, FK05, GPK00, GB02, GSC09, HLHZ06, HK02, KC03, Kar04, KVPA08, LAD06, MAdlPR02, MDH^{+}08, PRL07, Pyk04, RJDF08, RAI04, SC01, SMSR00, SMSV09, SvL04, SÅ06, TB08, TR03, VTB^{+}08, VATP05, VLA^{+}06, WFHP07, iWLSN00, XSY^{+}09, ZZZ06, ZCB08, ZLX09]. two-component [HK02]. two-dimensional [FK05, MDH^{+}08, ZZZ06, ZCB08]. two-locus [HLHZ06]. two-objective [CPV07]. two-predators [RJDF08]. two-prey [KVPA08, VATP05]. two-sex [SÅ06]. two-species [PRL07, SvL04]. Two-way [GDL06b, TSJ^{+}09]. type [BPJM00, GARB09, KJ08, KRvL^{+}02, LL07b, LS08a, LSHG08, Mit09, NM09, RJS^{+}06, SIS^{+}07, SFC05, SvdWB^{+}06, TS09, WH09]. type-2 [LS08a, LSHG08]. typed [OFK08]. types [BB08a, BG01a, BG01b, DAC^{+}09, EGE^{+}08, Jør08c, LAD06, New09, SWSF06, WO01b]. typical [SYCU09]. typographus [JPD^{+}06]. typology [FDCH08, SDA^{+}03].

U [Jør04a]. U.S. [Kir06b]. U.S.A. [CWS09, HTS^{+}07]. UK [Jør04g, Jør04a, Jor05k, KGJ08, DSD^{+}08, HXP^{+}09, WYMS07]. Ukraine [vdPLG^{+}00]. Ulanowicz [Bro09]. Ultracoelostoma [JDPI07]. ultrastructure [BLVC03]. ultraviolet [BCD^{+}05]. Ulva [AACIS^{+}08, dGdMPC05]. Umbria [LMPT05]. Unbiased [PBSOMG^{+}05, TN01]. unburned [OMHR06]. uncertain [GRHS00, OEK^{+}06]. Uncertainties [BBT06, SFM08, vHPS02, EBR02, RB06, TYZ^{+}05, WSZS08]. Uncertainty [BFOS08, BKC^{+}07, CW01, CK07b, HvI01, KTL^{+}05, LBBC08, Lek07, PVS^{+}09, PCP07, RFdV07, XG09, APCdlR07, Ann01, AMRS08, BT01, BO07, BSR04, BBB03, CSR08, DBBS03, DDF^{+}05, EB06, FJ05, FBDM09, Håk00, HBDA09, HMF00, IiTS^{+}04, JKS^{+}05, LBL^{+}08, LHC07, LHK00, Lin06, LFB07, MGV00, MC01, MBF09, MRE^{+}06, Mon09b, OBE^{+}07, OBR01, PCB07, PMLM08, Pra09, RA06, RLLB09, SA07a, SK01, SGF09, TIJ^{+}01, TNK04, TF09, WEW01, WGS^{+}02, WRC^{+}08, Wil09, WT04b, XRM08, XHH^{+}04, XHH^{+}05, YTT^{+}09, ZLW09, vNS03]. uncertainty/variability [WT04b]. undergone [VCM07]. underlying [PRRB09, TTHH07]. understand [PBC09, RHHM08, SHM07, YMD08]. Understanding [BMR06, NM09, WTST08, BU04, DPT09, Hos06, OJ02, YY06]. understory [HNF09, JAB^{+}06, MBM00]. underwater [CMM02]. undisturbed [BBD^{+}04]. unexpected [YJJG09]. unharvested [TK01]. unicolor [PSVH09]. uniform [LE04]. unintended [Lin03]. union [HSMN08]. union-find [HSMN08]. unique [KYH00]. Unit [BBC03, BVC^{+}01]. United [KRZ07, KR04, MCSG06, OB04, SHCS04, SWCO07, SS08b]. units [Zei04]. univariate [KŞE04]. universal [WGS^{+}02, Wil08]. Universalis [MC04]. universe [BOJ04]. University [Ano09-33, Jor05e, Jor05l, Nie08a]. unpredictability [Van04]. unreasonable [GJY07]. unsaturated [JWW^{+}07, KNZ04, VPSG05]. unstable [NHLPA06, ZHL07]. unsteady [ZVLP05]. Unsupervised [WO01a, CPP00, LKLK07]. Untangling [LLF02]. unusual [LC04]. unveiled [ABB06]. up-scaling [Wir00]. Update [Ano09g, SMG07]. Updating [VB06, CLHB^{+}08]. upgrowth [PBHGF07]. upland [LIS07, New06]. upper [HLR06, LWLZ06, AA05]. upscaling [BDK01]. upstream [GM05]. uptake [BA05, CC05, MDH^{+}08, NV03, ŠH09b, VL08, vdBDR02]. upwelling [BGF00, CSM^{+}06, DG04b, HB00a, PEE09, dGdMPC05]. Urban [FMdV01, zGsLcXwZ09, SYCU09, SHS08, Chu08, Gan06, HC05a, HKL07, LH01, LLS^{+}08, LYCU09, MEJ06, VLD02, ZYL06, ZYY06, ZYY09a, ZYY09b]. urbanisation [LGS03]. Urbanised [SHSP04]. urbanization [LKLK07]. urbanized [WHX^{+}03]. urchin [MMLR07, MHZ^{+}06]. urogallus [KK00a]. Ursus [FPS03]. US$ [Jør04c]. USA [Ano08a, Ano09-33, AB05b, HGD05, Jør04h, Jør06a, LBL^{+}08, AB05a, BJJ06, BWPC06, BO07, GSB^{+}06b, HP09, MNOS08, MMR06, MMF^{+}09b, MMF^{+}09a, PGLS03, Pyk04, RBW09, SR08, SGP^{+}06, SGP^{+}07, WHH^{+}08, WBP^{+}07, ZCKR07]. USD [Hey01a]. Use [Asp02, BKS05a, ÇDKK05, DGD03, GP07, KKA^{+}08, LP03a, LP03b, MCGO05, PWSS07, YMD08, ADSO08, ACE07, ASJD01, AM02, AMB07, ACJT08, BS08b, BHP05, BBGH07, Cal06, CV07, CKL^{+}06, CMB^{+}02, CBD^{+}09, CWBR01, CKBH00, CA04, dSDSM08, DS01, DCG01, Dun08, EBH^{+}01, ECP08, Esc05, FE04, FKIL08, GGB^{+}06, GS06, GW08, HDB^{+}06, HvI01, HHF05, HGR08, HBCL07, HHKH09, HC05b, Jor01c, JRBS02, KSÇ^{+}00, KWS^{+}07, KW02b, KFR06a, KJB07, KDW^{+}09b, Las06b, LHC09, LCY09, LMG08, LJ09, MHvIR00, MKM^{+}08, Mat03b, MSM^{+}08, MP04b, MLL^{+}05, MDGK01, MB07, MBPS04, Mur01, MG02, NT07, Oga09b, OIP^{+}08, OdKV03, OFK08, PK01, Paw00, PCL^{+}05, Pra08, RSC09, RGL^{+}07, RVRL05, RBEZ08, RMDC04, SB06b, SFM08, SV02, SV03a, SBS^{+}06, SGF08, TMP06, TEMJ06, THA02, TA05, WFM01, WG00]. use [WBvDHZ09, WBTC00, WBN^{+}03, XHC^{+}08, XZS^{+}07, XGL^{+}09, ZSKV05, ZH06, vODFS04, vdBDR02]. use/cover [SB06b]. used [AGM^{+}08, ABM^{+}06, BOB07, GBS00, Jon07, KM04, Oku09, RBSG06, RLBL01, SPTP01, War08]. Useful [PHD04]. useless [Nie08b]. user [GLS02]. Using [AMM04, AS03, Aru05, BS08b, BGMP06, BBP08, BDP^{+}02, CB07a, CM05a, CHHP02, Dei04, DDL^{+}06, DC08b, DD03, GGM06, GL06, GPC^{+}09, GSC09, HA08b, HvG07, IG02, IVC^{+}08, JLBS09, Jor05g, KRZ07, KDW^{+}09a, KHLS07, LHK00, LPCC05, MJR06, Nad07, Par00, RNKG03, SA07b, Sno08, SS08b, SHS08, SHM07, SWBH08, TDLP03, VBD06, WKL^{+}03, WDW00, WBS^{+}02, WWC^{+}07, AN06, AMSM06, ASG^{+}05, APHV08, AKMB01, AMSH08, AQS^{+}07, AP06, ADS^{+}07, ATD^{+}06, BCCR^{+}02, BPBF^{+}00, BB08a, BMB07, BGF00, BCM^{+}08, BPC04, BHFMG05, BLC^{+}07, BB04, BDI04, BJJ06, BRP^{+}06, BFOS08, BKC^{+}07, BCLR04, BCG08, CSSCC07, CK07a, CGC01, CLTH08, CPP00, CKP^{+}01, CEK08, CWL05, CG06a, CG07, CC05, DSA08, DDLB02, DDT07, Del04, DKL^{+}09, DGM08, DAR^{+}07, DFCM01, DDH^{+}09, ET04, FM07, FJ05, FCH04]. using [FCKH06, GBB^{+}09, GFGZ08, GWK^{+}06, GFG09, GIKS08, GNA^{+}06, Gra04, GBG^{+}03, GB08, GH07, GLS07, Håk09, HKB02, HD00a, HTS^{+}07, HXP^{+}09, HY07a, HTS00, HTA^{+}08, Hin09, HAS07, HB09b, Hör03, IVP08a, IVP08b, IGP^{+}03, IKS09, IMK^{+}07, JLAM09, JMN02, KPKP06, KMR^{+}07, KYL07, KFB09, KC05, KDK06, KMN^{+}07, KJ08, KH02, LN08, LR07a, LAD06, Las06a, LKLK07, LPP^{+}07, LAPAMD07, LELR02, LWLZ06, Lid01, LK03, LB01, LZ07, LAZ^{+}08, LLS^{+}08, LdVA06, LCH^{+}00, LMPT05, LLL^{+}07, MDGV09, MSL01, MVMA08, MMV08, MLH05, MKvdW^{+}09, MT02, MXC^{+}04, MCSG06, MRiI^{+}07, MFJM04, Met01, Met02, Met03, MLF^{+}06, MBD^{+}00, MF02a, MD06b, MLL^{+}05, MRE^{+}06, MLM06, MKiIK07, Mul07, NDD^{+}07, NDM00, NJF^{+}08, OLK^{+}04, OAL^{+}07, OJD04, OW06, OLB04, OFK08, PDL06, PCWP06, PLS^{+}06, PRL07, PT02, PF00b, PF00a, PMD^{+}09]. using [PWC^{+}09, PMC08, PP04, PDL^{+}08, PSVH09, PWH07, PS01, PFFM07, RHB06, RSF^{+}01, RSA04, RM09, RKH05, RNC08, RMA08, RLFB04, RPVR03, RF04, RKH^{+}07, RCL06, RHB04, RJS^{+}06, SSS^{+}09, SRW05, SIK07, SJ06, SV02, SL06, SB00, SKJvdHG06, SKCM07, SLZ09, SMET06, SGLH04, SGH04, SRR06, SBC07, TN09, TB06a, TYT^{+}09, TBP^{+}07, TYZC05, TK07, TSF^{+}05, TMM05, TSJ08, TN08, TKTB07, VMG05, WPBM06, WLG07, WZKL09, WGS^{+}02, WL06, Wan07, WYT09, Weg00, WM06, Whi00, WDR06, WSP08, WZC^{+}05, WA02, WT04b, WHZ06, YSM00, YSM^{+}06, YSG^{+}06, YSB08, ZVLP05, ZLO02, ZGF^{+}05, ZAM^{+}07, ZCQ^{+}09, vBBE^{+}08, MDH^{+}09]. Utah [MFB^{+}06, SRB06]. Utilisation [BGL01, KRvL^{+}02]. utility [Gof04, HO01]. utilization [CQ07, PCS01]. Utilizing [RLR05]. Utricularia [dCSB08]. Uzbekistan [RSM05].

V [Jor06b, LGR^{+}09, MC04, Ano04-94]. V.2 [MGCK^{+}03]. vaccination [BSTS^{+}02, CG06c, SKT00, SC02]. vadis [BP05]. vague [EBR02]. valid [SB07]. validate [OLB04]. Validating [OW08, HLR06]. Validation [DC08a, DDG^{+}05, DDH^{+}09, Jor05j, RPC^{+}05, SDL08, Asp02, EIRT00, IO02, JMVvDV02, JZC^{+}07, JKPL09, LHT^{+}08, LGR^{+}09, MZWM05, MLF^{+}06, NW06, NH07, PHD04, PHWH^{+}09, RE03, RF04, SM04b, SS08a, SWK^{+}05, SSKN08, WDBK08, WSF^{+}02, ZPD^{+}08]. validity [Dor07]. Valley [ZBW05]. Vallisneria [BB08b]. value [DSA08, For02, For03, GRR04, LC08, RCL06, WAB^{+}07]. values [BJK09, BEF03, BBB03, GSBN03, Håk09, IC02, LKP03, LAZ^{+}08, MK04, SDL08]. Vänern [DWH06]. Vantaanjoki [GBG^{+}03]. vapor [BW01a, KBE^{+}06]. VaR [WAB^{+}07]. var. [RAH07]. variabilis [LvGC^{+}04]. Variability [ASN02, BVD05, CEMS05, CPH00, ČKBB06, DC08b, GH00, GRPF07, HLK06, IMK^{+}07, KBB00, LZZA09, LG04, LC07b, LWBW05, MM06c, MC01, NHP^{+}06, RVWH06, RWW07, RC06, SI07, SCB^{+}09, SL04, TTJ^{+}09, WT04b, ZPK^{+}07]. Variable [NV04, AS03, Eza05, GBA08, Her08, JGL07, KE07a, KNB08b, LE04, MB07, OJ02, OJD04, TLW01, TW02, TMJ04, WBR07]. variables [AGM^{+}08, AKMB01, BB03a, CSSCC07, DGD03, dSDSM08, GDL03, GDL06b, IKS09, JAN^{+}03, LKP03, LRT^{+}08, LLL^{+}07, MCBA07, MFG^{+}06, PTGI09, SO06]. variance [MC08]. Variants [McK01, Xia07]. variation [AYK07, BWAM09, BA08, BWPC06, BPE^{+}07, BKMB08, GL08, Håk00, HMBG03, Jag01, Jag09, JL07, Jon07, JRBS02, KMM^{+}00, KPH02, KD05, MMM02, MF06, OWHS09, OzDBS07, PBE^{+}07, PL04, SDL08, WFB^{+}08, WKZP04]. Variational [FL07, Lev00]. variations [Ano04v, BMB07, CSU03, CSU04, CSKP08, CIM07, HY07b, MCD^{+}08, Rob05, WV05]. variegated [DMF07]. varies [BDLL06]. varieties [CMCD04]. various [AAKO^{+}08, BSTS^{+}02, SVB09, Weg00, YL07]. Varroa [WS02]. varying [Ann01, AP06, BDBS08, HS03, MCBA07, OG08]. VATLP [ZZLC06]. VC [CG07, CG06a]. VC-bounds [CG07, CG06a]. vec [HC05b]. vec-permutation [HC05b]. Vector [GH00, YM02, ZZLC06, CAG03, Gan06, Gri08, GKG05, LLL^{+}07, MAB01, RTB04]. vectors [CMSB07, PBS05]. Vegetation [AE02, Han02, KBE^{+}06, LLCL04, SIK07, AN06, AEK^{+}07, Ash06, Asp02, Bal00, BSG07, CZL05, CSHY08, CF09, CHW07, DVPS08, DGRU06, FGB08, GYY00, GYW^{+}04, GBG02, Gil08, GML05, GVC09, GTJ^{+}00, HRH^{+}05, HR03, HO01, HC03b, HVB06, KCD^{+}04, KBB00, KDW^{+}09a, KNM09, KRvL^{+}02, LLW^{+}06, Las06b, MCKNM09, MF02a, MFA07, MC01, MAHvD08, MWP07, MNEB01, Nak08, NHP^{+}06, OHH01, OWHS09, OS02, Pen00, PE07, PVS^{+}09, PWC^{+}09, PSBJ07, Pot04, Roe01, RJS^{+}06, SBDD04b, SBDD04a, Svi00a, SZ03, Svi04, Szi00, TSBH09, VA00, VBRS07, WGB^{+}08, WSZS08, XS08, ZW02, ZJBI03, ZWCL05, ZSZ06, Las06b, Las06a]. vegetational [STK00]. vegetative [DP03]. Vegoritis [AG03]. vehicular [NK06, SDB03]. veitchii [Oga09b]. velocity [KHT06]. Veluwe [CMM02, LvNMvdB04]. Venezuela [MHMHKA04, TSZdRR03]. Venice [Ano04v, BD05, CSU04, CUS01, CSU03, CCBB05, CCMT09, CPB^{+}08, CU06, LS09b, MM06b, PSC^{+}01, SCR03, SPC05, UZ01, ZSPV08]. Venus [GB08]. verge [GDL06a]. verification [AS00c, EB07a, ZAM^{+}03b]. Verlag [Jør04b, Jør04e, Jor05d, Jor05f, Jor05j]. Vermont [Ano08a, LBL^{+}08]. vernal [Pyk04]. versa [MRRJ06]. version [GTRF01, dBWG04, dBWG05]. versions [SC01]. versus [BLVC03, CBMP07, FWS^{+}05, FA03, LH05, LH06, LAPAMD07, MD06a, MVV07, MJ06, PG08, PBS05, RJFAA07, TSJ^{+}09, TA05, WW09, WHP03, WBN^{+}03]. vertebrate [KGA06, OPL^{+}09]. vertebrates [Mey04]. Vertical [GVDF07, BGMP06, BCD^{+}05, DW01, HJZ06, PRRB09, PvdBFJ02, RST05, RJGO00, SL02, WLJ00]. vertical-compressed [HJZ06]. Very [LJR06, KWS^{+}07]. vessels [DLG06]. VHR [LL07b]. VI [Ano04-95]. via [AF09, BMG08, KŞE04, LH07, PSCMMNS06, RD07, dSSGR00]. viability [CDMM08, DDFP07, GGS08, Hin09, KSG05, LTM^{+}04, MY02b, OI07, PT02, TDL^{+}07, ZLX09, vBBE^{+}08]. vice [MRRJ06]. Vietnam [ABV^{+}06, CV07]. VII [Ano04-96]. VIII [Ano04-97]. village [BLBT01]. vineyard [NK04]. violate [Ski03]. Viral [CP02, FSBD01, MSB07, MOJ01]. vireo [PT02]. virgin [Ray08]. virginianus [LGD01]. Virtual [PARH07, Ski04, HHM01, Xia07]. Virus [RMGR09, HTMO06, Mur06, PCL^{+}05, USK^{+}06]. viruses [GH00, LDM00]. Vistula [CT01]. visualisation [CKA^{+}02, RH09]. visualize [HMG06]. visualizing [Abe04]. vital [BDBS08, PMC08]. vitality [SAH03]. vitality-based [SAH03]. Voinov [Jor05i]. Vol [Ano01l, Ano01p, Ano01-54, Ano02k, Ano02n, Ano02-63]. volatiles [LLA^{+}09b]. vole [WG07a]. Volterra [MG09, FGFB05, HZ01, Leg08, LSAGF05, Svi00a, Svi08, TNO^{+}09a]. Volume [Ano01s, Ano01u, Ano01r, Ano01w, Ano02o, Ano02r, Ano02s, Ano02t, Ano02u, Ano02v, Ano02w, Ano02p, Ano02x, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03x, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04r, Ano04s, Ano04-91, Ano05k, Ano05l, Ano05m, Ano01t, Ano01q, Ano01v, Ano02l, Ano02q, Ano02-64, Ano03l, Ano03w, Ano03-78, Ano04p, Ano04q, Ano05g, Ano05h, Ano05i, Ano05n, Ano05o, Ano05p, Ano05-63, Ano05-64, Ano05-65, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06-49, Ano06-50, Ano06-51, Ano06-52, Ano06-53, Ano06-54, Ano06-55, MSL06, PS09]. volumes [Jør07a]. volumetric [DMBM05, KW05]. volunteer [BHM^{+}06, DSD^{+}08]. volunteers [CMCD04]. volutator [SKJvdHG06]. vortex [HPD09]. vorticity [WX09]. vs [Lud09, PPGP08, VPB08]. vulgare [BMD09]. vulgaris [FCP^{+}07]. vulnerability [LWLZ06, LSY^{+}09, Yos08]. Vulpes [Sel00, SHW04]. vulture [DGSBG09].

W [Jør04a, Jor05c, MC04]. Wada [Van04]. Wadden [EKBF04]. wading [MW03]. Wainwright [Jor05b]. waiting [SB00]. walk [iWLSN00, YMT03, ZJC^{+}07]. walleye [FHE06, Jen01]. Walsh [BHFMG05]. Wan [LTLH08]. Ware [KMiIW07]. warm [RvGC^{+}08, CABD09, FSE^{+}04c]. warm-monomictic [RvGC^{+}08]. Warm-water [FSE^{+}04c]. warmer [vLLv^{+}07]. warming [Ale07, FKIL08, GL08, HY07a, MBMP06, TS03]. warranted [JF00]. Wash [WYMS07]. Washington [AB05b, Jør04h, AB05a]. waste [KMM^{+}00, YLSH03]. wastewater [BK05b, SL06]. Water [LXP^{+}08, LT01, PTGI09, TYT^{+}09, TSJ08, vdBDR02, AG03, BW01a, BPW^{+}03, BD05, BB08b, BA08, BW01b, BBGH07, BKC^{+}07, BCLR04, BBKN03, BHP08, CLM^{+}09, Cal05, CJS^{+}07, CJ07, CBS09, CMB05, CWBR01, DFHP04, DPL^{+}04, DWH06, DDG^{+}05, DVPS08, DDF^{+}05, ETH^{+}04, EJG05, ESWG02, FSE^{+}04a, FSE^{+}04b, FSE^{+}04c, FC05b, FDP^{+}03, FGB08, GE05, GACO04, GZY^{+}06, GSM08, HPH00, HBM04, HPA00, HTS00, HHD01, HKPH08, HBRW07, HHKH09, JZY07, KLL^{+}07, KSH^{+}03, KKCC06, KFB09, KFS06, KFH07, KJB07, KNZ04, LHB08, LWC^{+}07, LCY09, Lin06, LFB07, MI01, Man00, MCJ^{+}04, MLL^{+}05, Mon02, Mon09b, MAG01, MS00, NP06, OIR^{+}08, OFK08, PLL04, PL02, PNU03, PCWP06, PCS03, PBCZ01, PCC^{+}07, PLB^{+}06, Rey03, RSW07, RMOA06, SRW05, Sal06, SC04, Sav00, SAR^{+}09, SL04, SRB06, SSCS06, SGY01, SMW02]. water [SBVB05, ŠH09b, SBMJ09, SHK^{+}07, SJM03, SSS00, SJG^{+}08, SGHG04, TN09, Tan02, TFTO07, VVF06, WKZP04, WBR08, WAB^{+}07, Weg00, WM06, WSS^{+}06, WHP03, WMM^{+}07, XS08, YSM00, YABM07, ZVLP05, ZRR^{+}08, ZCY09, ZCZ04, ZGL08, ZAM^{+}05, ZLW09, dSSGR00]. water-constrained [SGHG04]. water-limited [DPL^{+}04, DVPS08]. water-quality [BD05, Cal05, PBCZ01]. waterbird [BA08]. watercourses [AGD06]. waterfowl [Sil04]. waters [BB03a, CCC04, CSSCC07, CPH00, CM04, CM06b, DMH^{+}03, HSV07, Lin01, MSTK08, PCCL03, VADV06]. watershed [CMDP^{+}00, CB07c, GBG^{+}03, GB01, HBO07, TN06, VG08, WFM01, ZAM^{+}07]. watersheds [DBB^{+}08]. wave [Sil07, ZvBS05]. wave-front [Sil07]. wave-like [ZvBS05]. wavelets [KKCC06]. waves [PAF06]. way [Ano04-97, GDL06b, Ort04, TSJ^{+}09, Ulg04, WDBK08]. weak [PBM^{+}05]. weak-links [PBM^{+}05]. Weather [MP07, ČKBB06, LGD01, PvdBW^{+}02, SA07b, TM05b, YZS^{+}04, ZAM^{+}03a, ZAM^{+}03b, ZAM^{+}05]. weathering [vBBE^{+}08]. web [APHV08, AA05, ATD^{+}06, BMT07, CPTD08, CPT09, CN09, Fat07b, GPDF09, GE05, HFL07, JKJ06, JO09, KP09, KSvOO09, KFJ^{+}09, KKZK03, LCP06, LS09b, LCLC07, MMRLP06, MCE^{+}07, NKC04, OB04, OL09, PM06a, PG08, PA09b, PBM^{+}05, Rec03, RLLB09, RMWW07, San07, WVP08, WLB^{+}05, lXzL02b, Yos03, ZK08, dG02b]. webs [DRDD01, FH07a, JSM03, JBP07, Nie08c, PG08, PARH07, QHM05, RKS^{+}07, UDB07, Wil03, Wil05]. weed [CS01, DSD^{+}08, JBR07, PFFM07, SDDC07, WCH08]. weeds [GPB01, RS04, WKL^{+}03]. weevil [ZJG^{+}06]. Weibull [Fle01]. weighing [FMP^{+}00]. Weight [TTHH07]. Weights [RCL06]. Weights-of-Evidence [RCL06]. Weisse [WR08]. weissflogii [BLVC03]. welfare [Sav00]. well [LKKL09]. well-to-well [LKKL09]. wells [SRW05]. west [SL04, HJ02, LS09a, MAA^{+}09, Oke04a, SA07b]. Western [ECBD09, MMM02, RH04, SM07a, AMSH08, CMPO05b, DADGA06, Gar04, HY07a, KRK05, MWH00, NW06, OYiI09, VBFM^{+}08, VPR^{+}09, KKH^{+}09, LGS03, ZSKV05]. wet [AN06, BPC04, KRvL^{+}02]. Wetland [Jor05l, BCD^{+}05, CPJ06, CB07c, CGG^{+}05, FC05b, KM04, KR04, LXJ^{+}03, LBBR01, MTD^{+}09, MD07, MG02, Rai08, SAS06, SL06, TB06b, TAP07, VMR09, WTMG09, YLJY03, ZGL08, vdPvOV00]. wetlands [AM02, BA08, Del04, HKHB06, MLC05, MSA^{+}03, NSEDP06, PLC08, SCP05, SM03, WM00]. wetting [GPL05]. WGEN [TM05b]. Wheat [Kir06b, CT07a, EVF^{+}07, GTRF01, HTS00, JJWF07, MB08, MWD05, NMJ07, SMW02, XWB04, YLLW02, ZvBS05]. Where [Cam04, WGV^{+}08]. Which [HG02, MDH^{+}09, Eza05, Sai07, Ski03]. white [CK07b, HP09, Jen00, MGH^{+}05, MY02b, RBE^{+}08, SLL^{+}06, SJM03, VF07, XLD01, YTH03, ZJG^{+}06]. white-spotted [MY02b]. white-tailed [CK07b, SLL^{+}06, XLD01]. white-toothed [VF07]. white-winged [MGH^{+}05]. whole [DP03, DP07, MD04, WBTC00]. whole-ecosystem [MD04]. WhyWhere [Pet07]. wide [BPA08, DAC^{+}09, Gra04, Gri08, LAZ^{+}08, PMD^{+}09]. widgeon [dSA01]. width [KHT06, TBPF08]. wild [BLBT01, BW04, BNTK04, FSBD01, HAS07, Jen02, KVL^{+}09, LGS^{+}00, MLTN06, SvdWB^{+}06, TTA^{+}03]. wild-type [SvdWB^{+}06]. wildebeest [FG06]. wilderness [JM04]. Wildfire [KTL^{+}05, ES06, Gan06, PBV05]. wildfires [APCdlR07, OMHR06, RADD^{+}01, RMBM06]. wildland [MOLN06]. Wildlife [PSVH09, JJ00, MLPK01, MSM^{+}08, PGM08, Pra09, SC02, YEMZ03, Zha00]. Wiley [Jor05a, Jor05g, Jor05k]. William [Jør04h, Jør09b]. willow [AMSW07, PBC09]. Wind [Bye00, CWCH01, BS04, CCG07, CBS09, CGH^{+}05, GPK00, HLR06, SKP^{+}07, SJ05, BLDN00]. Wind-aided [Bye00]. Wind-driven [CWCH01]. wind-exposed [CBS09]. WINDA [BS04]. windborne [HC09]. windows [Xia05b]. windward [Bar04]. winged [MGH^{+}05, ÖTÖR06]. winter [CUS01, CT07a, HTS00, MLH05, MWD05, NMJ07, PvdBW^{+}02, SMW02, XWB04, YLLW02]. Wisconsin [RCZ^{+}06]. WIT [Jør04g, Jor05h, Zha06]. WIT-Press [Jor05h]. withdrawing [BPP09]. within [BMR07, BK07, BWAM09, BS04, DP07, DMH^{+}03, GG04a, KFN^{+}08, KD05, MNPJ03, MOJ01, New09, PFR09, PG08, POF08, SL04, SLT^{+}09, SHM07, TFTO07, WTMG09, WDW00, YSB08]. without [BM07, Den09, FL07, MTKM^{+}06, Sai07, SCBG09, lXzL02a]. WNMM [LWC^{+}07]. WOFOST [CABD09, ETH^{+}04]. wolf [FJ05, FCKH06, JM01, MJH02, RMS08, VB06, WG04]. wolf-elk [WG04]. wolves [JM04]. wood [BJJ06, DVdB^{+}08, LH09, YSM^{+}06]. wooded [GBG02]. woodland [APEA09, Gil08, HC03b, JG08, KIL^{+}03, MY02a, SMET06, WH04]. woodlands [RBE^{+}08, SVB09]. Woody [Haw00, GIKS08, HHL08, KKH^{+}09, MVPB02, Pet02, RKS^{+}07, TK01, ZKH09, ZBSA07]. words [Bro04b]. work [Den09, Hoy07, Phi04]. Workshop [Ano06-46, Ano08a, Ano09-33, DŽD06, JT09, LBL^{+}08]. world [ATDK08, CWA^{+}09, HN09, SBC^{+}09, SBS^{+}06, Svi08]. worm [WGV^{+}08]. WORMDYN [PBMRE08]. worst [BDE08]. worth [Bro04b]. Would [Svi08]. Wyoming [HLR06].

Xiaoxian [Bio03]. xxii [Log02]. xylostella [MG09, TNO^{+}09a].

yacaré [dCS05]. Yancheng [CL08b]. Yangtze [WWL^{+}05, WX09]. Yasso [LPPS05, PLTT05]. Yasso07 [TTJ^{+}09]. year [CCJ07, SA07a, VSG^{+}08, WV05, ZAM^{+}03a, ZAM^{+}03b]. year-round [ZAM^{+}03a, ZAM^{+}03b]. year-to-year [WV05]. yearly [LG04, LGD01]. years [Bir06, Jør00b, MP07]. yellow [BL01, FPDP06, SR08, TTHH07, LCY09, YLJY03]. yellowfin [DLK01, VREA06]. yellows [ZHMN02]. Yellowstone [VB06]. yield [CDMM08, GGPBEK07, GGH08, HHC02, Kir06b, LL02b, LOM06, LM07b, MHM^{+}03, MOP^{+}05, MLL^{+}05, PS01, PMJ08, VCRD06, WSF^{+}00, WSF^{+}02, Xia02, YSG^{+}06]. yields [Saa00, SMW02, YJJG09]. York [BS08b, Jor05f, Jor05g, Jor05j, MMF^{+}09a, SR08, Sin07]. young [SRG^{+}09, TW00]. youngs [BIS09]. Yoyo [TR03]. Yucatan [AS00b, VCAS01]. Yuri [Ano08-27, Jør08b].

Zea [NV03]. Zealand [BDP^{+}02, CS07, JDPI07, WZKL09, WCS00, ZLO02, ZVK05]. zebra [FG06, MMF^{+}09b, MMF^{+}09a, vNL^{+}08]. ZELIG [LAD06]. zero [BW02, FRM09, GSBM08, GARB09, SHN09]. zero-inflated [SHN09]. zero-inflation [FRM09]. zero-tillage [GARB09]. Zhang [Bio03]. Zhejiang [ZCG^{+}08]. zibethinus [OAAF07]. zinc [CPJ06]. Zn [VCD05]. Zone [Ano06-47, ABP05, CHL08, CWCH09, CF09, DCG01, HTS^{+}07, JWW^{+}07, LKKL09, LCM^{+}09b, MCA^{+}06, NKK^{+}09, PBA05, RFdV07, XGL^{+}09, YLJ^{+}01, ZJG^{+}06, ZCG^{+}08]. zones [BHSR01, HKHB06, MAdlPR02]. zoning [LL08]. Zooplankton [CUSZ09, ECBD09, BHI^{+}06, CP02, DR08, HB03b, KID^{+}07, KGSB01, MRK^{+}07, MRiI^{+}07, MF06, Mit09, PRRB09, RJGO00, RBSJ01, RMWW07, RWM^{+}07, SKCM07, UZSM05, ZRCA08]. Zooprophylactic [HNS08]. Zostera [AACIS^{+}08, PCM^{+}03]. Zürich [ORF01, OBR01]. Zwalm [DGGD04, DVGD07].

References

Angelini:2005:FWM

[AA05] Ronaldo Angelini and Angelo Antonio Agostinho. Food web model of the Upper Paraná River Floodplain: description and aggregation effects. Ecological Modelling, 181(2–3):109–121, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003606.

Abril:2000:MSH

[AAA00] J. M. Abril and M. M. Abdel-Aal. A modelling study on hydrodynamics and pollutant dispersion in the Suez Canal. Ecological Modelling, 128(1):1–17, April 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002227.

Aveytua-Alcazar:2008:MZM

[AACIS^{+}08] Leslie Aveytua-Alcázar, Victor F. Camacho-Ibar, Alejandro J. Souza, J. I. Allen, and Ricardo Torres. Modelling Zostera marina and Ulva spp. in a coastal lagoon. Ecological Modelling, 218(3–4):354–366, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003669.

Andreani-Aksoyoglu:2008:IVE

[AAKO^{+}08]  Şebnem Andreani-Aksoyoğlu, Johannes Keller, Carlos

Ordóñez, Michel Tinguely, Martin Schultz, and André S. H. Prévôt. Influence of various emission scenarios on ozone in Europe. Ecological Modelling, 217(3–4):209–218, October 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008002603.

Abia:2005:ASP

[AALM05] L. M. Abia, O. Angulo, and J. C. López-Marcos. Agestructured population models and their numerical solution. Ecological Modelling, 188(1):112–136, October 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005002450.

Apaydin:2006:TTC

[AAO06] Halit Apaydin, Alper S. Anli, and Ahmet Ozturk. The temporal transferability of calibrated parameters of a hydrological model. Ecological Modelling, 195(3–4):307–317, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005006083.

Abarca-Arenas:2002:ETA

[AAU02] Luis G. Abarca-Arenas and Robert E. Ulanowicz. The effects of taxonomic aggregation on network analysis. Ecological Modelling, 149(3):285–296, April 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004744.

Allesina:2003:SSE

[AB03] Stefano Allesina and Cristina Bondavalli. Steady state of ecosystem flow networks: a comparison between balancing procedures. Ecological Modelling, 165(2–3):221–229, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000759.

Arhonditsis:2005:EMLa

[AB05a] George B. Arhonditsis and Michael T. Brett. Eutrophication model for Lake Washington (USA): Part I. Model description and sensitivity analysis. Ecological Modelling, 187(2–3):140–178, September 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000773.

Arhonditsis:2005:EMLb

[AB05b] George B. Arhonditsis and Michael T. Brett. Eutrophication model for Lake Washington (USA): Part II — model calibration and system dynamics analysis. Ecological Modelling, 187(2–3):179–200, September 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000761.

Arora:2006:ACC

[AB06] Vivek Arora and George Boer. Achieving coexistence: Comment on “Modelling rainforest diversity: the role of competition” by Bampfylde et al. (2005). Ecological Modelling, 192(1–2):322–324, February 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003431.

Allesina:2006:SEE

[ABB06] Stefano Allesina, Antonio Bodini, and Cristina Bondavalli. Secondary extinctions in ecological networks: Bottlenecks unveiled. Ecological Modelling, 194(1–3):150–161, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005132.

Ardente:2003:ESE

[ABC03] Fulvio Ardente, Giorgio Beccali, and Maurizio Cellura. Eco-sustainable energy and environmental strategies in design for recycling: the software “ENDLESS”. Ecological Modelling, 163(1–2):101–118, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004180.

Ardente:2004:FCD

[ABC04] Fulvio Ardente, Marco Beccali, and Maurizio Cellura. F.A.L.C.A.D.E.: a fuzzy software for the energy and environmental balances of products. Ecological Modelling, 176(3–4):359–379, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000997.

Abel:2004:SDV

[Abe04] Thomas Abel. Systems diagrams for visualizing macroeconomics. Ecological Modelling, 178(1–2):189–194, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005143.

Austin:2006:ESM

[ABM^{+}06] M. P. Austin, L. Belbin, J. A. Meyers, M. D. Doherty, and M. Luoto. Evaluation of statistical models used for predicting plant species distributions: Role of artificial data and theory. Ecological Modelling, 199(2):197–216, November 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002584.

Artioli:2005:DMC

[ABP05] Y. Artioli, G. Bendoricchio, and L. Palmeri. Defining and modelling the coastal zone affected by the Po river (Italy). Ecological Modelling, 184(1):55–68, May 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005848.

Antoci:2005:BEG

[ABR05] Angelo Antoci, Simone Borghesi, and Paolo Russu. Biodiversity and economic growth: Trade-offs between stabilization of the ecological system and preservation of natural dynamics. Ecological Modelling, 189(3–4):333–346, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000500205X.

Allesina:2005:CAD

[ABS05] Stefano Allesina, Cristina Bondavalli, and Ursula M. Scharler. The consequences of the aggregation of detritus pools in ecological networks. Ecological Modelling, 189(1–2):221–232, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001997.

Anh:2006:BMN

[ABV^{+}06] Duc Trinh Anh, Marie Paule Bonnet, Georges Vachaud, Chau Van Minh, Nicolas Prieur, Loi Vu Duc, and Le Lan Anh. Biochemical modeling of the Nhue River (Hanoi, Vietnam): Practical identifiability analysis and parameters estimation. Ecological Modelling, 193(3–4):182–204, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004072.

Aguiar:2007:SSA

[ACE07] Ana Paula Dutra Aguiar, Gilberto Câmara, and Maria Isabel Sobral Escada. Spatial statistical analysis of landuse determinants in the Brazilian Amazonia: Exploring intra-regional heterogeneity. Ecological Modelling, 209(2–4):169–188, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003377.

Arii:2008:SHA

[ACJT08] Ken Arii, John P. Caspersen, Trevor A. Jones, and Sean C. Thomas. A selection harvesting algorithm for use in spatially explicit individual-based forest simulation models. Ecological Modelling, 211(3–4):251–266, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004589.

Auger:2000:AEE

[ACVP00] Pierre Auger, Sandrine Charles, Muriel Viala, and Jean-Christophe Poggiale. Aggregation and emergence in ecological modelling: integration of ecological levels. Ecological Modelling, 127(1):11–20, March 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009900201X.

Almeida:2006:NMS

[ADdSC06] Regina C. Almeida, Simone A. Delphim, and Michel I. da S. Costa. A numerical model to solve single-species invasion problems with Allee effects. Ecological Modelling, 192(3–4):601–617, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003741.

Arrignon:2007:MOS

[ADS^{+}07] Florent Arrignon, Marc Deconchat, Jean-Pierre Sarthou, Gérard Balent, and Claude Monteil. Modelling the overwintering strategy of a beneficial insect in a heterogeneous landscape using a multi-agent system. Ecological Modelling, 205(3–4):423–436, July 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700124X.

Agostinho:2008:UEA

[ADSO08] Feni Agostinho, Guaraci Diniz, Raúl Siche, and Enrique Ortega. The use of emergy assessment and the geographical information system in the diagnosis of small family farms in Brazil. Ecological Modelling, 210(1–2):37–57, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003614.

Aassine:2002:VDM

[AE02] Saadia Aassine and Marie C. El Jaï. Vegetation dynamics modelling: a method for coupling local and space dynamics. Ecological Modelling, 154(3):237–249, September 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002000613.

Alados:2007:SOS

[AEK^{+}07] C. L. Alados, A. El Aich, B. Komac, Y. Pueyo, and R. García-Gonzalez. Self-organized spatial patterns of vegetation in alpine grasslands. Ecological Modelling, 201 (2):233–242, February 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004248.

Alados:2004:CPS

[AEP^{+}04] Concepción L. Alados, Ahmed ElAich, Vasilios P. Papanastasis, Huseyin Ozbek, Teresa Navarro, Helena Freitas, Mihalis Vrahnakis, Driss Larrosi, and Baltasar Cabezudo. Change in plant spatial patterns and diversity along the successional gradient of Mediterranean grazing ecosystems. Ecological Modelling, 180(4):523–535, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003266.

Amemiya:2007:SDB

[AER^{+}07] Takashi Amemiya, Takatoshi Enomoto, Axel G. Rossberg, Tetsuya Yamamoto, Yuhei Inamori, and Kiminori Itoh. Stability and dynamical behavior in a lake-model and implications for regime shifts in real lakes. Ecological Modelling, 206(1–2):54–62, August 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001482.

Acerenza:2005:MRB

[AF05] Luis Acerenza and Hugo Fort. A model for the relationship between the interaction pattern of ecosystems and their fate. Ecological Modelling, 184(2–4):203–217, June 1, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004005472.

Acosta:2009:SMN

[AF09] Hernando Acosta and Barrie M. Forrest. The spread of marine non-indigenous species via recreational boating: a conceptual model for risk assessment based on fault tree analysis. Ecological Modelling, 220(13–14):1586–1598, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002245.

Aurambout:2009:CME

[AFLB09] J. P. Aurambout, K. J. Finlay, J. Luck, and G. A. C. Beattie. A concept model to estimate the potential distribution of the Asiatic citrus psyllid (Diaphorina citri kuwayama) in Australia under climate change — a means for assessing biosecurity risk. Ecological Modelling, 220(19):2512–2524, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009003330.

Anctil:2009:NNE

[AFT09] François Anctil, Mélanie Filion, and Julien Tournebize. A neural network experiment on the simulation of daily nitrate-nitrogen and suspended sediment fluxes from a small agricultural catchment. Ecological Modelling, 220 (6):879–887, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008005978.

Andersen:2007:EPF

[AFTB07] K. H. Andersen, K. D. Farnsworth, U. H. Thygesen, and J. E. Beyer. The evolutionary pressure from fishing on size at maturation of Baltic cod. Ecological Modelling, 204(1–2):246–252, May 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000087.

Ambroise:2001:POP

[AG01] C. Ambroise and Y. Grandvalet. Prediction of ozone peaks by mixture models. Ecological Modelling, 145(2– 3):275–289, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003994.

Antonopoulos:2003:SWT

[AG03] Vassilis Z. Antonopoulos and Soultana K. Gianniou. Simulation of water temperature and dissolved oxygen distribution in Lake Vegoritis, Greece. Ecological Modelling, 160(1–2):39–53, February 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002867.

Adriaenssens:2006:FKB

[AGD06] Veronique Adriaenssens, Peter L. M. Goethals, and Niels De Pauw. Fuzzy knowledge-based models for prediction of Asellus and Gammarus in watercourses in Flanders (Belgium). Ecological Modelling, 195(1–2):3–10, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005703.

Andere:2008:MVH

[AGM^{+}08] C. Andere, C. García, C. Marinelli, R. Cepeda, E. M. Rodríguez, and A. Palacio. Morphometric variables of honeybees Apis mellifera used in ecotypes characterization in Argentina. Ecological Modelling, 214(1):53–58, June 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000252.

Arias-Gonzalez:2004:TMI

[AGNLGSG04] J. Ernesto Arias-González, Enrique Nuñez-Lara, Carlos González-Salas, and René Galzin. Trophic models for investigation of fishing effect on coral reef ecosystems. Ecological Modelling, 172(2–4):197–212, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003697.

Anderson:2005:MFP

[AGZ05] James J. Anderson, Eliezer Gurarie, and Richard W. Zabel. Mean free-path length theory of predator–prey interactions: Application to juvenile salmon migration. Ecological Modelling, 186(2):196–211, August 15, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000268.

Antonic:2001:MGR

[AHKB01] Oleg Antonić, Dalibor Hatic, Josip Krian, and Dragan Bukovec. Modelling groundwater regime acceptable for the forest survival after the building of the hydro-electric power plant. Ecological Modelling, 138(1–3):277–288, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004087.

Antonic:2001:BDS

[AHP01] Oleg Antonić, Dalibor Hatic, and Renata Pernar. DEMbased depth in sink as an environmental estimator. Ecological Modelling, 138(1–3):247–254, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004051.

Anzhi:2005:STM

[AJJ^{+}05] Wang Anzhi, Li Jinzhong, Liu Jianmei, Pei Tiefan, and Jin Changjie. A semi-theoretical model of canopy rainfall interception for Pinus koraiensis nakai. Ecological Modelling, 184(2–4):355–361, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005757.

Aviron:2007:CBP

[AKB07] Stephanie Aviron, Pavel Kindlmann, and Francoise Burel. Conservation of butterfly populations in dynamic landscapes: the role of farming practices and landscape mosaic. Ecological Modelling, 205(1–2):135–145, July 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007000841.

Antonic:2001:STI

[AKMB01] Oleg Antonić, Josip Križan, Antun Marki, and Dragan Bukovec. Spatio-temporal interpolation of climatic variables over large region of complex terrain using neural networks. Ecological Modelling, 138(1–3):255–263, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004063.

Albanez-Lucero:2009:MSD

[ALAS09] M. O. Albañez-Lucero and F. Arreguín-Sánchez. Modelling the spatial distribution of red grouper (Epinephelus morio) at Campeche Bank, México, with respect substrate. Ecological Modelling, 220(20):2744–2750, October 24, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009004852.

Asgharbeygi:2006:IRQ

[ALBA06] Nima Asgharbeygi, Pat Langley, Stephen Bay, and Kevin Arrigo. Inductive revision of quantitative process models.

Ecological Modelling, 194(1–3):70–79, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004989.

Alexiadis:2007:GWH

[Ale07] Alessio Alexiadis. Global warming and human activity: a model for studying the potential instability of the carbon dioxide/temperature feedback mechanism. Ecological Modelling, 203(3–4):243–256, May 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600576X.

Alexandrov:2008:FGL

[Ale08] Georgii A. Alexandrov. Forest growth in the light of the thermodynamic theory of ecological systems. Ecological Modelling, 216(2):102–106, August 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001282.

An:2003:MMA

[ALJL03] Min An, D. L. Liu, I. R. Johnson, and J. V. Lovett. Mathematical modelling of allelopathy: II. The dynamics of allelochemicals from living plants in the environment. Ecological Modelling, 161(1–2):53–66, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002892.

Angulo:2000:NIN

[ALM00] Oscar Angulo and J. C. López-Marcos. Numerical integration of nonlinear size-structured population equations. Ecological Modelling, 133(1–2):3–14, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002763.

An:2001:SDS

[ALO^{+}01] Li An, Jianguo Liu, Zhiyun Ouyang, Marc Linderman, Shiqiang Zhou, and Hemin Zhang. Simulating demographic and socioeconomic processes on household level and implications for giant panda habitats. Ecological Modelling, 140(1–2):31–49, May 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002678.

Aloyan:2004:NMM

[Alo04] A. E. Aloyan. Numerical modeling of minor gas constituents and aerosols in the atmosphere. Ecological Modelling, 179(2):163–175, November 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002558.

Anderson:2003:EPM

[ALP03] Robert P. Anderson, Daniel Lew, and A. Townsend Peterson. Evaluating predictive models of species’ distributions: criteria for selecting optimal models. Ecological Modelling, 162(3):211–232, April 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003496.

Ahn:2002:EUR

[AM02] Changwoo Ahn and William J. Mitsch. Evaluating the use of recycled coal combustion products in constructed wetlands: an ecologic-economic modeling approach. Ecological Modelling, 150(1–2):117–140, April 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100477X.

Alemanno:2007:EIP

[AMB07] Sara Alemanno, Giorgio Mancinelli, and Alberto Basset. Effects of invertebrate patch use behaviour and detritus quality on reed leaf decomposition in aquatic systems: a modelling approach. Ecological Modelling, 205(3–4):492–506, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001287.

Aitkenhead:2003:NMT

[AMD^{+}03] M. J. Aitkenhead, A. J. S. McDonald, J. J. Dawson, G. Couper, R. P. Smart, M. Billett, D. Hope, and S. Palmer. A novel method for training neural networks for timeseries prediction in environmental systems. Ecological Modelling, 162(1–2):87–95, April 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004015.

Antonic:2000:GMM

[AMK00] Oleg Antonić, Antun Marki, and Josip Križan. A global model for monthly mean hourly direct solar radiation. Ecological Modelling, 129(2–3):113–118, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002295.

Aitkenhead:2004:UNN

[AMM04] M. J. Aitkenhead, M. J. Mustard, and A. J. S. McDonald. Using neural networks to predict spatial structure in ecological systems. Ecological Modelling, 179(3):393–403, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004002698.

Ascough:2008:FRC

[AMRS08] J. C. Ascough, H. R. Maier, J. K. Ravalico, and M. W. Strudley. Future research challenges for incorporation of uncertainty in environmental and ecological decisionmaking. Ecological Modelling, 219(3–4):383–399, December 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003554.

Araujo:2008:EFS

[AMSH08] Júlio N. Araújo, Steve Mackinson, Richard J. Stanford, and Paul J. B. Hart. Exploring fisheries strategies for the western English Channel using an ecosystem model. Ecological Modelling, 210(4):465–477, February 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004188.

Acharya:2006:PSR

[AMSM06] C. Acharya, S. Mohanty, L. B. Sukla, and V. N. Misra. Prediction of sulphur removal with Acidithiobacillus sp. using artificial neural networks. Ecological Modelling, 190 (1–2):223–230, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002255.

Ahn:2007:DDM

[AMSW07] Changwoo Ahn, Kurt F. Moser, Richard E. Sparks, and David C. White. Developing a dynamic model to predict the recruitment and early survival of black willow (Salix nigra) in response to different hydrologic conditions. Ecological Modelling, 204(3–4):315–325, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700021X.

Aota:2000:MAC

[AN00] Yasuaki Aota and Hisao Nakajima. Mathematical analysis on coexistence conditions of phytoplankton and bacteria systems with nutrient recycling. Ecological Modelling, 135(1):17–31, November 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003495.

Anderson:2002:PSS

[AN02] Kevin Anderson and Claudia Neuhauser. Patterns in spatial simulations — are they real? Ecological Modelling, 155(1):19–30, September 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000704.

Accad:2006:MPC

[AN06] Arnon Accad and David T. Neil. Modelling pre-clearing vegetation distribution using GIS-integrated statistical, ecological and data models: a case study from the wet tropics of Northeastern Australia. Ecological Modelling, 198(1–2):85–100, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001803.

Annan:2001:MUU

[Ann01] J. D. Annan. Modelling under uncertainty: Monte Carlo methods for temporally varying parameters. Ecological Modelling, 136(2–3):297–302, January 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004130.

Anonymous:2000:Ia

[Ano00a] Anonymous. Index. Ecological Modelling, 125(2–3):297, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099002100.

Anonymous:2000:Ib

[Ano00b] Anonymous. Index. Ecological Modelling, 125(2–3):299– 300, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380099002112.

Anonymous:2000:Ic

[Ano00c] Anonymous. Index. Ecological Modelling, 126(2–3):299– 317, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000003161.

Anonymous:2000:Id

[Ano00d] Anonymous. Index. Ecological Modelling, 126(2–3):319– 334, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000003173.

Anonymous:2000:Ie

[Ano00e] Anonymous. Index. Ecological Modelling, 126(2–3):335, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000003185.

Anonymous:2000:If

[Ano00f] Anonymous. Index. Ecological Modelling, 126(2–3):337– 338, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000003197.

Anonymous:2000:Ig

[Ano00g] Anonymous. Index. Ecological Modelling, 127(2–3): 313, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000002222.

Anonymous:2000:Ih

[Ano00h] Anonymous. Index. Ecological Modelling, 127(2–3):315– 316, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000002234.

Anonymous:2000:Ii

[Ano00i] Anonymous. Index. Ecological Modelling, 128(2–3):253, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000002441.

Anonymous:2000:Ij

[Ano00j] Anonymous. Index. Ecological Modelling, 128(2–3):255–256, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000002453.

Anonymous:2000:Ik

[Ano00k] Anonymous. Index. Ecological Modelling, 129(2–3):307, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003264.

Anonymous:2000:Il

[Ano00l] Anonymous. Index. Ecological Modelling, 129(2–3):309–310, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003276.

Anonymous:2000:Im

[Ano00m] Anonymous. Index. Ecological Modelling, 130(1–3):167, June 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003318.

Anonymous:2000:In

[Ano00n] Anonymous. Index. Ecological Modelling, 130(1–3):169–170, June 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000000332X.

Anonymous:2000:Io

[Ano00o] Anonymous. Index. Ecological Modelling, 131(2–3):319, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003343.

Anonymous:2000:Ip

[Ano00p] Anonymous. Index. Ecological Modelling, 131(2–3):321–322, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003355.

Anonymous:2000:Iq

[Ano00q] Anonymous. Index. Ecological Modelling, 132(3):305, August 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003409.

Anonymous:2000:Ir

[Ano00r] Anonymous. Index. Ecological Modelling, 132(3):307– 308, August 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000003410.

Anonymous:2000:Is

[Ano00s] Anonymous. Index. Ecological Modelling, 133(3):269, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000003616.

Anonymous:2000:It

[Ano00t] Anonymous. Index. Ecological Modelling, 133(3):271–272, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000003628.

Anonymous:2000:Iu

[Ano00u] Anonymous. Index. Ecological Modelling, 134(2–3):357, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003768.

Anonymous:2000:Iv

[Ano00v] Anonymous. Index. Ecological Modelling, 134(2–3):359– 360, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438000000377X.

Anonymous:2000:Iw

[Ano00w] Anonymous. Index. Ecological Modelling, 135(2–3):327, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000004142.

Anonymous:2000:Ix

[Ano00x] Anonymous. Index. Ecological Modelling, 135(2–3):329– 330, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000004154.

Anonymous:2000:PJa

[Ano00y] Anonymous. Pages 1–108 (1 January 2000). Ecological Modelling, 125(1):??, January 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PN

[Ano00z] Anonymous. Pages 1–110 (25 November 2000). Ecological Modelling, 135(1):??, November 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PMc

[Ano00-27] Anonymous. Pages 1–112 (3 May 2000). Ecological Modelling, 129(1):??, May 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PSb

[Ano00-28] Anonymous. Pages 1–116 (30 September 2000). Ecological Modelling, 134(1):??, September 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PAd

[Ano00-29] Anonymous. Pages 1–160 (25 August 2000). Ecological Modelling, 133(1–2):??, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PJc

[Ano00-30] Anonymous. Pages 1–172 (1 June 2000). Ecological Modelling, 130(1–3):??, June 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PJf

[Ano00-31] Anonymous. Pages 1–174 (30 July 2000). Ecological Modelling, 132(1–2):??, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PJd

[Ano00-32] Anonymous. Pages 1–78 (30 June 2000). Ecological Modelling, 131(1):??, June 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PAa

[Ano00-33] Anonymous. Pages 1–88 (3 April 2000). Ecological Modelling, 128(1):??, April 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PFa

[Ano00-34] Anonymous. Pages 1–94 (1 February 2000). Ecological Modelling, 126(1):??, February 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PMa

[Ano00-35] Anonymous. Pages 1–98 (1 March 2000). Ecological Modelling, 127(1):??, March 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PJb

[Ano00-36] Anonymous. Pages 109–302 (15 January 2000). Ecological Modelling, 125(2–3):??, January 15, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PD

[Ano00-37] Anonymous. Pages 111–332 (5 December 2000). Ecological Modelling, 135(2–3):??, December 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PMd

[Ano00-38] Anonymous. Pages 113–312 (30 May 2000). Ecological Modelling, 129(2–3):??, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PO

[Ano00-39] Anonymous. Pages 117–364 (30 October 2000). Ecological Modelling, 134(2–3):??, October 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PSa

[Ano00-40] Anonymous. Pages 161–274 (3 September 2000). Ecological Modelling, 133(3):??, September 3, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PAc

[Ano00-41] Anonymous. Pages 175–310 (5 August 2000). Ecological Modelling, 132(3):??, August 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PJe

[Ano00-42] Anonymous. Pages 79–324 (1 July 2000). Ecological Modelling, 131(2–3):??, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PAb

[Ano00-43] Anonymous. Pages 89–258 (20 April 2000). Ecological Modelling, 128(2–3):??, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PFb

[Ano00-44] Anonymous. Pages 95–340 (28 February 2000). Ecological Modelling, 126(2–3):??, February 28, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2000:PMb

[Ano00-45] Anonymous. Pages 99–317 (30 March 2000). Ecological Modelling, 127(2–3):??, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:AS

[Ano01a] Anonymous. Ad — SD. Ecological Modelling, 137(2–3): I, February 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001002526.

Anonymous:2001:AIa

[Ano01b] Anonymous. Author index. Ecological Modelling, 136(2–3): 307, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001002241.

Anonymous:2001:AIb

[Ano01c] Anonymous. Author index. Ecological Modelling, 137 (2–3):241, February 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002502.

Anonymous:2001:AIc

[Ano01d] Anonymous. Author index. Ecological Modelling, 138 (1–3):347–348, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002897.

Anonymous:2001:AId

[Ano01e] Anonymous. Author index. Ecological Modelling, 139(2– 3):309, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001003350.

Anonymous:2001:AIe

[Ano01f] Anonymous. Author index. Ecological Modelling, 140(3): 293, June 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001003465.

Anonymous:2001:AIf

[Ano01g] Anonymous. Author index. Ecological Modelling, 141(1– 3):313, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001003647.

Anonymous:2001:AIg

[Ano01h] Anonymous. Author index. Ecological Modelling, 142(3): 295, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001004100.

Anonymous:2001:AIh

[Ano01i] Anonymous. Author index. Ecological Modelling, 144(2–3): 315, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001004434.

Anonymous:2001:AIi

[Ano01j] Anonymous. Author index. Ecological Modelling, 145 (2–3):291, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004471.

Anonymous:2001:AIj

[Ano01k] Anonymous. Author index. Ecological Modelling, 146 (1–3):341, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004501.

Anonymous:2001:AIV

[Ano01l] Anonymous. Author index vol. 143. Ecological Modelling, 143(3):253, September 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004306.

Anonymous:2001:CAC

[Ano01m] Anonymous. Conference announcement/call for papers. Ecological Modelling, 146(1–3):I, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004537.

Anonymous:2001:Ca

[Ano01n] Anonymous. Contents. Ecological Modelling, 138(1–3):351– 352, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001003179.

Anonymous:2001:Cb

[Ano01o] Anonymous. Contents. Ecological Modelling, 139(2–3):313, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003374.

Anonymous:2001:CVe

[Ano01p] Anonymous. Contents of vol. 143. Ecological Modelling, 143(3):257–258, September 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100432X.

Anonymous:2001:CVd

[Ano01q] Anonymous. Contents of volume. Ecological Modelling, 142(3):299–300, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004124.

Anonymous:2001:CVg

[Ano01r] Anonymous. Contents of volume. Ecological Modelling, 145(2–3):295–296, November 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004495.

Anonymous:2001:CVa

[Ano01s] Anonymous. Contents of volume 137. Ecological Modelling, 137(2–3):245, February 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002496.

Anonymous:2001:CVb

[Ano01t] Anonymous. Contents of volume 140. Ecological Modelling, 140(3):297–298, June 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003489.

Anonymous:2001:CVc

[Ano01u] Anonymous. Contents of volume 141. Ecological Modelling, 141(1–3):317–318, July 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003660.

Anonymous:2001:CVf

[Ano01v] Anonymous. Contents of volume 144. Ecological Modelling, 144(2–3):319–320, October 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004458.

Anonymous:2001:CVh

[Ano01w] Anonymous. Contents of volume 146. Ecological Modelling, 146(1–3):345–346, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004525.

Anonymous:2001:PJa

[Ano01x] Anonymous. Pages 1–102 (1 January 2001). Ecological Modelling, 136(1):??, January 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PFa

[Ano01y] Anonymous. Pages 1–108 (1 February 2001). Ecological Modelling, 137(1):??, February 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:POa

[Ano01z] Anonymous. Pages 1–146 (15 October 2001). Ecological Modelling, 143(1–2):??, October 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PAb

[Ano01-27] Anonymous. Pages 1–187 (1 August 2001). Ecological Modelling, 142(1–2):??, August 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PMc

[Ano01-28] Anonymous. Pages 1–192 (30 May 2001). Ecological Modelling, 140(1–2):??, May 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PJd

[Ano01-29] Anonymous. Pages 1–318 (1 July 2001). Ecological Modelling, 141(1–3):??, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PD

[Ano01-30] Anonymous. Pages 1–346 (1 December 2001). Ecological Modelling, 146(1–3):??, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PMa

[Ano01-31] Anonymous. Pages 1–352 (15 March 2001). Ecological Modelling, 138(1–3):??, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:POb

[Ano01-32] Anonymous. Pages 1–98 (1 October 2001). Ecological Modelling, 144(1):??, October 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PNb

[Ano01-33] Anonymous. Pages 1–99 (1 November 2001). Ecological Modelling, 145(1):??, November 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PMb

[Ano01-34] Anonymous. Pages 1–99 (30 March 2001). Ecological Modelling, 139(1):??, March 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PNc

[Ano01-35] Anonymous. Pages 101–296 (15 November 2001). Ecological Modelling, 145(2–3):??, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PAa

[Ano01-36] Anonymous. Pages 101–314 (30 April 2001). Ecological Modelling, 139(2–3):??, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PJb

[Ano01-37] Anonymous. Pages 103–310 (20 January 2001). Ecological Modelling, 136(2–3):??, January 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PFb

[Ano01-38] Anonymous. Pages 109–245 (20 February 2001). Ecological Modelling, 137(2–3):??, February 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PS

[Ano01-39] Anonymous. Pages 147–258 (15 September 2001). Ecological Modelling, 143(3):??, September 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PAc

[Ano01-40] Anonymous. Pages 189–300 (15 August 2001). Ecological Modelling, 142(3):??, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PJc

[Ano01-41] Anonymous. Pages 193–298 (15 June 2001). Ecological Modelling, 140(3):??, June 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:POc

[Ano01-42] Anonymous. Pages 99–320 (15 October 2001). Ecological Modelling, 144(2–3):??, October 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2001:PNa

[Ano01-43] Anonymous. Publishers note. Ecological Modelling, 137 (2–3):iii–iv, February 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002484.

Anonymous:2001:SIa

[Ano01-44] Anonymous. Subject index. Ecological Modelling, 136 (2–3):309–310, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002253.

Anonymous:2001:SIb

[Ano01-45] Anonymous. Subject index. Ecological Modelling, 138 (1–3):349–350, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002903.

Anonymous:2001:SIc

[Ano01-46] Anonymous. Subject index. Ecological Modelling, 139 (2–3):311–312, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003362.

Anonymous:2001:SId

[Ano01-47]  Anonymous.  Subject index.  Ecological Modelling,

140(3):295–296, June 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003477.

Anonymous:2001:SIe

[Ano01-48]  Anonymous.  Subject index.  Ecological Modelling,

141(1–3):315–316, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003659.

Anonymous:2001:SIf

[Ano01-49] Anonymous. Subject index. Ecological Modelling, 142 (3):297–298, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004112.

Anonymous:2001:SIg

[Ano01-50]  Anonymous.  Subject index.  Ecological Modelling,

143(3):255, September 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004318.

Anonymous:2001:SIh

[Ano01-51] Anonymous. Subject index. Ecological Modelling, 144 (2–3):317–318, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004446.

Anonymous:2001:SIi

[Ano01-52] Anonymous. Subject index. Ecological Modelling, 145 (2–3):293–294, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004483.

Anonymous:2001:SIj

[Ano01-53] Anonymous. Subject index. Ecological Modelling, 146 (1–3):343–344, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004513.

Anonymous:2001:SIV

[Ano01-54] Anonymous. Subject index of vol 137. Ecological Modelling, 137(2–3):243–244, February 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002514.

Anonymous:2002:AIa

[Ano02a] Anonymous. Author index. Ecological Modelling, 148(3): 311, March 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380002000224.

Anonymous:2002:AIb

[Ano02b] Anonymous. Author index. Ecological Modelling, 149(3): 315, April 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380002000352.

Anonymous:2002:AIc

[Ano02c] Anonymous. Author index. Ecological Modelling, 150 (3):309, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438000200039X.

Anonymous:2002:AId

[Ano02d] Anonymous. Author index. Ecological Modelling, 151(2– 3):309, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380002000728.

Anonymous:2002:AIe

[Ano02e] Anonymous. Author index. Ecological Modelling, 152(2– 3):303, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380002000959.

Anonymous:2002:AIf

[Ano02f] Anonymous. Author index. Ecological Modelling, 153 (3):297–298, August 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001199.

Anonymous:2002:AIg

[Ano02g] Anonymous. Author index. Ecological Modelling, 154(3): 301, September 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380002001539.

Anonymous:2002:AIh

[Ano02h] Anonymous. Author index. Ecological Modelling, 155(2–3): 231, October 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380002001771.

Anonymous:2002:AIi

[Ano02i] Anonymous. Author index. Ecological Modelling, 156 (2–3):237, November 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002002387.

Anonymous:2002:AIj

[Ano02j] Anonymous. Author index. Ecological Modelling, 157 (2–3):343, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000200306X.

Anonymous:2002:AIVb

[Ano02k] Anonymous. Author index vol 158. Ecological Modelling, 158(3):249, December 31, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003368.

Anonymous:2002:AIVa

[Ano02l] Anonymous. Author index volume 147. Ecological Modelling, 147(3):309, January 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004951.

Anonymous:2002:C

[Ano02m] Anonymous. Contents. Ecological Modelling, 150(3):313, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002000418.

Anonymous:2002:CVc

[Ano02n] Anonymous. Contents of vol 149. Ecological Modelling, 149(3):319–320, April 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000376.

Anonymous:2002:CVb

[Ano02o] Anonymous. Contents of volume. Ecological Modelling, 148(3):315–316, March 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000248.

Anonymous:2002:CVj

[Ano02p] Anonymous. Contents of volume. Ecological Modelling, 157(2–3):347–348, November 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003083.

Anonymous:2002:CVa

[Ano02q] Anonymous. Contents of volume 147. Ecological Modelling, 147(3):313–314, January 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004975.

Anonymous:2002:CVd

[Ano02r] Anonymous. Contents of volume 151. Ecological Modelling, 151(2–3):313–314, June 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000741.

Anonymous:2002:CVe

[Ano02s] Anonymous. Contents of volume 152. Ecological Modelling, 152(2–3):307–308, July 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000972.

Anonymous:2002:CVf

[Ano02t] Anonymous. Contents of volume 153. Ecological Modelling, 153(3):301–302, August 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001217.

Anonymous:2002:CVg

[Ano02u] Anonymous. Contents of volume 154. Ecological Modelling, 154(3):305–306, September 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001552.

Anonymous:2002:CVh

[Ano02v] Anonymous. Contents of volume 155. Ecological Modelling, 155(2–3):235–236, October 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001795.

Anonymous:2002:CVi

[Ano02w] Anonymous. Contents of volume 156. Ecological Modelling, 156(2–3):241–242, November 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002405.

Anonymous:2002:CVk

[Ano02x] Anonymous. Contents of volume 158. Ecological Modelling, 158(3):253–254, December 31, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003381.

Anonymous:2002:ER

[Ano02y] Anonymous. Expert reviewers 2001. Ecological Modelling, 149(3):213–214, April 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000388.

Anonymous:2002:PJe

[Ano02z] Anonymous. Pages 1–102 (15 June 2002). Ecological Modelling, 152(1):??, June 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PJa

[Ano02-27] Anonymous. Pages 1–104 (1 January 2002). Ecological Modelling, 147(1):??, January 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PMd

[Ano02-28] Anonymous. Pages 1–108 (15 May 2002). Ecological Modelling, 151(1):??, May 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PFa

[Ano02-29] Anonymous. Pages 1–109 (1 February 2002). Ecological Modelling, 148(1):??, February 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PDa

[Ano02-30] Anonymous. Pages 1–165 (15 December 2002). Ecological Modelling, 158(1–2):??, December 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PJg

[Ano02-31] Anonymous. Pages 1–196 (15 July 2002). Ecological Modelling, 153(1–2):??, July 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PAd

[Ano02-32] Anonymous. Pages 1–202 (15 August 2002). Ecological Modelling, 154(1–2):??, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PAb

[Ano02-33] Anonymous. Pages 1–209 (15 April 2002). Ecological Modelling, 150(1–2):??, April 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PMb

[Ano02-34] Anonymous. Pages 1–211 (15 March 2002). Ecological Modelling, 149(1–2):??, March 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PNb

[Ano02-35] Anonymous. Pages 1–88 (15 November 2002). Ecological Modelling, 157(1):??, November 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:POb

[Ano02-36] Anonymous. Pages 1–91 (15 October 2002). Ecological Modelling, 156(1):??, October 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PSb

[Ano02-37] Anonymous. Pages 1–94 (15 September 2002). Ecological Modelling, 155(1):??, September 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PJf

[Ano02-38] Anonymous. Pages 103–308 (1 July 2002). Ecological Modelling, 152(2–3):??, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PJb

[Ano02-39] Anonymous. Pages 105–197 (15 January 2002). Ecological Modelling, 147(2):??, January 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PJd

[Ano02-40] Anonymous. Pages 109–314 (1 June 2002). Ecological Modelling, 151(2–3):??, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PFb

[Ano02-41] Anonymous. Pages 111–212 (15 February 2002). Ecological Modelling, 148(2):??, February 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PDb

[Ano02-42] Anonymous. Pages 167–254 (31 December 2002). Ecological Modelling, 158(3):??, December 31, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PAc

[Ano02-43] Anonymous. Pages 197–302 (1 August 2002). Ecological Modelling, 153(3):??, August 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PJc

[Ano02-44] Anonymous. Pages 199–314 (30 January 2002). Ecological Modelling, 147(3):??, January 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PSa

[Ano02-45] Anonymous. Pages 203–306 (1 September 2002). Ecological Modelling, 154(3):??, September 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PMc

[Ano02-46] Anonymous. Pages 211–314 (1 May 2002). Ecological Modelling, 150(3):??, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PMa

[Ano02-47] Anonymous. Pages 213–316 (1 March 2002). Ecological Modelling, 148(3):??, March 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PAa

[Ano02-48] Anonymous. Pages 213–320 (1 April 2002). Ecological Modelling, 149(3):??, April 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PNc

[Ano02-49] Anonymous. Pages 89–348 (30 November 2002). Ecological Modelling, 157(2–3):??, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PNa

[Ano02-50] Anonymous. Pages 93–242 (1 November 2002). Ecological Modelling, 156(2–3):??, November 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:POa

[Ano02-51] Anonymous. Pages 95–236 (1 October 2002). Ecological Modelling, 155(2–3):??, October 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2002:PTP

[Ano02-52] Anonymous. Prelim title page (iii). Ecological Modelling, 158(1–2):iii, December 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003162.

Anonymous:2002:SIa

[Ano02-53]  Anonymous.  Subject index.  Ecological Modelling,

148(3):313–314, March 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000236.

Anonymous:2002:SIb

[Ano02-54]  Anonymous.  Subject index.  Ecological Modelling,

149(3):317–318, April 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000364.

Anonymous:2002:SIc

[Ano02-55]  Anonymous.  Subject index.  Ecological Modelling,

150(3):311–312, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000406.

Anonymous:2002:SId

[Ano02-56]  Anonymous.  Subject index.  Ecological Modelling,

151(2–3):311–312, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000200073X.

Anonymous:2002:SIe

[Ano02-57]  Anonymous.  Subject index.  Ecological Modelling,

152(2–3):305–306, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000960.

Anonymous:2002:SIf

[Ano02-58] Anonymous. Subject index. Ecological Modelling, 153 (3):299–300, August 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001205.

Anonymous:2002:SIg

[Ano02-59] Anonymous. Subject index. Ecological Modelling, 154 (3):303–304, September 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001540.

Anonymous:2002:SIh

[Ano02-60] Anonymous. Subject index. Ecological Modelling, 155 (2–3):233–234, October 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001783.

Anonymous:2002:SIi

[Ano02-61] Anonymous. Subject index. Ecological Modelling, 156 (2–3):239–240, November 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002002399.

Anonymous:2002:SIj

[Ano02-62] Anonymous. Subject index. Ecological Modelling, 157 (2–3):345–346, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002003071.

Anonymous:2002:SIVb

[Ano02-63] Anonymous. Subject index vol 158. Ecological Modelling, 158(3):251–252, December 31, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200337X.

Anonymous:2002:SIVa

[Ano02-64] Anonymous. Subject index volume 147. Ecological Modelling, 147(3):311–312, January 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004963.

Anonymous:2003:AIa

[Ano03a] Anonymous. Author index. Ecological Modelling, 159(2–3): 297, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380002003757.

Anonymous:2003:AIb

[Ano03b] Anonymous. Author index. Ecological Modelling, 160(3):I, February 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003000024.

Anonymous:2003:AIc

[Ano03c] Anonymous. Author index. Ecological Modelling, 162(3): 267, April 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380003000589.

Anonymous:2003:AId

[Ano03d] Anonymous. Author index. Ecological Modelling, 163(3): 269, May 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380003000905.

Anonymous:2003:AIe

[Ano03e] Anonymous. Author index. Ecological Modelling, 164(2– 3):283, June 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380003001200.

Anonymous:2003:AIg

[Ano03f] Anonymous. Author index. Ecological Modelling, 166 (3):297–298, August 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002540.

Anonymous:2003:AIh

[Ano03g] Anonymous. Author index. Ecological Modelling, 167 (3):323–324, September 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003144.

Anonymous:2003:AIi

[Ano03h] Anonymous. Author index. Ecological Modelling, 168(3): 391–392, October 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438000300348X.

Anonymous:2003:AIj

[Ano03i] Anonymous. Author index. Ecological Modelling, 169 (2–3):367, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004058.

Anonymous:2003:AIk

[Ano03j] Anonymous. Author index. Ecological Modelling, 170 (2–3):481–482, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004435.

Anonymous:2003:AIf

[Ano03k] Anonymous. Author index 165. Ecological Modelling, 165 (2–3):303, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380003001777.

Anonymous:2003:AIV

[Ano03l] Anonymous. Author index of volume 161. Ecological Modelling, 161(3):249, March 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000358.

Anonymous:2003:CVa

[Ano03m] Anonymous. Contents of volume. Ecological Modelling, 159(2–3):301–302, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002003770.

Anonymous:2003:CVb

[Ano03n] Anonymous. Contents of volume. Ecological Modelling, 160(3):V–VI, February 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000048.

Anonymous:2003:CVd

[Ano03o] Anonymous. Contents of volume. Ecological Modelling, 162(3):271–272, April 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000607.

Anonymous:2003:CVe

[Ano03p] Anonymous. Contents of volume. Ecological Modelling, 163(3):273–274, May 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000929.

Anonymous:2003:CVf

[Ano03q] Anonymous. Contents of volume. Ecological Modelling, 164(2–3):287–288, June 15, 2003. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003001224.

Anonymous:2003:CVh

[Ano03r] Anonymous. Contents of volume. Ecological Modelling, 166(3):301–302, August 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002564.

Anonymous:2003:CVi

[Ano03s] Anonymous. Contents of volume. Ecological Modelling, 167(3):327–328, September 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003168.

Anonymous:2003:CVj

[Ano03t] Anonymous. Contents of volume. Ecological Modelling, 168(3):395–396, October 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003508.

Anonymous:2003:CVk

[Ano03u] Anonymous. Contents of volume. Ecological Modelling, 169(2–3):371–372, November 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004071.

Anonymous:2003:CVl

[Ano03v] Anonymous. Contents of volume. Ecological Modelling, 170(2–3):487–488, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004459.

Anonymous:2003:CVc

[Ano03w] Anonymous. Contents of volume 161. Ecological Modelling, 161(3):253–254, March 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000371.

Anonymous:2003:CVg

[Ano03x] Anonymous. Contents of volume 165. Ecological Modelling, 165(2–3):307–308, July 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001790.

Anonymous:2003:E

[Ano03y] Anonymous. Editorial. Ecological Modelling, 166(1–2): 1–2, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380003001637.

Anonymous:2003:EIT

[Ano03z] Anonymous. Editorial: ISEM: the Third European Ecological Modelling Conference 10–15 September, 2001, Dubrovnik, Croatia. Ecological Modelling, 170(2–3):105, December 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003004460.

Anonymous:2003:ERM

[Ano03-27] Anonymous. Erratum to “Regulation of mycorrhizal symbiosis: development of a structured nutritional dual model”: [Ecological Modelling 158 (2002) 121–142]. Ecological Modelling, 163(3):245, May 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000231.

Anonymous:2003:IFCa

[Ano03-28] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 162(3):ifc, April 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300053X.

Anonymous:2003:IFCb

[Ano03-29] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 163(1–2):ifc, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000796.

Anonymous:2003:IFCc

[Ano03-30] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 163(3):ifc, May 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000851.

Anonymous:2003:IFCd

[Ano03-31] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 164(2–3):ifc, June 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001248.

Anonymous:2003:IFCe

[Ano03-32] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 165(2–3):ifc, July 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001819.

Anonymous:2003:IFCf

[Ano03-33] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 166(3):ifc, August 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002588.

Anonymous:2003:IFCg

[Ano03-34] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 167(3):ifc, September 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003090.

Anonymous:2003:IFCh

[Ano03-35] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 168(3):ifc, October 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003429.

Anonymous:2003:IFCi

[Ano03-36] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 169(2–3):ifc, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004009.

Anonymous:2003:IFCj

[Ano03-37] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 170(2–3):ifc, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300437X.

Anonymous:2003:PJa

[Ano03-38] Anonymous. Pages 1–101 (1 January 2003). Ecological Modelling, 159(1):??, January 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PJc

[Ano03-39] Anonymous. Pages 1–102 (1 June 2003). Ecological Modelling, 164(1):??, June 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PD

[Ano03-40] Anonymous. Pages 1–104 (1 December 2003). Ecological Modelling, 170(1):??, December 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PJe

[Ano03-41] Anonymous. Pages 1–106 (1 July 2003). Ecological Modelling, 165(1):??, July 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PMa

[Ano03-42] Anonymous. Pages 1–149 (1 March 2003). Ecological Modelling, 161(1–2):??, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PMc

[Ano03-43] Anonymous. Pages 1–174 (1 May 2003). Ecological Modelling, 163(1–2):??, May 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PAa

[Ano03-44] Anonymous. Pages 1–176 (1 April 2003). Ecological Modelling, 162(1–2):??, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PF

[Ano03-45] Anonymous. Pages 1–183 (1 February 2003). Ecological Modelling, 160(1–2):??, February 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PAc

[Ano03-46] Anonymous. Pages 1–192 (1 August 2003). Ecological Modelling, 166(1–2):??, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PNa

[Ano03-47] Anonymous. Pages 1–212 (1 November 2003). Ecological Modelling, 169(1):??, November 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PSa

[Ano03-48] Anonymous. Pages 1–212 (1 September 2003). Ecological Modelling, 167(1–2):??, September 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PO

[Ano03-49] Anonymous. Pages 1–214 (1 October 2003). Ecological Modelling, 168(1–2):??, October 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PJd

[Ano03-50] Anonymous. Pages 103–290 (15 June 2003). Ecological Modelling, 164(2–3):??, June 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PJb

[Ano03-51] Anonymous. Pages 103–302 (15 January 2003). Ecological Modelling, 159(2–3):??, January 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PJf

[Ano03-52] Anonymous. Pages 107–308 (15 July 2003). Ecological Modelling, 165(2–3):??, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PMb

[Ano03-53] Anonymous. Pages 151–254 (March 2003). Ecological Modelling, 161(3):??, March 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PMd

[Ano03-54] Anonymous. Pages 175–274 (15 May 2003). Ecological Modelling, 163(3):??, May 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PAb

[Ano03-55] Anonymous. Pages 177–272 (15 April 2003). Ecological Modelling, 162(3):??, April 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PAd

[Ano03-56] Anonymous. Pages 193–302 (15 August 2003). Ecological Modelling, 166(3):??, August 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PSb

[Ano03-57] Anonymous. Pages 213–328 (15 September 2003). Ecological Modelling, 167(3):??, September 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PNb

[Ano03-58] Anonymous. Pages 213–372 (15 November 2003). Ecological Modelling, 169(2–3):??, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2003:PITa

[Ano03-59] Anonymous. prelim (iii) title page. Ecological Modelling, 162(1–2):iii, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000462.

Anonymous:2003:PITc

[Ano03-60] Anonymous. Prelim (iii) title page. Ecological Modelling, 166(1–2):iii, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002138.

Anonymous:2003:PITd

[Ano03-61] Anonymous. Prelim (iii) title page. Ecological Modelling, 167(1–2):iii, September 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002916.

Anonymous:2003:PITe

[Ano03-62] Anonymous. Prelim (iii) title page. Ecological Modelling, 168(1–2):iii, October 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003260.

Anonymous:2003:PITf

[Ano03-63] Anonymous. Prelim (iii) title page. Ecological Modelling, 169(1):iii, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003880.

Anonymous:2003:PITb

[Ano03-64] Anonymous. Prelim(iii) title page. Ecological Modelling, 164(1):iii, June 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000300108X.

Anonymous:2003:PITg

[Ano03-65] Anonymous. Prelim(iii) title page. Ecological Modelling, 170(1):iii, December 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004198.

Anonymous:2003:PNc

[Ano03-66] Anonymous. Publisher’s note. Ecological Modelling, 169(2–3):213, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004083.

Anonymous:2003:SIa

[Ano03-67] Anonymous. Subject index. Ecological Modelling, 159 (2–3):299–300, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002003769.

Anonymous:2003:SIb

[Ano03-68] Anonymous. Subject index. Ecological Modelling, 160(3): III–IV, February 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380003000036.

Anonymous:2003:SIc

[Ano03-69]  Anonymous.  Subject index.  Ecological Modelling,

162(3):269–270, April 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000590.

Anonymous:2003:SId

[Ano03-70]  Anonymous.  Subject index.  Ecological Modelling,

163(3):271–272, May 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000917.

Anonymous:2003:SIe

[Ano03-71] Anonymous. Subject index. Ecological Modelling, 164 (2–3):285–286, June 15, 2003. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003001212.

Anonymous:2003:SIg

[Ano03-72] Anonymous. Subject index. Ecological Modelling, 166 (3):299–300, August 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002552.

Anonymous:2003:SIh

[Ano03-73] Anonymous. Subject index. Ecological Modelling, 167 (3):325–326, September 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003156.

Anonymous:2003:SIi

[Ano03-74]  Anonymous.  Subject index.  Ecological Modelling,

168(3):393–394, October 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003491.

Anonymous:2003:SIj

[Ano03-75] Anonymous. Subject index. Ecological Modelling, 169 (2–3):369–370, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000300406X.

Anonymous:2003:SIk

[Ano03-76] Anonymous. Subject index. Ecological Modelling, 170 (2–3):483–485, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004447.

Anonymous:2003:SIf

[Ano03-77] Anonymous. Subject index 165. Ecological Modelling, 165(2–3):305–306, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003001789.

Anonymous:2003:SIV

[Ano03-78] Anonymous. Subject index of volume 161. Ecological Modelling, 161(3):251–252, March 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300036X.

Anonymous:2003:TPa

[Ano03-79] Anonymous. Title page. Ecological Modelling, 161(1–2):iii, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000300019X.

Anonymous:2003:TPb

[Ano03-80] Anonymous. Title page. Ecological Modelling, 165(1):iii, July 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001522.

Anonymous:2004:AIa

[Ano04a] Anonymous. Author index. Ecological Modelling, 171 (4):451–452, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000158.

Anonymous:2004:AIb

[Ano04b] Anonymous. Author index. Ecological Modelling, 172(2–4): 433–434, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380004000432.

Anonymous:2004:AIc

[Ano04c] Anonymous. Author index. Ecological Modelling, 174(4): 433–434, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380004001280.

Anonymous:2004:AId

[Ano04d] Anonymous. Author index. Ecological Modelling, 175 (4):425–426, August 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002145.

Anonymous:2004:AIe

[Ano04e] Anonymous. Author index. Ecological Modelling, 176 (3–4):419–420, September 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002297.

Anonymous:2004:AIf

[Ano04f] Anonymous. Author index. Ecological Modelling, 177 (3–4):427–428, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003060.

Anonymous:2004:AIg

[Ano04g] Anonymous. Author index. Ecological Modelling, 178 (3–4):503–504, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003424.

Anonymous:2004:AIh

[Ano04h] Anonymous. Author index. Ecological Modelling, 179 (4):575–576, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004521.

Anonymous:2004:AIi

[Ano04i] Anonymous. Author index. Ecological Modelling, 180 (4):557–558, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005113.

Anonymous:2004:AIV

[Ano04j] Anonymous. Author index to volume 171. Ecological Modelling, 171(3):315, January 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005040.

Anonymous:2004:CVa

[Ano04k] Anonymous. Contents of volume. Ecological Modelling, 171(3):319–320, January 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003005064.

Anonymous:2004:CVb

[Ano04l] Anonymous. Contents of volume. Ecological Modelling, 171(4):455–456, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000171.

Anonymous:2004:CVc

[Ano04m] Anonymous. Contents of volume. Ecological Modelling, 172(2–4):437–438, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000456.

Anonymous:2004:CVd

[Ano04n] Anonymous. Contents of volume. Ecological Modelling, 173(4):411–412, April 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000833.

Anonymous:2004:CVe

[Ano04o] Anonymous. Contents of volume. Ecological Modelling, 174(4):437–438, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001309.

Anonymous:2004:CVf

[Ano04p] Anonymous. Contents of volume. Ecological Modelling, 175(4):429–430, August 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002169.

Anonymous:2004:CVg

[Ano04q] Anonymous. Contents of volume. Ecological Modelling, 176(3–4):423–424, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002315.

Anonymous:2004:CVh

[Ano04r] Anonymous. Contents of volume. Ecological Modelling, 177(3–4):431–432, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003084.

Anonymous:2004:CVi

[Ano04s] Anonymous. Contents of volume. Ecological Modelling, 179(4):581–582, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004545.

Anonymous:2004:EBa

[Ano04t] Anonymous. Editorial Board. Ecological Modelling, 175 (1):iii, June 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380004001784.

Anonymous:2004:EBb

[Ano04u] Anonymous. Editorial Board. Ecological Modelling, 176(1–2):iii, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002224.

Anonymous:2004:EMR

[Ano04v] Anonymous. Erratum to “Modelling the responses of the Lagoon of Venice ecosystem to variations in physical forcings” [Ecol. Model. 170 (2003) 265–289]. Ecological Modelling, 175(2):195, July 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000663.

Anonymous:2004:ECI

[Ano04w] Anonymous. I. Early and continuing interests: biogeochemistry. Ecological Modelling, 178(1–2):29, October 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004002418.

Anonymous:2004:Ib

[Ano04x] Anonymous. IFC. Ecological Modelling, 174(3):ifc, May 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000400122X.

Anonymous:2004:Ic

[Ano04y] Anonymous. IFC. Ecological Modelling, 174(4):ifc, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004001322.

Anonymous:2004:Id

[Ano04z] Anonymous. IFC. Ecological Modelling, 175(4):ifc, August 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002091.

Anonymous:2004:IE

[Ano04-27] Anonymous. II.Ecosystemsandmetabolism. Ecological Modelling, 178(1–2):41–42, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002431.

Anonymous:2004:IGS

[Ano04-28] Anonymous. III. General systems, ecological modeling and symbols. Ecological Modelling, 178(1–2):81–82, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002443.

Anonymous:2004:IFCa

[Ano04-29] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 171(3):ifc, January 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300499X.

Anonymous:2004:IFCb

[Ano04-30] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 171(4):ifc, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000109.

Anonymous:2004:IFCc

[Ano04-31] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 172(2–4):ifc, March 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000377.

Anonymous:2004:IFCd

[Ano04-32] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 173(2–3):ifc, April 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000687.

Anonymous:2004:IFCe

[Ano04-33] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 173(4):ifc, April 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000766.

Anonymous:2004:IFCf

[Ano04-34] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 175(2):ifc, July 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001905.

Anonymous:2004:IFCg

[Ano04-35] Anonymous. Inside front cover/aims, scope editorial. Ecological Modelling, 175(3):ifc, July 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400198X.

Anonymous:2004:IFCh

[Ano04-36] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 176(3–4):ifc, September 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002339.

Anonymous:2004:IFCi

[Ano04-37] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 177(3–4):ifc, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003102.

Anonymous:2004:IFCj

[Ano04-38] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 178(3–4):ifc, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003370.

Anonymous:2004:IFCk

[Ano04-39] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 179(2):ifc, November 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400417X.

Anonymous:2004:IFCl

[Ano04-40] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 179(3):ifc, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004247.

Anonymous:2004:IFCm

[Ano04-41] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 179(4):ifc, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004569.

Anonymous:2004:IFCn

[Ano04-42] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 180(2–3):ifc, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004880.

Anonymous:2004:IFCo

[Ano04-43] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 180(4):ifc, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400506X.

Anonymous:2004:IA

[Ano04-44] Anonymous. Instructions to authors. Ecological Modelling, 171(3):321–327, January 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005076.

Anonymous:2004:IEE

[Ano04-45] Anonymous. IV. Ecological engineering. Ecological Modelling, 178(1–2):127, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002455.

Anonymous:2004:IPO

[Ano04-46] Anonymous. IX. Philosophical overview of the contributions of H. T. Odum. Ecological Modelling, 178(1–2):251, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002510.

Anonymous:2004:PFb

[Ano04-47] Anonymous. Pages 1–102 (15 February 2004). Ecological Modelling, 172(1):??, February 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PMa

[Ano04-48] Anonymous. Pages 1–108 (15 March 2004). Ecological Modelling, 173(1):??, March 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PJd

[Ano04-49] Anonymous. Pages 1–114 (15 June 2004). Ecological Modelling, 175(1):??, June 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PNc

[Ano04-50] Anonymous. Pages 1–150 (15 November 2004). Ecological Modelling, 179(1):??, November 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PJa

[Ano04-51] Anonymous. Pages 1–208 (1 January 2004). Ecological Modelling, 171(1–2):??, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PAd

[Ano04-52] Anonymous. Pages 1–210 (15 August 2004). Ecological Modelling, 176(1–2):??, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PSb

[Ano04-53] Anonymous. Pages 1–232 (15 September 2004). Ecological Modelling, 177(1–2):??, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PAa

[Ano04-54] Anonymous. Pages 109–312 (1 April 2004). Ecological Modelling, 173(2–3):??, April 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PJe

[Ano04-55] Anonymous. Pages 115–216 (1 July 2004). Ecological Modelling, 175(2):??, July 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PJb

[Ano04-56] Anonymous. Pages 209–328 (15 January 2004). Ecological Modelling, 171(3):??, January 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PSa

[Ano04-57] Anonymous. Pages 211–424 (1 September 2004). Ecological Modelling, 176(3–4):??, September 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PJf

[Ano04-58] Anonymous. Pages 217–318 (15 July 2004). Ecological Modelling, 175(3):??, July 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PMb

[Ano04-59] Anonymous. Pages 225–330 (15 May 2004). Ecological Modelling, 174(3):??, May 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PO

[Ano04-60] Anonymous. Pages 233–432 (1 October 2004). Ecological Modelling, 177(3–4):??, October 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PDb

[Ano04-61] Anonymous. Pages 233–444 (25 December 2004). Ecological Modelling, 180(2–3):??, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PNf

[Ano04-62] Anonymous. Pages 247–416 (30 November 2004). Ecological Modelling, 179(3):??, November 30, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PNb

[Ano04-63] Anonymous. Pages 295–512 (1 November 2004). Ecological Modelling, 178(3–4):??, November 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PAb

[Ano04-64] Anonymous. Pages 313–412 (15 April 2004). Ecological Modelling, 173(4):??, April 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PAc

[Ano04-65] Anonymous. Pages 319–430 (1 August 2004). Ecological Modelling, 175(4):??, August 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PFa

[Ano04-66] Anonymous. Pages 329–456 (1 February 2004). Ecological Modelling, 171(4):??, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PJc

[Ano04-67] Anonymous. Pages 331–438 (1 June 2004). Ecological Modelling, 174(4):??, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PDa

[Ano04-68] Anonymous. Pages 417–582 (1 December 2004). Ecological Modelling, 179(4):??, December 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PDc

[Ano04-69] Anonymous. Pages 445–562 (31 December 2004). Ecological Modelling, 180(4):??, December 31, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2004:PITb

[Ano04-70] Anonymous. Prelim (iii) title page. Ecological Modelling, 172(1):iii, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000304.

Anonymous:2004:PITc

[Ano04-71] Anonymous. Prelim (iii) title page. Ecological Modelling, 173(1):iii, March 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000584.

Anonymous:2004:PITd

[Ano04-72] Anonymous. Prelim (iii) title page. Ecological Modelling, 177(1–2):iii, September 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003023.

Anonymous:2004:PITe

[Ano04-73] Anonymous. Prelim (iii) title page. Ecological Modelling, 178(1–2):iii, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003205.

Anonymous:2004:PITf

[Ano04-74] Anonymous. Prelim (iii) title page. Ecological Modelling, 179(1):iii, November 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003886.

Anonymous:2004:PITg

[Ano04-75] Anonymous. Prelim (iii) title page. Ecological Modelling, 180(1):iii, December 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000400465X.

Anonymous:2004:PI

[Ano04-76] Anonymous. Prelim(iii). Ecological Modelling, 174(1–2):iii, May 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004000894.

Anonymous:2004:PITa

[Ano04-77] Anonymous. Prelim(iii) title page. Ecological Modelling, 171(1–2):iii, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004927.

Anonymous:2004:PNa

[Ano04-78] Anonymous. Publisher’s note. Ecological Modelling, 171 (1–2):1, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380003004964.

Anonymous:2004:PNd

[Ano04-79] Anonymous. Publisher’s note. Ecological Modelling, 179(1):1, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000400273X.

Anonymous:2004:PNe

[Ano04-80] Anonymous. Publisher’s note. Ecological Modelling, 179(1):149, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002285.

Anonymous:2004:SIa

[Ano04-81] Anonymous. Subject index. Ecological Modelling, 171 (4):453–454, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000400016X.

Anonymous:2004:SIb

[Ano04-82]  Anonymous.  Subject index.  Ecological Modelling,

172(2–4):435–436, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000444.

Anonymous:2004:SIc

[Ano04-83]  Anonymous.  Subject index.  Ecological Modelling,

173(4):409–410, April 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000821.

Anonymous:2004:SId

[Ano04-84]  Anonymous.  Subject index.  Ecological Modelling,

174(4):435–436, June 1, 2004. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001292.

Anonymous:2004:SIe

[Ano04-85] Anonymous. Subject index. Ecological Modelling, 175 (4):427–428, August 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002157.

Anonymous:2004:SIf

[Ano04-86] Anonymous. Subject index. Ecological Modelling, 176 (3–4):421–422, September 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002303.

Anonymous:2004:SIg

[Ano04-87] Anonymous. Subject index. Ecological Modelling, 177 (3–4):429–430, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003072.

Anonymous:2004:SIh

[Ano04-88] Anonymous. Subject index. Ecological Modelling, 178 (3–4):505–506, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003436.

Anonymous:2004:SIi

[Ano04-89] Anonymous. Subject index. Ecological Modelling, 179 (4):577–579, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004533.

Anonymous:2004:SIj

[Ano04-90] Anonymous. Subject index. Ecological Modelling, 180 (4):559–560, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005125.

Anonymous:2004:SIV

[Ano04-91] Anonymous. Subject index to volume 171. Ecological Modelling, 171(3):317–318, January 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005052.

Anonymous:2004:TCa

[Ano04-92] Anonymous. Table of contents. Ecological Modelling, 178 (3–4):507–510, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003448.

Anonymous:2004:TCb

[Ano04-93] Anonymous. Table of contents. Ecological Modelling, 180(4):561–562, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005137.

Anonymous:2004:VES

[Ano04-94] Anonymous. V. Environment and society. Ecological Modelling, 178(1–2):157, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002467.

Anonymous:2004:VEE

[Ano04-95] Anonymous. VI. Ecology and economics. Ecological Modelling, 178(1–2):177, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002479.

Anonymous:2004:VEA

[Ano04-96] Anonymous. VII. Emergy analysis. Ecological Modelling, 178(1–2):199, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002480.

Anonymous:2004:VPW

[Ano04-97] Anonymous. VIII. A prosperous way down. Ecological Modelling, 178(1–2):241, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002492.

Anonymous:2005:AIa

[Ano05a] Anonymous. Author index. Ecological Modelling, 181 (4):615–616, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005344.

Anonymous:2005:AIb

[Ano05b] Anonymous. Author index. Ecological Modelling, 182(3–4): 379–380, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380005000505.

Anonymous:2005:AIc

[Ano05c] Anonymous. Author index. Ecological Modelling, 183(4): 495–496, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380005001055.

Anonymous:2005:AId

[Ano05d] Anonymous. Author index. Ecological Modelling, 184 (2–4):413–414, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001584.

Anonymous:2005:AIe

[Ano05e] Anonymous. Author index. Ecological Modelling, 185 (2–4):551–552, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001833.

Anonymous:2005:AIf

[Ano05f] Anonymous. Author index. Ecological Modelling, 186(4): 502–503, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380005003212.

Anonymous:2005:AIVa

[Ano05g] Anonymous. Author index of volume 187. Ecological Modelling, 187(4):541–542, October 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004187.

Anonymous:2005:AIVb

[Ano05h] Anonymous. Author index of volume 188. Ecological Modelling, 188(2–4):494–495, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004722.

Anonymous:2005:AIVc

[Ano05i] Anonymous. Author index of volume 189. Ecological Modelling, 189(3–4):454–455, December 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500534X.

Anonymous:2005:CC

[Ano05j] Anonymous. Contents continued. Ecological Modelling, 188(2–4):501, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004862.

Anonymous:2005:CVa

[Ano05k] Anonymous. Contents of volume. Ecological Modelling, 182(3–4):383–384, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000529.

Anonymous:2005:CVb

[Ano05l] Anonymous. Contents of volume. Ecological Modelling, 185(2–4):557–558, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001857.

Anonymous:2005:CVc

[Ano05m] Anonymous. Contents of volume. Ecological Modelling, 186(4):507–509, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003236.

Anonymous:2005:CVd

[Ano05n] Anonymous. Contents of volume 187. Ecological Modelling, 187(4):546–547, October 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004205.

Anonymous:2005:CVe

[Ano05o] Anonymous. Contents of volume 188. Ecological Modelling, 188(2–4):499–500, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004746.

Anonymous:2005:CVf

[Ano05p] Anonymous. Contents of volume 189. Ecological Modelling, 189(3–4):459–460, December 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005363.

Anonymous:2005:GB

[Ano05q] Anonymous. Giuseppe Bendoricchio. Ecological Modelling, 184(1):1–3, May 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005794.

Anonymous:2005:IFCa

[Ano05r] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 181(2–3):ifc, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400523X.

Anonymous:2005:IFCb

[Ano05s] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 181(4):ifc, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005381.

Anonymous:2005:IFCc

[Ano05t] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 182(3–4):ifc, March 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500044X.

Anonymous:2005:IFCd

[Ano05u] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 183(2–3):ifc, April 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000943.

Anonymous:2005:IFCe

[Ano05v] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 183(4):ifc, May 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001092.

Anonymous:2005:IFCf

[Ano05w] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 184(2–4):ifc, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001535.

Anonymous:2005:IFCg

[Ano05x] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 185(2–4):CO2, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500178X.

Anonymous:2005:IFCh

[Ano05y] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 186(2):CO2, August 15, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002942.

Anonymous:2005:IFCi

[Ano05z] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 186(3):CO2, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003078.

Anonymous:2005:IFCj

[Ano05-27] Anonymous. Inside front cover/aims, scope, editorial. Ecological Modelling, 186(4):CO2, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003169.

Anonymous:2005:PFb

[Ano05-28] Anonymous. Pages 1–106 (25 February 2005). Ecological Modelling, 182(1):??, February 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PAa

[Ano05-29] Anonymous. Pages 1–146 (10 April 2005). Ecological Modelling, 183(1):??, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PJd

[Ano05-30] Anonymous. Pages 1–152 (10 June 2005). Ecological Modelling, 185(1):??, June 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PNc

[Ano05-31] Anonymous. Pages 1–240 (25 November 2005). Ecological Modelling, 189(1–2):??, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PJa

[Ano05-32] Anonymous. Pages 1–92 (10 January 2005). Ecological Modelling, 181(1):??, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PMa

[Ano05-33] Anonymous. Pages 107–216 (10 March 2005). Ecological Modelling, 182(2):??, March 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PNb

[Ano05-34] Anonymous. Pages 137–502 (10 November 2005). Ecological Modelling, 188(2–4):??, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PAc

[Ano05-35] Anonymous. Pages 143–270 (15 August 2005). Ecological Modelling, 186(2):??, August 15, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PAb

[Ano05-36] Anonymous. Pages 147–364 (25 April 2005). Ecological Modelling, 183(2–3):??, April 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PJe

[Ano05-37] Anonymous. Pages 153–558 (10 July 2005). Ecological Modelling, 185(2–4):??, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PJc

[Ano05-38] Anonymous. Pages 175–420 (1 June 2005). Ecological Modelling, 184(2–4):??, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PD

[Ano05-39] Anonymous. Pages 241–460 (10 December 2005). Ecological Modelling, 189(3–4):??, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PAd

[Ano05-40] Anonymous. Pages 271–374 (25 August 2005). Ecological Modelling, 186(3):??, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PMb

[Ano05-41] Anonymous. Pages 365–500 (10 May 2005). Ecological Modelling, 183(4):??, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PO

[Ano05-42] Anonymous. Pages 369–548 (10 October 2005). Ecological Modelling, 187(4):??, October 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PFa

[Ano05-43] Anonymous. Pages 445–620 (10 February 2005). Ecological Modelling, 181(4):??, February 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PJb

[Ano05-44] Anonymous. Pages 93–286 (20 January 2005). Ecological Modelling, 181(2–3):??, January 20, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PS

[Ano05-45] Anonymous. Pages 99–368 (25 September 2005). Ecological Modelling, 187(2–3):??, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2005:PITa

[Ano05-46] Anonymous. Prelim (iii) title page. Ecological Modelling, 181(1):iii, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005186.

Anonymous:2005:PITb

[Ano05-47] Anonymous. Prelim (iii) title page. Ecological Modelling, 182(1):iii, February 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005691.

Anonymous:2005:PITc

[Ano05-48] Anonymous. Prelim (iii) title page. Ecological Modelling, 183(1):iii, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000694.

Anonymous:2005:PITd

[Ano05-49] Anonymous. Prelim (iii) title page. Ecological Modelling, 184(1):iii, May 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380005001286.

Anonymous:2005:PITe

[Ano05-50] Anonymous. Prelim (iii) title page. Ecological Modelling, 185(1):iii, June 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001717.

Anonymous:2005:PITf

[Ano05-51] Anonymous. Prelim (iii) title page. Ecological Modelling, 186(1):iii, July 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380005002498.

Anonymous:2005:PITg

[Ano05-52] Anonymous. Prelim (iii) title page. Ecological Modelling, 187(1):iii, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003650.

Anonymous:2005:PITh

[Ano05-53] Anonymous. Prelim (iii) title page. Ecological Modelling, 188(1):iii, October 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004345.

Anonymous:2005:PITi

[Ano05-54] Anonymous. Prelim (iii) title page. Ecological Modelling, 189(1–2):iii, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005065.

Anonymous:2005:PAe

[Ano05-55] Anonymous. Publisher’s acknowlodgement. Ecological Modelling, 187(4):537–540, October 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000888.

Anonymous:2005:PNa

[Ano05-56] Anonymous. Publishers note. Ecological Modelling, 184 (2–4):175, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380005001614.

Anonymous:2005:SIa

[Ano05-57] Anonymous. Subject index. Ecological Modelling, 181 (4):617–618, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005356.

Anonymous:2005:SIb

[Ano05-58]  Anonymous.  Subject index.  Ecological Modelling,

182(3–4):381–382, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000517.

Anonymous:2005:SIc

[Ano05-59]  Anonymous.  Subject index.  Ecological Modelling,

183(4):497–498, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001067.

Anonymous:2005:SId

[Ano05-60]  Anonymous.  Subject index.  Ecological Modelling,

184(2–4):415–416, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001596.

Anonymous:2005:SIe

[Ano05-61] Anonymous. Subject index. Ecological Modelling, 185 (2–4):553–555, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001845.

Anonymous:2005:SIf

[Ano05-62]  Anonymous.  Subject index.  Ecological Modelling,

186(4):504–506, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003224.

Anonymous:2005:SIVa

[Ano05-63] Anonymous. Subject index of volume 187. Ecological Modelling, 187(4):543–545, October 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004199.

Anonymous:2005:SIVb

[Ano05-64] Anonymous. Subject index of volume 188. Ecological Modelling, 188(2–4):496–498, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004734.

Anonymous:2005:SIVc

[Ano05-65] Anonymous. Subject index of volume 189. Ecological Modelling, 189(3–4):456–458, December 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005351.

Anonymous:2005:TCa

[Ano05-66] Anonymous. Table of contents. Ecological Modelling, 181(4):619–620, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005368.

Anonymous:2005:TCb

[Ano05-67] Anonymous. Table of contents. Ecological Modelling, 183(4):499–500, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001079.

Anonymous:2005:TCc

[Ano05-68] Anonymous. Table of contents. Ecological Modelling, 184(2–4):417–418, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001602.

Anonymous:2006:AIVa

[Ano06a] Anonymous. Author index of volume 190. Ecological Modelling, 190(3–4):454–455, January 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005594.

Anonymous:2006:AIVb

[Ano06b] Anonymous. Author index of volume 191. Ecological Modelling, 191(3–4):557–558, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006356.

Anonymous:2006:AIVc

[Ano06c] Anonymous. Author index of volume 192. Ecological Modelling, 192(3–4):667–668, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600024X.

Anonymous:2006:AIVd

[Ano06d] Anonymous. Author index of volume 193. Ecological Modelling, 193(3–4):815–817, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000469.

Anonymous:2006:AIVe

[Ano06e] Anonymous. Author index of volume 195. Ecological Modelling, 195(3–4):411–412, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001918.

Anonymous:2006:AIVf

[Ano06f] Anonymous. Author index of volume 198. Ecological Modelling, 198(3–4):525–526, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003978.

Anonymous:2006:AIVg

[Ano06g] Anonymous. Author index of volume 199. Ecological Modelling, 199(4):517–518, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005163.

Anonymous:2006:CCa

[Ano06h] Anonymous. Contents continued. Ecological Modelling, 191(3–4):EX4, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006000044.

Anonymous:2006:CCb

[Ano06i] Anonymous. Contents (continued). Ecological Modelling, 192(1–2):325, February 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000159.

Anonymous:2006:CVa

[Ano06j] Anonymous. Contents of volume 190. Ecological Modelling, 190(3–4):458–459, January 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005612.

Anonymous:2006:CVb

[Ano06k] Anonymous. Contents of volume 191. Ecological Modelling, 191(3–4):562–564, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500637X.

Anonymous:2006:CVc

[Ano06l] Anonymous. Contents of volume 192. Ecological Modelling, 192(3–4):672–674, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000263.

Anonymous:2006:CVd

[Ano06m] Anonymous. Contents of volume 193. Ecological Modelling, 193(3–4):822–824, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000482.

Anonymous:2006:CVe

[Ano06n] Anonymous. Contents of volume 195. Ecological Modelling, 195(3–4):416–417, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001931.

Anonymous:2006:CVf

[Ano06o] Anonymous. Contents of volume 198. Ecological Modelling, 198(3–4):530–532, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003991.

Anonymous:2006:CVg

[Ano06p] Anonymous. Contents of volume 199. Ecological Modelling, 199(4):522–524, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005187.

Anonymous:2006:EABa

[Ano06q] Anonymous. Editorial Advisory Board. Ecological Modelling, 195(1–2):iii, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001578.

Anonymous:2006:EABb

[Ano06r] Anonymous. Editorial Advisory Board. Ecological Modelling, 196(1–2):iii, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006002249.

Anonymous:2006:EABc

[Ano06s] Anonymous. Editorial Advisory Board. Ecological Modelling, 197(1–2):iii, August 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600305X.

Anonymous:2006:EABd

[Ano06t] Anonymous. Editorial Advisory Board. Ecological Modelling, 198(1–2):iii, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003735.

Anonymous:2006:EABe

[Ano06u] Anonymous. Editorial Advisory Board. Ecological Modelling, 199(1):iii, November 1, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004376.

Anonymous:2006:OBC

[Ano06v] Anonymous. Outside back cover-contents. Ecological Modelling, 191(3–4):CO4, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006393.

Anonymous:2006:PN

[Ano06w] Anonymous. Pages 1–128 (1 November 2006). Ecological Modelling, 199(1):??, November 1, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PJa

[Ano06x] Anonymous. Pages 1–230 (10 January 2006). Ecological Modelling, 190(1–2):??, January 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PAb

[Ano06y] Anonymous. Pages 1–262 (10 August 2006). Ecological Modelling, 197(1–2):??, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PJe

[Ano06z] Anonymous. Pages 1–274 (10 July 2006). Ecological Modelling, 196(1–2):??, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PS

[Ano06-27] Anonymous. Pages 1–276 (15 September 2006). Ecological Modelling, 198(1–2):??, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PFb

[Ano06-28] Anonymous. Pages 1–326 (15 February 2006). Ecological Modelling, 192(1–2):??, February 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PM

[Ano06-29] Anonymous. Pages 141–824 (15 March 2006). Ecological Modelling, 193(3–4):??, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PJd

[Ano06-30] Anonymous. Pages 153–418 (15 June 2006). Ecological Modelling, 195(3–4):??, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PJc

[Ano06-31] Anonymous. Pages 197–314 (31 January 2006). Ecological Modelling, 191(2):??, January 31, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PJb

[Ano06-32] Anonymous. Pages 231–460 (25 January 2006). Ecological Modelling, 190(3–4):??, January 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PAc

[Ano06-33] Anonymous. Pages 263–528 (25 August 2006). Ecological Modelling, 197(3–4):??, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PJf

[Ano06-34] Anonymous. Pages 275–548 (25 July 2006). Ecological Modelling, 196(3–4):??, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PO

[Ano06-35] Anonymous. Pages 277–534 (15 October 2006). Ecological Modelling, 198(3–4):??, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PFa

[Ano06-36] Anonymous. Pages 315–564 (5 February 2006). Ecological Modelling, 191(3–4):??, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PFc

[Ano06-37] Anonymous. Pages 327–674 (25 February 2006). Ecological Modelling, 192(3–4):??, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PAa

[Ano06-38] Anonymous. Pages 329–434 (15 April 2006). Ecological Modelling, 194(4):??, April 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:PITa

[Ano06-39] Anonymous. Prelim (iii) title page. Ecological Modelling, 190(1–2):iii, January 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005491.

Anonymous:2006:PITb

[Ano06-40] Anonymous. Prelim (iii) title page. Ecological Modelling, 191(1):iii, January 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005005910.

Anonymous:2006:PITd

[Ano06-41] Anonymous. Prelim (iii) title page. Ecological Modelling, 192(1–2):iii, February 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000135.

Anonymous:2006:PITe

[Ano06-42] Anonymous. Prelim (iii) title page journal. Ecological Modelling, 193(1–2):iii, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000329.

Anonymous:2006:PITf

[Ano06-43] Anonymous. Prelim (iii) title page journal. Ecological Modelling, 194(1–3):iii, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000834.

Anonymous:2006:PITc

[Ano06-44] Anonymous. Prelim (iii) title page special issue. Ecological Modelling, 191(1):v, January 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005879.

Anonymous:2006:PCO

[Ano06-45] Anonymous. Proceedings of a conference organized by the International Society for Ecological Modelling (ISEM) in Quebec, Canada, August 22–24, 2004. Ecological Modelling, 199(3):??, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:RSW

[Ano06-46] Anonymous. Results from a Second Workshop on Advances in Predictive Species Distribution Models, held in Riederalp, Switzerland, 2004. Ecological Modelling, 199(2):??, November 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:SPC

[Ano06-47] Anonymous. Selected papers from the Conference on Southern European Coastal Lagoons: the Influence of River Basin–Coastal Zone Interactions, Ferrara, Italy, 10– 12 November 2003. Ecological Modelling, 193(1–2):??, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:SPF

[Ano06-48] Anonymous. Selected papers from the Fourth European Conference on Ecological Modelling, September 27–October 1, 2004, Bled, Slovenia. Ecological Modelling, 194(1–3):??, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2006:SIVa

[Ano06-49] Anonymous. Subject index of volume 190. Ecological Modelling, 190(3–4):456–457, January 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005600.

Anonymous:2006:SIVb

[Ano06-50] Anonymous. Subject index of volume 191. Ecological Modelling, 191(3–4):559–561, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006368.

Anonymous:2006:SIVc

[Ano06-51] Anonymous. Subject index of volume 192. Ecological Modelling, 192(3–4):669–671, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000251.

Anonymous:2006:SIVd

[Ano06-52] Anonymous. Subject index of volume 193. Ecological Modelling, 193(3–4):818–821, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000470.

Anonymous:2006:SIVe

[Ano06-53] Anonymous. Subject index of volume 195. Ecological Modelling, 195(3–4):413–415, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600192X.

Anonymous:2006:SIVf

[Ano06-54] Anonymous. Subject index of volume 198. Ecological Modelling, 198(3–4):527–529, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600398X.

Anonymous:2006:SIVg

[Ano06-55] Anonymous. Subject index of volume 199. Ecological Modelling, 199(4):519–521, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005175.

Anonymous:2007:CIS

[Ano07a] Anonymous. 4th Conference of the International Society for Ecological Informatics. Ecological Modelling, 203(1–2):??, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:EABa

[Ano07b] Anonymous. Editorial Advisory Board. Ecological Modelling, 200(1–2):iii, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600593X.

Anonymous:2007:EABb

[Ano07c] Anonymous. Editorial Advisory Board. Ecological Modelling, 201(1):iii, February 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006006533.

Anonymous:2007:EABc

[Ano07d] Anonymous. Editorial Advisory Board. Ecological Modelling, 203(1–2):iii, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001366.

Anonymous:2007:EBa

[Ano07e] Anonymous. Editorial Board. Ecological Modelling, 202 (1–2):iii, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007000737.

Anonymous:2007:EBb

[Ano07f] Anonymous. Editorial Board. Ecological Modelling, 204 (1–2):iii, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007002153.

Anonymous:2007:EBc

[Ano07g] Anonymous. Editorial Board. Ecological Modelling, 205 (1–2):iii, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007002700.

Anonymous:2007:EBd

[Ano07h] Anonymous. Editorial Board. Ecological Modelling, 206(1–2):iii, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007003328.

Anonymous:2007:EBe

[Ano07i] Anonymous. Editorial Board. Ecological Modelling, 207 (1):iii, September 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007003936.

Anonymous:2007:EBf

[Ano07j] Anonymous. Editorial Board. Ecological Modelling, 208 (1):iii, October 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007004681.

Anonymous:2007:EBg

[Ano07k] Anonymous. Editorial Board. Ecological Modelling, 209 (1):iii, November 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007005492.

Anonymous:2007:MNP

[Ano07l] Anonymous. Modeling of North Pacific marine ecosystems. Ecological Modelling, 202(1–2):??, March 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PAb

[Ano07m] Anonymous. Pages 1–212 (10 August 2007). Ecological Modelling, 206(1–2):??, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PJa

[Ano07n] Anonymous. Pages 1–276 (10 January 2007). Ecological Modelling, 200(1–2):??, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PJd

[Ano07o] Anonymous. Pages 1–276 (10 July 2007). Ecological Modelling, 205(1–2):??, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PMc

[Ano07p] Anonymous. Pages 1–278 (24 May 2007). Ecological Modelling, 204(1–2):??, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PMb

[Ano07q] Anonymous. Pages 185–532 (10 May 2007). Ecological Modelling, 203(3–4):??, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PAc

[Ano07r] Anonymous. Pages 213–438 (24 August 2007). Ecological Modelling, 206(3–4):??, August 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PAa

[Ano07s] Anonymous. Pages 225–560 (10 April 2007). Ecological Modelling, 202(3–4):??, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PMa

[Ano07t] Anonymous. Pages 243–564 (10 March 2007). Ecological Modelling, 201(3–4):??, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PJe

[Ano07u] Anonymous. Pages 277–526 (24 July 2007). Ecological Modelling, 205(3–4):??, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PJb

[Ano07v] Anonymous. Pages 277–530 (24 January 2007). Ecological Modelling, 200(3–4):??, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PJc

[Ano07w] Anonymous. Pages 279–558 (16 June 2007). Ecological Modelling, 204(3–4):??, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PO

[Ano07x] Anonymous. Pages 61–368 (10 October 2007). Ecological Modelling, 207(2–4):??, October 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PD

[Ano07y] Anonymous. Pages 65–416 (16 December 2007). Ecological Modelling, 209(2–4):??, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PN

[Ano07z] Anonymous. Pages 91–396 (10 November 2007). Ecological Modelling, 208(2–4):??, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2007:PF

[Ano07-27] Anonymous. Pages 97–242 (24 February 2007). Ecological Modelling, 201(2):??, February 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:SPW

[Ano08a] Anonymous. Selected Papers from a Workshop Organized by the International Society for Ecological Modelling (ISEM) at the Third Biennal Meeting of the International Environmental Modelling and Software Society (IEMSS) in Burlington, Vermont, USA, August 9–13, 2006. Ecological Modelling, 219(3–4):??, December 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:EBa

[Ano08b] Anonymous. Editorial Board. Ecological Modelling, 210(1–2):iii, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007005947.

Anonymous:2008:EBb

[Ano08c] Anonymous. Editorial Board. Ecological Modelling, 211(1–2):iii, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007006631.

Anonymous:2008:EBc

[Ano08d] Anonymous. Editorial Board. Ecological Modelling, 212 (1–2):i, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380008000471.

Anonymous:2008:EBd

[Ano08e] Anonymous. Editorial Board. Ecological Modelling, 213(1):i, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008001178.

Anonymous:2008:EBe

[Ano08f] Anonymous. Editorial Board. Ecological Modelling, 214(1): i, June 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008001920.

Anonymous:2008:EBf

[Ano08g] Anonymous. Editorial Board. Ecological Modelling, 216 (1):iii, August 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380008002846.

Anonymous:2008:EBg

[Ano08h] Anonymous. Editorial Board. Ecological Modelling, 217(1–2):i, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000800375X.

Anonymous:2008:EBh

[Ano08i] Anonymous. Editorial Board. Ecological Modelling, 219(1–2):i, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008004729.

Anonymous:2008:PAb

[Ano08j] Anonymous. Pages 1–142 (24 April 2008). Ecological Modelling, 213(1):??, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PO

[Ano08k] Anonymous. Pages 1–194 (24 October 2008). Ecological Modelling, 218(1–2):??, October 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PSb

[Ano08l] Anonymous. Pages 1–206 (24 September 2008). Ecological Modelling, 217(1–2):??, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PJa

[Ano08m] Anonymous. Pages 1–220 (10 January 2008). Ecological Modelling, 210(1–2):??, January 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PFb

[Ano08n] Anonymous. Pages 1–250 (24 February 2008). Ecological Modelling, 211(1–2):??, February 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PNb

[Ano08o] Anonymous. Pages 1–260 (24 November 2008). Ecological Modelling, 219(1–2):??, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PAc

[Ano08p] Anonymous. Pages 1–80 (10 August 2008). Ecological Modelling, 216(1):??, August 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PMb

[Ano08q] Anonymous. Pages 143–272 (10 May 2008). Ecological Modelling, 213(2):??, May 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PAa

[Ano08r] Anonymous. Pages 179–550 (10 April 2008). Ecological Modelling, 212(3–4):??, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PNa

[Ano08s] Anonymous. Pages 195–404 (10 November 2008). Ecological Modelling, 218(3–4):??, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PJb

[Ano08t] Anonymous. Pages 221–350 (24 January 2008). Ecological Modelling, 210(3):??, January 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PMa

[Ano08u] Anonymous. Pages 251–500 (10 March 2008). Ecological Modelling, 211(3–4):??, March 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PSa

[Ano08v] Anonymous. Pages 253–416 (10 September 2008). Ecological Modelling, 216(3–4):??, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PJd

[Ano08w] Anonymous. Pages 273–392 (24 July 2008). Ecological Modelling, 215(4):??, July 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PMc

[Ano08x] Anonymous. Pages 273–470 (24 May 2008). Ecological Modelling, 213(3–4):??, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PFa

[Ano08y] Anonymous. Pages 351–498 (10 February 2008). Ecological Modelling, 210(4):??, February 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:PJc

[Ano08z] Anonymous. Pages 75–420 (24 June 2008). Ecological Modelling, 214(2–4):??, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2008:YMS

[Ano08-27] Anonymous. Yuri Mikhailovich Svirezhev (22 September 1938–22 February 2007). Ecological Modelling, 212(1–2):3– 4, March 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005194.

Anonymous:2009:Aa

[Ano09a] Anonymous. Announcement. Ecological Modelling, 220(6): 755, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380009001008.

Anonymous:2009:Ab

[Ano09b]  Anonymous.  Announcement.  Ecological Modelling,

220(9–10):1181, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000900204X.

Anonymous:2009:AICa

[Ano09c] Anonymous. Announcement: ISEM 2009 conference. Ecological Modelling, 220(5):595, March 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900043X.

Anonymous:2009:AICb

[Ano09d] Anonymous. Announcement: ISEM 2009 conference. Ecological Modelling, 220(7):897, April 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001276.

Anonymous:2009:AICc

[Ano09e] Anonymous. Announcement: ISEM 2009 conference. Ecological Modelling, 220(8):1043, April 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001586.

Anonymous:2009:AICd

[Ano09f] Anonymous. Announcement: ISEM 2009 conference. Ecological Modelling, 220(11):1375, June 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002361.

Anonymous:2009:CU

[Ano09g] Anonymous. Conference update. Ecological Modelling, 220(17):1933–1934, September 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004773.

Anonymous:2009:EB

[Ano09h] Anonymous. Editorial Board. Ecological Modelling, 220 (1):i, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380008005292.

Anonymous:2009:ISE

[Ano09i] Anonymous. International Society for Ecological Modelling Conference (ISEM 2009): Ecological modelling for enhanced sustainability in management. Ecological Modelling, 220(1):7, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008005036.

Anonymous:2009:IAa

[Ano09j] Anonymous. ISEM 2009 annnouncement. Ecological Modelling, 220(2):85, January 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005498.

Anonymous:2009:IAb

[Ano09k] Anonymous. ISEM 2009 announcement. Ecological Modelling, 220(3):265, February 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005991.

Anonymous:2009:IAc

[Ano09l] Anonymous. ISEM 2009 announcement. Ecological Modelling, 220(4):437, February 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008006005.

Anonymous:2009:PJa

[Ano09m] Anonymous. Pages 1–84 (10 January 2009). Ecological Modelling, 220(1):??, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PAb

[Ano09n] Anonymous. Pages 1043–1180 (24 April 2009). Ecological Modelling, 220(8):??, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PMc

[Ano09o] Anonymous. Pages 1181–1374 (17 May 2009). Ecological Modelling, 220(9–10):??, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PJc

[Ano09p] Anonymous. Pages 1375–1460 (10 June 2009). Ecological Modelling, 220(11):??, June 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PJd

[Ano09q] Anonymous. Pages 1461–1550 (24 June 2009). Ecological Modelling, 220(12):??, June 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PJe

[Ano09r] Anonymous. Pages 1551–1734 (17 July 2009). Ecological Modelling, 220(13–14):??, July 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PAc

[Ano09s] Anonymous. Pages 1735–1854 (10 August 2009). Ecological Modelling, 220(15):??, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PSa

[Ano09t] Anonymous. Pages 1933–2120 (10 September 2009). Ecological Modelling, 220(17):??, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PSb

[Ano09u] Anonymous. Pages 2121–2348 (24 September 2009). Ecological Modelling, 220(18):??, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:POa

[Ano09v] Anonymous. Pages 2349–2602 (10 October 2009). Ecological Modelling, 220(19):??, October 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:POb

[Ano09w] Anonymous. Pages 2603–2824 (24 October 2009). Ecological Modelling, 220(20):??, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PFa

[Ano09x] Anonymous. Pages 265–436 (10 February 2009). Ecological Modelling, 220(3):??, February 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PD

[Ano09y] Anonymous. Pages 3241–3476 (10 December 2009). Ecological Modelling, 220(23):??, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PFb

[Ano09z] Anonymous. Pages 437–594 (24 February 2009). Ecological Modelling, 220(4):??, February 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PMa

[Ano09-27] Anonymous. Pages 595–754 (10 March 2009). Ecological Modelling, 220(5):??, March 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PMb

[Ano09-28] Anonymous. Pages 755–896 (24 March 2009). Ecological Modelling, 220(6):??, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PJb

[Ano09-29] Anonymous. Pages 85–264 (24 January 2009). Ecological Modelling, 220(2):??, January 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:PAa

[Ano09-30] Anonymous. Pages 897–1042 (10 April 2009). Ecological Modelling, 220(7):??, April 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2009:SAIa

[Ano09-31] Anonymous. Second announcement: ISEM 2009 Conference. Ecological Modelling, 220(13–14):1551, July 17, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000900338X.

Anonymous:2009:SAIb

[Ano09-32] Anonymous. Second announcement: ISEM 2009 Conference. Ecological Modelling, 220(16):1931–1932, August 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000900430X.

Anonymous:2009:SPT

[Ano09-33] Anonymous. Selected papers from the Third Workshop on Ecological Network Analysis, University of Georgia, Athens, GA, USA, April 2008. Ecological Modelling, 220(22):??, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anand:2000:HPE

[AO00] Madhur Anand and László Orlóci. On hierarchical partitioning of an ecological complexity function. Ecological Modelling, 132(1–2):51–61, July 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003045.

Albrecht:2008:ECG

[AO08] K.-F. Albrecht and D. Orlamünder. Electricity consumption: the growth pattern as an ecological indicator. Ecological Modelling, 216(2):127–133, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001312.

Altunkaynak:2005:FLM

[AÖÇ05] Abdüsselam Altunkaynak, Mehmet Özger, and Mehmet Çakmakcı. Fuzzy logic modeling of the dissolved oxygen fluctuations in Golden Horn. Ecological Modelling, 189(3–4):436–446, December 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002036.

Aoki:2008:ELA

[Aok08] Ichiro Aoki. Entropy law in aquatic communities and the general entropy principle for the development of living systems. Ecological Modelling, 215(1–3):89–92, July 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008000884.

Alexandrov:2002:SGP

[AOY02] G. A. Alexandrov, T. Oikawa, and Y. Yamagata. The scheme for globalization of a process-based model explaining gradations in terrestrial NPP and its application. Ecological Modelling, 148(3):293–306, March 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004562.

Angelini:2000:MPS

[AP00] Ronaldo Angelini and Miguel Petrere. A model for the plankton system of the Broa reservoir, São Carlos, Brazil. Ecological Modelling, 126(2–3):131–137, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002611.

An:2002:IIS

[AP02] Kwang-Guk An and Seok Soon Park. Indirect influence of the summer monsoon on chlorophyll–total phosphorus models in reservoirs: a case study. Ecological Modelling, 152(2–3):191–203, July 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000200.

Arii:2006:ECP

[AP06] Ken Arii and Lael Parrott. Examining the colonization process of exotic species varying in competitive abilities using a cellular automaton model. Ecological Modelling, 199(3):219–228, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002158.

Amatulli:2007:MLH

[APCdlR07] Giuseppe Amatulli, Fernando Peréz-Cabello, and Juan de la Riva. Mapping lightning/human-caused wildfires occurrence under ignition point location uncertainty. Ecological Modelling, 200(3–4):321–333, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003772.

Alados:2009:ESP

[APEA09] Concepción L. Alados, Yolanda Pueyo, Juan Escós, and Antonio Andujar. Effects of the spatial pattern of disturbance on the patch-occupancy dynamics of juniper–pine open woodland. Ecological Modelling, 220(12):1544–1550, June 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002282.

Alados:2003:QCR

[APG^{+}03] C. L. Alados, Y. Pueyo, M. L. Giner, T. Navarro, J. Escos, F. Barroso, B. Cabezudo, and J. M. Emlen. Quantitative characterization of the regressive ecological succession by fractal analysis of plant spatial patterns. Ecological Modelling, 163(1–2):1–17, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002946.

Ainsworth:2008:RHM

[APHV08] C. H. Ainsworth, T. J. Pitcher, J. J. Heymans, and M. Vasconcellos. Reconstructing historical marine ecosystems using food web models: Northern British Columbia from pre-European contact to present. Ecological Modelling, 216(3–4):354–368, September 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002299.

Antonic:2003:SDM

[APJ03] Oleg Antonić, Nikola Pernar, and Sven D. Jelaska. Spatial distribution of main forest soil groups in Croatia as a function of basic pedogenetic factors. Ecological Modelling, 170(2–3):363–371, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002394.

Arhonditsis:2007:ERA

[AQS^{+}07] George B. Arhonditsis, Song S. Qian, Craig A. Stow, E. Conrad Lamon, and Kenneth H. Reckhow. Eutrophication risk assessment using Bayesian calibration of process-based models: Application to a mesotrophic lake. Ecological Modelling, 208(2–4):215–229, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003055.

Acevedo:2002:GDT

[AR02] Miguel F. Acevedo and Josep Raventós. Growth dynamics of three tropical savanna grass species: an individualmodule model. Ecological Modelling, 154(1–2):45–60, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002000625.

Asaeda:2008:AOS

[ARF08] Takashi Asaeda, Lalith Rajapakse, and Takeshi Fujino. Applications of organ-specific growth models; modelling of resource translocation and the role of emergent aquatic plants in element cycles. Ecological Modelling, 215(1–3):170–179, July 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008001002.

Arino:2000:P

[Ari00] O. Arino. Preface. Ecological Modelling, 133(1–2):1, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002891.

Adiku:2006:PMS

[ARL^{+}06] S. G. K. Adiku, M. Reichstein, A. Lohila, N. Q. Dinh, M. Aurela, T. Laurila, J. Lueers, and J. D. Tenhunen. PIXGRO: a model for simulating the ecosystem CO_{2} exchange and growth of spring barley. Ecological Modelling, 190(3–4):260–276, January 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002656.

Armstrong:2004:BEE

[Arm04] Neal E. Armstrong. The beginnings of ecological engineering. Ecological Modelling, 178(1–2):129–131, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005155.

Arundel:2005:USM

[Aru05] Samantha T. Arundel. Using spatial models to establish climatic limiters of plant species’ distributions. Ecological Modelling, 182(2):159–181, March 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004119.

Alonso:2000:PSS

[AS00a] David Alonso and Ricard V. Solé. The =pkgDivGame Simulator: a stochastic cellular automata model of rainforest dynamics. Ecological Modelling, 133(1–2):131–141, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002878.

Arreguin-Sanchez:2000:ORG

[AS00b] Francisco Arreguín-Sánchez. Octopus–red grouper interaction in the exploited ecosystem of the northern continental shelf of Yucatan, Mexico. Ecological Modelling, 129(2–3):119–129, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002180.

Asaeda:2000:SSP

[AS00c] Takashi Asaeda and Duong Hong Son. Spatial structure and populations of a periphyton community: a model and verification. Ecological Modelling, 133(3):195–207, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000002933.

Asaeda:2001:MDP

[AS01] Takashi Asaeda and Duong Hong Son. A model of the development of a periphyton community: resource and flow dynamics. Ecological Modelling, 137(1):61–75, February 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004324.

Alexander:2003:ULC

[AS03] Robert R. Alexander and David W. Shields. Using land as a control variable in density-dependent bioeconomic models. Ecological Modelling, 170(2–3):193–201, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002266.

Arar:2009:ECR

[AS09] Joseph I. Arar and Douglas Southgate. Evaluating CO_{2} reduction strategies in the US. Ecological Modelling, 220 (4):582–588, February 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008005814.

Arreguin-Sanchez:2002:FBS

[ASAC02] Francisco Arreguín-Sánchez, Enrique Arcos, and Ernesto A. Chávez. Flows of biomass and structure in an exploited benthic ecosystem in the Gulf of California, Mexico. Ecological Modelling, 156(2–3):167–183, November 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200159X.

AAberg:2009:ESE

[ÅSCP09] Per Åberg, Carl Johan Svensson, Hal Caswell, and Henrik Pavia. Environment-specific elasticity and sensitivity analysis of the stochastic growth rate. Ecological Modelling, 220(5):605–610, March 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005784.

Agarwal:2005:TDM

[ASG^{+}05] Deepak K. Agarwal, John A. Silander, Alan E. Gelfand, Robert E. Dewar, and John G. Mickelson. Tropical deforestation in Madagascar: analysis using hierarchical, spatially explicit, Bayesian regression models. Ecological Modelling, 185(1):105–131, June 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400599X.

Ashcroft:2006:MIL

[Ash06] M. B. Ashcroft. A method for improving landscape scale temperature predictions and the implications for vegetation modelling. Ecological Modelling, 197(3–4):394–404, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006001372.

Arreguin-Sanchez:2004:OMS

[ASHHRRPE04] Francisco Arreguín-Sánchez, Agustín Hernández-Herrera, Mauricio Ramírez-Rodríguez, and Horacio Pérez-España. Optimal management scenarios for the artisanal fisheries in the ecosystem of La Paz Bay, Baja California Sur, Mexico. Ecological Modelling, 172(2–4):373–382, March 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003003806.

Akita:2008:SAM

[ASI^{+}08] Tetsuya Akita, Kenshi Sakai, Yuko Iwabuchi, Yoshinobu Hoshino, and Xujun Ye. Spatial autocorrelation in masting phenomena of Quercus serrata detected by multi-spectral imaging. Ecological Modelling, 215(1–3):217–224, July 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001014.

Ahearn:2001:TIB

[ASJD01] Sean C. Ahearn, James L. David Smith, Anup R. Joshi, and Ji Ding. TIGMOD: an individual-based spatially explicit model for simulating tiger/human interaction in multiple use forests. Ecological Modelling, 140(1–2):81–97, May 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002587.

Anisimov:2002:VST

[ASN02] O. A. Anisimov, N. I. Shiklomanov, and F. E. Nelson. Variability of seasonal thaw depth in permafrost regions: a stochastic modeling approach. Ecological Modelling, 153(3):217–227, August 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000169.

Aspinall:2002:ULR

[Asp02] Richard J. Aspinall. Use of logistic regression for validation of maps of the spatial distribution of vegetation species derived from high spatial resolution hyperspectral remotely sensed data. Ecological Modelling, 157 (2–3):301–312, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002002016.

Arhonditsis:2007:DRN

[ASP^{+}07] George B. Arhonditsis, Craig A. Stow, Hans W. Paerl, Lexia M. Valdes-Weaver, Laura J. Steinberg, and Kenneth H. Reckhow. Delineation of the role of nutrient dynamics and hydrologic forcing on phytoplankton patterns along a freshwater–marine continuum. Ecological Modelling, 208(2–4):230–246, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700316X.

Arhonditsis:2006:EEP

[ASS^{+}06] G. B. Arhonditsis, C. A. Stow, L. J. Steinberg, M. A. Kenney, R. C. Lathrop, S. J. McBride, and K. H. Reckhow. Exploring ecological patterns with structural equation modeling and Bayesian analysis. Ecological Modelling, 192(3–4):385–409, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003765.

Arreguin-Sanchez:2004:SRH

[ASZRMH^{+}04] Francisco Arreguin-Sanchez, Manuel Zetina-Rejón, Sherry Manickchand-Heileman, Mauricio Ramírez-Rodríguez, and Laura Vidal. Simulated response to harvesting strategies in an exploited ecosystem in the southwestern Gulf of Mexico. Ecological Modelling, 172(2–4):421–432, March 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003003788.

Arreguin-Sanchez:2008:EEB

[ASZRRR08] Francisco Arreguin-Sanchez, Manuel Zetina-Rejón, and Mauricio Ramírez-Rodríguez. Exploring ecosystem-based harvesting strategies to recover the collapsed pink shrimp (Farfantepenaeus duorarum) fishery in the southern Gulf of Mexico. Ecological Modelling, 214(2–4):83–94, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007006187.

Adioui:2003:AFS

[ATA03] M. Adioui, J. P. Treuil, and O. Arino. Alignment in a fish school: a mixed Lagrangian–Eulerian approach. Ecological Modelling, 167(1–2):19–32, September 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001017.

Arhonditsis:2000:QEN

[ATAK00] G. Arhonditsis, G. Tsirtsis, M. O. Angelidis, and M. Karydis. Quantification of the effects of nonpoint nutrient sources to coastal marine eutrophication: applications to a semienclosed gulf in the Mediterranean Sea. Ecological Modelling, 129(2–3):209–227, May 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002398.

Atanasova:2006:AMF

[ATD^{+}06] Nataša Atanasova, Ljupčo Todorovski, Sašo Džeroski, Špela Rekar Remec, Friedrich Recknagel, and Boris Kompare. Automated modelling of a food web in lake bled using measured data and a library of domain knowledge. Ecological Modelling, 194(1–3):37–48, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004953.

Atanasova:2006:CLD

[ATDK06] Nataša Atanasova, Ljupčo Todorovski, Sašo Džeroski, and Boris Kompare. Constructing a library of domain knowledge for automated modelling of aquatic ecosystems. Ecological Modelling, 194(1–3):14–36, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004941.

Atanasova:2008:AAM

[ATDK08] Nataša Atanasova, Ljupčo Todorovski, Sašo Džeroski, and Boris Kompare. Application of automated model discovery from data and expert knowledge to a real-world domain: Lake Glumsø. Ecological Modelling, 212(1–2):92–98, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005261.

Aubone:2004:LSO

[Aub04a] Aníbal Aubone. Loss of stability owing to a stable age structure skewed toward juveniles. Ecological Modelling, 175(1):55–64, June 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004678.

Aubone:2004:TSE

[Aub04b] Aníbal Aubone. Threshold for sustainable exploitation of an age-structured fishery stock. Ecological Modelling, 173(1):95–107, March 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003557.

Aumann:2007:MDS

[Aum07] Craig A. Aumann. A methodology for developing simulation models of complex systems. Ecological Modelling, 202(3–4):385–396, April 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005400.

Austin:2002:SPS

[Aus02] M. P. Austin. Spatial prediction of species distribution: an interface between ecological theory and statistical modelling. Ecological Modelling, 157(2–3):101–118, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002053.

Austin:2007:SDM

[Aus07] Mike Austin. Species distribution models and ecological theory: a critical assessment and some possible new approaches. Ecological Modelling, 200(1–2):1–19, January 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006003140.

Ainsworth:2008:ESS

[AVP08] C. H. Ainsworth, D. A. Varkey, and T. J. Pitcher. Ecosystem simulations supporting ecosystem-based fisheries management in the Coral Triangle, Indonesia. Ecological Modelling, 214(2–4):361–374, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001257.

Aggelis:2005:PPD

[AVTP05] George Aggelis, Dimitris V. Vayenas, Vasiliki Tsagou, and Stavros Pavlou. Prey–predator dynamics with predator switching regulated by a catabolic repression control mode. Ecological Modelling, 183(4):451–462, May 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004005034.

Acevedo:2000:MCS

[AW00] Miguel F. Acevedo and William T. Waller. Modelling and control of a simple trophic aquatic system. Ecological Modelling, 131(2–3):269–284, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002556.

Abdul-Wahab:2006:IFD

[AW06a] Sabah A. Abdul-Wahab. Impact of fugitive dust emissions from cement plants on nearby communities. Ecological Modelling, 195(3–4):338–348, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006125.

Abdul-Wahab:2006:RMP

[AW06b] Sabah A. Abdul-Wahab. The role of meteorology on predicting SO_{2} concentrations around a refinery: a case study from Oman. Ecological Modelling, 197(1–2):13–20, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000986.

Adams:2004:ASD

[AWL04] B. Adams, A. White, and T. M. Lenton. An analysis of some diverse approaches to modelling terrestrial net primary productivity. Ecological Modelling, 177(3–4):353–391, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004001462.

Alexandrov:2007:PFM

[AY07] Georgii A. Alexandrov and Yoshiki Yamagata. A peaked function for modeling temperature dependence of plant productivity. Ecological Modelling, 200(1–2):189–192, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003371.

Aydin:2008:MED

[Ayd08] Mehmet Aydin. A model for evaporation and drainage investigations at ground of ordinary rainfed-areas. Ecological Modelling, 217(1–2):148–156, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002986.

Aita:2007:IVL

[AYK07] Maki Noguchi Aita, Yasuhiro Yamanaka, and Michio J. Kishi. Interdecadal variation of the lower trophic ecosystem in the northern Pacific between 1948 and 2002, in a 3-D implementation of the NEMURO model. Ecological Modelling, 202(1–2):81–94, March 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004583.

Aydin:2005:TSM

[AYKY05] Mehmet Aydin, Sheng Li Yang, Nurten Kurt, and Tomohisa Yano. Test of a simple model for estimating evaporation from bare soils in different environments. Ecological Modelling, 182(1):91–105, February 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004089.

Avelino:2006:ICR

[AZM^{+}06] J. Avelino, H. Zelaya, A. Merlo, A. Pineda, M. Ordoñez, and S. Savary. The intensity of a coffee rust epidemic is dependent on production situations. Ecological Modelling, 197(3–4):431–447, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001360.

Barros:2005:CCS

[BA05] H. Barros and J. M. Abril. Constraints in the construction and/or selection of kinetic box models for the uptake of radionuclides and heavy metals by suspended particulate matter. Ecological Modelling, 185(2–4):371–385, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000074.

Bolduc:2008:MWA

[BA08] François Bolduc and Alan D. Afton. Monitoring waterbird abundance in wetlands: the importance of controlling results for variation in water depth. Ecological Modelling, 216(3–4):402–408, September 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002329.

Balzter:2000:MCM

[Bal00] Heiko Balzter. Markov chain models for vegetation dynamics. Ecological Modelling, 126(2–3):139–154, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002623.

Bandt:2005:OTS

[Ban05] Christoph Bandt. Ordinal time series analysis. Ecological Modelling, 182(3–4):229–238, March 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001607.

Benabdellah:2003:MCM

[BAP^{+}03] Benabdellah Benabdellah, Karl-Friedrich Albrecht, Valeri L. Pomaz, Elizabeth A. Denisenko, and Dmitrii O. Logofet. Markov chain models for forest successions in the Erzgebirge, Germany. Ecological Modelling, 159(2–3):145–160, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002855.

Bozic:2006:MDS

[BAP^{+}06] Mario Božić, Oleg Antonić, Renata Pernar, Sven D. Jelaska, Josip Križan, Juro Čavlović, and Vladimir Kušan. Modelling the damage status of silver fir trees (Abies alba mill.) on the basis of geomorphological, climatic and stand factors. Ecological Modelling, 194(1–3):202–208, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500517X.

Bartel:2000:ALP

[Bar00] A. Bartel. Analysis of landscape pattern: towards a ‘top down’ indicator for evaluation of landuse. Ecological Modelling, 130(1–3):87–94, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002143.

Barile:2004:CTS

[Bar04] Peter J. Barile. Comments on “Trophic structure and productivity of a windward coral reef community on Eniwetok Atoll” [Ecological Monographs 25 (3) (1955) 291–320]. Ecological Modelling, 178(1–2):51–57, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005167.

Bastianoni:2004:SGS

[Bas04] Simone Bastianoni. Sitting on a giant’s shoulders. Ecological Modelling, 178(1–2):253–254, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005179.

Behm:2001:MHC

[BB01] Pamela M. Behm and Roelof M. J. Boumans. Modeling herbivorous consumer consumption in the Great Bay Estuary, New Hampshire. Ecological Modelling, 143(1–2):71–94, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003556.

Barendregt:2003:RVP

[BB03a] Aat Barendregt and Ana M. F. Bio. Relevant variables to predict macrophyte communities in running waters. Ecological Modelling, 160(3):205–217, February 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002545.

Burnett:2003:MSS

[BB03b] C. Burnett and Thomas Blaschke. A multi-scale segmentation/object relationship modelling methodology for landscape analysis. Ecological Modelling, 168(3):233–249, October 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000300139X.

Bigler:2004:APT

[BB04] Christof Bigler and Harald Bugmann. Assessing the performance of theoretical and empirical tree mortality models using tree-ring series of Norway spruce. Ecological Modelling, 174(3):225–239, May 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004265.

Bisignanesi:2007:MIP

[BB07a] V. Bisignanesi and M. S. Borgas. Models for integrated pest management with chemicals in atmospheric surface layers.

Ecological Modelling, 201(1):2–10, February 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003504.

Bryhn:2007:CNG

[BB07b] Andreas C. Bryhn and Thorsten Blenckner. Can nitrogen gas be deficient for nitrogen fixation in lakes? Ecological Modelling, 202(3–4):362–372, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005357.

Ballyk:2008:RRT

[BB08a] Mary Ballyk and Ernest Barany. The role of resource types in the control of chemostats using feedback linearization. Ecological Modelling, 211(1–2):25–35, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700422X.

Best:2008:CFB

[BB08b] Elly P. H. Best and William A. Boyd. A carbon flowbased modelling approach to ecophysiological processes and biomass dynamics of Vallisneria americana, with applications to temperate and tropical water bodies. Ecological Modelling, 217(1–2):117–131, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002937.

Brugnach:2003:DST

[BBB03] Marcela Brugnach, John Bolte, and G. A. Bradshaw. Determining the significance of threshold values uncertainty in rule-based classification models. Ecological Modelling, 160(1–2):63–76, February 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003095.

Binder:2003:APL

[BBC03] Claudia Binder, Roelof M. Boumans, and Robert Costanza. Applying the Patuxent Landscape Unit Model to human dominated ecosystems: the case of agriculture. Ecological Modelling, 159(2–3):161–177, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002764.

Beulke:2004:EPM

[BBD^{+}04] Sabine Beulke, Colin D. Brown, Igor G. Dubus, Christopher J. Fryer, and Allan Walker. Evaluation of probabilistic modelling approaches against data on leaching of isoproturon through undisturbed lysimeters. Ecological Modelling, 179(1):131–144, November 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002674.

Bormann:2007:AES

[BBGH07] Helge Bormann, Lutz Breuer, Thomas Gräff, and Johan A. Huisman. Analysing the effects of soil properties changes associated with land use changes on the simulated water balance: a comparison of three hydrological catchment models for scenario analysis. Ecological Modelling, 209(1):29–40, November 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003560.

Braunisch:2008:LEM

[BBGH08] Veronika Braunisch, Kurt Bollmann, Roland F. Graf, and Alexandre H. Hirzel. Living on the edge — modelling habitat suitability for species at the edge of their fundamental niche. Ecological Modelling, 214(2–4):153–167, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000598.

Bampfylde:2005:MRF

[BBGM05] C. J. Bampfylde, N. D. Brown, D. J. Gavaghan, and P. K. Maini. Modelling rain forest diversity: the role of competition. Ecological Modelling, 188(2–4):253–278, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001249.

Bampfylde:2006:ACC

[BBGM06] C. J. Bampfylde, N. D. Brown, D. J. Gavaghan, and P. K. Maini. Achieving coexistence — comment on “Modelling rain forest diversity: the role of competition” by Bampfylde et al. (2005). Ecological Modelling, 195(3–4):407–408, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000020.

Brawley:2003:PAE

[BBKN03] John W. Brawley, Mark J. Brush, James N. Kremer, and Scott W. Nixon. Potential applications of an empirical phytoplankton production model to shallow water ecosystems. Ecological Modelling, 160(1–2):55–61, February 1, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002003101.

Bandyopadhyay:2004:DAH

[BBM04] M. Bandyopadhyay, R. Bhattacharyya, and B. Mukhopadhyay. Dynamics of an autotroph–herbivore ecosystem with nutrient recycling. Ecological Modelling, 176 (1–2):201–209, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000237.

Bald:2006:SDM

[BBM06] Juan Bald, Angel Borja, and Iñigo Muxika. A system dynamics model for the management of the gooseneck barnacle (Pollicipes pollicipes) in the marine reserve of Gaztelugatxe (Northern Spain). Ecological Modelling, 194(1–3):306–315, March 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005005284.

Bata:2007:ETS

[BBP^{+}07] Seth A. Bata, Stuart R. Borrett, Bernard C. Patten, Stuart J. Whipple, John R. Schramski, and David K. Gattie. Equivalence of throughflow- and storage-based environs. Ecological Modelling, 206(3–4):400–406, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001974.

Bode:2008:UCN

[BBP08] Michael Bode, Kevin Burrage, and Hugh P. Possingham. Using complex network metrics to predict the persistence of metapopulations with asymmetric connectivity patterns. Ecological Modelling, 214(2–4):201–209, June 24, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008000744.

Barros:2000:FMP

[BBT00] L. C. Barros, R. C. Bassanezi, and P. A. Tonelli. Fuzzy modelling in population dynamics. Ecological Modelling, 128(1):27–33, April 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380099002239.

Bergant:2006:UMC

[BBT06] Klemen Bergant, Lučka Kajfež Bogataj, and Stanislav Trdan. Uncertainties in modelling of climate change impact in future: an example of onion thrips (Thrips tabaci Lindeman) in Slovenia. Ecological Modelling, 194(1–3):244–255, March 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005005223.

Bartsch:2001:IBG

[BC01] J. Bartsch and S. H. Coombs. An individual-based growth and transport model of the early life-history stages of mackerel (Scomber scombrus) in the eastern North Atlantic. Ecological Modelling, 138(1–3):127–141, March 15, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000003987.

Bahamon:2003:MNF

[BC03] Nixon Bahamón and Antonio Cruzado. Modelling nitrogen fluxes in oligotrophic environments: NW Mediterranean and NE Atlantic. Ecological Modelling, 163(3):223–244, May 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000073.

BalaguerBallester:2002:EDA

[BCCR^{+}02] E. Balaguer Ballester, G. Camps i Valls, J. L. Carrasco-Rodriguez, E. Soria Olivas, and S. del Valle-Tascon. Effective 1-day ahead prediction of hourly surface ozone concentrations in eastern Spain using linear models and neural networks. Ecological Modelling, 156(1):27–41, October 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002001278.

Bracchini:2005:MCV

[BCD^{+}05] Luca Bracchini, Andrés Cózar, Arduino Massimo Dattilo, Maria Pia Picchi, Carmen Arena, Stefania Mazzuoli, and Steven A. Loiselle. Modelling the components of the vertical attenuation of ultraviolet radiation in a wetland lake ecosystem. Ecological Modelling, 186(1):43–54, July 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500150X.

Buzzelli:2000:SPP

[BCDJ00] Christopher P. Buzzelli, Daniel L. Childers, Quan Dong, and Ronald D. Jones. Simulation of periphyton phosphorus dynamics in Everglades National Park. Ecological Modelling, 134(1):103–115, September 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003392.

Brovkin:2008:ESC

[BCG08] Victor Brovkin, Alexander Cherkinsky, and Sergey Goryachkin. Estimating soil carbon turnover using radiocarbon data: a case-study for European Russia. Ecological Modelling, 216(2):178–187, August 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001385.

Buchi:2009:IES

[BCH09] Lucie Büchi, Pascal-Antoine Christin, and Alexandre H. Hirzel. The influence of environmental spatial structure on the life-history traits and diversity of species in a metacommunity. Ecological Modelling, 220(21):2857–2864, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009005365.

Bandelj:2009:MSD

[BCL^{+}09] Vinko Bandelj, Daniele Curiel, Sovan Lek, Andrea Rismondo, and Cosimo Solidoro. Modelling spatial distribution of hard bottom benthic communities and their functional response to environmental parameters. Ecological Modelling, 220(21):2838–2850, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002853.

Brando:2004:AEM

[BCLR04] Vittorio Ernesto Brando, Riccardo Ceccarelli, Simone Libralato, and Giampiero Ravagnan. Assessment of environmental management effects in a shallow water basin using mass-balance models. Ecological Modelling, 172(2–4):213–232, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003703.

Basille:2008:AHS

[BCM^{+}08] Mathieu Basille, Clément Calenge, Éric Marboutin, Reidar Andersen, and Jean-Michel Gaillard. Assessing habitat selection using multivariate statistics: Some refinements of the ecological-niche factor analysis. Ecological Modelling, 211(1–2):233–240, February 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004577.

Bonifazi:2003:MAQ

[BCMR03] Giuseppe Bonifazi, Laura Cutaia, Paolo Massacci, and Ivan Roselli. Monitoring of abandoned quarries by remote sensing and in situ surveying. Ecological Modelling, 170(2–3):213–218, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300228X.

Beeravolu:2009:SEC

[BCPM09] Champak R. Beeravolu, Pierre Couteron, Raphaël Pélissier, and François Munoz. Studying ecological communities from a neutral standpoint: a review of models’ structure and parameter estimation. Ecological Modelling, 220(20):2603–2610, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004499.

Bastianoni:2009:STP

[BCRP09] S. Bastianoni, D. E. Campbell, R. Ridolfi, and F. M. Pulselli. The solar transformity of petroleum fuels. Ecological Modelling, 220(1):40–50, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004407.

Benavent-Corai:2007:SPF

[BCRSTVG07] J. Benavent-Corai, C. Rojo, J. Suárez-Torres, and L. Velasco-García. Scaling properties in forest fire sequences: the human role in the order of nature. Ecological Modelling, 205(3–4):336–342, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001019.

Brown:2009:PNS

[BCS09] Mark T. Brown, Matthew J. Cohen, and Sharlynn Sweeney. Predicting national sustainability: the convergence of energetic, economic and environmental realities. Ecological Modelling, 220(23):3424–3438, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009006073.

Bastianoni:2005:STO

[BCST05] Simone Bastianoni, Daniel Campbell, Ludovico Susani, and Enzo Tiezzi. The solar transformity of oil and petroleum natural gas. Ecological Modelling, 186(2):212–220, August 15, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000500027X.

Bendoricchio:2005:WQM

[BD05] Giuseppe Bendoricchio and Gabriella De Boni. A waterquality model for the Lagoon of Venice, Italy. Ecological Modelling, 184(1):69–81, May 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400585X.

Bao-dong:2005:DED

[BdB05] Wang Bao-dong and Uwe Brockmann. On the deduction of equations for the determination of anthropogenic influence for assessment of estuarine trophic status. Ecological Modelling, 185(2–4):545–548, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004006131.

Banks:2008:TVV

[BDBS08] J. E. Banks, L. K. Dick, H. T. Banks, and J. D. Stark. Time-varying vital rates in ecotoxicology: Selective pesticides and aphid population dynamics. Ecological Modelling, 210(1–2):155–160, January 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003717.

Bergez:2001:MOO

[BDD^{+}01] J.-E. Bergez, P. Debaeke, J.-M. Deumier, B. Lacroix, D. Leenhardt, P. Leroy, and D. Wallach. MODERATO: an object-oriented decision tool for designing maize irrigation schedules. Ecological Modelling, 137(1):43–60, February 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000004312.

Bartsev:2008:PWS

[BDE08] Sergey I. Bartsev, Andrey G. Degermendzhi, and Dmitry V. Erokhin. Principle of the worst scenario in the modelling past and future of biosphere dynamics. Ecological Modelling, 216(2):160–171, August 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001361.

Betts:2006:ISA

[BDF^{+}06] Matthew G. Betts, A. W. Diamond, G. J. Forbes, M.-A. Villard, and J. S. Gunn. The importance of spatial autocorrelation, extent and resolution in predicting forest bird occurrence. Ecological Modelling, 191(2):197–224, January 31, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002759.

Buffoni:2001:LDT

[BDG01] G. Buffoni, G. Di Cola, and L. Garaventa. The local dynamics of a tri-trophic system. Ecological Modelling, 138(1–3):31–39, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003914.

Booker:2004:PJS

[BDI04] D. J. Booker, M. J. Dunbar, and A. Ibbotson. Predicting juvenile salmonid drift-feeding habitat quality using a three-dimensional hydraulic-bioenergetic model. Ecological Modelling, 177(1–2):157–177, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001140.

Brizhik:2009:REP

[BDJ^{+}09] Larissa Brizhik, Emilio Del Giudice, Sven E. Jørgensen, Nadia Marchettini, and Enzo Tiezzi. The role of electromagnetic potentials in the evolutionary dynamics of ecosystems. Ecological Modelling, 220(16):1865–1869, August 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009002750.

Borgesen:2001:EEL

[BDK01] Christen Duus Børgesen, Jørgen Djurhuus, and Arne Kyllingsbæk. Estimating the effect of legislation on nitrogen leaching by upscaling field simulations. Ecological Modelling, 136(1):31–48, January 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003860.

Bardos:2006:DRF

[BDLL06] David C. Bardos, Robert W. Day, Nicholas T. Lawson, and Nicholas A. Linacre. Dynamical response to fishing varies with compensatory mechanism: an abalone population model. Ecological Modelling, 192(3–4):523–542, February 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005003716.

Bradshaw:2002:UAN

[BDP^{+}02] Corey J. A. Bradshaw, Lloyd S. Davis, Martin Purvis, Qingqing Zhou, and George L. Benwell. Using artificial neural networks to model the suitability of coastline for breeding by New Zealand fur seals (Arctocephalus forsteri). Ecological Modelling, 148(2):111–131, February 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004252.

Beaujouan:2001:MES

[BDR01] Véronique Beaujouan, Patrick Durand, and Laurent Ruiz. Modelling the effect of the spatial distribution of agricultural practices on nitrogen fluxes in rural catchments. Ecological Modelling, 137(1):93–105, February 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000435X.

Bone:2006:FCC

[BDR06] Christopher Bone, Suzana Dragicevic, and Arthur Roberts. A fuzzy-constrained cellular automata model of forest insect infestations. Ecological Modelling, 192(1–2):107–125, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005004771.

Breuer:2003:PPV

[BEF03] Lutz Breuer, Klaus Eckhardt, and Hans-Georg Frede. Plant parameter values for models in temperate climates. Ecological Modelling, 169(2–3):237–293, November 15, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002746.

Bell:2000:MED

[BEM00] V. A. Bell, J. M. Elliott, and R. J. Moore. Modelling the effects of drought on the population of brown trout in black brows Beck. Ecological Modelling, 127(2–3):141–159, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002094.

Benhamou:2004:END

[Ben04] Simon Benhamou. On the expected net displacement of animals’ random movements. Ecological Modelling, 171 (1–2):207–208, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003375.

Bentley:2006:CDP

[Ben06] Richard P. Bentley. Consumption driven population dynamics (CDPD). Ecological Modelling, 192(1–2):1–24, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004096.

Berec:2002:TSE

[Ber02] Luděk Berec. Techniques of spatially explicit individualbased models: construction, simulation, and mean-field analysis. Ecological Modelling, 150(1–2):55–81, April 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000100463X.

Bey:2003:EMF

[Bey03] Ahmad Bey. Evapoclimatonomy modelling of four restoration stages following Krakatau’s 1883 destruction. Ecological Modelling, 169(2–3):327–337, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002783.

Brommer:2007:APD

[BF07] Jon E. Brommer and Marianne S. Fred. Accounting for possible detectable distances in a comparison of dispersal: Apollo dispersal in different habitats. Ecological Modelling, 209(2–4):407–411, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003651.

Basset:2002:OES

[BFD02] Alberto Basset, M. Fedele, and D. L. DeAngelis. Optimal exploitation of spatially distributed trophic resources and population stability. Ecological Modelling, 151(2–3):245–260, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001004902.

Bommarco:2007:OSP

[BFE07] Riccardo Bommarco, Sascha O. Firle, and Barbara Ekbom. Outbreak suppression by predators depends on spatial distribution of prey. Ecological Modelling, 201 (2):163–170, February 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004182.

Buonomo:2005:MMI

[BFHR05] Bruno Buonomo, Margherita Falcucci, Vincent Hull, and Salvatore Rionero. A mathematical model for an integrated experimental aquaculture plant. Ecological Modelling, 183(1):11–28, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004156.

Bottcher:2008:UAC

[BFOS08] Hannes Böttcher, Annette Freibauer, Michael Obersteiner, and Ernst-Detlef Schulze. Uncertainty analysis of climate change mitigation options in the forestry sector using a generic carbon budget model. Ecological Modelling, 213(1):45–62, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007006151.

Bricker:2003:IMA

[BFS03] S. B. Bricker, J. G. Ferreira, and T. Simas. An integrated methodology for assessment of estuarine trophic status. Ecological Modelling, 169(1):39–60, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001996.

Bricker:2005:RCD

[BFS05] Suzanne B. Bricker, Joao G. Ferreira, and Teresa Simas. Reply to “Comment on the deduction of equations for the determination of anthropogenic influence for assessment of estuarine trophic status”. Ecological Modelling, 185(2–4):549–550, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004006143.

Baird:2009:CAB

[BFU^{+}09] Dan Baird, Brian D. Fath, Robert E. Ulanowicz, Harald Asmus, and Ragnhild Asmus. On the consequences of aggregation and balancing of networks on system properties derived from ecological network analysis. Ecological Modelling, 220(23):3465–3471, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009006164.

Baum:2001:EHF

[BG01a] Kristen A. Baum and William E. Grant. Erratum to “Hummingbird foraging behavior in different patch types: simulation of alternative strategies”: [Ecol. Model. 137/2–3 (2001) 201–209]. Ecological Modelling, 142(1–2):185, August 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003684.

Baum:2001:HFB

[BG01b] Kristen A. Baum and William E. Grant. Hummingbird foraging behavior in different patch types: simulation of alternative strategies. Ecological Modelling, 137 (2–3):201–209, February 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000004361.

Buffoni:2003:LPM

[BG03] G. Buffoni and G. Gilioli. A lumped parameter model for acarine predator–prey population interactions. Ecological Modelling, 170(2–3):155–171, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002230.

Bouraoui:2008:IMF

[BG08] Fayçal Bouraoui and Bruna Grizzetti. An integrated modelling framework to estimate the fate of nutrients: Application to the Loire (France). Ecological Modelling, 212(3–4):450–459, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007005662.

Barciela:2000:CMP

[BGF00] Rosa M. Barciela, Emilio Garcia, and Emilio Fernández. Corrigendum to “Modelling primary production in a coastal embayment affected upwelling using dynamic ecosystem models and artificial neural networks”: [Ecol. Model. 120 (1999) 199–211]. Ecological Modelling, 128(2–3):251, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002288.

Brosse:2001:UNS

[BGL01] S. Brosse, J. L. Giraudel, and S. Lek. Utilisation of non-supervised neural networks and principal component analysis to study fish assemblages. Ecological Modelling, 146(1–3):159–166, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003039.

Bell:2006:UDM

[BGMP06] V. A. Bell, D. G. George, R. J. Moore, and J. Parker. Using a 1-D mixing model to simulate the vertical flux of heat and oxygen in a lake subject to episodic mixing. Ecological Modelling, 190(1–2):41–54, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002279.

Brophy:2007:MFA

[BGWC07] C. Brophy, D. J. Gibson, P. M. Wayne, and J. Connolly. A modelling framework for analysing the reproductive output of individual plants grown in monoculture. Ecological Modelling, 207(2–4):99–108, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002220.

Berger:2000:NAS

[BH00] Uta Berger and Hanno Hildenbrandt. A new approach to spatially explicit modelling of forest dynamics: spacing, ageing and neighbourhood competition of mangrove trees. Ecological Modelling, 132(3):287–302, August 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002982.

Beres:2001:PBT

[BH01] Diane L. Beres and Douglas M. Hawkins. Plackett–Burman technique for sensitivity analysis of many-parametered models. Ecological Modelling, 141(1–3):171–183, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000100271X.

Berjak:2002:ICA

[BH02] Stephen G. Berjak and John W. Hearne. An improved cellular automaton model for simulating fire in a spatially heterogeneous savanna system. Ecological Modelling, 148(2):133–151, February 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004239.

Boulion:2003:NGD

[BH03] Viktor V. Boulion and Lars Håkanson. A new general dynamic model to predict biomass and production of bacterioplankton in lakes. Ecological Modelling, 160 (1–2):91–114, February 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002003265.

Birkved:2006:PME

[BH06] Morten Birkved and Michael Z. Hauschild. PestLCI — a model for estimating field emissions of pesticides in agricultural LCA. Ecological Modelling, 198(3–4):433–451, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006002535.

Braak:2003:BMB

[BHAV03] Cajo J. F. Ter Braak, Herbert Hoijtink, Wies Akkermans, and Piet F. M. Verdonschot. Bayesian model-based cluster analysis for predicting macrofaunal communities. Ecological Modelling, 160(3):235–248, February 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002569.

Bannayan:2004:PIM

[BHC04a] M. Bannayan, Gerrit Hoogenboom, and N. M. J. Crout. Photothermal impact on maize performance: a simulation approach. Ecological Modelling, 180(2–3):277–290, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004002820.

Biber:2004:MDT

[BHC04b] Patrick D. Biber, Mark A. Harwell, and Wendell P. Cropper. Modeling the dynamics of three functional groups of macroalgae in tropical seagrass habitats. Ecological Modelling, 175(1):25–54, June 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004642.

Ben-Hamadou:2005:ISP

[BHFMG05] Radhouan Ben-Hamadou, Ibanez Frédéric, Picheral Marc, and Gorsky Gabriel. Identification of step pattern in ordered data sets using the Walsh transform algorithm. Ecological Modelling, 182(1):11–24, February 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400403X.

Bruce:2006:NSR

[BHI^{+}06] Louise C. Bruce, David Hamilton, Jörg Imberger, Gideon Gal, Moshe Gophen, Tamar Zohary, and K. David Hambright. A numerical simulation of the role of zooplankton in C, N and P cycling in Lake Kinneret, Israel. Ecological Modelling, 193(3–4):412–436, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004485.

Brown:2004:E

[BHJ04] M. T. Brown, C. A. S. Hall, and S. E. Jorgensen. Eulogy. Ecological Modelling, 178(1–2):1–10, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005118.

Benz:2001:EMD

[BHL01] J. Benz, R. Hoch, and T. Legović. ECOBAS — modelling and documentation. Ecological Modelling, 138(1–3):3–15, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003896.

Begg:2006:MPV

[BHM^{+}06] G. S. Begg, S. Hockaday, J. W. McNicol, M. Askew, and G. R. Squire. Modelling the persistence of volunteer oilseed rape (Brassica napus). Ecological Modelling, 198(1–2):195–207, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001827.

Beaumont:2005:PSD

[BHP05] Linda J. Beaumont, Lesley Hughes, and Michael Poulsen. Predicting species distributions: use of climatic parameters in BIOCLIM and its impact on predictions of species’ current and future distributions. Ecological Modelling, 186(2):251–270, August 15, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000530.

Burger:2008:MRI

[BHP08] David F. Burger, David P. Hamilton, and Conrad A. Pilditch. Modelling the relative importance of internal and external nutrient loads on water column nutrient concentrations and phytoplankton biomass in a shallow polymictic lake. Ecological Modelling, 211(3–4):411–423, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007004966.

Borsuk:2001:BHM

[BHSR01] Mark E. Borsuk, David Higdon, Craig A. Stow, and Kenneth H. Reckhow. A Bayesian hierarchical model to predict benthic oxygen demand from organic matter loading in estuaries and coastal zones. Ecological Modelling, 143(3):165–181, September 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003283.

Bui:2006:KDM

[BHV06] Elisabeth N. Bui, Brent L. Henderson, and Karin Viergever. Knowledge discovery from models of soil properties developed through data mining. Ecological Modelling, 191 (3–4):431–446, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002917.

Bennie:2008:SAC

[BHW^{+}08] Jonathan Bennie, Brian Huntley, Andrew Wiltshire, Mark O. Hill, and Robert Baxter. Slope, aspect and climate: Spatially explicit and implicit models of topographic microclimate in chalk grassland. Ecological Modelling, 216(1):47–59, August 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002056.

Bian:2003:REC

[Bia03] Ling Bian. The representation of the environment in the context of individual-based modeling. Ecological Modelling, 159(2–3):279–296, January 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002983.

Biondini:2001:TDS

[Bio01] Mario Biondini. A three-dimensional spatial model for plant competition in an heterogeneous soil environment. Ecological Modelling, 142(3):189–225, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100285X.

Biondini:2003:AXZ

[Bio03] Mario Biondini. Answer to Xiaoxian Zhang’s comment on “A three-dimensional spatial model for plant competition in an heterogeneous soil environment”. Ecological Modelling, 166(1–2):187–189, August 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001169.

Birky:2001:NSM

[Bir01] Alicia K. Birky. NDVI and a simple model of deciduous forest seasonal dynamics. Ecological Modelling, 143(1–2):43–58, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003544.

Birch:2006:DON

[Bir06] Colin P. D. Birch. Diagonal and orthogonal neighbours in grid-based simulations: Buffon’s stick after 200 years. Ecological Modelling, 192(3–4):637–644, February 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000500373X.

Bramanti:2009:MMC

[BIS09] Lorenzo Bramanti, Mimmo Iannelli, and Giovanni Santangelo. Mathematical modelling for conservation and management of gorgonians corals: youngs and olds, could they coexist? Ecological Modelling, 220(21):2851–2856, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000900074X.

Bruun:2002:ISO

[BJ02] Sander Bruun and Lars S. Jensen. Initialisation of the soil organic matter pools of the daisy model. Ecological Modelling, 153(3):291–295, August 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000170.

Branco:2006:MRI

[BJFM06] Manuela Branco, Hervé Jactel, José C. Franco, and Zvi Mendel. Modelling response of insect trap captures to pheromone dose. Ecological Modelling, 197 (1–2):247–257, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001141.

Blatt:2001:MSI

[BJH01] S. E. Blatt, J. A. Janmaat, and R. Harmsen. Modelling succession to include a herbivore effect. Ecological Modelling, 139(2–3):123–136, April 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002356.

Bettinger:2003:SFP

[BJJ03] Pete Bettinger, Debora L. Johnson, and K. Norman Johnson. Spatial forest plan development with ecological and economic goals. Ecological Modelling, 169(2– 3):215–236, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002710.

Boone:2006:SWF

[BJJ06] Randall B. Boone, Catherine M. Johnson, and Lucinda B. Johnson. Simulating wood frog movement in central Minnesota, USA using a diffusion model. Ecological Modelling, 198(1–2):255–262, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002079.

Baas:2009:ENE

[BJK09] J. Baas, T. Jager, and S. A. L. M. Kooijman. Estimation of no effect concentrations from exposure experiments when values scatter among individuals. Ecological Modelling, 220(3):411–418, February 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004997.

Baskent:2005:SFP

[BK05a] Emin Zeki Baskent and Sedat Keles. Spatial forest planning: a review. Ecological Modelling, 188(2–4):145–173, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001420.

Beran:2005:DMM

[BK05b] Bora Beran and Fikret Kargi. A dynamic mathematical model for wastewater stabilization ponds. Ecological Modelling, 181(1):39–57, January 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003576.

Batchelder:2007:EMN

[BK07] Harold P. Batchelder and Makoto Kashiwai. Ecosystem modeling with NEMURO within the PICES climate change and carrying capacity program. Ecological Modelling, 202(1–2):7–11, March 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004522.

Bolliger:2000:CMT

[BKB00] Janine Bolliger, Felix Kienast, and Harald Bugmann. Comparing models for tree distributions: concept, structures, and behavior. Ecological Modelling, 134(1):89–102, September 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003380.

Bourennane:2007:UAS

[BKC^{+}07] H. Bourennane, D. King, A. Couturier, B. Nicoullaud, B. Mary, and G. Richard. Uncertainty assessment of soil water content spatial patterns using geostatistical simulations: an empirical comparison of a simulation accounting for single attribute and a simulation accounting for secondary information. Ecological Modelling, 205(3–4):323–335, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001020.

Benie:2005:RSB

[BKGC05] G. B. Bénié, S. S. Kaboré, K. Goïta, and M.-F. Courel. Remote sensing-based spatio-temporal modeling to predict biomass in Sahelian grazing ecosystem. Ecological Modelling, 184(2–4):341–354, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005599.

Breton:2008:MMR

[BKMB08] André R. Breton, S. Dean Kildaw, K. Murra, and C. Loren Buck. Multilevel models reveal no cohort-level variation in time spent foraging to account for a collapse in kittiwake (Rissa tridactyla) breeding success. Ecological Modelling, 212(3–4):233–243, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005315.

Bahn:2008:DLS

[BKO08] Volker Bahn, William B. Krohn, and Raymond J. O’Connor. Dispersal leads to spatial autocorrelation in species distributions: a simulation model. Ecological Modelling, 213(3–4):285–292, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007006369.

Biondi:2008:NSM

[BKPS08] Franco Biondi, Tomasz J. Kozubowski, Anna K. Panorska, and Laurel Saito. A new stochastic model of episode peak and duration for eco-hydro-climatic applications. Ecological Modelling, 211(3–4):383–395, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004838.

Bartleson:2005:USM

[BKS05a] Richard D. Bartleson, W. Michael Kemp, and J. Court Stevenson. Use of a simulation model to examine effects of nutrient loading and grazing on Potamogeton perfoliatus L. communities in microcosms. Ecological Modelling, 185(2–4):483–512, July 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500013X.

Bogena:2005:DMG

[BKS^{+}05b] H. Bogena, R. Kunkel, T. Schöbel, H. P. Schrey, and F. Wendland. Distributed modeling of groundwater recharge at the macroscale. Ecological Modelling, 187(1): 15–26, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380005000372.

Brooks:2001:ORC

[BL01] Elizabeth N. Brooks and Jean-Dominique Lebreton. Optimizing removals to control a metapopulation: application to the yellow legged herring gull (Larus cachinnans). Ecological Modelling, 136(2–3):269–284, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004300.

Borcard:2002:ASS

[BL02] Daniel Borcard and Pierre Legendre. All-scale spatial analysis of ecological data by means of principal coordinates of neighbour matrices. Ecological Modelling, 153(1–2):51–68, July 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001005014.

Bousquet:2004:MAS

[BL04] F. Bousquet and C. Le Page. Multi-agent simulations and ecosystem management: a review. Ecological Modelling, 176(3–4):313–332, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000948.

Blackwell:2007:HPC

[Bla07a] P. G. Blackwell. Heterogeneity, patchiness and correlation of resources. Ecological Modelling, 207(2–4):349–355, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007002967.

Blanco:2007:RAE

[Bla07b] Juan A. Blanco. The representation of allelopathy in ecosystem-level forest models. Ecological Modelling, 209 (2–4):65–77, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007003225.

Blanchet:2008:MDS

[BLB08] F. Guillaume Blanchet, Pierre Legendre, and Daniel Borcard. Modelling directional spatial processes in ecological data. Ecological Modelling, 215(4):325–336, July 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001798.

Blanchet:2009:EMD

[BLB09] F. Guillaume Blanchet, Pierre Legendre, and Daniel Borcard. Erratum to “Modelling directional spatial processes in ecological data” [Ecol. Modell. 215 (2008) 325–336]. Ecological Modelling, 220(1):82–83, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003864.

Bousquet:2001:MSH

[BLBT01] F. Bousquet, C. Le Page, I. Bakam, and A. Takforyan. Multiagent simulations of hunting wild meat in a village in eastern Cameroon. Ecological Modelling, 138 (1–3):331–346, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000004129.

Bergamino:2007:EDP

[BLC^{+}07] Nadia Bergamino, Steven A. Loiselle, Andrés Cózar, Arduino M. Dattilo, Luca Bracchini, and Claudio Rossi. Examining the dynamics of phytoplankton biomass in Lake Tanganyika using Empirical Orthogonal Functions. Ecological Modelling, 204(1–2):156–162, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006715.

Boudjema:2006:ANM

[BLDCM06] Gérard Boudjema, Guy Lempérière, Magali Deschamps-Cottin, and David George Molland. Analysis and nonlinear modeling of the mound-building ant Formica lugubris spatial multi-scale dynamic in a larch-tree stand of the southern French Alps. Ecological Modelling, 190(1–2):147–158, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002401.

Bachelet:2000:IBF

[BLDN00] Dominique Bachelet, James M. Lenihan, Christopher Daly, and Ronald P. Neilson. Interactions between fire, grazing and climate change at Wind Cave National Park, SD. Ecological Modelling, 134(2–3):229–244, October 30, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000003434.

Barclay:2006:EFS

[BLHB06] Hugh J. Barclay, Chao Li, Brad Hawkes, and Laura Benson. Effects of fire size and frequency and habitat heterogeneity on forest age distribution. Ecological Modelling, 197(1–2):207–220, August 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001177.

Bettinger:2005:HSF

[BLJS05] Pete Bettinger, Marie Lennette, K. Norman Johnson, and Thomas A. Spies. A hierarchical spatial framework for forest landscape planning. Ecological Modelling, 182(1):25–48, February 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004041.

Bayraktaroglu:2003:DTW

[BLVC03] E. Bayraktaroğlu, T. Legović, Z. R. Velasquez, and A. Cruzado. Diatom Thalassiosira weissflogii in oligotrophic versus eutrophic culture: models and ultrastructure. Ecological Modelling, 170(2–3):237–243, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300231X.

Bastianoni:2000:PCP

[BM00] Simone Bastianoni and Nadia Marchettini. The problem of co-production in environmental accounting by emergy analysis. Ecological Modelling, 129(2–3):187–193, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002325.

Bithell:2007:ECS

[BM07] M. Bithell and W. D. Macmillan. Escape from the cell: Spatially explicit modelling with and without grids. Ecological Modelling, 200(1–2):59–78, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003425.

Boykin:2008:SPE

[BM08] Kenneth G. Boykin and Kirk C. McDaniel. Simulated potential effects of ecological factors on a hypothetical population of Chiricahua leopard frog (Rana chiricahuensis). Ecological Modelling, 218(1–2):175–181, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003098.

Banal:2007:SRV

[BMB07] S. Banal, D. J. Marceau, and A. Bouchard. Sapling responses to variations in gap densities and spatial configurations modeled using SORTIE. Ecological Modelling, 206(1–2):41–53, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001603.

Becerra-Munoz:2003:SPA

[BMBOCR03] Salvador Becerra-Muñoz, Héctor René Buelna-Osben, and Juan Manuel Catalán-Romero. Spatial patterns of ARIMA modeled rates of change of atherinids (Chirostoma spp.) and goodeid Chapalichthys encaustus from Lake Chapala, México. Ecological Modelling, 165(2–3):237–250, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000735.

Braune:2009:IEL

[BMD09] Henning Braune, Johannes Müller, and Wulf Diepenbrock. Integrating effects of leaf nitrogen, age, rank, and growth temperature into the photosynthesis-stomatal conductance model LEAFC3-N parameterised for barley (Hordeum vulgare L.). Ecological Modelling, 220(13–14):1599–1612, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002257.

BenDor:2006:MSE

[BMF^{+}06] Todd K. BenDor, Sara S. Metcalf, Lauren E. Fontenot, Brandi Sangunett, and Bruce Hannon. Modeling the spread of the emerald ash borer. Ecological Modelling, 197(1–2):221–236, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000600113X.

Beaudouin:2008:SPC

[BMG08] Rémy Beaudouin, Gilles Monod, and Vincent Ginot. Selecting parameters for calibration via sensitivity analysis: an individual-based model of mosquitofish population dynamics. Ecological Modelling, 218(1–2):29–48, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003013.

Breckling:2005:EPI

[BMR^{+}05] Broder Breckling, Felix Müller, Hauke Reuter, Franz Hölker, and Otto Fränzle. Emergent properties in individualbased ecological models — introducing case studies in an ecosystem research context. Ecological Modelling, 186(4):376–388, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001328.

Breckling:2006:IBM

[BMR06] Broder Breckling, Ulrike Middelhoff, and Hauke Reuter. Individual-based models as tools for ecological theory and application: Understanding the emergence of organisational properties in ecological systems. Ecological Modelling, 194(1–3):102–113, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005016.

Barnes:2007:AWG

[BMR07] Belinda Barnes, Karel Mokany, and Michael Roderick. Allocation within a generic scaling framework. Ecological Modelling, 201(2):223–232, February 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004212.

Bohanec:2008:QMA

[BMS^{+}08] Marko Bohanec, Antoine Messéan, Sara Scatasta, Frédérique Angevin, Bryan Griffiths, Paul Henning Krogh, Martin Žnidaršič, and Sašo Džeroski. A qualitative multi-attribute model for economic and ecological assessment of genetically modified crops. Ecological Modelling, 215(1–3):247–261, July 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008001063.

Bourret:2007:MRM

[BMT07] A. Bourret, Y. Martin, and M. Troussellier. Modelling the response of microbial food web to an increase of atmospheric CO_{2} partial pressure in a marine Mediterranean coastal ecosystem (Brusc Lagoon, France). Ecological Modelling, 208(2–4):189–204, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700302X.

Bedoya:2009:IOE

[BNM09] David Bedoya, Vladimir Novotny, and Elias S. Manolakos. Instream and offstream environmental conditions and stream biotic integrity: Importance of scale and site similarities for learning and prediction. Ecological Modelling, 220(19):2393–2406, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003901.

Buckland:2004:SSM

[BNTK04] S. T. Buckland, K. B. Newman, L. Thomas, and N. B. Koesters. State-space models for the dynamics of wild animal populations. Ecological Modelling, 171 (1–2):157–175, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003302.

Borrett:2007:EIS

[BO07] Stuart R. Borrett and Olufemi O. Osidele. Environ indicator sensitivity to flux uncertainty in a phosphorus model of Lake Sidney Lanier, USA. Ecological Modelling, 200(3–4):371–383, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003851.

Birch:2007:RHG

[BOB07] Colin P. D. Birch, Sander P. Oom, and Jonathan A. Beecham. Rectangular and hexagonal grids used for observation, experiment and simulation in ecology. Ecological Modelling, 206(3–4):347–359, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001949.

Boggess:2004:BEH

[Bog04] Carolyn Fonyo Boggess. Biogeoeconomics — energy hierarchy, biogeochemical cycles and money. Ecological Modelling, 178(1–2):39–40, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005180.

Brown:2004:EHT

[BOJ04] Mark T. Brown, Howard T. Odum, and S. E. Jorgensen. Energy hierarchy and transformity in the universe. Ecological Modelling, 178(1–2):17–28, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300512X.

Borer:2006:DAB

[Bor06] Elizabeth T. Borer. Does adding biological detail increase coexistence in an intraguild predation model? Ecological Modelling, 196(3–4):447–461, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000792.

Bormann:2007:RDH

[Bor07] H. Bormann. Recent developments in hydrological modelling towards sustainable catchment management. Ecological Modelling, 209(1):1, November 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003602.

Bonnet:2002:NMP

[BP02] M. P. Bonnet and M. Poulin. Numerical modelling of the planktonic succession in a nutrient-rich reservoir: environmental and physiological factors leading to Microcystis aeruginosa dominance. Ecological Modelling, 156 (2–3):93–112, November 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001321.

Borrett:2003:SPE

[BP03] Stuart R. Borrett and Bernard C. Patten. Structure of pathways in ecological networks: relationships between length and number. Ecological Modelling, 170(2–3):173–184, December 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003002242.

Bonnet:2004:DPB

[BP04] M. P. Bonnet and M. Poulin. DyLEM-1D: a 1D physical and biochemical model for planktonic succession, nutrients and dissolved oxygen cycling: Application to a hyper-eutrophic reservoir. Ecological Modelling, 180(2–3):317–344, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002856.

Bendoricchio:2005:QVE

[BP05] Giuseppe Bendoricchio and Luca Palmeri. Quo vadis ecosystem? Ecological Modelling, 184(1):5–17, May 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005812.

Bode:2007:CCT

[BP07] Michael Bode and Hugh Possingham. Can culling a threatened species increase its chance of persisting? Ecological Modelling, 201(1):11–18, February 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003516.

Bouchard:2008:SEC

[BP08] Mathieu Bouchard and David Pothier. Simulations of the effects of changes in mean fire return intervals on balsam fir abundance, and implications for spruce budworm outbreaks. Ecological Modelling, 218(3–4):207–218, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003281.

Balland:2008:MSH

[BPA08] Vincent Balland, Joseph A. P. Pollacco, and Paul A. Arp. Modeling soil hydraulic properties for a wide range of soil conditions. Ecological Modelling, 219(3–4):300–316, December 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008003505.

Bakran-Petricioli:2006:MSD

[BPAB^{+}06] Tatjana Bakran-Petricioli, Oleg Antonić, Dragan Bukovec, Donat Petricioli, Ivica Janeković, Josip Križan, Vladimir Kušan, and Sandro Dujmović. Modelling spatial distribution of the Croatian marine benthic habitats. Ecological Modelling, 191(1):96–105, January 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003522.

Ball:2000:INL

[BPBF^{+}00] G. R. Ball, D. Palmer-Brown, J. Fuhrer, L. Skärby, B. S. Gimeno, and G. Mills. Identification of non-linear influences on the seasonal ozone dose–response of sensitive and resistant clover clones using artificial neural networks. Ecological Modelling, 129(2–3):153–168, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002349.

Biesinger:2000:DIP

[BPBL00] Zy Biesinger, James Powell, Barbara Bentz, and Jesse Logan. Direct and indirect parametrization of a localized model for the mountain pine beetle — lodgepole pine system. Ecological Modelling, 129(2–3):273–296, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002404.

Belanger:2004:REE

[BPC04] Nicolas Bélanger, David Paré, and François Courchesne. Regression equations for estimating throughfall nutrient fluxes using wet deposition data and their applicability for simulating the soil acid–base status using the dynamic forest soil–atmosphere model SAFE. Ecological Modelling, 175(2):151–167, July 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004745.

Billoir:2007:ILS

[BPC07] Elise Billoir, Alexandre R. R. Péry, and Sandrine Charles. Integrating the lethal and sublethal effects of toxic compounds into the population dynamics of Daphnia magna: a combination of the DEBtox and matrix population models. Ecological Modelling, 203(3–4):204–214, May 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006005758.

Bown:2007:CIV

[BPE^{+}07] James L. Bown, Elizaveta Pachepsky, Alistair Eberst, Ursula Bausenwein, Peter Millard, Geoff R. Squire, and John W. Crawford. Consequences of intraspecific variation for the structure and function of ecological communities: Part 1. Model development and predicted patterns of diversity. Ecological Modelling, 207(2–4):264–276, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007002888.

Brown:2000:FDT

[BPJM00] T. N. Brown, J. Pastor, C. A. Johnston, and H. D. Mooers. A finite difference type algorithm with pro rata resource allocation. Ecological Modelling, 126(1):1–8, February 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001957.

Blackwell:2001:RPP

[BPM01] G. L. Blackwell, M. A. Potter, and E. O. Minot. Rodent and predator population dynamics in an eruptive system. Ecological Modelling, 142(3):227–245, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003271.

Bastianoni:2009:MWR

[BPP09] Simone Bastianoni, Riccardo M. Pulselli, and Federico M. Pulselli. Models of withdrawing renewable and nonrenewable resources based on Odum’s energy systems theory and Daly’s quasi-sustainability principle. Ecological Modelling, 220(16):1926–1930, August 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002671.

Bussenschutt:2000:DAA

[BPW00] Martin Büssenschütt and Claudia Pahl-Wostl. A discrete, allometric approach to the modeling of ecosystem dynamics. Ecological Modelling, 126(1):33–48, February 1, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001933.

Becu:2003:ABS

[BPW^{+}03] N. Becu, P. Perez, A. Walker, O. Barreteau, and C. Le Page. Agent based simulation of a small catchment water management in northern Thailand: Description of the CATCH-SCAPE model. Ecological Modelling, 170(2–3):319–331, December 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003002369.

Bobbin:2001:KDP

[BR01] Jason Bobbin and Friedrich Recknagel. Knowledge discovery for prediction and explanation of blue-green algal dynamics in lakes by evolutionary algorithms. Ecological Modelling, 146(1–3):253–262, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003118.

Bragg:2001:PRI

[Bra01a] Don C. Bragg. Potential relative increment (PRI): a new method to empirically derive optimal tree diameter growth. Ecological Modelling, 137(1):77–92, February 1, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000004336.

Brassil:2001:MTE

[Bra01b] Chad E. Brassil. Mean time to extinction of a metapopulation with an Allee effect. Ecological Modelling, 143 (1–2):9–16, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003519.

Bragg:2004:HAS

[BRC04] Don C. Bragg, David W. Roberts, and Thomas R. Crow. A hierarchical approach for simulating northern forest dynamics. Ecological Modelling, 173(1):31–94, March 15, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003003533.

Breckling:2005:P

[Bre05] B. Breckling. Preface. Ecological Modelling, 186(4): 375, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380005002243.

Billen:2009:CSS

[BRGS09] Norbert Billen, Clara Röder, Thomas Gaiser, and Karl Stahr. Carbon sequestration in soils of SW-Germany as affected by agricultural management — calibration of the EPIC model for regional simulations. Ecological Modelling, 220(1):71–80, January 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004067.

Biesinger:2007:HDM

[BRK07] Zy Biesinger, Edward B. Rastetter, and Bonnie L. Kwiatkowski. Hourly and daily models of active layer evolution in Arctic soils. Ecological Modelling, 206 (1–2):131–146, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001767.

Browder:2004:REM

[Bro04a] Joan A. Browder. Review of Ecological Microcosms, by Robert J. Beyers and Howard T. Odum. Ecological Modelling, 178(1–2):77–80, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005192.

Brown:2004:PWT

[Bro04b] Mark T. Brown. A picture is worth a thousand words: energy systems language and simulation. Ecological Modelling, 178(1–2):83–100, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005209.

Brown:2009:BRJ

[Bro09] Mark T. Brown. Book review: S. E. Jørgensen, B. Fath, S. Bastianoni, J. Marques, F. Muller, S. Nors Nielsen, B. Patten, E. Tiezzi, R. Ulanowicz. A New Ecology: Systems Perspective. 275 pp., Hardbound, illustrated, a few color figures. Ecological Modelling, 220(15):1852–1853, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002658.

Borsuk:2006:ADB

[BRP^{+}06] Mark E. Borsuk, Peter Reichert, Armin Peter, Eva Schager, and Patricia Burkhardt-Holm. Assessing the decline of brown trout (Salmo trutta) in Swiss rivers using a Bayesian probability network. Ecological Modelling, 192 (1–2):224–244, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000500339X.

Brugnach:2005:PLS

[Bru05] Marcela Brugnach. Process level sensitivity analysis for complex ecological models. Ecological Modelling, 187 (2–3):99–120, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000840.

Barbosa:2009:TEF

[BRV09] A. Márcia Barbosa, Raimundo Real, and J. Mario Vargas. Transferability of environmental favourability models in geographic space: the case of the Iberian desman (Galemys pyrenaicus) in Portugal and Spain. Ecological Modelling, 220(5):747–754, March 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005826.

Brown:2005:PEE

[BRW^{+}05] A. R. Brown, A. M. Riddle, I. J. Winfield, J. M. Fletcher, and J. B. James. Predicting the effects of endocrine disrupting chemicals on healthy and disease impacted populations of perch (Perca fluviatilis). Ecological Modelling, 189(3–4):377–395, December 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002073.

Blennow:2004:WSM

[BS04] Kristina Blennow and Ola Sallnäs. WINDA — a system of models for assessing the probability of wind damage to forest stands within a landscape. Ecological Modelling, 175(1):87–99, June 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004708.

Baird:2007:SRP

[BS07] Mark E. Baird and Iain M. Suthers. A size-resolved pelagic ecosystem model. Ecological Modelling, 203(3–4):185–203, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006005679.

Barker:2008:EEP

[BS08a] Daniel Barker and Richard M. Sibly. The effects of environmental perturbation and measurement error on estimates of the shape parameter in the theta-logistic model of population regulation. Ecological Modelling, 219(1–2):170–177, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004109.

Bean:2008:USE

[BS08b] William T. Bean and Eric W. Sanderson. Using a spatially explicit ecological model to test scenarios of fire use by Native Americans: an example from the Harlem Plains, New York, NY. Ecological Modelling, 211(3–4):301–308, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007004747.

Bobba:2000:AEM

[BSB00] A. Ghosh Bobba, Vijay P. Singh, and Lars Bengtsson. Application of environmental models to different hydrological systems. Ecological Modelling, 125(1):15–49, January 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001751.

Bald:2009:SDM

[BSB^{+}09] J. Bald, A. Sinquin, A. Borja, N. Caill-Milly, B. Duclercq, C. Dang, and X. de Montaudouin. A system dynamics model for the management of the Manila clam, Ruditapes philippinarum (Adams and Reeve, 1850) in the Bay of Arcachon (France). Ecological Modelling, 220(21):2828–2837, November 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002348.

Boulain:2007:CSE

[BSG07] Nicolas Boulain, Guillaume Simioni, and Jacques Gignoux. Changing scale in ecological modelling: a bottom up approach with an individual based vegetation model. Ecological Modelling, 203(3–4):257–269, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005795.

Boonen:2002:SSD

[BSJ^{+}02] C. Boonen, R. Samson, K. Janssens, H. Pien, R. Lemeur, and D. Berckmans. Scaling the spatial distribution of photosynthesis from leaf to canopy in a plant growth chamber. Ecological Modelling, 156(2–3):201–212, November 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002001710.

Bernard:2008:MAR

[BSM08] O. Bernard, A. Sciandra, and S. Madani. Multimodel analysis of the response of the coccolithophore Emiliania huxleyi to an elevation of pCO2 under nitrate limitation. Ecological Modelling, 211(3–4):324–338, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004760.

Borsuk:2004:BNE

[BSR04] Mark E. Borsuk, Craig A. Stow, and Kenneth H. Reckhow. A Bayesian network of eutrophication models for synthesis, prediction, and uncertainty analysis. Ecological Modelling, 173(2–3):219–239, April 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003958.

Barnes:2006:MIP

[BSR06] Belinda Barnes, H. S. Sidhu, and S. H. Roxburgh. A model integrating patch dynamics, competing species and the intermediate disturbance hypothesis. Ecological Modelling, 194(4):414–420, April 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005429.

Bohrer:2002:EVR

[BSTS^{+}02] Gil Bohrer, Shachar Shem-Tov, Eric Summer, Keren Or, and David Saltz. The effectiveness of various rabies spatial vaccination patterns in a simulated host population with clumped distribution. Ecological Modelling, 152(2–3):205–211, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002000030.

Barlund:2001:RPB

[BT01] Ilona Bärlund and Sirkka Tattari. Ranking of parameters on the basis of their contribution to model uncertainty. Ecological Modelling, 142(1–2):11–23, August 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002460.

Borga:2005:EIF

[BTdFC05] Marco Borga, Fabrizio Tonelli, Giancarlo dalla Fontana, and Federico Cazorzi. Evaluating the influence of forest roads on shallow landsliding. Ecological Modelling, 187(1):85–98, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000426.

Bach:2006:KCE

[BTPL06] Lars A. Bach, René Thomsen, Cino Pertoldi, and Volker Loeschcke. Kin competition and the evolution of dispersal in an individual-based model. Ecological Modelling, 192(3–4):658–666, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003868.

Brown:2004:EQE

[BU04] Mark T. Brown and Sergio Ulgiati. Energy quality, emergy, and transformity: H. T. Odum’s contributions to quantifying and understanding systems. Ecological Modelling, 178(1–2):201–213, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400119X.

Buldaev:2008:PMO

[Bul08] A. S. Buldaev. Perturbation methods in optimal control problems. Ecological Modelling, 216(2):157–159, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000800135X.

Buzzelli:2008:DAT

[Buz08] Christopher Buzzelli. Development and application of tidal creek ecosystem models. Ecological Modelling, 210 (1–2):127–143, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007003687.

Boumans:2001:NSC

[BVC^{+}01] Roelof M. Boumans, F. Villa, R. Costanza, A. Voinov, H. Voinov, and T. Maxwell. Non-spatial calibrations of a general unit model for ecosystem simulations. Ecological Modelling, 146(1–3):17–32, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002939.

Babu:2005:MTD

[BVD05] M. T. Babu, P. Vethamony, and Ehrlich Desa. Modelling tide-driven currents and residual eddies in the Gulf of Kachchh and their seasonal variability: a marine environmental planning perspective. Ecological Modelling, 184(2–4):299–312, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005551.

Bianchi:2004:MEF

[BvdW04] Felix J. J. A. Bianchi and Wopke van der Werf. Model evaluation of the function of prey in non-crop habitats for biological control by ladybeetles in agricultural landscapes. Ecological Modelling, 171(1–2):177–193, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003314.

Boyce:2002:ERS

[BVNS02] Mark S. Boyce, Pierre R. Vernier, Scott E. Nielsen, and Fiona K. A. Schmiegelow. Evaluating resource selection functions. Ecological Modelling, 157(2–3):281–300, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002002004.

Baldocchi:2001:MCW

[BW01a] Dennis D. Baldocchi and Kell B. Wilson. Modeling CO_{2} and water vapor exchange of a temperate broadleaved forest across hourly to decadal time scales. Ecological Modelling, 142(1–2):155–184, August 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002873.

Bonnet:2001:EDW

[BW01b] M. P. Bonnet and K. Wessen. ELMO, a 3-D water quality model for nutrients and chlorophyll: first application on a lacustrine ecosystem. Ecological Modelling, 141(1–3):19–33, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000004348.

Barry:2002:GAM

[BW02] Simon C. Barry and A. H. Welsh. Generalized additive modelling and zero inflated count data. Ecological Modelling, 157(2–3):179–188, November 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001941.

Brown:2004:CSS

[BW04] Paul Brown and Terence I. Walker. CARPSIM: stochastic simulation modelling of wild carp (Cyprinus carpio L.) population dynamics, with applications to pest control. Ecological Modelling, 176(1–2):83–97, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005581.

BenDor:2009:SPV

[BWAM09] Todd BenDor, James Westervelt, J. P. Aurambout, and William Meyer. Simulating population variation and movement within fragmented landscapes: an application to the gopher tortoise (Gopherus polyphemus). Ecological Modelling, 220(6):867–878, March 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900026X.

Baskent:2001:OOA

[BWJZ01] Emin Z. Baskent, Rick A. Wightman, Glenwood A. Jordan, and Younghua Zhai. Object-oriented abstraction of contemporary forest management design. Ecological Modelling, 143(3):147–164, September 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100254X.

Borrett:2006:IED

[BWPC06] Stuart R. Borrett, Stuart J. Whipple, Bernard C. Patten, and Robert R. Christian. Indirect effects and distributed control in ecosystems: Temporal variation of indirect effects in a seven-compartment model of nitrogen flow in the Neuse River Estuary, USA — time series analysis. Ecological Modelling, 194(1–3):178–188, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005156.

Byers:2000:WAD

[Bye00] John A. Byers. Wind-aided dispersal of simulated bark beetles flying through forests. Ecological Modelling, 125 (2–3):231–243, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380099001878.

Cropper:2004:PDT

[CA04] Wendell P. Cropper and Patti J. Anderson. Population dynamics of a tropical palm: use of a genetic algorithm for inverse parameter estimation. Ecological Modelling, 177(1–2):119–127, September 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001115.

Caplat:2008:SCC

[CAB08] Paul Caplat, Madhur Anand, and Chris Bauch. Symmetric competition causes population oscillations in an individual-based model of forest dynamics. Ecological Modelling, 211(3–4):491–500, March 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005303.

Confalonieri:2009:MME

[CABD09] Roberto Confalonieri, Marco Acutis, Gianni Bellocchi, and Marcello Donatelli. Multi-metric evaluation of the models WARM, CropSyst, and WOFOST for rice. Ecological Modelling, 220(11):1395–1410, June 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001495.

Castanera:2003:SSS

[CAG03] Mónica B. Castañera, Juan P. Aparicio, and Ricardo E. Gürtler. A stage-structured stochastic model of the population dynamics of Triatoma infestans, the main vector of Chagas disease. Ecological Modelling, 162(1–2):33–53, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003885.

Cairns:2005:SCB

[Cai05] David M. Cairns. Simulating carbon balance at treeline for Krummholz and dwarf tree growth forms. Ecological Modelling, 187(2–3):314–328, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000815.

Callies:2005:ISA

[Cal05] Ulrich Callies. Interaction structures analysed from waterquality data. Ecological Modelling, 187(4):475–490, October 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000852.

Calenge:2006:PPA

[Cal06] Clément Calenge. The package “adehabitat” for the R software: a tool for the analysis of space and habitat use by animals. Ecological Modelling, 197(3–4):516–519, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006001414.

Campbell:2004:PWA

[Cam04] Daniel E. Campbell. “Prophet where art thou?” Partnership with nature and other prophecies of H. T. Odum. Ecological Modelling, 178(1–2):255–262, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005222.

Chattopadhyay:2001:PRS

[CB01] J. Chattopadhyay and N. Bairagi. Pelicans at risk in Salton Sea — an eco-epidemiological model. Ecological Modelling, 136(2–3):103–112, January 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003501.

Chaloupka:2007:UBS

[CB07a] Milani Chaloupka and George Balazs. Using Bayesian statespace modelling to assess the recovery and harvest potential of the Hawaiian green sea turtle stock. Ecological Modelling, 205(1–2):93–109, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000634.

Chiang:2007:IBP

[CB07b] Jyh-Min Chiang and Kim J. Brown. Improving the budburst phenology subroutine in the forest carbon model PnET. Ecological Modelling, 205(3–4):515–526, July 24, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007001470.

Cohen:2007:MEH

[CB07c] Matthew J. Cohen and Mark T. Brown. A model examining hierarchical wetland networks for watershed stormwater management. Ecological Modelling, 201(2):179–193, February 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006004200.

Christian:2009:ENA

[CBD^{+}09] Robert R. Christian, Mark M. Brinson, James K. Dame, Galen Johnson, Charles H. Peterson, and Daniel Baird. Ecological network analyses and their use for establishing reference domain in functional assessment of an estuary. Ecological Modelling, 220(22):3113–3122, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004931.

Chaput-Bardy:2009:MES

[CBFLS09] A. Chaput-Bardy, C. Fleurant, C. Lemaire, and J. Secondi. Modelling the effect of in-stream and overland dispersal on gene flow in river networks. Ecological Modelling, 220(24):3589–3598, December 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004128.

Cambier:2007:NOV

[CBMP07] C. Cambier, M. Bousso, D. Masse, and E. Perrier. A new, offer versus demand, modelling approach to assess the impact of micro-organisms spatio-temporal population dynamics on soil organic matter decomposition rates. Ecological Modelling, 209(2–4):301–313, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003444.

Choi:2006:IBM

[CBP^{+}06] Y. H. Choi, D. A. Bohan, R. P. J. Potting, M. A. Semenov, and D. M. Glen. Individual based model of slug population and spatial dynamics. Ecological Modelling, 190(3–4):336–350, January 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002577.

Ceddia:2007:LGF

[CBP07] M. Graziano Ceddia, Mark Bartlett, and Charles Perrings. Landscape gene flow, coexistence and threshold effect: the case of genetically modified herbicide tolerant oilseed rape (Brassica napus). Ecological Modelling, 205(1–2):169–180, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007000889.

Chung:2009:MLB

[CBS09] Eu Gene Chung, Fabián A. Bombardelli, and S. Geoffrey Schladow. Modeling linkages between sediment resuspension and water quality in a shallow, eutrophic, wind-exposed lake. Ecological Modelling, 220(9–10):1251–1265, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009000933.

Ciffroy:2007:PNH

[CBSLS07] P. Ciffroy, K. Beaugelin-Seiller, M. Luck, and F. Siclet. Protection of non-human biota against radiation in freshwater — effect of time dependence in tiered exposure assessment. Ecological Modelling, 206(3–4):360–368, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001962.

Calsina:2000:MAD

[CC00] Àngel Calsina and Sílvia Cuadrado. A model for the adaptive dynamics of the maturation age. Ecological Modelling, 133(1–2):33–43, August 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002787.

Cropper:2005:OSF

[CC05] Wendell P. Cropper and N. B. Comerford. Optimizing simulated fertilizer additions using a genetic algorithm with a nutrient uptake model. Ecological Modelling, 185(2–4):271–281, July 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004006210.

Chen:2006:EFA

[CC06] B. Chen and G. Q. Chen. Ecological footprint accounting based on emergy — a case study of the Chinese society. Ecological Modelling, 198(1–2):101–114, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002018.

Carrer:2005:MPD

[CCBB05] S. Carrer, G. Coffaro, M. Bocci, and A. Barbanti. Modelling partitioning and distribution of micropollutants in the lagoon of Venice: a first step towards a comprehensive ecotoxicological model. Ecological Modelling, 184(1):83–101, May 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004005861.

Chen:2000:ITE

[CCC00] Wenjun Chen, Jing Chen, and Josef Cihlar. An integrated terrestrial ecosystem carbon-budget model based on changes in disturbance, climate, and atmospheric chemistry. Ecological Modelling, 135(1):55–79, November 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003719.

Cabecinha:2004:PSD

[CCC04] Edna Cabecinha, Rui Cortes, and João Alexandre Cabral. Performance of a stochastic-dynamic modelling methodology for running waters ecological assessment. Ecological Modelling, 175(3):303–317, July 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004836.

Cavallaro:2007:SID

[CCG07] Monia Cavallaro, Elisa Canepa, and Emilia Georgieva. The SAFE AIR II dispersion model: Description and statistical evaluation of its dispersion module against wind tunnel data from area sources. Ecological Modelling, 202(3–4):547–558, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006005710.

Colbach:2007:CEA

[CCGMJ07] N. Colbach, B. Chauvel, C. Gauvrit, and N. M. Munier-Jolain. Construction and evaluation of ALOMYSYS modelling the effects of cropping systems on the blackgrass lifecycle: From seedling to seed production. Ecological Modelling, 201(3–4):283–300, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004455.

Chen:2007:RSB

[CCJ07] Baozhang Chen, Jing M. Chen, and Weimin Ju. Remote sensing-based ecosystem–atmosphere simulation scheme (EASS) — model formulation and test with multiple-year data. Ecological Modelling, 209(2–4):277–300, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003420.

Chu:2006:PDS

[CCLS06] Cindy Chu, Nicholas C. Collins, Nigel P. Lester, and Brian J. Shuter. Population dynamics of smallmouth bass in response to habitat supply. Ecological Modelling, 195(3–4):349–362, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005006101.

Childress:2002:ACG

[CCM02] W. Michael Childress, Cade L. Coldren, and Terry McLendon. Applying a complex, general ecosystem model (EDYS) in large-scale land management. Ecological Modelling, 153(1–2):97–108, July 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100504X.

Cecconi:2009:SME

[CCMT09] Giovanni Cecconi, Claudia Cerasuolo, Nadia Marchettini, and Enzo Tiezzi. Salt-marshes as emergent novelties in the Venice lagoon. Ecological Modelling, 220(16):1870–1873, August 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002683.

Chen:2006:EBR

[CCY06] B. Chen, G. Q. Chen, and Z. F. Yang. Exergy-based resource accounting for China. Ecological Modelling, 196(3–4):313–328, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006000718.

Cao:2005:MCF

[CD05] Zhiping Cao and Richard Dawson. Modeling circulation function in agroecosystems. Ecological Modelling, 181(4):557–565, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003977.

Campbell:2007:HST

[CD07] K. A. Campbell and S. M. Dewhurst. A hierarchical simulation-through-optimization approach to forest disturbance modelling. Ecological Modelling, 202(3–4):281–296, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006005345.

Cornejo-Donoso:2008:PTM

[CDA08] Jorge Cornejo-Donoso and Tarsicio Antezana. Preliminary trophic model of the Antarctic Peninsula Ecosystem (Sub-area CCAMLR 48.1). Ecological Modelling, 218(1–2):1–17, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002901.

Chambon-Dubreuil:2006:EAB

[CDAGK06] Estelle Chambon-Dubreuil, Pierre Auger, Jean-Michel Gaillard, and Mohamed Khaladi. Effect of aggressive behaviour on age-structured population dynamics. Ecological Modelling, 193(3–4):777–786, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004497.

Camdevyren:2005:UPC

[ÇDKK05] Handan Çamdevýren, Nilsun Demýr, Arzu Kanik, and Sýddýk Keskýn. Use of principal component scores in multiple linear regression models for prediction of Chlorophyll — a in reservoirs. Ecological Modelling, 181(4):581–589, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004004004.

Charles:2000:ASM

[CDM^{+}00] S. Charles, R. Bravo De La Parra, J.-P. Mallet, H. Persat, and P. Auger. Annual spawning migrations in modelling brown trout population dynamics inside an arborescent river network. Ecological Modelling, 133(1–2):15–31, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002775.

Calenge:2005:KSA

[CDM05] Clément Calenge, A. B. Dufour, and D. Maillard. K-select analysis: a new method to analyse habitat selection in radiotracking studies. Ecological Modelling, 186(2):143–153, August 15, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004006167.

Chapel:2008:DYP

[CDMM08] Laetitia Chapel, Guillaume Deffuant, Sophie Martin, and Christian Mullon. Defining yield policies in a viability approach. Ecological Modelling, 212(1–2):10–15, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700525X.

Cavalett:2006:EAI

[CdQO06] Otávio Cavalett, Júlio Ferraz de Queiroz, and Enrique Ortega. Emergy assessment of integrated production systems of grains, pig and fish in small farms in the South Brazil. Ecological Modelling, 193(3–4):205–224, March 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005003972.

Chong:2005:DFP

[CDV05] Oscar A. G. Chong, Geraldo L. Diniz, and Francisco R. Villatoro. Dispersal of fish populations in dams: modelling and simulation. Ecological Modelling, 186(3):290–298, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005000141.

Conner:2008:ECM

[CEK08] Mary M. Conner, Michael R. Ebinger, and Frederick F. Knowlton. Evaluating coyote management strategies using a spatially explicit, individual-based, socially structured population model. Ecological Modelling, 219(1– 2):234–247, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008004328.

Caniego:2005:MSS

[CEMS05] F. J. Caniego, R. Espejo, M. A. Martín, and F. San José. Multifractal scaling of soil spatial variability. Ecological Modelling, 182(3–4):291–303, March 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001644.

Christiaans:2006:DTP

[CEP06] Thomas Christiaans, Thomas Eichner, and Rüdiger Pethig. On the dynamics of Tschirhart’s predator–prey model. Ecological Modelling, 197(1–2):52–58, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600130X.

Canziani:2008:F

[CF08] Graciela A. Canziani and Rosana Ferrati. Foreword. Ecological Modelling, 214(1):1–2, June 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000197.

Coates:2009:ELZ

[CF09] Michael J. Coates and Andrew M. Folkard. The effects of littoral zone vegetation on turbulent mixing in lakes. Ecological Modelling, 220(20):2714–2726, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004487.

Charles:2004:FAE

[CFCP04] S. Charles, M. Ferreol, A. Chaumot, and A. R. R. Péry. Food availability effect on population dynamics of the midge Chironomus riparius: a Leslie modeling approach. Ecological Modelling, 175(3):217–229, July 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004770.

Carnevale:2008:MAP

[CFPV08] Claudio Carnevale, Giovanna Finzi, Enrico Pisoni, and Marialuisa Volta. Modelling assessment of PM10 exposure control policies in Northern Italy. Ecological Modelling, 217(3–4):219–229, October 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002615.

Christensen:2009:SOP

[CFS09] Villy Christensen, Zach Ferdaña, and Jeroen Steenbeek. Spatial optimization of protected area placement incorporating ecological, social and economical criteria. Ecological Modelling, 220(19):2583–2593, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004384.

Cioffi:2000:RLP

[CG00] F. Cioffi and F. Gallerano. Response of Lake Piediluco to the change of hydrodynamic conditions and nutrient load reductions. Ecological Modelling, 135(2–3):199–229, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003665.

Corani:2006:MSD

[CG06a] Giorgio Corani and Marino Gatto. Model selection in demographic time series using VC-bounds. Ecological Modelling, 191(1):186–195, January 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003856.

Coudun:2006:DSR

[CG06b] Christophe Coudun and Jean-Claude Gégout. The derivation of species response curves with Gaussian logistic regression is sensitive to sampling intensity and curve characteristics. Ecological Modelling, 199(2):164–175, November 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002559.

Cross:2006:AVC

[CG06c] Paul C. Cross and Wayne M. Getz. Assessing vaccination as a control strategy in an ongoing epidemic: Bovine tuberculosis in African buffalo. Ecological Modelling, 196(3–4):494–504, July 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000706.

Corani:2007:EMS

[CG07] Giorgio Corani and Marino Gatto. Erratum to “Model selection in demographic time series using VC-bounds” [Ecol. Modell. 191 (1) (2006) 186–195]. Ecological Modelling, 200(1–2):273–274, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003905.

Chakrabarti:2009:NET

[CG09] C. G. Chakrabarti and Koyel Ghosh. Non-equilibrium thermodynamics of ecosystems: Entropic analysis of stability and diversity. Ecological Modelling, 220(17): 1950–1956, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003202.

Cereghino:2001:SAS

[CGC01] R. Céréghino, J. L. Giraudel, and A. Compin. Spatial analysis of stream invertebrates distribution in the Adour-Garonne drainage basin (France), using Kohonen self organizing maps. Ecological Modelling, 146 (1–3):167–180, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003040.

Chertov:2004:MMF

[CGD04] O. Chertov, A. Gorbushina, and B. Deventer. A model for microcolonial fungi growth on rock surfaces. Ecological Modelling, 177(3–4):415–426, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001486.

Caetano:2009:RCE

[CGdPRY09] Marco Antonio Leonel Caetano, Douglas Francisco Marcolino Gherardi, Gustavo de Paula Ribeiro, and Takashi Yoneyama. Reduction of CO_{2} emission by optimally tracking a pre-defined target. Ecological Modelling, 220 (19):2536–2542, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004050.

Cozar:2005:RSI

[CGG^{+}05] Andrés Cózar, Carlos M. García, José A. Gálvez, Steven A. Loiselle, Luca Bracchini, and Andrea Cognetta. Remote sensing imagery analysis of the lacustrine system of Ibera wetland (Argentina). Ecological Modelling, 186(1):29–41, July 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005000207.

Cozar:2008:SPT

[CGGE08] Andrés Cózar, Carlos M. García, José A. Gálvez, and Fidel Echevarría. Structuring pelagic trophic networks from the biomass size spectra. Ecological Modelling, 215(4):314–324, July 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008001245.

Cranford:2003:MPE

[CGH^{+}03] Peter J. Cranford, Donald C. Gordon, Charles G. Hannah, John W. Loder, Timothy G. Milligan, D. K. Muschenheim, and Y. Shen. Modelling potential effects of petroleum exploration drilling on northeastern Georges Bank scallop stocks. Ecological Modelling, 166(1–2):19–39, August 1, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003001005.

Cozar:2005:SRW

[CGH^{+}05] Andrés Cózar, José A. Gálvez, Vincent Hull, Carlos M. García, and Steven A. Loiselle. Sediment resuspension by wind in a shallow lake of Esteros del Iberá (Argentina): a model based on turbidimetry. Ecological Modelling, 186(1):63–76, July 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000335.

Chow:2005:HEM

[CGHW05] T. Edwin Chow, Karen F. Gaines, Michael E. Hodgson, and Machelle D. Wilson. Habitat and exposure modelling for ecological risk assessment: a case study for the raccoon on the Savannah River Site. Ecological Modelling, 189(1–2):151–167, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001912.

Cariboni:2007:RSA

[CGLS07] J. Cariboni, D. Gatelli, R. Liska, and A. Saltelli. The role of sensitivity analysis in ecological modelling. Ecological Modelling, 203(1–2):167–182, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005734.

Cooch:2003:ADS

[CGR03] Evan G. Cooch, Gilles Gauthier, and Robert F. Rockwell. Apparent differences in stochastic growth rates based on timing of census: a cautionary note. Ecological Modelling, 159(2–3):133–143, January 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002922.

Cartigny:2008:SDS

[CGS08] P. Cartigny, W. Gómez, and H. Salgado. The spatial distribution of small- and large-scale fisheries in a marine protected area. Ecological Modelling, 212(3–4):513–521, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700600X.

Cox:2006:TSS

[CGW^{+}06] G. M. Cox, J. M. Gibbons, A. T. A. Wood, J. Craigon, S. J. Ramsden, and N. M. J. Crout. Towards the systematic simplification of mechanistic models. Ecological Modelling, 198(1–2):240–246, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002031.

Caetano:2008:ORM

[CGY08] Marco Antonio Leonel Caetano, Douglas Francisco Marcolino Gherardi, and Takashi Yoneyama. Optimal resource management control for CO_{2} emission and reduction of the greenhouse effect. Ecological Modelling, 213(1):119–126, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007006229.

Cuddington:2004:IE

[CH04] Kim Cuddington and Alan Hastings. Invasive engineers. Ecological Modelling, 178(3–4):335–347, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004001528.

Chen:2006:NNF

[CH06] Ding-Geng Chen and Steven R. Hare. Neural network and fuzzy logic models for Pacific halibut recruitment analysis. Ecological Modelling, 195(1–2):11–19, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005715.

Chau:2000:DEP

[Cha00] Nguyen Phong Chau. Destabilising effect of periodic harvest on population dynamics. Ecological Modelling, 127(1):1–9, March 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099001908.

Chaloupka:2002:SSM

[Cha02] Milani Chaloupka. Stochastic simulation modelling of southern Great Barrier Reef green turtle population dynamics. Ecological Modelling, 148(1):79–109, February 1, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001004331.

Chan:2007:RCN

[Cha07] Victor Chan. Response to comments by S. Nadarajah. Ecological Modelling, 204(1–2):270, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000488.

Charnell:2008:IBM

[Cha08] Moshi Arthur Charnell. An individual-based model of a tritrophic ecology. Ecological Modelling, 218(3– 4):195–206, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000800327X.

Chang:2009:HEM

[Cha09] Ni-Bin Chang. Handbook of Ecological Modelling and Informatics, 2009, edited by Dr. Sven Erik Jørgensen. Ecological Modelling, 220(17):2119, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003214.

Chen:2005:ECE

[Che05] G. Q. Chen. Exergy consumption of the earth. Ecological Modelling, 184(2–4):363–380, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005769.

Chelle:2006:CPL

[Che06] Michaël Chelle. Could plant leaves be treated as Lambertian surfaces in dense crop canopies to estimate light absorption? Ecological Modelling, 198(1–2):219–228, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006001797.

Cromsigt:2002:UMM

[CHHP02] Joris P. G. M. Cromsigt, John Hearne, Ignas M. A. Heitkönig, and Herbert H. T. Prins. Using models in the management of black rhino populations. Ecological Modelling, 149(1–2):203–211, March 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005245.

Chang:2008:SDB

[CHL08] Y. C. Chang, F. W. Hong, and M. T. Lee. A system dynamic based DSS for sustainable coral reef management in kenting coastal zone, Taiwan. Ecological Modelling, 211(1–2):153–168, February 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004504.

Carrer:2000:MFD

[CHSB00] S. Carrer, B. Halling-Sørensen, and G. Bendoricchio. Modelling the fate of dioxins in a trophic network by coupling an ecotoxicological and an Ecopath model. Ecological Modelling, 126(2–3):201–223, February 28, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002660.

Churkina:2008:MCC

[Chu08] Galina Churkina. Modeling the carbon cycle of urban systems. Ecological Modelling, 216(2):107–113, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001294.

Colasanti:2007:SCA

[CHW07] R. L. Colasanti, R. Hunt, and L. Watrud. A simple cellular automaton model for high-level vegetation dynamics. Ecological Modelling, 203(3–4):363–374, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006259.

Chu:2007:NLS

[CIM07] Peter C. Chu, Leonid M. Ivanov, and Tetyana M. Margolina. On non-linear sensitivity of marine biological models to parameter variations. Ecological Modelling, 206 (3–4):369–382, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001986.

Chen:2007:CEB

[CJ07] G. Q. Chen and Xi Ji. Chemical exergy based evaluation of water quality. Ecological Modelling, 200(1–2):259–268, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006002973.

Chen:2002:MSC

[CJS^{+}02] Changsheng Chen, Rubao Ji, David J. Schwab, Dmitry Beletsky, Gary L. Fahnenstiel, Mingshun Jiang, Thomas H. Johengen, Henry Vanderploeg, Brian Eadie, Judith Wells Budd, Marie H. Bundy, Wayne Gardner, James Cotner, and Peter J. Lavrentyev. A model study of the coupled biological and physical dynamics in Lake Michigan. Ecological Modelling, 152(2–3):145–168, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000261.

Chao:2007:NMW

[CJS^{+}07] Xiaobo Chao, Yafei Jia, F. Douglas Shields, Sam S. Y. Wang, and Charles M. Cooper. Numerical modeling of water quality and sediment related processes. Ecological Modelling, 201(3–4):385–397, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004868.

Chen:2009:EAE

[CJY^{+}09] G. Q. Chen, M. M. Jiang, Z. F. Yang, B. Chen, Xi Ji, and J. B. Zhou. Exergetic assessment for ecological economic system: Chinese agriculture. Ecological Modelling, 220(3):397–410, February 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004808.

Carl:2007:ASA

[CK07a] G. Carl and I. Kühn. Analyzing spatial autocorrelation in species distributions using Gaussian and logit models. Ecological Modelling, 207(2–4):159–170, October 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007002566.

Collier:2007:UAS

[CK07b] Bret A. Collier and David G. Krementz. Uncertainty in age-specific harvest estimates and consequences for white-tailed deer management. Ecological Modelling, 201 (2):194–204, February 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004443.

Chertov:2002:IFS

[CKA^{+}02] O. Chertov, A. Komarov, G. Andrienko, N. Andrienko, and P. Gatalsky. Integrating forest simulation models and spatial–temporal interactive visualisation for decision making at landscape level. Ecological Modelling, 148(1):47–65, February 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004379.

Crepinsek:2006:MWV

[ČKBB06] Zalika Črepinšek, Lučka Kajfež-Bogataj, and Klemen Bergant. Modelling of weather variability effect on fitophenology. Ecological Modelling, 194(1–3):256–265, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005235.

Costa:2000:AGC

[CKBH00] M. I. S. Costa, E. Kaszkurewicz, A. Bhaya, and L. Hsu. Achieving global convergence to an equilibrium population in predator–prey systems by the use of a discontinuous harvesting policy. Ecological Modelling, 128(2–3):89–99, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002203.

Chertov:2006:UFE

[CKL^{+}06] O. Chertov, A. Komarov, A. Loukianov, A. Mikhailov, M. Nadporozhskaya, and E. Zubkova. The use of forest ecosystem model EFIMOD for research and practical implementation at forest stand, local and regional levels. Ecological Modelling, 194(1–3):227–232, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500520X.

Chertov:2001:RMF

[CKN^{+}01] O. G. Chertov, A. S. Komarov, M. Nadporozhskaya, S. S. Bykhovets, and S. L. Zudin. ROMUL — a model of forest soil organic matter dynamics as a substantial tool for forest ecosystem modeling. Ecological Modelling, 138(1–3):289–308, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004099.

Chon:2001:PST

[CKP^{+}01] Tae-Soo Chon, Inn-Sil Kwak, Young-Seuk Park, Tae-Hyung Kim, and YooShin Kim. Patterning and short-term predictions of benthic macroinvertebrate community dynamics by using a recurrent artificial neural network. Ecological Modelling, 146(1–3):181–193, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003052.

Chumachenko:2003:SML

[CKPP03] S. I. Chumachenko, V. N. Korotkov, M. M. Palenova, and D. V. Politov. Simulation modelling of long-term stand dynamics at different scenarios of forest management for coniferous–broad-leaved forests. Ecological Modelling, 170(2–3):345–361, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002382.

Cropper:2006:ISD

[CL06] Wendell P. Cropper and E. Louise Loudermilk. The interaction of seedling density dependence and fire in a matrix population model of longleaf pine (Pinus palustris). Ecological Modelling, 198(3–4):487–494, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002614.

Chefaoui:2008:AEP

[CL08a] Rosa M. Chefaoui and Jorge M. Lobo. Assessing the effects of pseudo-absences on predictive distribution model performance. Ecological Modelling, 210(4):478–486, February 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700419X.

Chen:2008:EHD

[CL08b] Ling-Ling Chen and Zhen-Shan Lin. The effect of habitat destruction on metapopulations with the Allee-like effect: a study case of Yancheng in Jiangsu Province, China. Ecological Modelling, 213(3–4):356–364, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000033.

Caron-Lormier:2009:HMW

[CLBH^{+}09] Geoffrey Caron-Lormier, David A. Bohan, Cathy Hawes, Alan Raybould, Alison J. Haughton, and Roger W. Humphry. How might we model an ecosystem? Ecological Modelling, 220(17):1935–1949, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002816.

Caron-Lormier:2008:ASU

[CLHB^{+}08] Geoffrey Caron-Lormier, Roger W. Humphry, David A. Bohan, Cathy Hawes, and Pernille Thorbek. Asynchronous and synchronous updating in individual-based models. Ecological Modelling, 212(3–4):522–527, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006011.

Cabecinha:2009:MSA

[CLM^{+}09] Edna Cabecinha, Martinho Lourenço, João Paulo Moura, Miguel Ângelo Pardal, and João Alexandre Cabral. A multiscale approach to modelling spatial and dynamic ecological patterns for reservoir’s water quality management. Ecological Modelling, 220(19):2559–2569, October 10, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009003950.

Chen:2008:ISP

[CLTH08] M. Chen, S. Liu, L. L. Tieszen, and D. Y. Hollinger. An improved state-parameter analysis of ecosystem models using data assimilation. Ecological Modelling, 219 (3–4):317–326, December 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008003530.

Chen:2003:IDM

[CM03] Qiuwen Chen and Arthur E. Mynett. Integration of data mining techniques and heuristic knowledge in fuzzy logic modelling of eutrophication in Taihu Lake. Ecological Modelling, 162(1–2):55–67, April 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003897.

Chen:2004:PPG

[CM04] Qiuwen Chen and Arthur E. Mynett. Predicting Phaeocystis globosa bloom in Dutch coastal waters by decision trees and nonlinear piecewise regression. Ecological Modelling, 176(3–4):277–290, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000493.

Chan:2005:UGD

[CM05a] Victor Chan and Robin A. Matthews. Using the generalized F distribution to model limnetic temperature profile and estimate thermocline depth. Ecological Modelling, 188(2–4):374–385, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002231.

Chen:2005:OHS

[CM05b] Sanping Chen and Randall S. Morley. Observed herd size and animal association. Ecological Modelling, 189 (3–4):425–435, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002103.

Caruso:2006:NFG

[CM06a] Tancredi Caruso and Massimo Migliorini. A new formulation of the geometric series with applications to oribatid (Acari, Oribatida) species assemblages from humandisturbed Mediterranean areas. Ecological Modelling, 195(3–4):402–406, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005006289.

Chen:2006:MAB

[CM06b] Qiuwen Chen and Arthur E. Mynett. Modelling algal blooms in the Dutch coastal waters by integrated numerical and fuzzy cellular automata approaches. Ecological Modelling, 199(1):73–81, November 1, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600281X.

Caruso:2007:ENF

[CM07a] Tancredi Caruso and Massimo Migliorini. Erratum to “A new formulation of the geometric series with applications to oribatid (Acari, Oribatida) species assemblages from humandisturbed Mediterranean areas” [Ecol. Model. 195 (3–4) (2006) 402–406]. Ecological Modelling, 204(1–2):277, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006006636.

Clark:2007:MTA

[CM07b] M. E. Clark and T. E. Martin. Modeling tradeoffs in avian life history traits and consequences for population growth. Ecological Modelling, 209(2–4):110–120, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003195.

Couto:2008:IBS

[CM08] Paula R. L. Couto and Sandra M. C. Malta. Interaction between sorption and biodegradation processes in the contaminant transport. Ecological Modelling, 214(1):65–73, June 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000276.

Coates:2009:EEF

[CM09] Michael J. Coates and Julie Mondon. Effect of environmental flows on deep, anoxic pools. Ecological Modelling, 220(13–14):1643–1651, July 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001896.

Chaumot:2006:FSM

[CMA^{+}06] A. Chaumot, N. Milioni, A. Abdoli, D. Pont, and S. Charles. First step of a modeling approach to evaluate spatial heterogeneity in a fish (Cottus gobio) population dynamics. Ecological Modelling, 197(3–4):263–273, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001451.

Chiesi:2002:CAF

[CMB^{+}02] M. Chiesi, F. Maselli, M. Bindi, L. Fibbi, L. Bonora, A. Raschi, R. Tognetti, J. Cermak, and N. Nadezhdina. Calibration and application of FOREST-BGC in a Mediterranean area by the use of conventional and remote sensing data. Ecological Modelling, 154(3):251–262, September 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002000571.

Confalonieri:2005:AMW

[CMB05] R. Confalonieri, L. Mariani, and S. Bocchi. Analysis and modelling of water and near water temperatures in flooded rice (Oryza sativa L.). Ecological Modelling, 183(2–3):269–280, April 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004831.

Colbach:2004:SAM

[CMCD04] N. Colbach, N. Molinari, and C. Clermont-Dauphin. Sensitivity analyses for a model simulating demography and genotype evolutions with time: Application to GeneSys modelling gene flow between rape seed varieties and volunteers. Ecological Modelling, 179(1):91–113, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002649.

Cai:2006:MPP

[CMD06] T. T. Cai, C. L. Montague, and J. S. Davis. The maximum power principle: an empirical investigation. Ecological Modelling, 190(3–4):317–335, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002553.

Chapelle:2000:MNP

[CMDP^{+}00] Annie Chapelle, Alain Ménesguen, Jean-Marc Deslous-Paoli, Philippe Souchu, Nabila Mazouni, André Vaquer, and Bertrand Millet. Modelling nitrogen, primary production and oxygen in a Mediterranean lagoon. Impact of oysters farming and inputs from the watershed. Ecological Modelling, 127(2–3):161–181, March 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002069.

Ciffroy:2000:MST

[CMG00] P. Ciffroy, C. Moulin, and J. Gailhard. A model simulating the transport of dissolved and particulate copper in the Seine Rivero. Ecological Modelling, 127(2–3):99–117, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001829.

Colbach:2009:SSM

[CML09] Nathalie Colbach, Hervé Monod, and Claire Lavigne. A simulation study of the medium-term effects of field patterns on cross-pollination rates in oilseed rape (Brassica napus L.). Ecological Modelling, 220(5):662–672, March 10, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008005760.

Chen:2002:ACA

[CMM02] Q. Chen, A. E. Mynett, and A. W. Minns. Application of cellular automata to modelling competitive growths of two underwater species Chara aspera and Potamogeton pectinatus in Lake Veluwe. Ecological Modelling, 147(3):253–265, January 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001004288.

Chiesi:2007:ABB

[CMM^{+}07] M. Chiesi, F. Maselli, M. Moriondo, L. Fibbi, M. Bindi, and S. W. Running. Application of BIOME-BGC to simulate Mediterranean forest processes. Ecological Modelling, 206(1–2):179–190, August 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001780.

Corbeels:2005:PBMa

[CMPO05a] M. Corbeels, R. E. McMurtrie, D. A. Pepper, and A. M. O’Connell. A process-based model of nitrogen cycling in forest plantations: Part I. Structure, calibration and analysis of the decomposition model. Ecological Modelling, 187(4):426–448, October 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000785.

Corbeels:2005:PBMb

[CMPO05b] M. Corbeels, R. E. McMurtrie, D. A. Pepper, and A. M. O’Connell. A process-based model of nitrogen cycling in forest plantations: Part II. Simulating growth and nitrogen mineralisation of Eucalyptus globulus plantations in south-western Australia. Ecological Modelling, 187 (4):449–474, October 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003376.

Carusela:2009:CPC

[CMR09] M. Florencia Carusela, Fernando R. Momo, and Lilia Romanelli. Competition, predation and coexistence in a three trophic system. Ecological Modelling, 220(19): 2349–2352, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004001.

Chalvet-Monfray:2007:DMA

[CMSB07] Karine Chalvet-Monfray, Philippe Sabatier, and Dominique J. Bicout. Downscaling modeling of the aggressiveness of mosquitoes vectors of diseases. Ecological Modelling, 204(3–4):540–546, June 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000476.

Crispi:2001:NCM

[CMSC01] G. Crispi, R. Mosetti, C. Solidoro, and A. Crise. Nutrients cycling in Mediterranean basins: the role of the biological pump in the trophic regime. Ecological Modelling, 138(1–3):101–114, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003963.

Chave:2007:CSD

[CN07a] Jérôme Chave and Natalia Norden. Changes of species diversity in a simulated fragmented neutral landscape. Ecological Modelling, 207(1):3–10, September 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001500.

Cropp:2007:IPP

[CN07b] Roger Cropp and John Norbury. Investigations into a plankton population model: Mortality and its importance in climate change scenarios. Ecological Modelling, 201(2):97–117, February 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004108.

Cropp:2009:SPP

[CN09] Roger Cropp and John Norbury. Simple predator–prey interactions control dynamics in a plankton food web model. Ecological Modelling, 220(13–14):1552–1565, July 17, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009002324.

Cornick:2006:ACB

[CNG06] Leslie A. Cornick, William Neill, and William E. Grant. Assessing competition between Steller sea lions and the commercial groundfishery in Alaska: a bioenergetics modelling approach. Ecological Modelling, 199(1):107–114, November 1, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003115.

Chuang:2006:PMM

[COB^{+}06] Yao-Li Chuang, Ram Oren, Andrea L. Bertozzi, Nathan Phillips, and Gabriel G. Katul. The porous media model for the hydraulic system of a conifer tree: Linking sap flux data to transpiration rate. Ecological Modelling, 191(3–4):447–468, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003054.

Cai:2004:MEP

[COC04] T. T. Cai, T. W. Olsen, and D. E. Campbell. Maximum (em)power: a foundational principle linking man and nature. Ecological Modelling, 178(1–2):115–119, October 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003005210.

Cohen:2009:FPM

[Coh09] Yosef Cohen. Fitting probability models to population dynamics data. Ecological Modelling, 220(13–14):1613–1619, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002269.

Comar:2004:AHO

[Com04] Vito Comar. Applying H. T. Odum’s concepts and principles in developing countries. Ecological Modelling, 178(1–2):171–175, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003005234.

Corani:2005:AQP

[Cor05] Giorgio Corani. Air quality prediction in Milan: feedforward neural networks, pruned neural networks and lazy learning. Ecological Modelling, 185(2–4):513–529, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000165.

Couto:2003:AAS

[Cou03] Paula Couto. Assessing the accuracy of spatial simulation models. Ecological Modelling, 167(1–2):181–198, September 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001765.

Cramer:2001:MFP

[CP01] Patricia C. Cramer and Kenneth M. Portier. Modeling Florida panther movements in response to human attributes of the landscape and ecological settings. Ecological Modelling, 140(1–2):51–80, May 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100268X.

Chattopadhyay:2002:VIP

[CP02] J. Chattopadhyay and S. Pal. Viral infection on phytoplankton– zooplankton system — a mathematical model. Ecological Modelling, 151(1):15–28, May 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100415X.

Christensen:2004:PFT

[CP04] Villy Christensen and Daniel Pauly. Placing fisheries in their ecosystem context, an introduction. Ecological Modelling, 172(2–4):103–107, March 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003648.

Cowden:2009:MMS

[CP09] Charles C. Cowden and Chris J. Peterson. A multimutualist simulation: Applying biological market models to diverse mycorrhizal communities. Ecological Modelling, 220(12):1522–1533, June 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002270.

Ciavatta:2008:EPP

[CPB^{+}08] Stefano Ciavatta, Roberto Pastres, Christian Badetti, Giorgio Ferrari, and M. B. Beck. Estimation of phytoplanktonic production and system respiration from data collected by a real-time monitoring network in the Lagoon of Venice. Ecological Modelling, 212(1–2):28–36, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700508X.

Carlsson:2000:CMM

[CPH00] Lennart Carlsson, Johan Persson, and Lars Håkanson. Corrigendum to “A management model to predict seasonal variability in oxygen concentration and oxygen consumption in thermally stratified coastal waters”: [Ecol. Model. 119 (1999) 117–134]. Ecological Modelling, 127 (2–3):309–311, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438009900215X.

Choi:2006:SDS

[CPJ06] Jung Hyun Choi, Seok Soon Park, and Peter R. Jaffé. Simulating the dynamics of sulfur species and zinc in wetland sediments. Ecological Modelling, 199(3):315–323, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002328.

Chon:2000:DTP

[CPP00] Tae-Soo Chon, Young-Seuk Park, and June Ho Park. Determining temporal pattern of community dynamics by using unsupervised learning algorithms. Ecological Modelling, 132(1–2):151–166, July 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003124.

Coll:2009:DCN

[CPT09] Marta Coll, Isabel Palomera, and Sergi Tudela. Decadal changes in a NW Mediterranean Sea food web in relation to fishing exploitation. Ecological Modelling, 220 (17):2088–2102, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003093.

Coll:2008:FWD

[CPTD08] Marta Coll, Isabel Palomera, Sergi Tudela, and Michael Dowd. Food-web dynamics in the South Catalan Sea ecosystem (NW Mediterranean) for 1978–2003. Ecological Modelling, 217(1–2):95–116, September 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002925.

Carnevale:2007:SEO

[CPV07] Claudio Carnevale, Enrico Pisoni, and Marialuisa Volta. Selecting effective ozone exposure control policies solving a two-objective problem. Ecological Modelling, 204(1–2): 93–103, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380006006624.

Chen:2007:SAS

[CQ07] G. Q. Chen and Z. H. Qi. Systems account of societal exergy utilization: China 2003. Ecological Modelling, 208(2–4):102–118, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002426.

Choi:2003:MMP

[CR03] Won Il Choi and Mun Il Ryoo. A matrix model for predicting seasonal fluctuations in field populations of Paronychiurus kimi (Collembola: Onychiruidae). Ecological Modelling, 162(3):259–265, April 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003848.

Cuadra:2006:EEP

[CR06] M. Cuadra and T. Rydberg. Emergy evaluation on the production, processing and export of coffee in Nicaragua. Ecological Modelling, 196(3–4):421–433, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600072X.

Chakraborty:2008:NLT

[CRC08] Subhendu Chakraborty, Shovonlal Roy, and J. Chattopadhyay. Nutrient-limited toxin production and the dynamics of two phytoplankton in culture media: a mathematical model. Ecological Modelling, 213(2):191–201, May 10, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000700628X.

Chubaty:2009:DHS

[CRL09] Alex M. Chubaty, Bernard D. Roitberg, and Chao Li. A dynamic host selection model for mountain pine beetle, Dendroctonus ponderosae Hopkins. Ecological Modelling, 220(9–10):1241–1250, May 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900088X.

Christensen:2009:BFB

[CRM09] Morten Christensen, Santosh Rayamajhi, and Henrik Meilby. Balancing fuelwood and biodiversity concerns in rural Nepal. Ecological Modelling, 220(4):522–532, February 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000800505X.

Cairns:2007:CMP

[CRT07] B. J. Cairns, J. V. Ross, and T. Taimre. A comparison of models for predicting population persistence. Ecological Modelling, 201(1):19–26, February 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003528.

Colbach:2001:BAM

[CS01] N. Colbach and I. Sache. Blackgrass (Alopecurus myosuroides huds.) seed dispersal from a single plant and its consequences on weed infestation. Ecological Modelling, 139(2–3):201–219, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001002368.

Chattopadhyay:2003:COP

[CS03] J. Chattopadhyay and R. R. Sarkar. Chaos to order: preliminary experiments with a population dynamics models of three trophic levels. Ecological Modelling, 163(1–2):45–50, May 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002003812.

Chikumbo:2007:SBA

[CS07] Oliver Chikumbo and Greg A. Steward. A stand basal area model for plantation grown New Zealand kauri. Ecological Modelling, 209(2–4):367–376, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700350X.

Cossarini:2008:GSA

[CS08] Gianpiero Cossarini and Cosimo Solidoro. Global sensitivity analysis of a trophodynamic model of the Gulf of Trieste.

Ecological Modelling, 212(1–2):16–27, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005224.

Chattopadhyay:2003:PRS

[CSB03] J. Chattopadhyay, P. D. N. Srinivasu, and N. Bairagi. Pelicans at risk in Salton Sea — an eco-epidemiological model — II. Ecological Modelling, 167(1–2):199–211, September 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000300187X.

Chattopadhyay:2004:EEF

[CSCB04] J. Chattopadhyay, R. R. Sarkar, S. Chaki, and S. Bhattacharya. Effects of environmental fluctuations on the occurrence of malignant malaria — a model based study. Ecological Modelling, 177(1–2):179–192, September 15, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004001152.

Cianelli:2009:IBA

[CSdZ09] Daniela Cianelli, Luciana Sabia, Maurizio Ribera d’Alcalà, and Enrico Zambianchi. An individual-based analysis of the dynamics of two coexisting phytoplankton species in the mixed layer. Ecological Modelling, 220(19):2380–2392, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009004177.

Colbert:2004:ITB

[CSF^{+}04] J. J. Colbert, Michael Schuckers, Desta Fekedulegn, James Rentch, Máirtín MacSiúrtáin, and Kurt Gottschalk. Individual tree basal-area growth parameter estimates for four models. Ecological Modelling, 174(1–2):115–126, May 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000400002X.

Chattopadhyay:2002:RIP

[CSG02] J. Chattopadhyay, R. R. Sarkar, and G. Ghosal. Removal of infected prey prevent limit cycle oscillations in an infected prey–predator system — a mathematical study. Ecological Modelling, 156(2–3):113–121, November 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001333.

Chen:2008:FDM

[CSHY08] Jun Chen, Masae Shiyomi, Yoshimichi Hori, and Yasuo Yamamura. Frequency distribution models for spatial patterns of vegetation abundance. Ecological Modelling, 211(3–4):403–410, March 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004826.

Charles:2008:IBM

[CSKP08] Sandrine Charles, Fabien Subtil, Janice Kielbassa, and Didier Pont. An individual-based model to describe a bullhead population dynamics including temperature variations. Ecological Modelling, 215(4):377–392, July 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800183X.

Coll:2006:CTF

[CSM^{+}06] Marta Coll, Lynne J. Shannon, Coleen L. Moloney, Isabel Palomera, and Sergi Tudela. Comparing trophic flows and fishing impacts of a NW Mediterranean ecosystem with coastal upwelling systems by means of standardized models and indicators. Ecological Modelling, 198 (1–2):53–70, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001724.

Callies:2008:CUA

[CSR08] U. Callies, M. Scharfe, and M. Ratto. Calibration and uncertainty analysis of a simple model of silicalimited diatom growth in the Elbe River. Ecological Modelling, 213(2):229–244, May 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006503.

Cabecinha:2007:ASD

[CSSCC07] Edna Cabecinha, Pedro Silva-Santos, Rui Cortes, and João Alexandre Cabral. Applying a stochastic-dynamic methodology (StDM) to facilitate ecological monitoring of running waters, using selected trophic and taxonomic metrics as state variables. Ecological Modelling, 207 (2–4):109–127, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007002335.

Canu:2003:MRL

[CSU03] Donata Melaku Canu, Cosimo Solidoro, and Georg Umgiesser. Modelling the responses of the Lagoon of Venice ecosystem to variations in physical forcings. Ecological Modelling, 170(2–3):265–289, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002333.

Canu:2004:EMR

[CSU04] Donata Melaku Canu, Cosimo Solidoro, and Georg Umgiesser. Erratum to “Modelling the responses of the Lagoon of Venice ecosystem to variations in physical forcings”: [Ecol. Model. 170 (2003) 265–289]. Ecological Modelling, 175(2):197–216, July 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400050X.

Chubarenko:2001:MMM

[CT01] Irina Chubarenko and Irina Tchepikova. Modelling of man-made contribution to salinity increase into the Vistula Lagoon (Baltic Sea). Ecological Modelling, 138 (1–3):87–100, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003951.

Changjie:2007:AMD

[CT07a] Jin Changjie and Pei Tiefan. An application of Markov decision programming to ecologico-cybernetics — a decision method applied to the antifrost of winter wheat. Ecological Modelling, 200(3–4):452–458, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003887.

Courtois:2007:GIC

[CT07b] Pierre Courtois and Tarik Tazdaït. Games of influence in climate change negotiations: Modelling interactions. Ecological Modelling, 204(3–4):301–314, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000221.

Corson:2003:SDC

[CTG03] M. S. Corson, P. D. Teel, and W. E. Grant. Simulating detection of cattle-fever tick (Boophilus spp.) infestations in rotational grazing systems. Ecological Modelling, 167(3):277–286, September 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001947.

Corson:2004:MIP

[CTG04] Michael S. Corson, Pete D. Teel, and William E. Grant. Microclimate influence in a physiological model of cattlefever tick (Boophilus spp.) population dynamics. Ecological Modelling, 180(4):487–514, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002972.

Coquillard:2000:SMA

[CTH^{+}00] P. Coquillard, T. Thibaut, D. R. C. Hill, J. Gueugnot, C. Mazel, and Y. Coquillard. Simulation of the mollusc ascoglossa Elysia subornata population dynamics: application to the potential biocontrol of Caulerpa taxifolia growth in the Mediterranean Sea. Ecological Modelling, 135(1):1–16, November 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003422.

Cucco:2006:MVL

[CU06] Andrea Cucco and Georg Umgiesser. Modeling the Venice lagoon residence time. Ecological Modelling, 193(1–2):34–51, March 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005004552.

Cucco:2009:ELT

[CUF^{+}09] Andrea Cucco, Georg Umgiesser, Cristian Ferrarin, Angelo Perilli, Donata Melaku Canu, and Cosimo Solidoro. Eulerian and Lagrangian transport time scales of a tidal active coastal basin. Ecological Modelling, 220(7):913–922, April 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009000301.

Canu:2001:STS

[CUS01] D. Melaku Canu, G. Umgiesser, and C. Solidoro. Shortterm simulations under winter conditions in the lagoon of Venice: a contribution to the environmental impact assessment of temporary closure of the inlets. Ecological Modelling, 138(1–3):215–230, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004038.

Cianelli:2009:ZEP

[CUSZ09] Daniela Cianelli, Marco Uttieri, J. Rudi Strickler, and Enrico Zambianchi. Zooplankton encounters in patchy particle distributions. Ecological Modelling, 220(5):596–604, March 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005048.

Costanza:2001:MEE

[CV01] Robert Costanza and Alexey Voinov. Modeling ecological and economic systems with STELLA: Part III. Ecological Modelling, 143(1–2):1–7, October 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003581.

Castella:2007:CPO

[CV07] Jean-Christophe Castella and Peter H. Verburg. Combination of process-oriented and pattern-oriented models of land-use change in a mountain area of Vietnam. Ecological Modelling, 202(3–4):410–420, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005436.

Carroll:2001:UAC

[CW01] R. W. H. Carroll and J. J. Warwick. Uncertainty analysis of the Carson River mercury transport model. Ecological Modelling, 137(2–3):211–224, February 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004385.

Christensen:2004:EEM

[CW04] Villy Christensen and Carl J. Walters. Ecopath with Ecosim: methods, capabilities and limitations. Ecological Modelling, 172(2–4):109–139, March 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300365X.

Christensen:2009:DDM

[CWA^{+}09] Villy Christensen, Carl J. Walters, Robert Ahrens, Jacqueline Alder, Joe Buszowski, Line Bang Christensen, William W. L. Cheung, John Dunne, Rainer Froese, Vasiliki Karpouzi, Kristin Kaschner, Kelly Kearney, Sherman Lai, Vicki Lam, Maria L. D. Palomares, Aja Peters-Mason, Chiara Piroddi, Jorge L. Sarmiento, Jeroen Steenbeek, Rashid Sumaila, Reg Watson, Dirk Zeller, and Daniel Pauly. Database-driven models of the world’s large marine ecosystems. Ecological Modelling, 220(17):1984–1996, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002841.

Coops:2001:CPN

[CWBR01] N. C. Coops, R. H. Waring, S. R. Brown, and S. W. Running. Comparisons of predictions of net primary production and seasonal patterns in water use derived with two forest growth models in Southwestern Oregon. Ecological Modelling, 142(1–2):61–81, August 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002794.

Chubarenko:2001:WDC

[CWCH01] Boris V. Chubarenko, Yongqi Wang, Irina P. Chubarenko, and Kolumban Hutter. Wind-driven current simulations around the Island Mainau (Lake Constance). Ecological Modelling, 138(1–3):55–73, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003938.

Chen:2009:AEL

[CWCH09] Weiping Chen, Laosheng Wu, Andrew C. Chang, and Zhenan Hou. Assessing the effect of long-term crop cultivation on distribution of Cd in the root zone. Ecological Modelling, 220(15):1836–1843, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002804.

Clarke:2003:RMP

[CWF03] Ralph T. Clarke, John F. Wright, and Mike T. Furse. RIV-PACS models for predicting the expected macroinvertebrate fauna and assessing the ecological quality of rivers. Ecological Modelling, 160(3):219–233, February 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002557.

Carroll:2000:SMT

[CWH^{+}00] R. W. H. Carroll, J. J. Warwick, K. J. Heim, J. C. Bonzongo, J. R. Miller, and W. B. Lyons. Simulation of mercury transport and fate in the Carson River, Nevada. Ecological Modelling, 125(2–3):255–278, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001866.

Coops:2005:APF

[CWL05] Nicholas C. Coops, Richard H. Waring, and Beverly E. Law. Assessing the past and future distribution and productivity of ponderosa pine in the Pacific northwest using a process model, 3-PG. Ecological Modelling, 183(1):107–124, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000400434X.

Coops:2009:CGP

[CWS09] Nicholas C. Coops, Richard H. Waring, and Todd A. Schroeder. Combining a generic process-based productivity model and a statistical classification method to predict the presence and absence of tree species in the Pacific Northwest, U.S.A. Ecological Modelling, 220(15):1787–1796, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002907.

Cornwell:2001:PEL

[CWWS01] Charles F. Cornwell, Luc T. Wille, Yegang Wu, and Fred H. Sklar. Parallelization of an ecological landscape model by functional decomposition. Ecological Modelling, 144(1):13–20, October 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003441.

Cao:2004:EGM

[CYHK04] Y. Q. Cao, Y. S. Yang, K. R. Hu, and R. M. Kalin. Ecological and geochemical modelling of hydrogeological system with particular connection to human health. Ecological Modelling, 174(4):375–385, June 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004587.

Chuanyan:2005:MMT

[CZG05] Zhao Chuanyan, Nan Zhongren, and Cheng Guodong. Methods for modelling of temporal and spatial distribution of air temperature at landscape scale in the southern qilian mountains, China. Ecological Modelling, 189 (1–2):209–220, November 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001961.

Chen:2005:ITP

[CZL05] Xiongwen Chen, Xin-Shi Zhang, and Bai-Lian Li. Influence of Tibetan Plateau on vegetation distributions in East Asia: a modeling perspective. Ecological Modelling, 181(1):79–86, January 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400359X.

Duru:2009:MAG

[DAC^{+}09] M. Duru, M. Adam, P. Cruz, G. Martin, P. Ansquer, C. Ducourtieux, C. Jouany, J. P. Theau, and J. Viegas. Modelling above-ground herbage mass for a wide range of grassland community types. Ecological Modelling, 220(2):209–225, January 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004560.

Dieguez-Aranda:2006:DGM

[DADGA06] Ulises Diéguez-Aranda, Fernando Castedo Dorado, Juan Gabriel Álvarez González, and Alberto Rojo Alboreca. Dynamic growth model for Scots pine (Pinus sylvestris L.) plantations in Galicia (north-western Spain). Ecological Modelling, 191(2):225–242, January 31, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002747.

Doce:2008:FEM

[DAdSC08] Ana Paula C. Rio Doce, Regina C. Almeida, and Michel I. da S. Costa. A finite element method to solve a population dynamics stage-structured model of intertidal barnacles. Ecological Modelling, 214(1):26–38, June 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000227.

Damgaard:2003:MPC

[Dam03] Christian Damgaard. Modeling plant competition along an environmental gradient. Ecological Modelling, 170 (1):45–53, December 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002990.

Damgaard:2008:MPP

[Dam08] Christian Damgaard. Modelling pin-point plant cover data along an environmental gradient. Ecological Modelling, 214(2–4):404–410, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001531.

Driezen:2007:ELC

[DAR^{+}07] Kassandra Driezen, Frank Adriaensen, Carlo Rondinini, C. Patrick Doncaster, and Erik Matthysen. Evaluating leastcost model predictions with empirical dispersal data: a casestudy using radiotracking data of hedgehogs (Erinaceus europaeus). Ecological Modelling, 209(2–4):314–322, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007003468.

Dumont:2008:FIA

[DBB^{+}08] Egon Dumont, Evert Jan Bakker, Lex Bouwman, Carolien Kroeze, Rik Leemans, and Alfred Stein. A framework to identify appropriate spatial and temporal scales for modeling N flows from watersheds. Ecological Modelling, 212(3–4):256–272, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005340.

Duchesne:2003:SUA

[DBBS03] S. Duchesne, P. Boyer, and K. Beaugelin-Seiller. Sensitivity and uncertainty analysis of a model computing radionuclides transfers in fluvial ecosystems (CASTEAUR): application to ^{137}Cs accumulation in chubs. Ecological Modelling, 166(3):257–276, August 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001650.

Davi:2008:MLM

[DBD^{+}08] H. Davi, C. Barbaroux, E. Dufrêne, C. François, P. Montpied, N. Bréda, and F. Badeck. Modelling leaf mass per area in forest canopy as affected by prevailing radiation conditions. Ecological Modelling, 211(3–4):339–349, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007004772.

deBarros:2004:MSN

[dBWG04] Inácio de Barros, Jimmy R. Williams, and Thomas Gaiser. Modeling soil nutrient limitations to crop production in semiarid NE of Brazil with a modified EPIC version: I. Changes in the source code of the model. Ecological Modelling, 178(3–4):441–456, November 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002030.

deBarros:2005:MSN

[dBWG05] Inácio de Barros, Jimmy R. Williams, and Thomas Gaiser. Modeling soil nutrient limitations to crop production in semiarid NE of Brazil with a modified EPIC version: II: Field test of the model. Ecological Modelling, 181 (4):567–580, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003990.

Dame:2008:EEN

[DC08a] James K. Dame and Robert R. Christian. Evaluation of ecological network analysis: Validation of output. Ecological Modelling, 210(3):327–338, January 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004115.

Drury:2008:UMI

[DC08b] Kevin L. S. Drury and J. Fabian Candelaria. Using model identification to analyze spatially explicit data with habitat, and temporal, variability. Ecological Modelling, 214(2–4):305–315, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000835.

de-Camino-Beck:2009:ISS

[dCBL09] T. de Camino-Beck and M. A. Lewis. Invasion with stage-structured coupled map lattices: Application to the spread of scentless chamomile. Ecological Modelling, 220(23):3394–3403, December 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009006115.

Duffy:2001:SLU

[DCG01] Sheila Bliss Duffy, Michael S. Corson, and W. E. Grant. Simulating land-use decisions in the La Amistad Biosphere Reserve buffer zone in Costa Rica and Panama. Ecological Modelling, 140(1–2):9–29, May 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002666.

Drago:2001:TDN

[DCI01] Michele Drago, Bruno Cescon, and Luigi Iovenitti. A threedimensional numerical model for eutrophication and pollutant transport. Ecological Modelling, 145(1):17–34, November 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003842.

Deng:2007:EBS

[DCP^{+}07] Zhiqun Deng, Thomas J. Carlson, Gene R. Ploskey, Marshall C. Richmond, and Dennis D. Dauble. Evaluation of blade-strike models for estimating the biological performance of kaplan turbines. Ecological Modelling, 208(2–4):165–176, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003006.

deCastro:2001:SBF

[dCPC01] Luiz Arnaud Britto de Castro, Miguel Petrere, and Antonio Evaldo Comune. Sensitivity of the BEAM4 fisheries bioeconomic model to the main biological input parameters. Ecological Modelling, 141(1–3):53–66, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002411.

deCastro:2005:MMI

[dCS05] M. L. de Castro and J. A. L. Silva. Mathematical modelling of the Ibera Caiman yacaré. Ecological Modelling, 186(1):99–109, July 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000244.

Santino:2008:CCP

[dCSB08] Marcela Bianchessi da Cunha Santino and Irineu Bianchini. Carbon cycling potential from Utricularia breviscapa decomposition in a tropical oxbow lake (São Paulo, Brazil). Ecological Modelling, 218(3–4):375–382, November 10, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008003682.

Dilao:2000:GAM

[DD00] Rui Dilão and Tiago Domingos. A general approach to the modelling of trophic chains. Ecological Modelling, 132(3):191–202, August 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002726.

DeBlasio:2002:RDS

[DD02] Birgitte Freiesleben De Blasio and Fabio Vittorio De Blasio. Role of differential survival in a model of species–area curve. Ecological Modelling, 147(2):153–159, January 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004045.

Dzeroski:2003:URT

[DD03] Sašo Džeroski and Darrin Drumm. Using regression trees to identify the habitat preference of the sea cucumber (Holothuria leucospilota) on Rarotonga, Cook Islands. Ecological Modelling, 170(2–3):219–226, December 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002291.

Dufrene:2005:MCW

[DDF^{+}05] E. Dufrêne, H. Davi, C. François, G. le Maire, V. Le Dantec, and A. Granier. Modelling carbon and water cycles in a beech forest: Part i: Model description and uncertainty analysis on modelled NEE. Ecological Modelling, 185(2–4):407–436, July 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000098.

Doyen:2007:SEM

[DDFP07] L. Doyen, M. De Lara, J. Ferraris, and D. Pelletier. Sustainability of exploited marine ecosystems through protected areas: a viability model and a coral reef case study. Ecological Modelling, 208(2–4):353–366, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003146.

Davi:2005:MCW

[DDG^{+}05] H. Davi, E. Dufrêne, A. Granier, V. Le Dantec, C. Barbaroux, C. François, and N. Bréda. Modelling carbon and water cycles in a beech forest: Part II.: Validation of the main processes from organ to stand scale. Ecological Modelling, 185(2–4):387–405, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000086.

Dubey:2001:PPI

[DDH01] B. Dubey, B. Das, and J. Hussain. A predator–prey interaction model with self and cross-diffusion. Ecological Modelling, 141(1–3):67–76, July 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002551.

Dueri:2009:MCE

[DDH^{+}09] S. Dueri, I. Dahllöf, M. Hjorth, D. Marinov, and J. M. Zaldívar. Modeling the combined effect of nutrients and pyrene on the plankton population: Validation using mesocosm experiment data and scenario analysis. Ecological Modelling, 220(17):2060–2067, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003123.

Debeljak:2001:HSM

[DDJ^{+}01] Marko Debeljak, Sašo Džeroski, Klemen Jerina, Andrej Kobler, and Miha Adamič. Habitat suitability modelling for red deer (Cervus elaphus L.) in South-central Slovenia with classification trees. Ecological Modelling, 138(1–3):321–330, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000004117.

Demsar:2006:UMO

[DDL^{+}06] Damjan Demšar, Sašo Džeroski, Thomas Larsen, Jan Struyf, Jørgen Axelsen, Marianne Bruus Pedersen, and Paul Henning Krogh. Using multi-objective classification to model communities of soil microarthropods. Ecological Modelling, 191(1):131–143, January 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003819.

Dale:2002:AIS

[DDLB02] Michael B. Dale, Patricia E. R. Dale, Cen Li, and Gautam Biswas. Assessing impacts of small perturbations using a model-based approach. Ecological Modelling, 156(2–3):185–199, November 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001588.

Drury:2007:IED

[DDLD07] K. L. S. Drury, J. M. Drake, D. M. Lodge, and G. Dwyer. Immigration events dispersed in space and time: Factors affecting invasion success. Ecological Modelling, 206(1–2):63–78, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001585.

Durand:2004:MMS

[DDS^{+}04] B. Durand, M. A. Dubois, P. Sabatier, D. Calavas, C. Ducrot, and A. Van de Wielle. Multiscale modelling of scrapie epidemiology: II. Geographical level: hierarchical transfer of the herd model to the regional disease spread. Ecological Modelling, 179(4):515–531, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002777.

Dorneles:2001:MHE

[DdSG01] M. R. F. Dorneles, C. M. da Silva, and A. A. Gomes. A model for hyphae effects in phosphorus absorption by plants. Ecological Modelling, 142(1–2):83–89, August 1, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001002800.

Dorneles:2004:MHI

[DdSG04] M. R. F. Dorneles, C. M. da Silva, and A. A. Gomes. A model for hyphae-induced interaction between plants. Ecological Modelling, 174(3):323–330, May 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004332.

Dale:2007:SCU

[DDT07] M. B. Dale, P. E. R. Dale, and P. Tan. Supervised clustering using decision trees and decision graphs: an ecological comparison. Ecological Modelling, 204(1–2):70–78, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006006612.

Dzeroski:2006:FEC

[DDŽ06] Sašo Džeroski, Marko Debeljak, and Bernard Ženko. The Fourth European Conference on Ecological Modelling, September 27–October 1, 2004, Bled, Slovenia. Ecological Modelling, 194(1–3):1–2, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004928.

Debeljak:2002:AGS

[Deb02] Marko Debeljak. Applicability of genome size in exergy calculation. Ecological Modelling, 152(2–3):103–107, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002000108.

Deister:2004:UEF

[Dei04] Ane D. Deister. Using energy flows to track and influence organizational dynamics and decision making. Ecological Modelling, 178(1–2):163–166, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005246.

Delfino:2004:RHM

[Del04] Joseph J. Delfino. Revisiting heavy metals in the environment: using wetlands for their removal. Ecological Modelling, 178(1–2):35–38, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005258.

Deng:2008:TIP

[Den08] Bo Deng. The time invariance principle, the absence of ecological chaos, and a fundamental pitfall of discrete modeling. Ecological Modelling, 215(4):287–292, July 24, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008001543.

Deng:2009:CDM

[Den09] Bo Deng. Can discrete modellers work without the TIP? Ecological Modelling, 220(19):2600–2601, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004840.

Descy:2003:F

[Des03] J.-P. Descy. Foreword. Ecological Modelling, 160(3):189–190, February 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002002533.

Dew:2001:TMN

[Dew01] Ian M. Dew. Theoretical model of a new fishery under a simple quota management system. Ecological Modelling, 143(1–2):59–70, October 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003568.

Dube:2001:QGD

[DFCM01] P. Dubé, M. J. Fortin, C. D. Canham, and D. J. Marceau. Quantifying gap dynamics at the patch mosaic level using a spatially-explicit model of a northern hardwood forest ecosystem. Ecological Modelling, 142(1–2):39–60, August 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002381.

Didion:2007:FAS

[DFF07] M. Didion, M.-J. Fortin, and A. Fall. Forest age structure as indicator of boreal forest sustainability under alternative management and fire regimes: a landscape level sensitivity analysis. Ecological Modelling, 200(1–2):45–58, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003383.

DeLeo:2004:DGP

[DFGC04] Giulio A. De Leo, Stefano Focardi, Marino Gatto, and Isabella M. Cattadori. The decline of the grey partridge in Europe: comparing demographies in traditional and modern agricultural landscapes. Ecological Modelling, 177(3–4):313–335, October 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001437.

DAutilia:2004:STD

[DFHP04] Roberto D’Autilia, Margherita Falcucci, Vincent Hull, and Luisa Parrella. Short time dissolved oxygen dynamics in shallow water ecosystems. Ecological Modelling, 179(3):297–306, November 30, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001449.

Drielsma:2007:RBT

[DFM07] Michael Drielsma, Simon Ferrier, and Glenn Manion. A raster-based technique for analysing habitat configuration: the cost–benefit approach. Ecological Modelling, 202(3–4):324–332, April 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005291.

deFrutos:2007:AMS

[dFOV07] Ángel de Frutos, Pedro P. Olea, and Rubén Vera. Analyzing and modelling spatial distribution of summering lesser kestrel: the role of spatial autocorrelation. Ecological Modelling, 200(1–2):33–44, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003188.

dOultremont:2002:MMRa

[dG02a] Thibaud d’Oultremont and Andrew Paul Gutierrez. A multitrophic model of a rice–fish agroecosystem: II. Linking the flooded rice–fishpond systems. Ecological Modelling, 155(2–3):159–176, October 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001308.

dOultremont:2002:MMRb

[dG02b] Thibaud d’Oultremont and Andrew Paul Gutierrez. A multitrophic model of a rice-fish agroecosystem: I. A tropical fishpond food web. Ecological Modelling, 156 (2–3):123–142, November 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001291.

deBruyn:2004:MDF

[dG04a] Adrian M. H. deBruyn and Frank A. P. C. Gobas. Modelling the diagenetic fate of persistent organic pollutants in organically enriched sediments. Ecological Modelling, 179(3):405–416, November 30, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002704.

Duarte:2004:TRS

[DG04b] Luis O. Duarte and Camilo B. García. Trophic role of small pelagic fishes in a tropical upwelling ecosystem. Ecological Modelling, 172(2–4):323–338, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003764.

Dheygere:2003:UGA

[DGD03] Tom D’heygere, Peter L. M. Goethals, and Niels De Pauw. Use of genetic algorithms to select input variables in decision tree models for the prediction of benthic macroinvertebrates. Ecological Modelling, 160(3):291–300, February 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002002600.

Dheygere:2006:GAO

[DGD06] Tom D’heygere, Peter L. M. Goethals, and Niels De Pauw. Genetic algorithms for optimisation of predictive ecosystems models based on decision trees and neural networks. Ecological Modelling, 195(1–2):20–29, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005727.

deGuimaraens:2005:MUS

[dGdMPC05] Mariana Alves de Guimaraens, Afonso de Moraes Paiva, and Ricardo Coutinho. Modeling Ulva spp. dynamics in a tropical upwelling region. Ecological Modelling, 188(2–4):448–460, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002644.

Dedecker:2004:OAN

[DGGD04] Andy P. Dedecker, Peter L. M. Goethals, Wim Gabriels, and Niels De Pauw. Optimization of Artificial Neural Network (ANN) model design for prediction of macroinvertebrates in the Zwalm river basin (Flanders, Belgium). Ecological Modelling, 174(1–2):161–173, May 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000055.

Downs:2008:ECC

[DGM08] Joni A. Downs, Robert J. Gates, and Alan T. Murray. Estimating carrying capacity for sandhill cranes using habitat suitability and spatial optimization models. Ecological Modelling, 214(2–4):284–292, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800080X.

Dimov:2004:CCN

[DGOZ04] I. Dimov, K. Georgiev, Tz. Ostromsky, and Z. Zlatev. Computational challenges in the numerical treatment of large air pollution models. Ecological Modelling, 179(2):187–203, November 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002583.

Dobrowski:2006:IID

[DGRU06] S. Z. Dobrowski, J. A. Greenberg, C. M. Ramirez, and S. L. Ustin. Improving image derived vegetation maps with regression based distribution modeling. Ecological Modelling, 192(1–2):126–142, February 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004503.

Deygout:2009:MIF

[DGSBG09] C. Deygout, A. Gault, F. Sarrazin, and C. Bessa-Gomes. Modeling the impact of feeding stations on vulture scavenging service efficiency. Ecological Modelling, 220(15):1826–1835, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009002889.

Dubey:2000:MAE

[DH00] B. Dubey and J. Hussain. A model for the allelopathic effect on two competing species. Ecological Modelling, 129(2–3):195–207, May 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002283.

Dumont:2001:MAS

[DH01] Bertrand Dumont and David R. C. Hill. Multi-agent simulation of group foraging in sheep: effects of spatial memory, conspecific attraction and plot size. Ecological Modelling, 141(1–3):201–215, July 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002745.

DeAngelis:2006:ERD

[DH06] Donald L. DeAngelis and J. Nathaniel Holland. Emergence of ratio-dependent and predator-dependent functional responses for pollination mutualism and seed parasitism. Ecological Modelling, 191(3–4):551–556, February 5, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005003030.

Dunstan:2005:PGD

[DJ05] Piers K. Dunstan and Craig R. Johnson. Predicting global dynamics from local interactions: individual-based models predict complex features of marine epibenthic communities.

Ecological Modelling, 186(2):221–233, August 15, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005000281.

deJong:2002:MES

[dJBPG02] Meindert D. de Jong, Graeme W. Bourdôt, James Powell, and Jan Goudriaan. A model of the escape of Sclerotinia sclerotiorum ascospores from pasture. Ecological Modelling, 150(1–2):83–105, April 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004628.

Torres-Meza:2009:GBM

[dJTMBGMP^{+}09] María de Jesús Torres-Meza, Alma Delia Báez-González, Luis Humberto Maciel-Pérez, Esperanza Quezada-Guzmán, and J. Santos Sierra-Tristán. GIS-based modeling of the geographic distribution of Quercus emoryi Torr. (Fagaceae) in México and identification of significant environmental factors influencing the species’ distribution. Ecological Modelling, 220(24):3599–3611, December 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000702.

Didion:2009:PLU

[DKL^{+}09] M. Didion, A. D. Kupferschmid, M. J. Lexer, W. Rammer, R. Seidl, and H. Bugmann. Potentials and limitations of using large-scale forest inventory data for evaluating forest succession models. Ecological Modelling, 220(2):133–147, January 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008004626.

Dalaka:2000:MEE

[DKRŠS00] A. Dalaka, B. Kompare, M. Robnik-Šikonja, and S. P. Sgardelis. Modelling the effects of environmental conditions on apparent photosynthesis of Stipa bromoides by machine learning tools. Ecological Modelling, 129(2–3):245–257, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002374.

Dreyfus-Leon:2007:RPG

[DLC07] Michel Dreyfus-León and D. G. Chen. Recruitment prediction with genetic algorithms with application to the Pacific herring fishery. Ecological Modelling, 203(1–2):141–146, April 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006005655.

delaFuente:2008:PEM

[dlFN08] Alberto de la Fuente and Yarko Niño. Pseudo 2D ecosystem model for a dendritic reservoir. Ecological Modelling, 213(3–4):389–401, May 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000331.

Dreyfus-Leon:2006:MPI

[DLG06] Michel Dreyfus-Leon and Daniel Gaertner. Modeling performance and information exchange between fishing vessels with artificial neural networks. Ecological Modelling, 195(1–2):30–36, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005005739.

Dreyfus-Leon:2001:SIB

[DLK01] Michel Dreyfus-Leon and Pierre Kleiber. A spatial individual behaviour-based model approach of the yellowfin tuna fishery in the eastern Pacific Ocean. Ecological Modelling, 146(1–3):47–56, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002952.

Duan:2009:MCC

[DLL^{+}09] L. J. Duan, S. Y. Li, Y. Liu, J. Moreau, and V. Christensen. Modeling changes in the coastal ecosystem of the Pearl River Estuary from 1981 to 1998. Ecological Modelling, 220(20):2802–2818, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005006.

Degenne:2009:DDS

[DLP^{+}09a] P. Degenne, D. Lo Seen, D. Parigot, R. Forax, A. Tran, A. Ait Lahcen, O. Curé, and R. Jeansoulin. Design of a domain specific language for modelling processes in landscapes. Ecological Modelling, 220(24):3527–3535, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000900413X.

Drechsler:2009:CPS

[DLP09b] Martin Drechsler, Reinaldo Lourival, and Hugh P. Possingham. Conservation planning for successional landscapes. Ecological Modelling, 220(4):438–450, February 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005425.

Dray:2006:SMC

[DLPN06] Stéphane Dray, Pierre Legendre, and Pedro R. Peres-Neto. Spatial modelling: a comprehensive framework for principal coordinate analysis of neighbour matrices (PCNM). Ecological Modelling, 196(3–4):483–493, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000925.

delaParra:2005:TEM

[dlPP05] Rafael Bravo de la Parra and Jean-Christophe Poggiale. Theoretical ecology and mathematical modelling: Problems and methods. Ecological Modelling, 188(1):1–2, October 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002322.

Dambacher:2003:QPM

[DLR03] Jeffrey M. Dambacher, Hiram W. Li, and Philippe A. Rossignol. Qualitative predictions in model ecosystems. Ecological Modelling, 161(1–2):79–93, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002958.

delaSen:2007:ECC

[dlS07] M. de la Sen. The environment carrying capacity is not independent of the intrinsic growth rate for subcritical spawning stock biomass in the Beverton–Holt equation. Ecological Modelling, 204(1–2):271–273, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006041.

Dowd:2003:BAE

[DM03] Michael Dowd and Renate Meyer. A Bayesian approach to the ecosystem inverse problem. Ecological Modelling, 168(1–2):39–55, October 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001868.

Dube:2005:SIS

[DMBM05] P. Dubé, A. Ménard, A. Bouchard, and D. J. Marceau. Simulating the impact of small-scale extrinsic disturbances over forest species volumetric light environment. Ecological Modelling, 182(2):113–129, March 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002935.

Dagorn:2000:CEM

[DMBO00] Laurent Dagorn, Filippo Menczer, Pascal Bach, and Robert J. Olson. Co-evolution of movement behaviours by tropical pelagic predatory fishes in response to prey environment: a simulation model. Ecological Modelling, 134(2–3):325–341, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003744.

Drielsma:2007:SLT

[DMF07] Michael Drielsma, Glenn Manion, and Simon Ferrier. The spatial links tool: Automated mapping of habitat linkages in variegated landscapes. Ecological Modelling, 200(3–4):403–411, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600384X.

Duarte:2003:MMA

[DMH^{+}03] P. Duarte, R. Meneses, A. J. S. Hawkins, M. Zhu, J. Fang, and J. Grant. Mathematical modelling to assess the carrying capacity for multi-species culture within coastal waters. Ecological Modelling, 168(1–2):109–143, October 1, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002059.

deMattos:2007:MQB

[dMPVNO07] Sérgio Henrique Vannucchi Leme de Mattos, José Roberto Castilho Piqueira, João Vasconcelos-Neto, and Fernando Moya Orsatti. Measuring q-bits in three-trophic level systems. Ecological Modelling, 200(1–2):183–188, January 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006003164.

Duursma:2007:DTS

[DMRP07] R. A. Duursma, J. D. Marshall, A. P. Robinson, and R. E. Pangle. Description and test of a simple process-based model of forest growth for mixed-species stands. Ecological Modelling, 203(3–4):297–311, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006090.

Dolgonosov:2006:IFH

[DN06] B. M. Dolgonosov and V. I. Naidenov. An informational framework for human population dynamics. Ecological Modelling, 198(3–4):375–386, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002390.

Donovan:2006:DOM

[Don06] Geoffrey H. Donovan. Determining the optimal mix of federal and contract fire crews: a case study from the Pacific Northwest. Ecological Modelling, 194(4):372–378, April 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005417.

Dormann:2007:AVA

[Dor07] Carsten F. Dormann. Assessing the validity of autologistic regression. Ecological Modelling, 207(2–4):234–242, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007002748.

Dowd:2005:BPC

[Dow05] Michael Dowd. A bio-physical coastal ecosystem model for assessing environmental effects of marine bivalve aquaculture. Ecological Modelling, 183(2–3):323–346, April 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004004867.

Drouet:2003:GMG

[DP03] J.-L. Drouet and L. Pagès. GRAAL: a model of GRowth, architecture and carbon ALlocation during the vegetative phase of the whole maize plant: Model description and parameterisation. Ecological Modelling, 165(2–3):147–173, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003000723.

Drouet:2007:GCM

[DP07] Jean-Louis Drouet and Loïc Pagès. GRAAL-CN: a model of GRowth, architecture and ALlocation for carbon and nitrogen dynamics within whole plants formalised at the organ level. Ecological Modelling, 206(3–4):231–249, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001810.

DOdorico:2004:PMN

[DPL^{+}04] Paolo D’Odorico, Amilcare Porporato, Francesco Laio, Luca Ridolfi, and Ignacio Rodriguez-Iturbe. Probabilistic modeling of nitrogen and carbon dynamics in water-limited ecosystems. Ecological Modelling, 179(2):205–219, November 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004002595.

DelGiudice:2009:TIP

[DPT09] Emilio Del Giudice, Riccardo M. Pulselli, and Enzo Tiezzi. Thermodynamics of irreversible processes and quantum field theory: an interplay for the understanding of ecosystem dynamics. Ecological Modelling, 220(16):1874–1879, August 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002762.

Das:2008:EDN

[DR08] Kalyan Das and Santanu Ray. Effect of delay on nutrient cycling in phytoplankton–zooplankton interactions in estuarine system. Ecological Modelling, 215(1–3):69–76, July 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001026.

Diffendorfer:2001:ALP

[DRDD01] James E. Diffendorfer, Paul M. Richards, George H. Dalrymple, and Donald L. DeAngelis. Applying linear programming to estimate fluxes in ecosystems or food webs: an example from the herpetological assemblage of the freshwater Everglades. Ecological Modelling, 144(2–3):99–120, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003428.

Drechsler:2001:PPI

[DS01] Martin Drechsler and Josef Settele. Predator–prey interactions in rice ecosystems: effects of guild composition, trophic relationships, and land use changes — a model study exemplified for Philippine rice terraces. Ecological Modelling, 137(2–3):135–159, February 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004233.

Deng:2002:OCP

[DS02] Zhi-Qiang Deng and Vijay P. Singh. Optimum channel pattern for environmentally sound training and management of alluvial rivers. Ecological Modelling, 154(1–2):61–74, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002000534.

daSilva:2001:DSM

[dSA01] Eduardo Teixeira da Silva and Milton L. Asmus. A dynamic simulation model of the widgeon grass Ruppia maritima and its epiphytes in the estuary of the Patos Lagoon, RS, Brazil. Ecological Modelling, 137(2–3):161–179, February 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000004257.

DaConceicaoPrates-Clark:2008:PGD

[DSA08] Cássia Da Conceição Prates-Clark, Sassan S. Saatchi, and Donat Agosti. Predicting geographical distribution models of high-value timber trees in the Amazon Basin using remotely sensed data. Ecological Modelling, 211 (3–4):309–323, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004759.

Debeljak:2008:RBO

[DSD^{+}08] Marko Debeljak, Geoff R. Squire, Damjan Demšar, Mark W. Young, and Sašo Džeroski. Relations between the oilseed rape volunteer seedbank, and soil factors, weed functional groups and geographical location in the UK. Ecological Modelling, 212(1–2):138–146, March 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005145.

Dur:2009:IBM

[DSD^{+}09] Gael Dur, Sami Souissi, David Devreker, Vincent Ginot, François G. Schmitt, and Jiang-Shiou Hwang. An individual-based model to study the reproduction of egg bearing copepods: Application to Eurytemora affinis (Copepoda Calanoida) from the Seine estuary, France. Ecological Modelling, 220(8):1073–1089, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005917.

Dias:2008:UCB

[dSDSM08] Carlos Tadeu dos Santos Dias, Ari Samaranayaka, and Bryan Manly. On the use of correlated beta random variables with animal population modelling. Ecological Modelling, 215(4):293–300, July 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001646.

Sternberg:2002:CIC

[dSLSD02] Leonel da Silveira Lobo Sternberg and Donald L. DeAngelis. Carbon isotope composition of ambient CO_{2} and recycling: a matrix simulation model. Ecological Modelling, 154(1–2):179–192, August 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000558.

deSouza:2009:CSA

[dSMZ09] A. A. de Souza, S. G. F. Martins, and M. S. Zacarias. Computer simulation applied to the biological control of the insect Aphis gossypii for the parasitoid Lysiphlebus testaceipes. Ecological Modelling, 220(6):756–763, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005796.

Peixoto:2008:MCA

[dSPdBB08a] Magda da Silva Peixoto, Laécio Carvalho de Barros, and Rodney Carlos Bassanezi. A model of cellular automata for the spatial and temporal analysis of citrus sudden death with the fuzzy parameter. Ecological Modelling, 214(1):45– 52, June 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008000240.

Peixoto:2008:PPF

[dSPdBB08b] Magda da Silva Peixoto, Laécio Carvalho de Barros, and Rodney Carlos Bassanezi. Predator–prey fuzzy model. Ecological Modelling, 214(1):39–44, June 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000239.

deSevereyn:2000:EWT

[dSSGR00] Y. García de Severeyn, H. Severeyn, W. Grant, and Y. Reverol. Effect of water temperature on larval development of the bivalve mollusk Tivela mactroides: evaluation in the laboratory and via simulation. Ecological Modelling, 129(2–3):143–151, May 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002313.

Dzeroski:2003:LPD

[DT03] Sašo Džeroski and Ljupčo Todorovski. Learning population dynamics models from data and domain knowledge. Ecological Modelling, 170(2–3):129–140, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002217.

Diaz-Uribe:2007:MPS

[DUASCM07] J. Gabriel Díaz-Uribe, Francisco Arreguín-Sánchez, and Miguel A. Cisneros-Mata. Multispecies perspective for small-scale fisheries management: a trophic analysis of La Paz Bay in the Gulf of California, Mexico. Ecological Modelling, 201(2):205–222, February 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600442X.

Dubey:2003:EIP

[DUH03] B. Dubey, R. K. Upadhyay, and J. Hussain. Effects of industrialization and pollution on resource biomass: a mathematical model. Ecological Modelling, 167(1–2):83–95, September 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001686.

Dunne:2008:URS

[Dun08] R. P. Dunne. The use of remotely sensed solar radiation data in modelling susceptibility of coral reefs to environmental stress: Comment on Maina et al. [Ecol. Model. 212 (2008) 180–199]. Ecological Modelling, 218 (1–2):188–191, October 24, 2008. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0304380008001841. See [MVMA08, GB08].

Deckmyn:2008:ASS

[DVdB^{+}08] G. Deckmyn, H. Verbeeck, M. Op de Beeck, D. Vansteenkiste, K. Steppe, and R. Ceulemans. ANAFORE: a stand-scale process-based forest model that includes wood tissue development and labile carbon storage in trees. Ecological Modelling, 215(4):345–368, July 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001816.

Dedecker:2007:DMM

[DVGD07] Andy P. Dedecker, Koen Van Melckebeke, Peter L. M. Goethals, and Niels De Pauw. Development of migration models for macroinvertebrates in the Zwalm river basin (Flanders, Belgium) as tools for restoration management. Ecological Modelling, 203(1–2):72–86, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600559X.

Deutsch:2008:MCA

[DVJ^{+}08] Felix Deutsch, Jean Vankerkom, Liliane Janssen, Stijn Janssen, László Bencs, René Van Grieken, Frans Fierens, Gerwin Dumont, and Clemens Mensink. Modelling concentrations of airborne primary and secondary PM_{10} and PM_{2.5} with the BelEUROS-model in Belgium. Ecological Modelling, 217(3–4):230–239, October 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002561.

DeMichele:2008:MMS

[DVPS08] C. De Michele, R. Vezzoli, H. Pavlopoulos, and R. J. Scholes. A minimal model of soil water–vegetation interactions forced by stochastic rainfall in water-limited ecosystems. Ecological Modelling, 212(3–4):397–407, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005613.

Drake:2001:SVH

[DW01] J. B. Drake and J. F. Weishampel. Simulating vertical and horizontal multifractal patterns of a longleaf pine savanna. Ecological Modelling, 145(2–3):129–142, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003982.

Dornbusch:2007:MEM

[DWD07] Tino Dornbusch, Peter Wernecke, and Wulf Diepenbrock. A method to extract morphological traits of plant organs from 3D point clouds as a database for an architectural plant model. Ecological Modelling, 200(1–2):119–129, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003486.

Dahl:2006:CSP

[DWH06] Magnus Dahl, David I. Wilson, and Lars Håkanson. A combined suspended particle and phosphorus water quality model: Application to Lake Vänern. Ecological Modelling, 190(1–2):55–71, January 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002267.

Dittrich:2009:LSD

[DWR09] M. Dittrich, B. Wehrli, and P. Reichert. Lake sediments during the transient eutrophication period: Reactivetransport model and identifiability study. Ecological Modelling, 220(20):2751–2769, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005018.

Dzeroski:2006:FIW

[DŽD06] Sašo Džeroski, Bernard Ženko, and Marko Debeljak. The Fourth International Workshop on Environmental Applications of Machine Learning, 27 September–1 October 2004, Bled, Slovenia. Ecological Modelling, 191(1):1–3, January 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003443.

Dzeroski:2001:ASM

[Dže01] Sašo Džeroski. Applications of symbolic machine learning to ecological modelling. Ecological Modelling, 146 (1–3):263–273, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000100312X.

Etterson:2006:EUA

[EB06] Matthew A. Etterson and Richard S. Bennett. The effects of uncertainty about age at transition on bias in the Mayfield family of estimators. Ecological Modelling, 199(3):253–260, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002742.

Edgar:2007:DVM

[EB07a] Christopher B. Edgar and Thomas E. Burk. Demonstration and verification of a model that generates defoliation patterns in forested landscapes. Ecological Modelling, 205(3–4):301–313, July 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001007.

ElSaadi:2007:IBM

[EB07b] N. El Saadi and A. Bah. An individual-based model for studying the aggregation behavior in phytoplankton. Ecological Modelling, 204(1–2):193–212, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000099.

Easterling:2001:SIH

[EBH^{+}01] William E. Easterling, James R. Brandle, Cynthia J. Hays, Qinfeng Guo, and David S. Guertin. Simulating the impact of human land use change on forest composition in the Great Plains agroecosystems with the Seedscape model. Ecological Modelling, 140(1–2):163–176, May 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002630.

Eppstein:2006:STC

[EBM06] Margaret J. Eppstein, James D. Bever, and Jane Molofsky. Spatio-temporal community dynamics induced by frequency dependent interactions. Ecological Modelling, 197(1–2):133–147, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001050.

Elith:2002:MEU

[EBR02] Jane Elith, Mark A. Burgman, and Helen M. Regan. Mapping epistemic uncertainties and vague concepts in predictions of species distribution. Ecological Modelling, 157(2–3):313–329, November 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002028.

Eisenhauer:2009:ZPM

[ECBD09] L. Eisenhauer, F. Carlotti, M. Baklouti, and F. Diaz. Zooplankton population model coupled to a biogeochemical model of the North Western Mediterranean Sea ecosystem. Ecological Modelling, 220(21):2865–2876, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004141.

Echeverria:2008:SEM

[ECHN08] Cristian Echeverria, David A. Coomes, Myrna Hall, and Adrian C. Newton. Spatially explicit models to analyze forest loss and fragmentation between 1976 and 2020 in southern Chile. Ecological Modelling, 212(3–4):439–449, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005650.

Euskirchen:2002:MLN

[ECL^{+}02] Eugénie S. Euskirchen, Jiquan Chen, Harbin Li, Eric J. Gustafson, and Thomas R. Crow. Modeling landscape net ecosystem productivity (LandNEP) under alternative management regimes. Ecological Modelling, 154(1–2):75–91, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002000522.

ElMaayar:2008:UFM

[ECP08] Mustapha El Maayar, Jing M. Chen, and David T. Price. On the use of field measurements of energy fluxes to evaluate land surface models. Ecological Modelling, 214(2–4):293–304, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000823.

Edwards:2006:ESS

[ECZ^{+}06] Thomas C. Edwards, D. Richard Cutler, Niklaus E. Zimmermann, Linda Geiser, and Gretchen G. Moisen. Effects of sample survey design on the accuracy of classification tree models in species distribution models. Ecological Modelling, 199(2):132–141, November 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002456.

Emlen:2006:IAM

[EDKF06] John M. Emlen, Jeffrey J. Duda, Matt D. Kirchhoff, and D. Carl Freeman. Interaction assessment: a modeling tool for predicting population dynamics from field data. Ecological Modelling, 192(3–4):557–570, February 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005003595.

Elith:2005:ESN

[EFHL05] Jane Elith, Simon Ferrier, Falk Huettmann, and John Leathwick. The evaluation strip: a new and robust method for plotting predicted responses from species distribution models. Ecological Modelling, 186(3):280–289, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004006180.

Ebel:2008:STC

[EFJ^{+}08] A. Ebel, H. Feldmann, H. J. Jakobs, M. Memmesheimer, D. Offermann, V. Kuell, and B. Schäler. Simulation of transport and composition changes during a blocking episode over the East Atlantic and North Europe. Ecological Modelling, 217(3–4):240–254, October 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002627.

Emlen:2003:FPM

[EFK^{+}03] John M. Emlen, D. C. Freeman, M. D. Kirchhoff, C. L. Alados, J. Escos, and J. J. Duda. Fitting population models from field data. Ecological Modelling, 162(1–2):119–143, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002003903.

Elkalay:2003:MSD

[EFS^{+}03] Khalid Elkalay, Constantin Frangoulis, Nikos Skliris, Anne Goffart, Sylvie Gobert, Gilles Lepoint, and Jean-Henri Hecq. A model of the seasonal dynamics of biomass and production of the seagrass Posidonia oceanica in the Bay of Calvi (Northwestern Mediterranean). Ecological Modelling, 167(1–2):1–18, September 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000747.

Eguchi:2009:BAL

[EG09] Tomoharu Eguchi and Tim Gerrodette. A Bayesian approach to line-transect analysis for estimating abundance. Ecological Modelling, 220(13–14):1620–1630, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002300.

Esther:2008:AIS

[EGE^{+}08] Alexandra Esther, Jürgen Groeneveld, Neal J. Enright, Ben P. Miller, Byron B. Lamont, George L. W. Perry, Frank M. Schurr, and Florian Jeltsch. Assessing the importance of seed immigration on coexistence of plant functional types in a species-rich ecosystem. Ecological Modelling, 213(3–4):402–416, May 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000355.

Elliott:2000:MFP

[EIRT00] J. A. Elliott, A. E. Irish, C. S. Reynolds, and P. Tett. Modelling freshwater phytoplankton communities: an exercise in validation. Ecological Modelling, 128(1):19–26, April 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002215.

El-Jaick:2005:EDW

[EJG05] Léa J. El-Jaick and A. A. Gomes. Estimate of the drainage water behaviour in shallow lakes. Ecological Modelling, 184(2–4):219–227, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005484.

Ebenhoh:2004:SMM

[EKBF04] W. Ebenhöh, C. Kohlmeier, J. W. Baretta, and G. Flöser. Shallowness may be a major factor generating nutrient gradients in the Wadden Sea. Ecological Modelling, 174(3):241–252, May 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004277.

Eliseev:2008:ESC

[EM08] Alexey V. Eliseev and Igor I. Mokhov. Eventual saturation of the climate–carbon cycle feedback studied with a conceptual model. Ecological Modelling, 213(1):127–132, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007006230.

Evans:2001:DMH

[EMdC^{+}01] Tom P. Evans, Aaron Manire, Fabio de Castro, Eduardo Brondizio, and Stephen McCracken. A dynamic model of household decision-making and parcel level landcover change in the eastern Amazon. Ecological Modelling, 143(1–2):95–113, October 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100357X.

Ezoe:2006:SDS

[EN06] Hideo Ezoe and Sayaka Nakamura. Size distribution and spatial autocorrelation of subpopulations in a size structured metapopulation model. Ecological Modelling, 198 (3–4):293–300, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006002122.

Etterson:2008:MSE

[EN08] Matthew A. Etterson and Laura R. Nagy. Is mean squared error a consistent indicator of accuracy for spatially structured demographic models? Ecological Modelling, 211(1–2):202–208, February 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004528.

Even:2004:MOD

[EPM^{+}04] Stéphanie Even, Michel Poulin, Jean-Marie Mouchel, Martin Seidl, and Pierre Servais. Modelling oxygen deficits in the Seine River downstream of combined sewer overflows. Ecological Modelling, 173(2–3):177–196, April 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003934.

Epperson:2000:SST

[Epp00] Bryan K. Epperson. Spatial and space–time correlations in ecological models. Ecological Modelling, 132(1–2):63–76, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003057.

Estrada-Pena:2006:FMA

[EPQdCA06] Agustín Estrada-Peña, Joaquín Quílez, Emilio del Cacho, and Caridad Sánchez Acedo. A framework to map abundance of tick metapopulations. Ecological Modelling, 193(3–4):663–674, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004850.

Edwards:2004:HNL

[EPS04] Andrew M. Edwards, Trevor Platt, and Shubha Sathyendranath. The high-nutrient, low-chlorophyll regime of the ocean: limits on biomass and nitrate before and after iron enrichment. Ecological Modelling, 171(1–2):103–125, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003003028.

Elliott:2007:PML

[EPTB07] J. A. Elliott, I. Persson, S. J. Thackeray, and T. Blenckner. Phytoplankton modelling of Lake Erken, Sweden by linking the models PROBE and PROTECH. Ecological Modelling, 202(3–4):421–426, April 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005412.

Enea:2001:FAE

[ES01] Mario Enea and Giuseppe Salemi. Fuzzy approach to the environmental impact evaluation. Ecological Modelling, 136(2–3):131–147, January 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000380X.

Eiswerth:2006:EPW

[ES06] Mark E. Eiswerth and J. Scott Shonkwiler. Examining post-wildfire reseeding on arid rangeland: a multivariate tobit modelling approach. Ecological Modelling, 192 (1–2):286–298, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003364.

Eschenbach:2005:EPM

[Esc05] Christiane Eschenbach. Emergent properties modelled with the functional structural tree growth model ALMIS: Computer experiments on resource gain and use. Ecological Modelling, 186(4):470–488, September 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001377.

Erunova:2006:GAS

[ESG06] Marina G. Erunova, Michael G. Sadovsky, and Anna A. Gosteva. GIS-aided simulation of spatially distributed environmental processes at “Stolby” state reservation. Ecological Modelling, 195(3–4):296–306, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500623X.

Espana:2003:EIS

[Esp03] Horacio Pérez España. Ecological importance of snappers in the stability of modeled coastal ecosystems. Ecological Modelling, 168(1–2):13–24, October 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002011.

Engel:2002:FCH

[ESWG02] V. C. Engel, M. Stieglitz, M. Williams, and K. L. Griffin. Forest canopy hydraulic properties and catchment water balance: observations and modeling. Ecological Modelling, 154(3):263–288, September 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000686.

Engelbrecht:2000:MPA

[ESZ^{+}00] Johann P. Engelbrecht, Leon Swanepoel, Mark Zunckel, Judith C. Chow, John G. Watson, and Richard T. Egami. Modelling PM _{10} aerosol data from the qalabotjha lowsmoke fuels macro-scale experiment in South Africa. Ecological Modelling, 127(2–3):235–244, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002124.

Elliott:2004:SPS

[ET04] J. A. Elliott and S. J. Thackeray. The simulation of phytoplankton in shallow and deep lakes using PROTECH. Ecological Modelling, 178(3–4):357–369, November 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004001541.

Eitzinger:2004:CPC

[ETH^{+}04] J. Eitzinger, M. Trnka, J. Hösch, Z. Žalud, and M. Dubrovský. Comparison of CERES, WOFOST and SWAP models in simulating soil water content during growing season under different soil conditions. Ecological Modelling, 171(3):223–246, January 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003003351.

Etienne:2004:OCI

[Eti04] Rampal S. Etienne. On optimal choices in increase of patch area and reduction of interpatch distance for metapopulation persistence. Ecological Modelling, 179(1):77–90, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004002637.

Evers:2007:AMS

[EVF^{+}07] Jochem B. Evers, Jan Vos, Christian Fournier, Bruno Andrieu, Michaël Chelle, and Paul C. Struik. An architectural model of spring wheat: Evaluation of the effects of population density and shading on model parameterization and performance. Ecological Modelling, 200(3–4):308–320, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006003681.

Etienne:2002:IBD

[EWH^{+}02] Rampal Etienne, Bregje Wertheim, Lia Hemerik, Petra Schneider, and James Powell. The interaction between dispersal, the Allee effect and scramble competition affects population dynamics. Ecological Modelling, 148(2):153–168, February 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004173.

Ewing:2002:EDC

[EYB^{+}02] Bland Ewing, Brian S. Yandell, James F. Barbieri, Robert F. Luck, and Lisa D. Forster. Event-driven competing risks. Ecological Modelling, 158(1–2):35–50, December 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002181.

Ezanno:2005:DTC

[Eza05] P. Ezanno. Dynamics of a tropical cattle herd in a variable environment: a modelling approach in order to identify the target period and animals on which concentrating management efforts to improve productivity. Ecological Modelling, 188(2–4):470–482, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001651.

Fleming:2003:SSV

[FA03] Christopher M. Fleming and Robert R. Alexander. Singlespecies versus multiple-species models: the economic implications. Ecological Modelling, 170(2–3):203–211, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002278.

Fernandez:2002:CEE

[FAA^{+}02] Clemente Fernández, Francisco J. Acosta, Gerardo Abellá, Francisco López, and Mario Díaz. Complex edge effect fields as additive processes in patches of ecological systems. Ecological Modelling, 149(3):273–283, April 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004641.

Ford:2007:SEO

[FAB^{+}07] R. Glenn Ford, David G. Ainley, Evelyn D. Brown, Robert M. Suryan, and David B. Irons. A spatially explicit optimal foraging model of black-legged kittiwake behavior based on prey density, travel distances, and colony size. Ecological Modelling, 204(3–4):335–348, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000257.

Fath:2004:CDS

[Fat04a] Brian D. Fath. Control of distributed systems and environmental applications. Ecological Modelling, 179(2):151– 152, November 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380004002522.

Fath:2004:DCE

[Fat04b] Brian D. Fath. Distributed control in ecological networks. Ecological Modelling, 179(2):235–245, November 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004002613.

Fath:2004:NAA

[Fat04c] Brian D. Fath. Network analysis applied to large-scale cyberecosystems. Ecological Modelling, 171(4):329–337, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002758.

Fath:2007:NMP

[Fat07a] Brian D. Fath. Network mutualism: Positive communitylevel relations in ecosystems. Ecological Modelling, 208(1): 56–67, October 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007002529.

Fath:2007:SFW

[Fat07b] Brian D. Fath. Structural food web regimes. Ecological Modelling, 208(2–4):391–394, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003122.

Fath:2009:ELN

[Fat09a] Brian D. Fath. Editorial: Letter from the new Editorin-Chief. Ecological Modelling, 220(1):1–3, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008004614.

Fath:2009:EEL

[Fat09b] Brian D. Fath. Erratum to “Editorial: Letter from the New Editor-in-Chief” [Ecol. Model. 220 (2009) 1–3]. Ecological Modelling, 220(11):1459, June 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001562.

Fleming:2002:SAT

[FBC02] Richard A. Fleming, Hugh J. Barclay, and Jean-Noël Candau. Scaling-up an autoregressive time-series model (of spruce budworm population dynamics) changes its qualitative behavior. Ecological Modelling, 149(1–2):127– 142, March 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380001005191.

Foltete:2008:CDD

[FBC08] J. C. Foltête, K. Berthier, and J. F. Cosson. Cost distance defined by a topological function of landscape. Ecological Modelling, 210(1–2):104–114, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700364X.

Fortin:2009:ATP

[FBDM09] Mathieu Fortin, Steve Bédard, Josianne DeBlois, and Sébastien Meunier. Assessing and testing prediction uncertainty for single tree-based models: a case study applied to northern hardwood stands in southern Québec, Canada. Ecological Modelling, 220(20):2770–2781, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004451.

Ferguson:2006:SMD

[FBF^{+}06] Megan C. Ferguson, Jay Barlow, Paul Fiedler, Stephen B. Reilly, and Tim Gerrodette. Spatial models of delphinid (family Delphinidae) encounter rate and group size in the eastern tropical Pacific Ocean. Ecological Modelling, 193(3–4):645–662, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004898.

Fath:2004:EFI

[FC04] Brian D. Fath and Heriberto Cabezas. Exergy and Fisher information as ecological indices. Ecological Modelling, 174(1–2):25–35, May 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003005660.

Federico:2005:MPD

[FC05a] Paula Federico and Graciela Ana Canziani. Modeling the population dynamics of capybara Hydrochaeris hydrochaeris: a first step towards a management plan. Ecological Modelling, 186(1):111–121, July 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000190.

Ferrati:2005:AWL

[FC05b] Rosana Ferrati and Graciela A. Canziani. An analysis of water level dynamics in Esteros del Ibera wetland. Ecological Modelling, 186(1):17–27, July 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000323.

Foody:2006:MSR

[FC06] Giles M. Foody and Mark E. J. Cutler. Mapping the species richness and composition of tropical forests from remotely sensed data with neural networks. Ecological Modelling, 195(1–2):37–42, May 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005740.

Fu:2000:SCD

[FCC^{+}00] Shenglei Fu, Miguel L. Cabrera, David C. Coleman, Keith W. Kisselle, Carol J. Garrett, Paul F. Hendrix, and D. A. Crossley. Soil carbon dynamics of conventional tillage and no-till agroecosystems at Georgia Piedmont — HSB-C models. Ecological Modelling, 131(2–3):229–248, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002507.

Fule:2004:CCF

[FCC^{+}04] Peter Z. Fulé, Joseph E. Crouse, Allison E. Cocke, Margaret M. Moore, and W. Wallace Covington. Changes in canopy fuels and potential fire behavior 1880–2040: Grand Canyon, Arizona. Ecological Modelling, 175(3):231–248, July 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003004782.

Favier:2004:MFS

[FCF^{+}04] Charly Favier, Jérôme Chave, Aline Fabing, Dominique Schwartz, and Marc A. Dubois. Modelling forest–savanna mosaic dynamics in man-influenced environments: effects of fire, climate and soil heterogeneity. Ecological Modelling, 171(1–2):85–102, January 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003016.

Franke:2004:AMB

[FCH04] Alastair Franke, Terry Caelli, and Robert J. Hudson. Analysis of movements and behavior of caribou (Rangifer tarandus) using hidden Markov models. Ecological Modelling, 173(2–3):259–270, April 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003983.

Franke:2006:PWC

[FCKH06] Alastair Franke, Terry Caelli, Gerald Kuzyk, and Robert J. Hudson. Prediction of wolf (Canis lupus) kill-sites using hidden Markov models. Ecological Modelling, 197 (1–2):237–246, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001104.

Ferrati:2005:EIH

[FCM05] Rosana Ferrati, Graciela Ana Canziani, and Diego Ruiz Moreno. Esteros del Ibera: hydrometeorological and hydrological characterization. Ecological Modelling, 186(1):3–15, July 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005000347.

Fergola:2007:ACB

[FCP^{+}07] P. Fergola, M. Cerasuolo, A. Pollio, G. Pinto, and M. DellaGreca. Allelopathy and competition between Chlorella vulgaris and Pseudokirchneriella subcapitata: Experiments and mathematical model. Ecological Modelling, 208(2– 4):205–214, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007003031.

Ferreol:2008:ETF

[FDCH08] M. Ferréol, A. Dohet, H.-M. Cauchie, and L. Hoffmann. An environmental typology of freshwater sites in Luxembourg as a tool for predicting macroinvertebrate fauna under nonpolluted conditions. Ecological Modelling, 212(1–2):99–108, March 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005200.

Ferreyra:2003:NEW

[FDP^{+}03] R. A. Ferreyra, J. L. Dardanelli, L. B. Pachepsky, D. J. Collino, P. C. Faustinelli, G. Giambastiani, V. R. Reddy, and J. W. Jones. Nonlinear effects of water stress on peanut photosynthesis at crop and leaf scales. Ecological Modelling, 168(1–2):57–76, October 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001959.

Fankhauser:2004:SAC

[FE04] Regine Fankhauser and Peter Enggist. Simulation of alpine chamois Rupicapra r. rupicapra habitat use. Ecological Modelling, 175(3):291–302, July 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004824.

Flipo:2004:BMR

[FEP^{+}04] Nicolas Flipo, Stephanie Even, Michel Poulin, Marie-Hélène Tusseau-Vuillemin, Thierry Ameziane, and Alain Dauta. Biogeochemical modelling at the river scale: plankton and periphyton dynamics: Grand Morin case study, France. Ecological Modelling, 176(3–4):333–347, September 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004000973.

Fall:2001:DSL

[FF01] Andrew Fall and Joseph Fall. A domain-specific language for models of landscape dynamics. Ecological Modelling, 141(1–3):1–18, July 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003349.

Fenlon:2006:MPF

[FF06] John S. Fenlon and Malcolm J. Faddy. Modelling predation in functional response. Ecological Modelling, 198 (1–2):154–162, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001694.

Fay:2006:LWZ

[FG06] Temple H. Fay and Johanna C. Greeff. Lion, wildebeest and zebra: a predator–prey model. Ecological Modelling, 196(1–2):237–244, July 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000652.

Frenod:2007:FST

[FG07] E. Frénod and E. Goubert. A first step towards modelling confinement of paralic ecosystems. Ecological Modelling, 200(1–2):139–148, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003139.

Fay:2008:TSC

[FG08] Temple H. Fay and Johanna C. Greeff. A three species competition model as a decision support tool. Ecological Modelling, 211(1–2):142–152, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004425.

Fischer:2008:ECS

[FGB08] Björn Fischer, Valeri Goldberg, and Christian Bernhofer. Effect of a coupled soil water–plant gas exchange on forest energy fluxes: Simulations with the coupled vegetation– boundary layer model HIRVAC. Ecological Modelling, 214(2–4):75–82, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000800121X.

Fay:2006:TMO

[FGEG06] Temple H. Fay, Johanna C. Greeff, Benjamin E. Eisenberg, and Hendrik T. Groeneveld. Testing the model for one predator and two mutualistic prey species. Ecological Modelling, 196(1–2):245–255, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000640.

Filho:2005:LVC

[FGFB05] Tarcísio M. Rocha Filho, Iram M. Gléria, Annibal Figueiredo, and Léon Brenig. The Lotka–Volterra canonical format. Ecological Modelling, 183(1):95–106, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004004338.

Fath:2007:CEP

[FH07a] Brian D. Fath and Geir Halnes. Cyclic energy pathways in ecological food webs. Ecological Modelling, 208(1):17–24, October 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007002487.

Fletcher:2007:RLE

[FH07b] C. S. Fletcher and David W. Hilbert. Resilience in landscape exploitation systems. Ecological Modelling, 201 (3–4):440–452, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004960.

Fukuda:2008:PAS

[FH08] Shinji Fukuda and Kazuaki Hiramatsu. Prediction ability and sensitivity of artificial intelligence-based habitat preference models for predicting spatial distribution of Japanese medaka (Oryzias latipes). Ecological Modelling, 215(4):301–313, July 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001786.

Fayram:2006:CCM

[FHE06] Andrew H. Fayram, Michael J. Hansen, and Timothy J. Ehlinger. Characterizing changes in maturity of lakes resulting from supplementation of walleye populations. Ecological Modelling, 197(1–2):103–115, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001049.

Frost:2009:PMM

[FHT^{+}09] Charles J. Frost, Scott E. Hygnstrom, Andrew J. Tyre, Kent M. Eskridge, David M. Baasch, Justin R. Boner, Gregory M. Clements, Jason M. Gilsdorf, Travis C. Kinsell, and Kurt C. Vercauteren. Probabilistic movement model with emigration simulates movements of deer in Nebraska, 1990– 2006. Ecological Modelling, 220(19):2481–2490, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004396.

Fielding:2004:ICS

[Fie04] Dennis J. Fielding. Intraspecific competition and spatial heterogeneity alter life history traits in an individualbased model of grasshoppers. Ecological Modelling, 175(2):169–187, July 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004757.

Figueira:2009:CDD

[Fig09] Will F. Figueira. Connectivity or demography: Defining sources and sinks in coral reef fish metapopulations. Ecological Modelling, 220(8):1126–1137, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000726.

Fiscus:2009:CNA

[Fis09] Daniel A. Fiscus. Comparative network analysis toward characterization of systemic organization for human– environmental sustainability. Ecological Modelling, 220 (22):3123–3132, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003287.

Fieberg:2005:AUE

[FJ05] John Fieberg and Kurt J. Jenkins. Assessing uncertainty in ecological systems using global sensitivity analyses: a case example of simulated wolf reintroduction effects on elk. Ecological Modelling, 187(2–3):259–280, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000827.

Farber:2003:AAA

[FK03] Oren Farber and Ronen Kadmon. Assessment of alternative approaches for bioclimatic modeling with special emphasis on the Mahalanobis distance. Ecological Modelling, 160(1–2):115–130, February 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003277.

Fink:2005:TDS

[FK05] M. Fink and A. Kofoet. A two-dimensional stochastic model of downy mildew of radish. Ecological Modelling, 181(2–3):139–148, January 20, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400362X.

Fath:2007:REP

[FK07] Brian D. Fath and Megan C. Killian. The relevance of ecological pyramids in community assemblages. Ecological Modelling, 208(2–4):286–294, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003079.

Freitas:2008:PHU

[FKIL08] Carla Freitas, Kit M. Kovacs, Rolf A. Ims, and Christian Lydersen. Predicting habitat use by ringed seals (Phoca hispida) in a warming Arctic. Ecological Modelling, 217(1–2):19–32, September 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002354.

Fursova:2007:VMM

[FL07] P. V. Fursova and A. P. Levich. Variational model of microorganism polyculture development without resupply of mutually irreplaceable resources. Ecological Modelling, 200(1–2):160–170, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003395.

Fan:2009:DAS

[FL09] Wei Fan and Xianqing Lv. Data assimilation in a simple marine ecosystem model based on spatial biological parameterizations. Ecological Modelling, 220(17):1997–2008, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000900310X.

Fleming:2001:WME

[Fle01] Richard A. Fleming. The Weibull model and an ecological application: describing the dynamics of foliage biomass on Scots pine. Ecological Modelling, 138(1–3):309–319, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004105.

Frisman:2006:PDH

[FLS06] E. Ya. Frisman, E. V. Last, and E. I. Skaletskaya. Population dynamics of harvested species with complex age structure (for Pacific salmons fish stocks as an example). Ecological Modelling, 198(3–4):463–472, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002511.

Freeman:2009:NCL

[FLS09] Angelina M. Freeman, E. Conrad Lamon, and Craig A. Stow. Nutrient criteria for lakes, ponds, and reservoirs: a Bayesian TREED model approach. Ecological Modelling, 220(5):630–639, March 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005887.

Fetahi:2007:TAL

[FM07] Tadesse Fetahi and Seyoum Mengistou. Trophic analysis of Lake Awassa (Ethiopia) using mass-balance Ecopath model. Ecological Modelling, 201(3–4):398–408, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004881.

Freeman:2008:CPT

[FM08] Elizabeth A. Freeman and Gretchen G. Moisen. A comparison of the performance of threshold criteria for binary classification in terms of predicted prevalence and kappa. Ecological Modelling, 217(1–2):48–58, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002275.

Fragoso:2008:MSH

[FMC^{+}08] Carlos R. Fragoso, David M. L. Motta Marques, Walter Collischonn, Carlos E. M. Tucci, and Egbert H. van Nes. Modelling spatial heterogeneity of phytoplankton in Lake Mangueira, a large shallow subtropical lake in South Brazil. Ecological Modelling, 219(1–2):125–137, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008004055.

Faerge:2001:UNB

[FMdV01] Jens Færge, Jakob Magid, and Frits W. T. Penning de Vries. Urban nutrient balance for Bangkok. Ecological Modelling, 139(1):63–74, March 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002332.

Facchini:2007:NTS

[FMM^{+}07] Angelo Facchini, Chiara Mocenni, Norbert Marwan, Antonio Vicino, and Enzo Tiezzi. Nonlinear time series analysis of dissolved oxygen in the Orbetello Lagoon (Italy). Ecological Modelling, 203(3–4):339–348, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006223.

Fonseca:2000:NDD

[FMP^{+}00] Júlio C. Fonseca, João C. Marques, Artur A. Paiva, Albertina M. Freitas, Vítor M. C. Madeira, and Sven E. Jørgensen. Nuclear DNA in the determination of weighing factors to estimate exergy from organisms biomass. Ecological Modelling, 126(2–3):179–189, February 28, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000002647.

Forcada:2009:MPT

[FMRS09] Jaume Forcada, Donald Malone, J. Andrew Royle, and Iain J. Staniland. Modelling predation by transient leopard seals for an ecosystem-based management of Southern Ocean fisheries. Ecological Modelling, 220(12):1513–1521, June 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002002.

Fohrer:2005:ASR

[Foh05] N. Fohrer. Advances in sustainable river basin management. Ecological Modelling, 187(1):1, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000359.

Fortin:2002:OSB

[For02] Daniel Fortin. Optimal searching behaviour: the value of sampling information. Ecological Modelling, 153 (3):279–290, August 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000315.

Fortin:2003:EOS

[For03] Daniel Fortin. Erratum to “Optimal searching behaviour: the value of sampling information”: [Ecological Modelling 153 (2002) 279–290]. Ecological Modelling, 159(1):101, January 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002002740.

Ferreyra:2000:MPL

[FPCA00] R. A. Ferreyra, Ludmila B. Pachepsky, D. Collino, and B. Acock. Modeling peanut leaf gas exchange for the calibration of crop models for different cultivars. Ecological Modelling, 131(2–3):285–298, July 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002520.

Fabre:2006:BYD

[FPDP06] F. Fabre, J. S. Pierre, C. A. Dedryver, and M. Plantegenest. Barley yellow dwarf disease risk assessment based on Bayesian modelling of aphid population dynamics. Ecological Modelling, 193(3–4):457–466, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004047.

Fyllas:2007:DPG

[FPK^{+}07] Nikolaos M. Fyllas, Oliver L. Phillips, William E. Kunin, Yiannis G. Matsinos, and Andreas I. Troumbis. Development and parameterization of a general forest gap dynamics simulator for the North-eastern Mediterranean Basin (GREek FOrest species). Ecological Modelling, 204(3–4):439–456, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007000531.

Freedman:2003:LHA

[FPS03] Adam H. Freedman, Kenneth M. Portier, and Melvin E. Sunquist. Life history analysis for black bears (Ursus americanus) in a changing demographic landscape. Ecological Modelling, 167(1–2):47–64, September 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001716.

Fulton:2004:BMEa

[FPSJ04] Elizabeth A. Fulton, John S. Parslow, Anthony D. M. Smith, and Craig R. Johnson. Biogeochemical marine ecosystem models II: the effect of physiological detail on model performance. Ecological Modelling, 173(4):371–406, April 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003004253.

Peng:2007:MTRb

[fPzWhSY07] Jian feng Peng, Bao zhen Wang, Yong hui Song, and Peng Yuan. Modeling N transformation and removal in a duckweed pond: Model application. Ecological Modelling, 206(3–4):294–300, August 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001858.

Falge:2005:CSE

[FRB^{+}05] Eva Falge, Sascha Reth, Nicolas Brüggemann, Klaus Butterbach-Bahl, Valeri Goldberg, Alexander Oltchev, Stefan Schaaf, Gerald Spindler, Bernd Stiller, Ronald Queck, Barbara Köstner, and Christian Bernhofer. Comparison of surface energy exchange models with eddy flux data in forest and grassland ecosystems of Germany. Ecological Modelling, 188(2–4):174–216, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001390.

Flores:2009:AOZ

[FRM09] O. Flores, V. Rossi, and F. Mortier. Autocorrelation offsets zero-inflation in models of tropical saplings density. Ecological Modelling, 220(15):1797–1809, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000271.

Freyer:2000:PIM

[FRZ00] Bernhard Freyer, Yvonne Reisner, and Dieter Zuberbühler. Potential impact model to assess agricultural pressure to landscape ecological functions. Ecological Modelling, 130(1–3):121–129, June 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002088.

Fa:2001:IBM

[FSBD01] John E. Fa, Colin M. Sharples, Diana J. Bell, and Don DeAngelis. An individual-based model of rabbit viral haemorrhagic disease in European wild rabbits (Oryctolagus cuniculus). Ecological Modelling, 144(2–3):121–138, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003635.

Fang:2004:STDa

[FSE^{+}04a] Xing Fang, Heinz G. Stefan, John G. Eaton, J. Howard McCormick, and Shoeb R. Alam. Simulation of thermal/ dissolved oxygen habitat for fishes in lakes under different climate scenarios:Part 1. Cool-water fish in the contiguous US. Ecological Modelling, 172(1):13–37, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002825.

Fang:2004:STDb

[FSE^{+}04b] Xing Fang, Heinz G. Stefan, John G. Eaton, J. Howard McCormick, and Shoeb R. Alam. Simulation of thermal/ dissolved oxygen habitat for fishes in lakes under different climate scenarios:Part 2. Cold-water fish in the contiguous US. Ecological Modelling, 172(1):39–54, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002850.

Fang:2004:STDc

[FSE^{+}04c] Xing Fang, Heinz G. Stefan, John G. Eaton, J. Howard McCormick, and Shoeb R. Alam. Simulation of thermal/ dissolved oxygen habitat for fishes in lakes under different climate scenarios:Part 3. Warm-water fish in the contiguous US. Ecological Modelling, 172(1):55–68, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002862.

Fulton:2003:MPE

[FSJ03] Elizabeth A. Fulton, Anthony D. M. Smith, and Craig R. Johnson. Mortality and predation in ecosystem models: is it important how these are expressed? Ecological Modelling, 169(1):157–178, November 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002680.

Fulton:2004:BMEb

[FSJ04a] Elizabeth A. Fulton, Anthony D. M. Smith, and Craig R. Johnson. Biogeochemical marine ecosystem models i: IGBEM — a model of marine bay ecosystems. Ecological Modelling, 174(3):267–307, May 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004290.

Fulton:2004:ESR

[FSJ04b] Elizabeth A. Fulton, Anthony D. M. Smith, and Craig R. Johnson. Effects of spatial resolution on the performance and interpretation of marine ecosystem models. Ecological Modelling, 176(1–2):27–42, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005416.

Franklin:2001:SED

[FSM^{+}01] Janet Franklin, Alexandra D. Syphard, David J. Mladenoff, Hong S. He, Dena K. Simons, Ross P. Martin, Douglas Deutschman, and John F. O’Leary. Simulating the effects of different fire regimes on plant functional groups in Southern California. Ecological Modelling, 142(3):261–283, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001002861.

Fath:2007:ENA

[FSUH07] Brian D. Fath, Ursula M. Scharler, Robert E. Ulanowicz, and Bruce Hannon. Ecological network analysis: network construction. Ecological Modelling, 208(1):49–55, October 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007002517.

Ferrarin:2005:HMC

[FU05] Christian Ferrarin and Georg Umgiesser. Hydrodynamic modeling of a coastal lagoon: the Cabras lagoon in Sardinia, Italy. Ecological Modelling, 188(2–4):340–357, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001675.

Fukuda:2009:CFI

[Fuk09] Shinji Fukuda. Consideration of fuzziness: Is it necessary in modelling fish habitat preference of Japanese medaka (Oryzias latipes)? Ecological Modelling, 220 (21):2877–2884, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009000222.

Feagin:2007:EEL

[FWF07] R. A. Feagin, X. B. Wu, and T. Feagin. Edge effects in lacunarity analysis. Ecological Modelling, 201 (3–4):262–268, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004467.

Feng:2006:AAM

[FWHL06] Jianfeng Feng, Hongli Wang, Dongwei Huang, and Shengpeng Li. Alternative attractors in marine ecosystems: a comparative analysis of fishing effects. Ecological Modelling, 195(3–4):377–384, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006137.

Feagin:2005:IVC

[FWS^{+}05] R. A. Feagin, X. B. Wu, F. E. Smeins, S. G. Whisenant, and W. E. Grant. Individual versus community level processes and pattern formation in a model of sand dune plant succession. Ecological Modelling, 183(4):435–449, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004005022.

Ferreira:2005:DBE

[FWSB05] J. G. Ferreira, W. J. Wolff, T. C. Simas, and S. B. Bricker. Does biodiversity of estuarine phytoplankton depend on hydrology? Ecological Modelling, 187(4):513–523, October 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001973.

Fujii:2007:CSC

[FYN^{+}07] Masahiko Fujii, Yasuhiro Yamanaka, Yukihiro Nojiri, Michio J. Kishi, and Fei Chai. Comparison of seasonal characteristics in biogeochemistry among the subarctic North Pacific stations described with a NEMURO-based marine ecosystem model. Ecological Modelling, 202(1–2):52–67, March 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000600456X.

Grant:2005:ITM

[GAA^{+}05] R. F. Grant, A. Arain, V. Arora, A. Barr, T. A. Black, J. Chen, S. Wang, F. Yuan, and Y. Zhang. Intercomparison of techniques to model high temperature effects on CO_{2} and energy exchange in temperate and boreal coniferous forests. Ecological Modelling, 188(2–4):217–252, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001407.

Groeneveld:2009:IFD

[GAB^{+}09] J. Groeneveld, L. F. Alves, L. C. Bernacci, E. L. M. Catharino, C. Knogge, J. P. Metzger, S. Pütz, and A. Huth. The impact of fragmentation and density regulation on forest succession in the Atlantic rain forest. Ecological Modelling, 220(19):2450–2459, October 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004153.

GomesLopes:2004:HWQ

[GACO04] Luis Filipe Gomes Lopes, José S. Antunes Do Carmo, Rui Manuel Vitor Cortes, and Daniel Oliveira. Hydrodynamics and water quality modelling in a regulated river segment: application on the instream flow definition. Ecological Modelling, 173(2–3):197–218, April 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003946.

Gan:2006:CAW

[Gan06] Jianbang Gan. Causasilty among wildfire, ENSO, timber harvest, and urban sprawl: the vector autoregression approach. Ecological Modelling, 191(2):304–314, January 31, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002784.

Gutierrez:2006:PBDb

[GAPE06] Andrew Paul Gutierrez, John J. Adamczyk, Sergine Ponsard, and C. K. Ellis. Physiologically based demographics of Bt cotton–pest interactions: II. Temporal refuges, natural enemy interactions. Ecological Modelling, 191(3–4):360–382, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003029.

Garman:2004:DEF

[Gar04] Steven L. Garman. Design and evaluation of a forest landscape change model for western Oregon. Ecological Modelling, 175(4):319–337, August 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004769.

Gaiser:2009:MCS

[GARB09] Thomas Gaiser, Mohammad Abdel-Razek, and Heike Bakara. Modeling carbon sequestration under zero-tillage at the regional scale. II. The influence of crop rotation and soil type. Ecological Modelling, 220(23):3372–3379, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009005316.

Gascuel:2005:TLB

[Gas05] Didier Gascuel. The trophic-level based model: a theoretical approach of fishing effects on marine ecosystems. Ecological Modelling, 189(3–4):315–332, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002115.

Gaucherel:2006:ISP

[Gau06] Cédric Gaucherel. Influence of spatial patterns on ecological applications of extremal principles. Ecological Modelling, 193(3–4):531–542, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004230.

Gaucherel:2008:NMP

[Gau08]  C. Gaucherel.  Neutral models for polygonal land-

scapes with linear networks. Ecological Modelling, 219 (1–2):39–48, November 24, 2008. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008003803.

Gullison:2001:SPT

[GB01] Jeremy J. Gullison and Charles P.-A. Bourque. Spatial prediction of tree and shrub succession in a small watershed in Northern Cape Breton Island, Nova Scotia, Canada. Ecological Modelling, 137(2–3):181–199, February 20, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000004282.

Gill:2002:AMA

[GB02] Samantha J. Gill and Gregory S. Biging. Autoregressive moving average models of crown profiles for two California hardwood species. Ecological Modelling, 152(2–3):213–226, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002000194.

Guneralp:2003:DMS

[GB03] Burak Güneralp and Yaman Barlas. Dynamic modelling of a shallow freshwater lake for ecological and economic sustainability. Ecological Modelling, 167(1–2):115–138, September 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001728.

Guinotte:2008:CMS

[GB08] John M. Guinotte and Robert W. Buddemeier. Comment on “Modelling susceptibility of coral reefs to environmental stress-using remote sensing data and GIS models”, authors Maina, Venus, McClanahan, and Ateweberhan. Ecological Modelling, 218(3–4):400–402, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002962. See [Dun08, MVMA08].

Golinski:2008:EEE

[GBA08] Michael Golinski, Chris Bauch, and Madhur Anand. The effects of endogenous ecological memory on population stability and resilience in a variable environment. Ecological Modelling, 212(3–4):334–341, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700587X.

Grimm:2006:SPD

[GBB^{+}06] Volker Grimm, Uta Berger, Finn Bastiansen, Sigrunn Eliassen, Vincent Ginot, Jarl Giske, John Goss-Custard, Tamara Grand, Simone K. Heinz, Geir Huse, Andreas Huth, Jane U. Jepsen, Christian Jørgensen, Wolf M. Mooij, Birgit Müller, Guy Pe’er, Cyril Piou, Steven F. Railsback, Andrew M. Robbins, Martha M. Robbins, Eva Rossmanith, Nadja Rüger, Espen Strand, Sami Souissi, Richard A. Stillman, Rune Vabø, Ute Visser, and Donald L. DeAngelis. A standard protocol for describing individual-based and agent-based models. Ecological Modelling, 198(1–2):115–126, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002043.

Galparsoro:2009:PSH

[GBB^{+}09] Ibon Galparsoro, Ángel Borja, Juan Bald, Pedro Liria, and Guillem Chust. Predicting suitable habitat for the European lobster (Homarus gammarus), on the Basque continental shelf (Bay of Biscay), using Ecological-Niche Factor Analysis. Ecological Modelling, 220(4):556–567, February 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005395.

Galbiati:2006:MDP

[GBEB06] L. Galbiati, F. Bouraoui, F. J. Elorza, and G. Bidoglio. Modeling diffuse pollution loading into a Mediterranean lagoon: Development and application of an integrated surface– subsurface model tool. Ecological Modelling, 193(1–2):4–18, March 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005004539.

Gillet:2002:PCM

[GBG02] François Gillet, Olivier Besson, and Jean-Michel Gobat. PA-TUMOD: a compartment model of vegetation dynamics in wooded pastures. Ecological Modelling, 147(3):267–290, January 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001004276.

Grizzetti:2003:MDE

[GBG^{+}03] B. Grizzetti, F. Bouraoui, K. Granlund, S. Rekolainen, and G. Bidoglio. Modelling diffuse emission and retention of nutrients in the Vantaanjoki watershed (Finland) using the SWAT model. Ecological Modelling, 169(1):25–38, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001984.

Garzon:2006:PHS

[GBN^{+}06] Marta Benito Garzón, Radim Blazek, Markus Neteler, Rut Sánchez de Dios, Helios Sainz Ollero, and Cesare Furlanello. Predicting habitat suitability with machine learning models: the potential area of Pinus sylvestris L. in the Iberian Peninsula. Ecological Modelling, 197 (3–4):383–393, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001384.

Genard:2000:PCM

[GBS00] M. Génard, F. Baret, and D. Simon. A 3D peach canopy model used to evaluate the effect of tree architecture and density on photosynthesis at a range of scales. Ecological Modelling, 128(2–3):197–209, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009900232X.

Garcia-Barrios:2008:EST

[GBSP08] L. E. García-Barrios, E. N. Speelman, and M. S. Pimm. An educational simulation tool for negotiating sustainable natural resource management strategies among stakeholders with conflicting interests. Ecological Modelling, 210 (1–2):115–126, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007003663.

Green:2002:ETD

[GC02] Edwin J. Green and Michael Clutter. Estimating tree diameter class frequencies. Ecological Modelling, 147(1):1–9, January 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001004021.

Grant:2007:BMC

[GCG^{+}07] Jon Grant, Kristian J. Curran, Thomas L. Guyondet, Guglielmo Tita, Cédric Bacher, Vladimir Koutitonsky, and Michael Dowd. A box model of carrying capacity for suspended mussel aquaculture in Lagune de la Grande– Entrée, Îles-de-la-Madeleine, Québec. Ecological Modelling, 200(1–2):193–206, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003620.

Goldstein:2003:MFM

[GCLG03] Michael I. Goldstein, Michael S. Corson, Thomas E. Lacher, and William E. Grant. Managed forests and migratory bird populations: evaluating spatial configurations through simulation. Ecological Modelling, 162(1–2):155–175, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000300005X.

Gantes:2005:ANP

[GCM^{+}05] P. Gantes, A. Sánchez Caro, F. Momo, M. A. Casset, and A. Torremorel. An approximation to the nitrogen and phosphorus budgets in floating soils of a subtropical peatland (Iberá, Argentina). Ecological Modelling, 186(1):77–83, July 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000500030X.

Gonzalez:2008:AFL

[GdBD08] Jesus Rodriguez Gonzalez, Gabriel del Barrio, and Beatriz Duguy. Assessing functional landscape connectivity for disturbance propagation on regional scales — a cost-surface model approach applied to surface fire spread. Ecological Modelling, 211(1–2):121–141, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004437.

Gaff:2000:DLM

[GDG^{+}00] Holly Gaff, Donald L. DeAngelis, Louis J. Gross, Rene Salinas, and Moris Shorrosh. A dynamic landscape model for fish in the Everglades and its application to restoration. Ecological Modelling, 127(1):33–52, March 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002021.

Gevrey:2003:RCM

[GDL03] Muriel Gevrey, Ioannis Dimopoulos, and Sovan Lek. Review and comparison of methods to study the contribution of variables in artificial neural network models. Ecological Modelling, 160(3):249–264, February 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002570.

Garnier:2006:SMF

[GDL06a] A. Garnier, A. Deville, and J. Lecomte. Stochastic modelling of feral plant populations with seed immigration and road verge management. Ecological Modelling, 197(3–4):373–382, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001621.

Gevrey:2006:TWI

[GDL06b] Muriel Gevrey, Ioannis Dimopoulos, and Sovan Lek. Two-way interaction of input variables in the sensitivity analysis of neural network models. Ecological Modelling, 195(1–2):43–50, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005005752.

Giacomini:2009:ECA

[GDP09] Henrique Corrêa Giacomini, Paulo De Marco, and Miguel Petrere. Exploring community assembly through an individual-based model for trophic interactions. Ecological Modelling, 220(1):23–39, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004389.

Gamito:2005:TFW

[GE05] Sofia Gamito and Karim Erzini. Trophic food web and ecosystem attributes of a water reservoir of the Ria Formosa (south Portugal). Ecological Modelling, 181 (4):509–520, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003849.

Guisan:2002:GLG

[GEH02] Antoine Guisan, Thomas C. Edwards, and Trevor Hastie. Generalized linear and generalized additive models in studies of species distributions: setting the scene. Ecological Modelling, 157(2–3):89–100, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002041.

Gaucherel:2006:NMP

[GFAD06] C. Gaucherel, D. Fleury, D. Auclair, and P. Dreyfus. Neutral models for patchy landscapes. Ecological Modelling, 197(1–2):159–170, August 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001074.

Gilboa:2009:AEE

[GFG09] Yael Gilboa, Eran Friedler, and Gideon Gal. Adapting empirical equations to Lake Kinneret data by using three calibration methods. Ecological Modelling, 220 (23):3291–3300, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006152.

Gehrig-Fasel:2008:ETT

[GFGZ08] Jacqueline Gehrig-Fasel, Antoine Guisan, and Niklaus E. Zimmermann. Evaluating thermal treeline indicators based on air and soil temperature using an air-to-soil temperature transfer model. Ecological Modelling, 213(3–4):345–355, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008000100.

Griebeler:2000:ITM

[GG00] Eva Maria Griebeler and Eckhard Gottschalk. The influence of temperature model assumptions on the prognosis accuracy of extinction risk. Ecological Modelling, 134 (2–3):343–356, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003732.

Gertseva:2002:MFP

[GG02] V. V. Gertseva and V. I. Gertsev. A model of fish population distribution in the space of inhabitation. Ecological Modelling, 147(2):161–170, January 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004082.

Gardner:2004:SDR

[GG04a] Robert H. Gardner and Eric J. Gustafson. Simulating dispersal of reintroduced species within heterogeneous landscapes. Ecological Modelling, 171(4):339–358, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003003387.

Gertsev:2004:CMM

[GG04b] V. I. Gertsev and V. V. Gertseva. Classification of mathematical models in ecology. Ecological Modelling, 178 (3–4):329–334, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001516.

Gui:2005:EHP

[GG05] Zhanji Gui and Weigao Ge. The effect of harvesting on a predator–prey system with stage structure. Ecological Modelling, 187(2–3):329–340, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500089X.

Gonzalez-Guzman:2008:SNS

[GG08a] Jorge González-Guzmán. Sexual and natural selection in a continuous time model. Ecological Modelling, 214(1):17–25, June 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008000215.

Gregorius:2008:GSD

[GG08b] Hans-Rolf Gregorius and Elizabeth M. Gillet. Generalized Simpson-diversity. Ecological Modelling, 211 (1–2):90–96, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004255.

Ginot:2006:CUL

[GGB^{+}06] Vincent Ginot, Sabrina Gaba, Rémy Beaudouin, Franck Aries, and Hervé Monod. Combined use of local and ANOVA-based global sensitivity analyses for the investigation of a stochastic dynamic model: Application to the case study of an individual-based model of a fish population. Ecological Modelling, 193(3–4):479–491, March 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005004084.

GarciayGarcia:2008:IGS

[GGH08] Axel Garcia y Garcia, Larry C. Guerra, and Gerrit Hoogenboom. Impact of generated solar radiation on simulated crop growth and yield. Ecological Modelling, 210(3):312–326, January 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007004103.

Garbey:2006:UMI

[GGM06] C. Garbey, M. Garbey, and S. Muller. Using modeling to improve models. Ecological Modelling, 197(3–4):303–319, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006001098.

Gertsev:2008:POS

[GGM08] Vladimir I. Gertsev, Vladlena V. Gertseva, and Sean E. Matson. Population as an oscillating system. Ecological Modelling, 210(3):242–246, January 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003705.

Gertseva:2003:TEP

[GGP03] V. V. Gertseva, V. I. Gertsev, and N. Y. Ponomarev. Tropho-ethological polymorphism of fish as a strategy of habitat development: a simulation model. Ecological Modelling, 167(1–2):159–164, September 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001741.

Gertseva:2006:IMP

[GGP06] Vladlena V. Gertseva, Vladimir I. Gertsev, and Nikolay Y. Ponomarev. Integrative model of a population distribution in a habitat. Ecological Modelling, 193(3–4):575–588, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004254.

Garcia-Gonzalo:2007:ECC

[GGPBEK07] Jordi Garcia-Gonzalo, Heli Peltola, Elemer Briceño-Elizondo, and Seppo Kellomäki. Effects of climate change and management on timber yield in boreal forests, with economic implications: a case study. Ecological Modelling, 209(2–4):220–234, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003390.

Gilad:2008:SDA

[GGS08] Oranit Gilad, William E. Grant, and David Saltz. Simulated dynamics of Arabian oryx (Oryx leucoryx) in the Israeli Negev: Effects of migration corridors and postreintroduction changes in natality on population viability. Ecological Modelling, 210(1–2):169–178, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003730.

Gaucherel:2006:DSL

[GGV^{+}06] Cédric. Gaucherel, Nathalie Giboire, Valérie Viaud, Thomas Houet, Jacques Baudry, and Françoise Burel. A domainspecific language for patchy landscape modelling: the Brittany agricultural mosaic as a case study. Ecological Modelling, 194(1–3):233–243, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005211.

Gilioli:2009:EME

[GGV^{+}09] G. Gilioli, M. Groppi, M. P. Vesperoni, J. Baumgärtner, and A. P. Gutierrez. An epidemiological model of East Coast fever in African livestock. Ecological Modelling, 220 (13–14):1652–1662, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009001938.

Grilli:2000:VFP

[GH00] M. P. Grilli and J. Holt. Vector feeding period variability in epidemiological models of persistent plant viruses. Ecological Modelling, 126(1):49–57, February 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001945.

Guven:2007:ICP

[GH07] Basak Guven and Alan Howard. Identifying the critical parameters of a cyanobacterial growth and movement model by using generalised sensitivity analysis. Ecological Modelling, 207(1):11–21, September 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001512.

Gaucherel:2009:PSP

[GH09] C. Gaucherel and T. Houet. Preface to the selected papers on spatially explicit landscape modelling: Current practices and challenges. Ecological Modelling, 220 (24):3477–3480, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004165.

Gros:2008:ESB

[GHP08] Andreas Gros, Thomas Hovestadt, and Hans Joachim Poethke. Evolution of sex-biased dispersal: the role of sexspecific dispersal costs, demographic stochasticity, and inbreeding. Ecological Modelling, 219(1–2):226–233, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008004158.

Gal:2009:IEM

[GHP^{+}09] G. Gal, M. R. Hipsey, A. Parparov, U. Wagner, V. Makler, and T. Zohary. Implementation of ecological modeling as an effective management and investigation tool: Lake Kinneret as a case study. Ecological Modelling, 220(13–14):1697–1718, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002336.

Giannantoni:2004:HD

[Gia04] Corrado Giannantoni. A harmonious dissonance. Ecological Modelling, 178(1–2):263–265, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300526X.

Guidi:2009:NPO

[GICB09] Lionel Guidi, Frédéric Ibanez, Vincent Calcagno, and Grégory Beaugrand. A new procedure to optimize the selection of groups in a classification tree: Applications for ecological data. Ecological Modelling, 220(4):451–461, February 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008005437.

Glenz:2008:MIF

[GIKS08] C. Glenz, I. Iorgulescu, F. Kienast, and R. Schlaepfer. Modelling the impact of flooding stress on the growth performance of woody species using fuzzy logic. Ecological Modelling, 218(1–2):18–28, October 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002974.

Gillet:2008:MVD

[Gil08] François Gillet. Modelling vegetation dynamics in heterogeneous pasture-woodland landscapes. Ecological Modelling, 217(1–2):1–18, September 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002548.

Gal:2003:STD

[GIZ^{+}03] Gideon Gal, Jorg Imberger, Tamar Zohary, Jason Antenucci, Ayal Anis, and Tzahi Rosenberg. Simulating the thermal dynamics of Lake Kinneret. Ecological Modelling, 162(1–2):69–86, April 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003800.

Gyldenkaerne:2000:MBP

[GJ00] Steen Gyldenkærne and Sven Erik Jørgensen. Modelling the bioavailability of pesticides to soil-dwelling organisms. Ecological Modelling, 132(3):203–230, August 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002416.

Guttal:2007:INB

[GJ07] Vishwesha Guttal and C. Jayaprakash. Impact of noise on bistable ecological systems. Ecological Modelling, 201 (3–4):420–428, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000600490X.

Ginzburg:2007:AUE

[GJY07] Lev R. Ginzburg, Christopher X. J. Jensen, and Jeffrey V. Yule. Aiming the “unreasonable effectiveness of mathematics” at ecological theory. Ecological Modelling, 207(2–4):356–362, October 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002876.

Guo:2005:SVM

[GKG05] Qinghua Guo, Maggi Kelly, and Catherine H. Graham. Support vector machines for predicting distribution of sudden oak death in California. Ecological Modelling, 182(1):75–90, February 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004077.

Gattie:2007:IEE

[GKT07] David K. Gattie, Nadia N. Kellam, and H. Jeff Turk. Informing ecological engineering through ecological network analysis, ecological modelling, and concepts of systems and engineering ecology. Ecological Modelling, 208(1):25–40, October 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007002499.

Gecek:2001:NGM

[GL01a] Sunčana Geček and Tarzan Legović. Nutrients and grazing in modeling the deep chlorophyll maximum. Ecological Modelling, 138(1–3):143–152, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003999.

Giraudel:2001:CSO

[GL01b] J. L. Giraudel and S. Lek. A comparison of self-organizing map algorithm and some conventional statistical methods for ecological community ordination. Ecological Modelling, 146(1–3):329–339, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003246.

Groot:2004:OOM

[GL04] Jeroen C. J. Groot and Egbert A. Lantinga. An objectoriented model of the morphological development and digestibility of perennial ryegrass. Ecological Modelling, 177(3–4):297–312, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001413.

Garnier:2006:USS

[GL06] A. Garnier and J. Lecomte. Using a spatial and stagestructured invasion model to assess the spread of feral populations of transgenic oilseed rape. Ecological Modelling, 194(1–3):141–149, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005120.

Guedon:2008:ATC

[GL08] Yann Guédon and Jean Michel Legave. Analyzing the timecourse variation of apple and pear tree dates of flowering stages in the global warming context. Ecological Modelling, 219(1–2):189–199, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004146.

Gamble:2009:ATA

[GL09] Robert J. Gamble and Jason S. Link. Analyzing the tradeoffs among ecological and fishing effects on an example fish community: a multispecies (fisheries) production model. Ecological Modelling, 220(19):2570–2582, October 10, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009003998.

Gao:2007:ESC

[GLD07] Meng Gao, Zizhen Li, and Huawei Dai. Effects on spectral color of a spatially-structured population: Environmental noise, dispersal, spatial heterogeneity. Ecological Modelling, 201(3–4):326–330, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004492.

Gasca-Leyva:2008:OHT

[GLHV08] Eucario Gasca-Leyva, Juan M. Hernández, and Vladimir M. Veliov. Optimal harvesting time in a size-heterogeneous population. Ecological Modelling, 210(1–2):161–168, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003729.

Gao:2007:SSH

[GLL^{+}07] Meng Gao, Wenlong Li, Zizhen Li, Huawei Dai, and Hongtao Liu. Spatial synchrony in host–parasitoid populations. Ecological Modelling, 204(1–2):29–39, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006405.

Goreaud:2002:SSS

[GLM02] François Goreaud, Michel Loreau, and Claude Millier. Spatial structure and the survival of an inferior competitor: a theoretical model of neighbourhood competition in plants. Ecological Modelling, 158(1–2):1–19, December 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002000583.

Gaucel:2005:IIS

[GLP05] Sebastien Gaucel, Michel Langlais, and Dominique Pontier. Invading introduced species in insular heterogeneous environments. Ecological Modelling, 188(1):62–75, October 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002425.

Ginot:2002:MAA

[GLS02] Vincent Ginot, Christophe Le Page, and Sami Souissi. A multi-agents architecture to enhance end-user individualbased modelling. Ecological Modelling, 157(1):23–41, November 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380002002119.

Gwaltney:2007:CES

[GLS07] C. Ryan Gwaltney, WenTao Luo, and Mark A. Stadtherr. Computation of equilibrium states and bifurcations using interval analysis: Application to food chain models. Ecological Modelling, 203(3–4):495–510, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006430.

Gomez-Mourelo:2005:IBM

[GM05] Pablo Gómez-Mourelo. From individual-based models to partial differential equations: an application to the upstream movement of elvers. Ecological Modelling, 188(1):93–111, October 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002449.

Giusti:2005:MIB

[GML05] Elisabetta Giusti and Stefano Marsili-Libelli. Modelling the interactions between nutrients and the submersed vegetation in the Orbetello Lagoon. Ecological Modelling, 184(1):141– 161, May 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004005897.

Giusti:2006:IMO

[GML06] Elisabetta Giusti and Stefano Marsili-Libelli. An integrated model for the Orbetello lagoon ecosystem. Ecological Modelling, 196(3–4):379–394, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000767.

Giusti:2009:STD

[GML09] Elisabetta Giusti and Stefano Marsili-Libelli. Spatiotemporal dissolved oxygen dynamics in the Orbetello lagoon by fuzzy pattern recognition. Ecological Modelling, 220(19):2415–2426, October 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004013.

Gascuel:2008:TFK

[GMPC08] Didier Gascuel, Lyne Morissette, Maria Lourdes D. Palomares, and Villy Christensen. Trophic flow kinetics in marine ecosystems: Toward a theoretical approach to ecosystem functioning. Ecological Modelling, 217(1– 2):33–47, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002263.

Goodwin:2006:FDF

[GNA^{+}06] R. Andrew Goodwin, John M. Nestler, James J. Anderson, Larry J. Weber, and Daniel P. Loucks. Forecasting 3-D fish movement behavior using a Eulerian– Lagrangian-agent method (ELAM). Ecological Modelling, 192(1–2):197–223, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003327.

Goforth:2004:CSD

[Gof04] Gary Goforth. Cypress Swamps: demonstrating the utility of swamps. Ecological Modelling, 178(1–2):75–76, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005271.

Gonzalez-Olivares:2003:DCP

[GORJ03] Eduardo González-Olivares and Rodrigo Ramos-Jiliberto. Dynamic consequences of prey refuges in a simple model system: more prey, fewer predators and enhanced stability. Ecological Modelling, 166(1–2):135–146, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001315.

Garcet:2006:MTC

[GORV06] J. D. Piñeros Garcet, A. Ordoñez, J. Roosen, and M. Vanclooster. Metamodelling: Theory, concepts and application to nitrate leaching modelling. Ecological Modelling, 193(3–4):629–644, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004461.

Givens:2002:PLF

[GP02] Geof H. Givens and David Poole. Problematic likelihood functions from sensible population dynamics models: a case study. Ecological Modelling, 151(2–3):109–124, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004161.

Gutierrez:2006:PBDa

[GP06] Andrew Paul Gutierrez and Sergine Ponsard. Physiologically based demographics of Bt cotton–pest interactions: I. Pink bollworm resistance, refuge and risk. Ecological Modelling, 191(3–4):346–359, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003017.

Gilioli:2007:UIB

[GP07] Gianni Gilioli and Sara Pasquali. Use of individual-based models for population parameters estimation. Ecological Modelling, 200(1–2):109–118, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003474.

Gascuel:2009:PEM

[GP09] Didier Gascuel and Daniel Pauly. EcoTroph: Modelling marine ecosystem functioning and impact of fishing. Ecological Modelling, 220(21):2885–2898, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005286.

Goslee:2001:MIW

[GPB01] S. C. Goslee, D. P. C. Peters, and K. G. Beck. Modeling invasive weeds in grasslands: the role of allelopathy in Acroptilon repens invasion. Ecological Modelling, 139(1):31–45, March 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002319.

Guo:2009:UAH

[GPC^{+}09] Y. Guo, G. Poulton, P. Corke, G. J. Bishop-Hurley, T. Wark, and D. L. Swain. Using accelerometer, high sample rate GPS and magnetometer data to develop a cattle movement and behaviour model. Ecological Modelling, 220(17):2068–2075, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003135.

Gall:2009:PMF

[GPDF09] Anne Roué-Le Gall, Michel Poulin, Didier Debroas, and Nicolas Flipo. A physical-microbial food web coupled model to study the evolution of the ecological functioning of a new reservoir after its flooding (Sep, Puy de Dôme). Ecological Modelling, 220(6):841–856, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005802.

Gardiner:2000:CTM

[GPK00] Barry Gardiner, Heli Peltola, and Seppo Kellomäki. Comparison of two models for predicting the critical wind speeds required to damage coniferous trees. Ecological Modelling, 129(1):1–23, May 3, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002209.

Guber:2005:FMS

[GPL05] A. K. Guber, Ya. A. Pachepsky, and E. V. Levkovsky. Fractal mass–size scaling of wetting soil aggregates. Ecological Modelling, 182(3–4):317–322, March 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001668.

Grant:2002:QMC

[GPP02] W. E. Grant, Tarla Rai Peterson, and Markus J. Peterson. Quantitative modeling of coupled natural/human systems: simulation of societal constraints on environmental action drawing on Luhmann’s social theory. Ecological Modelling, 158(1–2):143–165, December 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002193.

Ganev:2008:APT

[GPSM08] Kostadin Ganev, Maria Prodanova, Dimiter Syrakov, and Nikolai Miloshev. Air pollution transport in the Balkan region and country-to-country pollution exchange between Romania, Bulgaria and Greece. Ecological Modelling, 217(3–4):255–269, October 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000800255X.

Gabusi:2008:FSA

[GPV08] Veronica Gabusi, Enrico Pisoni, and Marialuisa Volta. Factor separation in air quality simulations. Ecological Modelling, 218(3–4):383–392, November 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003797.

Garcia-Quijano:2005:ICP

[GQB05] Juan F. Garcia-Quijano and Ana P. Barros. Incorporating canopy physiology into a hydrological model: photosynthesis, dynamic respiration, and stomatal sensitivity. Ecological Modelling, 185(1):29–49, June 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005927.

Gergs:2009:PFR

[GR09] André Gergs and Hans Toni Ratte. Predicting functional response and size selectivity of juvenile Notonecta maculata foraging on Daphnia magna. Ecological Modelling, 220(23):3331–3341, December 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005729.

Gray:2004:GML

[Gra04] David R. Gray. The gypsy moth life stage model: landscape-wide estimates of gypsy moth establishment using a multi-generational phenology model. Ecological Modelling, 176(1–2):155–171, August 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005623.

Gray:2005:SDA

[Gra05] Brian R. Gray. Selecting a distributional assumption for modelling relative densities of benthic macroinvertebrates. Ecological Modelling, 185(1):1–12, June 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005800.

Graniero:2007:ILH

[Gra07] Phil A. Graniero. The influence of landscape heterogeneity and local habitat effects on the response to competitive pressures in metapopulations. Ecological Modelling, 203(3–4):349–362, May 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006235.

Girardin:2008:RTG

[GRBT08] Martin P. Girardin, Frédéric Raulier, Pierre Y. Bernier, and Jacques C. Tardif. Response of tree growth to a changing climate in boreal central Canada: a comparison of empirical, process-based, and hybrid modelling approaches. Ecological Modelling, 213(2):209–228, May 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700631X.

Gimenez:2007:SSM

[GRC^{+}07] Olivier Gimenez, Vivien Rossi, Rémi Choquet, Camille Dehais, Blaise Doris, Hubert Varella, Jean-Pierre Vila, and Roger Pradel. State-space modelling of data on marked individuals. Ecological Modelling, 206(3–4):431–438, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001925.

Gregorius:2006:RSG

[Gre06] Hans-Rolf Gregorius. Revealing spatial genetic structure through cluster analyses. Ecological Modelling, 198 (3–4):312–320, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006002110.

Gouze:2000:IOU

[GRHS00] J. L. Gouzé, A. Rapaport, and M. Z. Hadj-Sadok. Interval observers for uncertain biological systems. Ecological Modelling, 133(1–2):45–56, August 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002799.

Gribble:2004:SEM

[Gri04] N. A. Gribble. A spatially explicit multi-competitor coexistence model of penaeid (shrimp) distribution on the Australian Great Barrier Reef. Ecological Modelling, 177(1–2):61–74, September 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001073.

Grilli:2008:AWM

[Gri08] Mariano P. Grilli. An area-wide model approach for the management of a disease vector planthopper in an extensive agricultural system. Ecological Modelling, 213(3–4):308–318, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007006345.

Grunwald:2007:MSV

[GRPF07] S. Grunwald, K. R. Reddy, J. P. Prenger, and M. M. Fisher. Modeling of the spatial variability of biogeochemical soil properties in a freshwater ecosystem. Ecological Modelling, 201(3–4):521–535, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600528X.

Gerber:2004:FHF

[GRR04] L. R. Gerber, O. J. Reichman, and J. Roughgarden. Food hoarding: future value in optimal foraging decisions. Ecological Modelling, 175(1):77–85, June 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004691.

Gasalla:2004:CEA

[GRW04] M. A. Gasalla and C. L. D. B. Rossi-Wongtschowski. Contribution of ecosystem analysis to investigating the effects of changes in fishing strategies in the South Brazil Bight coastal ecosystem. Ecological Modelling, 172(2–4):283–306, March 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003003740.

Griebeler:2002:IBM

[GS02] Eva Maria Griebeler and Alfred Seitz. An individual based model for the conservation of the endangered Large Blue Butterfly, Maculinea arion (Lepidoptera: Lycaenidae). Ecological Modelling, 156(1):43–60, October 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200131X.

Griebeler:2006:UMM

[GS06] Eva Maria Griebeler and Alfred Seitz. The use of Markovian metapopulation models: Reducing the dimensionality of transition matrices by self-organizing Kohonen networks. Ecological Modelling, 192(1–2):271–285, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003042.

Golubyatnikov:2008:LCM

[GS08] Leonid L. Golubyatnikov and Yuri M. Svirezhev. Lifecycle model of terrestrial carbon exchange. Ecological Modelling, 213(2):202–208, May 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006308.

Grabarnik:2009:MSS

[GS09] Pavel Grabarnik and Aila Särkkä. Modelling the spatial structure of forest stands by multivariate point processes with hierarchical interactions. Ecological Modelling, 220(9–10):1232–1240, May 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001537.

Gabriel:2005:PLE

[GSB05] Jean-Pierre Gabriel, Francis Saucy, and Louis-Félix Bersier. Paradoxes in the logistic equation? Ecological Modelling, 185(1):147–151, June 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400554X.

Gabriel:2006:AH

[GSB06a] Jean-Pierre Gabriel, Francis Saucy, and Louis-Félix Bersier. Answer to Hui. Ecological Modelling, 192(1–2):321, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003303.

Gattie:2006:IED

[GSB^{+}06b] D. K. Gattie, J. R. Schramski, S. R. Borrett, B. C. Patten, S. A. Bata, and S. J. Whipple. Indirect effects and distributed control in ecosystems: Network environ analysis of a seven-compartment model of nitrogen flow in the Neuse River Estuary, USA — steady-state analysis. Ecological Modelling, 194(1–3):162–177, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005144.

Gaiser:2008:MCS

[GSBM08] Thomas Gaiser, Karl Stahr, Norbert Billen, and Mohammad Abdel-Razek Mohammad. Modeling carbon sequestration under zero tillage at the regional scale. I. The effect of soil erosion. Ecological Modelling, 218(1–2):110–120, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003074.

Gomez-Sal:2003:ALV

[GSBN03] Antonio Gómez-Sal, Juan-Alfonso Belmontes, and José-Manuel Nicolau. Assessing landscape values: a proposal for a multidimensional conceptual model. Ecological Modelling, 168(3):319–341, October 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001443.

Gutierrez:2009:UCT

[GSC09] Álvaro Gómez Gutiérrez, Susanne Schnabel, and J. Francisco Lavado Contador. Using and comparing two nonparametric methods (CART and MARS) to model the potential distribution of gullies. Ecological Modelling, 220(24):3630–3637, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004104.

Gertseva:2004:SMD

[GSG^{+}04] V. V. Gertseva, J. E. Schindler, V. I. Gertsev, N. Y. Ponomarev, and W. R. English. A simulation model of the dynamics of aquatic macroinvertebrate communities. Ecological Modelling, 176(1–2):173–186, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000201.

Guarini:2008:MDM

[GSM08] Jean-Marc Guarini, Nadir Sari, and Charlotte Moritz. Modelling the dynamics of the microalgal biomass in semi-enclosed shallow-water ecosystems. Ecological Modelling, 211(3–4):267–278, March 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004590.

Ganev:2008:NPS

[GSZ08] Kostadin Ganev, Dimiter Syrakov, and Zahari Zlatev. New parameterization scheme for effective indices for emissions from road transport. Ecological Modelling, 217(3–4):270–278, October 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002585.

Guiot:2000:IVM

[GTJ^{+}00] J. Guiot, F. Torre, D. Jolly, O. Peyron, J. J. Boreux, and R. Cheddadi. Inverse vegetation modeling by Monte Carlo sampling to reconstruct palaeoclimates under changed precipitation seasonality and CO_{2} conditions: application to glacial climate in Mediterranean region. Ecological Modelling, 127(2–3):119–140, March 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002197.

Gosselke:2001:GLV

[GTRF01] U. Gosselke, H. Triltsch, D. Roßberg, and B. Freier. GET-LAUS01 — the latest version of a model for simulating aphid population dynamics in dependence on antagonists in wheat. Ecological Modelling, 145(2–3):143–157, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003866.

Gutschick:2007:PAE

[Gut07] Vincent P. Gutschick. Plant acclimation to elevated CO_{2} — from simple regularities to biogeographic chaos. Ecological Modelling, 200(3–4):433–451, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003917.

Guala:2009:MSV

[GVC09] Sebastián Guala, Flora A. Vega, and Emma F. Covelo. Modification of a soil–vegetation nonlinear interaction model with acid deposition for simplified experimental applicability. Ecological Modelling, 220(18):2137–2141, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009003317.

Goulart:2007:VLL

[GVDF07] Antonio G. Goulart, Marco T. Vilhena, Gervásio A. Degrazia, and Denise T. Flores. Vertical, lateral and longitudinal eddy diffusivities for a decaying turbulence in the convective boundary layer. Ecological Modelling, 204(3–4):516–522, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007000579.

Grasman:2009:DDM

[GvNK09] Johan Grasman, Egbert H. van Nes, and Kees Kersting. Data-directed modelling of Daphnia dynamics in a long-term micro-ecosystem experiment. Ecological Modelling, 220(3):343–350, February 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800495X.

Grund:2004:DEA

[GW04] Marrett D. Grund and Alan Woolf. Development and evaluation of an accounting model for estimating deer population sizes. Ecological Modelling, 180(2–3):345–357, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002868.

Growns:2008:CAB

[GW08] Ivor Growns and Greg West. Classification of aquatic bioregions through the use of distributional modelling of freshwater fish. Ecological Modelling, 217(1–2):79–86, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008002883.

Getz:2003:CAT

[GWHS03] Wayne M. Getz, Hans V. Westerhoff, Jan-Hendrik S. Hofmeyr, and Jacky L. Snoep. Control analysis of trophic chains. Ecological Modelling, 168(1–2):153–171, October 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002084.

Gevrey:2006:ERE

[GWK^{+}06] Muriel Gevrey, Sue Worner, Nikola Kasabov, Joel Pitt, and Jean-Luc Giraudel. Estimating risk of events using SOM models: a case study on invasive species establishment. Ecological Modelling, 197(3–4):361–372, August 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006001633.

Gao:2004:RBR

[GYW^{+}04] Qiong Gao, Mei Yu, Junhua Wang, Haikun Jia, and Kun Wang. Relationships between regional primary production and vegetation patterns. Ecological Modelling, 172(1):1–12, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002503.

Gao:2000:SAR

[GYY00] Qiong Gao, Mei Yu, and Xiusheng Yang. A simulation analysis of the relationship between regional primary production and vegetation structure under climatic change scenarios. Ecological Modelling, 131(1):33–45, June 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002477.

Guisan:2000:PHD

[GZ00] Antoine Guisan and Niklaus E. Zimmermann. Predictive habitat distribution models in ecology. Ecological Modelling, 135(2–3):147–186, December 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003549.

Guo:2005:DMP

[GZ05] K. Guo and I. Zoccarato. A dynamic model to predict the nitrogen excretion in growing-finishing cattle. Ecological Modelling, 187(2–3):219–231, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000621.

Gurkan:2006:DSD

[GZJ06] Zeren Gurkan, Jingjie Zhang, and Sven Erik Jørgensen. Development of a structurally dynamic model for forecasting the effects of restoration of Lake Fure, Denmark. Ecological Modelling, 197(1–2):89–102, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001165.

Grant:2006:ITM

[GZY^{+}06] R. F. Grant, Y. Zhang, F. Yuan, S. Wang, P. J. Hanson, D. Gaumont-Guay, J. Chen, T. A. Black, A. Barr, D. D. Baldocchi, and A. Arain. Intercomparison of techniques to model water stress effects on CO_{2} and energy exchange in temperate and boreal deciduous forests. Ecological Modelling, 196(3–4):289–312, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000974.

Haltar:2008:OMA

[HA08a] D. Haltar and B. Altansuvd. Optimization models of an agespecific herd. Ecological Modelling, 216(2):145–149, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001336.

Haltar:2008:UMP

[HA08b] D. Haltar and G. Ankhbayar. Using the maximum principle of impulse control for ecology-economical models. Ecological Modelling, 216(2):150–156, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001348.

Haltar:2008:OHM

[HAA08] D. Haltar, G. Ankhbayar, and B. Altansuvd. The optimal harvest and management in the models of animal populations. Ecological Modelling, 216(2):240–244, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001476.

Han:2000:TSS

[HAG^{+}00] Bo-Ping Han, Joan Armengol, Juan Carlos Garcia, Marta Comerma, Montse Roura, Josep Dolz, and Milan Straskraba. The thermal structure of Sau Reservoir (NE: Spain): a simulation approach. Ecological Modelling, 125(2–3):109–122, January 15, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001763.

Haakanson:2000:RCC

[Håk00] Lars Håkanson. The role of characteristic coefficients of variation in uncertainty and sensitivity analyses, with examples related to the structuring of lake eutrophication models. Ecological Modelling, 131(1):1–20, June 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002192.

Haakanson:2009:GPB

[Håk09] Lars Håkanson. A general process-based mass-balance model for phosphorus/eutrophication as a tool to estimate historical reference values for key bioindicators, as exemplified using data for the Gulf of Riga. Ecological Modelling, 220(2):226–244, January 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004341.

Hall:2004:CIM

[Hal04] Charles A. S. Hall. The continuing importance of maximum power. Ecological Modelling, 178(1–2):107–113, October 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004001206.

Hannunen:2002:VAR

[Han02] Salla Hannunen. Vegetation architecture and redistribution of insects moving on the plant surface. Ecological Modelling, 155(2–3):149–157, October 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001254.

Hanson:2009:PLE

[Han09] Niklas Hanson. Population level effects of reduced fecundity in the fish species perch (Perca fluviatilis) and the implications for environmental monitoring. Ecological Modelling, 220(17):2051–2059, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003081.

Hartbauer:2008:CMS

[Har08] Manfred Hartbauer. Chorus model of the synchronizing bushcricket species Mecopoda elongata. Ecological Modelling, 213(1):105–118, April 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006205.

Holland:2007:SSG

[HAS07] E. P. Holland, J. N. Aegerter, and G. C. Smith. Spatial sensitivity of a generic population model, using wild boar (Sus scrofa) as a test case. Ecological Modelling, 205(1–2):146–158, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007000853.

Hawkes:2000:WPM

[Haw00] Corinna Hawkes. Woody plant mortality algorithms: description, problems and progress. Ecological Modelling, 126(2–3):225–248, February 28, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002672.

Heymans:2000:CFM

[HB00a] J. J. Heymans and D. Baird. A carbon flow model and network analysis of the northern Benguela upwelling system, Namibia. Ecological Modelling, 126(1):9–32, February 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099001921.

Heymans:2000:NAN

[HB00b] J. J. Heymans and D. Baird. Network analysis of the northern Benguela ecosystem by means of Netwrk and Ecopath. Ecological Modelling, 131(2–3):97–119, July 1, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000002751.

Haakanson:2001:PAP

[HB01] Lars Håkanson and Viktor V. Boulion. A practical approach to predict the duration of the growing season for European lakes. Ecological Modelling, 140(3):235–245, June 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003192.

Haefner:2002:PBM

[HB02a] James W. Haefner and Mark D. Bowen. Physical-based model of fish movement in fish extraction facilities. Ecological Modelling, 152(2–3):227–245, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000066.

Haakanson:2002:EDM

[HB02b] Lars Håkanson and Viktor V. Boulion. Empirical and dynamical models to predict the cover, biomass and production of macrophytes in lakes. Ecological Modelling, 151(2–3):213–243, June 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004586.

Haakanson:2003:GDM

[HB03a] Lars Håkanson and Viktor V. Boulion. A general dynamic model to predict biomass and production of phytoplankton in lakes. Ecological Modelling, 165(2–3):285–301, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000966.

Haakanson:2003:MPB

[HB03b] Lars Håkanson and Viktor V. Boulion. Modelling production and biomasses of herbivorous and predatory zooplankton in lakes. Ecological Modelling, 161(1–2):1–33, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002107.

Hau:2004:PPE

[HB04] Jorge L. Hau and Bhavik R. Bakshi. Promise and problems of emergy analysis. Ecological Modelling, 178(1–2):215–225, October 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005283.

Holker:2005:SIB

[HB05] Franz Hölker and Broder Breckling. A spatiotemporal individual-based fish model to investigate emergent properties at the organismal and the population level. Ecological Modelling, 186(4):406–426, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001341.

Hense:2006:TMC

[HB06] Inga Hense and Aike Beckmann. Towards a model of cyanobacteria life cycle — effects of growing and resting stages on bloom formation of N_{2}-fixing species. Ecological Modelling, 195(3–4):205–218, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005661.

Hellweger:2009:BTI

[HB09a] Ferdi L. Hellweger and Vanni Bucci. A bunch of tiny individuals — individual-based modeling for microbes. Ecological Modelling, 220(1):8–22, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004390.

Hopkins:2009:MFF

[HB09b] Robert L. Hopkins and Brooks M. Burr. Modeling freshwater fish distributions using multiscale landscape data: a case study of six narrow range endemics. Ecological Modelling, 220(17):2024–2034, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002920.

Huang:2007:SOH

[HBCL07] Shu-Li Huang, William W. Budd, Shih-Liang Chan, and Ying-Chen Lin. Stream order, hierarchy, and energy convergence of land use. Ecological Modelling, 205(1–2):255–264, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000700097X.

Holland:2009:MUI

[HBDA09] E. P. Holland, J. F. Burrow, C. Dytham, and J. N. Aegerter. Modelling with uncertainty: Introducing a probabilistic framework to predict animal population dynamics. Ecological Modelling, 220(9–10):1203–1217, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001239.

Haakanson:2004:NGM

[HBM04] Lars Håkanson, Thorsten Blenckner, and J. Mikael Malmaeus. New, general methods to define the depth separating surface water from deep water, outflow and internal loading for mass-balance models for lakes. Ecological Modelling, 175(4):339–352, August 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004848.

Hartman:2007:ACE

[HBO07] Melannie D. Hartman, Jill S. Baron, and Dennis S. Ojima. Application of a coupled ecosystem-chemical equilibrium model, DayCent-Chem, to stream and soil chemistry in a Rocky Mountain watershed. Ecological Modelling, 200(3–4):493–510, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003942.

Howitt:2007:MBP

[HBRW07] Julia A. Howitt, Darren S. Baldwin, Gavin N. Rees, and Janice L. Williams. Modelling blackwater: Predicting water quality during flooding of lowland river forests. Ecological Modelling, 203(3–4):229–242, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005692.

Hubbard:2004:MFF

[HBSM04] Simon Hubbard, Petro Babak, Sven Th. Sigurdsson, and Kjartan G. Magnússon. A model of the formation of fish schools and migrations of fish. Ecological Modelling, 174(4):359–374, June 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004356.

Hof:2002:OHL

[HBUS02] John Hof, Michael Bevers, Daniel W. Uresk, and Gregory L. Schenbeck. Optimizing habitat location for blacktailed prairie dogs in southwestern South Dakota. Ecological Modelling, 147(1):11–21, January 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100401X.

Hean:2003:GMG

[HC03a] Robyn L. Hean and Oscar J. Cacho. A growth model for giant clams Tridacna crocea and T. derasa. Ecological Modelling, 163(1–2):87–100, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004003.

Hooley:2003:MFL

[HC03b] J. L. Hooley and E. V. J. Cohn. Models of field layer vegetation interactions in an experimental secondary woodland. Ecological Modelling, 169(1):89–102, November 1, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002643.

Huang:2005:TUE

[HC05a] Shu-Li Huang and Chia-Wen Chen. Theory of urban energetics and mechanisms of urban development. Ecological Modelling, 189(1–2):49–71, November 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002000.

Hunter:2005:UVP

[HC05b] Christine M. Hunter and Hal Caswell. The use of the vecpermutation matrix in spatial matrix population models. Ecological Modelling, 188(1):15–21, October 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002346.

Holt:2006:MCS

[HC06] J. Holt and J. F. Cooper. A model to compare the suitability of locust hopper targets for control by insecticide barriers. Ecological Modelling, 195(3–4):273–280, June 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005005776.

Hoyle:2009:MFD

[HC09] Martin Hoyle and James E. Cresswell. Maximum feasible distance of windborne cross-pollination in Brassica napus: a ‘mass budget’ model. Ecological Modelling, 220(8):1090–1097, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009000404.

Han:2009:SDM

[HCJ^{+}09] Hongjuan Han, Yonggen Chen, Sven Erik Jørgensen, Søren Nors Nielsen, and Weiping Hu. A system-dynamic model on the competitive growth between Potamogeton malaianus Miq. and Spirogyra sp. Ecological Modelling, 220(18):2206–2217, September 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003299.

Huntingford:2000:CRS

[HCL00] C. Huntingford, P. M. Cox, and T. M. Lenton. Contrasting responses of a simple terrestrial ecosystem model to global change. Ecological Modelling, 134(1):41–58, September 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003306.

Hauser:2006:CSP

[HCL06] C. E. Hauser, E. G. Cooch, and J.-D. Lebreton. Control of structured populations by harvest. Ecological Modelling, 196(3–4):462–470, July 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000743.

Hull:2005:P

[HCLR05] Vincent Hull, Graciela Canziani, Steven Loiselle, and Claudio Rossi. Preface. Ecological Modelling, 186(1):1–2, July 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000189.

Haefner:2000:IBM

[HD00a] James W. Haefner and Christopher J. Dugaw. Individualbased models solved using fast Fourier transforms. Ecological Modelling, 125(2–3):159–172, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001799.

Huth:2000:SGL

[HD00b] A. Huth and T. Ditzer. Simulation of the growth of a lowland dipterocarp rain forest with FORMIX3. Ecological Modelling, 134(1):1–25, September 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003288.

Huettmann:2001:SCL

[HD01] F. Huettmann and A. W. Diamond. Seabird colony locations and environmental determination of seabird distribution: a spatially explicit breeding seabird model for the Northwest Atlantic. Ecological Modelling, 141(1–3):261–298, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002782.

Hedger:2006:IMJ

[HDB^{+}06] R. D. Hedger, J. J. Dodson, J-F. Bourque, N. E. Bergeron, and P. E. Carbonneau. Improving models of juvenile Atlantic salmon habitat use through high resolution remote sensing. Ecological Modelling, 197(3–4):505–511, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006001517.

Hay:2002:ESS

[HDBM02a] G. J. Hay, P. Dubé, A. Bouchard, and D. J. Marceau. Erratum to “A scale-space primer for exploring and quantifying complex landscapes” [Ecological Modelling 153 (2002) 27–49]. Ecological Modelling, 157(1):87–88, November 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000200265X.

Hay:2002:SSP

[HDBM02b] G. J. Hay, P. Dubé, A. Bouchard, and D. J. Marceau. A scale-space primer for exploring and quantifying complex landscapes. Ecological Modelling, 153(1–2):27–49, July 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001005002.

Hardi:2000:IAD

[HDH00] Peter Hardi and Juanita Ama DeSouza-Huletey. Issues in analyzing data and indicators for sustainable development. Ecological Modelling, 130(1–3):59–65, June 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002027.

Haakanson:2005:SPMb

[HE05] Lars Håkanson and Jan Eckhéll. Suspended particulate matter (SPM) in the Baltic Sea — new empirical data and models. Ecological Modelling, 189(1–2):130–150, November 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001985.

Hosack:2009:DMP

[HE09] Geoffrey R. Hosack and Peter M. Eldridge. Do microbial processes regulate the stability of a coral atoll’s enclosed pelagic ecosystem? Ecological Modelling, 220 (20):2665–2682, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004888.

Heegaard:2002:OBC

[Hee02] Einar Heegaard. The outer border and central border for species–environmental relationships estimated by nonparametric generalised additive models. Ecological Modelling, 157(2–3):131–139, November 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001916.

Hellweger:2008:RIG

[Hel08] Ferdi L. Hellweger. The role of inter-generation memory in diel phytoplankton division patterns. Ecological Modelling, 212(3–4):382–396, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005601.

Hendriks:2007:PSM

[Hen07] A. Jan Hendriks. The power of size: a meta-analysis reveals consistency of allometric regressions. Ecological Modelling, 205(1–2):196–208, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000919.

Hering:2003:SFT

[Her03] Klaus Guenther Hering. A scientific formulation of tropical forest management. Ecological Modelling, 166 (3):211–238, August 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003001194.

Herendeen:2004:BTE

[Her04a] Robert A. Herendeen. Bottom-up and top-down effects in food chains depend on functional dependence: an explicit framework. Ecological Modelling, 171(1–2):21–33, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002734.

Herendeen:2004:DTC

[Her04b] Robert A. Herendeen. Dynamic trophic cascade. Ecological Modelling, 177(1–2):129–142, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001127.

Herendeen:2004:EAE

[Her04c] Robert A. Herendeen. Energy analysis and EMERGY analysis — a comparison. Ecological Modelling, 178(1–2):227–237, October 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005295.

Hernandez:2008:SDV

[Her08] Maria-Josefina Hernandez. Spatiotemporal dynamics in variable population interactions with density-dependent interaction coefficients. Ecological Modelling, 214(1):3–16, June 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008000203.

Heymans:2001:EEM

[Hey01a] Johanna Jacomina Heymans. ECOTOX: Ecological Modelling and Ecotoxicology: Edited by L. A. Jørgensen, S. E. Jørgensen, S. N. Nielsen, Elsevier Science, Amsterdam, 2000. CD-ROM. ISBN 0-444-50240-8; NLG 1,200 or USD 609. Ecological Modelling, 144(1):97–98, October 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003490.

Heymans:2001:HEP

[Hey01b] Johanna Jacomina Heymans. Handbook of Ecological Parameters and Ecotoxicology, Edited by L. A. Jørgensen, S. E. Jørgensen and S. N. Nielsen, Elsevier, Amsterdam, 2000, First Electronic Edition. Ecological Modelling, 137(1):107–108, February 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004415.

Halnes:2007:MNM

[HFL07] Geir Halnes, Brian D. Fath, and Hans Liljenström. The modified niche model: Including detritus in simple structural food web models. Ecological Modelling, 208(1):9–16, October 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007002475.

Holmberg:2006:AAN

[HFSH06] Maria Holmberg, Martin Forsius, Michael Starr, and Markus Huttunen. An application of artificial neural networks to carbon, nitrogen and phosphorus concentrations in three boreal streams and impacts of climate change. Ecological Modelling, 195(1–2):51–60, May 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005764.

Hardegree:2003:HGR

[HFV03] S. P. Hardegree, G. N. Flerchinger, and S. S. Van Vactor. Hydrothermal germination response and the development of probabilistic germination profiles. Ecological Modelling, 167(3):305–322, September 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001923.

Hutchings:2001:DMH

[HG01] N. J. Hutchings and I. J. Gordon. A dynamic model of herbivore–plant interactions on grasslands. Ecological Modelling, 136(2–3):209–222, January 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004269.

Hirzel:2002:WOS

[HG02] Alexandre Hirzel and Antoine Guisan. Which is the optimal sampling strategy for habitat suitability modelling. Ecological Modelling, 157(2–3):331–341, November 30, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000200203X.

Harmand:2007:DDK

[HG07] J. Harmand and J. J. Godon. Density-dependent kinetics models for a simple description of complex phenomena in macroscopic mass-balance modeling of bioreactors. Ecological Modelling, 200(3–4):393–402, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003838.

Haakanson:2004:DCM

[HGB04] Lars Håkanson, Andreas Gyllenhammar, and Angelica Brolin. A dynamic compartment model to predict sedimentation and suspended particulate matter in coastal areas. Ecological Modelling, 175(4):353–384, August 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000300509X.

Halbert:2005:GDC

[HGD05] Natalie D. Halbert, William E. Grant, and James N. Derr. Genetic and demographic consequences of importing animals into a small population: a simulation model of the Texas State Bison Herd (USA). Ecological Modelling, 181(2–3):263–276, January 20, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003710.

Harveson:2006:RDF

[HGL^{+}06] Patricia M. Harveson, William E. Grant, Roel R. Lopez, Nova J. Silvy, and Philip A. Frank. The role of dispersal in Florida key deer metapopulation dynamics. Ecological Modelling, 195(3–4):393–401, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006071.

Hernandez:2003:GMG

[HGLLV03] Juan M. Hernández, Eucario Gasca-Leyva, Carmelo J. León, and J. M. Vergara. A growth model for gilthead seabream (Sparus aurata). Ecological Modelling, 165(2–3):265–283, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003000954.

Horne:2008:SMA

[HGR08] Jon S. Horne, Edward O. Garton, and Janet L. Rachlow. A synoptic model of animal space use: Simultaneous estimation of home range, habitat selection, and inter/intra-specific relationships. Ecological Modelling, 214 (2–4):338–348, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008001233.

Hargrove:2000:SFP

[HGT^{+}00] W. W. Hargrove, R. H. Gardner, M. G. Turner, W. H. Romme, and D. G. Despain. Simulating fire patterns in heterogeneous landscapes. Ecological Modelling, 135 (2–3):243–263, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003689.

Herendeen:2004:GDM

[HH04] Robert A. Herendeen and Walter R. Hill. Growth dilution in multilevel food chains. Ecological Modelling, 178 (3–4):349–356, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000400153X.

Herbst:2008:MHS

[HHB^{+}08] M. Herbst, H. J. Hellebrand, J. Bauer, J. A. Huisman, J. Šimůnek, L. Weihermüller, A. Graf, J. Vanderborght, and H. Vereecken. Multiyear heterotrophic soil respiration: Evaluation of a coupled CO_{2} transport and carbon turnover model. Ecological Modelling, 214(2–4):271–283, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000811.

Heinemann:2002:IPE

[HHC02] Alexandre B. Heinemann, Gerrit Hoogenboom, and Bogdan Chojnicki. The impact of potential errors in rainfall observation on the simulation of crop growth, development and yield. Ecological Modelling, 157(1):1–21, November 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002090.

Herzfeld:2001:CMS

[HHD01] M. Herzfeld, D. P. Hamilton, and G. B. Douglas. Comparison of a mechanistic sediment model and a water column model for hindcasting oxygen decay in benthic chambers. Ecological Modelling, 136(2–3):255–267, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004294.

Hormann:2005:ELU

[HHF05] Georg Hörmann, Andreas Horn, and Nicola Fohrer. The evaluation of land-use options in mesoscale catchments: Prospects and limitations of eco-hydrological models. Ecological Modelling, 187(1):3–14, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000360.

Hayes:2007:PBM

[HHK07] John W. Hayes, Nicholas F. Hughes, and Lon H. Kelly. Process-based modelling of invertebrate drift transport, net energy intake and reach carrying capacity for drift-feeding salmonids. Ecological Modelling, 207(2–4):171–188, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000700258X.

Huang:2009:MMS

[HHKH09] Shaochun Huang, Cornelia Hesse, Valentina Krysanova, and Fred Hattermann. From meso- to macro-scale dynamic water quality modelling for the assessment of land use change scenarios. Ecological Modelling, 220(19):2543–2558, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009004359.

He:2002:SSL

[HHL^{+}02] Hong S. He, Zhanqing Hao, David R. Larsen, Limin Dai, Yuanman Hu, and Yu Chang. A simulation study of landscape scale forest succession in northeastern China. Ecological Modelling, 156(2–3):153–166, November 1, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002001047.

Harris:2008:ESE

[HHL08] Nancy L. Harris, Charles A. S. Hall, and Ariel E. Lugo. Estimates of species- and ecosystem-level respiration of woody stems along an elevational gradient in the Luquillo Mountains, Puerto Rico. Ecological Modelling, 216(3– 4):253–264, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002007.

Hirzel:2001:AHS

[HHM01] A. H. Hirzel, V. Helfer, and F. Metral. Assessing habitatsuitability models with a virtual species. Ecological Modelling, 145(2–3):111–121, November 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003969.

Hilker:2006:PEC

[HHP06] Frank M. Hilker, Martin Hinsch, and Hans Joachim Poethke. Parameterizing, evaluating and comparing metapopulation models with data from individual-based simulations. Ecological Modelling, 199(4):476–485, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002912.

Hinrichsen:2009:PVA

[Hin09] Richard A. Hinrichsen. Population viability analysis for several populations using multivariate state-space models. Ecological Modelling, 220(9–10):1197–1202, May 17, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009001227.

Hongping:2002:SAD

[HJ02] Pei Hongping and Ma Jianyi. Study on the algal dynamic model for West Lake, Hangzhou. Ecological Modelling, 148(1):67–77, February 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004355.

Hu:2006:VCT

[HJZ06] Weiping Hu, Sven E. Jørgensen, and Fabing Zhang. A vertical-compressed three-dimensional ecological model in Lake Taihu, China. Ecological Modelling, 190(3– 4):367–398, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002565.

Haag:2001:PPP

[HK01] Daniel Haag and Martin Kaupenjohann. Parameters, prediction, post-normal science and the precautionary principle — a roadmap for modelling for decision-making. Ecological Modelling, 144(1):45–60, October 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003611.

Hanegraaf:2002:DTT

[HK02] P. P. F. Hanegraaf and B. W. Kooi. The dynamics of a tritrophic food chain with two-component populations from a biochemical perspective. Ecological Modelling, 152(1):47–64, June 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001004823.

Haberlandt:2002:ANL

[HKB02] Uwe Haberlandt, Valentina Krysanova, and András Bárdossy. Assessment of nitrogen leaching from arable land in large river basins: Part II: regionalisation using fuzzy rule based modelling. Ecological Modelling, 150(3):277–294, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001005269.

Hattermann:2006:IWR

[HKHB06] F. F. Hattermann, V. Krysanova, A. Habeck, and A. Bronstert. Integrating wetlands and riparian zones in river basin modelling. Ecological Modelling, 199(4):379–392, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002833.

Huang:2007:EMD

[HKL07] Shu-Li Huang, Wei-Chieh Kao, and Chun-Lin Lee. Energetic mechanisms and development of an urban landscape system. Ecological Modelling, 201(3–4):495–506, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006005308.

Hesse:2008:EHM

[HKPH08] Cornelia Hesse, Valentina Krysanova, Jens Päzolt, and Fred F. Hattermann. Eco-hydrological modelling in a highly regulated lowland catchment to find measures for improving water quality. Ecological Modelling, 218(1–2):135–148, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008003128.

Hui:2003:DCM

[HL03] Cang Hui and Zizhen Li. Dynamical complexity and metapopulation persistence. Ecological Modelling, 164 (2–3):201–209, June 15, 2003. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000255.

Helser:2004:BHM

[HL04] Thomas E. Helser and Han-Lin Lai. A Bayesian hierarchical meta-analysis of fish growth: with an example for North American largemouth bass, Micropterus salmoides. Ecological Modelling, 178(3–4):399–416, November 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004001577.

Hirzel:2006:EAH

[HLH^{+}06] Alexandre H. Hirzel, Gwenaëlle Le Lay, Véronique Helfer, Christophe Randin, and Antoine Guisan. Evaluating the ability of habitat suitability models to predict species presences. Ecological Modelling, 199(2):142–152, November 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002468.

Han:2006:PMS

[HLHZ06] Xiaozhuo Han, Zizhen Li, Cang Hui, and Feng Zhang. Polymorphism maintenance in a spatially structured population: a two-locus genetic model of niche construction. Ecological Modelling, 192(1–2):160–174, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003340.

Hiltunen:2006:IBE

[HLK06] Teppo Hiltunen, Jouni Laakso, and Veijo Kaitala. Interactions between environmental variability and immigration rate control patterns of species diversity. Ecological Modelling, 194(1–3):125–131, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005107.

Hiemstra:2006:OMV

[HLR06] Christopher A. Hiemstra, Glen E. Liston, and William A. Reiners. Observing, modelling, and validating snow redistribution by wind in a Wyoming upper treeline landscape. Ecological Modelling, 197(1–2):35–51, August 10, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006001153.

Hosack:2009:SSS

[HLR09] Geoffrey R. Hosack, Hiram W. Li, and Philippe A. Rossignol. Sensitivity of system stability to model structure. Ecological Modelling, 220(8):1054–1062, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000684.

Holzkamper:2006:OLC

[HLS06] Annelie Holzkämper, Angela Lausch, and Ralf Seppelt. Optimizing landscape configuration to enhance habitat suitability for species with contrasting habitat requirements. Ecological Modelling, 198(3–4):277–292, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002109.

Hui:2004:MDD

[HLxY04] Cang Hui, Zizhen Li, and Dong xia Yue. Metapopulation dynamics and distribution, and environmental heterogeneity induced by niche construction. Ecological Modelling, 177(1–2):107–118, September 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001103.

Hesaaraki:2001:ELC

[HM01] M. Hesaaraki and S. M. Moghadas. Existence of limit cycles for predator–prey systems with a class of functional responses. Ecological Modelling, 142(1–2):1–9, August 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004427.

Haakanson:2003:CVC

[HMBG03] L. Håkanson, J. M. Malmaeus, U. Bodemer, and V. Gerhardt. Coefficients of variation for chlorophyll, green algae, diatoms, cryptophytes and blue-greens in rivers as a basis for predictive modelling and aquatic management. Ecological Modelling, 169(1):179–196, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002692.

Hunter:2000:PUE

[HMF00] C. M. Hunter, H. Moller, and D. Fletcher. Parameter uncertainty and elasticity analyses of a population model: setting research priorities for shearwaters. Ecological Modelling, 134(2–3):299–324, October 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003446.

Hull:2000:TML

[HMFM00] Vincent Hull, Chiara Mocenni, Margherita Falcucci, and Nadia Marchettini. A trophodynamic model for the lagoon of Fogliano (Italy) with ecological dependent modifying parameters. Ecological Modelling, 134(2–3):153–167, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003586.

Hildenbrandt:2006:HDV

[HMG06] Hanno Hildenbrandt, Michael S. Müller, and Volker Grimm. How to detect and visualize extinction thresholds for structured PVA models. Ecological Modelling, 191(3–4):545–550, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000500284X.

Hancock:2005:IBM

[HMGK05] P. A. Hancock, E. J. Milner-Gulland, and M. J. Keeling. An individual based model of bearded pig abundance. Ecological Modelling, 181(2–3):123–137, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003618.

Haakanson:2005:SPMa

[HMPF05] Lars Håkanson, Maya Mikrenska, Krassimir Petrov, and Ian Foster. Suspended particulate matter (SPM) in rivers: empirical data and models. Ecological Modelling, 183(2–3):251–267, April 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400482X.

Ma:2008:SCP

[hMzL08] Zhi hui Ma and Zi zhen Li. Stability condition for the predator-free equilibrium point of predator–prey systems with a class of functional response. Ecological Modelling, 216(3–4):415–416, September 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800224X.

Heegaard:2009:MWP

[HN09] Einar Heegaard and Trygve Nilsen. In the mixed world of predictive biology — contradiction or support. Ecological Modelling, 220(12):1461–1468, June 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900180X.

Hurteau:2009:MIP

[HNF09] Matthew D. Hurteau, Malcolm North, and Theodore Foin. Modeling the influence of precipitation and nitrogen deposition on forest understory fuel connectivity in Sierra Nevada mixed-conifer forest. Ecological Modelling, 220(19):2460–2468, October 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004347.

Hassanali:2008:ZDM

[HNS08] A. Hassanali, L. V. Nedorezov, and A. M. Sadykov. Zooprophylactic diversion of mosquitoes from human to alternative hosts: a static simulation model. Ecological Modelling, 212(1–2):155–161, March 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005066.

Hilbert:2001:UAN

[HO01] David W. Hilbert and Bertram Ostendorf. The utility of artificial neural networks for modelling the distribution of vegetation in past, present and future climates. Ecological Modelling, 146(1–3):311–327, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003234.

He:2009:IBM

[HOK^{+}09] Bin He, Taikan Oki, Shinjiro Kanae, Goro Mouri, Ken Kodama, Daisuke Komori, and Shinta Seto. Integrated biogeochemical modelling of nitrogen load from anthropogenic and natural sources in Japan. Ecological Modelling, 220(18):2325–2334, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003846.

Horsch:2003:MSD

[Hör03] Bianca Hörsch. Modelling the spatial distribution of montane and subalpine forests in the central Alps using digital elevation models. Ecological Modelling, 168(3):267–282, October 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001418.

Hosseini:2006:PFI

[Hos06] Parviez R. Hosseini. Pattern formation and individual-based models: the importance of understanding individual-based movement. Ecological Modelling, 194(4):357–371, April 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005442.

Hoyle:2007:WCW

[Hoy07] Martin Hoyle. When corridors work: Insights from a microecosystem. Ecological Modelling, 202(3–4):441–453, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006005473.

Hovestadt:2006:CEC

[HP06] Thomas Hovestadt and Hans Joachim Poethke. The control of emigration and its consequences for the survival of populations. Ecological Modelling, 190(3–4):443–453, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002590.

Hatten:2009:SMW

[HP09] James R. Hatten and Michael J. Parsley. A spatial model of white sturgeon rearing habitat in the lower Columbia River, USA. Ecological Modelling, 220(24):3638–3646, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001926.

Haskett:2000:ECA

[HPA00] Jonathan D. Haskett, Yakov A. Pachepsky, and Basil Acock. Effect of climate and atmospheric change on soybean water stress: a study of Iowa. Ecological Modelling, 135(2–3):265–277, December 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003690.

Hein:2009:WSD

[HPD09] Silke Hein, Hans-Joachim Poethke, and Silvia Dorn. What stops the ‘diploid male vortex’? — a simulation study for species with single locus complementary sex determination. Ecological Modelling, 220(13–14):1663–1669, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002312.

Hull:2008:MDO

[HPF08] Vincent Hull, Luisa Parrella, and Margherita Falcucci. Modelling dissolved oxygen dynamics in coastal lagoons. Ecological Modelling, 211(3–4):468–480, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005029.

Haakanson:2000:MIW

[HPH00] L. Håkanson, A. Parparov, and K. D. Hambright. Modelling the impact of water level fluctuations on water quality (suspended particulate matter) in Lake Kinneret, Israel. Ecological Modelling, 128(2–3):101–125, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002003.

Hein:2004:PDM

[HPHP04] S. Hein, B. Pfenning, T. Hovestadt, and H.-J. Poethke. Patch density, movement pattern, and realised dispersal distances in a patch-matrix landscape a simulation study. Ecological Modelling, 174(4):411–420, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004654.

Hettrich:2003:MST

[HR03] Anke Hettrich and Stephan Rosenzweig. Multivariate statistics as a tool for model-based prediction of floodplain vegetation and fauna. Ecological Modelling, 169(1):73–87, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003002631.

Hauser:2007:OCA

[HRC^{+}07] C. E. Hauser, M. C. Runge, E. G. Cooch, F. A. Johnson, and W. F. Harvey. Optimal control of Atlantic population Canada geese. Ecological Modelling, 201(1):27–36, February 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000600353X.

Hahn:2005:SML

[HRH^{+}05] B. D. Hahn, F. D. Richardson, M. T. Hoffman, R. Roberts, S. W. Todd, and P. J. Carrick. A simulation model of long-term climate, livestock and vegetation interactions on communal rangelands in the semi-arid Succulent Karoo, Namaqualand, South Africa. Ecological Modelling, 183(2–3):211–230, April 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004752.

Holmberg:2000:SSA

[HRJ^{+}00] Maria Holmberg, Katri Rankinen, Matti Johansson, Martin Forsius, Sirpa Kleemola, Johanna Ahonen, and Sanna Syri. Sensitivity of soil acidification model to deposition and forest growth. Ecological Modelling, 135(2–3):311–325, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003872.

Hoang:2001:PMM

[HRMC01] Huong Hoang, Friedrich Recknagel, Jonathan Marshall, and Satish Choy. Predictive modelling of macroinvertebrate assemblages for stream habitat assessments in Queensland (Australia). Ecological Modelling, 146(1–3):195–206, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003064.

Han:2001:SDB

[HS01] Bo-Ping Han and Milan Straskraba. Size dependence of biomass spectra and abundance spectra: the optimal distributions. Ecological Modelling, 145(2–3):175–187, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003891.

Herb:2003:IGS

[HS03] William R. Herb and Heinz G. Stefan. Integral growth of submersed macrophytes in varying light regimes. Ecological Modelling, 168(1–2):77–100, October 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002060.

Herb:2006:SGS

[HS06] William R. Herb and Heinz G. Stefan. Seasonal growth of submersed macrophytes in lakes: the effects of biomass density and light competition. Ecological Modelling, 193 (3–4):560–574, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004126.

He:2004:SFF

[HSC^{+}04] Hong S. He, Bo Z. Shang, Thomas R. Crow, Eric J. Gustafson, and Stephen R. Shifley. Simulating forest fuel and fire risk dynamics across landscapes — LANDIS fuel module design. Ecological Modelling, 180(1):135–151, December 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004003795.

Heymans:2004:CNB

[HSJ04] Johanna J. Heymans, Lynne J. Shannon, and Astrid Jarre. Changes in the northern Benguela ecosystem over three decades: 1970s, 1980s, and 1990s. Ecological Modelling, 172(2–4):175–195, March 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003685.

Harrison:2008:DRD

[HSMN08] Paul Harrison, Daniel Spring, Michael MacKenzie, and Ralph Mac Nally. Dynamic reserve design with the unionfind algorithm. Ecological Modelling, 215(4):369–376, July 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001828.

Hengl:2009:SPS

[HSRD09] Tomislav Hengl, Henk Sierdsema, Andreja Radović, and Arta Dilo. Spatial prediction of species’ distributions from occurrence-only records: combining point pattern analysis, ENFA and regression-kriging. Ecological Modelling, 220(24):3499–3511, December 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004438.

Holvoet:2007:MMP

[HSV07] Katrijn M. A. Holvoet, Piet Seuntjens, and Peter A. Vanrolleghem. Monitoring and modeling pesticide fate in surface waters at the catchment scale. Ecological Modelling, 209(1):53–64, November 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003584.

Heidmann:2008:CPP

[HTA^{+}08] T. Heidmann, C. Tofteng, P. Abrahamsen, F. Plauborg, S. Hansen, A. Battilani, J. Coutinho, F. Doležal, W. Mazurczyk, J. D. R. Ruiz, J. Takáč, and J. Vacek. Calibration procedure for a potato crop growth model using information from across Europe. Ecological Modelling, 211(1–2):209–223, February 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004553.

Holtkamp:2007:IIT

[HTK07] Remko Holtkamp and Maria A. Tobor-Kapłon. Information indices as a tool for quantifying development of below-ground terrestrial ecosystems. Ecological Modelling, 208(1):41–48, October 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007002505.

Hu:2006:RMD

[HTMO06] Wenbiao Hu, Shilu Tong, Kerrie Mengersen, and Brian Oldenburg. Rainfall, mosquito density and the transmission of Ross River virus: a time-series forecasting model. Ecological Modelling, 196(3–4):505–514, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000937.

Heidmann:2000:MSW

[HTS00] T. Heidmann, A. Thomsen, and K. Schelde. Modelling soil water dynamics in winter wheat using different estimates of canopy development. Ecological Modelling, 129(2–3):229–243, May 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002386.

Harp:2007:EEM

[HTS^{+}07] D. R. Harp, M. M. Reda Taha, J. C. Stormont, E. Farfan, and J. Coonrod. An evaporation estimation model using optimized fuzzy learning from example algorithm with an application to the riparian zone of the Middle Rio Grande in New Mexico, U.S.A. Ecological Modelling, 208(2–4):119–128, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002591.

Huang:2003:HDS

[Hua03] Xi Huang. How do simple energy activities comprise complex behavior of life systems?: a conceptual synthesis and decomposition of the energy structure of life systems. Ecological Modelling, 165(1):79–89, July 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000632.

Huang:2008:ETC

[Hua08] Xi Huang. On the energy trait control phenomena in ecosystems. Ecological Modelling, 211(1–2):36–46, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007004231.

Heymans:2002:NAS

[HUB02] J. J. Heymans, R. E. Ulanowicz, and C. Bondavalli. Network analysis of the South Florida Everglades graminoid marshes and comparison with nearby cypress ecosystems. Ecological Modelling, 149(1–2):5–23, March 15, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001005117.

Hui:2006:CCP

[Hui06] Cang Hui. Carrying capacity, population equilibrium, and environment’s maximal load. Ecological Modelling, 192 (1–2):317–320, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003339.

Hui:2007:NCB

[Hui07] C. Hui. Negative correlation between dynamical complexity and metapopulation persistence: a reply. Ecological Modelling, 200(1–2):271–272, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003152.

Hulburt:2001:NIR

[Hul01] Edward M. Hulburt. Non-interference and reciprocal adaptation. Ecological Modelling, 136(1):1–13, January 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003513.

Hulburt:2002:FPA

[Hul02] Edward M. Hulburt. The four principles of adaptation. Ecological Modelling, 156(1):61–84, October 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001242.

Hulburt:2004:FPA

[Hul04] Edward M. Hulburt. The four principles of adaptation and their set theory foundation. Ecological Modelling, 180(2–3):253–276, December 25, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002789.

Hughes:2006:DGS

[HVB06] John K. Hughes, Paul J. Valdes, and Richard Betts. Dynamics of a global-scale vegetation model. Ecological Modelling, 198(3–4):452–462, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002523.

Hobbelen:2007:UDE

[HvG07] P. H. F. Hobbelen and C. A. M. van Gestel. Using dynamic energy budget modeling to predict the influence of temperature and food density on the effect of Cu on earthworm mediated litter consumption. Ecological Modelling, 202(3–4):373–384, April 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005394.

Hengsdijk:2001:UTC

[HvI01] H. Hengsdijk and M. K. van Ittersum. Uncertainty in technical coefficients for future-oriented land use studies: a case study for N-relationships in cropping systems. Ecological Modelling, 144(1):31–44, October 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100360X.

Hein:2006:ESM

[HvI06] Lars Hein and Ekko van Ierland. Efficient and sustainable management of complex forest ecosystems. Ecological Modelling, 190(3–4):351–366, January 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500270X.

Hillmer:2008:PPT

[HvRIZ08] Ingrid Hillmer, Penelope van Reenen, Jörg Imberger, and Tamar Zohary. Phytoplankton patchiness and their role in the modelled productivity of a large, seasonally stratified lake. Ecological Modelling, 218(1–2):49–59, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003001.

Holsten:2009:ICC

[HVVK09] Anne Holsten, Tobias Vetter, Katrin Vohland, and Valentina Krysanova. Impact of climate change on soil moisture dynamics in brandenburg with a focus on nature conservation areas. Ecological Modelling, 220(17):2076–2087, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002890.

Harwood:2009:ERA

[HXP^{+}09] Thomas D. Harwood, Xiangming Xu, Marco Pautasso, Mike J. Jeger, and Michael W. Shaw. Epidemiological risk assessment using linked network and grid based modelling: Phytophthora ramorum and Phytophthora kernoviae in the UK. Ecological Modelling, 220(23):3353–3361, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009005754.

Hashioka:2007:ECW

[HY07a] Taketo Hashioka and Yasuhiro Yamanaka. Ecosystem change in the western North Pacific associated with global warming using 3D-NEMURO. Ecological Modelling, 202(1–2):95–104, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004595.

Hashioka:2007:SRV

[HY07b] Taketo Hashioka and Yasuhiro Yamanaka. Seasonal and regional variations of phytoplankton groups by top–down and bottom–up controls obtained by a 3D ecosystem model. Ecological Modelling, 202(1–2):68–80, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004571.

Huang:2001:PME

[HZ01] Xi Huang and Yuangang Zu. The LES population model: essentials and relationship to the Lotka–Volterra model. Ecological Modelling, 143(3):215–225, September 15, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001003404.

Hormann:2007:CSS

[HZF07] G. Hörmann, X. Zhang, and N. Fohrer. Comparison of a simple and a spatially distributed hydrologic model for the simulation of a lowland catchment in Northern Germany. Ecological Modelling, 209(1):21–28, November 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007003596.

Hui:2005:CES

[HZHL05] Cang Hui, Feng Zhang, Xiaozhuo Han, and Zizhen Li. Cooperation evolution and self-regulation dynamics in metapopulation: Stage-equilibrium hypothesis. Ecological Modelling, 184(2–4):397–412, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005782.

Ichinose:2008:RMF

[IA08] Genki Ichinose and Takaya Arita. The role of migration and founder effect for the evolution of cooperation in a multilevel selection context. Ecological Modelling, 210(3):221–230, January 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007004024.

Ibanez:2002:PMV

[IC02] Frédéric Ibanez and Alessandra Conversi. Prediction of missing values and detection of ‘exceptional events’ in a chronological planktonic series: a single algorithm. Ecological Modelling, 154(1–2):9–23, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000339.

Ibanez:2005:NSA

[ICGÁ05] J. J. Ibáñez, J. Caniego, and A. García-Álvarez. Nested subset analysis and taxa-range size distributions of pedological assemblages: implications for biodiversity studies. Ecological Modelling, 182(3–4):239–256, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001619.

Ibanez:2005:PAR

[ICSC05] J. J. Ibáñez, J. Caniego, F. San José, and C. Carrera. Pedodiversity-area relationships for islands. Ecological Modelling, 182(3–4):257–269, March 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001620.

Ibendahl:2007:CAN

[IF07] Gregory Ibendahl and Ronald A. Fleming. Controlling aquifer nitrogen levels when fertilizing crops: a study of groundwater contamination and denitrification. Ecological Modelling, 205(3–4):507–514, July 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001433.

Inchausti:2002:UPS

[IG02] Pablo Inchausti and Lev Ginzburg. Using the phase shift for assessing the causation of population cycles. Ecological Modelling, 152(1):89–102, June 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004884.

Ibarra:2003:MFI

[IGP^{+}03] Alonso Aguilar Ibarra, Muriel Gevrey, Young-Seuk Park, Puy Lim, and Sovan Lek. Modelling the factors that influence fish guilds composition using a back-propagation network: Assessment of metrics for indices of biotic integrity. Ecological Modelling, 160(3):281–290, February 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002002594.

Itoh:2004:SEP

[IiTS^{+}04] Yu Itoh, Kei ichi Tainaka, Tomoyuki Sakata, Tomomi Tao, and Nariyuki Nakagiri. Spatial enhancement of population uncertainty near the extinction threshold. Ecological Modelling, 174(1–2):191–201, May 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000079.

Imam:2009:EST

[IKS09] Ekwal Imam, S. P. S. Kushwaha, and Aditya Singh. Evaluation of suitable tiger habitat in Chandoli National Park, India, using spatial modelling of environmental variables. Ecological Modelling, 220(24):3621–3629, December 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009004402.

Ilichev:2008:ACM

[Il’08] Vitaly G. Il’ichev. Analysis of competition models in periodic medium. Ecological Modelling, 216(2):188–196, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001397.

Ito:2007:IVG

[IMK^{+}07] Shin-Ichi Ito, Bernard A. Megrey, Michio J. Kishi, Daiki Mukai, Yutaka Kurita, Yasuhiro Ueno, and Yasuhiro Yamanaka. On the interannual variability of the growth of Pacific saury (Cololabis saira): a simple 3-box model using NEMURO.FISH. Ecological Modelling, 202(1–2):174–183, March 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006004698.

Ibanez:2007:DDO

[IMS07] Javier Ibáñez, Jaime Martínez, and Susanne Schnabel. Desertification due to overgrazing in a dynamic commercial livestock–grass–soil system. Ecological Modelling, 205(3–4):277–288, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007000993.

Ito:2002:SMC

[IO02] Akihiko Ito and Takehisa Oikawa. A simulation model of the carbon cycle in land ecosystems (Sim-CYCLE): a description based on dry-matter production theory and plot-scale validation. Ecological Modelling, 151(2–3):143–176, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004732.

Ito:2008:NNA

[IOIA08] Eriko Ito, Kenji Ono, Yoichi M. Ito, and Makoto Araki. A neural network approach to simple prediction of soil nitrification potential: a case study in Japanese temperate forests. Ecological Modelling, 219(1–2):200–211, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008004122.

Izsak:2000:LBE

[IP00] János Izsák and László Papp. A link between ecological diversity indices and measures of biodiversity. Ecological Modelling, 130(1–3):151–156, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002039.

Tainaka:2003:PEO

[iT03] Kei ichi Tainaka. Perturbation expansion and optimized death rate in a lattice ecosystem. Ecological Modelling, 163(1–2):73–85, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004143.

Ivanovska:2009:MOB

[ITDD09] Aneta Ivanovska, Ljupčo Todorovski, Marko Debeljak, and Sašo Džeroski. Modelling the outcrossing between genetically modified and conventional maize with equation discovery. Ecological Modelling, 220(8):1063–1072, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009000957.

Tainaka:2002:PDB

[iTI02] Kei ichi Tainaka and Yu Itoh. Patch dynamics based on Prisoner’s dilemma game: superiority of golden rule. Ecological Modelling, 150(3):295–307, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004811.

Ito:2005:MCC

[Ito05] Akihiko Ito. Modelling of carbon cycle and fire regime in an east Siberian larch forest. Ecological Modelling, 187(2–3):121–139, September 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500061X.

Ito:2007:ASB

[Ito07] Kazuo Ito. Application of the simplified biogeochemical process (SIMPLB) model to determine the impact of acidic deposition on a Japanese cedar forest soil. Ecological Modelling, 205(1–2):265–269, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000865.

Ivanovska:2008:FCE

[IVC^{+}08] Aneta Ivanovska, Celine Vens, Nathalie Colbach, Marko Debeljak, and Sašo Džeroski. The feasibility of co-existence between conventional and genetically modified crops: Using machine learning to analyse the output of simulation models. Ecological Modelling, 215(1–3):262–271, July 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001075.

Ibanez:2008:ADR

[IVP08a] Javier Ibáñez, Jaime Martínez Valderrama, and Juan Puigdefábregas. Assessing desertification risk using system stability condition analysis. Ecological Modelling, 213(2):180–190, May 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007006254.

Ibanez:2008:AOM

[IVP08b] Javier Ibáñez, Jaime Martínez Valderrama, and Juan Puigdefábregas. Assessing overexploitation in Mediterranean aquifers using system stability condition analysis. Ecological Modelling, 218(3–4):260–266, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003463.

Wu:2000:MAM

[iWLSN00] Hsin i Wu, Bai-Lian Li, Timothy A. Springer, and William H. Neill. Modelling animal movement as a persistent random walk in two dimensions: expected magnitude of net displacement. Ecological Modelling, 132(1–2):115–124, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003094.

Izaurralde:2006:SSC

[IWM^{+}06] R. C. Izaurralde, J. R. Williams, W. B. McGill, N. J. Rosenberg, and M. C. Quiroga Jakas. Simulating soil C dynamics with EPIC: Model description and testing against long-term data. Ecological Modelling, 192(3–4):362–384, February 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003571.

Ikegami:2008:OBA

[IWW08] Makihiko Ikegami, Dennis F. Whigham, and Marinus J. A. Werger. Optimal biomass allocation in heterogeneous environments in a clonal plant — spatial division of labor. Ecological Modelling, 213(2):156–164, May 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006242.

Jelaska:2006:RFH

[JAB^{+}06] Sven D. Jelaska, Oleg Antonić, Mario Božić, Josip Križan, and Vladimir Kušan. Responses of forest herbs to available understory light measured with hemispherical photographs in silver fir–beech forest in Croatia. Ecological Modelling, 194(1–3):209–218, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005181.

Jager:2001:IVL

[Jag01] Henriëtte I. Jager. Individual variation in life history characteristics can influence extinction risk. Ecological Modelling, 144(1):61–76, October 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003623.

Jager:2009:EIV

[Jag09] Henriëtte I. Jager. Erratum to “Individual variation in life history characteristics can influence extinction risk” [Ecol. Modell. 144 (2001) 61–76]. Ecological Modelling, 220(1):81, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008003839.

Janekovic:2006:MBP

[JAK^{+}06] Ivica Janeković, Oleg Antonić, Josip Križan, Dragan Bukovec, and Tatjana Bakran-Petricioli. Modelling basic physical parameters in the Adriatic Sea as the basis for marine benthic habitats mapping. Ecological Modelling, 194(1–3):62–69, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004977.

Janssen:2001:EIM

[Jan01] Marco A. Janssen. An exploratory integrated model to assess management of lake eutrophication. Ecological Modelling, 140(1–2):111–124, May 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002605.

Jelaska:2003:EPS

[JAN^{+}03] Sven D. Jelaska, Oleg Antonić, Toni Nikolić, Vladimir Hršak, Miško Plazibat, and Josip Križan. Estimating plant species occurrence in MTB/64 quadrants as a function of DEM-based variables — a case study for Medvednica Nature Park, Croatia. Ecological Modelling, 170(2–3):333–343, December 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003002370.

Jaeger:2005:PWA

[JBB^{+}05] Jochen A. G. Jaeger, Jeff Bowman, Julie Brennan, Lenore Fahrig, Dan Bert, Julie Bouchard, Neil Charbonneau, Karin Frank, Bernd Gruber, and Katharina Tluk von Toschanowitz. Predicting when animal populations are at risk from roads: an interactive model of road avoidance behavior. Ecological Modelling, 185(2–4):329–348, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000050.

Jolicoeur:2002:RMS

[JBMBP02] Mario Jolicoeur, E. Bouchard-Marchand, G. Bécard, and M. Perrier. Regulation of mycorrhizal symbiosis: development of a structured nutritional dual model. Ecological Modelling, 158(1–2):121–142, December 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001680.

Jolicoeur:2003:RMS

[JBMBP03] Mario Jolicoeur, E. Bouchard-Marchand, G. Bécard, and M. Perrier. Regulation of mycorrhizal symbiosis: development of a structured nutritional dual model. Ecological Modelling, 163(3):247–267, May 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000243.

Jordan:2007:QPI

[JBP07] Ferenc Jordán, Zsófia Benedek, and János Podani. Quantifying positional importance in food webs: a comparison of centrality indices. Ecological Modelling, 205(1–2):270–275, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007001184.

Jonsen:2007:IDS

[JBR07] Ian D. Jonsen, Robert S. Bourchier, and Jens Roland. Influence of dispersal, stochasticity, and an Allee effect on the persistence of weed biocontrol introductions. Ecological Modelling, 203(3–4):521–526, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006284.

Jepsen:2005:EEC

[JBT^{+}05] J. U. Jepsen, J. M. Baveco, C. J. Topping, J. Verboom, and C. C. Vos. Evaluating the effect of corridors and landscape heterogeneity on dispersal probability: a comparison of three spatially explicit modelling approaches. Ecological Modelling, 181(4):445–459, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003254.

Ji:2002:ISS

[JCB^{+}02] Rubao Ji, Changsheng Chen, Judith Wells Budd, David J. Schwab, Dmitry Beletsky, Gary L. Fahnenstiel, Thomas H. Johengen, Henry Vanderploeg, Brian Eadie, James Cotner, Wayne Gardner, and Marie Bundy. Influences of suspended sediments on the ecosystem in Lake Michigan: a 3-D coupled bio-physical modeling experiment. Ecological Modelling, 152(2–3):169–190, July 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000273.

Jager:2006:SEH

[JCE06] Henriette I. Jager, Eric A. Carr, and Rebecca A. Efroymson. Simulated effects of habitat loss and fragmentation on a solitary mustelid predator. Ecological Modelling, 191(3–4):416–430, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002826.

Jordan:2009:TFO

[JcLM09] Ferenc Jordán, Wei chung Liu, and Ágnes Mike. Trophic field overlap: a new approach to quantify keystone species. Ecological Modelling, 220(21):2899–2907, November 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008005838.

Jiang:2008:TCJ

[JCX^{+}08] Hong Jiang, He-Qin Cheng, Hai-Gen Xu, Francisco Arreguín-Sánchez, Manuel J. Zetina-Rejón, Pablo Del Monte Luna, and William J. F. Le Quesne. Trophic controls of jellyfish blooms and links with fisheries in the East China Sea. Ecological Modelling, 212(3–4):492–503, April 10, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007005996.

Jerina:2003:MBB

[JDD^{+}03] Klemen Jerina, Marko Debeljak, Sašo Džeroski, Andrej Kobler, and Miha Adamič. Modeling the brown bear population in Slovenia: a tool in the conservation management of a threatened species. Ecological Modelling, 170(2–3):453–469, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000300245X.

James:2007:DMH

[JDPI07] Alex James, Roger Dungan, Michael Plank, and Ryoko Ito. A dynamical model of honeydew droplet production by sooty-beech scale insects (Ultracoelostoma spp.) in New Zealand Nothofagus forest. Ecological Modelling, 209(2–4):323–332, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003456.

Jeraj:2006:AML

[JDTD06] Marjeta Jeraj, Sašo Džeroski, Ljupčo Todorovski, and Marko Debeljak. Application of machine learning methods to palaeoecological data. Ecological Modelling, 191(1):159–169, January 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003832.

Jensen:2000:SAS

[Jen00] A. L. Jensen. Sex and age structured matrix model applied to harvesting a white tailed deer population. Ecological Modelling, 128(2–3):245–249, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000001988.

Jensen:2001:MEC

[Jen01] A. L. Jensen. Modelling the effect of cormorant predation on stocked walleye. Ecological Modelling, 145(2– 3):123–127, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003970.

Jensen:2002:AHE

[Jen02] A. L. Jensen. Analysis of harvest and effort data for wild populations in fluctuating environments. Ecological Modelling, 157(1):43–49, November 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001758.

Jensen:2005:HFE

[Jen05] A. L. Jensen. Harvest in a fluctuating environment and conservative harvest for the Fox surplus production model. Ecological Modelling, 182(1):1–9, February 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003326.

John:2000:MPT

[JF00] Eurgain H. John and Kevin J. Flynn. Modelling phosphate transport and assimilation in microalgae; how much complexity is warranted? Ecological Modelling, 125 (2–3):145–157, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380099001787.

Jorgensen:2004:MSA

[JF04] Sven E. Jørgensen and Brian D. Fath. Modelling the selective adaptation of Darwin’s finches. Ecological Modelling, 176(3–4):409–418, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001024.

Jorgensen:2006:EEN

[JF06] S. E. Jørgensen and B. D. Fath. Examination of ecological networks. Ecological Modelling, 196(3–4):283–288, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000998.

Jorgensen:2008:NIT

[JFG^{+}08] Sven Erik Jørgensen, Brian D. Fath, William E. Grant, Tarzan Legovic, and Søren Nors Nielsen. New initiative for thematic issues: an invitation. Ecological Modelling, 215(4):273–275, July 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002226.

Jorgensen:2006:EPE

[JFGN06] Sven Erik Jørgensen, Brian D. Fath, William Grant, and Søren Nors Nielsen. The editorial policy of ecological modelling. Ecological Modelling, 199(1):1–3, November 1, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002997.

Jensen:2005:PTF

[JG05] Christopher X. J. Jensen and Lev R. Ginzburg. Paradoxes or theoretical failures? The jury is still out. Ecological Modelling, 188(1):3–14, October 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002334.

Johnson:2008:SSD

[JG08] Chris J. Johnson and Michael P. Gillingham. Sensitivity of species-distribution models to error, bias, and model design: an application to resource selection functions for woodland caribou. Ecological Modelling, 213(2):143–155, May 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007006217.

Jusup:2007:IAM

[JGL07] Marko Jusup, Sunčana Geček, and Tarzan Legović. Impact of aquacultures on the marine ecosystem: Modelling benthic carbon loading over variable depth. Ecological Modelling, 200(3–4):459–466, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003899.

Ji:2000:SMG

[JJ00] Wei Ji and Clinton Jeske. Spatial modeling of the geographic distribution of wildlife populations: a case study in the lower Mississippi River region. Ecological Modelling, 132(1–2):95–104, July 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003070.

Jeong:2001:PEP

[JJK^{+}01] Kwang-Seuk Jeong, Gea-Jae Joo, Hyun-Woo Kim, Kyong Ha, and Friedrich Recknagel. Prediction and elucidation of phytoplankton dynamics in the Nakdong River (Korea) by means of a recurrent artificial neural network. Ecological Modelling, 146(1–3):115–129, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003003.

Jian:2007:DDM

[JJWF07] Fuji Jian, Digvir S. Jayas, Noel D. G. White, and Paul G. Fields. A distributed-delay model to predict ageing and survival rates of adults of Cryptolestes ferrugineus (Stephens) (Coleoptera: Laemophloeidae) in granaries filled with wheat. Ecological Modelling, 200(3–4):412–420, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003875.

Jian:2008:NAP

[JJWS08] Fuji Jian, Digvir S. Jayas, Noel D. G. White, and E. A. Smith. Numerical analysis and parameter estimation technique for insect population redistribution models. Ecological Modelling, 211(1–2):47–56, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004243.

Jonsson:2006:FWS

[JKJ06] Tomas Jonsson, Patrik Karlsson, and Annie Jonsson. Food web structure affects the extinction risk of species in ecological communities. Ecological Modelling, 199 (1):93–106, November 1, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006002948.

Jeong:2008:NLA

[JKJ^{+}08] Kwang-Seuk Jeong, Dong-Kyun Kim, Jong-Mun Jung, Myoung-Chul Kim, and Gea-Jae Joo. Non-linear autoregressive modelling by temporal recurrent neural networks for the prediction of freshwater phytoplankton dynamics. Ecological Modelling, 211(3–4):292–300, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004516.

Jusup:2009:PAD

[JKPL09] Marko Jusup, Jasminka Klanjšček, Donat Petricioli, and Tarzan Legović. Predicting aquaculture-derived benthic organic enrichment: Model validation. Ecological Modelling, 220(19):2407–2414, October 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003962.

Jager:2005:SUA

[JKS^{+}05] Henriette I. Jager, Anthony W. King, Nathan H. Schumaker, Tom L. Ashwood, and Barbara L. Jackson. Spatial uncertainty analysis of population models. Ecological Modelling, 185(1):13–27, June 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005915.

Jeong:2003:MMA

[JKWJ03] Kwang-Seuk Jeong, Dong-Kyun Kim, Peter Whigham, and Gea-Jae Joo. Modelling Microcystis aeruginosa bloom dynamics in the Nakdong River by means of evolutionary computation and statistical approach. Ecological Modelling, 161(1–2):67–78, March 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002806.

Jenerette:2007:MCS

[JL07] G. Darrel Jenerette and Rattan Lal. Modeled carbon sequestration variation in a linked erosion–deposition system. Ecological Modelling, 200(1–2):207–216, January 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006003632.

Jiao:2009:HDA

[JLAM09] Yan Jiao, Nicolas W. R. Lapointe, Paul L. Angermeier, and Brian R. Murphy. Hierarchical demographic approaches for assessing invasion dynamics of non-indigenous species: an example using northern snakehead (Channa argus). Ecological Modelling, 220(13–14):1681–1689, July 17, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009002592.

Johnson:2009:USN

[JLBS09] Jeffrey C. Johnson, Joseph J. Luczkovich, Stephen P. Borgatti, and Tom A. B. Snijders. Using social network analysis tools in ecology: Markov process transition models applied to the seasonal trophic network dynamics of the Chesapeake Bay. Ecological Modelling, 220(22):3133–3140, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004414.

Jorgensen:2005:CEO

[JLDM05] Sven Erik Jørgensen, Niels Ladegaard, Marko Debeljak, and Joao Carlos Marques. Calculations of exergy for organisms. Ecological Modelling, 185(2–4):165–175, July 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004006015.

Jorgensen:2008:RPE

[JLF08] Sven E. Jørgensen, Tarzan Legović, and Brian D. Fath. Recent progress in ecological modeling activity. Ecological Modelling, 212(3–4):179, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000070.

Johansson:2001:MPD

[JLH01] Håkan Johansson, Martin Lindström, and Lars Håkanson. On the modelling of the particulate and dissolved fractions of substances in aquatic ecosystems — sedimentological and ecological interactions. Ecological Modelling, 137(2–3):225–240, February 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004397.

Jensen:2001:ASM

[JM01] A. L. Jensen and D. H. Miller. Age structured matrix predation model for the dynamics of wolf and deer populations. Ecological Modelling, 141(1–3):299–305, July 1, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001002307.

Jensen:2004:MEW

[JM04] A. L. Jensen and D. H. Miller. Modeling emigration of wolves from a wilderness area into adjacent agricultural regions. Ecological Modelling, 175(2):115–120, July 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003003971.

Jorgensen:2002:SCE

[JMN02] Sven Erik Jørgensen, Joao Marques, and Søren Nors Nielsen. Structural changes in an estuary, described by models and using exergy as orientor. Ecological Modelling, 158(3):233–240, December 31, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200234X.

Jimenez-Montealegre:2002:CVD

[JMVvDV02] R. Jiménez-Montealegre, M. C. J. Verdegem, A. van Dam, and J. A. J. Verreth. Conceptualization and validation of a dynamic model for the simulation of nitrogen transformations and fluxes in fish ponds. Ecological Modelling, 147(2):123–152, January 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004033.

Johnson:2009:EAP

[JNA^{+}09] Galen A. Johnson, Nathalie Niquil, Harald Asmus, Cédric Bacher, Ragnhild Asmus, and Daniel Baird. The effects of aggregation on the performance of the inverse method and indicators of network analysis. Ecological Modelling, 220(23):3448–3464, December 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005341.

Jonckheere:2006:FDB

[JNM^{+}06] I. Jonckheere, K. Nackaerts, B. Muys, J. van Aardt, and P. Coppin. A fractal dimension-based modelling approach for studying the effect of leaf distribution on LAI retrieval in forest canopies. Ecological Modelling, 197(1–2):179–195, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006001001.

Jonzen:2002:SHM

[JNSS02] Niclas Jonzén, Bart A. Nolet, Luis Santamaría, and Mats G. E. Svensson. Seasonal herbivory and mortality compensation in a swan–pondweed system. Ecological Modelling, 147(3):209–219, January 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004264.

Jordan:2009:SFW

[JO09] Ferenc Jordán and Györgyi Osváth. The sensitivity of food web topology to temporal data aggregation. Ecological Modelling, 220(22):3141–3146, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003238.

Jorgensen:2004:EES

[JOB04] S. E. Jorgensen, H. T. Odum, and M. T. Brown. Emergy and exergy stored in genetic information. Ecological Modelling, 178(1–2):11–16, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005131.

Jordan:2008:IIS

[JOBL08] Ferenc Jordán, Thomas A. Okey, Barbara Bauer, and Simone Libralato. Identifying important species: Linking structure and function in ecological networks. Ecological Modelling, 216(1):75–80, August 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001981.

Jongschaap:2007:SCG

[Jon07] Raymond E. E. Jongschaap. Sensitivity of a crop growth simulation model to variation in LAI and canopy nitrogen used for run-time calibration. Ecological Modelling, 200(1–2):89–98, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003450.

Jordan:2000:RTA

[Jor00a] Ferenc Jordán. A reliability-theory approach to corridor design. Ecological Modelling, 128(2–3):211–220, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000001976.

Jorgensen:2000:YEM

[Jør00b] Sven Erik Jørgensen. 25 years of ecological modelling by ecological modelling. Ecological Modelling, 126 (2–3):95–99, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003203.

Jorgensen:2001:BR

[Jor01a] S. Jorgensen. Book reviews. Ecological Modelling, 136 (2–3):303–306, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003884.

Jorgensen:2001:MS

[Jør01b] S. E. Jørgensen. Milan Straskraba. Ecological Modelling, 140(3):193–194, June 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003398.

Jorgensen:2001:PEC

[Jor01c] Sven Erik Jorgensen. Parameter estimation and calibration by use of exergy. Ecological Modelling, 146 (1–3):299–302, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003155.

Jorgensen:2002:BR

[Jør02a] S. E. Jørgensen. Book reviews. Ecological Modelling, 156(1):85–91, October 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002002065.

Jorgensen:2002:EER

[Jør02b] Sven Erik Jørgensen. Explanation of ecological rules and observation by application of ecosystem theory and ecological models. Ecological Modelling, 158(3):241–248, December 31, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002002363.

Jorgensen:2004:HIE

[Jør04a] S. E. Jørgensen. A Handbook of Industrial Ecology: Robert U. Ayres, Leslie W. Ayres (Eds.), Edward Elgar, Cheltenham, UK, 690 pp., hardbound, 175 British pounds. Ecological Modelling, 172(1):97–98, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300454X.

Jorgensen:2004:DMM

[Jør04b] S. E. Jørgensen. Dynamic Modeling for Marine Conservation: Matthias Ruth, James Lingholm (Eds.), Springer-Verlag, Heidelberg, 450 pages, hardbound, 212 illustrations. Ecological Modelling, 172(1):97, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004538.

Jorgensen:2004:EEE

[Jør04c] S. E. Jørgensen. Economics–Energy–Environment Simulation: Beyond the Kyoto, Number 20 in the Series Economy and Environment, Kimio Uno (Ed.), Kluwer Academic Publishers, Dordrecht, 2002, 340 pages, hardbound 125 euro, (US$ 115) or 79 British Pounds. Ecological Modelling, 172(1):96, February 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004514.

Jorgensen:2004:ESE

[Jør04d] S. E. Jørgensen. Economics of Sustainable Energy in Agriculture: Number 24 in the Series Economy and Environment, Ekko C. van Ierland, Alfred Oude Lansink (Eds.), Kluwer Academic Publishers, Dordrecht, 2002, 250 pages, hardbound 84 euro, (US $73) or 51 British Pounds. Ecological Modelling, 172(1):95–96, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004502.

Jorgensen:2004:MSM

[Jør04e] S. E. Jørgensen. Model Selection and Multimodel Inference: a Practical Information — Theoretic Approach, Second Edition, Kenneth P. Brunham, David R. Anderson, Springer-Verlag, Heidelberg, 2002, 490 pages, hardbound, 31 illustrations. Ecological Modelling, 172(1):96–97, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003004526.

Jorgensen:2004:SLD

[Jør04f] Sven Erik Jørgensen. Sustainability, Local Democracy and the Future: the Swedish Model: Uno Svedin, Britt Hägerhäll Aniansson (Eds.), Kluwer Academic Publishers, 2002, 232 pages. Ecological Modelling, 172(1):98–99, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003004563.

Jorgensen:2004:ETP

[Jør04g] Sven Erik Jørgensen. The End of Time (199 pages), The Essence of Time (125 pages): Enzo Tiezzi, WIT Press, Southhampton, UK, fall 2002. Ecological Modelling, 172(1):95, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004496.

Jorgensen:2004:TCE

[Jør04h] Sven Erik Jørgensen. The Technological Change and the Environment: Resources for the Future. Washington, Arnulf Grübler, Nebojsa Nakicenovic and William D. Nordhaus (Eds.), DC, USA, International Institute of Applied System Analysis, Laxenburg, Austria, 2002, 405 pages, Hardbound. Ecological Modelling, 172(1):98, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004551.

Jorgensen:2005:ASS

[Jor05a] S. E. Jorgensen. Andrea Saltelli, Stefano Tarantola, Francesca Campololongo, Marco Ratto, Sensitivity Analysis in Practice. A Guide to Assessing Scientific Models, John Wiley and Sons, Ltd., England, 39.95 Brit. Pounds, Hard Bound, 220 pp., ISBN 0-470-87093-1. Ecological Modelling, 181(1):92, January 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400451X.

Jorgensen:2005:BRE

[Jor05b] S. E. Jorgensen. Book review: Environmental Modelling — Finding Simplicity in Complexity, John Wainwright, Mark Mulligan (Eds.) (2004), 408 pp., Paper back, £29.95, ISBN 0-471-49618-9. Ecological Modelling, 181(1):88–89, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004004442.

Jorgensen:2005:BRR

[Jor05c] S. E. Jorgensen. Book review: R. B. Matthews, W. Stephens (Eds.), Crop-Soil Simulation Models, Application in Developing Countries, CABI Publishing, Hard bound, £55.00, 278 pp., ISBN 0-85199-563-2. Ecological Modelling, 181(1):91–92, January 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004508.

Jorgensen:2005:EMP

[Jor05d] S. E. Jorgensen. Environmental Modelling and Prediction, G. Peng, L. M. Leslie, Y. Shao (Eds.). Springer Verlag, Heidelberg (2002), hard bound, 480 pp. Ecological Modelling, 181(1):87–88, January 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004429.

Jorgensen:2005:FES

[Jor05e] S. E. Jorgensen. First Ecology, second ed.; Beeby, Alan; Brennan, Anne-Marie. Oxford University Press (2003), 318 pp. + 36 pp. with colour photos, Softbound, ISBN 0-19-926124-5. Ecological Modelling, 181(2–3):277–283, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003989.

Jorgensen:2005:GEC

[Jor05f] S. E. Jorgensen. Global Environmental Change — Modelling and Monitoring. Springer-Verlag, Berlin, Heidelberg, New York (2002), Hard bound, 316 pp., ISBN 3-540-43373-2. Ecological Modelling, 181(1):89–90, January 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004004466.

Jorgensen:2005:GMU

[Jor05g] S. E. Jorgensen. Groundwater Modelling, Using Geographical Information Systems, Georg F. Pinder, L. M. Leslie, Y. Shao (Eds.). John Wiley & Sons, Inc., New York (2002), includes CDE-ROM with model solutions, hard bound, 233 pp., £59.5, ISBN 0-471-08498-0. Ecological Modelling, 181(1):88, January 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004430.

Jorgensen:2005:IMD

[Jor05h] S. E. Jorgensen. Introduction to Multi-Disciplinary Model-Building, H. G. Natke. WIT-Press, Southampton (2003), 380 pp., hardbound, £119, ISBN 1-85312-954-2. Ecological Modelling, 181(1):90, January 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004478.

Jorgensen:2005:LSM

[Jor05i] S. E. Jorgensen. Landscape Simulation Modeling, A. Voinov, R. Costanza (Eds.). Springer, Heidelberg (2004), hardbound, 330 pp., includes a CD-ROM with Stella models, US $69.95, ISBN 0-387-00835-7. Ecological Modelling, 181(1):87, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004417.

Jorgensen:2005:MNS

[Jor05j] S. E. Jorgensen. Modelling in Natural Sciences, Design, Validation and Case Studies, T. Müller, H. Müller (Eds.). Springer Verlag, Berlin, Heidelberg, New York (2003), 460 pp., Hard bound, ISBN 3-540-00153-0. Ecological Modelling, 181(1):91, January 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004491.

Jorgensen:2005:RRM

[Jor05k] S. E. Jorgensen. Rainfall-Runoff Modelling, Keith J. Beven. John Wiley & Sons, UK (2003), 360 pp., Paper back, £27.50, ISBN: 0-470-86671-3. Ecological Modelling, 181(1):89, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004004454.

Jorgensen:2005:SEE

[Jor05l] S. E. Jorgensen. Spatial Ecological–Economic Analysis For Wetland Management, Jeroen C. J. M. van der Bergh, Aat Barendregt, J. Alison Gilbert (Eds.). Cambridge University Press (2004), 240 pp., hardbound, ISBN 0-521-82230-0. Ecological Modelling, 181(1):90–91, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400448X.

Jorgensen:2005:EME

[Jør05m] Sven Erik Jørgensen. Ecological modelling: editorial overview 2000–2005. Ecological Modelling, 188(2– 4):137–144, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003601.

Jorgensen:2006:EDC

[Jør06a] S. E. Jørgensen. Ecological Dynamics and Cybernetic Principle, P. Tiefan, J. Changjie, G. Dexin (Eds.). Science Press, Science Press USA Inc. (2004), (400 pages, Hardbound).

Ecological Modelling, 196(1–2):272, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001323.

Jorgensen:2006:ESM

[Jor06b] Sven Erik Jorgensen. Exactly Solvable Models of Biological Invasion, Sergei V. Petrovskii, Bai-Lian Li (Eds.). Chapman & Hall/CRC (2006), (216 pp., Hardbound, £49.99). Ecological Modelling, 196(1–2):271, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001335.

Jorgensen:2007:THV

[Jør07a] S. E. Jørgensen. Two hundred volumes of ecological modelling. Ecological Modelling, 200(3–4):277–278, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004911.

Jorgensen:2007:DAE

[Jør07b] Sven Erik Jørgensen. Description of aquatic ecosystem’s development by eco-exergy and exergy destruction. Ecological Modelling, 204(1–2):22–28, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006387.

Jorgensen:2007:ENT

[Jør07c] Sven Erik Jørgensen. Ecological network theory. Ecological Modelling, 208(1):1–2, October 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002451.

Jorgensen:2007:EE

[Jør07d] Sven Erik Jørgensen. Evolution and exergy. Ecological Modelling, 203(3–4):490–494, May 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006399.

Jorgensen:2008:IGD

[Jør08a] Sven Erik Jørgensen. Info-gap, decision theory. Ecological Modelling, 211(1–2):249, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005625.

Jorgensen:2008:OYS

[Jør08b] Sven Erik Jørgensen. Obituary for Yuri Svirezhev. Ecological Modelling, 216(2):81–88, August 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001269.

Jorgensen:2008:OMT

[Jør08c] Sven Erik Jørgensen. Overview of the model types available for development of ecological models. Ecological Modelling, 215(1–3):3–9, July 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001105.

Jorgensen:2009:BHI

[Jør09a] S. E. Jørgensen. A brief history of the international journal of ecological modelling. Ecological Modelling, 220(1):4–6, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008004596.

Jorgensen:2009:EMC

[Jør09b] Sven Erik Jørgensen. Ecological Modeling: A Common-Sense Approach to Theory and Practice, William E. Grant, Todd M. Swannack. Blackwell Publishing (2008), 155 pp., Softbound, Price: US $17.99. Ecological Modelling, 220(3):436, February 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005085.

James:2003:RBP

[JPB03] Alex James, Jonathan W. Pitchford, and John Brindley. The relationship between plankton blooms, the hatching of fish larvae, and recruitment. Ecological Modelling, 160(1–2):77–90, February 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003113.

Jurc:2006:SBB

[JPD^{+}06] Maja Jurc, Marko Perko, Sašo Džeroski, Damjan Demšar, and Boris Hrašovec. Spruce bark beetles (Ips typographus, Pityogenes chalcographus, Col.: Scolytidae) in the Dinaric mountain forests of Slovenia: Monitoring and modeling. Ecological Modelling, 194(1–3):219–226, March 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005005193.

Jorgensen:2000:EEG

[JPS00] Sven E. Jørgensen, Bernard C. Patten, and Milan Straškraba. Ecosystems emerging: 4. Growth. Ecological Modelling, 126(2–3):249–284, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002684.

Jopp:2005:DCB

[JR05] Fred Jopp and Hauke Reuter. Dispersal of carabid beetles — emergence of distribution patterns. Ecological Modelling, 186(4):389–405, September 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500133X.

Jorgensen:2002:ICE

[JRBS02] S. E. Jørgensen, S. Ray, L. Berec, and M. Straskraba. Improved calibration of a eutrophication model by use of the size variation due to succession. Ecological Modelling, 153(3):269–277, August 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000297.

Justic:2002:MID

[JRT02] Dubravko Justić, Nancy N. Rabalais, and R. Eugene Turner. Modeling the impacts of decadal changes in riverine nutrient fluxes on coastal eutrophication near the Mississippi River Delta. Ecological Modelling, 152(1):33–46, June 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004720.

Jordan:2003:PFR

[JSM03] Ferenc Jordán, István Scheuring, and István Molnár. Persistence and flow reliability in simple food webs. Ecological Modelling, 161(1–2):117–124, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200296X.

Jager:2001:DMR

[JT01] H. I. Jager and J. A. Tyler. Discussion on Movement rules for individual-based models of stream fish: [S. F. Railsback et al. 123 (1999) 73–89]. Ecological Modelling, 143(3):245–248, September 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003325.

Jorgensen:2009:PWE

[JT09] Sven Erik Jørgensen and Enzo Tiezzi. Preface to workshop on “Emergence of Novelties”, 9–16 October 2008, Pacina, Siena, Italy. Ecological Modelling, 220(16):1855–1856, August 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000900266X.

Joehnk:2001:MMO

[JU01] K. D. Joehnk and L. Umlauf. Modelling the metalimnetic oxygen minimum in a medium sized alpine lake. Ecological Modelling, 136(1):67–80, January 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003811.

Jorgensen:2009:NCA

[JU09] Sven Erik Jørgensen and Robert Ulanowicz. Network calculations and ascendency based on eco-exergy. Ecological Modelling, 220(16):1893–1896, August 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002701.

Juanico:2009:HTA

[Jua09] Dranreb Earl O. Juanico. Herding tendency as an aggregating factor in a binary mixture of social entities. Ecological Modelling, 220(24):3521–3526, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004372.

Jorgensen:2002:EOS

[JVL02] S. E. Jørgensen, Piet Verdonschot, and Sovan Lek. Explanation of the observed structure of functional feeding groups of aquatic macro-invertebrates by an ecological model and the maximum exergy principle. Ecological Modelling, 158(3):223–231, December 31, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002338.

Jiang:2009:MEA

[JW09] Jiexian Jiang and Nianfeng Wan. A model for ecological assessment to pesticide pollution management. Ecological Modelling, 220(15):1844–1851, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002919.

Janssen:2000:AAM

[JWLA00] Marco A. Janssen, Brian H. Walker, Jenny Langridge, and Nick Abel. An adaptive agent model for analysing coevolution of management and policies in a complex rangeland system. Ecological Modelling, 131(2–3):249–268, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002568.

Jackson:2007:CSM

[JWW^{+}07] B. M. Jackson, H. S. Wheater, A. J. Wade, D. Butterfield, S. A. Mathias, A. M. Ireson, A. P. Butler, N. R. McIntyre, and P. G. Whitehead. Catchment-scale modelling of flow and nutrient transport in the chalk unsaturated zone. Ecological Modelling, 209(1):41–52, November 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003572.

Johnson:2007:EHS

[JZC^{+}07] Scott N. Johnson, Xiaoxian Zhang, John W. Crawford, Peter J. Gregory, and Iain M. Young. Egg hatching and survival time of soil-dwelling insect larvae: a partial differential equation model and experimental validation. Ecological Modelling, 202(3–4):493–502, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005515.

Jia:2007:MCS

[JZY07] Bingrui Jia, Guangsheng Zhou, and Wenping Yuan. Modeling and coupling of soil respiration and soil water content in fenced Leymus chinensis steppe, inner Mongolia. Ecological Modelling, 201(2):157–162, February 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004170.

Kobler:2000:IBB

[KA00] Andrej Kobler and Miha Adamic. Identifying brown bear habitat by a combined GIS and machine learning method. Ecological Modelling, 135(2–3):291–300, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003847.

Katterer:2001:IFA

[KA01] Thomas Kätterer and Olof Andrén. The ICBM family of analytically solved models of soil carbon, nitrogen and microbial biomass dynamics — descriptions and application examples. Ecological Modelling, 136(2–3):191–207, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004208.

Kaitaniemi:2000:CMT

[Kai00] Pekka Kaitaniemi. A canonical model of tree resource allocation after defoliation and bud consumption. Ecological Modelling, 129(2–3):259–272, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002362.

Kangas:2004:EEB

[Kan04a] Patrick Kangas. Ecological economics began on the Texas bays during the 1950s. Ecological Modelling, 178(1–2):179–181, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003005301.

Kangas:2004:RPE

[Kan04b] Patrick Kangas. The role of passive electrical analogs in H. T. Odum’s systems thinking. Ecological Modelling, 178(1–2):101–106, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005313.

Kumazawa:2009:PDM

[KAN^{+}09] Fuga Kumazawa, Takahiro Asami, Nariyuki Nakagiri, Kei ichi Tainaka, Tatsuya Togashi, Tatsuo Miyazaki, and Jin Yoshimura. Population dynamics of Müllerian mimicry under interspecific competition. Ecological Modelling, 220 (3):424–429, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008005413.

Karev:2003:IMT

[Kar03] G. P. Karev. Inhomogeneous models of tree stand selfthinning. Ecological Modelling, 160(1–2):23–37, February 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002879.

Kar:2004:IEN

[Kar04] Tapan Kumar Kar. Influence of environmental noises on the Gompertz model of two species fishery. Ecological Modelling, 173(2–3):283–293, April 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004125.

Kazanci:2007:PEN

[Kaz07] Caner Kazancı. EcoNet: a new software for ecological modeling, simulation and network analysis. Ecological Modelling, 208(1):3–8, October 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002463.

Kemp:2004:PTS

[KB04] W. M. Kemp and W. R. Boynton. Productivity, trophic structure, and energy flow in the steady-state ecosystems of Silver Springs, Florida. Ecological Modelling, 178(1–2):43–49, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003005325.

Kirilenko:2000:GMV

[KBB00] Andrei P. Kirilenko, Nickolay V. Belotelov, and Boris G. Bogatyrev. Global model of vegetation migration: incorporation of climatic variability. Ecological Modelling, 132(1–2):125–133, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003100.

Kucharik:2006:MED

[KBE^{+}06] Christopher J. Kucharik, Carol C. Barford, Mustapha El Maayar, Steven C. Wofsy, Russell K. Monson, and Dennis D. Baldocchi. A multiyear evaluation of a dynamic global vegetation model at three AmeriFlux forest sites: Vegetation structure, phenology, soil temperature, and CO_{2} and H_{2}O vapor exchange. Ecological Modelling, 196(1–2):1–31, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005006186.

Kramer:2008:BGB

[KBF^{+}08] K. Kramer, J. Buiteveld, M. Forstreuter, T. Geburek, S. Leonardi, P. Menozzi, F. Povillon, M. J. Schelhaas, E. Teissier du Cros, G. G. Vendramin, and D. C. van der Werf. Bridging the gap between ecophysiological and genetic knowledge to assess the adaptive potential of European beech. Ecological Modelling, 216(3–4):333–353, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008002287.

Karlberg:2006:MTG

[KBGJS06] Louise Karlberg, Alon Ben-Gal, Per-Erik Jansson, and Uri Shani. Modelling transpiration and growth in salinitystressed tomato under different climatic conditions. Ecological Modelling, 190(1–2):15–40, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002206.

Kessler:2007:MTH

[KBK07] Michael Kessler, Jürgen Böhner, and Jürgen Kluge. Modelling tree height to assess climatic conditions at tree lines in the Bolivian Andes. Ecological Modelling, 207 (2–4):223–233, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000700275X.

Kriticos:2003:SPB

[KBM^{+}03] Darren J. Kriticos, Joel R. Brown, Gunter F. Maywald, Ian D. Radford, D. Mike Nicholas, Robert W. Sutherst, and Steve W. Adkins. SPAnDX: a process-based population dynamics model to explore management and climate change impacts on an invasive alien plant, Acacia nilotica. Ecological Modelling, 163(3):187–208, May 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000097.

Kooi:2008:STE

[KBvVK08] B. W. Kooi, D. Bontje, G. A. K. van Voorn, and S. A. L. M. Kooijman. Sublethal toxic effects in a simple aquatic food chain. Ecological Modelling, 212(3–4):304–318, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005698.

Kampichler:2000:SDF

[KBW00] Christian Kampichler, Jutta Barthel, and Ralf Wieland. Species density of foliage-dwelling spiders in field margins: a simple, fuzzy rule-based model. Ecological Modelling, 129(1):87–99, May 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002246.

Kercher:2001:PEG

[KC01] J. R. Kercher and J. Q. Chambers. Parameter estimation for a global model of terrestrial biogeochemical cycling by an iterative method. Ecological Modelling, 139(2–3):137–175, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001002344.

Kar:2003:NSH

[KC03] T. K. Kar and K. S. Chaudhuri. On non-selective harvesting of two competing fish species in the presence of toxicity. Ecological Modelling, 161(1–2):125–137, March 1, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000200323X.

Knightes:2005:EPE

[KC05] Christopher D. Knightes and Michael Cyterski. Evaluating predictive errors of a complex environmental model using a general linear model and least square means. Ecological Modelling, 186(3):366–374, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000591.

Kallay:2008:SPP

[KC08] Michael Kallay and Yosef Cohen. The simplest predator– prey model. Ecological Modelling, 218(3–4):398–399, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438000800344X.

Krivtsov:2000:IRR

[KCBS00] V. Krivtsov, J. Corliss, E. Bellinger, and D. Sigee. Indirect regulation rule for consecutive stages of ecological succession. Ecological Modelling, 133(1–2):73–81, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002817.

Keane:2004:CLF

[KCD^{+}04] Robert E. Keane, Geoffrey J. Cary, Ian D. Davies, Michael D. Flannigan, Robert H. Gardner, Sandra Lavorel, James M. Lenihan, Chao Li, and T. Scott Rupp. A classification of landscape fire succession models: spatial simulations of fire and vegetation dynamics. Ecological Modelling, 179(1):3–27, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002066.

Kim:2007:PFR

[KCJ^{+}07] Dong-Kyun Kim, Hongqing Cao, Kwang-Seuk Jeong, Friedrich Recknagel, and Gea-Jae Joo. Predictive function and rules for population dynamics of Microcystis aeruginosa in the regulated Nakdong River (South Korea), discovered by evolutionary algorithms. Ecological Modelling, 203(1–2):147–156, April 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005667.

Klanjscek:2006:IDE

[KCNN06] Tin Klanjscek, Hal Caswell, Michael G. Neubert, and Roger M. Nisbet. Integrating dynamic energy budgets into matrix population models. Ecological Modelling, 196(3–4):407–420, July 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600069X.

Kasih:2008:MED

[KCY^{+}08] G. A. Anggara Kasih, S. Chiba, Y. Yamagata, Y. Shimizu, and K. Haraguchi. Modeling early diagenesis of sediment in Ago Bay, Japan: a comparison of steady state and dynamic calculations. Ecological Modelling, 215(1–3):40–54, July 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008000951.

Komarov:2003:EMG

[KCZ^{+}03] A. Komarov, O. Chertov, S. Zudin, M. Nadporozhskaya, A. Mikhailov, S. Bykhovets, E. Zudina, and E. Zoubkova. EFIMOD 2 — a model of growth and cycling of elements in boreal forest ecosystems. Ecological Modelling, 170(2–3):373–392, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002400.

Kritzer:2005:DVW

[KD05] Jacob P. Kritzer and Campbell R. Davies. Demographic variation within spatially structured reef fish populations: when are larger-bodied subpopulations more important? Ecological Modelling, 182(1):49–65, February 25, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004004053.

Kobler:2006:HMU

[KDK06] Andrej Kobler, Sašo Džeroski, and Iphigenia Keramitsoglou. Habitat mapping using machine learning-extended kernelbased reclassification of an Ikonos satellite image. Ecological Modelling, 191(1):83–95, January 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003510.

Kent:2003:CPR

[KDS03] Adam Kent, C. Patrick Doncaster, and Tim Sluckin. Consequences for predators of rescue and Allee effects on prey.

Ecological Modelling, 162(3):233–245, April 15, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002003435.

Kocev:2009:USM

[KDW^{+}09a] Dragi Kocev, Sašo Džeroski, Matt D. White, Graeme R. Newell, and Peter Griffioen. Using single- and multitarget regression trees and ensembles to model a compound index of vegetation condition. Ecological Modelling, 220(8):1159–1168, April 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000775.

Kurz:2009:CCM

[KDW^{+}09b] W. A. Kurz, C. C. Dymond, T. M. White, G. Stinson, C. H. Shaw, G. J. Rampley, C. Smyth, B. N. Simpson, E. T. Neilson, J. A. Trofymow, J. Metsaranta, and M. J. Apps. CBM-CFS3: a model of carbon-dynamics in forestry and land-use change implementing IPCC standards. Ecological Modelling, 220(4):480–504, February 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005012.

Kohlmeier:2007:MCM

[KE07a] C. Kohlmeier and W. Ebenhöh. Mass conserving modelling of aquatic ecosystems with a variable tide level — why extrapolation from point tracers is inadmissible and how to solve the problem with a semi-Lagrangian approach. Ecological Modelling, 207(2–4):293–303, October 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000700292X.

Kohlmeier:2007:MED

[KE07b] C. Kohlmeier and W. Ebenhöh. Modelling the ecosystem dynamics and nutrient cycling of the spiekeroog back barrier system with a coupled Euler–Lagrange model on the base of ERSEM. Ecological Modelling, 202(3–4):297–310, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006005321.

Kendal:2002:SAC

[Ken02] Wayne S. Kendal. Spatial aggregation of the Colorado potato beetle described by an exponential dispersion model. Ecological Modelling, 151(2–3):261–269, June 1, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000100494X.

Kneis:2009:SWQ

[KFB09] D. Kneis, S. Förster, and A. Bronstert. Simulation of water quality in a flood detention area using models of different spatial discretization. Ecological Modelling, 220(13–14):1631–1642, July 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002610.

Komatsu:2007:MAF

[KFH07] Eiji Komatsu, Takehiko Fukushima, and Hideo Harasawa. A modeling approach to forecast the effect of long-term climate change on lake water quality. Ecological Modelling, 209(2–4):351–366, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003481.

Krause:2009:AHF

[KFJ^{+}09] Ann E. Krause, Ken A. Frank, Michael L. Jones, Thomas F. Nalepa, Richard P. Barbiero, Charles P. Madenjian, Megan Agy, Marlene S. Evans, William W. Taylor, Doran M. Mason, and Nancy J. Leonard. Adaptations in a hierarchical food web of southeastern Lake Michigan. Ecological Modelling, 220(22):3147–3162, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004955.

Klar:2008:MSN

[KFN^{+}08] C. W. Klar, P. Fiener, P. Neuhaus, V. I. S. Lenz-Wiedemann, and K. Schneider. Modelling of soil nitrogen dynamics within the decision support system DANU-BIA. Ecological Modelling, 217(1–2):181–196, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003037.

Komuro:2006:UMC

[KFR06a] Rié Komuro, E. David Ford, and Joel H. Reynolds. The use of multi-criteria assessment in developing a process model. Ecological Modelling, 197(3–4):320–330, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006001645.

Krasnopolsky:2006:NSP

[KFR06b] Vladimir M. Krasnopolsky and Michael S. Fox-Rabinovitz. A new synergetic paradigm in environmental numerical modeling: Hybrid models combining deterministic and machine learning components. Ecological Modelling, 191(1):5–18, January 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003455.

Kuka:2007:MIP

[KFR07] K. Kuka, U. Franko, and J. Rühlmann. Modelling the impact of pore space distribution on carbon turnover. Ecological Modelling, 208(2–4):295–306, November 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007003080.

Komatsu:2006:MPD

[KFS06] Eiji Komatsu, Takehiko Fukushima, and Hiroaki Shiraishi. Modeling of P-dynamics and algal growth in a stratified reservoir — mechanisms of P-cycle in water and interaction between overlying water and sediment. Ecological Modelling, 197(3–4):331–349, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001396.

Koons:2006:PMA

[KGA06] David N. Koons, James B. Grand, and Jennifer M. Arnold. Population momentum across vertebrate life histories. Ecological Modelling, 197(3–4):418–430, August 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006001657.

Kristensen:2003:MRC

[KGBC03] Nadiah Pardede Kristensen, Albert Gabric, Roger Braddock, and Roger Cropp. Is maximizing resilience compatible with established ecological goal functions? Ecological Modelling, 169(1):61–71, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002485.

Krivtsov:2008:HAS

[KGJ08] V. Krivtsov, J. Gascoigne, and S. E. Jones. Harmonic analysis of suspended particulate matter in the Menai Strait (UK). Ecological Modelling, 212(1–2):53–67, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005108.

Kumblad:2003:EME

[KGNK03] Linda Kumblad, Michael Gilek, Björn Næslund, and Ulrik Kautsky. An ecosystem model of the environmental transport and fate of carbon-14 in a bay of the Baltic Sea, Sweden. Ecological Modelling, 166(3):193–210, August 15, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003001352.

Krivtsov:2001:EMR

[KGSB01] V. Krivtsov, C. Goldspink, D. C. Sigee, and E. G. Bellinger. Expansion of the model ‘rostherne’ for fish and zooplankton: role of top-down effects in modifying the prevailing pattern of ecosystem functioning. Ecological Modelling, 138(1–3):153–171, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004002.

Koons:2005:TPD

[KGZR05] David N. Koons, James B. Grand, Bertram Zinner, and Robert F. Rockwell. Transient population dynamics: Relations to life history and initial population state. Ecological Modelling, 185(2–4):283–297, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000025.

Krivan:2000:LMP

[KH00] Vlastimil Křivan and Jan Havelka. Leslie model for predatory gall-midge population. Ecological Modelling, 126(1):73–77, February 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438009900191X.

Krysanova:2002:ANL

[KH02] Valentina Krysanova and Uwe Haberlandt. Assessment of nitrogen leaching from arable land in large river basins: Part I. Simulation experiments using a process-based model. Ecological Modelling, 150(3):255–275, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005257.

Kohler:2007:IRL

[KH07] Peter Köhler and Andreas Huth. Impacts of recruitment limitation and canopy disturbance on tropical tree species richness. Ecological Modelling, 203(3–4):511–517, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006005771.

Krivtsov:2008:MMI

[KHJ^{+}08] V. Krivtsov, M. J. Howarth, S. E. Jones, A. J. Souza, and C. F. Jago. Monitoring and modelling of the Irish Sea and Liverpool Bay: an overview and an SPM case study. Ecological Modelling, 212(1–2):37–52, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005121.

Kuo:2007:UAN

[KHLS07] Jan-Tai Kuo, Ming-Han Hsieh, Wu-Seng Lung, and Nian She. Using artificial neural network for reservoir eutrophication prediction. Ecological Modelling, 200 (1–2):171–177, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006002985.

Kratzer:2006:MIS

[KHT06] Jud F. Kratzer, Daniel B. Hayes, and Bradley E. Thompson. Methods for interpolating stream width, depth, and current velocity. Ecological Modelling, 196(1–2):256–264, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000627.

Karouby:2007:SMD

[KID^{+}07] Oren Karouby, Athanassios Iliadis, Jean-Pierre Durbec, Virginie Riandey, and François Carlotti. Stochastic modelling for determining zooplankton abundance. Ecological Modelling, 204(3–4):372–378, June 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000269.

Krivtsov:2003:SAC

[KIL^{+}03] V. Krivtsov, J. B. Illian, K. Liddell, A. Garside, T. Bezginova, R. Salmond, J. Thompson, B. Griffiths, H. J. Staines, R. Watling, A. Brendler, and J. W. Palfreyman. Some aspects of complex interactions involving soil mesofauna: analysis of the results from a Scottish woodland. Ecological Modelling, 170(2–3):441–452, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002448.

King:2004:CNC

[Kin04] Robert J. King. Creating a new conceptual framework: environment power and society. Ecological Modelling, 178(1–2):159–162, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005337.

King:2005:LTF

[Kin05] David A. King. Linking tree form, allocation and growth with an allometrically explicit model. Ecological Modelling, 185(1):77–91, June 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005976.

Kirsta:2001:IHO

[Kir01] Yu. B. Kirsta. Information-hierarchical organization of biosphere and problems of its sustainable development. Ecological Modelling, 145(1):49–59, November 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003829.

Kirsta:2006:SAMa

[Kir06a] Yuri B. Kirsta. System-analytical modelling — Part I: General principles and theoretically best accuracies of ecological models. Soil-moisture exchange in agroecosystems. Ecological Modelling, 191(3–4):315–330, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002796.

Kirsta:2006:SAMb

[Kir06b] Yuri B. Kirsta. System-analytical modelling — part II: Wheat biotime run and yield formation. Agroclimatic potential, the Le Chatelier principle, and changes in agroclimatic potential and climate in Russia and the U.S. Ecological Modelling, 191(3–4):331–345, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002802.

Korner:2008:DGP

[KJ08] Katrin Körner and Florian Jeltsch. Detecting general plant functional type responses in fragmented landscapes using spatially-explicit simulations. Ecological Modelling, 210(3):287–300, January 24, 2008. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004085.

Krause:2007:MIL

[KJB07] Stefan Krause, Jörg Jacobs, and Axel Bronstert. Modelling the impacts of land-use and drainage density on the water balance of a lowland–floodplain landscape in northeast Germany. Ecological Modelling, 200(3–4):475–492, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003930.

Kangas:2000:PFC

[KK00a] A. Kangas and S. Kurki. Predicting the future of capercaillie (Tetrao urogallus) in Finland. Ecological Modelling, 134(1):73–87, September 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003379.

Kooi:2000:ISC

[KK00b] B. W. Kooi and S. A. L. M. Kooijman. Invading species can stabilize simple trophic systems. Ecological Modelling, 133(1–2):57–72, August 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002805.

Kuroda:2004:DAT

[KK04] Hiroshi Kuroda and Michio J. Kishi. A data assimilation technique applied to estimate parameters for the NEMURO marine ecosystem model. Ecological Modelling, 172(1):69–85, February 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003624.

Kubo:2007:MSB

[KK07] Takuya Kubo and Tsuyoshi Kobayashi. Modelling of stem biomass accumulation in Pinus densiflora seedlings exposed to aqueous-phase OH radicals generating mist. Ecological Modelling, 200(3–4):467–474, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003929.

Karev:2008:ESP

[KK08] Georgy P. Karev and Vladimir N. Korotkov. Ergodicity and successions in Prioksko–Terrasnyi Biosphere Reserve. Ecological Modelling, 212(1–2):116–121, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005212.

Kennen:2008:UIF

[KKA^{+}08] Jonathan G. Kennen, Leon J. Kauffman, Mark A. Ayers, David M. Wolock, and Susan J. Colarullo. Use of an integrated flow model to estimate ecologically relevant hydrologic characteristics at stream biomonitoring sites. Ecological Modelling, 211(1–2):57–76, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004279.

Kim:2006:IWA

[KKCC06] Cheol-Ki Kim, Inn-Sil Kwak, Eui-Young Cha, and Tae-Soo Chon. Implementation of wavelets and artificial neural networks to detection of toxic response behavior of chironomids (Chironomidae: Diptera) for water quality monitoring. Ecological Modelling, 195(1–2):61–71, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005788.

Kumada:2008:LCD

[KKH^{+}08] Shungo Kumada, Takuya Kawanishi, Yoshishige Hayashi, Kenichi Ogomori, Yuka Kobayashi, Nobuhide Takahashi, Masahiro Saito, Hiroyuki Hamano, Toshinori Kojima, and Koichi Yamada. Litter carbon dynamics analysis in forests in an arid ecosystem with a model incorporating the physical removal of litter. Ecological Modelling, 215(1–3):190–199, July 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008001051.

Kumada:2009:EDM

[KKH^{+}09] Shungo Kumada, Takuya Kawanishi, Yoshishige Hayashi, Hiroyuki Hamano, Satoko Kawarasaki, Shin ichi Aikawa, Nobuhide Takahashi, Yasuyuki Egashira, Hiroyuki Tanouchi, Toshinori Kojima, Adrianne Kinnear, and Koichi Yamada. Effects of different mobilities of leaf and woody litters on litter carbon dynamics in arid ecosystems in Western Australia. Ecological Modelling, 220(20):2792–2801, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004967.

Kokkonen:2006:IPS

[KKL^{+}06] T. Kokkonen, H. Koivusalo, A. Laurén, S. Penttinen, M. Starr, S. Kellomäki, and L. Finér. Implications of processing spatial data from a forested catchment for a hillslope hydrological model. Ecological Modelling, 199(4):393–408, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006002845.

Kishi:2007:NLT

[KKW^{+}07] Michio J. Kishi, Makoto Kashiwai, Daniel M. Ware, Bernard A. Megrey, David L. Eslinger, Francisco E. Werner, Maki Noguchi-Aita, Tomonori Azumaya, Masahiko Fujii, Shinji Hashimoto, Daji Huang, Hitoshi Iizumi, Yukimasa Ishida, Sukyung Kang, Gennady A. Kantakov, Hyun cheol Kim, Kosei Komatsu, Vadim V. Navrotsky, S. Lan Smith, Kazuaki Tadokoro, Atsushi Tsuda, Orio Yamamura, Yasuhiro Yamanaka, Katsumi Yokouchi, Naoki Yoshie, Jing Zhang, Yury I. Zuenko, and Vladimir I. Zvalinsky. NE-MURO — a lower trophic level model for the North Pacific marine ecosystem. Ecological Modelling, 202(1–2):12–25, March 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006004534.

Kasimov:2008:MGF

[KKW08] Nikolaj Kasimov, Natalia Kosheleva, and Victor Wagner. Modeling geochemical fields based on landscape-guided interpolation. Ecological Modelling, 212(1–2):109–115, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005248.

Kuijper:2003:OFW

[KKZK03] Lothar D. J. Kuijper, Bob W. Kooi, Cor Zonneveld, and Sebastiaan A. L. M. Kooijman. Omnivory and food web dynamics. Ecological Modelling, 163(1–2):19–32, May 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003514.

Klanjscek:2001:TCL

[KL01] Tin Klanjšček and Tarzan Legović. Toward a closed life support system for interplanetary missions. Ecological Modelling, 138(1–3):41–54, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003926.

Kim:2003:OMC

[KL03] Dong-Soon Kim and Joon-Ho Lee. Oviposition model of Carposina sasakii (Lepidoptera: Carposinidae). Ecological Modelling, 162(1–2):145–153, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004027.

Klanjscek:2007:AEE

[KL07] J. Klanjšček and T. Legović. Is anchovy (Engraulis encrasicolus, L.) overfished in the Adriatic Sea? Ecological Modelling, 201(3–4):312–316, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004479.

Komarov:2008:EFE

[KL08] Alexander S. Komarov and Tarzan Legović. ECEM 2005: the Fifth European Conference on Ecological Modelling, 19–23 September 2005, Pushchino, Russia. Ecological Modelling, 212(1–2):1–2, March 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005030.

Klanjscek:2009:RSC

[KL09] Jasminka Klanjšček and Tarzan Legović. Reply to Santojanni’s comments on “Is anchovy (Engraulis encrasicolus, L.) overfished in the Adriatic Sea?”, Ecol. Model. 201: 312–316. Ecological Modelling, 220(3):434–435, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005061.

Korotkov:2001:SMB

[KLL01] Vladimir N. Korotkov, Dmitrii O. Logofet, and Michel Loreau. Succession in mixed boreal forest of Russia: Markov models and non-Markov effects. Ecological Modelling, 142(1–2):25–38, August 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002472.

Kannel:2007:AAQ

[KLL^{+}07] Prakash Raj Kannel, S. Lee, Y.-S. Lee, S. R. Kanel, and G. J. Pelletier. Application of automated QUAL2Kw for water quality modeling and management in the Bagmati River, Nepal. Ecological Modelling, 202(3–4):503–517, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007000051.

Keller:2002:ECF

[KLM^{+}02] Franziska Keller, Heike Lischke, Thomas Mathis, Adrian Möhl, Lucia Wick, Brigitta Ammann, and Felix Kienast. Effects of climate, fire, and humans on forest dynamics: forest simulations compared to the palaeological record. Ecological Modelling, 152(2–3):109–127, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200011X.

Kruk:2007:FAL

[KLPP07] A. Kruk, S. Lek, Y.-S. Park, and T. Penczak. Fish assemblages in the large lowland Narew River system (Poland):

Application of the self-organizing map algorithm. Ecological Modelling, 203(1–2):45–61, April 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005576.

Kuwata:2000:EAS

[KM00] Akira Kuwata and Tatsuo Miyazaki. Effects of ammonium supply rates on competition between Microcystis novacekii (Cyanobacteria) and Scenedesmus quadricauda (Chlorophyta): simulation study. Ecological Modelling, 135(1):81–87, November 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000363X.

Kincanon:2004:ECU

[KM04] Renee Kincanon and A. Steve McAnally. Enhancing commonly used model predictions for constructed wetland performance: as-built design considerations. Ecological Modelling, 174(3):309–322, May 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004307.

Katajisto:2006:KBH

[KM06] Jonna Katajisto and Atte Moilanen. Kernel-based home range method for data with irregular sampling intervals. Ecological Modelling, 194(4):405–413, April 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005005648.

Kim:2008:MSM

[KMB08] Eunhee Kim, Robert P. Mason, and Christine M. Bergeron. A modeling study on methylmercury bioaccumulation and its controlling factors. Ecological Modelling, 218(3–4):267–289, November 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003499.

Kishi:2007:MDM

[KMiIW07] Michio J. Kishi, Bernard A. Megrey, Shin ichi Ito, and Francisco E. Werner. In memoriam Daniel M. Ware. Ecological Modelling, 202(1–2):1, March 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004509.

Kishi:2007:P

[KMIW07] Michio J. Kishi, Bernard A. Megrey, Shin-Ichi Ito, and Francisco E. Werner. Preface. Ecological Modelling, 202(1–2):3– 6, March 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006004510.

Kayombo:2000:MDV

[KMM^{+}00] S. Kayombo, T. S. A. Mbwette, A. W. Mayo, J. H. Y. Katima, and S. E. Jorgensen. Modelling diurnal variation of dissolved oxygen in waste stabilization ponds. Ecological Modelling, 127(1):21–31, March 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001969.

Komatsu:2007:EAP

[KMN^{+}07] Kosei Komatsu, Yasuo Matsukawa, Kaoru Nakata, Tadafumi Ichikawa, and Katsuyuki Sasaki. Effects of advective processes on planktonic distributions in the Kuroshio region using a 3-D lower trophic model and a data assimilative OGCM. Ecological Modelling, 202(1–2):105–119, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004601.

Kuper:2003:IME

[KMPB03] M. Kuper, C. Mullon, Y. Poncet, and E. Benga. Integrated modelling of the ecosystem of the Niger river inland delta in Mali. Ecological Modelling, 164(1):83–102, June 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000061.

Kazama:2007:CFH

[KMR^{+}07] So Kazama, Satoru Matsumoto, S. Priyantha Ranjan, Hiroshi Hamamoto, and Masaki Sawamoto. Characterization of firefly habitat using a geographical information system with hydrological simulation. Ecological Modelling, 209(2–4):392–400, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003511.

Kuparinen:2007:MAM

[KMRV07] Anna Kuparinen, Tiina Markkanen, Hermanni Riikonen, and Timo Vesala. Modeling air-mediated dispersal of spores, pollen and seeds in forested areas. Ecological Modelling, 208(2–4):177–188, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003018.

Kazanci:2009:CEI

[KMT09] C. Kazanci, L. Matamba, and E. W. Tollner. Cycling in ecosystems: an individual based approach. Ecological Modelling, 220(21):2908–2914, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004316.

Karev:2008:MDI

[KNB08a] Georgy P. Karev, Artem S. Novozhilov, and Faina S. Berezovskaya. Modeling the dynamics of inhomogeneous natural rotifer populations under toxicant exposure. Ecological Modelling, 212(1–2):80–85, March 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005170.

Kuznetsov:2008:MSE

[KNB08b] I. Kuznetsov, T. Neumann, and H. Burchard. Model study on the ecosystem impact of a variable C:N:P ratio for cyanobacteria in the Baltic proper. Ecological Modelling, 219(1–2):107–114, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800402X.

Kavanagh:2004:APO

[KNCP04] Patricia Kavanagh, Nathaniel Newlands, Villy Christensen, and Daniel Pauly. Automated parameter optimization for Ecopath ecosystem models. Ecological Modelling, 172(2–4):141–149, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003661.

Koniak:2009:HMS

[KNM09] Gili Koniak and Imanuel Noy-Meir. A hierarchical, multiscale, management-responsive model of Mediterranean vegetation dynamics. Ecological Modelling, 220(8):1148–1158, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009000878.

Knoke:2003:PRH

[Kno03] Thomas Knoke. Predicting red heartwood formation in beech trees (Fagus sylvatica L.). Ecological Modelling, 169(2–3):295–312, November 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300276X.

Kuechler:2004:IGD

[KNZ04] Roland Kuechler, Klaus Noack, and Torsten Zorn. Investigation of gypsum dissolution under saturated and unsaturated water conditions. Ecological Modelling, 176(1–2):1–14, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005374.

Kindvall:2000:CMI

[KP00] Oskar Kindvall and Anders Petersson. Consequences of modelling interpatch migration as a function of patch geometry when predicting metapopulation extinction risk. Ecological Modelling, 129(1):101–109, May 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002155.

Khan:2009:AII

[KP09] M. Feroz Khan and Preetha Panikkar. Assessment of impacts of invasive fishes on the food web structure and ecosystem properties of a tropical reservoir in India. Ecological Modelling, 220(18):2281–2290, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003822.

Kooi:2002:AMF

[KPAK02] B. W. Kooi, J. C. Poggiale, P. Auger, and S. A. L. M. Kooijman. Aggregation methods in food chains with nutrient recycling. Ecological Modelling, 157(1):69–86, November 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000200217X.

Keane:2002:EHR

[KPH02] Robert E. Keane, Russell A. Parsons, and Paul F. Hessburg. Estimating historical range and variation of landscape patch dynamics: limitations of the simulation approach. Ecological Modelling, 151(1):29–49, May 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004707.

Kangur:2007:PLT

[KPK^{+}07] Külli Kangur, Young-Seuk Park, Andu Kangur, Peeter Kangur, and Sovan Lek. Patterning long-term changes of fish community in large shallow Lake Peipsi. Ecological Modelling, 203(1–2):34–44, April 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005564.

Karakitsios:2006:APB

[KPKP06] Spyros P. Karakitsios, Costas L. Papaloukas, Pavlos A. Kassomenos, and Georgios A. Pilidis. Assessment and prediction of benzene concentrations in a street canyon using artificial neural networks and deterministic models: Their response to “what if” scenarios. Ecological Modelling, 193(3–4):253–270, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003984.

Komarov:2003:CDD

[KPS03] A. S. Komarov, M. M. Palenova, and O. V. Smirnova. The concept of discrete description of plant ontogenesis and cellular automata models of plant populations. Ecological Modelling, 170(2–3):427–439, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002436.

Kallimanis:2009:DPP

[KPT^{+}09] A. S. Kallimanis, T. Petanidou, J. Tzanopoulos, J. D. Pantis, and S. P. Sgardelis. Do plant–pollinator interaction networks result from stochastic processes? Ecological Modelling, 220(5):684–693, March 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005474.

Koneff:2004:MWC

[KR04] Mark D. Koneff and J. Andrew Royle. Modeling wetland change along the United States Atlantic Coast. Ecological Modelling, 177(1–2):41–59, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001061.

Kristensen:2003:MAL

[Kri03] Søren P. Kristensen. Multivariate analysis of landscape changes and farm characteristics in a study area in central Jutland, Denmark. Ecological Modelling, 168(3):303–318, October 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001431.

Krivtsov:2004:IIR

[Kri04a] V. Krivtsov. Investigations of indirect relationships in ecology and environmental sciences: a review and the implications for comparative theoretical ecosystem analysis. Ecological Modelling, 174(1–2):37–54, May 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005672.

Krivtsov:2004:LAC

[Kri04b] V. Krivtsov. Lugano aftermath and contemporary advances in ecological modelling. Ecological Modelling, 174(1–2):1–3, May 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005635.

Karlsson:2005:MOM

[KRK05] O. Magnus Karlsson, John S. Richardson, and Peter M. Kiffney. Modelling organic matter dynamics in headwater streams of south-western British Columbia, Canada. Ecological Modelling, 183(4):463–476, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005046.

Kint:2004:ESM

[KRN04] Vincent Kint, De Wulf Robert, and Lust Noël. Evaluation of sampling methods for the estimation of structural indices in forest stands. Ecological Modelling, 180(4):461–476, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004002959.

Kumar:2002:RBV

[KRvL^{+}02] Lalit Kumar, Max Rietkerk, Frank van Langevelde, Johan van de Koppel, Jelte van Andel, John Hearne, Nico de Ridder, Leo Stroosnijder, Andrew K. Skidmore, and Herbert H. T. Prins. Relationship between vegetation growth rates at the onset of the wet season and soil type in the Sahel of Burkina Faso: implications for resource utilisation at large scales. Ecological Modelling, 149(1–2):143–152, March 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001005208.

Keane:2007:USH

[KRZ07] Robert E. Keane, Matthew Rollins, and Zhi-Liang Zhu. Using simulated historical time series to prioritize fuel treatments on landscapes across the United States: the LANDFIRE prototype project. Ecological Modelling, 204(3–4):485–502, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007000580.

Kindlmann:2004:TFT

[KS04a] Pavel Kindlmann and Bernhard Stadler. Temporal fluctuations in throughfall carbon concentrations in a spruce forest. Ecological Modelling, 176(3–4):381–388, September 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004001000.

Kmet:2004:FAF

[KS04b] T. Kmeť and M. Straškraba. Feeding adaptations of filter feeders: Daphnia. Ecological Modelling, 178(3– 4):313–327, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001504.

Kuparinen:2007:FMF

[KS07] Anna Kuparinen and Frank M. Schurr. A flexible modelling framework linking the spatio-temporal dynamics of plant genotypes and populations: Application to gene flow from transgenic forests. Ecological Modelling, 202(3–4):476–486, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006005503.

Knoke:2008:ISE

[KS08] Thomas Knoke and Thomas Seifert. Integrating selected ecological effects of mixed European beech–Norway spruce stands in bioeconomic modelling. Ecological Modelling, 210(4):487–498, February 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004206.

Karul:2000:CSU

[KSÇ^{+}00] Cüneyt Karul, Selçuk Soyupak, Ahmet F. Çilesiz, Nihat Akbay, and Emin Germen. Case studies on the use of neural networks in eutrophication modeling. Ecological Modelling, 134(2–3):145–152, October 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003604.

Kocak:2004:NPN

[KŞE04] Kasım Koçak, Levent Şaylan, and Josef Eitzinger. Nonlinear prediction of near-surface temperature via univariate and multivariate time series embedding. Ecological Modelling, 173(1):1–7, March 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002497.

Kohlmann:2005:PVA

[KSG05] Stephan G. Kohlmann, Gregory A. Schmidt, and David K. Garcelon. A population viability analysis for the Island Fox on Santa Catalina Island, California. Ecological Modelling, 183(1):77–94, April 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004326.

Karim:2003:SFB

[KSH^{+}03] Md. Rezaul Karim, Masahiko Sekine, Takaya Higuchi, Tsuyoshi Imai, and Masao Ukita. Simulation of fish behavior and mortality in hypoxic water in an enclosed bay. Ecological Modelling, 159(1):27–42, January 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200282X.

Kramer-Schadt:2007:PPS

[KSRWG07] Stephanie Kramer-Schadt, Eloy Revilla, Thorsten Wiegand, and Volker Grimm. Patterns for parameters in simulation models. Ecological Modelling, 204(3–4):553–556, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007000452.

Kilic:2007:ANB

[KST^{+}07] Hürevren Kılıç, Selçuk Soyupak, İlhami Tüzün, Özlem İnce, and Gökben Başaran. An automata networks based preprocessing technique for artificial neural network modelling of primary production levels in reservoirs. Ecological Modelling, 201(3–4):359–368, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004650.

Kones:2009:NIR

[KSvOO09] Julius K. Kones, Karline Soetaert, Dick van Oevelen, and John O. Owino. Are network indices robust indicators of food web functioning? A Monte Carlo approach. Ecological Modelling, 220(3):370–382, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005024.

Kaewmanee:2003:CAS

[KT03] C. Kaewmanee and I. M. Tang. Cannibalism in an age-structured predator–prey system. Ecological Modelling, 167(3):213–220, September 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300190X.

Kim:2007:RPP

[KTB07] Seong-Hee Kim, John Tschirhart, and Steven W. Buskirk. Reconstructing past population processes with general equilibrium models: House mice in Kern County, California, 1926–1927. Ecological Modelling, 209(2–4):235–248, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007003407.

Kitazawa:2008:CSB

[KTKR08] Daisuke Kitazawa, Shigeru Tabeta, Takayoshi Kato, and Piet Ruardij. A comparative study of the biomass-based and individual-based models of blue mussels. Ecological Modelling, 215(1–3):93–104, July 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000926.

Kaloudis:2005:AWD

[KTL^{+}05] Spiros Kaloudis, Athena Tocatlidou, Nikos A. Lorentzos, Alexander B. Sideridis, and Michael Karteris. Assessing wildfire destruction danger: a decision support system incorporating uncertainty. Ecological Modelling, 181(1):25–38, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003564.

Krcmar:2005:MFM

[KvKV05] Emina Krcmar, G. Cornelis van Kooten, and Ilan Vertinsky. Managing forest and marginal agricultural land for multiple tradeoffs: compromising on economic, carbon and structural diversity objectives. Ecological Modelling, 185(2–4):451–468, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005000116.

Krivtsov:2009:FMT

[KVL^{+}09] V. Krivtsov, O. Vigy, C. Legg, T. Curt, E. Rigolot, I. Lecomte, M. Jappiot, C. Lampin-Maillet, P. Fernandes, and G. B. Pezzatti. Fuel modelling in terrestrial ecosystems: an overview in the context of the development of an object-orientated database for wild fire analysis. Ecological Modelling, 220(21):2915–2926, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005778.

Kavadia:2007:DFL

[KVPA07] A. Kavadia, D. V. Vayenas, S. Pavlou, and G. Aggelis. Dynamics of free-living nitrogen-fixing bacterial populations in antagonistic conditions. Ecological Modelling, 200(1–2):243–253, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000600367X.

Kavadia:2008:DFL

[KVPA08] A. Kavadia, D. V. Vayenas, S. Pavlou, and G. Aggelis. Dynamics of free-living nitrogen-fixing bacterial populations and nitrogen fixation in a two-prey–one-predator system. Ecological Modelling, 218(3–4):323–338, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003657.

Kizner:2001:GFH

[KVV01] Z. Kizner, R. Vago, and L. Vaky. Growth forms of hermatypic corals: stable states and noise-induced transitions. Ecological Modelling, 141(1–3):227–239, July 1, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001002769.

Kellomaki:2000:STE

[KW00] S. Kellomäki and K.-Y. Wang. Short-term environmental controls on carbon dioxide flux in a boreal coniferous forest: model computation compared with measurements by eddy covariance. Ecological Modelling, 128(1):63–88, April 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002240.

King:2002:DSS

[KW02a] Anthony W. King and Kimberly A. With. Dispersal success on spatially structured landscapes: when do spatial pattern and dispersal behavior really matter? Ecological Modelling, 147(1):23–39, January 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004008.

Kok:2002:MLU

[KW02b] Kasper Kok and Manuel Winograd. Modelling landuse change for Central America, with special reference to the impact of Hurricane Mitch. Ecological Modelling, 149(1–2):53–69, March 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005142.

Kirkpatrick:2005:QSS

[KW05] Lee Anne Kirkpatrick and John F. Weishampel. Quantifying spatial structure of volumetric neutral models. Ecological Modelling, 186(3):312–325, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000554.

Krivtsov:2004:AEF

[KWD^{+}04] V. Krivtsov, P. A. Wäger, P. Dacombe, P. W. Gilgen, S. Heaven, L. M. Hilty, and C. J. Banks. Analysis of energy footprints associated with recycling of glass and plastic — case studies for industrial ecology. Ecological Modelling, 174(1–2):175–189, May 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000067.

Kathuroju:2007:UAV

[KWS^{+}07] Naven Kathuroju, Michael A. White, Jürgen Symanzik, Mark D. Schwartz, James A. Powell, and Ramakrishna R. Nemani. On the use of the advanced very high resolution radiometer for development of prognostic land surface phenology models. Ecological Modelling, 201(2):144–156, February 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006004169.

Krivtsov:2003:AFF

[KWW^{+}03] V. Krivtsov, R. Watling, S. J. J. Walker, D. Knott, J. W. Palfreyman, and H. J. Staines. Analysis of fungal fruiting patterns at the dawyck botanic garden. Ecological Modelling, 170(2–3):393–406, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002412.

Kriticos:2009:PBP

[KWW^{+}09] Darren J. Kriticos, Michael S. Watt, Toni M. Withers, Agathe Leriche, and Michelle C. Watson. A process-based population dynamics model to explore target and non-target impacts of a biological control agent. Ecological Modelling, 220(17):2035–2050, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900307X.

Kaitala:2000:NUP

[KYH00] Veijo Kaitala, Janica Ylikarjula, and Mikko Heino. Nonunique population dynamics: basic patterns. Ecological Modelling, 135(2–3):127–134, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003574.

Keats:2007:ECO

[KYL07] Andrew Keats, Eugene Yee, and Fue-Sang Lien. Efficiently characterizing the origin and decay rate of a nonconservative scalar using probability theory. Ecological Modelling, 205(3–4):437–452, July 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001275.

Kemp:2007:ADR

[KZH07] Stanley J. Kemp, Patricia Zaradic, and Frank Hansen. An approach for determining relative input parameter importance and significance in artificial neural networks. Ecological Modelling, 204(3–4):326–334, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000245.

Linacre:2004:EID

[LA04] Nicholas A. Linacre and Peter K. Ades. Estimating isolation distances for genetically modified trees in plantation forestry. Ecological Modelling, 179(3):247–257, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004000249.

Laughlin:2007:ABF

[LA07] Daniel C. Laughlin and Scott R. Abella. Abiotic and biotic factors explain independent gradients of plant community composition in ponderosa pine forests. Ecological Modelling, 205(1–2):231–240, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000944.

Linderman:2005:MST

[LAB^{+}05] Marc A. Linderman, Li An, Scott Bearer, Guangming He, Zhiyun Ouyang, and Jianguo Liu. Modeling the spatio-temporal dynamics and interactions of households, landscapes, and giant panda habitat. Ecological Modelling, 183(1):47–65, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004302.

Larocque:2006:MFS

[LAD06] Guy R. Larocque, Louis Archambault, and Claude Delisle. Modelling forest succession in two southeastern Canadian mixedwood ecosystem types using the ZELIG model. Ecological Modelling, 199(3):350–362, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002353.

Lek-Ang:2007:CCP

[LAPAMD07] Sithan Lek-Ang, Young-Seuk Park, Salah Ait-Mouloud, and Louis Deharveng. Collembolan communities in a peat bog versus surrounding forest analyzed by using self-organizing map. Ecological Modelling, 203(1–2):9–17, April 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006005539.

Larocque:2002:CDP

[Lar02] Guy R. Larocque. Coupling a detailed photosynthetic model with foliage distribution and light attenuation functions to compute daily gross photosynthesis in sugar maple (Acer saccharum Marsh.) stands. Ecological Modelling, 148(3):213–232, March 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004422.

Lasaponara:2006:ESS

[Las06a] R. Lasaponara. Estimating spectral separability of satellite derived parameters for burned areas mapping in the Calabria region by using SPOT-Vegetation data. Ecological Modelling, 196(1–2):265–270, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000676.

Lasaponara:2006:UPC

[Las06b] R. Lasaponara. On the use of principal component analysis (PCA) for evaluating interannual vegetation anomalies from SPOT/VEGETATION NDVI temporal series. Ecological Modelling, 194(4):429–434, April 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005454.

Latham:2006:NFA

[Lat06] Luke G. Latham. Network flow analysis algorithms. Ecological Modelling, 192(3–4):586–600, February 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005003935.

Lewis:2007:MPF

[LAXP07] Gilbert N. Lewis, Martin T. Auer, Xinyu Xiang, and Michael R. Penn. Modeling phosphorus flux in the sediments of Onondaga Lake: Insights on the timing of lake response and recovery. Ecological Modelling, 209(2–4):121–135, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003237.

Liu:2008:RMP

[LAZ^{+}08] Shuguang Liu, Pamela Anderson, Guoyi Zhou, Boone Kauffman, Flint Hughes, David Schimel, Vicente Watson, and Joseph Tosi. Resolving model parameter values from carbon and nitrogen stock measurements in a wide range of tropical mature forests using nonlinear inversion and regression trees. Ecological Modelling, 219(3–4):327–341, December 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008003578.

Liu:2001:SBS

[LB01] D. L. Liu and T. A. Bull. Simulation of biomass and sugar accumulation in sugarcane using a process-based model. Ecological Modelling, 144(2–3):181–211, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003726.

Larocque:2008:UAC

[LBBC08] Guy R. Larocque, Jagtar S. Bhatti, Robert Boutin, and Oleg Chertov. Uncertainty analysis in carbon cycle models of forest ecosystems: Research needs and development of a theoretical framework to estimate error propagation. Ecological Modelling, 219(3–4):400–412, December 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003542.

Leguerrier:2006:PAF

[LBBN06] Delphine Leguerrier, Cédric Bacher, Eric Benoît, and Nathalie Niquil. A probabilistic approach of flow-balanced network based on Markov chains. Ecological Modelling, 193(3–4):295–314, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004242.

Loiselle:2001:MEF

[LBBR01] Steven Loiselle, Luca Bracchini, Claudia Bonechi, and Claudio Rossi. Modelling energy fluxes in remote wetland ecosystems with the help of remote sensing. Ecological Modelling, 145(2–3):243–261, November 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003945.

Larocque:2008:IUS

[LBL^{+}08] Guy R. Larocque, Jagtar S. Bhatti, Jinxun Liu, James C. Ascough, Nancy Luckai, and Andrew M. Gordon. The importance of uncertainty and sensitivity analyses in processbased models of carbon and nitrogen cycling in terrestrial ecosystems with particular emphasis on forest ecosystems: Selected papers from a workshop organized by the International Society for Ecological Modelling (ISEM) at the third biennial meeting of the International Environmental Modelling and Software Society (IEMSS) in Burlington, Vermont, USA, August 9–13, 2006. Ecological Modelling, 219(3–4):261–263, December 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800358X.

LeBer:2006:SCS

[LBS^{+}06] F. Le Ber, M. Benoît, C. Schott, J.-F. Mari, and C. Mignolet. Studying crop sequences with C arrotAge, a HMM-based data mining software. Ecological Modelling, 191(1):170–185, January 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005003844.

Lee:2008:TSO

[LBS08a] S.-H. Lee, P. Bardunias, and N.-Y. Su. Two strategies for optimizing the food encounter rate of termite tunnels simulated by a lattice model. Ecological Modelling, 213(3–4):381–388, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008000082.

Lopez:2008:MAS

[LBS08b] Bruno Díaz López, Mandy Bunke, and Julia Andrea Bernal Shirai. Marine aquaculture off Sardinia Island (Italy): Ecosystem effects evaluated through a trophic mass-balance model. Ecological Modelling, 212(3–4):292–303, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005571.

Li:2001:DSR

[LC01] Bai-Lian Li and Eric L. Charnov. Diversity–stability relationships revisited: scaling rules for biological communities near equilibrium. Ecological Modelling, 140(3):247–254, June 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003209.

Lewison:2004:EBU

[LC04] Rebecca L. Lewison and Jacoby Carter. Exploring behavior of an unusual megaherbivore: a spatially explicit foraging model of the hippopotamus. Ecological Modelling, 171(1–2):127–138, January 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003053.

Lenser:2007:NAM

[LC07a] Thorsten Lenser and Andrew Constable. A nonparametric algorithm to model movement between polygon subdomains in a spatially explicit ecosystem model. Ecological Modelling, 206(1–2):79–92, August 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001597.

Liedloff:2007:MER

[LC07b] Adam C. Liedloff and Garry D. Cook. Modelling the effects of rainfall variability and fire on tree populations in an Australian tropical savanna with the flames simulation model. Ecological Modelling, 201(3–4):269–282, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004236.

Lin:2008:PES

[LC08] Hua Lin and Min Cao. Plant energy storage strategy and caloric value. Ecological Modelling, 217(1– 2):132–138, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002913.

Lazzarotto:2009:DGC

[LCF09] P. Lazzarotto, P. Calanca, and J. Fuhrer. Dynamics of grass–clover mixtures — an analysis of the response to management with the PROductive GRASsland simulator (PRO-GRASS). Ecological Modelling, 220(5):703–724, March 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005747.

Link:2009:RBN

[LCG^{+}09] Jason Link, Laurel Col, Vincent Guida, David Dow, John O’Reilly, Jack Green, William Overholtz, Debra Palka, Chris Legault, Joseph Vitaliano, Carolyn Griswold, Michael Fogarty, and Kevin Friedland. Response of balanced network models to large-scale perturbation: Implications for evaluating the role of small pelagics in the Gulf of Maine. Ecological Modelling, 220(3):351–369, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004973.

Loiselle:2000:FAR

[LCH^{+}00] Steven Loiselle, Giuseppe M. Carpaneto, Vincent Hull, Tomas Waller, and Claudio Rossi. Feedback analysis in reserve management: studying local myths using qualitative models. Ecological Modelling, 129(1):25–37, May 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002179.

Liu:2007:FWS

[LCLC07] Qi-Gen Liu, Yong Chen, Jia-Le Li, and Li-Qiao Chen. The food web structure and ecosystem properties of a filterfeeding carps dominated deep reservoir ecosystem. Ecological Modelling, 203(3–4):279–289, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600603X.

Llorens:2009:IEH

[LCM^{+}09a] Esther Llorens, Joaquim Comas, Eugènia Martí, Joan Lluís Riera, Francesc Sabater, and Manel Poch. Integrating empirical and heuristic knowledge in a KBS to approach stream eutrophication. Ecological Modelling, 220 (18):2162–2172, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003949.

Lopes:2009:MTP

[LCM^{+}09b] José Fortes Lopes, Ana Cristina Cardoso, Maria Teresa Moita, Alfredo C. Rocha, and Juan A. Ferreira. Modelling the temperature and the phytoplankton distributions at the aveiro near coastal zone, Portugal. Ecological Modelling, 220(7):940–961, April 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005371.

Libralato:2006:MIK

[LCP06] Simone Libralato, Villy Christensen, and Daniel Pauly. A method for identifying keystone species in food web models. Ecological Modelling, 195(3–4):153–171, June 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005006149.

Lophaven:2006:SMD

[LCR06] Søren Lophaven, Jacob Carstensen, and Helle Rootzén. Stochastic modelling of dissolved inorganic nitrogen in space and time. Ecological Modelling, 193(3–4):467–478, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004059.

Lin:2009:ASO

[LCSZ09] Hua Lin, Min Cao, Paul C. Stoy, and Yiping Zhang. Assessing self-organization of plant communities — a thermodynamic approach. Ecological Modelling, 220 (6):784–790, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009000295.

Li:2009:ENA

[LCY09] Y. Li, B. Chen, and Z. F. Yang. Ecological network analysis for water use systems — a case study of the Yellow River Basin. Ecological Modelling, 220(22):3163–3173, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000900564X.

Laganis:2006:SAE

[LD06a] Jana Laganis and Marko Debeljak. Sensitivity analysis of the emergy flows at the solar salt production process in Slovenia. Ecological Modelling, 194(1–3):287–295, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005260.

Leung:2006:MSD

[LD06b] Brian Leung and David G. Delaney. Managing sparse data in biological invasions: a simulation study. Ecological Modelling, 198(1–2):229–239, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600202X.

Liu:2000:SSS

[LDM00] Yir-Chung Liu, Richard Durrett, and Michael G. Milgroom. A spatially-structured stochastic model to simulate heterogeneous transmission of viruses in fungal populations. Ecological Modelling, 127(2–3):291–308, March 30, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099002161.

Loibel:2006:IRG

[LdVA06] Selene Loibel, João B. R. do Val, and Marinho G. Andrade. Inference for the Richards growth model using Box and Cox transformation and bootstrap techniques. Ecological Modelling, 191(3–4):501–512, February 5, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002814.

Ledermann:2004:MAU

[LE04] Thomas Ledermann and Otto Eckmüllner. A method to attain uniform resolution of the competition variable basalarea-in-larger trees (BAL) during forest growth projections of small plots. Ecological Modelling, 171(1–2):195–206, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003003338.

Lebreton:2005:ASR

[Leb05] J. D. Lebreton. Age, stages, and the role of generation time in matrix models. Ecological Modelling, 188(1):22–29, October 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002358.

Legovic:2001:ISE

[Leg01] Tarzan Legovic. ISEM: the Second European Ecological Modelling Conference, 20–24 September 1999, Pula, Croatia. Ecological Modelling, 138(1–3):1, March 15, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001002885.

Legovic:2003:D

[Leg03a] Tarzan Legovic. Dedication. Ecological Modelling, 170(2–3):479, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002473.

Legovic:2003:PSQ

[Leg03b] Tarzan Legović. Prediction of seawater quality around Island Rab (Adriatic Sea). Ecological Modelling, 160 (1–2):131–143, February 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002003253.

Legovic:2008:IDF

[Leg08] Tarzan Legović. Impact of demersal fishery and evidence of the Volterra principle to the extreme in the Adriatic Sea. Ecological Modelling, 212(1–2):68–73, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005091.

Leathwick:2006:CPG

[LEH06] J. R. Leathwick, J. Elith, and T. Hastie. Comparative performance of generalized additive models and multivariate adaptive regression splines for statistical modelling of species distributions. Ecological Modelling, 199(2):188–196, November 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006002572.

Lek:2007:UEM

[Lek07] Sovan Lek. Uncertainty in ecological models. Ecological Modelling, 207(1):1–2, September 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001494.

Lewis:2002:SAR

[LELR02] D. M. Lewis, J. A. Elliott, M. F. Lambert, and C. S. Reynolds. The simulation of an Australian reservoir using a phytoplankton community model: PROTECH. Ecological Modelling, 150(1–2):107–116, April 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004665.

Lenzen:2007:SPA

[Len07] Manfred Lenzen. Structural path analysis of ecosystem networks. Ecological Modelling, 200(3–4):334–342, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003693.

Levich:2000:VMT

[Lev00] A. P. Levich. Variational modelling theorems and algocoenoses functioning principles. Ecological Modelling, 131(2–3):207–227, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002593.

Lindenschmidt:2007:SUR

[LFB07] Karl-Erich Lindenschmidt, Katrin Fleischbein, and Martina Baborowski. Structural uncertainty in a river water quality modelling system. Ecological Modelling, 204(3–4):289–300, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007000208.

Li:2006:MNL

[LFJ^{+}06] Changsheng Li, Neda Farahbakhshazad, Dan B. Jaynes, Dana L. Dinnes, William Salas, and Dennis McLaughlin. Modeling nitrate leaching with a biogeochemical model modified based on observations in a row-crop field in Iowa. Ecological Modelling, 196(1–2):116–130, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000664.

Li:2000:SAE

[LG00] Dianmo Li and Zhongwei Guo. Some aspects of ecological modeling developments in China. Ecological Modelling, 132(1–2):3–10, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003008.

Leung:2003:SSM

[LG03] Brian Leung and Bryan T. Grenfell. A spatial stochastic model simulating a scabies epidemic and coyote population dynamics. Ecological Modelling, 166(1–2):41–52, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001170.

Lemos:2004:DLE

[LG04] Ricardo T. Lemos and João F. Gomes. Do local environmental factors induce daily and yearly variability in bluefin tuna (Thunnus thynnus) trap catches? Ecological Modelling, 177(1–2):143–156, September 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001139.

Labiod:2007:RBS

[LGC07] C. Labiod, R. Godillot, and B. Caussade. The relationship between stream periphyton dynamics and near-bed turbulence in rough open-channel flow. Ecological Modelling, 209(2–4):78–96, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003158.

Lusk:2001:NBC

[LGD01] Jeffrey J. Lusk, Fred S. Guthery, and Stephen J. DeMaso. Northern bobwhite (Colinus virginianus) abundance in relation to yearly weather and long-term climate patterns. Ecological Modelling, 146(1–3):3–15, December 1, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001002927.

Loot:2002:NDM

[LGL02] Géraldine Loot, Jean-Luc Giraudel, and Sovan Lek. A nondestructive morphometric technique to predict Ligula intestinalis L. plerocercoid load in roach (Rutilus rutilus L.) abdominal cavity. Ecological Modelling, 156(1):1–11, October 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002000303.

Liang:2009:SSO

[LGR^{+}09] Y. Liang, H. T. Gollany, R. W. Rickman, S. L. Albrecht, R. F. Follett, W. W. Wilhelm, J. M. Novak, and C. L. Douglas. Simulating soil organic matter with CQESTR (v. 2.0): Model description and validation against longterm experiments across North America. Ecological Modelling, 220(4):568–581, February 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005449.

Lopez:2000:RWT

[LGS^{+}00] Roel R. Lopez, William E. Grant, Nova J. Silvy, Markus J. Peterson, Charles K. Feuerbacher, and Michael S. Corson. Restoration of the wild turkey in east Texas: simulation of alternative restocking strategies. Ecological Modelling, 132(3):275–285, August 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002957.

Lundstrom-Gillieron:2003:HAI

[LGS03] Corinne Lundström-Gilliéron and Rodolphe Schlaepfer. Hare abundance as an indicator for urbanisation and intensification of agriculture in Western Europe. Ecological Modelling, 168(3):283–301, October 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300142X.

Lindstrom:2001:MCH

[LH01] Martin Lindström and Lars Håkanson. A model to calculate heavy metal load to lakes dominated by urban runoff and diffuse inflow. Ecological Modelling, 137(1):1–21, February 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004403.

Lexer:2004:EEM

[LH04] Manfred J. Lexer and Karl Hönninger. Effects of error in model input: experiments with a forest patch model. Ecological Modelling, 173(2–3):159–176, April 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003922.

LeCorff:2005:PGV

[LH05] Josiane Le Corff and Carol C. Horvitz. Population growth versus population spread of an ant-dispersed neotropical herb with a mixed reproductive strategy. Ecological Modelling, 188(1):41–51, October 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002383.

LeCorff:2006:EPG

[LH06] Josiane Le Corff and Carol C. Horvitz. Erratum to “Population growth versus population spread of an antdispersed neotropical herb with a mixed reproductive strategy” [Ecol. Model. 188 (1) (2005) 41–51]. Ecological Modelling, 195(3–4):409, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600041X.

Lawrie:2007:RMC

[LH07] Jock Lawrie and John Hearne. Reducing model complexity via output sensitivity. Ecological Modelling, 207 (2–4):137–144, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007002372.

Liu:2009:LTF

[LH09] Guoliang Liu and Shijie Han. Long-term forest management and timely transfer of carbon into wood products help reduce atmospheric carbon. Ecological Modelling, 220(13–14):1719–1723, July 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002622.

Lemasson:2008:EAB

[LHB08] Bertrand H. Lemasson, James W. Haefner, and Mark D. Bowen. The effect of avoidance behavior on predicting fish passage rates through water diversion structures. Ecological Modelling, 219(1–2):178–188, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004110.

Larssen:2007:ITS

[LHC07] Thorjørn Larssen, Tore Høgåsen, and B. Jack Cosby. Impact of time series data on calibration and prediction uncertainty for a deterministic hydrogeochemical model. Ecological Modelling, 207(1):22–33, September 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001524.

Lee:2009:SSD

[LHC09] Chun-Lin Lee, Shu-Li Huang, and Shih-Liang Chan. Synthesis and spatial dynamics of socio-economic metabolism and land use change of Taipei metropolitan region. Ecological Modelling, 220(21):2940–2959, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003895.

Lee:2003:NNM

[LHDJ03] Joseph H. W. Lee, Yan Huang, Mike Dickman, and A. W. Jayawardena. Neural network modelling of coastal algal blooms. Ecological Modelling, 159(2–3):179–201, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002818.

Levy:2000:UEI

[LHK00] Jason K. Levy, Keith W. Hipel, and D. Marc Kilgour. Using environmental indicators to quantify the robustness of policy alternatives to uncertainty. Ecological Modelling, 130(1–3):79–86, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000226X.

Leitinger:2008:DVS

[LHT^{+}08] Georg Leitinger, Peter Höller, Erich Tasser, Janette Walde, and Ulrike Tappeiner. Development and validation of a spatial snow-glide model. Ecological Modelling, 211 (3–4):363–374, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004796.

Liu:2006:OAS

[LHZ^{+}06] Guoliang Liu, Shijie Han, Xiuhai Zhao, John D. Nelson, Hongshu Wang, and Weiying Wang. Optimisation algorithms for spatially constrained forest planning. Ecological Modelling, 194(4):421–428, April 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004916.

Li:2000:FGA

[Li00a] Bai-Lian Li. Fractal geometry applications in description and analysis of patch patterns and patch dynamics. Ecological Modelling, 132(1–2):33–50, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003033.

Li:2000:P

[Li00b] Bai-Lian Li. Preface. Ecological Modelling, 132(1–2):1, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002994.

Li:2000:RNF

[Li00c] Chao Li. Reconstruction of natural fire regimes through ecological modelling. Ecological Modelling, 134(2–3):129– 144, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000002908.

Li:2002:EFF

[Li02] Chao Li. Estimation of fire frequency and fire cycle: a computational perspective. Ecological Modelling, 154 (1–2):103–120, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000698.

Liddel:2001:SSC

[Lid01] Michael Liddel. A simple space competition model using stochastic and episodic disturbance. Ecological Modelling, 143(1–2):33–41, October 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003532.

Limburg:2004:BSR

[Lim04] Karin E. Limburg. The biogeochemistry of strontium: a review of H. T. Odum’s contributions. Ecological Modelling, 178(1–2):31–33, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005350.

Lindstrom:2001:DPR

[Lin01] Martin Lindström. Distribution of particulate and reactive mercury in surface waters of Swedish forest lakes — an empirically based predictive model. Ecological Modelling, 136(1):81–93, January 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003823.

Lin:2003:SUE

[Lin03] Zhen-Shan Lin. Simulating unintended effects restoration. Ecological Modelling, 164(2–3):169–175, June 15, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003000140.

Lin:2005:EOP

[Lin05]  Zhen-Shan Lin.  The ecological order of persisting

species during habitat destruction. Ecological Modelling, 184(2–4):249–256, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004005502.

Lindenschmidt:2006:ECP

[Lin06] Karl-Erich Lindenschmidt. The effect of complexity on parameter sensitivity and model uncertainty in river water quality modelling. Ecological Modelling, 190 (1–2):72–86, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002218.

Lirman:2003:SMP

[Lir03] Diego Lirman. A simulation model of the population dynamics of the branching coral Acropora palmata effects of storm intensity and frequency. Ecological Modelling, 161(3):169–182, March 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002003460.

Lamers:2007:MEF

[LIS07] Marc Lamers, Joachim Ingwersen, and Thilo Streck. Modelling N_{2}O emission from a forest upland soil: a procedure for an automatic calibration of the biogeochemical model Forest-DNDC. Ecological Modelling, 205(1–2):52–58, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007000592.

Liu:2001:IEH

[Liu01] Jianguo Liu. Integrating ecology with human demography, behavior, and socioeconomics: Needs and approaches. Ecological Modelling, 140(1–2):1–8, May 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002654.

Li:2000:STR

[LiWZ00] Bai-Lian Li, Hsin i Wu, and Guangzhou Zou. Self-thinning rule: a causal interpretation from ecological field theory. Ecological Modelling, 132(1–2):167–173, July 30, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000003136.

Laird:2007:CSE

[LJ07] Simon Laird and Henrik Jeldtoft Jensen. Correlation, selection and the evolution of species networks. Ecological Modelling, 209(2–4):149–156, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003249.

Ludovisi:2009:CEF

[LJ09] Alessandro Ludovisi and Sven Erik Jørgensen. Comparison of exergy found by a classical thermodynamic approach and by the use of the information stored in the genome. Ecological Modelling, 220(16):1897–1903, August 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002725.

Lassueur:2006:VHR

[LJR06] Thierry Lassueur, Stéphane Joost, and Christophe F. Randin. Very high resolution digital elevation models: Do they improve models of plant species distribution? Ecological Modelling, 198(1–2):139–153, September 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006001712.

Logan:2003:CRM

[LJW03] J. David Logan, Anthony Joern, and William Wolesensky. Chemical reactor models of optimal digestion efficiency with constant foraging costs. Ecological Modelling, 168(1–2):25–38, October 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002023.

Logofet:2002:HOH

[LK02] Dmitrii O. Logofet and Vladimir N. Korotkov. ‘hybrid’ optimisation: a heuristic solution to the Markov-chain calibration problem. Ecological Modelling, 151(1):51–61, May 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004860.

Linacre:2003:DEF

[LK03] Nicholas A. Linacre and Michael J. Keough. Demographic effects of fragmentation history in modular organisms: illustrated using the bryozoan Mucropetraliella ellerii (MacGillivray). Ecological Modelling, 170(1):61–71, December 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000300303X.

Lehsten:2007:TPT

[LK07] Veiko Lehsten and Michael Kleyer. Turnover of plant trait hierarchies in simulated community assembly in response to fertility and disturbance. Ecological Modelling, 203(3–4):270–278, May 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006028.

LeMaitre:2008:HSD

[LKH^{+}08] David C. Le Maitre, Rainer M. Krug, John H. Hoffmann, Anthony J. Gordon, and Theresa N. Mgidi. Hakea sericea: Development of a model of the impacts of biological control on population dynamics and rates of spread of an invasive species. Ecological Modelling, 212(3–4):342–358, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007006059.

Lee:2009:NFI

[LKKL09] Eun Jung Lee, Meejeong Kim, Young Kim, and Kang-Kun Lee. Numerical and field investigation of enhanced in situ denitrification in a shallow-zone well-to-well recirculation system. Ecological Modelling, 220(19):2441–2449, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009003913.

Lee:2007:CBB

[LKLK07] Jangho Lee, Inn-Sil Kwak, Eungkyoung Lee, and Kyung A. Kim. Classification of breeding bird communities along an urbanization gradient using an unsupervised artificial neural network. Ecological Modelling, 203(1–2):62–71, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006005588.

Leskinen:2003:AEV

[LKP03] Pekka Leskinen, Jyrki Kangas, and Anna-Mari Pasanen. Assessing ecological values with dependent explanatory variables in multi-criteria forest ecosystem management. Ecological Modelling, 170(1):1–12, December 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002837.

Laakso:2003:NLB

[LKR03] Jouni Laakso, Veijo Kaitala, and Esa Ranta. Non-linear biological responses to disturbance: consequences on population dynamics. Ecological Modelling, 162(3):247–258, April 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000200385X.

Laakso:2006:PDC

[LKR06] Jouni Laakso, Veijo Kaitala, and Esa Ranta. Population dynamic consequences of adaptive growth rate in fluctuating environment. Ecological Modelling, 194(1–3):132–140, March 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005005119.

Lee:2006:MAS

[LKY^{+}06] Who-Seung Lee, Young-Soo Kwon, Jeong-Chil Yoo, Mi-Young Song, and Tae-Soo Chon. Multivariate analysis and self-organizing mapping applied to analysis of nest-site selection in black-tailed gulls. Ecological Modelling, 193(3–4):602–614, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004424.

Logofet:2000:MMM

[LL00] Dmitrii O. Logofet and Ekaterina V. Lesnaya. The mathematics of Markov models: what Markov chains can really predict in forest successions. Ecological Modelling, 126(2–3):285–298, February 28, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002696.

Langeveld:2002:MBP

[LL02a] Cor A. Langeveld and Peter A. Leffelaar. Modelling belowground processes to explain field-scale emissions of nitrous oxide. Ecological Modelling, 149(1–2):97–112, March 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001005178.

Lin:2002:MSY

[LL02b] Zhen-Shan Lin and Bai-Lian Li. The maximum sustainable yield of Allee dynamic system. Ecological Modelling, 154(1–2):1–7, August 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004793.

Lin:2003:TSL

[LL03a] Zhen-Shan Lin and Bai-Lian Li. A twin-scale law about the minimum cost. Ecological Modelling, 165(2–3):231–235, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003000760.

Lu:2003:MEH

[LL03b] Fan Lu and Zizhen Li. A model of ecosystem health and its application. Ecological Modelling, 170(1):55–59, December 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003003004.

Lischke:2006:ISD

[LL06] Heike Lischke and Thomas J. Löffler. Intra-specific density dependence is required to maintain species diversity in spatio-temporal forest simulations with reproduction. Ecological Modelling, 198(3–4):341–361, October 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006002377.

Lamour:2007:ITD

[LL07a] Angélique Lamour and Lambertus A. P. Lotz. The importance of tillage depth in relation to seedling emergence in stale seedbeds. Ecological Modelling, 201 (3–4):536–546, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006005266.

Lasaponara:2007:CSV

[LL07b] Rosa Lasaponara and Antonio Lanorte. On the capability of satellite VHR QuickBird data for fuel type characterization in fragmented landscape. Ecological Modelling, 204(1–2):79–84, May 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006454.

Law:2007:OSR

[LL07c] Peter R. Law and Wayne L. Linklater. Optimising the sex ratio of translocation for genetic rescue as a function of invested resources. Ecological Modelling, 208(2– 4):317–341, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007003110.

Liu:2008:SSF

[LL08] Xuehua Liu and Jihong Li. Scientific solutions for the functional zoning of nature reserves in China. Ecological Modelling, 215(1–3):237–246, July 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001038.

LeBer:2009:NMA

[LLA^{+}09a] F. Le Ber, C. Lavigne, K. Adamczyk, F. Angevin, N. Colbach, J.-F. Mari, and H. Monod. Neutral modelling of agricultural landscapes by tessellation methods — application for gene flow simulation. Ecological Modelling, 220(24):3536–3545, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004116.

Liu:2009:MTI

[LLA^{+}09b] Ying Hu Liu, De Li Liu, Min An, Yin Lian Fu, Ren Sen Zeng, Shi Ming Luo, Hanwen Wu, and Jim Pratley. Modelling tritrophic interactions mediated by induced defence volatiles. Ecological Modelling, 220(23):3241–3247, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009004827.

Lin:2004:VRA

[LLCL04] Chao-Yuan Lin, Huang-Mu Lo, Wen-Chieh Chou, and Wen-Tzu Lin. Vegetation recovery assessment at the Jou-Jou Mountain landslide area caused by the 921 Earthquake in Central Taiwan. Ecological Modelling, 176(1–2):75–81, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000300557X.

Lischke:2002:UHP

[LLF02] Heike Lischke, André F. Lotter, and Andreas Fischlin. Untangling a Holocene pollen record with forest model simulations and independent climate data. Ecological Modelling, 150(1–2):1–21, April 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000133.

Lee:2006:NTT

[LLH^{+}06] Mee-Sun Lee, Kang-Kun Lee, Yunjung Hyun, T. Prabhakar Clement, and David Hamilton. Nitrogen transformation and transport modeling in groundwater aquifers. Ecological Modelling, 192(1–2):143–159, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003418.

Lui:2007:MAB

[LLL^{+}07] Gilbert C. S. Lui, W. K. Li, Kenneth M. Y. Leung, Joseph H. W. Lee, and A. W. Jayawardena. Modelling algal blooms using vector autoregressive model with exogenous variables and long memory filter. Ecological Modelling, 200(1–2):130–138, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002961.

Leibowitz:2000:MLF

[LLLP00] Scott G. Leibowitz, Craig Loehle, Bai-Lian Li, and Eric M. Preston. Modeling landscape functions and effects: a network approach. Ecological Modelling, 132(1–2):77–94, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003069.

Liu:2008:IRS

[LLLT08] Jinxun Liu, Shuguang Liu, Thomas R. Loveland, and Larry L. Tieszen. Integrating remotely sensed land cover observations and a biogeochemical model for estimating forest ecosystem carbon dynamics. Ecological Modelling, 219(3–4):361–372, December 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003566.

Liu:2008:SCU

[LLS^{+}08] Xiaoping Liu, Xia Li, Xun Shi, Shaokun Wu, and Tao Liu. Simulating complex urban development using kernelbased non-linear cellular automata. Ecological Modelling, 211(1–2):169–181, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004449.

Lacerte:2006:CFV

[LLW^{+}06] V. Lacerte, G. R. Larocque, M. Woods, W. J. Parton, and M. Penner. Calibration of the forest vegetation simulator (FVS) model for the main forest species of Ontario, Canada. Ecological Modelling, 199(3):336–349, December 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006002341.

Liu:2008:EPC

[LLZ^{+}08] HuiYu Liu, JiaJia Lin, Mingyang Zhang, ZhenShan Lin, and Ten Wen. Extinction of poorest competitors and temporal heterogeneity of habitat destruction. Ecological Modelling, 219(1–2):30–38, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002949.

Little:2007:SAS

[LM07a] L. R. Little and A. D. McDonald. Simulations of agents in social networks harvesting a resource. Ecological Modelling, 204(3–4):379–386, June 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000294.

Long:2007:AES

[LM07b] Lin Hui Long and Zhuang Qi Ming. Adaptive extent and seed yield predictions for Microula sikkimensis grown in the Qinghai-Tibet Plateau, China. Ecological Modelling, 201(3–4):507–520, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005278.

Ling:2008:DAE

[LMG08] Stephen Ling and E. J. Milner-Gulland. Developing an artificial ecology for use as a strategic management tool: a case study of ibex hunting in the North Tien Shan. Ecological Modelling, 210(1–2):15–36, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003535.

Langmead:2009:RDN

[LMGM^{+}09] Olivia Langmead, Abigail McQuatters-Gollop, Laurence D. Mee, Jana Friedrich, Alison J. Gilbert, Marian-Traian Gomoiu, Emma L. Jackson, Ståle Knudsen, Galina Minicheva, and Valentina Todorova. Recovery or decline of the northwestern Black Sea: a societal choice revealed by socio-ecological modelling. Ecological Modelling, 220 (21):2927–2939, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008004420.

Lafarge:2005:MTC

[LMH05] Michel Lafarge, Claude Mazel, and David R. C. Hill. A modelling of the tillering capable of reproducing the finescale horizontal heterogeneity of a pure grass sward and its dynamics. Ecological Modelling, 183(1):125–141, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004004351.

Lencioni:2007:MAG

[LMM^{+}07] Valeria Lencioni, Bruno Maiolini, Laura Marziali, Sovan Lek, and Bruno Rossaro. Macroinvertebrate assemblages in glacial stream systems: a comparison of linear multivariate methods with artificial neural networks. Ecological Modelling, 203(1–2):119–131, April 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005631.

Leskinen:2009:MCS

[LMMK09] Pekka Leskinen, Jari Miina, Lauri Mehtätalo, and Annika Kangas. Model correlation in stochastic forest simulators — a case of multilevel multivariate model for seedling establishment. Ecological Modelling, 220(4):545–555, February 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000800536X.

Lenz:2000:IO

[LMPP00] Roman Lenz, Irina G. Malkina-Pykh, and Yuri Pykh. Introduction and overview. Ecological Modelling, 130(1–3):1–11, June 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000002258.

Larocque:2006:EMD

[LMPR06] Guy R. Larocque, Dave A. Mauriello, Richard A. Park, and Edward J. Rykiel. Ecological Models as Decision Tools in the 21st Century: Proceedings of a conference organized by the International Society for Ecological Modelling (ISEM) in Quebec, Canada, August 22–24, 2004. Ecological Modelling, 199(3):217–218, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002146.

Ludovisi:2005:MHS

[LMPT05] Alessandro Ludovisi, Marco Minozzo, Piera Pandolfi, and M. Illuminata Taticchi. Modelling the horizontal spatial structure of planktonic community in Lake Trasimeno (Umbria, Italy) using multivariate geostatistical methods. Ecological Modelling, 181(2–3):247–262, January 20, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004003709.

LaRue:2008:MPD

[LN08] Michelle A. LaRue and Clayton K. Nielsen. Modelling potential dispersal corridors for cougars in midwestern North America using least-cost path methods. Ecological Modelling, 212(3–4):372–381, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005595.

Liu:2000:TOA

[LNW00] Guoliang Liu, John D. Nelson, and Chad W. Wardman. A target-oriented approach to forest ecosystem design — changing the rules of forest planning. Ecological Modelling, 127(2–3):269–281, March 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002173.

Loehle:2000:OCS

[Loe00] Craig Loehle. Optimal control of spatially distributed process models. Ecological Modelling, 131(2–3):79–95, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000000274X.

Loehle:2004:CCD

[Loe04] Craig Loehle. Climate change: detection and attribution of trends from long-term geologic data. Ecological Modelling, 171(4):433–450, February 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003600.

Logofet:2002:MPM

[Log02] Dmitrii O. Logofet. Matrix Population Models: Construction, Analysis, and Interpretation: Hal Caswell, 2nd ed., Sinauer Associates, Sunderland, Masssachusetts, 2001, xxii + 722 pp. ISBN 0-87893-096-5; $64.95. Ecological Modelling, 148(3):307–310, March 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000029.

Logofet:2008:CPM

[Log08] Dmitrii O. Logofet. Convexity in projection matrices: Projection to a calibration problem. Ecological Modelling, 216(2):217–228, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008001439.

Lehmann:2002:GGR

[LOL02] Anthony Lehmann, Jacob McC. Overton, and John R. Leathwick. GRASP: generalized regression analysis and spatial prediction. Ecological Modelling, 157(2– 3):189–207, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001953.

Lehmann:2003:EGG

[LOL03a] Anthony Lehmann, Jacob McC. Overton, and John R. Leathwick. Erratum to “GRASP: generalized regression analysis and spatial prediction”: [Ecological Modelling 157 (2002) 189–207]. Ecological Modelling, 160(1–2):163, February 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002003526.

Lehmann:2003:GGR

[LOL03b] Anthony Lehmann, Jacob McC. Overton, and John R. Leathwick. GRASP: generalized regression analysis and spatial prediction. Ecological Modelling, 160(1– 2):165–183, February 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000200354X.

Lobell:2006:RIC

[LOM06] David B. Lobell and J. Ivan Ortiz-Monasterio. Regional importance of crop yield constraints: Linking simulation models and geostatistics to interpret spatial patterns. Ecological Modelling, 196(1–2):173–182, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006162.

Ludovisi:2003:UTIa

[LP03a] Alessandro Ludovisi and Antonio Poletti. Use of thermodynamic indices as ecological indicators of the development state of lake ecosystems. 1. Entropy production indices. Ecological Modelling, 159(2–3):203–222, January 15, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002002776.

Ludovisi:2003:UTIb

[LP03b] Alessandro Ludovisi and Antonio Poletti. Use of thermodynamic indices as ecological indicators of the development state of lake ecosystems: 2. Exergy and specific exergy indices. Ecological Modelling, 159(2–3):223–238, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002843.

Liu:2005:NCE

[LPC05] Jinxun Liu, David T. Price, and Jing M. Chen. Nitrogen controls on ecosystem carbon sequestration: a model implementation and application to Saskatchewan, Canada. Ecological Modelling, 186(2):178–195, August 15, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000256.

Lopes:2005:EPD

[LPCC05] C. Lopes, A. R. R. Péry, A. Chaumot, and S. Charles. Ecotoxicology and population dynamics: Using DEBtox models in a Leslie modeling approach. Ecological Modelling, 188(1):30–40, October 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000500236X.

Lacitignola:2007:MSE

[LPCZ07] D. Lacitignola, I. Petrosillo, M. Cataldi, and G. Zurlini. Modelling socio-ecological tourism-based systems for sustainability. Ecological Modelling, 206(1–2):191–204, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001809.

Laganis:2008:MRG

[LPD08] Jana Laganis, Aleksandar Pečkov, and Marko Debeljak. Modeling radial growth increment of black alder (Alnus glutionsa (L.) Gaertn.) tree. Ecological Modelling, 215(1–3):180–189, July 2008. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008000938.

Linard:2009:MAS

[LPFL09] Catherine Linard, Nicolas Ponçon, Didier Fontenille, and Eric F. Lambin. A multi-agent simulation to assess the risk of malaria re-emergence in southern France. Ecological Modelling, 220(2):160–174, January 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004304.

Little:2007:EME

[LPM^{+}07] L. R. Little, A. E. Punt, B. D. Mapstone, F. Pantus, A. D. M. Smith, C. R. Davies, and A. D. McDonald. ELFSim — a model for evaluating management options for spatially structured reef fish populations: an illustration of the “larval subsidy” effect. Ecological Modelling, 205(3–4):381–396, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007001238.

Little:2009:ABM

[LPM^{+}09] L. Richard Little, André E. Punt, Bruce D. Mapstone, Gavin A. Begg, Barry Goldman, and Ashley J. Williams. An agent-based model for simulating trading of multispecies fisheries quota. Ecological Modelling, 220(23): 3404–3412, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009005353.

Lee:2007:REF

[LPP^{+}07] Sang Dong Lee, Sohyun Park, Young-Seuk Park, Yeong-Jin Chung, Buom-Young Lee, and Tae-Soo Chon. Range expansion of forest pest populations by using the lattice model. Ecological Modelling, 203(1–2):157–166, April 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006005722.

Liski:2005:CDM

[LPPS05] Jari Liski, Taru Palosuo, Mikko Peltoniemi, and Risto Sievänen. Carbon and decomposition model Yasso for forest soils. Ecological Modelling, 189(1–2):168–182, November 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002012.

Larsson:2007:DPM

[LPU^{+}07] Martin H. Larsson, Kristian Persson, Barbro Ulén, Anders Lindsjö, and Nicholas J. Jarvis. A dual porosity model to quantify phosphorus losses from macroporous soils. Ecological Modelling, 205(1–2):123–134, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700083X.

Lin:2005:CBC

[LQL05] Zhen-Shan Lin, Xiang-Zhen Qi, and Bai-Lian Li. Can best competitors avoid extinction as habitat destruction? Ecological Modelling, 182(2):107–112, March 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001425.

Lefroy:2003:EET

[LR03] Edward Lefroy and Torbjörn Rydberg. Emergy evaluation of three cropping systems in southwestern Australia. Ecological Modelling, 161(3):195–211, March 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003411.

Lambert:2007:IIP

[LR07a] Patrick Lambert and Eric Rochard. Identification of the inland population dynamics of the European eel using pattern-oriented modelling. Ecological Modelling, 206 (1–2):166–178, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001779.

Lessin:2007:MSD

[LR07b] Gennadi Lessin and Urmas Raudsepp. Modelling the spatial distribution of phytoplankton and inorganic nitrogen in Narva Bay, southeastern Gulf of Finland, in the biologically active period. Ecological Modelling, 201 (3–4):348–358, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004649.

Lin:2009:DSC

[LR09] Jianyi Lin and Sergio Rinaldi. A derivation of the statistical characteristics of forest fires. Ecological Modelling, 220(7):898–903, April 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000350.

Laine:2009:ASH

[LRJ^{+}09] Anna Laine, Terhi Riutta, Sari Juutinen, Minna Väliranta, and Eeva-Stiina Tuittila. Acknowledging the spatial heterogeneity in modelling/reconstructing carbon dioxide exchange in a northern aapa mire. Ecological Modelling, 220(20):2646–2655, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004505.

Lippitt:2008:IAV

[LRT^{+}08] Christopher D. Lippitt, John Rogan, James Toledano, Florencia Sangermano, J. Ronald Eastman, Victor Mastro, and Alan Sawyer. Incorporating anthropogenic variables into a species distribution model to map gypsy moth risk. Ecological Modelling, 210(3):339–350, January 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004139.

Laterra:2001:DSS

[LS01] P. Laterra and O. T. Solbrig. Dispersal strategies, spatial heterogeneity and colonization success in fire-managed grasslands. Ecological Modelling, 139(1):17–29, March 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002277.

Latham:2002:QCA

[LS02a] Luke G. Latham and Erik P. Scully. Quantifying constraint to assess development in ecological networks. Ecological Modelling, 154(1–2):25–44, August 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000327.

Liu:2002:FCA

[LS02b] Mingqin Liu and Ashok Samal. A fuzzy clustering approach to delineate agroecozones. Ecological Modelling, 149(3):215–228, April 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100446X.

Langmead:2004:CRC

[LS04a] O. Langmead and C. Sheppard. Coral reef community dynamics and disturbance: a simulation model. Ecological Modelling, 175(3):271–290, July 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004812.

Latham:2004:NOM

[LS04b] Luke G. Latham and Erik P. Scully. Network optimization model implies strength of average mutual information in ascendency. Ecological Modelling, 179(3):373–392, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002686.

Lopes:2006:TSD

[LS06] José Fortes Lopes and Catarina Silva. Temporal and spatial distribution of dissolved oxygen in the Ria de Aveiro lagoon. Ecological Modelling, 197(1–2):67–88, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006001311.

Li:2008:MDN

[LS08a] Yimin Li and Xihao Sun. Modelling dynamic niche and community model by type-2 fuzzy set. Ecological Modelling, 211(3–4):375–382, March 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004802.

Libralato:2008:BGM

[LS08b] Simone Libralato and Cosimo Solidoro. A bioenergetic growth model for comparing Sparus aurata ’s feeding experiments. Ecological Modelling, 214(2–4):325–337, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001087.

LeFur:2009:NHC

[LS09a] Jean Le Fur and Patrick Simon. A new hypothesis concerning the nature of small pelagic fish clusters: an individualbased modelling study of Sardinella aurita dynamics off West Africa. Ecological Modelling, 220(9–10):1291–1304, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009001240.

Libralato:2009:BBF

[LS09b] Simone Libralato and Cosimo Solidoro. Bridging biogeochemical and food web models for an end-to-end representation of marine ecosystem dynamics: the Venice lagoon case study. Ecological Modelling, 220(21):2960–2971, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009005791.

Lopez-Sanchez:2005:IPS

[LSAGF05] J. F. López-Sánchez, F. Alhama, and C. F. González-Fernández. Introduction and permanence of species in a diffusive Lotka–Volterra system with time-dependent coefficients. Ecological Modelling, 183(1):1–9, April 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004004144.

Li:2008:MRS

[LSHG08] Yimin Li, Xihao Sun, Jing Hua, and Chongying Gong. Modelling redundant structure in ecosystem by type-2 fuzzy logic system. Ecological Modelling, 211(1–2):113–120, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007004292.

Lehtonen:2004:PLS

[LSM^{+}04] Aleksi Lehtonen, Risto Sievänen, Annikki Mäkelä, Raisa Mäkipää, Kari T. Korhonen, and Tatu Hokkanen. Potential litterfall of Scots pine branches in southern Finland. Ecological Modelling, 180(2–3):305–315, December 25, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004002844.

Lin:2000:MGB

[LSS^{+}00] Bin-Le Lin, Akiyoshi Sakoda, Ryosuke Shibasaki, Naohiro Goto, and Motoyuki Suzuki. Modelling a global biogeochemical nitrogen cycle in terrestrial ecosystems. Ecological Modelling, 135(1):89–110, November 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003720.

Li:2009:FAH

[LSY^{+}09] Lu Li, Zhi-Hua Shi, Wei Yin, Dun Zhu, Sai Leung Ng, Chong-Fa Cai, and A-Lin Lei. A fuzzy analytic hierarchy process (FAHP) approach to eco-environmental vulnerability assessment for the Danjiangkou Reservoir area, China. Ecological Modelling, 220(23):3439–3447, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006139.

Lonin:2001:WQM

[LT01] Serguei A. Lonin and Yuri S. Tuchkovenko. Water quality modelling for the ecosystem of the Ciénaga de Tesca coastal lagoon. Ecological Modelling, 144(2–3):279–293, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003763.

Linacre:2004:DIR

[LT04] Nicholas A. Linacre and Colin J. Thompson. Dynamics of insect resistance in Bt-corn. Ecological Modelling, 171(3):271–278, January 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003399.

Ludovisi:2006:IBD

[LT06] Alessandro Ludovisi and Maria Illuminata Taticchi. Investigating beta diversity by Kullback–Leibler information measures. Ecological Modelling, 192(1–2):299–313, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002929.

Lin:2008:ARP

[LTLH08] Wen-Tzu Lin, Jing-Shyan Tsai, Chao-Yuan Lin, and Pi-Hui Huang. Assessing reforestation placement and benefit for erosion control: a case study on the Chi-Jia-Wan Stream, Taiwan. Ecological Modelling, 211(3–4):444–452, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007004991.

Larson:2004:LPV

[LTM^{+}04] Michael A. Larson, Frank R. Thompson, Joshua J. Millspaugh, William D. Dijak, and Stephen R. Shifley. Linking population viability, habitat suitability, and landscape simulation models for conservation planning. Ecological Modelling, 180(1):103–118, December 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003771.

Libralato:2006:EEI

[LTP06] Simone Libralato, Patrizia Torricelli, and Fabio Pranovi. Exergy as ecosystem indicator: an application to the recovery process of marine benthic communities. Ecological Modelling, 192(3–4):571–585, February 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003960.

Ludovisi:2004:BAE

[Lud04] Alessandro Ludovisi. Biotic and abiotic entropy production in lake ecosystems. Ecological Modelling, 179 (1):145–147, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002078.

Ludovisi:2009:EVI

[Lud09] A. Ludovisi. Exergy vs. information in ecological successions: Interpreting community changes by a classical thermodynamic approach. Ecological Modelling, 220 (13–14):1566–1577, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009002580.

Lugo:2004:HOL

[Lug04] Ariel E. Lugo. H. T. Odum and the Luquillo Experimental Forest. Ecological Modelling, 178(1–2):65–74, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005386.

Logofet:2006:SDC

[LUKD06] Dmitrii O. Logofet, Nina G. Ulanova, Iya N. Klochkova, and Anna N. Demidova. Structure and dynamics of a clonal plant population: Classical model results in a non-classic formulation. Ecological Modelling, 192(1–2):95–106, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005003947.

Lee:2004:DGM

[LvGC^{+}04] Woo-Kyun Lee, Klaus von Gadow, Dong-Jun Chung, Jong-Lak Lee, and Man-Yong Shin. DBH growth model for Pinus densiflora and Quercus variabilis mixed forests in central Korea. Ecological Modelling, 176(1–2):187–200, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004000225.

Lammens:2004:ECF

[LvNMvdB04] Eddy H. R. R. Lammens, Egbert H. van Nes, Marie-Louise Meijer, and Marcel S. van den Berg. Effects of commercial fishery on the bream population and the expansion of Chara aspera in Lake Veluwe. Ecological Modelling, 177(3–4):233–244, October 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004575.

Liu:2009:SAS

[LW09] Meng Liu and Ke Wang. Survival analysis of stochastic single-species population models in polluted environments. Ecological Modelling, 220(9–10):1347–1357, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001550.

Lin:2005:ASS

[LWBW05] Hangsheng Lin, Dan Wheeler, Jay Bell, and Larry Wilding. Assessment of soil spatial variability at multiple scales. Ecological Modelling, 182(3–4):271–290, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001632.

Li:2007:SRW

[LWC^{+}07] Yong Li, Robert White, Deli Chen, Jiabao Zhang, Baoguo Li, Yuming Zhang, Yuanfang Huang, and Robert Edis. A spatially referenced water and nitrogen management model (WNMM) for (irrigated) intensive cropping systems in the North China plain. Ecological Modelling, 203(3–4):395–423, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006006296.

Logan:2006:TDP

[LWJ06] J. David Logan, William Wolesensky, and Anthony Joern. Temperature-dependent phenology and predation in arthropod systems. Ecological Modelling, 196(3–4):471–482, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000913.

Louie:2002:DMG

[LWL^{+}02] Ken Louie, Graeme C. Wake, Greg Lambert, Alec MacKay, and David Barker. A delay model for the growth of ryegrass–clover mixtures: formulation and preliminary simulations. Ecological Modelling, 155(1):31–42, September 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002000637.

Li:2006:EEV

[LWLZ06] Ainong Li, Angsheng Wang, Shunlin Liang, and Wancun Zhou. Eco-environmental vulnerability evaluation in mountainous region using remote sensing and GIS — a case study in the upper reaches of Minjiang River, China. Ecological Modelling, 192(1–2):175–187, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003388.

Li:2003:SMN

[LXJ^{+}03] Xiuzhen Li, Duning Xiao, Rob H. Jongman, W. Bert Harms, and Arnold K. Bregt. Spatial modeling on the nutrient retention of an estuary wetland. Ecological Modelling, 167(1–2):33–46, September 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001704.

Lin:2008:WQR

[LXP^{+}08] Jing Lin, Lian Xie, Leonard J. Pietrafesa, Hongzhou Xu, Wendy Woods, Michael A. Mallin, and Michael J. Durako. Water quality responses to simulated flow and nutrient reductions in the Cape Fear River Estuary and adjacent coastal region, North Carolina. Ecological Modelling, 212(3–4):200–217, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005297.

Xu:2002:IID

[lXzL02a] Cai lin Xu and Zi zhen Li. Influence of intraspecific density dependence on a three-species food chain with and without external stochastic disturbances. Ecological Modelling, 155(1):71–83, September 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000674.

Xu:2002:PRE

[lXzL02b] Cai lin Xu and Zi zhen Li. Population’s response to environmental noise: the influence of food web structure. Ecological Modelling, 154(1–2):193–202, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000510.

Xu:2002:SER

[lXzL02c] Cai lin Xu and Zi zhen Li. Stochastic ecosystem resilience and productivity: seeking a relationship. Ecological Modelling, 156(2–3):143–152, November 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200145X.

Liao:2008:PAS

[LYC08] Chung-Min Liao, Ching-Hung Yeh, and Szu-Chieh Chen. Predation affects the susceptibility of hard clam Meretrix lusori a to Hg-stressed birnavirus. Ecological Modelling, 210(3):253–262, January 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700405X.

Liu:2009:EBU

[LYCU09] G. Y. Liu, Z. F. Yang, B. Chen, and S. Ulgiati. Emergybased urban health evaluation and development pattern analysis. Ecological Modelling, 220(18):2291–2301, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009003834.

Liu:2007:MDP

[LZ07] Hanwu Liu and Li Zhou. Modeling dispersal of the plateau pika (Ochotona curzoniae) using a cellular automata model.

Ecological Modelling, 202(3–4):487–492, April 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006005497.

Lischke:2006:TFL

[LZB^{+}06] Heike Lischke, Niklaus E. Zimmermann, Janine Bolliger, Sophie Rickebusch, and Thomas J. Löffler. TreeMig: a forest-landscape model for simulating spatio-temporal patterns from stand to landscape scale. Ecological Modelling, 199(4):409–420, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002857.

Li:2007:IEE

[LZZ^{+}07] Zhong-Wu Li, Guang-Ming Zeng, Hua Zhang, Bin Yang, and Sheng Jiao. The integrated eco-environment assessment of the red soil hilly region based on GIS — a case study in Changsha City, China. Ecological Modelling, 202(3–4):540–546, April 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005448.

Law:2009:SCL

[LZZA09] Tony Law, Weitao Zhang, Jingyang Zhao, and George B. Arhonditsis. Structural changes in lake functioning induced from nutrient loading and climate variability. Ecological Modelling, 220(7):979–997, April 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000337.

Mette:2009:EFG

[MAA^{+}09] Tobias Mette, Axel Albrecht, Christian Ammer, Peter Biber, Ulrich Kohnle, and Hans Pretzsch. Evaluation of the forest growth simulator SILVA on dominant trees in mature mixed silver fir–Norway spruce stands in South-West Germany. Ecological Modelling, 220(13–14):1670–1680, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002026.

Morris:2001:SVM

[MAB01] Colin W. Morris, Arnaud Autret, and Lynne Boddy. Support vector machines for identifying organisms — a comparison with strongly partitioned radial basis function networks. Ecological Modelling, 146(1–3):57–67, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002964.

Mackinson:2000:AFE

[Mac00] Steven Mackinson. An adaptive fuzzy expert system for predicting structure, dynamics and distribution of herring shoals. Ecological Modelling, 126(2–3):155–178, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002635.

Mchich:2002:DFT

[MAdlPR02] R. Mchich, P. M. Auger, R. Bravo de la Parra, and N. Raissi. Dynamics of a fishery on two fishing zones with fish stock dependent migrations: aggregation and control. Ecological Modelling, 158(1–2):51–62, December 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002375.

Mpimpas:2001:MWP

[MAG01] Hercules Mpimpas, Petros Anagnostopoulos, and Jacques Ganoulis. Modelling of water pollution in the Thermaikos Gulf with fuzzy parameters. Ecological Modelling, 142(1–2):91–104, August 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002812.

Mouissie:2008:CPV

[MAHvD08] A. Maarten Mouissie, M. Emile F. Apol, Gerrit W. Heil, and Rudy van Diggelen. Creation and preservation of vegetation patterns by grazing. Ecological Modelling, 218(1–2):60–72, October 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003025.

Maina:2008:RLE

[Mai08] Joseph Maina. Reply to the letter to the editor by John Guinotte. Ecological Modelling, 218(3–4):403, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003840.

Musio:2004:PMC

[MAK^{+}04] Monica Musio, Nicole Augustin, Hans-Peter Kahle, Andreas Krall, Edgar Kublin, Rüdiger Unseld, and Klaus von Wilpert. Predicting magnesium concentration in needles of silver fir and Norway spruce — a case study. Ecological Modelling, 179(3):307–316, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001735.

Malacic:2001:NMI

[Mal01] Vlado Malačič. Numerical modelling of the initial spread of sewage from diffusers in the Bay of Piran (northern Adriatic). Ecological Modelling, 138(1–3):173–191, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004014.

Malanson:2003:DAC

[Mal03] George P. Malanson. Dispersal across continuous and binary representations of landscapes. Ecological Modelling, 169(1):17–24, November 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002047.

Mchich:2006:EAS

[MAL06] Rachid Mchich, Pierre Auger, and Christophe Lett. Effects of aggregative and solitary individual behaviors on the dynamics of predator–prey game models. Ecological Modelling, 197(3–4):281–289, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001475.

Mirabet:2007:SSS

[MAL07] Vincent Mirabet, Pierre Auger, and Christophe Lett. Spatial structures in simulations of animal grouping. Ecological Modelling, 201(3–4):468–476, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005102.

Mankin:2000:IAM

[Man00] K. R. Mankin. An integrated approach for modelling and managing golf course water quality and ecosystem diversity. Ecological Modelling, 133(3):259–267, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003331.

Martins:2006:ESL

[Mar06] M. L. Martins. Exact solution for the An–Liu–Johnson– Lovett model related to the dynamics of allelochemicals in the environment. Ecological Modelling, 193 (3–4):809–814, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004837.

Matter:2001:SED

[Mat01] Stephen F. Matter. Synchrony, extinction, and dynamics of spatially segregated, heterogeneous populations. Ecological Modelling, 141(1–3):217–226, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002757.

Matter:2003:MDA

[Mat03a] Stephen F. Matter. Modeling the density–area relationship in a dynamic landscape: an examination for the beetle Tetraopes tetraophthalmus and a generalized model. Ecological Modelling, 169(1):103–117, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002655.

Matthiopoulos:2003:USA

[Mat03b] Jason Matthiopoulos. The use of space by animals as a function of accessibility and preference. Ecological Modelling, 159(2–3):239–268, January 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002934.

Matthews:2006:PLM

[Mat06] Robin Matthews. The people and landscape model (PALM): Towards full integration of human decision-making and biophysical simulation models. Ecological Modelling, 194(4):329–343, April 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005005405.

Maud:2007:DSS

[Mau07] Sholto Maud. A DirectScience: the synthesis of ScienceDirect and EmSim. Ecological Modelling, 203(3–4):518–520, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006006053.

Moen:2000:TCA

[MB00] Aaron N. Moen and G. Scott Boomer. A tandem cosine algorithm for modeling rhythmic change. Ecological Modelling, 134(2–3):275–282, October 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003525.

Meunier:2002:ILM

[MB02a] Julien Meunier and Carlos Bernstein. The influence of local mate competition on host–parasitoid dynamics. Ecological Modelling, 152(1):77–88, June 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004914.

Mummery:2002:DIQ

[MB02b] D. Mummery and M. Battaglia. Data input quality and resolution effects on regional and local scale Eucalyptus globulus productivity predictions in north-east Tasmania. Ecological Modelling, 156(1):13–25, October 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200042X.

Middelhoff:2005:SFR

[MB05a] Ulrike Middelhoff and Broder Breckling. From single fine roots to a black alder forest ecosystem: How system behaviour emerges from single component activities. Ecological Modelling, 186(4):447–469, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001365.

Moen:2005:MAE

[MB05b] Aaron N. Moen and G. Scott Boomer. Modeling annual energy metabolism rhythms in mammals. Ecological Modelling, 184(2–4):193–202, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005460.

McArthur:2006:ECS

[MB06a] Lynne C. McArthur and John W. Boland. The economic contribution of seagrass to secondary production in South Australia. Ecological Modelling, 196(1–2):163–172, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000603.

Mukhopadhyay:2006:MPA

[MB06b] B. Mukhopadhyay and R. Bhattacharyya. Modelling phytoplankton allelopathy in a nutrient-plankton model with spatial heterogeneity. Ecological Modelling, 198(1– 2):163–173, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001748.

Mulder:2007:OSA

[MB07] Kenneth Mulder and William Breck Bowden. Organismal stoichiometry and the adaptive advantage of variable nutrient use and production efficiency in Daphnia. Ecological Modelling, 202(3–4):427–440, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600545X.

Ma:2008:SAB

[MB08] Zhanshan (Sam) Ma and Edward J. Bechinski. A survival-analysis-based simulation model for Russian wheat aphid population dynamics. Ecological Modelling, 216 (3–4):323–332, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002081.

Mukhopadhyay:2009:RPS

[MB09] B. Mukhopadhyay and R. Bhattacharyya. Role of predator switching in an eco-epidemiological model with disease in the prey. Ecological Modelling, 220(7):931–939, April 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009000398.

McCarthy:2007:MPF

[MBCS07] Annette M. McCarthy, Jerad D. Bales, W. Gregory Cope, and Damian Shea. Modeling pesticide fate in a small tidal estuary. Ecological Modelling, 200(1–2):149–159, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003176.

Mikhailova:2000:MSO

[MBD^{+}00] E. A. Mikhailova, R. B. Bryant, S. D. DeGloria, C. J. Post, and I. I. Vassenev. Modeling soil organic matter dynamics after conversion of native grassland to long-term continuous fallow using the CENTURY model. Ecological Modelling, 132(3):247–257, August 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002738.

Mari:2009:CEH

[MBDB09] Lorenzo Mari, Cristian Biotto, Astrid Decoene, and Luca Bonaventura. A coupled ecological–hydrodynamic model for the spatial distribution of sessile aquatic species in thermally forced basins. Ecological Modelling, 220(18): 2310–2324, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003366.

Mitchell:2009:MSP

[MBF09] Stephen Mitchell, Keith Beven, and Jim Freer. Multiple sources of predictive uncertainty in modeled estimates of net ecosystem CO_{2} exchange. Ecological Modelling, 220(23):3259–3270, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009006000.

Monga:2008:GSB

[MBGP08] Olivier Monga, Mamadou Bousso, Patricia Garnier, and Valérie Pot. 3D geometric structures and biological activity: Application to microbial soil organic matter decomposition in pore space. Ecological Modelling, 216(3– 4):291–302, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002044.

Manter:2003:MIF

[MBK^{+}03] Daniel K. Manter, Barbara J. Bond, Kathleen L. Kavanagh, Jeffrey K. Stone, and Gregory M. Filip. Modelling the impacts of the foliar pathogen, Phaeocryptopus gaeumannii, on Douglas-fir physiology: net canopy carbon assimilation, needle abscission and growth. Ecological Modelling, 164(2–3):211–226, June 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000267.

Mooij:2002:EED

[MBKD02] Wolf M. Mooij, Robert E. Bennetts, Wiley M. Kitchens, and Donald L. DeAngelis. Exploring the effect of drought extent and interval on the Florida snail kite: interplay between spatial and temporal scales. Ecological Modelling, 149(1–2):25–39, March 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005129.

Mathevet:2003:ABS

[MBLA03] Raphaël Mathevet, François Bousquet, Christophe Le Page, and Martine Antona. Agent-based simulations of interactions between duck population, farming decisions and leasing of hunting rights in the Camargue (Southern France). Ecological Modelling, 165(2–3):107–126, July 15, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000300098X.

Martens:2000:SDU

[MBM00] Scott N. Martens, David D. Breshears, and Clifton W. Meyer. Spatial distributions of understory light along the grassland/forest continuum: effects of cover, height, and spatial pattern of tree canopies. Ecological Modelling, 126(1):79–93, February 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009900188X.

Mazaris:2006:IBMb

[MBM06] Antonios D. Mazaris, Breckling Broder, and Yiannis G. Matsinos. An individual based model of a sea turtle population to analyze effects of age dependent mortality. Ecological Modelling, 198(1–2):174–182, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001736.

Malmaeus:2006:LPD

[MBMP06] J. M. Malmaeus, T. Blenckner, H. Markensten, and I. Persson. Lake phosphorus dynamics and climate warming: a mechanistic model approach. Ecological Modelling, 190(1–2):1–14, January 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002309.

Marthews:2008:SDT

[MBP^{+}08] T. R. Marthews, D. F. R. P. Burslem, S. R. Paton, F. Yangüez, and C. E. Mullins. Soil drying in a tropical forest: Three distinct environments controlled by gap size. Ecological Modelling, 216(3–4):369–384, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008002317.

Munier:2004:CEE

[MBPS04] B. Münier, K. Birr-Pedersen, and J. S. Schou. Combined ecological and economic modelling in agricultural land use scenarios. Ecological Modelling, 174(1–2):5–18, May 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005647.

Miehle:2009:CFP

[MBS^{+}09] Peter Miehle, Michael Battaglia, Peter J. Sands, David I. Forrester, Paul M. Feikema, Stephen J. Livesley, Jim D. Morris, and Stefan K. Arndt. A comparison of four processbased models and a statistical regression model to predict growth of Eucalyptus globulus plantations. Ecological Modelling, 220(5):734–746, March 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005875.

Matejicek:2003:EMN

[MBT03] Luboš Matějíček, Libuše Benešová, and Jaroslav Tonika. Ecological modelling of nitrate pollution in small river basins by spreadsheets and GIS. Ecological Modelling, 170(2–3):245–263, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002321.

Meixler:2009:PBP

[MBW09a] Marcia S. Meixler, Mark B. Bain, and M. Todd Walter. Predicting barrier passage and habitat suitability for migratory fish species. Ecological Modelling, 220(20):2782–2791, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000900502X.

Merico:2009:TBA

[MBW09b] Agostino Merico, Jorn Bruggeman, and Kai Wirtz. A trait-based approach for downscaling complexity in plankton ecosystem models. Ecological Modelling, 220 (21):3001–3010, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003275.

Mitchell:2001:ASU

[MC01] Scott W. Mitchell and Ferenc Csillag. Assessing the stability and uncertainty of predicted vegetation growth under climatic variability: northern mixed grass prairie. Ecological Modelling, 139(2–3):101–121, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002290.

Maud:2004:REH

[MC04] Sholto Maud and Dino Cevolatti. Realising the enlightenment: H. T. Odum’s energy systems language qua G. W. v Leibniz’s Characteristica Universalis. Ecological Modelling, 178(1–2):279–292, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005404.

Massada:2008:IOV

[MC08] Avi Bar Massada and Yohay Carmel. Incorporating output variance in local sensitivity analysis for stochastic models. Ecological Modelling, 213(3–4):463–467, May 24, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008000550.

Mchich:2006:OSD

[MCA^{+}06] Rachid Mchich, Najib Charouki, Pierre Auger, Nadia Raïssi, and Omar Ettahiri. Optimal spatial distribution of the fishing effort in a multi fishing zone model. Ecological Modelling, 197(3–4):274–280, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001463.

Monteiro:2006:BCS

[MCB06] L. H. A. Monteiro, H. D. B. Chimara, and J. G. Chaui Berlinck. Big cities: Shelters for contagious diseases. Ecological Modelling, 197(1–2):258–262, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001086.

Martin:2007:PSS

[MCBA07] Manuel Pascal Martin, Stéphane Cordier, Jérôme Balesdent, and Dominique Arrouays. Periodic solutions for soil carbon dynamics equilibriums with time-varying forcing variables. Ecological Modelling, 204(3–4):523–530, June 16, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006006703.

Martins:2008:SDV

[MCD^{+}08] Irene Martins, Ana Colaço, Paul R. Dando, Inês Martins, Daniel Desbruyères, Pierre-Marie Sarradin, João Carlos Marques, and Ricardo Serrão-Santos. Size-dependent variations on the nutritional pathway of Bathymodiolus azoricus demonstrated by a C-flux model. Ecological Modelling, 217(1–2):59–71, September 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002330.

Moniz:2007:AIT

[MCE^{+}07] L. J. Moniz, E. G. Cooch, S. P. Ellner, J. D. Nichols, and J. M. Nichols. Application of information theory methods to food web reconstruction. Ecological Modelling, 208(2–4):145–158, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002980.

McGill:2005:MMM

[McG05] Brian McGill. A mechanistic model of a mutualism and its ecological and evolutionary dynamics. Ecological Modelling, 187(4):413–425, October 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001195.

Menendez:2005:UFS

[MCGO05] I. Menéndez, J. Caniego, J. F. Gallardo, and K. Olechko. Use of fractal scaling to discriminate between and macroand meso-pore sizes in forest soils. Ecological Modelling, 182(3–4):323–335, March 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400167X.

Merritt:2004:BTA

[MCJ^{+}04] W. S. Merritt, B. F. W. Croke, A. J. Jakeman, R. A. Letcher, and P. Perez. A biophysical toolbox for assessment and management of land and water resources in rural catchments in Northern Thailand. Ecological Modelling, 171(3):279–300, January 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003405.

Mo:2008:OEM

[MCJB08] Xingguo Mo, Jing M. Chen, Weimin Ju, and T. Andrew Black. Optimization of ecosystem model parameters through assimilating eddy covariance flux data with an ensemble Kalman filter. Ecological Modelling, 217(1– 2):157–173, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002998.

McKenna:2000:FHS

[McK00] James E. McKenna. FITPOP, a heuristic simulation model of population dynamics and genetics with special reference to fisheries. Ecological Modelling, 127(1):81–95, March 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002057.

McKay:2001:VGP

[McK01] R. I. (Bob) McKay. Variants of genetic programming for species distribution modelling — fitness sharing, partial functions, population evaluation. Ecological Modelling, 146(1–3):231–241, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100309X.

Munoz:2007:SAS

[MCK07] François Munoz, Pierre-Olivier Cheptou, and Finn Kjellberg. Spectral analysis of simulated species distribution maps provides insights into metapopulation dynamics. Ecological Modelling, 205(3–4):314–322, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000968.

Massada:2009:EDB

[MCKNM09] Avi Bar Massada, Yohay Carmel, Gili Koniak, and Imanuel Noy-Meir. The effects of disturbance based management on the dynamics of Mediterranean vegetation: a hierarchical and spatially explicit modeling approach. Ecological Modelling, 220(19):2525–2535, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004074.

Maselli:2009:MFC

[MCM^{+}09] F. Maselli, M. Chiesi, M. Moriondo, L. Fibbi, M. Bindi, and S. W. Running. Modelling the forest carbon budget of a Mediterranean region through the integration of ground and satellite data. Ecological Modelling, 220(3):330–342, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000800464X.

Martin:2008:IMC

[MCQA08] Jodie Martin, Clément Calenge, Pierre-Yves Quenette, and Dominique Allainé. Importance of movement constraints in habitat selection studies. Ecological Modelling, 213(2):257–262, May 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006321.

Mau-Crimmins:2006:CIR

[MCSG06] Theresa M. Mau-Crimmins, Heather R. Schussman, and Erika L. Geiger. Can the invaded range of a species be predicted sufficiently using only native-range data?: Lehmann lovegrass (Eragrostis lehmanniana) in the southwestern United States. Ecological Modelling, 193(3–4):736– 746, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438000500445X.

Mitsch:2004:TBW

[MD04] William J. Mitsch and John W. Day. Thinking big with whole-ecosystem studies and ecosystem restoration

— a legacy of H. T. Odum. Ecological Modelling, 178(1–2):133–155, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003005441.

Macedo:2006:PPM

[MD06a] M. F. Macedo and P. Duarte. Phytoplankton production modelling in three marine ecosystems — static versus dynamic approach. Ecological Modelling, 190 (3–4):299–316, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002668.

Mishra:2006:DFU

[MD06b] A. K. Mishra and V. R. Desai. Drought forecasting using feed-forward recursive neural network. Ecological Modelling, 198(1–2):127–138, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002055.

Meyer:2007:PDW

[MD07] João Frederico C. A. Meyer and Geraldo L. Diniz. Pollutant dispersion in wetland systems: Mathematical modelling and numerical simulation. Ecological Modelling, 200 (3–4):360–370, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006003814.

Menard:2002:EPS

[MDB^{+}02] A. Ménard, P. Dubé, A. Bouchard, C. D. Canham, and D. J. Marceau. Evaluating the potential of the SORTIE forest succession model for spatio-temporal analysis of small-scale disturbances. Ecological Modelling, 153(1–2):81–96, July 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001005038.

Monestiez:2006:GMS

[MDB^{+}06] P. Monestiez, L. Dubroca, E. Bonnin, J.-P. Durbec, and C. Guinet. Geostatistical modelling of spatial distribution of Balaenoptera physalus in the Northwestern Mediterranean Sea from sparse count data and heterogeneous observation efforts. Ecological Modelling, 193(3–4):615–628, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004436.

Moxnes:2001:OSU

[MDGK01] Erling Moxnes, Öje Danell, Eldar Gaare, and Jouko Kumpula. Optimal strategies for the use of reindeer rangelands. Ecological Modelling, 145(2–3):225–241, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003933.

Ma:2009:IBS

[MDGV09] Brian O. Ma, Brad H. Davis, David R. Gillespie, and Sherah L. VanLaerhoven. Incorporating behaviour into simple models of dispersal using the biological control agent Dicyphus hesperus. Ecological Modelling, 220(23): 3271–3279, December 10, 2009. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006061.

Mollier:2008:TDS

[MDH^{+}08] Alain Mollier, Peter De Willigen, Marius Heinen, Christian Morel, André Schneider, and Sylvain Pellerin. A two-dimensional simulation model of phosphorus uptake including crop growth and P-response. Ecological Modelling, 210(4):453–464, February 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004164.

Mackinson:2009:WFF

[MDH^{+}09] S. Mackinson, G. Daskalov, J. J. Heymans, S. Neira, H. Arancibia, M. Zetina-Rejón, H. Jiang, H. Q. Cheng, M. Coll, F. Arreguin-Sanchez, K. Keeble, and L. Shannon. Which forcing factors fit? Using ecosystem models to investigate the relative influence of fishing and changes in primary productivity on the dynamics of marine ecosystems. Ecological Modelling, 220(21):2972–2987, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004961.

Mooij:2009:LSE

[MDJ09] W. M. Mooij, L. N. De Senerpont Domis, and J. H. Janse. Linking species- and ecosystem-level impacts of climate change in lakes with a complex and a minimal model. Ecological Modelling, 220(21):3011–3020, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000921.

Moravie:2006:BFG

[MDPC06] M.-A. Moravie, A. C. Davison, D. Pasquier, and P.-J. Charmillot. Bayesian forecasting of grape moth emergence. Ecological Modelling, 197(3–4):478–489, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006001529.

Maionchi:2006:CPF

[MdRdA06] Daniela O. Maionchi, S. F. dos Reis, and M. A. M. de Aguiar. Chaos and pattern formation in a spatial tritrophic food chain. Ecological Modelling, 191 (2):291–303, January 31, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002723.

Mouton:2009:PAO

[MDVG09] Ans M. Mouton, Bernard De Baets, Ester Van Broekhoven, and Peter L. M. Goethals. Prevalence-adjusted optimisation of fuzzy models for species distribution. Ecological Modelling, 220(15):1776–1786, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002828.

Mihailovic:2007:MTC

[ME07] Dragutin T. Mihailovic and Josef Eitzinger. Modelling temperatures of crop environment. Ecological Modelling, 202(3–4):465–475, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006005485.

Medved:2006:PFE

[Med06] Sašo Medved. Present and future ecological footprint of Slovenia — the influence of energy demand scenarios. Ecological Modelling, 192(1–2):25–36, February 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005003297.

Matejicek:2006:GBA

[MEJ06] Luboš Matějíček, Pavel Engst, and Zbyněk Jaňour. A GIS-based approach to spatio-temporal analysis of environmental pollution in urban areas: a case study of Prague’s environment extended by LIDAR data. Ecological Modelling, 199(3):261–277, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002298.

Malmaeus:2008:OSD

[MEKL08] J. M. Malmaeus, J. M. Eklund, O. M. Karlsson, and D. Lindgren. The optimal size of dynamic phosphorus models for Baltic coastal areas. Ecological Modelling, 216(3–4):303–315, September 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002068.

Moisen:2006:FAP

[MEO06] Gretchen G. Moisen, Thomas C. Edwards, and Patrick E. Osborne. Further advances in predicting species distributions. Ecological Modelling, 199(2):129–131, November 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002444.

Metternicht:2001:ATS

[Met01] G. Metternicht. Assessing temporal and spatial changes of salinity using fuzzy logic, remote sensing and GIS. foundations of an expert system. Ecological Modelling, 144 (2–3):163–179, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003714.

Metternicht:2002:EAT

[Met02] G. Metternicht. Erratum to “Assessing temporal and spatial changes of salinity using fuzzy logic, remote sensing and GIS. Foundations of an expert system”: [Ecological Modelling 144 (2001) 163–179]. Ecological Modelling, 149(3):313, April 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002000091.

Metternicht:2003:CFC

[Met03] G. I. Metternicht. Categorical fuzziness: a comparison between crisp and fuzzy class boundary modelling for mapping salt-affected soils using Landsat TM data and a classification based on anion ratios. Ecological Modelling, 168(3):371–389, October 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001479.

Meynecke:2004:EGC

[Mey04] Jan-Olaf Meynecke. Effects of global climate change on geographic distributions of vertebrates in North Queensland. Ecological Modelling, 174(4):347–357, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004344.

Miller:2002:MDF

[MF02a] Jennifer Miller and Janet Franklin. Modeling the distribution of four vegetation alliances using generalized linear models and classification trees with spatial dependence. Ecological Modelling, 157(2–3):227–247, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002001965.

Moisen:2002:CFM

[MF02b] Gretchen G. Moisen and Tracey S. Frescino. Comparing five modelling techniques for predicting forest characteristics. Ecological Modelling, 157(2–3):209–225, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002001977.

Mitra:2006:AVP

[MF06] Aditee Mitra and Kevin J. Flynn. Accounting for variation in prey selectivity by zooplankton. Ecological Modelling, 199(1):82–92, November 1, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600295X.

Miller:2007:ISD

[MFA07] Jennifer Miller, Janet Franklin, and Richard Aspinall. Incorporating spatial dependence in predictive vegetation models. Ecological Modelling, 202(3–4):225–242, April 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006006247.

Mazzuoli:2005:STC

[MFB^{+}05] Stefania Mazzuoli, Silvia Focardi, Luca Bracchini, Margherita Falcucci, Steven A. Loiselle, and Claudio Rossi. Spatial and temporal characterisations of the degradation of dissolved humic substances in freshwater lake. Ecological Modelling, 186(1):55–61, July 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000232.

Moisen:2006:PTS

[MFB^{+}06] Gretchen G. Moisen, Elizabeth A. Freeman, Jock A. Blackard, Tracey S. Frescino, Niklaus E. Zimmermann, and Thomas C. Edwards. Predicting tree species presence and basal area in Utah: a comparison of stochastic gradient boosting, generalized additive models, and tree-based methods. Ecological Modelling, 199(2):176–187, November 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002560.

Mellin:2006:DCR

[MFG^{+}06] C. Mellin, J. Ferraris, R. Galzin, M. Kulbicki, and D. Ponton. Diversity of coral reef fish assemblages: Modelling of the species richness spectra from multi-scale environmental variables in the Tuamotu Archipelago (French Polynesia). Ecological Modelling, 198(3–4):409–425, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002419.

Mellin:2008:NAI

[MFG^{+}08] C. Mellin, J. Ferraris, R. Galzin, M. Harmelin-Vivien, M. Kulbicki, and T. Lison de Loma. Natural and anthropogenic influences on the diversity structure of reef fish communities in the Tuamotu Archipelago (French Polynesia). Ecological Modelling, 218(1–2):182–187, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003700.

Mehta:2004:EEA

[MFJM04] Smita Mehta, Lee E. Frelich, Malcolm T. Jones, and Jim Manolis. Examining the effects of alternative management strategies on landscape-scale forest patterns in northeastern Minnesota using LANDIS. Ecological Modelling, 180(1):73–87, December 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003503.

Milner-Gulland:2001:DGM

[MG01] E. J. Milner-Gulland. A dynamic game model for the decision to join an aggregation. Ecological Modelling, 145(1):85–99, November 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003817.

Musacchio:2002:APW

[MG02] Laura R. Musacchio and William E. Grant. Agricultural production and wetland habitat quality in a coastal prairie ecosystem: simulated effects of alternative resource policies on land-use decisions. Ecological Modelling, 150(1–2):23–43, April 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001004598.

Melia:2005:SBM

[MG05] Paco Melià and Marino Gatto. A stochastic bioeconomic model for the management of clam farming. Ecological Modelling, 184(1):163–174, May 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005903.

Mesterton-Gibbons:2009:EAI

[MG09] Mike Mesterton-Gibbons. Erratum to “Assessing the impact of biological control of Plutella xylostella through the application of Lotka–Volterra model” [Ecol. Model. 220 (2009) 60–70]. Ecological Modelling, 220(19):2602, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000900489X.

Masera:2003:MCS

[MGCK^{+}03] Omar R. Masera, J. F. Garza-Caligaris, M. Kanninen, T. Karjalainen, J. Liski, G. J. Nabuurs, A. Pussinen, B. H. J. de Jong, and G. M. J. Mohren. Modeling carbon sequestration in afforestation, agroforestry and forest management projects: the CO2FIX V.2 approach. Ecological Modelling, 164(2–3):177–199, June 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004192.

Martinez:2005:SAP

[MGH^{+}05] Cristina A. Martinez, W. E. Grant, Sallie J. Hejl, Markus J. Peterson, Alfonso Martinez, and Gary L. Waggerman. Simulation of annual productivity and long-term population trends of white-winged doves in the Tamaulipan Biotic Province. Ecological Modelling, 181(2–3):149–159, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003631.

Makarieva:2004:OGM

[MGL04] Anastassia M. Makarieva, Victor G. Gorshkov, and Bai-Lian Li. Ontogenetic growth: models and theory. Ecological Modelling, 176(1–2):15–26, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005398.

Merckx:2009:PMN

[MGS^{+}09] Bea Merckx, Peter Goethals, Maaike Steyaert, Ann Vanreusel, Magda Vincx, and Jan Vanaverbeke. Predictability of marine nematode biodiversity. Ecological Modelling, 220(11):1449–1458, June 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900194X.

Moolenaar:2007:TMA

[MGSdG07] Hanneke E. Moolenaar, Johan Grasman, Frank M. Selten, and Maarten de Gee. Testing a method for analyzing the effect of parameter change in climate driven ecological systems. Ecological Modelling, 205(3–4):289–300, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007000981.

MacFarlane:2000:IUP

[MGV00] David W. MacFarlane, Edwin J. Green, and Harry T. Valentine. Incorporating uncertainty into the parameters of a forest process model. Ecological Modelling, 134(1):27–40, September 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000000329X.

Machiwa:2002:EMF

[MH02] John F. Machiwa and Rolf O. Hallberg. An empirical model of the fate of organic carbon in a mangrove forest partly affected by anthropogenic activity. Ecological Modelling, 147(1):69–83, January 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004070.

Malmaeus:2003:DMP

[MH03] J. M. Malmaeus and L. Håkanson. A dynamic model to predict suspended particulate matter in lakes. Ecological Modelling, 167(3):247–262, September 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001662.

Malmaeus:2004:DLE

[MH04] J. M. Malmaeus and L. Håkanson. Development of a lake eutrophication model. Ecological Modelling, 171 (1–2):35–63, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002977.

Melesse:2005:ANN

[MH05] Assefa M. Melesse and Rodney S. Hanley. Artificial neural network application for multi-ecosystem carbon flux simulation. Ecological Modelling, 189(3–4):305–314, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002061.

Moxnes:2007:PDP

[MH07] John F. Moxnes and Kjell Hausken. The population dynamics of potato cyst nematodes. Ecological Modelling, 207(2–4):339–348, October 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003213.

Muller:2000:IEI

[MHKW00] Felix Müller, Regina Hoffmann-Kroll, and Hubert Wiggering. Indicating ecosystem integrity — theoretical concepts and environmental requirements. Ecological Modelling, 130(1–3):13–23, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002106.

Matala:2003:CPM

[MHM^{+}03] J. Matala, J. Hynynen, J. Miina, R. Ojansuu, H. Peltola, R. Sievänen, H. Väisänen, and S. Kellomäki. Comparison of a physiological model and a statistical model for prediction of growth and yield in boreal forests. Ecological Modelling, 161(1–2):95–116, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002971.

Manickchand-Heileman:2004:TME

[MHMHKA04] Sherry Manickchand-Heileman, Jeremy Mendoza-Hill, Amoy Lum Kong, and Freddy Arocha. A trophic model for exploring possible ecosystem impacts of fishing in the Gulf of Paria, between Venezuela and Trinidad. Ecological Modelling, 172(2–4):307–322, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003752.

Muller-Herold:2000:CAF

[MHN00] Ulrich Müller-Herold and Gregor Nickel. A closed analytical formula for the characteristic spatial range of persistent organic pollutants. Ecological Modelling, 126 (2–3):191–200, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002659.

Monte:2006:ECS

[MHP^{+}06] Luigi Monte, Lars Håkanson, Raul Periañez, Gennady Laptev, Mark Zheleznyak, Vladimir Maderich, Giacomo Angeli, and Vladimir Koshebutsky. Experiences from a case study of multi-model application to assess the behaviour of pollutants in the Dnieper–Bug Estuary. Ecological Modelling, 195(3–4):247–263, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006265.

Muller-Herold:2006:SMP

[MHSP^{+}06] Ulrich Müller-Herold, Timo Smieszek, Philippe Peter, Martin Scheringer, and Marco Morosini. A simple measure for precautionary assessment of organic chemicals with respect to global cold condensation. Ecological Modelling, 194(1–3):266–273, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005247.

Makowski:2000:FSN

[MHvIR00] David Makowski, Eligius M. T. Hendrix, Martin K. van Ittersum, and Walter A. H. Rossing. A framework to study nearly optimal solutions of linear programming models developed for agricultural land use exploration. Ecological Modelling, 131(1):65–77, June 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002490.

Mumby:2006:RCU

[MHZ^{+}06] Peter J. Mumby, John D. Hedley, Kamila Zychaluk, Alastair R. Harborne, and Paul G. Blackwell. Revisiting the catastrophic die-off of the urchin Diadema antillarum on Caribbean coral reefs: Fresh insights on resilience from a simulation model. Ecological Modelling, 196(1–2):131–148, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005006174.

Mahujchariyawong:2001:MEP

[MI01] Jukkrit Mahujchariyawong and Saburo Ikeda. Modelling of environmental phytoremediation in eutrophic river — the case of water hyacinth harvest in Tha-chin River, Thailand. Ecological Modelling, 142(1–2):121–134, August 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002836.

Miina:2000:DTR

[Mii00] Jari Miina. Dependence of tree-ring, earlywood and latewood indices of Scots pine and Norway spruce on climatic factors in eastern Finland. Ecological Modelling, 132(3):259–273, August 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002969.

Mitra:2009:CTA

[Mit09] Aditee Mitra. Are closure terms appropriate or necessary descriptors of zooplankton loss in nutrient–phytoplankton– zooplankton type models? Ecological Modelling, 220 (5):611–620, March 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008005899.

Marques:2002:TSE

[MJ02] João Carlos Marques and Sven Erik Jørgensen. Three selected ecological observations interpreted in terms of a thermodynamic hypothesis. Contribution to a general theoretical framework. Ecological Modelling, 158 (3):213–221, December 31, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002002351.

Munkemuller:2006:CVC

[MJ06] Tamara Münkemüller and Karin Johst. Compensatory versus over-compensatory density regulation: Implications for metapopulation persistence in dynamic landscapes. Ecological Modelling, 197(1–2):171–178, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001037.

Miller:2002:DDM

[MJH02] David H. Miller, Al L. Jensen, and James H. Hammill. Density dependent matrix model for gray wolf population projection. Ecological Modelling, 151(2–3):271–278, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004938.

McPherson:2006:UCG

[MJR06] Jana M. McPherson, Walter Jetz, and David J. Rogers. Using coarse-grained occurrence data to predict species distributions at finer spatial resolutions — possibilities and limitations. Ecological Modelling, 192(3–4):499–522, February 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003789.

Merkle:2000:DEE

[MK00] Andrea Merkle and Martin Kaupenjohann. Derivation of ecosystemic effect indicators method. Ecological Modelling, 130(1–3):39–46, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002131.

McLachlan-Karr:2004:EEH

[MK04] John McLachlan-Karr. The ecosystem, energy and human values. Ecological Modelling, 178(1–2):267–274, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005428.

Mailly:2000:DSC

[MKB00] D. Mailly, J. P. Kimmins, and R. T. Busing. Disturbance and succession in a coniferous forest of northwestern North America: simulations with dryades, a spatial gap model. Ecological Modelling, 127(2–3):183–205, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002082.

Mathey:2008:OOC

[MKDV08] Anne-Hélène Mathey, Emina Krcmar, Suzana Dragicevic, and Ilan Vertinsky. An object-oriented cellular automata model for forest planning problems. Ecological Modelling, 212(3–4):359–371, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006060.

Mukai:2007:ISS

[MKiIK07] Daiki Mukai, Michio J. Kishi, Shin ichi Ito, and Yutaka Kurita. The importance of spawning season on the growth of Pacific saury: a model-based study using NEMURO.FISH. Ecological Modelling, 202(1–2):165–173, March 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006004686.

Matis:2007:PSM

[MKM^{+}07] James H. Matis, Thomas R. Kiffe, Timothy I. Matis, John A. Jackman, and Harvir Singh. Population size models based on cumulative size, with application to aphids. Ecological Modelling, 205(1–2):81–92, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000610.

Matis:2008:UGR

[MKM^{+}08] James H. Matis, Thomas R. Kiffe, Timothy I. Matis, John A. Jackman, William E. Grant, and Harvir Singh. On the use of growth rate parameters for projecting population sizes: Application to aphids. Ecological Modelling, 213(1):133–142, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007006291.

Matis:2003:SSA

[MKRH03] James H. Matis, Thomas R. Kiffe, Eric Renshaw, and Janet Hassan. A simple saddlepoint approximation for the equilibrium distribution of the stochastic logistic model of population growth. Ecological Modelling, 161(3):239–248, March 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003447.

Momo:2002:RLE

[MKS^{+}02] Fernando Momo, Jens Kowalke, Irene Schloss, Guillermo Mercuri, and Gustavo Ferreyra. The role of Laternula elliptica in the energy budget of Potter Cove (King George Island, Antarctica). Ecological Modelling, 155 (1):43–51, September 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000819.

Matamba:2009:TAS

[MKS^{+}09] L. Matamba, C. Kazanci, J. R. Schramski, M. Blessing, P. Alexander, and B. C. Patten. Throughflow analysis: a stochastic approach. Ecological Modelling, 220 (22):3174–3181, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004815.

Matis:2009:PDM

[MKvdW^{+}09] James H. Matis, Thomas R. Kiffe, Wopke van der Werf, Alejandro C. Costamagna, Timothy I. Matis, and William E. Grant. Population dynamics models based on cumulative density dependent feedback: a link to the logistic growth curve and a test for symmetry using aphid data. Ecological Modelling, 220(15):1745–1751, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900283X.

Marsili-Libelli:2004:FPR

[ML04] S. Marsili-Libelli. Fuzzy pattern recognition of circadian cycles in ecosystems. Ecological Modelling, 174(1–2):67–84, May 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005696.

Muttil:2005:GPA

[ML05] Nitin Muttil and Joseph H. W. Lee. Genetic programming for analysis and real-time prediction of coastal algal blooms. Ecological Modelling, 189(3–4):363–376, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002127.

Marks:2006:HTS

[ML06] Christian O. Marks and Martin J. Lechowicz. A holistic tree seedling model for the investigation of functional trait diversity. Ecological Modelling, 193(3–4):141–181, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004825.

Mihailovic:2002:SAT

[MLA^{+}02] D. T. Mihailović, B. Lalić, I. Arsenić, J. Eitzinger, and N. Dušanić. Simulation of air temperature inside the canopy by the LAPS surface scheme. Ecological Modelling, 147(3):199–207, January 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004227.

Mladenoff:2004:LFL

[Mla04] David J. Mladenoff. LANDIS and forest landscape models. Ecological Modelling, 180(1):7–19, December 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003461.

Marsili-Libelli:2005:IDM

[MLC05] Stefano Marsili-Libelli and Nicola Checchi. Identification of dynamic models for horizontal subsurface constructed wetlands. Ecological Modelling, 187(2–3):201–218, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000839.

Miehle:2006:APC

[MLF^{+}06] P. Miehle, S. J. Livesley, P. M. Feikema, C. Li, and S. K. Arndt. Assessing productivity and carbon sequestration capacity of Eucalyptus globulus plantations using the process model Forest-DNDC: Calibration and validation. Ecological Modelling, 192(1–2):83–94, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003959.

Marsili-Libelli:2003:CRE

[MLGC03] S. Marsili-Libelli, S. Guerrizio, and N. Checchi. Confidence regions of estimated parameters for ecological systems. Ecological Modelling, 165(2–3):127–146, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000681.

Moritz:2009:QDM

[MLGG09] Charlotte Moritz, Nicolas Loeuille, Jean-Marc Guarini, and Katell Guizien. Quantifying the dynamics of marine invertebrate metacommunities: What processes can maintain high diversity with low densities in the Mediterranean Sea? Ecological Modelling, 220(21):3021–3032, November 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008005450.

Montoya:2000:ESN

[MLGV00a] Ruby A. Montoya, A. L. Lawrence, W. E. Grant, and M. Velasco. Erratum to “Simulation of nitrogen dynamics and shrimp growth in an intensive shrimp culture system: effects of feed and feeding parameters”: [Ecol. Model. 122 (1999) 81–95]. Ecological Modelling, 129(1):111, May 3, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000002271.

Montoya:2000:SPD

[MLGV00b] Ruby A. Montoya, A. L. Lawrence, W. E. Grant, and M. Velasco. Simulation of phosphorus dynamics in an intensive shrimp culture system: effects of feed formulations and feeding strategies. Ecological Modelling, 129(2–3):131–142, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002301.

Malve:2005:EWR

[MLH05] Olli Malve, Marko Laine, and Heikki Haario. Estimation of winter respiration rates and prediction of oxygen regime in a lake using Bayesian inference. Ecological Modelling, 182(2):183–197, March 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004120.

Marmion:2009:PSA

[MLHT09] Mathieu Marmion, Miska Luoto, Risto K. Heikkinen, and Wilfried Thuiller. The performance of state-of-the-art modelling techniques depends on geographical distribution of species. Ecological Modelling, 220(24):3512–3520, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008004948.

Mo:2005:PCY

[MLL^{+}05] X. Mo, S. Liu, Z. Lin, Y. Xu, Y. Xiang, and T. R. McVicar. Prediction of crop yield, water consumption and water use efficiency with a SVAT-crop growth model using remotely sensed data on the North China plain. Ecological Modelling, 183(2–3):301–322, April 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004855.

Moret:2006:LPC

[MLM06] Stephanie L. Moret, William T. Langford, and Dragos D. Margineantu. Learning to predict channel stability using biogeomorphic features. Ecological Modelling, 191(1):47–57, January 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005003480.

Li:2005:FAA

[mLMfT05] Xiang mei Li, Min Min, and Chuan feng Tan. The functional assessment of agricultural ecosystems in Hubei Province, China. Ecological Modelling, 187(2–3):352–360, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000797.

McDonald:2001:SEE

[MLPK01] Amanda A. McDonald, Jianguo Liu, Harold Prince, and Kiersten Kress. A socio-economic-ecological simulation model of land acquisition to expand a national wildlife refuge. Ecological Modelling, 140(1–2):99–110, May 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002599.

Monty:2008:IDI

[MLR08] Arnaud Monty, Philippe Lejeune, and Jacques Rondeux. Individual distance-independent girth increment model for Douglas-fir in southern Belgium. Ecological Modelling, 212(3–4):472–479, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007005686.

Meng:2006:SPL

[MLTN06] Yaobin Meng, Bin-Le Lin, Mamoru Tominaga, and Junko Nakanishi. Simulation of the population-level effects of 4nonylphenol on wild Japanese medaka (Oryzias latipes). Ecological Modelling, 197(3–4):350–360, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006001359.

Murty:2000:DFP

[MM00] Danuse Murty and Ross E. McMurtrie. The decline of forest productivity as stands age: a model-based method for analysing causes for the decline. Ecological Modelling, 134(2–3):185–205, October 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003458.

Malins:2006:ASE

[MM06a] David Malins and Graciela Metternicht. Assessing the spatial extent of dryland salinity through fuzzy modeling. Ecological Modelling, 193(3–4):387–411, March 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005004448.

Marchini:2006:FLM

[MM06b] Agnese Marchini and Carlo Marchini. A fuzzy logic model to recognise ecological sectors in the lagoon of Venice based on the benthic community. Ecological Modelling, 193(1–2):105–118, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500459X.

Mazaris:2006:IBMa

[MM06c] Antonios D. Mazaris and Yiannis G. Matsinos. An individual based model of sea turtles: Investigating the effect of temporal variability on population dynamics. Ecological Modelling, 194(1–3):114–124, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005028.

Moisan:2002:MET

[MMA02] John R. Moisan, Tiffany A. Moisan, and Mark R. Abbott. Modelling the effect of temperature on the maximum growth rates of phytoplankton populations. Ecological Modelling, 153(3):197–215, August 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200008X.

Miller:2007:NMI

[MMB07] Harlan L. Miller, Christof Meile, and Adrian B. Burd. A novel 2D model of internal O_{2} dynamics and H_{2}S intrusion in seagrasses. Ecological Modelling, 205(3–4):365–380, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001214.

Miehls:2009:ISIb

[MMF^{+}09a] Andrea L. Jaeger Miehls, Doran M. Mason, Kenneth A. Frank, Ann E. Krause, Scott D. Peacor, and William W. Taylor. Invasive species impacts on ecosystem structure and function: a comparison of Oneida Lake, New York, USA, before and after zebra mussel invasion. Ecological Modelling, 220(22):3194–3209, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004992.

Miehls:2009:ISIa

[MMF^{+}09b] Andrea L. Jaeger Miehls, Doran M. Mason, Kenneth A. Frank, Ann E. Krause, Scott D. Peacor, and William W. Taylor. Invasive species impacts on ecosystem structure and function: a comparison of the Bay of Quinte, Canada, and Oneida Lake, USA, before and after zebra mussel invasion. Ecological Modelling, 220(22):3182–3193, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004943.

Milden:2006:MDP

[MMHE06] Mikael Mildén, Zuzana Münzbergová, Tomáš Herben, and Johan Ehrlén. Metapopulation dynamics of a perennial plant, Succisa pratensis, in an agricultural landscape. Ecological Modelling, 199(4):464–475, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002900.

Muller:2003:DPR

[MMJN03] Torsten Müller, Jakob Magid, Lars Stoumann Jensen, and Niels Erik Nielsen. Decomposition of plant residues of different quality in soil — DAISY model calibration and simulation based on experimental data. Ecological Modelling, 166(1–2):3–18, August 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001145.

Matthiopoulos:2000:KFH

[MML00] Jason Matthiopoulos, Robert Moss, and Xavier Lambin. The kin-facilitation hypothesis for red grouse population cycles: territory sharing between relatives. Ecological Modelling, 127(1):53–63, March 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001994.

Matthiopoulos:2002:KFH

[MML02] Jason Matthiopoulos, Robert Moss, and Xavier Lambin. The kin facilitation hypothesis for red grouse population cycles: territorial dynamics of the family cluster. Ecological Modelling, 147(3):291–307, January 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004203.

Montano-Moctezuma:2007:ACS

[MMLR07] Gabriela Montaño-Moctezuma, Hiram W. Li, and Philippe A. Rossignol. Alternative community structures in a kelpurchin community: a qualitative modeling approach. Ecological Modelling, 205(3–4):343–354, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001160.

Martins:2002:MES

[MMM02] Irene Martins, Paulo Maranhão, and João Carlos Marques. Modelling the effects of salinity variation on Echinogammarus marinus Leach (Amphipoda, Gammaridae) density and biomass in the Mondego Estuary (Western Portugal). Ecological Modelling, 152(2–3):247–260, July 1, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002000121.

Mukherjee:2008:MCN

[MMN08] B. Mukherjee, D. Mukherjee, and M. Nivedita. Modelling carbon and nutrient cycling in a simulated pond system at ranchi. Ecological Modelling, 213(3–4):437–448, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000343.

Mukherjee:2007:TRI

[MMPN07] B. Mukherjee, D. Mukherjee, A. Prasad, and M. Nivedita. The transfer rates of inorganic carbon and budgetary analysis in a simulated aquatic system. Ecological Modelling, 204(3–4):279–288, June 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000191.

McCausland:2006:SMS

[MMPT06] W. D. McCausland, E. Mente, G. J. Pierce, and I. Theodossiou. A simulation model of sustainability of coastal communities: Aquaculture, fishing, environment and labour markets. Ecological Modelling, 193(3–4):271–294, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004060.

McLeod:2002:MLS

[MMR02] P. McLeod, A. P. Martin, and K. J. Richards. Minimum length scale for growth-limited oceanic plankton distributions. Ecological Modelling, 158(1–2):111–120, December 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000200176X.

McNeil:2006:GBM

[MMR06] Brenden E. McNeil, Richard E. Martell, and Jane M. Read. GIS and biogeochemical models for examining the legacy of forest disturbance in the Adirondack Park, NY, USA. Ecological Modelling, 195(3–4):281–295, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006034.

McKenna:2006:SPC

[MMRLP06] K. C. McKenna, D. L. Moorhead, E. C. Roberts, and J. Laybourn-Parry. Simulated patterns of carbon flow in the pelagic food web of Lake Fryxell, Antarctica: Little evidence of top-down control. Ecological Modelling, 192(3–4):457–472, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003558.

Morgan:2007:PTM

[MMT^{+}07] E. R. Morgan, G. F. Medley, P. R. Torgerson, B. S. Shaikenov, and E. J. Milner-Gulland. Parasite transmission in a migratory multiple host system. Ecological Modelling, 200(3–4):511–520, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004091.

Maina:2008:MSM

[MMV08] Joseph Maina, Timothy R. McClanahan, and Valentijn Venus. Meso-scale modelling of coral’s susceptibility to environmental stress using remotely sensed data: Reply to comments by Dunne (2008). Ecological Modelling, 218(1–2):192–194, October 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003104.

Meyer:2009:PSE

[MMV^{+}09] Katrin M. Meyer, Wolf M. Mooij, Matthijs Vos, W. H. Gera Hol, and Wim H. van der Putten. The power of simulating experiments. Ecological Modelling, 220(19):2594–2597, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009004086.

Mladenov:2007:SDF

[MMWMH07] N. Mladenov, D. M. McKnight, P. Wolski, and M. Murray-Hudson. Simulation of DOM fluxes in a seasonal floodplain of the Okavango Delta, Botswana. Ecological Modelling, 205(1–2):181–195, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000890.

Munier:2001:BLM

[MNEB01] Bernd Münier, B. Nygaard, R. Ejrnæs, and H. G. Bruun. A biotope landscape model for prediction of semi-natural vegetation in Denmark. Ecological Modelling, 139(2–3):221–233, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001002447.

McDonald:2008:MHO

[MNOS08] Patrick T. McDonald, Clayton K. Nielsen, Tonny J. Oyana, and Wanxiao Sun. Modelling habitat overlap among sympatric mesocarnivores in southern Illinois, USA. Ecological Modelling, 215(4):276–286, July 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001658.

Marques:2003:IER

[MNPJ03] João C. Marques, Søren N. Nielsen, Miguel A. Pardal, and Sven E. Jørgensen. Impact of eutrophication and river management within a framework of ecosystem theories. Ecological Modelling, 166(1–2):147–168, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001340.

Marinov:2006:ACM

[MNZ06] Dimitar Marinov, Alain Norro, and José-Manuel Zaldívar. Application of COHERENS model for hydrodynamic investigation of Sacca di Goro coastal lagoon (Italian Adriatic Sea shore). Ecological Modelling, 193(1–2):52–68, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004564.

Muzy:2008:MSF

[MNZC08] A. Muzy, J. J. Nutaro, B. P. Zeigler, and P. Coquillard. Modeling and simulation of fire spreading through the activity tracking paradigm. Ecological Modelling, 219(1–2):212–225, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004134.

Moghadas:2002:TCG

[Mog02] S. M. Moghadas. Two core group models for sexual transmission of disease. Ecological Modelling, 148(1):15–26, February 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000100429X.

Moilanen:2004:SSS

[Moi04] Atte Moilanen. SPOMSIM: software for stochastic patch occupancy models of metapopulation dynamics. Ecological Modelling, 179(4):533–550, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002790.

Murray:2001:MTV

[MOJ01] Alexander G. Murray, Michael O’Callaghan, and Brian Jones. A model of transmission of a viral epidemic among schools within a shoal of pilchards. Ecological Modelling, 144(2–3):245–259, October 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003751.

McKenzie:2006:IMP

[MOLN06] Donald McKenzie, Susan M. O’Neill, Narasimhan K. Larkin, and Robert A. Norheim. Integrating models to predict regional haze from wildland fire. Ecological Modelling, 199(3):278–288, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002304.

Monte:2002:MMC

[Mon02] Luigi Monte. A methodology for modelling the contamination of moving organisms in water bodies with spatial and time dependent pollution levels. Ecological Modelling, 158(1–2):21–33, December 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002314.

Montero:2005:RDA

[Mon05] Eloísa Montero. Rényi dimensions analysis of soil particlesize distributions. Ecological Modelling, 182(3–4):305–315, March 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004001656.

Monte:2009:MAD

[Mon09a] Luigi Monte. A methodological approach to develop “contaminant migration — population effects” models. Ecological Modelling, 220(23):3280–3290, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009006085.

Monte:2009:MMA

[Mon09b] Luigi Monte. Multi-model approach and evaluation of the uncertainty of model results. Rationale and applications to predict the behaviour of contaminants in the abiotic components of the fresh water environment. Ecological Modelling, 220(12):1469–1480, June 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001987.

Matala:2005:IET

[MOP^{+}05] J. Matala, R. Ojansuu, H. Peltola, R. Sievänen, and S. Kellomäki. Introducing effects of temperature and CO_{2} elevation on tree growth into a statistical growth and yield model. Ecological Modelling, 181(2–3):173–190, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003655.

Matala:2006:MRT

[MOP^{+}06] J. Matala, R. Ojansuu, H. Peltola, H. Raitio, and S. Kellomäki. Modelling the response of tree growth to temperature and CO_{2} elevation as related to the fertility and current temperature sum of a site. Ecological Modelling, 199(1):39–52, November 1, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002791.

Malkina-Pykh:2000:DTE

[MP00] Irina G. Malkina-Pykh. From data and theory to environmental models and indices formation. Ecological Modelling, 130(1–3):67–77, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002064.

Mahevas:2004:IFG

[MP04a] Stéphanie Mahévas and Dominique Pelletier. ISIS-Fish, a generic and spatially explicit simulation tool for evaluating the impact of management measures on fisheries dynamics. Ecological Modelling, 171(1–2):65–84, January 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002989.

Mitchell:2004:MHR

[MP04b] Michael S. Mitchell and Roger A. Powell. A mechanistic home range model for optimal use of spatially distributed resources. Ecological Modelling, 177(1–2):209–232, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004001176.

Markensten:2007:WDI

[MP07] Hampus Markensten and Donald C. Pierson. Weather driven influences on phytoplankton succession in a shallow lake during contrasting years: Application of PROTBAS. Ecological Modelling, 207(2–4):128–136, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002554.

Mitchell:2008:EHR

[MP08] Michael S. Mitchell and Roger A. Powell. Estimated home ranges can misrepresent habitat relationships on patchy landscapes. Ecological Modelling, 216(3–4):409–414, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380008002238.

Mayer:2006:FID

[MPC06] Audrey L. Mayer, Christopher W. Pawlowski, and Heriberto Cabezas. Fisher information and dynamic regime changes in ecological systems. Ecological Modelling, 195(1–2):72–82, May 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000500579X.

Malchow:2002:NSP

[MPM02] Horst Malchow, Sergei V. Petrovskii, and Alexander B. Medvinsky. Numerical study of plankton–fish dynamics in a spatially structured and noisy environment. Ecological Modelling, 149(3):247–255, April 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004677.

Martin:2005:SFD

[MPP05] Miguel Angel Martín, Yakov A. Pachepsky, and Edmund Perfect. Scaling, fractals and diversity in soils and ecohydrology. Ecological Modelling, 182(3–4):217–220, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004001589.

Martins:2004:MED

[MPRJ04] Irene Martins, José Costa Pereira, Jaime A. Ramos, and Sven Erik Jørgensen. Modelling the effects of different quality prey fish species and of food supply reduction on growth performance of roseate tern chicks. Ecological Modelling, 177(1–2):95–106, September 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001097.

Mukherjee:2002:MMS

[MPS02] B. Mukherjee, P. N. Pandey, and S. N. Singh. Mathematical modelling and system analysis of inorganic carbon in the aquatic environment. Ecological Modelling, 152(2–3):129–143, July 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004574.

McDermot:2000:IBM

[MR00] Dennis McDermot and Kenneth A. Rose. An individualbased model of lake fish communities: application to piscivore stocking in Lake Mendota. Ecological Modelling, 125(1):67–102, January 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001726.

Mieleitner:2006:ATB

[MR06] Johanna Mieleitner and Peter Reichert. Analysis of the transferability of a biogeochemical lake model to lakes of different trophic state. Ecological Modelling, 194(1–3):49–61, March 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005004965.

Mieleitner:2008:MFG

[MR08] Johanna Mieleitner and Peter Reichert. Modelling functional groups of phytoplankton in three lakes of different trophic state. Ecological Modelling, 211(3–4):279– 291, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007004735.

Moilanen:2006:PRR

[MRE^{+}06] Atte Moilanen, Michael C. Runge, Jane Elith, Andrew Tyre, Yohay Carmel, Eric Fegraus, Brendan A. Wintle, Mark Burgman, and Yakov Ben-Haim. Planning for robust reserve networks using uncertainty analysis. Ecological Modelling, 199(1):115–124, November 1, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003127.

Mandal:2009:MCD

[MRG09] Sudipto Mandal, Santanu Ray, and Phani Bhusan Ghosh. Modelling of the contribution of dissolved inorganic nitrogen (DIN) from litterfall of adjacent mangrove forest to Hooghly–Matla estuary, India. Ecological Modelling, 220(21):2988–3000, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000829.

Megrey:2007:NPB

[MRiI^{+}07] Bernard A. Megrey, Kenneth A. Rose, Shin ichi Ito, Douglas E. Hay, Francisco E. Werner, Yasuhiro Yamanaka, and Maki Noguchi Aita. North Pacific basin-scale differences in lower and higher trophic level marine ecosystem responses to climate impacts using a nutrient-phytoplankton– zooplankton model coupled to a fish bioenergetics model. Ecological Modelling, 202(1–2):196–210, March 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006004832.

Megrey:2007:BBP

[MRK^{+}07] Bernard A. Megrey, Kenneth A. Rose, Robert A. Klumb, Douglas E. Hay, Francisco E. Werner, David L. Eslinger, and S. Lan Smith. A bioenergetics-based population dynamics model of Pacific herring (Clupea harengus pallasi) coupled to a lower trophic level nutrient–phytoplankton– zooplankton model: Description, calibration, and sensitivity analysis. Ecological Modelling, 202(1–2):144–164, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004674.

Mandal:2006:OCV

[MRRJ06] Sandip Mandal, Santanu Ray, Samar Roy, and Sven Erik Jørgensen. Order to chaos and vice versa in an aquatic ecosystem. Ecological Modelling, 197(3–4):498–504, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000600144X.

Mandal:2007:ITP

[MRRJ07] Sandip Mandal, Santanu Ray, Samar Roy, and Sven Erik Jørgensen. Investigation of thermodynamic properties in an ecological model developing from ordered to chaotic states. Ecological Modelling, 204(1–2):40–46, May 24, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006006429.

Mistro:2005:MMD

[MRS05] Diomar C. Mistro, Luiz Alberto D. Rodrigues, and Andréia B. Schmid. A mathematical model for dispersal of an annual plant population with a seed bank. Ecological Modelling, 188(1):52–61, October 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002413.

Martin:2005:EBH

[MRT05] Miguel Angel Martín, José-Manuel Rey, and Fco. Javier Taguas. An entropy-based heterogeneity index for mass– size distributions in Earth science. Ecological Modelling, 182(3–4):221–228, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001590.

Muhammetoglu:2000:TDW

[MS00] Ayşe (Bayar) Muhammetoğlu and Selçuk Soyupak. A threedimensional water quality-macrophyte interaction model for shallow lakes. Ecological Modelling, 133(3):161–180, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000002970.

Muller:2003:LML

[MS03] Felix Müller and Uta Steinhardt. Landscape modelling and landscape analysis. Ecological Modelling, 168(3):215–216, October 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001376.

Martineau:2007:PBM

[MS07] Yann Martineau and Bernard Saugier. A process-based model of old field succession linking ecosystem and community ecology. Ecological Modelling, 204(3–4):399–419, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700052X.

Merot:2003:TCT

[MSA^{+}03] Ph. Merot, H. Squividant, P. Aurousseau, M. Hefting, T. Burt, V. Maitre, M. Kruk, A. Butturini, C. Thenail, and V. Viaud. Testing a climato-topographic index for predicting wetlands distribution along an European climate gradient. Ecological Modelling, 163(1–2):51–71, May 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003873.

Monteiro:2007:CDA

[MSB07] L. H. A. Monteiro, J. B. Sasso, and J. G. Chaui Berlinck. Continuous and discrete approaches to the epidemiology of viral spreading in populations taking into account the delay of incubation time. Ecological Modelling, 201(3–4):553–557, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004662.

Moriondo:2008:ROT

[MSB08] M. Moriondo, F. M. Stefanini, and M. Bindi. Reproduction of olive tree habitat suitability for global change impact assessment. Ecological Modelling, 218(1–2):95–109, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008003050.

Mullon:2009:NNE

[MSC09] Christian Mullon, Yunne Shin, and Philippe Cury. NEATS: a network economics approach to trophic systems. Ecological Modelling, 220(21):3033–3045, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000891.

Moustakas:2006:SEL

[MSD06] Aristides Moustakas, William Silvert, and Apostolos Dimitromanolakis. A spatially explicit learning model of migratory fish and fishers for evaluating closed areas. Ecological Modelling, 192(1–2):245–258, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003406.

Martin:2000:PBM

[MSHL00] M. J. Martin, C. M. Stirling, S. W. Humphries, and S. P. Long. A process-based model to predict the effects of climatic change on leaf isoprene emission rates. Ecological Modelling, 131(2–3):161–174, July 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002581.

Mayaka:2004:PDM

[MSHP04] Theodore B. Mayaka, Johannes D. Stigter, Ignas M. A. Heitkönig, and Herbert H. T. Prins. A population dynamics model for the management of Buffon’s kob (Kobus kob kob) in the Bénoué National Park Complex, Cameroon. Ecological Modelling, 176(1–2):135–153, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005611.

Maier:2001:FCA

[MSL01] Holger R. Maier, Tarek Sayed, and Barbara J. Lence. Forecasting cyanobacterium Anabaena spp. in the River Murray, South Australia, using B-spline neurofuzzy models. Ecological Modelling, 146(1–3):85–96, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002988.

Mottus:2006:ECV

[MSL06] Matti Mõttus, Madis Sulev, and Mait Lang. Estimation of crown volume for a geometric radiation model from detailed measurements of tree structure. Ecological Modelling, 198(3–4):506–514, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002602.

Mantzouni:2007:PSS

[MSM^{+}07] Irene Mantzouni, Stylianos Somarakis, Dimitrios K. Moutopoulos, Argyris Kallianiotis, and Constantin Koutsikopoulos. Periodic, spatially structured matrix model for the study of anchovy (Engraulis encrasicolus) population dynamics in N Aegean Sea (E. Mediterranean). Ecological Modelling, 208(2–4):367–377, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003043.

McRae:2008:MMF

[MSM^{+}08] Brad H. McRae, Nathan H. Schumaker, Robert B. McKane, Richard T. Busing, Allen M. Solomon, and Connie A. Burdick. A multi-model framework for simulating wildlife population response to land-use and climate change. Ecological Modelling, 219(1–2):77–91, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004018.

Mouton:2008:OFP

[MSP^{+}08] Ans M. Mouton, Matthias Schneider, Armin Peter, Georg Holzer, Rudolf Müller, Peter L. M. Goethals, and Niels De Pauw. Optimisation of a fuzzy physical habitat model for spawning European grayling (Thymallus thymallus L.) in the Aare river (Thun, Switzerland). Ecological Modelling, 215(1–3):122–132, July 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000859.

McDonald:2002:INS

[MSS02] A. David McDonald, Leif K. Sandal, and Stein Ivar Steinshamn. Implications of a nested stochastic/deterministic bio-economic model for a pelagic fishery. Ecological Modelling, 149(1–2):193–201, March 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005233.

Mizumukai:2008:NSE

[MSTK08] Kentaro Mizumukai, Toru Sato, Sigeru Tabeta, and Daisuke Kitazawa. Numerical studies on ecological effects of artificial mixing of surface and bottom waters in density stratification in semi-enclosed bay and open sea. Ecological Modelling, 214(2–4):251–270, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000793.

Moustakas:2009:SPD

[MSW^{+}09] Aristides Moustakas, Konstantinos Sakkos, Kerstin Wiegand, David Ward, Katrin M. Meyer, and Dirk Eisinger. Are savannas patch-dynamic systems? A landscape model. Ecological Modelling, 220(24):3576–3588, December 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009004426.

Matsinos:2002:MCD

[MT02] Yiannis G. Matsinos and Andreas Y. Troumbis. Modeling competition, dispersal and effects of disturbance in the dynamics of a grassland community using a cellular automaton model. Ecological Modelling, 149(1–2):71–83, March 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001005154.

Marin:2009:SDE

[MTD^{+}09] Víctor H. Marín, Antonio Tironi, Luisa E. Delgado, Manuel Contreras, Fernando Novoa, Marcela Torres-Gómez, René Garreaud, Irma Vila, and Italo Serey. On the sudden disappearance of Egeria densa from a Ramsar wetland site of Southern Chile: a climatic event trigger model. Ecological Modelling, 220(15):1752–1763, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002798.

Muller:2006:CCA

[MTKM^{+}06] Torsten Müller, Kristian Thorup-Kristensen, Jakob Magid, Lars Stoumann Jensen, and Søren Hansen. Catch crops affect nitrogen dynamics in organic farming systems without livestock husbandry — simulations with the DAISY model. Ecological Modelling, 191(3–4):538–544, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002838.

Moreira:2005:SAM

[MTVC05] Davidson M. Moreira, Tiziano Tirabassi, Marco T. Vilhena, and Jonas C. Carvalho. A semi-analytical model for the tritium dispersion simulation in the PBL from the Angra I nuclear power plant. Ecological Modelling, 189(3–4):413–424, December 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002164.

Mamedov:2002:CTD

[MU02] Alexandre Mamedov and Sergey Udalov. A computer tool to develop individual-based models for simulation of population interactions. Ecological Modelling, 147(1):53–68, January 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001004069.

Mulder:2007:MDN

[Mul07] Kenneth Mulder. Modeling the dynamics of nutrient limited consumer populations using constant elasticity production functions. Ecological Modelling, 207(2–4):319–326, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007002943.

Munoz:2009:DBE

[Mun09] François Munoz. Distance-based eigenvector maps (DBEM) to analyse metapopulation structure with irregular sampling. Ecological Modelling, 220(20):2683–2689, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004980.

Murray:2001:USM

[Mur01] A. G. Murray. The use of simple models in the design and calibration of a dynamic 2D model of a semi-enclosed Australian bay. Ecological Modelling, 136(1):15–30, January 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003756.

Murray:2006:MEI

[Mur06] Alexander G. Murray. A model of the emergence of infectious pancreatic necrosis virus in Scottish salmon farms 1996–2003. Ecological Modelling, 199(1):64–72, November 1, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002808.

Moreno-Valdez:2000:SMM

[MVGH00] Arnulfo Moreno-Valdez, William E. Grant, and Rodney L. Honeycutt. A simulation model of Mexican long-nosed bat (Leptonycteris nivalis) migration. Ecological Modelling, 134(2–3):117–127, October 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002532.

Maina:2008:MSC

[MVMA08] Joseph Maina, Valentijn Venus, Timothy R. McClanahan, and Mebrahtu Ateweberhan. Modelling susceptibility of coral reefs to environmental stress using remote sensing data and GIS models. Ecological Modelling, 212(3–4):180–199, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005285. See com-

ments [GB08, Dun08].

Martinez-Vilalta:2002:HMP

[MVPB02] Jordi Martínez-Vilalta, Josep Piñol, and Keith Beven. A hydraulic model to predict drought-induced mortality in woody plants: an application to climate change in the Mediterranean. Ecological Modelling, 155(2–3):127–147, October 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000200025X.

Mulderij:2007:MPI

[MVV07] Gabi Mulderij, Egbert H. Van Nes, and Ellen Van Donk. Macrophyte–phytoplankton interactions: the relative importance of allelopathy versus other factors. Ecological Modelling, 204(1–2):85–92, May 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006600.

Maffei:2005:DPC

[MVZM05] Carlotta Maffei, Daria Vagaggini, Paola Zarattini, and Graziella Mura. The dormancy problem for Crustacea Anostraca: a rigorous model connecting hatching strategies and environmental conditions. Ecological Modelling, 185(2–4):469–481, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000128.

Matsinos:2003:IOM

[MW03] Yiannis G. Matsinos and W. F. Wolff. An individualoriented model for ecological risk assessment of wading birds. Ecological Modelling, 170(2–3):471–478, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002461.

Muller:2005:LNS

[MWD05] Johannes Müller, Peter Wernecke, and Wulf Diepenbrock. LEAFC3-N: a nitrogen-sensitive extension of the CO_{2} and H_{2}O gas exchange model LEAFC3 parameterised and tested for winter wheat (Triticum aestivum L.). Ecological Modelling, 183(2–3):183–210, April 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004612.

Morrison:2000:RIS

[MWH00] Heather A. Morrison, D. Michael Whittle, and G. Douglas Haffner. The relative importance of species invasions and sediment disturbance in regulating chemical dynamics in western Lake Erie. Ecological Modelling, 125(2–3):279–294, January 15, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001891.

Malanson:2007:EPS

[MWK07] George P. Malanson, Qian Wang, and John A. Kupfer. Ecological processes and spatial patterns before, during and after simulated deforestation. Ecological Modelling, 202(3–4):397–409, April 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005424.

Marin:2002:SAM

[MWM02] Sandra L. Marín, Renato Westermeier, and Jorge Melipillán. Simulation of alternative management strategies for red algae, luga roja, (Gigartina skottsbergii Setchell and Gardner) in southern Chile. Ecological Modelling, 154(1–2):121–133, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000546.

Morales:2006:MDM

[MWMN06] Y. Morales, L. J. Weber, A. E. Mynett, and T. J. Newton. Mussel dynamics model: a hydroinformatics tool for analyzing the effects of different stressors on the dynamics of freshwater mussel communities. Ecological Modelling, 197(3–4):448–460, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001426.

Mueller:2007:SVB

[MWP07] Eva Nora Mueller, John Wainwright, and Anthony J. Parsons. The stability of vegetation boundaries and the propagation of desertification in the American southwest: a modelling approach. Ecological Modelling, 208 (2–4):91–101, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007002347.

Meyer:2007:SSS

[MWWM07] Katrin M. Meyer, Kerstin Wiegand, David Ward, and Aristides Moustakas. SATCHMO: a spatial simulation model of growth, competition, and mortality in cycling savanna patches. Ecological Modelling, 209(2–4):377–391, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003547.

Madon:2001:IMA

[MWWZ01] Sharook P. Madon, Gregory D. Williams, Janelle M. West, and Joy B. Zedler. The importance of marsh access to growth of the California killifish, Fundulus parvipinnis, evaluated through bioenergetics modeling. Ecological Modelling, 136(2–3):149–165, January 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004166.

Matsushita:2004:ERN

[MXC^{+}04] Bunkei Matsushita, Ming Xu, Jin Chen, Satoshi Kameyama, and Masayuki Tamura. Estimation of regional net primary productivity (NPP) using a process-based ecosystem model: How important is the accuracy of climate data? Ecological Modelling, 178(3–4):371–388, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001553.

Miller:2002:MFS

[MY02a] Jay D. Miller and Stephen R. Yool. Modeling fire in semidesert grassland/oak woodland: the spatial implications. Ecological Modelling, 153(3):229–245, August 1, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002000157.

Morita:2002:PVS

[MY02b] Kentaro Morita and Akira Yokota. Population viability of stream-resident salmonids after habitat fragmentation: a case study with white-spotted charr (Salvelinus leucomaenis) by an individual based model. Ecological Modelling, 155(1):85–94, September 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200128X.

Miki:2005:EAF

[MY05] Takeshi Miki and Norio Yamamura. Effects of asynchronous fluctuations in DOC supply and bacterial growth on biodegradation efficiency. Ecological Modelling, 183 (2–3):281–299, April 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004843.

Morales-Zarate:2004:ETS

[MZASLMLC04] M. V. Morales-Zárate, F. Arreguín-Sánchez, J. López-Martínez, and S. E. Lluch-Cota. Ecosystem trophic structure and energy flux in the Northern Gulf of California, México. Ecological Modelling, 174(4):331–345, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003004320.

Mi:2005:TGA

[MZWM05] Xiangcheng Mi, Yingbin Zou, Wei Wei, and Keping Ma. Testing the generalization of artificial neural networks with cross-validation and independent-validation in modelling rice tillering dynamics. Ecological Modelling, 181 (4):493–508, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003746.

Neira:2004:CAT

[NAC04] Sergio Neira, Hugo Arancibia, and Luis Cubillos. Comparative analysis of trophic structure of commercial fishery species off Central Chile in 1992 and 1998. Ecological Modelling, 172(2–4):233–248, March 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003715.

Nadarajah:2007:CUG

[Nad07] Saralees Nadarajah. Comments on “Using the generalized F distribution to model limnetic temperature profile and estimate thermocline depth”. Ecological Modelling, 204(1–2):269, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006005783.

Nakayama:2008:FCV

[Nak08] Tadanobu Nakayama. Factors controlling vegetation succession in kushiro mire. Ecological Modelling, 215(1–3):225– 236, July 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000800104X.

Nally:2000:MCE

[Nal00] Ralph Mac Nally. Modelling confinement experiments in community ecology: differential mobility among competitors. Ecological Modelling, 129(1):65–85, May 3, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000002167.

Nally:2001:ISS

[Nal01] Ralph Mac Nally. Interaction strengths and spatial scale in community ecology: simulated quadrat-sampling and confinement experiments involving animals of different mobilities. Ecological Modelling, 144(2–3):139–152, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003672.

Nemeth:2005:DDE

[NBP05] Erwin Nemeth, Peter Bossew, and Christoph Plutzar. A distance-dependent estimation of foraging ranges of neighbouring bird colonies. Ecological Modelling, 182 (1):67–73, February 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004004065.

Niccolucci:2009:HDF

[NBT^{+}09] V. Niccolucci, S. Bastianoni, E. B. P. Tiezzi, M. Wackernagel, and N. Marchettini. How deep is the footprint? A 3D representation. Ecological Modelling, 220 (20):2819–2823, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009005213.

Nishida:2007:FLA

[NCM07] Tom Nishida, Ding-Geng Chen, and Masahiko Mohri. Fuzzy logic analyses for the spawner-recruitment relationship of bigeye tuna (Thunnus obesus) in the Indian Ocean incorporating the environmental regime shift. Ecological Modelling, 203(1–2):132–140, April 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005643.

Naithani:2007:SNP

[NDD^{+}07] Jaya Naithani, François Darchambeau, Eric Deleersnijder, Jean-Pierre Descy, and Eric Wolanski. Study of the nutrient and plankton dynamics in Lake Tanganyika using a reducedgravity model. Ecological Modelling, 200(1–2):225–233, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006003656.

Nerini:2000:AOR

[NDM00] David Nerini, Jean Pierre Durbec, and Claude Manté. Analysis of oxygen rate time series in a strongly polluted lagoon using a regression tree method. Ecological Modelling, 133(1–2):95–105, August 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002830.

Nedorezov:2009:CRD

[Ned09] Lev V. Nedorezov. Comments regarding Deng, B. 2008. The Time Invariance Principle, the absence of ecological chaos, and a fundamental pitfall of discrete modeling, Ecological Modelling 215, 287–292. Ecological Modelling, 220(19):2598–2599, October 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003986.

Newton:2006:FPM

[New06] Peter F. Newton. Forest production model for upland black spruce stands — optimal site occupancy levels for maximizing net production. Ecological Modelling, 190 (1–2):190–204, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002371.

Newton:2009:DID

[New09] Peter F. Newton. Development of an integrated decisionsupport model for density management within jack pine stand-types. Ecological Modelling, 220(23):3301–3324, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380009005237.

Neuhauser:2004:MPC

[NF04] Claudia Neuhauser and Joseph E. Fargione. A mutualism– parasitism continuum model and its application to plant– mycorrhizae interactions. Ecological Modelling, 177(3– 4):337–352, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001450.

Negi:2007:DBD

[NG07] Kuldeep Negi and Sunita Gakkhar. Dynamics in a Beddington–DeAngelis prey–predator system with impulsive harvesting. Ecological Modelling, 206(3–4):421–430, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007002219.

Ng:2009:TBA

[NG09] C. A. Ng and K. A. Gray. Tracking bioaccumulation in aquatic organisms: a dynamic model integrating life history characteristics and environmental change. Ecological Modelling, 220(9–10):1266–1273, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000945.

Nuttle:2007:DVS

[NH07] Tim Nuttle and James W. Haefner. Design and validation of a spatially explicit simulation model for bottomland hardwood forests. Ecological Modelling, 200(1–2):20–32, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000600336X.

Nguyen-Huu:2006:EMF

[NHLPA06] T. Nguyen-Huu, C. Lett, J. C. Poggiale, and P. Auger. Effect of movement frequency on global host–parasitoid spatial dynamics with unstable local dynamics. Ecological Modelling, 197(3–4):290–295, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001487.

Ni:2006:ICV

[NHP^{+}06] Jian Ni, Sandy P. Harrison, I. Colin Prentice, John E. Kutzbach, and Stephen Sitch. Impact of climate variability on present and Holocene vegetation: a model-based study. Ecological Modelling, 191(3–4):469–486, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002851.

Nitschke:2008:TCA

[NI08] Craig R. Nitschke and John L. Innes. A tree and climate assessment tool for modelling ecosystem response to climate change. Ecological Modelling, 210(3):263–277, January 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007004061.

Nielsen:2000:NEL

[Nie00a] S. N. Nielsen. Numerical Ecology. Legendre P. and Legendre L., second ed., Elsevier, Amsterdam, p. 853, 1998. Ecological Modelling, 132(3):303–304, August 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000291X.

Nielsen:2000:TEI

[Nie00b] S. N. Nielsen. Thermodynamics of an ecosystem interpreted as a hierarchy of embedded systems. Ecological Modelling, 135(2–3):279–289, December 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003793.

Nielsen:2006:EPL

[Nie06] Søren Nors Nielsen. Ecological Paradigms Lost. Routes of Theory Change, K. Cuddington, B. Beisner (Eds.). Elsevier Academic Press (2005), 900 references, Paperback, Price: US $79.95, EUR 74.95, £49.95, ISBN: 0-12-088459-3. Ecological Modelling, 199(1):125–126, November 1, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006003103.

Nielsen:2008:BRT

[Nie08a] Søren Nors Nielsen. Book review: Topics in Mathematical Modelling, K. K. Tung. Princeton University Press (2007), 300 pages, £29.95. Ecological Modelling, 211 (1–2):247–248, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004309.

Nieminen:2008:BMU

[Nie08b] Jouni K. Nieminen. Is bad mathematics useless? Ecological Modelling, 213(3–4):468–470, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000562.

Nieminen:2008:LFW

[Nie08c] Jouni K. Nieminen. Linking food webs to ecosystem processes: Piecewise linear models of soil microcosms. Ecological Modelling, 217(1–2):87–94, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002895.

Nielsen:2009:TMF

[Nie09a] Søren Nors Nielsen. Teaching models in future. Ecological Modelling, 220(23):3475–3476, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005225.

Nieminen:2009:EPD

[Nie09b] Jouni K. Nieminen. Enchytraeid population dynamics: Resource limitation and size-dependent mortality. Ecological Modelling, 220(11):1425–1430, June 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001847.

Nakagiri:2004:IEH

[NiT04] Nariyuki Nakagiri and Kei ichi Tainaka. Indirect effects of habitat destruction in model ecosystems. Ecological Modelling, 174(1–2):103–114, May 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005714.

Nakagiri:2001:IRB

[NiTT01] Nariyuki Nakagiri, Kei ichi Tainaka, and Tomomi Tao. Indirect relation between species extinction and habitat destruction. Ecological Modelling, 137(2–3):109–118, February 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004178.

Niggemann:2009:DPP

[NJB^{+}09] Marc Niggemann, Jens Jetzkowitz, Stefan Brunzel, Matthias C. Wichmann, and Ronald Bialozyt. Distribution patterns of plants explained by human movement behaviour. Ecological Modelling, 220(9–10):1339–1346, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001513.

Norman:2008:SNE

[NJF^{+}08] Josefine Norman, Per-Erik Jansson, Neda Farahbakhshazad, Klaus Butterbach-Bahl, Changsheng Li, and Leif Klemedtsson. Simulation of NO and N_{2}O emissions from a spruce forest during a freeze/thaw event using an N-flux submodel from the PnET-N-DNDC model integrated to CoupModel. Ecological Modelling, 216(1):18–30, August 10, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008002019.

Nendel:2004:SMA

[NK04] C. Nendel and K. C. Kersebaum. A simple model approach to simulate nitrogen dynamics in vineyard soils. Ecological Modelling, 177(1–2):1–15, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001048.

Nagendra:2006:ANN

[NK06] S. M. Shiva Nagendra and Mukesh Khare. Artificial neural network approach for modelling nitrogen dioxide dispersion from vehicular exhaust emissions. Ecological Modelling, 190(1–2):99–115, January 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002280.

Neubert:2004:RTD

[NKC04] Michael G. Neubert, Tin Klanjscek, and Hal Caswell. Reactivity and transient dynamics of predator–prey and food web models. Ecological Modelling, 179(1):29–38, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002509.

Navratil:2007:ARS

[NKK^{+}07] T. Navrátil, D. Kurz, P. Krám, J. Hofmeister, and J. Hruška. Acidification and recovery of soil at a heavily impacted forest catchment (Lysina, Czech Republic) — SAFE modeling and field results. Ecological Modelling, 205(3–4):464–474, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001263.

Nikolaidis:2009:MNE

[NKK^{+}09] Nikolaos P. Nikolaidis, Aristomenis P. Karageorgis, Vasilios Kapsimalis, Paraskevi Drakopoulou, Nikolaos Skoulikidis, Horst Behrendt, and Zlatko Levkov. Management of nutrient emissions of Axios River catchment: Their effect in the coastal zone of Thermaikos Gulf, Greece. Ecological Modelling, 220(3):383–396, February 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004638.

Nielsen:2009:UFP

[NM09] S. N. Nielsen and F. Müller. Understanding the functional principles of nature — proposing another type of ecosystem services. Ecological Modelling, 220(16):1913–1925, August 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002749.

Nadporozhskaya:2006:DSO

[NMC^{+}06] M. A. Nadporozhskaya, G. M. J. Mohren, O. G. Chertov, A. S. Komarov, and A. V. Mikhailov. Dynamics of soil organic matter in primary and secondary forest succession on sandy soils in the Netherlands: an application of the RO-MUL model. Ecological Modelling, 190(3–4):399–418, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002619.

Naud:2007:AIP

[NMJ07] Cédric Naud, David Makowski, and Marie-Hélène Jeuffroy. Application of an interacting particle filter to improve nitrogen nutrition index predictions for winter wheat. Ecological Modelling, 207(2–4):251–263, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002773.

Na:2006:HWQ

[NP06] Eun Hye Na and Seok Soon Park. A hydrodynamic and water quality modeling study of spatial and temporal patterns of phytoplankton growth in a stratified lake with buoyant incoming flow. Ecological Modelling, 199(3):298–314, December 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006002316.

Nunes:2007:SER

[NS07] Carla Nunes and Amilcar Soares. Stochastic environmental research risk assessment. Ecological Modelling, 201(2):118–126, February 24, 2007. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000600411X.

Nedorezov:2008:GOL

[NS08] Lev V. Nedorezov and Dinara L. Sadykova. Green oak leaf roller moth dynamics: an application of discrete time mathematical models. Ecological Modelling, 212(1–2):162–170, March 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005078.

Nour:2006:AAN

[NSEDP06] Mohamed H. Nour, Daniel W. Smith, Mohamed Gamal El-Din, and Ellie E. Prepas. The application of artificial neural networks to flow and phosphorus dynamics in small streams on the boreal plain, with emphasis on the role of wetlands. Ecological Modelling, 191(1):19–32, January 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003467.

Nagaya:2008:ESH

[NSO^{+}08] Takayuki Nagaya, Yoshiki Shiraishi, Kouki Onitsuka, Makoto Higashino, Tohru Takami, Noriharu Otsuka, Juichiro Akiyama, and Hiroaki Ozeki. Evaluation of suitable hydraulic conditions for spawning of ayu with horizontal 2D numerical simulation and PHABSIM. Ecological Modelling, 215(1–3):133–143, July 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000860.

Noble:2007:NEU

[NT07] Peter A. Noble and Erik H. Tribou. Neuroet: an easy-to-use artificial neural network for ecological and biological modeling. Ecological Modelling, 203(1–2):87–98, April 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006005606.

Nielsen:2000:CBT

[NU00] Søren N. Nielsen and Robert E. Ulanowicz. On the consistency between thermodynamical and network approaches to ecosystems. Ecological Modelling, 132(1–2):23–31, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003021.

Nungesser:2003:MMB

[Nun03] Martha K. Nungesser. Modelling microtopography in boreal peatlands: hummocks and hollows. Ecological Modelling, 165(2–3):175–207, July 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300067X.

Novak:2003:TNU

[NV03] Viliam Novák and Jozef Vidovič. Transpiration and nutrient uptake dynamics in maize (Zea mays L.). Ecological Modelling, 166(1–2):99–107, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001029.

Nijboer:2004:VSM

[NV04] Rebi C. Nijboer and Piet F. M. Verdonschot. Variable selection for modelling effects of eutrophication on stream and river ecosystems. Ecological Modelling, 177 (1–2):17–39, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000400105X.

Nalder:2006:MIL

[NW06] Ian Anthony Nalder and Ross W. Wein. A model for the investigation of long-term carbon dynamics in boreal forests of western Canada: I. Model development and validation. Ecological Modelling, 192(1–2):37–66, February 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000500342X.

Nehrbass:2007:GHS

[NW07] Nana Nehrbass and Eckart Winkler. Is the Giant Hogweed still a threat? An individual-based modelling approach for local invasion dynamics of Heracleum mantegazzianum. Ecological Modelling, 201(3–4):377–384, March 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006004856.

Nogueira:2006:PCE

[NWH^{+}06] E. Nogueira, J. D. Woods, C. Harris, A. J. Field, and S. Talbot. Phytoplankton co-existence: Results from an individual-based simulation model. Ecological Modelling, 198(1–2):1–22, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001670.

Na:2006:MME

[NWP06] Yu-Mee Na, Sookyun Wang, and Seok Soon Park. A mathematical model to estimate nitrate release from ocher pellets applied to anaerobic benthic sediment. Ecological Modelling, 199(3):324–335, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600233X.

Ogawa:2007:RBF

[OAAF07] Kazuharu Ogawa, Ahmad Makmom Abdullah, Muhamad Awang, and Akio Furukawa. Relationship between fruit growth and peduncle cross-sectional area in durian (Durio zibethinus Murray). Ecological Modelling, 200(1– 2):254–258, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006003784.

Ortiz:2009:SMB

[OACB09] Marco Ortiz, Miguel Avendaño, Leonardo Campos, and Fernando Berrios. Spatial and mass balanced trophic models of La Rinconada Marine Reserve (SE Pacific coast), a protected benthic ecosystem: Management strategy assessment. Ecological Modelling, 220(23):3413–3423, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006012.

Oh:2007:CPI

[OAL^{+}07] Hee-Mock Oh, Chi-Yong Ahn, Jae-Won Lee, Tae-Soo Chon, Kyung Hee Choi, and Young-Seuk Park. Community patterning and identification of predominant factors in algal bloom in Daechung Reservoir (Korea) using artificial neural networks. Ecological Modelling, 203(1–2):109–118, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000600562X.

Osidele:2004:FWM

[OB04] O. O. Osidele and M. B. Beck. Food web modelling for investigating ecosystem behaviour in large reservoirs of the southeastern United States: lessons from Lake Lanier, Georgia. Ecological Modelling, 173(2–3):129–158, April 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003003831.

Ogle:2007:EBA

[OBE^{+}07] Stephen M. Ogle, F. Jay Breidt, Mark Easter, Steve Williams, and Keith Paustian. An empirically based approach for estimating uncertainty associated with modelling carbon sequestration in soils. Ecological Modelling, 205(3–4):453–463, July 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001251.

Omlin:2001:BMLb

[OBR01] Martin Omlin, Roland Brun, and Peter Reichert. Biogeochemical model of Lake Zürich: sensitivity, identifiability and uncertainty analysis. Ecological Modelling, 141(1–3):105–123, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002575.

Oborny:2001:EER

[OCK01] Beáta Oborny, Tamás Czárán, and Ádám Kun. Exploration and exploitation of resource patches by clonal growth: a spatial model on the effect of transport between modules. Ecological Modelling, 141(1–3):151–169, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002708.

Overmars:2003:SAM

[OdKV03] K. P. Overmars, G. H. J. de Koning, and A. Veldkamp. Spatial autocorrelation in multi-scale land use models. Ecological Modelling, 164(2–3):257–270, June 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300070X.

Odum:2002:EER

[Odu02] Howard T. Odum. Explanations of ecological relationships with energy systems concepts. Ecological Modelling, 158(3):201–211, December 31, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002326.

Odum:2004:EAS

[Odu04] Elisabeth Odum. Emergy analysis of shrimp mariculture in Ecuador: a review. Ecological Modelling, 178(1–2):239–240, October 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005453.

Owega:2006:ACT

[OEK^{+}06] Sandy Owega, Greg J. Evans, Badi-Uz-Zaman Khan, Robert E. Jervis, and Mike Fila. Application of a counting technique to determine certain and uncertain geographic regions of emission sources. Ecological Modelling, 192(3–4):627–636, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003704.

Ozaki:2008:SEA

[OFK08] Noriatsu Ozaki, Takehiko Fukushima, and Toshiharu Kojiri. Simulation of the effects of the alteration of the river basin land use on river water temperature using the multi-layer mesh-typed runoff model. Ecological Modelling, 215(1–3):159–169, July 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000999.

OCallaghan:2008:IBS

[OG08] Michael O’Callaghan and G. N. G. Gordon. An individualbased stochastic hazard model of eastern king prawn (Melicertus plebejus) migration with spatially and temporally varying fishing effort. Ecological Modelling, 218 (1–2):73–94, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008003116.

Ogawa:2009:CTG

[Oga09a] Kazuharu Ogawa. Consideration of translocation into a growth model of a plant organ. Ecological Modelling, 220(12):1492–1494, June 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002014.

Ogawa:2009:MAC

[Oga09b] Kazuharu Ogawa. Mathematical analysis of change in forest carbon use efficiency with stand development: a case study on Abies veitchii lindl. Ecological Modelling, 220(11):1419–1424, June 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001835.

Ostendorf:2001:ECC

[OHH01] Bertram Ostendorf, David W. Hilbert, and Mike S. Hopkins. The effect of climate change on tropical rainforest vegetation pattern. Ecological Modelling, 145(2–3):211–224, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003921.

Oulehle:2007:MLT

[OHH07] Filip Oulehle, Jeňýk Hofmeister, and Jakub Hruška. Modeling of the long-term effect of tree species (Norway spruce and European beech) on soil acidification in the Ore Mountains. Ecological Modelling, 204(3–4):359–371, June 16, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007000282.

Ooba:2006:CGF

[OHM^{+}06] Makoto Ooba, Takashi Hirano, Jun-Ichi Mogami, Ryuichi Hirata, and Yasumi Fujinuma. Comparisons of gap-filling methods for carbon flux dataset: a combination of a genetic algorithm and an artificial neural network. Ecological Modelling, 198(3–4):473–486, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002766.

Obioh:2007:PVA

[OI07] G. I. B. Obioh (née G.U.Ahuama) and A. O. Isichei. A population viability analysis of serendipity berry (Dioscoreophyllum cumminsii) in a semi-deciduous forest in Nigeria. Ecological Modelling, 201(3–4):558–562, March 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006004947.

Olchev:2008:ELU

[OIP^{+}08] A. Olchev, A. Ibrom, J. Priess, S. Erasmi, C. Leemhuis, A. Twele, K. Radler, H. Kreilein, O. Panferov, and G. Gravenhorst. Effects of land-use changes on evapotranspiration of tropical rain forest margin area in Central Sulawesi (Indonesia): Modelling study with a regional SVAT model. Ecological Modelling, 212(1–2):131–137, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005054.

Olchev:2008:MAS

[OIR^{+}08] A. Olchev, A. Ibrom, T. Ross, U. Falk, G. Rakkibu, K. Radler, S. Grote, H. Kreilein, and G. Gravenhorst. A modelling approach for simulation of water and carbon dioxide exchange between multi-species tropical rain forest and the atmosphere. Ecological Modelling, 212(1–2):122–130, March 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005042.

Olden:2002:IBB

[OJ02] Julian D. Olden and Donald A. Jackson. Illuminating the “black box”: a randomization approach for understanding variable contributions in artificial neural networks. Ecological Modelling, 154(1–2):135–150, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000649.

Olden:2004:ACM

[OJD04] Julian D. Olden, Michael K. Joy, and Russell G. Death. An accurate comparison of methods for quantifying variable importance in artificial neural networks using simulated data. Ecological Modelling, 178(3–4):389–397, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004001565.

Okey:2004:SCE

[Oke04a] T. Okey. Simulating community effects of sea floor shading by plankton blooms over the West Florida Shelf. Ecological Modelling, 172(2–4):339–359, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003776.

Okey:2004:TMG

[Oke04b] T. Okey. A trophic model of a Galápagos subtidal rocky reef for evaluating fisheries and conservation strategies. Ecological Modelling, 172(2–4):383–401, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003818.

Okuyama:2009:LIB

[Oku09] Toshinori Okuyama. Local interactions between predators and prey call into question commonly used functional responses. Ecological Modelling, 220(9–10):1182–1188, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009000969.

Overholtz:2009:SME

[OL09] William Overholtz and Jason Link. A simulation model to explore the response of the Gulf of Maine food web to large-scale environmental and ecological changes. Ecological Modelling, 220(19):2491–2502, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004463.

Ottaviani:2004:TSM

[OLB04] Daniela Ottaviani, Giovanna Jona Lasinio, and Luigi Boitani. Two statistical methods to validate habitat suitability models using presence-only data. Ecological Modelling, 179(4):417–443, December 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002406.

Oli:2003:PLCa

[Oli03a] Madan K. Oli. Partial life-cycle analysis: a simplified model for post-breeding census data. Ecological Modelling, 168(1–2):101–108, October 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001935.

Oli:2003:PLCb

[Oli03b] Madan K. Oli. Partial life-cycle models: how good are they? Ecological Modelling, 169(2–3):313–325, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002771.

Oh:2004:MCP

[OLK^{+}04] E. J. Oh, H. W. Lee, A. Kondo, A. Kaga, and K. Yamaguchi. Micro-climate prediction in a residential development region using a numerical model. Ecological Modelling, 177(3–4):283–295, October 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001401.

Oksanen:2002:CTR

[OM02] Jari Oksanen and Peter R. Minchin. Continuum theory revisited: what shape are species responses along ecological gradients? Ecological Modelling, 157(2–3):119–129, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002001904.

Odulaja:2001:MTT

[OMA01] Adedapo Odulaja and Mohamed M. Mohamed-Ahmed. Modelling the trappability of tsetse, Glossina fuscipes fuscipes, in relation to distance from their natural habitats. Ecological Modelling, 143(3):183–189, September 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003295.

Ordonez:2006:MRP

[OMHR06] José Luis Ordóñez, Roberto Molowny-Horas, and Javier Retana. A model of the recruitment of Pinus nigra from unburned edges after large wildfires. Ecological Modelling, 197(3–4):405–417, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001505.

Oguz:2001:SPS

[OMRD01] Temel Oguz, Paola Malanotte-Rizzoli, and Hugh W. Ducklow. Simulations of phytoplankton seasonal cycle with multilevel and multi-layer physical-ecosystem models: the Black Sea example. Ecological Modelling, 144(2–3):295–314, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003787.

Onderka:2007:CBS

[Ond07] Milan Onderka. Correlations between several environmental factors affecting the bloom events of cyanobacteria in Liptovska Mara reservoir (Slovakia) — a simple regression model. Ecological Modelling, 209(2–4):412–416, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007004048.

Ozesmi:2004:EMB

[ÖÖ04] Uygar Özesmi and Stacy L. Özesmi. Ecological models based on people’s knowledge: a multi-step fuzzy cognitive mapping approach. Ecological Modelling, 176(1–2):43–64, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000300543X.

Ottaviani:2009:MSA

[OPL^{+}09] D. Ottaviani, M. Panzacchi, G. Jona Lasinio, P. Genovesi, and L. Boitani. Modelling semi-aquatic vertebrates’ distribution at the drainage basin scale: the case of the otter Lutra lutra in Italy. Ecological Modelling, 220(2):111–121, January 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008004377.

Omlin:2001:BMLa

[ORF01] Martin Omlin, Peter Reichert, and Richard Forster. Biogeochemical model of Lake Zürich: model equations and results. Ecological Modelling, 141(1–3):77–103, July 1, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001002563.

Olchev:2009:ATD

[ORS^{+}09] A. Olchev, K. Radler, A. Sogachev, O. Panferov, and G. Gravenhorst. Application of a three-dimensional model for assessing effects of small clear-cuttings on radiation and soil temperature. Ecological Modelling, 220 (21):3046–3056, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000900091X.

Ortega:2004:PWP

[Ort04] Enrique Ortega. The prosperous way down: a proposal for action. Ecological Modelling, 178(1–2):243–246, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005465.

Ortiz:2008:MBD

[Ort08] Marco Ortiz. Mass balanced and dynamic simulations of trophic models of kelp ecosystems near the Mejillones Peninsula of northern Chile (SE Pacific): Comparative network structure and assessment of harvest strategies. Ecological Modelling, 216(1):31–46, August 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001993.

Owen-Smith:2002:MMA

[OS02] Norman Owen-Smith. A metaphysiological modelling approach to stability in herbivore–vegetation systems. Ecological Modelling, 149(1–2):153–178, March 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100521X.

Omori:2009:SRB

[OSHO09] Koji Omori, Atsushi Sogabe, Hideki Hamaoka, and Hidejiro Ohnishi. Storage and the regulation of body mass in animals: a general growth equation based on an energy balance model. Ecological Modelling, 220(20):2618–2623, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009004876.

Osborne:2002:SDP

[OSS02] Patrick E. Osborne and Susana Suárez-Seoane. Should data be partitioned spatially before building large-scale distribution models? Ecological Modelling, 157(2– 3):249–259, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001989.

Ooba:2003:EAS

[OT03] Makoto Ooba and Hidenori Takahashi. Effect of asymmetric stomatal response on gas-exchange dynamics. Ecological Modelling, 164(1):65–82, June 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000127.

Ozesmi:2006:MIB

[ÖTÖ06] Stacy L. Özesmi, Can O. Tan, and Uygar Özesmi. Methodological issues in building, training, and testing artificial neural networks in ecological applications. Ecological Modelling, 195(1–2):83–93, May 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005806.

Ozesmi:2006:GAN

[ÖTÖR06] Uygar Özesmi, Can O. Tan, Stacy L. Özesmi, and Raleigh J. Robertson. Generalizability of artificial neural network models in ecological applications: Predicting nest occurrence and breeding success of the red-winged blackbird Agelaius phoeniceus. Ecological Modelling, 195(1–2):94–104, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005818.

Oli:2006:PDI

[OVK^{+}06] Madan K. Oli, Meenakshi Venkataraman, Paul A. Klein, Lori D. Wendland, and Mary B. Brown. Population dynamics of infectious diseases: a discrete time model. Ecological Modelling, 198(1–2):183–194, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001815.

Ortiz:2002:DSM

[OW02] Marco Ortiz and Matthias Wolff. Dynamical simulation of mass-balance trophic models for benthic communities of north-central Chile: assessment of resilience time under alternative management scenarios. Ecological Modelling, 148(3):277–291, March 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004549.

Olivier:2005:GBA

[OW05] Frédérique Olivier and Simon J. Wotherspoon. GIS-based application of resource selection functions to the prediction of snow petrel distribution and abundance in East Antarctica: Comparing models at multiple scales. Ecological Modelling, 189(1–2):105–129, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001924.

Olivier:2006:MHS

[OW06] Frédérique Olivier and Simon J. Wotherspoon. Modelling habitat selection using presence-only data: Case study of a colonial hollow nesting bird, the snow petrel. Ecological Modelling, 195(3–4):187–204, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005673.

Olivier:2008:NSS

[OW08] Frédérique Olivier and Simon J. Wotherspoon. Nest selection by snow petrels Pagodroma nivea in East Antarctica: Validating predictive habitat selection models at the continental scale. Ecological Modelling, 210(4):414–430, February 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007004140.

Ouyang:2009:TSD

[OWHS09] Wei Ouyang, Xuelei Wang, Fanghua Hao, and R. Srinivasan. Temporal-spatial dynamics of vegetation variation on non-point source nutrient pollution. Ecological Modelling, 220(20):2702–2713, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004475.

Obach:2001:MPD

[OWWS01] Michael Obach, Rüdiger Wagner, Heinrich Werner, and Hans-Heinrich Schmidt. Modelling population dynamics of aquatic insects with artificial neural networks. Ecological Modelling, 146(1–3):207–217, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003076.

Okunishi:2009:SMJ

[OYiI09] Takeshi Okunishi, Yasuhiro Yamanaka, and Shin ichi Ito. A simulation model for Japanese sardine (Sardinops melanostictus) migrations in the western North Pacific. Ecological Modelling, 220(4):462–479, February 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004936.

Ovadia:2007:CBS

[OzDBS07] Ofer Ovadia, Heinrich zu Dohna, Ginger Booth, and Oswald J. Schmitz. Consequences of body size variation among herbivores on the strength of plant–herbivore interactions in a seasonal environment. Ecological Modelling, 206(1–2):119–130, August 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001627.

Ouyang:2007:DDD

[OZL07] Ying Ouyang, Jia En Zhang, and Shi Ming Luo. Dynamic data driven application system: Recent development and future perspective. Ecological Modelling, 204(1–2):1–8, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006006375.

Pauer:2009:FTN

[PA09a] James J. Pauer and Martin T. Auer. Formulation and testing of a novel river nitrification model. Ecological Modelling, 220(6):857–866, March 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005929.

Perhar:2009:ESF

[PA09b] Gurbir Perhar and George B. Arhonditsis. The effects of seston food quality on planktonic food web patterns. Ecological Modelling, 220(6):805–820, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005954.

Pontier:2000:IBP

[PAdlPS00] Dominique Pontier, Pierre Auger, Rafael Bravo de la Parra, and Eva Sánchez. The impact of behavioral plasticity at individual level on domestic cat population dynamics. Ecological Modelling, 133(1–2):117–124, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002854.

Price:2006:STW

[PAF06] Bronwyn Price, Britta Allgöwer, and Andreas Fischlin. Synchrony and travelling waves of larch bud moth? Time series analysis with changing scale. Ecological Modelling, 199(4):433–441, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002870.

Parchevsky:2000:URA

[Par00] Konstantin V. Parchevsky. Using regularizing algorithms for the reconstruction of growth rate from the experimental data. Ecological Modelling, 133(1–2):107–115, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002842.

Parris:2002:MBY

[Par02] Kirsten M. Parris. More bang for your buck: the effect of caller position, habitat and chorus noise on the efficiency of calling in the spring peeper. Ecological Modelling, 156(2–3):213–224, November 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001709.

Peacor:2007:NCS

[PARH07] Scott D. Peacor, Stefano Allesina, Rick L. Riolo, and Tim S. Hunter. A new computational system, DOVE (digital organisms in a virtual ecosystem), to study phenotypic plasticity and its effects in food webs. Ecological Modelling, 205(1–2):13–28, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000518.

Phillips:2006:MEM

[PAS06] Steven J. Phillips, Robert P. Anderson, and Robert E. Schapire. Maximum entropy modeling of species geographic distributions. Ecological Modelling, 190(3–4):231–259, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000500267X.

Patten:2007:ECL

[Pat07] Bernard C. Patten. Ecodynamics: Coimplication of life and environment in a coevolving ecosphere. Ecological Modelling, 204(1–2):274–276, May 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000920.

Pawlowski:2000:CUP

[Paw00] Christopher W. Pawlowski. Comments on the use of propositional logic to examine sustainability concepts. Ecological Modelling, 127(2–3):207–219, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002070.

Porte:2002:MMF

[PB02] A. Porté and H. H. Bartelink. Modelling mixed forest growth: a review of models for forest management. Ecological Modelling, 150(1–2):141–188, April 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004768.

Palmeri:2005:MNE

[PBA05] L. Palmeri, G. Bendoricchio, and Y. Artioli. Modelling nutrient emissions from river systems and loads to the coastal zone: Po River case study, Italy. Ecological Modelling, 184(1):37–53, May 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005836.

Pichancourt:2006:AEH

[PBA06] Jean-Baptiste Pichancourt, Françoise Burel, and Pierre Auger. Assessing the effect of habitat fragmentation on population dynamics: an implicit modelling approach. Ecological Modelling, 192(3–4):543–556, February 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005003546.

Papajorgji:2004:ADS

[PBB04] Petraq Papajorgji, Howard W. Beck, and Jose Luis Braga. An architecture for developing service-oriented and component-based environmental models. Ecological Modelling, 179(1):61–76, November 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002625.

Paraiba:2002:LIF

[PBC02] Lourival Costa Paraiba, Rafael Bru, and José María Carrasco. Level IV fugacity model depending on temperature by a periodic control system. Ecological Modelling, 147(3):221–232, January 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004215.

Peinetti:2009:SMU

[PBC09] H. Raul Peinetti, Bruce W. Baker, and Michael B. Coughenour. Simulation modeling to understand how selective foraging by beaver can drive the structure and function of a willow community. Ecological Modelling, 220(7):998–1012, April 10, 2009. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009000374.

Petersen:2001:PIP

[PBCZ01] W. Petersen, L. Bertino, U. Callies, and E. Zorita. Process identification by principal component analysis of river waterquality data. Ecological Modelling, 138(1–3):193–213, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004026.

Prommer:2000:NMD

[PBD00] H. Prommer, D. A. Barry, and G. B. Davis. Numerical modelling for design and evaluation of groundwater remediation schemes. Ecological Modelling, 128(2–3):181–195, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002306.

Pachepsky:2007:CIV

[PBE^{+}07] Elizaveta Pachepsky, James L. Bown, Alistair Eberst, Ursula Bausenwein, Peter Millard, Geoff R. Squire, and John W. Crawford. Consequences of intraspecific variation for the structure and function of ecological communities. Part 2: Linking diversity and function. Ecological Modelling, 207(2–4):277–285, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700289X.

Piou:2009:PIC

[PBG09] Cyril Piou, Uta Berger, and Volker Grimm. Proposing an information criterion for individual-based models developed in a pattern-oriented modelling framework. Ecological Modelling, 220(17):1957–1967, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900324X.

Piou:2007:SCM

[PBH^{+}07] Cyril Piou, Uta Berger, Hanno Hildenbrandt, Volker Grimm, Karen Diele, and Coralie D’Lima. Simulating cryptic movements of a mangrove crab: Recovery phenomena after small scale fishery. Ecological Modelling, 205(1–2):110–122, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000609.

Picard:2007:EUT

[PBHGF07] Nicolas Picard, Avner Bar-Hen, and Sylvie Gourlet-Fleury. Estimator of upgrowth transition rates for sizeclassified matrix from small samples. Ecological Modelling, 204(1–2):59–69, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006006442.

Pitt:2003:IBM

[PBK03] William C. Pitt, Paul W. Box, and Frederick. F. Knowlton. An individual-based model of canid populations: modelling territoriality and social structure. Ecological Modelling, 166(1–2):109–121, August 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001303.

Pinnegar:2005:ARW

[PBM^{+}05] John K. Pinnegar, Julia L. Blanchard, Steve Mackinson, Robert D. Scott, and Daniel E. Duplisea. Aggregation and removal of weak-links in food-web models: system stability and recovery from disturbance. Ecological Modelling, 184(2–4):229–248, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005496.

Pringle:2008:SAE

[PBML08] M. J. Pringle, S. J. Baxter, B. P. Marchant, and R. M. Lark. Spatial analysis of the error in a model of soil nitrogen. Ecological Modelling, 211(3–4):453–467, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005005.

Pelosi:2008:WML

[PBMRE08] Céline Pelosi, Michel Bertrand, David Makowski, and Jean Roger-Estrade. WORMDYN: a model of Lumbricus terrestris population dynamics in agricultural fields. Ecological Modelling, 218(3–4):219–234, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003293.

Pastres:2004:TRP

[PBPZ04] R. Pastres, D. Brigolin, A. Petrizzo, and M. Zucchetta. Testing the robustness of primary production models in shallow coastal areas: a case study. Ecological Modelling, 179(2):221–233, November 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002601.

Paradis:2002:MLS

[PBS02] Emmanuel Paradis, Stephen R. Baillie, and William J. Sutherland. Modeling large-scale dispersal distances. Ecological Modelling, 151(2–3):279–292, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004872.

Porphyre:2005:MAM

[PBS05] T. Porphyre, D. J. Bicout, and P. Sabatier. Modelling the abundance of mosquito vectors versus flooding dynamics. Ecological Modelling, 183(2–3):173–181, April 25, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004004387.

Pastor-Barcenas:2005:USA

[PBSOMG^{+}05] O. Pastor-Bárcenas, E. Soria-Olivas, J. D. Martín-Guerrero, G. Camps-Valls, J. L. Carrasco-Rodríguez, and S. del Valle-Tascón. Unbiased sensitivity analysis and pruning techniques in neural networks for surface ozone modelling. Ecological Modelling, 182(2):149–158, March 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004107.

Pinol:2005:MEF

[PBV05] J. Piñol, K. Beven, and D. X. Viegas. Modelling the effect of fire-exclusion and prescribed fire on wildfire size in Mediterranean ecosystems. Ecological Modelling, 183(4):397–409, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004004995.

Phillips:2003:IBS

[PBY^{+}03] Paul D. Phillips, T. E. Brash, Irsyal Yasman, Prapto Subagyo, and P. R. van Gardingen. An individualbased spatially explicit tree growth model for forests in East Kalimantan (Indonesian Borneo). Ecological Modelling, 159(1):1–26, January 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001266.

Park:2007:BIM

[PC07] Young-Seuk Park and Tae-Soo Chon. Biologically-inspired machine learning implemented to ecological informatics. Ecological Modelling, 203(1–2):1–7, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005527.

Pinol:2007:CUE

[PCB07] Josep Piñol, Marc Castellnou, and Keith J. Beven. Conditioning uncertainty in ecological models: Assessing the impact of fire management strategies. Ecological Modelling, 207(1):34–44, September 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001536.

Pittman:2007:PMF

[PCC^{+}07] S. J. Pittman, J. D. Christensen, C. Caldow, C. Menza, and M. E. Monaco. Predictive mapping of fish species richness across shallow-water seascapes in the Caribbean. Ecological Modelling, 204(1–2):9–21, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006338.

Park:2003:AAN

[PCCL03] Young-Seuk Park, Régis Céréghino, Arthur Compin, and Sovan Lek. Applications of artificial neural networks for patterning and predicting aquatic insect species richness in running waters. Ecological Modelling, 160(3):265–280, February 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002582.

Pedersen:2003:ESG

[PCHG03] E. K. Pedersen, J. W. Connelly, J. R. Hendrickson, and W. E. Grant. Effect of sheep grazing and fire on sage grouse populations in southeastern Idaho. Ecological Modelling, 165(1):23–47, July 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003824.

Porcasi:2005:USD

[PCL^{+}05] X. Porcasi, G. Calderón, M. Lamfri, N. Gardenal, J. Polop, M. Sabattini, and C. M. Scavuzzo. The use of satellite data in modeling population dynamics and prevalence of infection in the rodent reservoir of junin virus. Ecological Modelling, 185(2–4):437–449, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000104.

Plus:2003:MSD

[PCM^{+}03] Martin Plus, Annie Chapelle, Alain Ménesguen, Jean-Marc Deslous-Paoli, and Isabelle Auby. Modelling seasonal dynamics of biomasses and nitrogen contents in a seagrass meadow (Zostera noltii hornem.): application to the Thau lagoon (French Mediterranean coast). Ecological Modelling, 161(3):213–238, March 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003502.

Pitchford:2007:USF

[PCP07] Jonathan W. Pitchford, Edward A. Codling, and Despina Psarra. Uncertainty and sustainability in fisheries and the benefit of marine protected areas. Ecological Modelling, 207(2–4):286–292, October 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002906.

Pal:2009:RCP

[PCpDC09] S. Pal, Samrat Chatterjee, Krishna pada Das, and J. Chattopadhyay. Role of competition in phytoplankton population for the occurrence and control of plankton bloom in the presence of environmental fluctuations. Ecological Modelling, 220(2):96–110, January 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004651.

Pierson:2001:PBH

[PCS01] F. B. Pierson, D. H. Carlson, and K. E. Spaeth. A process-based hydrology submodel dynamically linked to the plant component of the simulation of production and utilization on rangelands SPUR model. Ecological Modelling, 141(1–3):241–260, July 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002770.

Pastres:2003:EKF

[PCS03] R. Pastres, S. Ciavatta, and C. Solidoro. The Extended Kalman Filter (EKF) as a tool for the assimilation of high frequency water quality data. Ecological Modelling, 170(2–3):227–235, December 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002308.

Park:2008:PAM

[PCW08] Richard A. Park, Jonathan S. Clough, and Marjorie Coombs Wellman. AQUATOX: Modeling environmental fate and ecological effects in aquatic ecosystems. Ecological Modelling, 213(1):1–15, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008000057.

Park:2006:DWQ

[PCWP06] Su-Young Park, Jung Hyun Choi, Sookyun Wang, and Seok Soon Park. Design of a water quality monitoring network in a large river system using the genetic algorithm. Ecological Modelling, 199(3):289–297, December 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006002596.

Phillips:2004:IBS

[PdAD^{+}04] P. D. Phillips, C. P. de Azevedo, B. Degen, I. S. Thompson, J. N. M. Silva, and P. R. van Gardingen. An individualbased spatially explicit simulation model for strategic forest management planning in the eastern Amazon. Ecological Modelling, 173(4):335–354, April 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300423X.

Pearson:2002:SSE

[PDBH02] R. G. Pearson, T. P. Dawson, P. M. Berry, and P. A. Harrison. SPECIES: a spatial evaluation of climate impact on the envelope of species. Ecological Modelling, 154(3):289–300, September 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200056X.

Popp:2009:LED

[PDBJ09] Alexander Popp, Stephanie Domptail, Niels Blaum, and Florian Jeltsch. Landuse experience does qualify for adaptation to climate change. Ecological Modelling, 220(5):694–702, March 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008005772.

Palmieri:2006:QMC

[PdC06] Vivian Palmieri and Roberto José de Carvalho. Qual2e model for the Corumbataí River. Ecological Modelling, 198(1–2):269–275, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001785.

Papatheodorou:2006:LTS

[PDL06] George Papatheodorou, Gerasimoula Demopoulou, and Nicolaos Lambrakis. A long-term study of temporal hydrochemical data in a shallow lake using multivariate statistical techniques. Ecological Modelling, 193(3–4):759–776, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004515.

Pivard:2008:CPO

[PDL^{+}08] Sandrine Pivard, Damjan Demšar, Jane Lecomte, Marko Debeljak, and Sašo Džeroski. Characterizing the presence of oilseed rape feral populations on field margins using machine learning. Ecological Modelling, 212(1–2):147–154, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005169.

Piqueira:2009:MCT

[PdMVN09] J. R. C. Piqueira, S. H. V. L. de Mattos, and J. Vasconcelos-Neto. Measuring complexity in three-trophic level systems. Ecological Modelling, 220(3):266–271, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005000.

Pronk:2007:CTP

[PDS07] T. E. Pronk, H. J. During, and F. Schieving. Coexistence by temporal partitioning of the available light in plants with different height and leaf investments. Ecological Modelling, 204(3–4):349–358, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000270.

Peters:2007:RFT

[PDV^{+}07] Jan Peters, Bernard De Baets, Niko E. C. Verhoest, Roeland Samson, Sven Degroeve, Piet De Becker, and Willy Huybrechts. Random forests as a tool for ecohydrological distribution modelling. Ecological Modelling, 207 (2–4):304–318, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007002931.

Perry:2002:SML

[PE02] George L. W. Perry and N. J. Enright. Spatial modelling of landscape composition and pattern in a maquis– forest complex, Mont Do, New Caledonia. Ecological Modelling, 152(2–3):279–302, July 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000042.

Potts:2006:CSA

[PE06] Joanne M. Potts and Jane Elith. Comparing species abundance models. Ecological Modelling, 199(2):153–163, November 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000600247X.

Perry:2007:COS

[PE07] George L. W. Perry and Neal J. Enright. Contrasting outcomes of spatially implicit and spatially explicit models of vegetation dynamics in a forest-shrubland mosaic. Ecological Modelling, 207(2–4):327–338, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002955.

Parry:2008:CAP

[PE08] Hazel R. Parry and Andrew J. Evans. A comparative analysis of parallel processing and super-individual methods for improving the computational performance of a large individual-based model. Ecological Modelling, 214(2–4):141–152, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008000574.

Perez-Espana:2001:IRB

[PEAS01] Horacio Pérez-España and Francisco Arreguín-Sánchez. An inverse relationship between stability and maturity in models of aquatic ecosystems. Ecological Modelling, 145(2–3):189–196, November 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003908.

Pereira:2009:FMC

[PEE09] Gilberto C. Pereira, Alexandre Evsukoff, and Nelson F. F. Ebecken. Fuzzy modelling of chlorophyll production in a Brazilian upwelling system. Ecological Modelling, 220(12):1506–1512, June 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001951.

Parry:2006:APR

[PEM06] Hazel R. Parry, Andrew J. Evans, and Derek Morgan. Aphid population response to agricultural landscape change: a spatially explicit, individual-based model. Ecological Modelling, 199(4):451–463, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002894.

Peng:2000:SBM

[Pen00] Changhui Peng. From static biogeographical model to dynamic global vegetation model: a global perspective on modelling vegetation dynamics. Ecological Modelling, 135(1):33–54, November 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000003483.

Penczak:2009:FAC

[Pen09] T. Penczak. Fish assemblage compositions after implementation of the IndVal method on the Narew River system. Ecological Modelling, 220(3):419–423, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005401.

Pinol:2009:TNM

[PEPB09] Josep Piñol, Xavier Espadaler, Nicolás Pérez, and Keith Beven. Testing a new model of aphid abundance with sedentary and non-sedentary predators. Ecological Modelling, 220(19):2469–2480, October 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004360.

Perianez:2007:COS

[Per07] R. Periáñez. Chemical and oil spill rapid response modelling in the Strait of Gibraltar–Alborán Sea. Ecological Modelling, 207(2–4):210–222, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700261X.

Peters:2002:PSD

[Pet02] Debra P. C. Peters. Plant species dominance at a grassland–shrubland ecotone: an individual-based gap dynamics model of herbaceous and woody species. Ecological Modelling, 152(1):5–32, June 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004604.

Peterson:2004:CRE

[Pet04a] Larry L. Peterson. Comments on “Relationship of energy and complexity and planning architectural design”: [Architectural Design 10 (1972) 624–629]. Ecological Modelling, 178(1–2):167–168, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005477.

Peterson:2004:ORI

[Pet04b] Larry L. Peterson. The Oregon resource interaction model: an application of systems energetics principles to decisionmaking in Oregon state government — 1974. Ecological Modelling, 178(1–2):169–170, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005489.

Peterson:2007:WWN

[Pet07] A. Townsend Peterson. Why not WhyWhere: the need for more complex models of simpler environmental spaces. Ecological Modelling, 203(3–4):527–530, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006351.

Pearce:2000:EPP

[PF00a] Jennie Pearce and Simon Ferrier. Evaluating the predictive performance of habitat models developed using logistic regression. Ecological Modelling, 133(3):225–245, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003227.

Pearce:2000:EAA

[PF00b] Jennie Pearce and Simon Ferrier. An evaluation of alternative algorithms for fitting species distribution models using logistic regression. Ecological Modelling, 128(2–3):127–147, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099002276.

Picard:2001:AIB

[PF01] Nicolas Picard and Alain Franc. Aggregation of an individual-based space-dependent model of forest dynamics into distribution-based and space-independent models. Ecological Modelling, 145(1):69–84, November 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003805.

Picard:2003:EGS

[PF03] Nicolas Picard and Alain Franc. Are ecological groups of species optimal for forest dynamics modelling? Ecological Modelling, 163(3):175–186, May 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000103.

Pontier:2008:LNS

[PFBBJ08] Dominique Pontier, David Fouchet, Joël Bried, and Nargès Bahi-Jaber. Limited nest site availability helps seabirds to survive cat predation on islands. Ecological Modelling, 214(2–4):316–324, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000847.

Provencher:2007:CAM

[PFFM07] Louis Provencher, Tara A. Forbis, Leonardo Frid, and Gary Medlyn. Comparing alternative management strategies of fire, grazing, and weed control using spatial modeling. Ecological Modelling, 209(2–4):249–263, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003432.

Parolo:2009:OTI

[PFR09] Gilberto Parolo, Alessandro Ferrarini, and Graziano Rossi. Optimization of tourism impacts within protected areas by means of genetic algorithms. Ecological Modelling, 220(8):1138–1147, April 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000416.

Plotnick:2002:GMS

[PG02] Roy E. Plotnick and Robert H. Gardner. A general model for simulating the effects of landscape heterogeneity and disturbance on community patterns. Ecological Modelling, 147(2):171–197, January 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004185.

Perez-Garcia:2004:SMN

[PG04] M. Pérez-García. Simplified modelling of the nocturnal clear sky atmospheric radiation for environmental applications. Ecological Modelling, 180(2–3):395–406, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000400290X.

Paves:2008:CFW

[PG08] Héctor J. Pavés and Humberto E. González. Carbon fluxes within the pelagic food web in the coastal area off Antofagasta (23^{o}S), Chile: the significance of the microbial versus classical food webs. Ecological Modelling, 212(3–4):218–232, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005327.

Piana:2006:CPP

[PGA06] P. A. Piana, L. C. Gomes, and A. A. Agostinho. Comparison of predator–prey interaction models for fish assemblages from the neotropical region. Ecological Modelling, 192(1–2):259–270, February 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003352.

Patterson:2004:IES

[PGCK04] Trista Patterson, Tim Gulden, Ken Cousins, and Egor Kraev. Integrating environmental, social and economic systems: a dynamic model of tourism in Dominica. Ecological Modelling, 175(2):121–136, July 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004605.

Pennanen:2004:SES

[PGFM04] Juho Pennanen, David F. Greene, Marie-Josée Fortin, and Christian Messier. Spatially explicit simulation of long-term boreal forest landscape dynamics: incorporating quantitative stand attributes. Ecological Modelling, 180(1):195–209, December 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003825.

Prokopkin:2006:MEP

[PGG06] I. G. Prokopkin, V. G. Gubanov, and M. I. Gladyshev. Modelling the effect of planktivorous fish removal in a reservoir on the biomass of cyanobacteria. Ecological Modelling, 190(3–4):419–431, January 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002681.

Peterson:2003:IFS

[PGLS03] M. Nils Peterson, William E. Grant, Roel R. Lopez, and Nova J. Silvy. Implications of fetal sex ratio hypotheses in endangered populations: simulated dynamics of Florida Key deer, Florida, USA. Ecological Modelling, 165(2–3):209–220, July 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003000711.

Pisani:2008:SIC

[PGM08] Jorge M. Pisani, William E. Grant, and Miguel A. Mora. Simulating the impact of cholinesterase-inhibiting pesticides on non-target wildlife in irrigated crops. Ecological Modelling, 210(1–2):179–192, January 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003742.

Parra-Guevara:2003:EMM

[PGS03] David Parra-Guevara and Yuri N. Skiba. Elements of the mathematical modeling in the control of pollutants emissions. Ecological Modelling, 167(3):263–275, September 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001911.

Parysow:2000:EAB

[PGW00] Pablo Parysow, George Gertner, and James Westervelt. Efficient approximation for building error budgets for process models. Ecological Modelling, 135(2–3):111–125, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003471.

Pohlert:2007:IDB

[PHBF07] T. Pohlert, J. A. Huisman, L. Breuer, and H.-G. Frede. Integration of a detailed biogeochemical model into SWAT for improved nitrogen predictions — model development, sensitivity, and GLUE analysis. Ecological Modelling, 203(3–4):215–228, May 10, 2007. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006005746.

Pontius:2004:UTV

[PHD04] Robert Gilmore Pontius, Diana Huffaker, and Kevin Denman. Useful techniques of validation for spatially explicit land-change models. Ecological Modelling, 179 (4):445–461, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002716.

Popp:2000:MEC

[PHH00] Jennie Hughes Popp, D. Eric Hyatt, and Dana Hoag. Modeling environmental condition with indices: a case study of sustainability and soil resources. Ecological Modelling, 130(1–3):131–143, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002015.

Phillips:2002:GLF

[Phi02] Jonathan D. Phillips. Global and local factors in earth surface systems. Ecological Modelling, 149(3):257–272, April 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004616.

Philomena:2004:HTO

[Phi04] Antônio L. Philomena. Howard Thomas Odum: his early work and influences in Brazil. Ecological Modelling, 178(1–2):275–278, October 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000213.

Pfenning:2004:EPC

[PHP04] Brenda Pfenning, Thomas Hovestadt, and Hans Joachim Poethke. The effect of patch constellation on the exchange of individuals between habitat-islands. Ecological Modelling, 180(4):515–522, December 31, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002984.

Politano:2008:NST

[PHW08] Marcela Politano, M. D. M. Haque, and Larry J. Weber. A numerical study of the temperature dynamics at Mc-Nary Dam. Ecological Modelling, 212(3–4):408–421, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005637.

Preuss:2009:DVI

[PHWH^{+}09] Thomas Günter Preuss, Monika Hammers-Wirtz, Udo Hommen, Mascha Nadine Rubach, and Hans Toni Ratte. Development and validation of an individual based Daphnia magna population model: the influence of crowding on population dynamics. Ecological Modelling, 220 (3):310–329, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008004602.

Pillet:2004:ECE

[Pil04] Gonzague Pillet. Emternalities as counterpart to economic externalities … ! Ecological Modelling, 178(1–2):183–187, October 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005490.

Peng:2002:EHR

[PJAZ02] Changhui Peng, Hong Jiang, Michael J. Apps, and Yanli Zhang. Effects of harvesting regimes on carbon and nitrogen dynamics of boreal forests in central Canada: a process model simulation. Ecological Modelling, 155 (2–3):177–189, October 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002001345.

Potthoff:2006:CDS

[PJGW06] Michael Potthoff, Karin Johst, Julian Gutt, and Christian Wissel. Clumped dispersal and species coexistence. Ecological Modelling, 198(1–2):247–254, September 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006001700.

Parrott:2001:GPP

[PK01] Lael Parrott and Robert Kok. A generic primary producer model for use in ecosystem simulation. Ecological Modelling, 139(1):75–99, March 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002320.

Parrott:2002:GIB

[PK02] Lael Parrott and Robert Kok. A generic, individualbased approach to modelling higher trophic levels in simulation of terrestrial ecosystems. Ecological Modelling, 154(1–2):151–178, August 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000650.

Panikkar:2008:CMB

[PK08] Preetha Panikkar and M. Feroz Khan. Comparative massbalanced trophic models to assess the impact of environmental management measures in a tropical reservoir ecosystem. Ecological Modelling, 212(3–4):280–291, April 10, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000700556X.

Park:2001:IAN

[PKC^{+}01] Young-Seuk Park, Inn-Sil Kwak, Tae-Soo Chon, Jwa-Kwan Kim, and Sven Erik Jørgensen. Implementation of artificial neural networks in patterning and prediction of exergy in response to temporal dynamics of benthic macroinvertebrate communities in streams. Ecological Modelling, 146(1–3):143–157, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003027.

Pal:2006:RSI

[PKC06] S. Pal, Kusumika Kundu, and J. Chattopadhyay. Role of standard incidence in an eco-epidemiological system: a mathematical study. Ecological Modelling, 199(3):229–239, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000600216X.

Pykh:2000:OSA

[PKG00] Yuri A. Pykh, Elizabeth T. Kennedy, and W. E. Grant. An overview of systems analysis methods in delineating environmental quality indices. Ecological Modelling, 130(1–3):25–38, June 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002118.

Post:2007:IEH

[PKS^{+}07] Joachim Post, Valentina Krysanova, Felicitas Suckow, Wilfried Mirschel, Jutta Rogasik, and Ines Merbach. Integrated eco-hydrological modelling of soil organic matter dynamics for the assessment of environmental change impacts in meso- to macro-scale river basins. Ecological Modelling, 206(1–2):93–109, August 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001731.

Peer:2008:IPR

[PKS08] Guy Pe’er and Stephanie Kramer-Schadt. Incorporating the perceptual range of animals into connectivity models. Ecological Modelling, 213(1):73–85, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006199.

Park:2002:WQM

[PL02] Seok Soon Park and Yong Seok Lee. A water quality modeling study of the Nakdong River, Korea. Ecological Modelling, 152(1):65–75, June 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004896.

Pape:2004:MST

[PL04] Roland Pape and Jörg Löffler. Modelling spatio-temporal near-surface temperature variation in high mountain landscapes. Ecological Modelling, 178(3–4):483–501, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002054.

Pasini:2006:NNM

[PLA06] Antonello Pasini, Massimo Lorè, and Fabrizio Ameli. Neural network modelling for the analysis of forcings/temperatures relationships at different scales in the climate system. Ecological Modelling, 191(1):58–67, January 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003492.

Plus:2006:MWD

[PLB^{+}06] Martin Plus, Isabelle La Jeunesse, Fayçal Bouraoui, Jose-Manuel Zaldívar, Annie Chapelle, and Pascal Lazure. Modelling water discharges and nitrogen inputs into a Mediterranean lagoon: Impact on the primary production. Ecological Modelling, 193(1–2):69–89, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004576.

Powell:2008:MFI

[PLC08] S. J. Powell, R. A. Letcher, and B. F. W. Croke. Modelling floodplain inundation for environmental flows: Gwydir wetlands, Australia. Ecological Modelling, 211 (3–4):350–362, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004784.

Peng:2002:TGH

[PLD^{+}02] Changhui Peng, Jinxun Liu, Qinglai Dang, Michael J. Apps, and Hong Jiang. TRIPLEX: a generic hybrid model for predicting forest growth and carbon and nitrogen dynamics.

Ecological Modelling, 153(1–2):109–130, July 15, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001005051.

Parviainen:2009:RLL

[PLH09] Miia Parviainen, Miska Luoto, and Risto K. Heikkinen. The role of local and landscape level measures of greenness in modelling boreal plant species richness. Ecological Modelling, 220(20):2690–2701, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004979.

Parinet:2004:PCA

[PLL04] Bernard Parinet, Antoine Lhote, and Bernard Legube. Principal component analysis: an appropriate tool for water quality evaluation and management — application to a tropical lake system. Ecological Modelling, 178 (3–4):295–311, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001498.

Pielaat:2006:SSD

[PLLdCB06] A. Pielaat, M. A. Lewis, S. Lele, and T. de Camino-Beck. Sequential sampling designs for catching the tail of dispersal kernels. Ecological Modelling, 190(1–2):205–222, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002395.

Plante:2004:SDH

[PLP^{+}04] Martin Plante, Kim Lowell, François Potvin, Barry Boots, and Marie-Josée Fortin. Studying deer habitat on Anticosti Island, Québec: relating animal occurrences and forest map information. Ecological Modelling, 174(4):387–399, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003004599.

Park:2006:PEB

[PLS^{+}06] Young-Seuk Park, Sovan Lek, Michele Scardi, Piet F. M. Verdonschot, and Sven E. Jørgensen. Patterning exergy of benthic macroinvertebrate communities using selforganizing maps. Ecological Modelling, 195(1–2):105–113, May 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000500582X.

Palosuo:2005:LDA

[PLTT05] T. Palosuo, J. Liski, J. A. Trofymow, and B. D. Titus. Litter decomposition affected by climate and litter quality — testing the Yasso model with litterbag data from the Canadian intersite decomposition experiment. Ecological Modelling, 189(1–2):183–198, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002024.

Plummer:2000:PCE

[Plu00] S. E. Plummer. Perspectives on combining ecological process models and remotely sensed data. Ecological Modelling, 129(2–3):169–186, May 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002337.

Patricio:2006:MBM

[PM06a] Joana Patrício and João Carlos Marques. Mass balanced models of the food web in three areas along a gradient of eutrophication symptoms in the south arm of the Mondego estuary (Portugal). Ecological Modelling, 197(1–2):21–34, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006001189.

Pike:2006:BCS

[PM06b] Nathan Pike and Andrea Manica. The basis of cowardice in social defenders. Ecological Modelling, 196(3–4):275–282, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006001128.

Pawlowski:2009:MMA

[PM09] Christopher W. Pawlowski and Christopher McCord. A Markov model for assessing ecological stability properties. Ecological Modelling, 220(2):86–95, January 24, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008004766.

Podur:2003:SPL

[PMC03] Justin Podur, David L. Martell, and Ferenc Csillag. Spatial patterns of lightning-caused forest fires in Ontario, 1976– 1998. Ecological Modelling, 164(1):1–20, June 1, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002003861.

Picard:2008:IEV

[PMC08] Nicolas Picard, Frédéric Mortier, and Pierrette Chagneau. Influence of estimators of the vital rates in the stock recovery rate when using matrix models for tropical rainforests. Ecological Modelling, 214(2–4):349–360, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001221.

Pelletier:2009:EBS

[PMD^{+}09] Dominique Pelletier, Stéphanie Mahevas, Hilaire Drouineau, Youen Vermard, Olivier Thebaud, Olivier Guyader, and Benjamin Poussin. Evaluation of the bioeconomic sustainability of multi-species multi-fleet fisheries under a wide range of policy options using ISIS-Fish. Ecological Modelling, 220(7):1013–1033, April 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000349.

Panzieri:2000:IBJ

[PMH00] M. Panzieri, N. Marchettini, and T. G. Hallam. Importance of the Bradhyrizobium japonicum symbiosis for the sustainability of a soybean cultivation. Ecological Modelling, 135(2–3):301–310, December 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003835.

Prost:2008:CSS

[PMJ08] Lorène Prost, David Makowski, and Marie-Hélène Jeuffroy. Comparison of stepwise selection and Bayesian model averaging for yield gap analysis. Ecological Modelling, 219(1–2):66–76, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003815.

Platts:2008:PTD

[PMLM08] Philip J. Platts, Colin J. McClean, Jon C. Lovett, and Rob Marchant. Predicting tree distributions in an East African biodiversity hotspot: model selection, data bias and envelope uncertainty. Ecological Modelling, 218 (1–2):121–134, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008003062.

Pedersen:2008:TML

[PNN^{+}08] Torstein Pedersen, Marianne Nilsen, Einar M. Nilssen, Erik Berg, and Marit Reigstad. Trophic model of a lightly exploited cod-dominated ecosystem. Ecological Modelling, 214(2–4):95–111, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006357.

Park:2003:OEM

[PNU03] Seok S. Park, Yumee Na, and Christopher G. Uchrin. An oxygen equivalent model for water quality dynamics in a macrophyte dominated river. Ecological Modelling, 168(1–2):1–12, October 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001893.

Penteriani:2008:FMW

[POF08] Vincenzo Penteriani, Fermín Otalora, and Miguel Ferrer. Floater mortality within settlement areas can explain the Allee effect in breeding populations. Ecological Modelling, 213(1):98–104, April 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006175.

Petersen:2002:FTS

[POH02] Bjørn M. Petersen, Jørgen E. Olesen, and Tove Heidmann. A flexible tool for simulation of soil carbon turnover. Ecological Modelling, 151(1):1–14, May 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000340.

Potter:2004:PCC

[Pot04] Christopher Potter. Predicting climate change effects on vegetation, soil thermal dynamics, and carbon cycling in ecosystems of interior Alaska. Ecological Modelling, 175(1):1–24, June 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003004319.

Pinnegar:2004:PIE

[PP04] John. K. Pinnegar and Nicholas V. C. Polunin. Predicting indirect effects of fishing in Mediterranean rocky littoral communities using a dynamic simulation model. Ecological Modelling, 172(2–4):249–267, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003727.

Pedersen:2007:MSA

[PPE^{+}07] Anders Pedersen, Bjørn M. Petersen, Jørgen Eriksen, Søren Hansen, and Lars S. Jensen. A model simulation analysis of soil nitrate concentrations — does soil organic matter pool structure or catch crop growth parameters matter most? Ecological Modelling, 205(1–2):209–220, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007000907.

Pineiro:2008:HEM

[PPGP08] Gervasio Piñeiro, Susana Perelman, Juan P. Guerschman, and José M. Paruelo. How to evaluate models: Observed vs. predicted or predicted vs. observed? Ecological Modelling, 216(3–4):316–322, September 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002305.

Potting:2005:IBE

[PPP05] R. P. J. Potting, J. N. Perry, and W. Powell. Insect behavioural ecology and other factors affecting the control efficacy of agro-ecosystem diversification strategies. Ecological Modelling, 182(2):199–216, March 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004132.

Paul:2003:SAP

[PPR03] Keryn I. Paul, Phil J. Polglase, and Gary P. Richards. Sensitivity analysis of predicted change in soil carbon following afforestation. Ecological Modelling, 164(2–3):137–152, June 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000279.

Peterson:2008:RRO

[PPS08] A. Townsend Peterson, Monica Papeş, and Jorge Soberón. Rethinking receiver operating characteristic analysis applications in ecological niche modeling. Ecological Modelling, 213(1):63–72, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007006163.

Portela:2001:DMP

[PR01] Rosimeiry Portela and Ida Rademacher. A dynamic model of patterns of deforestation and their effect on the ability of the Brazilian Amazonia to provide ecosystem services. Ecological Modelling, 143(1–2):115–146, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003593.

Pollett:2007:MCD

[PR07] Philip K. Pollett and Joshua V. Ross. Management, control and decision making for ecological systems. Ecological Modelling, 201(1):1, February 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003498.

Prato:2000:MAB

[Pra00] Tony Prato. Multiple attribute Bayesian analysis of adaptive ecosystem management. Ecological Modelling, 133 (3):181–193, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002921.

Prato:2005:BAM

[Pra05a] Tony Prato. Bayesian adaptive management of ecosystems. Ecological Modelling, 183(2–3):147–156, April 25, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004004363.

Prato:2005:FLA

[Pra05b] Tony Prato. A fuzzy logic approach for evaluating ecosystem sustainability. Ecological Modelling, 187(2– 3):361–368, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000608.

Prato:2007:SEP

[Pra07] Tony Prato. Selection and evaluation of projects to conserve ecosystem services. Ecological Modelling, 203(3–4):290–296, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006006065.

Prato:2008:SMA

[Pra08] Tony Prato. Stochastic multiple attribute evaluation of land use policies. Ecological Modelling, 219(1–2): 115–124, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008004043.

Prato:2009:EMW

[Pra09] Tony Prato. Evaluating and managing wildlife impacts of climate change under uncertainty. Ecological Modelling, 220(7):923–930, April 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000362.

Park:2007:SAS

[PRL07] Young-Seuk Park, Jorge Rabinovich, and Sovan Lek. Sensitivity analysis and stability patterns of two-species pest models using artificial neural networks. Ecological Modelling, 204(3–4):427–438, June 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000506.

Petzoldt:2009:EZD

[PRRB09] Thomas Petzoldt, Lars Rudolf, Karsten Rinke, and Jürgen Benndorf. Effects of zooplankton diel vertical migration on a phytoplankton community: a scenario analysis of the underlying mechanisms. Ecological Modelling, 220(9–10):1358–1368, May 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000866.

Priya:2001:NSC

[PS01] Satya Priya and Ryosuke Shibasaki. National spatial crop yield simulation using GIS-based crop production model. Ecological Modelling, 136(2–3):113–129, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003641.

Perttunen:2005:ILS

[PS05] Jari Perttunen and Risto Sievänen. Incorporating Lindenmayer systems for architectural development in a functional-structural tree model. Ecological Modelling, 181(4):479–491, February 10, 2005. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003734.

Pedersen:2009:IBP

[PS09] Rune Østergaard Pedersen and Jens Peter Skovsgaard. Impact of bias in predicted height on tree volume estimation: a case-study of intrinsic nonlinearity. Ecological Modelling, 220(20):2656–2664, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004529.

Popp:2007:SIV

[PSBJ07] Alexander Popp, Monika Schwager, Niels Blaum, and Florian Jeltsch. Simulating the impacts of vegetation structure on the occurrence of a small mammalian carnivore in semi-arid savanna rangelands. Ecological Modelling, 209(2–4):136–148, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003250.

Pastres:2001:MRT

[PSC^{+}01] R. Pastres, C. Solidoro, G. Cossarini, D. Melaku Canu, and C. Dejak. Managing the rearing of Tapes philippinarum in the lagoon of Venice: a decision support system. Ecological Modelling, 138(1–3):231–245, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000404X.

Payne:2004:MOM

[PSC04] Adam G. Payne, Carl Smith, and Andrew C. Campbell. A model of oxygen-mediated filial cannibalism in fishes. Ecological Modelling, 174(3):253–266, May 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004289.

Peterson:2006:TPE

[PSCMMNS06] A. Townsend Peterson, Victor Sánchez-Cordero, Enrique Martínez-Meyer, and Adolfo G. Navarro-Sigüenza. Tracking population extirpations via melding ecological niche modeling with land-cover information. Ecological Modelling, 195(3–4):229–236, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500606X.

Peterson:2001:EGC

[PSCS^{+}01] A. Townsend Peterson, Victor Sánchez-Cordero, Jorge Soberón, Jeremy Bartley, Robert W. Buddemeier, and Adolfo G. Navarro-Sigüenza. Effects of global climate change on geographic distributions of Mexican cracidae. Ecological Modelling, 144(1):21–30, October 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003453.

Park:2007:CPB

[PSP^{+}07] Young-Seuk Park, Mi-Young Song, Young-Cheol Park, Kyung-Hee Oh, Eungchun Cho, and Tae-Soo Chon. Community patterns of benthic macroinvertebrates collected on the national scale in Korea. Ecological Modelling, 203(1–2):26–33, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006005552.

Pettijohn:2009:MMS

[PSPD09] Justin C. Pettijohn, Guido D. Salvucci, Nathan G. Phillips, and Michael J. Daley. Mechanisms of moisture stress in a mid-latitude temperate forest: Implications for feedforward and feedback controls from an irrigation experiment. Ecological Modelling, 220(7):968–978, April 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005930.

Podchong:2009:IAI

[PSVH09] Suchart Podchong, Dietrich Schmidt-Vogt, and Kiyoshi Honda. An improved approach for identifying suitable habitat of sambar deer (Cervus unicolor Kerr) using ecological niche analysis and environmental categorization: Case study at Phu-Khieo Wildlife Sanctuary, Thailand. Ecological Modelling, 220(17):2103–2114, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003160.

Parysow:2002:AEE

[PT02] Pablo Parysow and David J. Tazik. Assessing the effect of estimation error on population viability analysis: an example using the black-capped vireo. Ecological Modelling, 155(2–3):217–229, October 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001461.

Penenko:2008:ODM

[PT08] Vladimir Penenko and Elena Tsvetova. Orthogonal decomposition methods for inclusion of climatic data into environmental studies. Ecological Modelling, 217(3–4):279–291, October 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008002597.

Portilla:2009:DSA

[PTGI09] E. Portilla, P. Tett, P. A. Gillibrand, and Mark Inall. Description and sensitivity analysis for the LESV model: Water quality variables and the balance of organisms in a fjordic region of restricted exchange. Ecological Modelling, 220(18):2187–2205, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003263.

Phillips:2004:SMT

[PTS^{+}04] P. D. Phillips, I. S. Thompson, J. N. M. Silva, P. R. van Gardingen, and B. Degen. Scaling up models of tree competition for tropical forest population genetics simulation. Ecological Modelling, 180(2–3):419–434, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004002923.

Preuss:2004:SGD

[PV04] Ilana Preuss and Amanda W. Vemuri. “Smart growth” and dynamic modeling: implications for quality of life in Montgomery County, Maryland. Ecological Modelling, 171(4):415–432, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003582.

Pielaat:2002:SVS

[PvdBFJ02] A. Pielaat, F. van den Bosch, B. D. L. Fitt, and M. J. Jeger. Simulation of vertical spread of plant diseases in a crop canopy by stem extension and splash dispersal. Ecological Modelling, 151(2–3):195–212, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004847.

Papastamati:2002:MDP

[PvdBW^{+}02] K. Papastamati, F. van den Bosch, S. J. Welham, B. D. L. Fitt, N. Evans, and J. M. Steed. Modelling the daily progress of light leaf spot epidemics on winter oilseed rape (Brassica napus), in relation to Pyrenopeziza brassicae inoculum concentrations and weather factors. Ecological Modelling, 148(2):169–189, February 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004197.

Peters:2009:UPV

[PVS^{+}09] Jan Peters, Niko E. C. Verhoest, Roeland Samson, Marc Van Meirvenne, Liesbet Cockx, and Bernard De Baets. Uncertainty propagation in vegetation distribution models based on ensemble classifiers. Ecological Modelling, 220(6):791–804, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008005942.

Petropoulos:2009:GBS

[PWC^{+}09] G. Petropoulos, M. J. Wooster, T. N. Carlson, M. C. Kennedy, and M. Scholze. A global Bayesian sensitivity analysis of the 1d SimSphere soil–vegetation–atmospheric transfer (SVAT) model using Gaussian model emulation. Ecological Modelling, 220(19):2427–2440, October 10, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009004025.

Pollino:2007:ECI

[PWH07] Carmel A. Pollino, Andrea K. White, and Barry T. Hart. Examination of conflicts and improved strategies for the management of an endangered eucalypt species using Bayesian networks. Ecological Modelling, 201(1):37–59, February 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006003541.

Purucker:2007:UHC

[PWSS07] S. T. Purucker, C. J. E. Welsh, R. N. Stewart, and P. Starzec. Use of habitat-contamination spatial correlation to determine when to perform a spatially explicit ecological risk assessment. Ecological Modelling, 204(1–2):180–192, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007000038.

Peng:2007:MTRa

[PWSY07] Jian-Feng Peng, Bao-Zhen Wang, Yong-Hui Song, and Peng Yuan. Modeling N transformation and removal in a duckweed pond: Model development and calibration. Ecological Modelling, 206(1–2):147–152, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001743.

Platt:2009:PPB

[PWZ^{+}09] Trevor Platt, George N. White, Li Zhai, Shubha Sathyendranath, and Shovonlal Roy. The phenology of phytoplankton blooms: Ecosystem indicators from remote sensing. Ecological Modelling, 220(21):3057–3069, November 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008005462.

Pyke:2004:SVP

[Pyk04] Christopher R. Pyke. Simulating vernal pool hydrologic regimes for two locations in California, USA. Ecological Modelling, 173(2–3):109–127, April 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003612.

Quince:2005:TSI

[QHM05] Christopher Quince, Paul G. Higgs, and Alan J. McKane. Topological structure and interaction strengths in model food webs. Ecological Modelling, 187(4):389–412, October 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001031.

Qian:2003:TSM

[QKR03] Song S. Qian, Ryan S. King, and Curtis J. Richardson. Two statistical methods for the detection of environmental thresholds. Ecological Modelling, 166(1–2):87–97, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003000978.

Quartier:2000:SSR

[QMH00] Robin Quartier and Ulrich Müller-Herold. On secondary spatial ranges of transformation products in the environment. Ecological Modelling, 135(2–3):187–198, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003653.

Qian:2003:MCM

[QSB03] Song S. Qian, Craig A. Stow, and Mark E. Borsuk. On Monte Carlo methods for Bayesian inference. Ecological Modelling, 159(2–3):269–277, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002995.

Qi:2002:TSS

[QXW02] Ye Qi, Ming Xu, and Jianguo Wu. Temperature sensitivity of soil respiration and its effects on ecosystem carbon budget: nonlinearity begets surprises. Ecological Modelling, 153(1–2):131–142, July 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005063.

Ricotta:2006:SCE

[RA06] Carlo Ricotta and Madhur Anand. Spatial complexity of ecological communities: Bridging the gap between probabilistic and non-probabilistic uncertainty measures. Ecological Modelling, 197(1–2):59–66, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001116.

Ricotta:2001:SOC

[RADD^{+}01] Carlo Ricotta, Margarita Arianoutsou, Ricardo Díaz-Delgado, Beatriz Duguy, Francisco Lloret, Eleni Maroudi, Stefano Mazzoleni, José Manuel Moreno, Serge Rambal, Ramon Vallejo, and Antonio Vázquez. Self-organized criticality of wildfires ecologically revisited. Ecological Modelling, 141(1–3):307–311, July 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002721.

Rautiainen:2007:GDS

[RAH07] Pirjo Rautiainen, Sami Aikio, and Marko Hyvärinen. Growth and disturbance in the spatial dynamics of an endangered seashore grass Arctophila fulva var. pendulina. Ecological Modelling, 207(2–4):145–154, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007002438.

Railsback:2001:CCA

[Rai01] Steven F. Railsback. Concepts from complex adaptive systems as a framework for individual-based modelling. Ecological Modelling, 139(1):47–62, March 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002289.

Romero:2004:OTD

[RAI04] J. R. Romero, J. P. Antenucci, and J. Imberger. One- and three-dimensional biogeochemical simulations of two differing reservoirs. Ecological Modelling, 174(1–2):143–160, May 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004000043.

Rai:2008:MWS

[Rai08] Vikas Rai. Modeling a wetland system: the case of Keoladeo National Park (KNP), India. Ecological Modelling, 210(3):247–252, January 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003997.

Rukhovets:2003:DLL

[RAM^{+}03] L. A. Rukhovets, G. P. Astrakhantsev, V. V. Menshutkin, T. R. Minina, N. A. Petrova, and V. N. Poloskov. Development of Lake Ladoga ecosystem models: modeling of the phytoplankton succession in the eutrophication process. I. Ecological Modelling, 165(1):49–77, July 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000619.

Ray:2008:CSV

[Ray08] Santanu Ray. Comparative study of virgin and reclaimed islands of Sundarban mangrove ecosystem through network analysis. Ecological Modelling, 215(1–3):207–216, July 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000872.

Repsilber:2002:SGP

[RB02] Dirk Repsilber and F. Bialozyt. Spatial genetic patterns systematically accelerate and bias drift-based genetic erosion. Ecological Modelling, 148(3):251–261, March 1, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001004410.

Ray:2006:SUH

[RB06] Nicolas Ray and Mark A. Burgman. Subjective uncertainties in habitat suitability maps. Ecological Modelling, 195(3–4):172–186, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006277.

Ruzzenenti:2009:CCS

[RB09] F. Ruzzenenti and R. Basosi. Complexity change and space symmetry rupture. Ecological Modelling, 220(16):1880–1885, August 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002737.

Rinaldo:2005:NST

[RBB05] Andrea Rinaldo, Enrico Bertuzzo, and Gianluca Botter. Nonpoint source transport models from empiricism to coherent theoretical frameworks. Ecological Modelling, 184(1):19– 35, May 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004005824.

Robinson:2009:MCS

[RBC09] D. T. Robinson, D. G. Brown, and W. S. Currie. Modelling carbon storage in highly fragmented and human-dominated landscapes: Linking land-cover patterns and ecosystem models. Ecological Modelling, 220(9–10):1325–1338, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001501.

Ross:2008:MDW

[RBE^{+}08] Karen A. Ross, Michael Bedward, Murray V. Ellis, Andrew Deane, Christopher C. Simpson, and Ross A. Bradstock. Modelling the dynamics of white cypress pine Callitris glaucophylla woodlands in inland south-eastern Australia. Ecological Modelling, 211(1–2):11–24, February 24, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007004218.

Rutherford:2008:ALU

[RBEZ08] Gillian N. Rutherford, Peter Bebi, Peter J. Edwards, and Niklaus E. Zimmermann. Assessing land-use statistics to model land cover change in a mountainous landscape in the European Alps. Ecological Modelling, 212(3–4):460–471, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005674.

Richmond:2005:REG

[RBR05] Courtney E. Richmond, Denise L. Breitburg, and Kenneth A. Rose. The role of environmental generalist species in ecosystem function. Ecological Modelling, 188 (2–4):279–295, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005001882.

Rogers-Bennett:2006:SEG

[RBR06] Laura Rogers-Bennett and Donald W. Rogers. A semiempirical growth estimation method for matrix models of endangered species. Ecological Modelling, 195(3–4):237–246, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005006046.

Raick:2006:CPC

[RBSG06] C. Raick, J.-M. Beckers, K. Soetaert, and M. Grégoire. Can principal component analysis be used to predict the dynamics of a strongly non-linear marine biogeochemical model? Ecological Modelling, 196(3–4):345–364, July 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006000780.

Ray:2001:OEI

[RBSJ01] Santanu Ray, Ludĕk Berec, Milan Straškraba, and Sven Erik Jørgensen. Optimization of exergy and implications of body sizes of phytoplankton and zooplankton in an aquatic ecosystem model. Ecological Modelling, 140(3):219–234, June 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003222.

Rashleigh:2009:FMP

[RBW09] Brenda Rashleigh, M. Craig Barber, and David M. Walters. Foodweb modeling for polychlorinated biphenyls (PCBs) in the Twelvemile Creek Arm of Lake Hartwell, South Carolina, USA. Ecological Modelling, 220(2):254–264, January 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008004353.

Ripley:2006:RVS

[RC06] Bonnie J. Ripley and Hal Caswell. Recruitment variability and stochastic population growth of the soft-shell clam, Mya arenaria. Ecological Modelling, 193(3–4):517–530, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004114.

Reid:2008:TMF

[RC08] Tim Reid and Neil Crout. A thermodynamic model of freshwater Antarctic lake ice. Ecological Modelling, 210(3):231–241, January 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004036.

Rossi:2008:MMS

[RCGB08] Vittorio Rossi, Tito Caffi, Simona Giosuè, and Riccardo Bugiani. A mechanistic model simulating primary infections of downy mildew in grapevine. Ecological Modelling, 212(3–4):480–491, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005881.

Romero-Calcerrada:2006:HQA

[RCL06] Raul Romero-Calcerrada and Sandra Luque. Habitat quality assessment using weights-of-evidence based GIS modelling: the case of Picoides tridactylus as species indicator of the biodiversity value of the Finnish forest. Ecological Modelling, 196(1–2):62–76, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000600055X.

Ryu:2006:SEP

[RCZ^{+}06] Soung-Ryoul Ryu, Jiquan Chen, Daolan Zheng, Mary K. Bresee, and Thomas R. Crow. Simulating the effects of prescribed burning on fuel loading and timber production (EcoFL) in managed northern Wisconsin forests. Ecological Modelling, 196(3–4):395–406, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000779.

Reineking:2006:CPR

[Rd06] Björn Reineking and Boris Schrö der. Constrain to perform: Regularization of habitat models. Ecological Modelling, 193(3–4):675–690, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004849.

Rashleigh:2007:CCF

[RD07] Brenda Rashleigh and Donald L. DeAngelis. Conditions for coexistence of freshwater mussel species via partitioning of fish host resources. Ecological Modelling, 201 (2):171–178, February 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004194.

Rivera:2007:SME

[RdQFO07] Elmer Ccopa Rivera, Julio Ferraz de Queiroz, José Maria Ferraz, and Enrique Ortega. Systems models to evaluate eutrophication in the Broa Reservoir, São Carlos, Brazil. Ecological Modelling, 202(3–4):518–526, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005680.

Raghu:2007:PRB

[RDS07] S. Raghu, K. Dhileepan, and J. C. Scanlan. Predicting risk and benefit a priori in biological control of invasive plant species: a systems modelling approach. Ecological Modelling, 208(2–4):247–262, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003067.

Robinson:2003:DVH

[RE03] Andrew P. Robinson and Alan R. Ek. Description and validation of a hybrid model of forest growth and stand dynamics for the Great Lakes region. Ecological Modelling, 170(1):73–104, December 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003041.

Recknagel:2001:AML

[Rec01a] Friedrich Recknagel. Applications of machine learning to ecological modelling. Ecological Modelling, 146 (1–3):303–310, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003167.

Recknagel:2001:P

[Rec01b] Friedrich Recknagel. Preface. Ecological Modelling, 146(1–3):1–2, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002915.

Recknagel:2003:SAF

[Rec03] Friedrich Recknagel. Simulation of aquatic food web and species interactions by adaptive agents embodied with evolutionary computation: a conceptual framework. Ecological Modelling, 170(2–3):291–302, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002345.

Remm:2004:CBP

[Rem04] Kalle Remm. Case-based predictions for species and habitat mapping. Ecological Modelling, 177(3–4):259–281, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004001383.

Reuter:2005:CPE

[Reu05] Hauke Reuter. Community processes as emergent properties: Modelling multilevel interaction in small mammals communities. Ecological Modelling, 186(4):427–446, September 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005001353.

Reynolds:2002:EPE

[Rey02] C. S. Reynolds. Ecological pattern and ecosystem theory. Ecological Modelling, 158(3):181–200, December 31, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002302.

Reynolds:2003:PCA

[Rey03] C. S. Reynolds. Planktic community assembly in flowing water and the ecosystem health of rivers. Ecological Modelling, 160(3):191–203, February 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002521.

Robinson:2004:MVU

[RF04] Andrew P. Robinson and Robert E. Froese. Model validation using equivalence tests. Ecological Modelling, 176(3–4):349–358, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000985.

Rammig:2009:SFS

[RF09] Anja Rammig and Lorenz Fahse. Simulating forest succession after blowdown events: the crucial role of space for a realistic management. Ecological Modelling, 220 (24):3555–3564, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000900444X.

Rasse:2001:MST

[RFA^{+}01] Daniel P. Rasse, Louis François, Marc Aubinet, Andrew S. Kowalski, Inge Vande Walle, Eric Laitat, and Jean-Claude Gérard. Modelling short-term CO_{2} fluxes and long-term tree growth in temperate forests with ASPECTS. Ecological Modelling, 141(1–3):35–52, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002393.

Rueda:2007:PRN

[RFdV07] Francisco J. Rueda, William E. Fleenor, and Inmaculada de Vicente. Pathways of river nutrients towards the euphotic zone in a deep-reservoir of small size: Uncertainty analysis. Ecological Modelling, 202(3–4):345–361, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006005382.

Rinaldi:2004:DFS

[RG04] Sergio Rinaldi and Alessandra Gragnani. Destabilizing factors in slow–fast systems. Ecological Modelling, 180 (4):445–460, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003077.

Rashleigh:2005:IBS

[RG05] Brenda Rashleigh and Gary D. Grossman. An individualbased simulation model for mottled sculpin (Cottus bairdi) in a southern Appalachian stream. Ecological Modelling, 187(2–3):247–258, September 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000645.

Robinson:2006:MDM

[RG06] Donald C. E. Robinson and Brian W. Geils. Modelling dwarf mistletoe at three scales: life history, ballistics and contagion. Ecological Modelling, 199(1):23–38, November 1, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000600278X.

Riedo:2000:NPP

[RGF00] Marcel Riedo, Dimitrios Gyalistras, and Jürg Fuhrer. Net primary production and carbon stocks in differently managed grasslands: simulation of site-specific sensitivity to an increase in atmospheric CO_{2} and to climate change. Ecological Modelling, 134(2–3):207–227, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003562.

Rickebusch:2007:CPL

[RGL^{+}07] Sophie Rickebusch, Mario Gellrich, Heike Lischke, Antoine Guisan, and Niklaus E. Zimmermann. Combining probabilistic land-use change and tree population dynamics modelling to simulate responses in mountain forests. Ecological Modelling, 209(2–4):157–168, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003262.

Richards:2006:GBA

[RGO^{+}06] M. Richards, M. Ghanem, M. Osmond, Y. Guo, and J. Hassard. Grid-based analysis of air pollution data. Ecological Modelling, 194(1–3):274–286, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005259.

Robson:2004:TDM

[RH04] B. J. Robson and D. P. Hamilton. Three-dimensional modelling of a Microcystis bloom event in the Swan River estuary, Western Australia. Ecological Modelling, 174(1–2):203– 222, May 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004000080.

Raymond:2009:NBE

[RH09] Ben Raymond and Graham Hosie. Network-based exploration and visualisation of ecological data. Ecological Modelling, 220(5):673–683, March 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005863.

Roy:2004:EFP

[RHB04] Denis Roy, G. D. Haffner, and Stephen B. Brandt. Estimating fish production potentials using a temporally explicit model. Ecological Modelling, 173(2–3):241–257, April 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000300396X.

Risch:2005:SSF

[RHB05] Anita C. Risch, Caroline Heiri, and Harald Bugmann. Simulating structural forest patterns with a forest gap model: a model evaluation. Ecological Modelling, 181 (2–3):161–172, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003643.

Rachowicz:2006:ESS

[RHB06] Lara J. Rachowicz, Alan E. Hubbard, and Steven R. Beissinger. Evaluating at-sea sampling designs for marbled murrelets using a spatially explicit model. Ecological Modelling, 196(3–4):329–344, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000731.

Ramzi:2006:LDR

[RHE06] A. Ramzi, My. L. Hbid, and O. Ettahiri. Larval dynamics and recruitment modelling of the Moroccan Atlantic coast sardine (Sardina pilchardus). Ecological Modelling, 197(3–4):296–302, August 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001499.

Richardson:2005:DGS

[RHH05] F. D. Richardson, B. D. Hahn, and M. T. Hoffman. On the dynamics of grazing systems in the semi-arid succulent karoo: the relevance of equilibrium and non-equilibrium concepts to the sustainability of semi-arid pastoral systems. Ecological Modelling, 187(4):491–512, October 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000500102X.

Reynolds-Hogland:2008:EID

[RHHM08] Melissa J. Reynolds-Hogland, John S. Hogland, and Michael S. Mitchell. Evaluating intercepts from demographic models to understand resource limitation and resource thresholds. Ecological Modelling, 211(3–4):424– 432, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380007004978.

Reuter:2005:CEC

[RHM^{+}05] Hauke Reuter, Franz Hölker, Ulrike Middelhoff, Fred Jopp, Christiane Eschenbach, and Broder Breckling. The concepts of emergent and collective properties in individualbased models — summary and outlook of the Bornhöved case studies. Ecological Modelling, 186(4):489–501, September 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005001389.

Rout:2007:MLT

[RHP07] T. M. Rout, C. E. Hauser, and H. P. Possingham. Minimise long-term loss or maximise short-term gain?: Optimal translocation strategies for threatened species. Ecological Modelling, 201(1):67–74, February 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003565.

Ricotta:2000:TES

[Ric00] Carlo Ricotta. From theoretical ecology to statistical physics and back: self-similar landscape metrics as a synthesis of ecological diversity and geometrical complexity. Ecological Modelling, 125(2–3):245–253, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001854.

Ricotta:2002:BGB

[Ric02] Carlo Ricotta. Bridging the gap between ecological diversity indices and measures of biodiversity with Shannon’s entropy: comment to Izsák and papp. Ecological Modelling, 152(1):1–3, June 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004689.

Ricotta:2005:APR

[Ric05] Carlo Ricotta. Additive partitioning of Rao’s quadratic diversity: a hierarchical approach. Ecological Modelling, 183(4):365–371, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000400496X.

Reynolds:2001:EBS

[RIE01] C. S. Reynolds, A. E. Irish, and J. A. Elliott. The ecological basis for simulating phytoplankton responses to environmental change (PROTECH). Ecological Modelling, 140(3):271–291, June 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003301.

Rios-Insua:2006:MAD

[RIGJM06] Sixto Ríos-Insua, Eduardo Gallego, Antonio Jiménez, and Alfonso Mateos. A multi-attribute decision support system for selecting intervention strategies for radionuclide contaminated freshwater ecosystems. Ecological Modelling, 196(1–2):195–208, July 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006150.

Ramos-Jiliberto:2008:DEI

[RJDF08] Rodrigo Ramos-Jiliberto, Hector Duarte, and Ernesto Frodden. Dynamic effects of inducible defenses in a one-prey twopredators system. Ecological Modelling, 214(2–4):242–250, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000800077X.

Ramos-Jiliberto:2007:PEV

[RJFAA07] Rodrigo Ramos-Jiliberto, Ernesto Frodden, and Adriana Aránguiz-Acuña. Pre-encounter versus post-encounter inducible defenses in predator–prey model systems. Ecological Modelling, 200(1–2):99–108, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003462.

Ramos-Jiliberto:2000:RBP

[RJGO00] Rodrigo Ramos-Jiliberto and Eduardo González-Olivares. Relating behavior to population dynamics: a predator–prey metaphysiological model emphasizing zooplankton diel vertical migration as an inducible response. Ecological Modelling, 127(2–3):221–233, March 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002148.

Reich:2004:PLN

[RJR04] Robin M. Reich, Suzanne M. Joy, and Richard T. Reynolds. Predicting the location of northern goshawk nests: modeling the spatial dependency between nest locations and forest structure. Ecological Modelling, 176(1–2):109–133, August 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000300560X.

Roy:2006:BMI

[RJS^{+}06] P. S. Roy, P. K. Joshi, S. Singh, S. Agarwal, D. Yadav, and C. Jegannathan. Biome mapping in India using vegetation type map derived using temporal satellite data and environmental parameters. Ecological Modelling, 197(1–2):148–158, August 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001062.

Renton:2005:FSP

[RKH05] Michael Renton, Pekka Kaitaniemi, and Jim Hanan. Functional–structural plant modelling using a combination of architectural analysis, L-systems and a canonical model of function. Ecological Modelling, 184(2–4):277–298, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004005538.

Roder:2007:AGG

[RKH^{+}07] A. Röder, T. Kuemmerle, J. Hill, V. P. Papanastasis, and G. M. Tsiourlis. Adaptation of a grazing gradient concept to heterogeneous Mediterranean rangelands using cost surface modelling. Ecological Modelling, 204(3–4):387–398, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007000312.

Roth:2007:LTA

[RKS^{+}07] Brian M. Roth, Isaac C. Kaplan, Greg G. Sass, Pieter T. Johnson, Anna E. Marburg, Anthony C. Yannarell, Tanya D. Havlicek, Theodore V. Willis, Monica G. Turner, and Stephen R. Carpenter. Linking terrestrial and aquatic ecosystems: the role of woody habitat in lake food webs. Ecological Modelling, 203(3–4):439–452, May 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006006302.

Raynaud:2005:SBC

[RL05a] Xavier Raynaud and Paul W. Leadley. Symmetry of belowground competition in a spatially explicit model of nutrient competition. Ecological Modelling, 189(3– 4):447–453, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002048.

Reed:2005:MBR

[RL05b] J. Michael Reed and Stephen H. Levine. A model for behavioral regulation of metapopulation dynamics. Ecological Modelling, 183(4):411–423, May 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005009.

Rossi:2009:CSO

[RL09] Federico Rossi and Maria Liria Turco Liveri. Chemical selforganization in self-assembling biomimetic systems. Ecological Modelling, 220(16):1857–1864, August 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002713.

Reyjol:2001:MMU

[RLBL01] Yorick Reyjol, Puy Lim, Alain Belaud, and Sovan Lek. Modelling of microhabitat used by fish in natural and regulated flows in the river Garonne (France). Ecological Modelling, 146(1–3):131–142, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003015.

Russell:2009:IPM

[RLDL09] James C. Russell, Vincent Lecomte, Yves Dumont, and Matthieu Le Corre. Intraguild predation and mesopredator release effect on long-lived prey. Ecological Modelling, 220(8):1098–1104, April 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000714.

Richter:2004:SMC

[RLFB04] Hartmut Richter, Christian Lückstädt, Ulfert Focken, and Klaus Becker. Some mathematical considerations in estimating daily ration in fish using food consumption models. Ecological Modelling, 171(4):381–393, February 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003003569.

Railsback:2001:RDM

[RLHD01] S. F. Railsback, R. H. Lamberson, B. C. Harvey, and W. E. Duffy. Reply to ‘Discussion on ‘Movement rules for Individual-based models of stream fish’: S. F. Railsback et al. Ecol. Model. 123 (1999) 73–89. Ecological Modelling, 143(3):249–251, September 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003337.

Rochette:2009:DUQ

[RLLB09] S. Rochette, J. Lobry, M. Lepage, and Ph. Boët. Dealing with uncertainty in qualitative models with a semiquantitative approach based on simulations. Application to the Gironde estuarine food web (France). Ecological Modelling, 220(2):122–132, January 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004572.

Robin:2005:USI

[RLR05] Jessica Robin, Elissa Levine, and Susan Riha. Utilizing satellite imagery and GLOBE student data to model soil dynamics. Ecological Modelling, 185(1):133–145, June 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004006003.

Remm:2009:RDF

[RLR09] Kalle Remm, Madli Linder, and Liina Remm. Relative density of finds for assessing similarity-based maps of orchid occurrence. Ecological Modelling, 220(3):294–309, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008004985.

Rosas-Luis:2008:IJS

[RLSZK^{+}08] R. Rosas-Luis, C. A. Salinas-Zavala, V. Koch, P. Del Monte Luna, and M. V. Morales-Zárate. Importance of jumbo squid Dosidicus gigas (Orbigny, 1835) in the pelagic ecosystem of the central Gulf of California. Ecological Modelling, 218(1–2):149–161, October 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800313X.

Reed:2002:PLB

[RM02] William J. Reed and Kevin S. McKelvey. Power-law behaviour and parametric models for the size-distribution of forest fires. Ecological Modelling, 150(3):239–254, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004835.

Renken:2009:MDC

[RM09] Henk Renken and Peter J. Mumby. Modelling the dynamics of coral reef macroalgae using a Bayesian belief network approach. Ecological Modelling, 220(9–10):1305–1314, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001525.

Richards:2008:OCI

[RMA08] M. Richards, A. J. S. McDonald, and M. J. Aitkenhead. Optimisation of competition indices using simulated annealing and artificial neural networks. Ecological Modelling, 214(2–4):375–384, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800149X.

Ricotta:2006:CSS

[RMBM06] Carlo Ricotta, Laura Micozzi, Maria Bellelli, and Stefano Mazzoleni. Characterizing self-similar temporal clustering of wildfires in the Cilento National Park (Southern Italy). Ecological Modelling, 197(3–4):512–515, August 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006001530.

Richards:2004:EES

[RMD04] Paul M. Richards, Wolf M. Mooij, and Donald L. DeAngelis. Evaluating the effect of salinity on a simulated American crocodile (Crocodylus acutus) population with applications to conservation and Everglades restoration. Ecological Modelling, 180(2–3):371–394, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002881.

Ryan:2004:UAN

[RMDC04] Matthew Ryan, Christoph Müller, Hong J. Di, and Keith C. Cameron. The use of artificial neural networks (ANNs) to simulate N_{2}O emissions from a temperate grassland ecosystem. Ecological Modelling, 175(2):189–194, July 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003004721.

Rayfield:2009:ICR

[RMF09] Bronwyn Rayfield, Atte Moilanen, and Marie-Josée Fortin. Incorporating consumer–resource spatial interactions in reserve design. Ecological Modelling, 220(5):725–733, March 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008005759.

Ruane:2009:MIS

[RMGR09] Neil M. Ruane, Alexander G. Murray, Fiona Geoghegan, and Rob S. Raynard. Modelling the initiation and spread of infectious pancreatic necrosis virus (IPNV) in the Irish salmon farming industry: the role of inputs. Ecological Modelling, 220(9–10):1369–1374, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001471.

Rueda:2006:RTR

[RMOA06] Francisco Rueda, Enrique Moreno-Ostos, and Joan Armengol. The residence time of river water in reservoirs. Ecological Modelling, 191(2):260–274, January 31, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002760.

Roy:2008:DOM

[RMR08] Madhumita Roy, Sudipto Mandal, and Santanu Ray. Detrital ontogenic model including decomposer diversity. Ecological Modelling, 215(1–3):200–206, July 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000896.

Roth:2008:SDS

[RMS08] James D. Roth, Dennis L. Murray, and Todd D. Steury. Spatial dynamics of sympatric canids: Modeling the impact of coyotes on red wolf recovery. Ecological Modelling, 214(2–4):391–403, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800152X.

Riedo:2002:CSP

[RMSS02] Marcel Riedo, Celia Milford, Martin Schmid, and Mark A. Sutton. Coupling soil–plant–atmosphere exchange of ammonia with ecosystem functioning in grasslands. Ecological Modelling, 158(1–2):83–110, December 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001692.

Rose:2007:CNN

[RMWW07] Kenneth A. Rose, Bernard A. Megrey, Francisco E. Werner, and Dan M. Ware. Calibration of the NEMURO nutrient– phytoplankton–zooplankton food web model to a coastal ecosystem: Evaluation of an automated calibration approach. Ecological Modelling, 202(1–2):38–51, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004558.

Ribarova:2008:MNP

[RNC08] Irina Ribarova, Plamen Ninov, and David Cooper. Modeling nutrient pollution during a first flood event using HSPF software: Iskar River case study, Bulgaria. Ecological Modelling, 211(1–2):241–246, February 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005017.

Rathgeber:2003:UBM

[RNKG03] Cyrille Rathgeber, Antoine Nicault, Jed O. Kaplan, and Joël Guiot. Using a biogeochemistry model in simulating forests productivity responses to climatic change and [CO_{2}] increase: example of Pinus halepensis in Provence (southeast France). Ecological Modelling, 166(3):239–255, August 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001613.

Robert:2003:SET

[Rob03] Audrey Robert. Simulation of the effect of topography and tree falls on stand dynamics and stand structure of tropical forests. Ecological Modelling, 167(3):287–303, September 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000300200X.

Robson:2005:RED

[Rob05] B. J. Robson. Representing the effects of diurnal variations in light on primary production on a seasonal time-scale. Ecological Modelling, 186(3):358–365, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500058X.

Rodrigues:2007:HAD

[Rod07] João Rodrigues. How ants determine the number of potential recruits. Ecological Modelling, 200(3–4):384–392, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003826.

Roelke:2000:CFQ

[Roe00] D. L. Roelke. Copepod food-quality threshold as a mechanism influencing phytoplankton succession and accumulation of biomass, and secondary productivity: a modeling study with management implications. Ecological Modelling, 134(2–3):245–274, October 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000346X.

Roelandt:2001:CSP

[Roe01] Caroline Roelandt. Coupled simulation of potential natural vegetation, terrestrial carbon balance and physical landsurface properties with the ALBIOC model. Ecological Modelling, 143(3):191–214, September 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003313.

Rodrigues:2009:TDM

[ROQ^{+}09] M. Rodrigues, A. Oliveira, H. Queiroga, A. B. Fortunato, and Y. J. Zhang. Three-dimensional modeling of the lower trophic levels in the Ria de Aveiro (Portugal). Ecological Modelling, 220(9–10):1274–1290, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000982.

Ross:2009:NDD

[Ros09] J. V. Ross. A note on density dependence in population models. Ecological Modelling, 220(23):3472–3474, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006097.

Ross:2007:CDP

[RP07] J. V. Ross and P. K. Pollett. On costs and decisions in population management. Ecological Modelling, 201(1):60–66, February 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006003553.

Riegl:2009:MCP

[RP09] Bernhard M. Riegl and Samuel J. Purkis. Model of coral population response to accelerated bleaching and mass mortality in a changing climate. Ecological Modelling, 220(2):192–208, January 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004559.

Roy:2005:VGM

[RPC^{+}05] P. S. Roy, Hitendra Padalia, Nidhi Chauhan, M. C. Porwal, S. Gupta, S. Biswas, and R. Jagdale. Validation of geospatial model for biodiversity characterization at landscape level — a study in Andaman & Nicobar Islands, India. Ecological Modelling, 185(2–4):349–369, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000062.

Rivot:2004:BSS

[RPPB04] E. Rivot, E. Prévost, E. Parent, and J. L. Baglinière. A Bayesian state-space modelling framework for fitting a salmon stage-structured population dynamic model to multiple time series of field data. Ecological Modelling, 179(4):463–485, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002728.

Reimoser:2009:RDH

[RPRV09] Susanne Reimoser, Ernst Partl, Friedrich Reimoser, and Sonja Vospernik. Roe-deer habitat suitability and predisposition of forest to browsing damage in its dependence on forest growth — model sensitivity in an alpine forest region. Ecological Modelling, 220(18):2231–2243, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009003871.

Robertson:2003:CMP

[RPVR03] Mark P. Robertson, Craig I. Peter, Martin H. Villet, and Brad S. Ripley. Comparing models for predicting species’ potential distributions: a case study using correlative and mechanistic predictive modelling techniques. Ecological Modelling, 164(2–3):153–167, June 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000280.

Ren:2001:DEB

[RR01] Jeffrey S. Ren and Alex H. Ross. A dynamic energy budget model of the Pacific oyster Crassostrea gigas. Ecological Modelling, 142(1–2):105–120, August 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002824.

Ren:2005:EIM

[RR05a] Jeffrey S. Ren and Alex H. Ross. Environmental influence on mussel growth: a dynamic energy budget model and its application to the greenshell mussel Perna canaliculus. Ecological Modelling, 189(3–4):347–362, December 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005002152.

Retzer:2005:MCC

[RR05b] V. Retzer and C. Reudenbach. Modelling the carrying capacity and coexistence of pika and livestock in the mountain steppe of the South Gobi, Mongolia. Ecological Modelling, 189(1–2):89–104, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001894.

Radtke:2006:BSC

[RR06] Philip J. Radtke and Andrew P. Robinson. A Bayesian strategy for combining predictions from empirical and processbased models. Ecological Modelling, 190(3–4):287–298, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002632.

Rocchini:2007:LCW

[RR07] Duccio Rocchini and Carlo Ricotta. Are landscapes as crisp as we may think? Ecological Modelling, 204(3–4):535–539, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006006661.

Regan:2001:MIT

[RRB^{+}01] T. J. Regan, H. M. Regan, K. Bonham, R. J. Taylor, and M. A. Burgman. Modelling the impact of timber harvesting on a rare carnivorous land snail (Tasmaphena lamproides) in northwest Tasmania, Australia. Ecological Modelling, 139(2–3):253–264, April 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100240X.

Ray:2001:IDF

[RS01] Santanu Ray and Milan Straškraba. The impact of detritivorous fishes on a mangrove estuarine system. Ecological Modelling, 140(3):207–218, June 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003210.

Richter:2004:FGI

[RS04] O. Richter and R. Seppelt. Flow of genetic information through agricultural ecosystems: a generic modelling framework with application to pesticide-resistance weeds and genetically modified crops. Ecological Modelling, 174(1–2):55– 66, May 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005684.

Raventos:2004:GDT

[RSA04] J. Raventós, J. Segarra, and M. F. Acevedo. Growth dynamics of tropical savanna grass species using projection matrices. Ecological Modelling, 174(1–2):85–101, May 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003005702.

Roberts:2002:MCH

[RSB02] Catherine A. Roberts, Doug Stallman, and Joanna A. Bieri. Modeling complex human–environment interactions: the Grand Canyon river trip simulator. Ecological Modelling, 153(1–2):181–196, July 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005099.

Risch:2009:PLT

[RSB09] Anita C. Risch, Martin Schütz, and Harald Bugmann. Predicting long-term development of abandoned subalpine conifer forests in the Swiss National Park. Ecological Modelling, 220(13–14):1578–1585, July 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002609.

Rahman:2009:SEH

[RSC09] Md. Rejaur Rahman, Z. H. Shi, and Cai Chongfa. Soil erosion hazard evaluation — an integrated use of remote sensing, GIS and statistical approaches with biophysical parameters towards management strategies. Ecological Modelling, 220(13–14):1724–1734, July 17, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002634.

Raghunathan:2001:EDF

[RSF^{+}01] Ramesh Raghunathan, Tad Slawecki, Thomas D. Fontaine, Zhenquan Chen, David W. Dilks, Victor J. Bierman, and Scott Wade. Exploring the dynamics and fate of total phosphorus in the Florida Everglades using a calibrated mass balance model. Ecological Modelling, 142(3):247–259, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003258.

Ruger:2005:FHS

[RSM05] Nadja Rüger, Maja Schlüter, and Michael Matthies. A fuzzy habitat suitability index for Populus euphratica in the Northern Amudarya delta (Uzbekistan). Ecological Modelling, 184(2–4):313–328, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005575.

Rabouille:2005:FAM

[RST05] Sophie Rabouille, Marie-José Salençon, and Jean-Marc Thébault. Functional analysis of Microcystis vertical migration: a dynamic model as a prospecting tool: I — processes analysis. Ecological Modelling, 188(2–4):386–403, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005001201.

Rode:2007:MOC

[RSW07] Michael Rode, Ursula Suhr, and Gunter Wriedt. Multiobjective calibration of a river water quality model — information content of calibration data. Ecological Modelling, 204(1–2):129–142, May 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006697.

Ribeiro:2008:CSP

[RT08] Rita Ribeiro and Luís Torgo. A comparative study on predicting algae blooms in Douro River, Portugal. Ecological Modelling, 212(1–2):86–91, March 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005133.

Roy:2004:FDV

[RTB04] A. Roy, S. K. Tripathi, and S. K. Basu. Formulating diversity vector for ecosystem comparison. Ecological Modelling, 179(4):499–513, December 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002753.

Rinke:2005:MAE

[RV05] Karsten Rinke and Jacobus Vijverberg. A model approach to evaluate the effect of temperature and food concentration on individual life-history and population dynamics of Daphnia. Ecological Modelling, 186(3):326–344, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000566.

Rietkerk:2002:ES

[RvdKK^{+}02] Max Rietkerk, Johan van de Koppel, Lalit Kumar, Herbert H. T. Langevelde, and Prins. The ecology of scale. Ecological Modelling, 149(1–2):1–4, March 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005105.

Recknagel:2008:GLM

[RvGC^{+}08] Friedrich Recknagel, Carin van Ginkel, Hongqing Cao, Lydia Cetin, and Byron Zhang. Generic limnological models on the touchstone: Testing the lake simulation library SALMO-OO and the rule-based Microcystis agent for warmmonomictic hypertrophic lakes in South Africa. Ecological Modelling, 215(1–3):144–158, July 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000987.

Romanowicz:2005:SSM

[RVRL05] A. A. Romanowicz, M. Vanclooster, M. Rounsevell, and I. La Jeunesse. Sensitivity of the SWAT model to the soil and land use data parametrisation: a case study in the Thyle catchment, Belgium. Ecological Modelling, 187(1):27–39, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000396.

Reineking:2006:EVA

[RVWH06] Björn Reineking, Maik Veste, Christian Wissel, and Andreas Huth. Environmental variability and allocation trade-offs maintain species diversity in a process-based model of succulent plant communities. Ecological Modelling, 199(4):486–504, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002924.

Rose:2007:SHG

[RWM^{+}07] Kenneth A. Rose, Francisco E. Werner, Bernard A. Megrey, Maki Noguchi Aita, Yasuhiro Yamanaka, Douglas E. Hay, Jake F. Schweigert, and Matthew Birch Foster. Simulated herring growth responses in the Northeastern Pacific to historic temperature and zooplankton conditions generated by the 3-dimensional NEMURO nutrient–phytoplankton– zooplankton model. Ecological Modelling, 202(1–2):184–195, March 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006004820.

Richmond:2007:CCI

[RWW07] Courtney E. Richmond, David S. Wethey, and Sarah A. Woodin. Climate change and increased environmental variability: Demographic responses in an estuarine harpacticoid copepod. Ecological Modelling, 209(2–4):189–202, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007003389.

Rajar:2004:ATD

[RŽČ^{+}04] Rudi Rajar, Dušan Žagar, Matjaž Četina, Hirokatsu Akagi, Shinichiro Yano, Takashi Tomiyasu, and Milena Horvat. Application of three-dimensional mercury cycling model to coastal seas. Ecological Modelling, 171(1–2):139–155, January 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003003296.

Sundelof:2006:BFS

[SÅ06] Andreas Sundelöf and Per Åberg. Birth functions in stage structured two-sex models. Ecological Modelling, 193(3–4):787–795, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003996.

Saloranta:2007:MMY

[SA07a] Tuomo M. Saloranta and Tom Andersen. MyLake — a multi-year lake simulation model code suitable for uncertainty and sensitivity analysis simulations. Ecological Modelling, 207(1):45–60, September 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001548.

Schuol:2007:UMW

[SA07b] J. Schuol and K. C. Abbaspour. Using monthly weather statistics to generate daily data in a SWAT model application to West Africa. Ecological Modelling, 201 (3–4):301–311, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004637.

Saarikko:2000:ASB

[Saa00] Riitta A. Saarikko. Applying a site based crop model to estimate regional yields under current and changed climates. Ecological Modelling, 131(2–3):191–206, July 1, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000000257X.

Sharma:2003:FSD

[SAB03] Mahadev Sharma, Ralph L. Amateis, and Harold E. Burkhart. Forest stand dynamics and similarity theory. Ecological Modelling, 167(1–2):165–180, September 1, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003001753.

Scotti:2006:ETP

[SAB^{+}06] Marco Scotti, Stefano Allesina, Cristina Bondavalli, Antonio Bodini, and Luis Gerardo Abarca-Arenas. Effective trophic positions in ecological acyclic networks. Ecological Modelling, 198(3–4):495–505, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002638.

Salinger:2003:PER

[SAH03] David H. Salinger, James J. Anderson, and Owen S. Hamel. A parameter estimation routine for the vitality-based survival model. Ecological Modelling, 166(3):287–294, August 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001625.

Saisho:2007:SMD

[Sai07] Yasumasa Saisho. A stochastic model describing the number of male insects which are left without partners. Ecological Modelling, 207(2–4):155–158, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700244X.

Serizawa:2009:PBC

[SAI09] H. Serizawa, T. Amemiya, and K. Itoh. Patchiness and bistability in the comprehensive cyanobacterial model (CCM). Ecological Modelling, 220(6):764–773, March 24, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000800598X.

Sazykina:2000:SOT

[SAK00] T. G. Sazykina, V. V. Alekseev, and A. I. Kryshev. The selforganization of trophic structure in ecosystem models: the succession phenomena, trigger regimes and hysteresis. Ecological Modelling, 133(1–2):83–94, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002829.

Sakanoue:2007:ELM

[Sak07] Seiichi Sakanoue. Extended logistic model for growth of single-species populations. Ecological Modelling, 205 (1–2):159–168, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007000877.

Sakanoue:2009:RBA

[Sak09] Seiichi Sakanoue. A resource-based approach to modelling the dynamics of interacting populations. Ecological Modelling, 220(11):1383–1394, June 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001859.

Salthe:2002:ENP

[Sal02] S. N. Salthe. An exercise in the natural philosophy of ecology. Ecological Modelling, 158(3):167–179, December 31, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002211.

Saloranta:2006:HMC

[Sal06] Tuomo M. Saloranta. Highlighting the model code selection and application process in policy-relevant water quality modelling. Ecological Modelling, 194(1–3):316–327, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005296.

Scopelitis:2007:MCR

[SAL07] Julie Scopélitis, Serge Andréfouët, and Christine Largouët. Modelling coral reef habitat trajectories: Evaluation of an integrated timed automata and remote sensing approach. Ecological Modelling, 205(1–2):59–80, July 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007000622.

Sandberg:2007:CEA

[San07] J. Sandberg. Cross-ecosystem analyses of pelagic food web structure and processes in the Baltic Sea. Ecological Modelling, 201(3–4):243–261, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004224.

Santojanni:2009:CAE

[San09] Alberto Santojanni. Comments on “Is anchovy (Engraulis encrasicolus L.) overfished in the Adriatic Sea?” by Klanjšček and Legović [Ecol. Model. 201 (2007) 312– 316]. Ecological Modelling, 220(3):430–433, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008002251.

Sarkar:2004:SMA

[Sar04] R. R. Sarkar. A stochastic model for autotroph–herbivore system with nutrient reclycing. Ecological Modelling, 178 (3–4):429–440, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001863.

Segarra:2009:CSW

[SAR^{+}09] J. Segarra, M. Acevedo, J. Raventós, C. Garcia-Núñez, and J. F. Silva. Coupling soil water and shoot dynamics in three grass species: a spatial stochastic model on water competition in neotropical savanna. Ecological Modelling, 220(20):2734–2743, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004530.

Shanmugam:2006:CCW

[SAS06] Palanisamy Shanmugam, Yu-Hwan Ahn, and Shanmugam Sanjeevi. A comparison of the classification of wetland characteristics by linear spectral mixture modelling and traditional hard classifiers on multispectral remotely sensed imagery in southern India. Ecological Modelling, 194(4):379–394, April 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005005430.

Savard:2000:MRT

[Sav00] Marielle Savard. Modelling risk, trade, agricultural and environmental policies to assess trade-offs between water quality and welfare in the hog industry. Ecological Modelling, 125(1):51–66, January 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001738.

Slavtchova-Bojkova:2000:CWT

[SB00] Maroussia Slavtchova-Bojkova. Computation of waiting time to successful experiment using age-dependent branching model. Ecological Modelling, 133(1–2):125–130, August 25, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002866.

Saysel:2001:DMS

[SB01] Ali Kerem Saysel and Yaman Barlas. A dynamic model of salinization on irrigated lands. Ecological Modelling, 139(2–3):177–199, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002423.

Smallegange:2002:FSD

[SB02] Isabel M. Smallegange and Arend M. H. Brunsting. Food supply and demand, a simulation model of the functional response of grazing ruminants. Ecological Modelling, 149(1–2):179–192, March 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005221.

Schley:2003:DDS

[SB03] D. Schley and M. A. Bees. Delay dynamics of the slug Deroceras reticulatum, an agricultural pest. Ecological Modelling, 162(3):177–198, April 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003587.

Salles:2006:MPC

[SB06a] Paulo Salles and Bert Bredeweg. Modelling population and community dynamics with qualitative reasoning. Ecological Modelling, 195(1–2):114–128, May 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005831.

Samaniego:2006:SIL

[SB06b] Luis Samaniego and András Bárdossy. Simulation of the impacts of land use/cover and climatic changes on the runoff characteristics at the mesoscale. Ecological Modelling, 196(1–2):45–61, July 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000548.

Schley:2006:RTD

[SB06c] David Schley and Martin A. Bees. The role of time delays in a non-autonomous host–parasitoid model of slug biocontrol with nematodes. Ecological Modelling, 193 (3–4):543–559, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004278.

Strauss:2007:ETS

[SB07] B. Strauss and R. Biedermann. Evaluating temporal and spatial generality: How valid are species–habitat relationship models? Ecological Modelling, 204(1–2):104–114, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000600665X.

Salles:2006:QMA

[SBA06] Paulo Salles, Bert Bredeweg, and Symone Araújo. Qualitative models about stream ecosystem recovery: Exploratory studies. Ecological Modelling, 194(1–3):80–89, March 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005004990.

Salles:2006:AGQ

[SBB06] Paulo Salles, Bert Bredeweg, and Nurit Bensusan. The ants’ garden: Qualitative models of complex interactions between populations. Ecological Modelling, 194(1–3):90–101, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005004.

Scotti:2009:LTP

[SBB09] Marco Scotti, Cristina Bondavalli, and Antonio Bodini. Linking trophic positions and flow structure constraints in ecological networks: Energy transfer efficiency or topology effect? Ecological Modelling, 220(21):3070–3080, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003852.

Sweeney:2007:AEF

[SBC07] A. W. Sweeney, N. W. Beebe, and R. D. Cooper. Analysis of environmental factors influencing the range of anopheline mosquitoes in northern Australia using a genetic algorithm and data mining methods. Ecological Modelling, 203(3–4):375–386, May 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006260.

Solidoro:2009:CEM

[SBC^{+}09] Cosimo Solidoro, Vinko Bandelj, Gianpiero Cossarini, Simone Libralato, and Donata Melaku Canu. Challenges for ecological modelling in a changing world: Global changes, sustainability and ecosystem based management. Ecological Modelling, 220(21):2825–2827, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900578X.

Seoane:2004:EVM

[SBDD04a] Javier Seoane, Javier Bustamante, and Ricardo Díaz-Delgado. Are existing vegetation maps adequate to predict bird distributions? Ecological Modelling, 175(2):137–149, July 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003004733.

Seoane:2004:CRL

[SBDD04b] Javier Seoane, Javier Bustamante, and Ricardo Díaz-Delgado. Competing roles for landscape, vegetation, topography and climate in predictive models of bird distribution. Ecological Modelling, 171(3):209–222, January 15, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000300334X.

Scanlan:2006:PDE

[SBG06] J. C. Scanlan, D. M. Berman, and W. E. Grant. Population dynamics of the European rabbit (Oryctolagus cuniculus) in north eastern Australia: Simulated responses to control. Ecological Modelling, 196(1–2):221–236, July 10, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006000615.

Strandesen:2007:FDM

[SBHH07] M. Strandesen, M. Birkved, P. E. Holm, and M. Z. Hauschild. Fate and distribution modelling of metals in life cycle impact assessment. Ecological Modelling, 203(3–4):327–338, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006006211.

Spijkerboer:2002:AGP

[SBJ^{+}02] H. P. Spijkerboer, J. E. Beniers, D. Jaspers, H. J. Schouten, J. Goudriaan, R. Rabbinge, and W. van der Werf. Ability of the Gaussian plume model to predict and describe spore dispersal over a potato crop. Ecological Modelling, 155(1):1–18, September 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004756.

Severini:2003:PED

[SBL03] Maurizio Severini, Johann Baumgärtner, and Lidia Limonta. Parameter estimation for distributed delay based population models from laboratory data: egg hatching of Oulema duftschmidi Redthenbacher (Coleoptera, Chrysomelidae) as an example. Ecological Modelling, 167(3):233–246, September 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001881.

Schumacher:2004:IFT

[SBM04] Sabine Schumacher, Harald Bugmann, and David J. Mladenoff. Improving the formulation of tree growth and succession in a spatially explicit landscape model. Ecological Modelling, 180(1):175–194, December 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003813.

Singh:2009:ANN

[SBMJ09] Kunwar P. Singh, Ankita Basant, Amrita Malik, and Gunja Jain. Artificial neural network modeling of the river water quality — a case study. Ecological Modelling, 220(6):888–895, March 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000283.

Seidl:2007:MTM

[SBR^{+}07] Rupert Seidl, Peter Baier, Werner Rammer, Axel Schopf, and Manfred J. Lexer. Modelling tree mortality by bark beetle infestation in Norway spruce forests. Ecological Modelling, 206(3–4):383–399, August 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001998.

Stockwell:2006:UGG

[SBS^{+}06] David R. B. Stockwell, James H. Beach, Aimee Stewart, Gregory Vorontsov, David Vieglais, and Ricardo Scachetti Pereira. The use of the GARP genetic algorithm and Internet grid computing in the Lifemapper world atlas of species biodiversity. Ecological Modelling, 195(1–2):139–145, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005855.

Silva:2005:PEI

[SBVB05] Carla Silva, Cristina Boia, Joana Valente, and Carlos Borrego. Pesticides in Esteros del Ibera (AR): evaluation of impacts and proposal of guidelines for water quality protection. Ecological Modelling, 186(1):85–97, July 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000220.

Schomberg:2001:MSM

[SC01] Harry H. Schomberg and Miguel L. Cabrera. Modeling in situ N mineralization in conservation tillage fields: comparison of two versions of the CERES nitrogen submodel. Ecological Modelling, 145(1):1–15, November 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003799.

Smith:2002:MMC

[SC02] G. C. Smith and C. L. Cheeseman. A mathematical model for the control of diseases in wildlife populations: culling, vaccination and fertility control. Ecological Modelling, 150(1–2):45–53, April 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004719.

Santos:2004:STE

[SC04] Silvia N. Monteiro Santos and Marcos Heil Costa. A simple tropical ecosystem model of carbon, water and energy fluxes. Ecological Modelling, 176(3–4):291–312, September 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004000936.

Schmickl:2007:HMH

[SC07] Thomas Schmickl and Karl Crailsheim. HoPoMo: a model of honeybee intracolonial population dynamics and resource management. Ecological Modelling, 204(1–2):219–245, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007000075.

Scardi:2001:ANN

[Sca01] Michele Scardi. Advances in neural network modeling of phytoplankton primary production. Ecological Modelling, 146(1–3):33–45, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002940.

Scardi:2006:P

[Sca06] Michele Scardi. Preface. Ecological Modelling, 195(1–2):1, May 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005005685.

Seoane:2005:SST

[SCAP05] Javier Seoane, Luis M. Carrascal, César Luis Alonso, and David Palomino. Species-specific traits associated to prediction errors in bird habitat suitability modelling. Ecological Modelling, 185(2–4):299–308, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000037.

Scharfe:2009:SLM

[SCB^{+}09] Mirco Scharfe, Ulrich Callies, Gerd Blöcker, Wilhelm Petersen, and Friedhelm Schroeder. A simple Lagrangian model to simulate temporal variability of algae in the Elbe River. Ecological Modelling, 220(18):2173–2186, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009003147.

Saltre:2009:PMD

[SCBG09] F. Saltré, I. Chuine, S. Brewer, and C. Gaucherel. A phenomenological model without dispersal kernel to model species migration. Ecological Modelling, 220 (24):3546–3554, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004098.

Schultink:2000:CEI

[Sch00] Gerhardus Schultink. Critical environmental indicators: performance indices and assessment models for sustainable rural development planning. Ecological Modelling, 130(1–3):47–58, June 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000000212X.

Schulte:2003:APSb

[Sch03] R. P. O. Schulte. Analysis of the production stability of mixed grasslands: II: a mathematical framework for the quantification of production stability in grassland ecosystems. Ecological Modelling, 159(1):71–99, January 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002624.

Sharma:2005:QFA

[SCH05] Rishi Sharma, Andrew B. Cooper, and Ray Hilborn. A quantitative framework for the analysis of habitat and hatchery practices on Pacific salmon. Ecological Modelling, 183(2–3):231–250, April 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004818.

Simonit:2005:MHE

[SCP05] Silvio Simonit, Franco Cattaneo, and Charles Perrings. Modelling the hydrological externalities of agriculture in wetlands: the case of rice in Esteros del Iberà, Argentina. Ecological Modelling, 186(1):123–141, July 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000311.

Sebert-Cuvillier:2007:LPD

[SCPC^{+}07] Emmanuelle Sebert-Cuvillier, Frédéric Paccaut, Olivier Chabrerie, Patrick Endels, Olivier Goubet, and Guillaume Decocq. Local population dynamics of an invasive tree species with a complex life-history cycle: a stochastic matrix model. Ecological Modelling, 201(2):127–143, February 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004157.

Solidoro:2003:EEC

[SCR03] Cosimo Solidoro, D. Melaku Canu, and R. Rossi. Ecological and economic considerations on fishing and rearing of Tapes phillipinarum in the lagoon of Venice. Ecological Modelling, 170(2–3):303–318, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002357.

Shastri:2008:OCL

[SD08] Y. Shastri and U. Diwekar. Optimal control of lake pH for mercury bioaccumulation control. Ecological Modelling, 216(1):1–17, August 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001518.

Schmitz:2003:RBL

[SDA^{+}03] M. F. Schmitz, I. De Aranzabal, P. Aguilera, A. Rescia, and F. D. Pineda. Relationship between landscape typology and socioeconomic structure: Scenarios of change in Spanish cultural landscapes. Ecological Modelling, 168(3):343–356, October 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001455.

Skiba:2003:EIV

[SDB03] Yuri N. Skiba and Valentina Davydova-Belitskaya. On the estimation of impact of vehicular emissions. Ecological Modelling, 166(1–2):169–184, August 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001339.

Sabatier:2004:MMS

[SDD^{+}04] P. Sabatier, B. Durand, M. A. Dubois, C. Ducrot, D. Calavas, and A. Van de Wielle. Multiscale modelling of scrapie epidemiology: I. Herd level: a discrete model of disease transmission in a sheep flock. Ecological Modelling, 180(2–3):233–252, December 25, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002765.

Sester:2007:MEC

[SDDC07] Mathilde Sester, Carolyne Dürr, Henri Darmency, and Nathalie Colbach. Modelling the effects of cropping systems on the seed bank dynamics and the emergence of weed beet. Ecological Modelling, 204(1–2):47–58, May 24, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006006417.

Soetaert:2002:FFE

[SdH02] Karline Soetaert, Vanni deClippele, and Peter Herman. FEMME, a flexible environment for mathematically modelling the environment. Ecological Modelling, 151(2–3):177–193, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004690.

Steppe:2008:VDS

[SDL08] Kathy Steppe, Dirk J. W. De Pauw, and Raoul Lemeur. Validation of a dynamic stem diameter variation model and the resulting seasonal changes in calibrated parameter values. Ecological Modelling, 218(3–4):247–259, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003475.

Stollenwerk:2001:TNS

[SDS01] Nico Stollenwerk, Friedhelm R. Drepper, and Helge Siegel. Testing nonlinear stochastic models on phytoplankton biomass time series. Ecological Modelling, 144 (2–3):261–277, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003775.

Scheller:2007:DDA

[SDS^{+}07] Robert M. Scheller, James B. Domingo, Brian R. Sturtevant, Jeremy S. Williams, Arnold Rudy, Eric J. Gustafson, and David J. Mladenoff. Design, development, and application of LANDIS-II, a spatial landscape simulation model with flexible temporal and spatial resolution. Ecological Modelling, 201(3–4):409–419, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004893.

Schwalm:2004:PBM

[SE04] Christopher R. Schwalm and Alan R. Ek. A process-based model of forest ecosystems driven by meteorology. Ecological Modelling, 179(3):317–348, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400242X.

Sekine:2008:P

[Sek08] Masahiko Sekine. Preface. Ecological Modelling, 215(1–3):1– 2, July 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008001099.

Selhorst:2000:IOI

[Sel00] T. Selhorst. Improving the oral immunization of foxes (Vulpes vulpes) against rabies with the help of an evolutionary algorithm. Ecological Modelling, 129(2–3):297–305, May 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000002350.

Semenov:2008:BCG

[Sem08] D. A. Semenov. Biological contribution to global climate dynamics on a “geological” timescale. Ecological Modelling, 212(1–2):171–177, March 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005157.

Seppelt:2000:ROC

[Sep00] Ralf Seppelt. Regionalised optimum control problems for agroecosystem management. Ecological Modelling, 131(2–3):121–132, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002702.

Syphard:2004:SAE

[SF04] Alexandra D. Syphard and Janet Franklin. Spatial aggregation effects on the simulation of landscape pattern and ecological processes in southern California plant communities. Ecological Modelling, 180(1):21–40, December 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004003473.

Sanchez-Flores:2007:GMN

[SF07] Erick Sánchez-Flores. GARP modeling of natural and human factors affecting the potential distribution of the invasives Schismus arabicus and Brassica tournefortii in ‘El Pinacate y Gran Desierto de Altar’ Biosphere Reserve. Ecological Modelling, 204(3–4):457–474, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000555.

Scharler:2009:CNA

[SF09] Ursula M. Scharler and Brian D. Fath. Comparing network analysis methodologies for consumer–resource relations at species and ecosystems scales. Ecological Modelling, 220(22):3210–3218, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001252.

Scull:2005:ACT

[SFC05] P. Scull, J. Franklin, and O. A. Chadwick. The application of classification tree analysis to soil type prediction in a desert landscape. Ecological Modelling, 181(1):1–15, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003540.

Soares-Filho:2002:DSC

[SFCP02] Britaldo Silveira Soares-Filho, Gustavo Coutinho Cerqueira, and Cássio Lopes Pennachin. dinamica — a stochastic cellular automata model designed to simulate the landscape dynamics in an Amazonian colonization frontier. Ecological Modelling, 154(3):217–235, September 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000595.

Shannon:2004:SAS

[SFM04] Lynne J. Shannon, John G. Field, and Coleen L. Moloney. Simulating anchovy–sardine regime shifts in the southern Benguela ecosystem. Ecological Modelling, 172(2–4):269–281, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003003739.

Schmidt:2008:ULS

[SFM08] Thomas G. Schmidt, Uwe Franko, and Ralph Meissner. Uncertainties in large-scale analysis of agricultural land use — a case study for simulation of nitrate leaching. Ecological Modelling, 217(1–2):174–180, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003049.

Sanderson:2001:SEM

[SFU01] Eric W. Sanderson, Theodore C. Foin, and Susan L. Ustin. A simple empirical model of salt marsh plant spatial distributions with respect to a tidal channel network. Ecological Modelling, 139(2–3):293–307, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002538.

Srinivasu:2005:IPR

[SG05] P. D. N. Srinivasu and I. L. Gayatri. Influence of prey reserve capacity on predator–prey dynamics. Ecological Modelling, 181(2–3):191–202, January 20, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003667.

Schramski:2009:CDI

[SG09] John R. Schramski and David K. Gattie. Cross-disciplinary informed ecological network theory. Ecological Modelling, 220(22):3111–3112, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006188.

Swanack:2008:UMS

[SGF08] Todd M. Swanack, William E. Grant, and Brian D. Fath. On the use of multi-species NK models to explore ecosystem development. Ecological Modelling, 218(3–4):367–374, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008003670.

Swanack:2009:PPT

[SGF09] Todd M. Swanack, William E. Grant, and Michael R. J. Forstner. Projecting population trends of endangered amphibian species in the face of uncertainty: a pattern-oriented approach. Ecological Modelling, 220(2):148–159, January 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008004365.

Sturtevant:2004:MDS

[SGH04] Brian R. Sturtevant, Eric J. Gustafson, and Hong S. He. Modeling disturbance and succession in forest landscapes using LANDIS: introduction. Ecological Modelling, 180(1):1–5, December 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000400345X.

Skjoth:2008:ITS

[SGH^{+}08] Carsten A. Skjøth, Camilla Geels, Martin Hvidberg, Ole Hertel, Jørgen Brandt, Lise M. Frohn, Kaj M. Hansen, Gitte B. Hedegaard, Jesper H. Christensen, and L. Moseholm. An inventory of tree species in Europe — an essential data input for air pollution modelling. Ecological Modelling, 217(3–4):292–304, October 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002640.

Svoray:2004:AHP

[SGHG04] Tal Svoray, Shiri Bar-Yamin Gancharski, Zalmen Henkin, and Mario Gutman. Assessment of herbaceous plant habitats in water-constrained environments: predicting indirect effects with fuzzy logic. Ecological Modelling, 180 (4):537–556, December 31, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003527.

Scanlan:2001:HEF

[SGHM01] J. C. Scanlan, W. E. Grant, D. M. Hunter, and R. J. Milner. Habitat and environmental factors influencing the control of migratory locusts (Locusta migratoria) with an entomopathogenic fungus (Metarhizium anisopliae). Ecological Modelling, 136(2–3):223–236, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004245.

Sturtevant:2004:MBD

[SGLH04] Brian R. Sturtevant, Eric J. Gustafson, Wei Li, and Hong S. He. Modeling biological disturbances in LANDIS: a module description and demonstration using spruce budworm. Ecological Modelling, 180(1):153–174, December 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003801.

Sole:2005:MAA

[SGLRE05] Jordi Solé, Emilio García-Ladona, Piet Ruardij, and Marta Estrada. Modelling allelopathy among marine algae. Ecological Modelling, 183(4):373–384, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004971.

Schramski:2006:IED

[SGP^{+}06] John R. Schramski, D. K. Gattie, B. C. Patten, S. R. Borrett, B. D. Fath, C. R. Thomas, and S. J. Whipple. Indirect effects and distributed control in ecosystems: Distributed control in the environ networks of a seven-compartment model of nitrogen flow in the Neuse River Estuary, USA — steady-state analysis. Ecological Modelling, 194(1–3):189–201, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005168.

Schramski:2007:IED

[SGP^{+}07] J. R. Schramski, D. K. Gattie, B. C. Patten, S. R. Borrett, B. D. Fath, and S. J. Whipple. Indirect effects and distributed control in ecosystems: Distributed control in the environ networks of a seven-compartment model of nitrogen flow in the Neuse River Estuary, USA — time series analysis. Ecological Modelling, 206(1–2):18–30, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001445.

Shang:2001:DEM

[SGY01] Zongbo Shang, Qiong Gao, and Dianan Yang. Digital ecological model and case study on China water condition. Ecological Modelling, 139(2–3):235–252, April 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002435.

Swaney:2004:OTB

[SH04] Dennis P. Swaney and Charles A. S. Hall. Odum in Texas: a brief review of H. T. Odum’s Texas Bays studies. Ecological Modelling, 178(1–2):59–63, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005507.

Santika:2009:ESR

[SH09a] Truly Santika and Michael F. Hutchinson. The effect of species response form on species distribution model prediction and inference. Ecological Modelling, 220(19): 2365–2379, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004049.

Simunek:2009:MCR

[ŠH09b] Jiří Šimůnek and Jan W. Hopmans. Modeling compensated root water and nutrient uptake. Ecological Modelling, 220(4):505–521, February 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005383.

Schauser:2004:SMP

[SHB04] Inke Schauser, Michael Hupfer, and Rainer Brüggemann. SPIEL — a model for phosphorus diagenesis and its application to lake restoration. Ecological Modelling, 176(3–4):389–407, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001012.

Schauser:2006:SAP

[SHB06] Inke Schauser, Michael Hupfer, and Rainer Brüggemann. Sensitivity analysis with a phosphorus diagenesis model (SPIEL). Ecological Modelling, 190(1–2):87–98, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002607.

Shang:2004:FLR

[SHCS04] Bo Z. Shang, Hong S. He, Thomas R. Crow, and Stephen R. Shifley. Fuel load reductions and fire risk in central hardwood forests of the United States: a spatial simulation study. Ecological Modelling, 180(1):89–102, December 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004003515.

Shen:2006:OEP

[She06] Jian Shen. Optimal estimation of parameters for a estuarine eutrophication model. Ecological Modelling, 191 (3–4):521–537, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002887.

Skov:2008:AHS

[SHG^{+}08] Henrik Skov, Elizabeth Humphreys, Stefan Garthe, Kerstin Geitner, David Grémillet, Keith C. Hamer, Janos Hennicke, Hjalte Parner, and Sarah Wanless. Application of habitat suitability modelling to tracking data of marine animals as a means of analyzing their feeding habitats. Ecological Modelling, 212(3–4):504–512, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005984.

Shimatani:2004:SME

[Shi04a] Kenichiro Shimatani. Spatial molecular ecological models for genotyped adults and offspring. Ecological Modelling, 174(4):401–410, June 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004617.

Shiyomi:2004:HDB

[Shi04b] Masae Shiyomi. How are distances between individuals of grazing cows explained by a statistical model? Ecological Modelling, 172(1):87–94, February 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003636.

Song:2007:SOM

[SHK^{+}07] Mi-Young Song, Hyun-Ju Hwang, Inn-Sil Kwak, Chang Woo Ji, Yong-Nam Oh, Byung Jin Youn, and Tae-Soo Chon. Selforganizing mapping of benthic macroinvertebrate communities implemented to community assessment and water quality evaluation. Ecological Modelling, 203(1–2):18–25, April 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006005540.

Swain:2007:USE

[SHM07] D. L. Swain, M. R. Hutchings, and G. Marion. Using a spatially explicit model to understand the impact of search rate and search distance on spatial heterogeneity within an herbivore grazing system. Ecological Modelling, 203(3–4):319–326, May 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006077.

Sileshi:2009:TOA

[SHN09] Gudeta Sileshi, Girma Hailu, and Gerson I. Nyadzi. Traditional occupancy–abundance models are inadequate for zero-inflated ecological count data. Ecological Modelling, 220(15):1764–1775, August 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001963.

Stehfest:2007:SGC

[SHP^{+}07] Elke Stehfest, Maik Heistermann, Joerg A. Priess, Dennis S. Ojima, and Joseph Alcamo. Simulation of global crop production with the ecosystem model DayCent. Ecological Modelling, 209(2–4):203–219, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003365.

Stickney:2000:IMP

[SHS00] H. L. Stickney, R. R. Hood, and D. K. Stoecker. The impact of mixotrophy on planktonic marine ecosystems. Ecological Modelling, 125(2–3):203–230, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001817.

Svirejeva-Hopkins:2008:UEC

[SHS08] A. Svirejeva-Hopkins and H.-J. Schellnhuber. Urban expansion and its contribution to the regional carbon emissions: Using the model based on the population density distribution. Ecological Modelling, 216(2):208–216, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001427.

Svirejeva-Hopkins:2004:UTS

[SHSP04] Anastasia Svirejeva-Hopkins, Hans J. Schellnhuber, and Valeri L. Pomaz. Urbanised territories as a specific component of the global carbon cycle. Ecological Modelling, 173(2–3):295–312, April 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004149.

Seymour:2004:PER

[SHW04] A. S. Seymour, S. Harris, and P. C. L. White. Potential effects of reserve size on incidental nest predation by red foxes Vulpes vulpes. Ecological Modelling, 175(1):101–114, June 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000300471X.

Sivakumar:2005:FIS

[SHZ05] B. Sivakumar, T. Harter, and H. Zhang. A fractal investigation of solute travel time in a heterogeneous aquifer: transition probability/Markov chain representation. Ecological Modelling, 182(3–4):355–370, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001693.

Su:2009:HDS

[SHZL09] Min Su, Cang Hui, Yanyu Zhang, and Zizhen Li. How does the spatial structure of habitat loss affect the eco-epidemic dynamics? Ecological Modelling, 220(1):51–59, January 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000800433X.

Schlicht:2006:DPL

[SI06] Robert Schlicht and Yoh Iwasa. Deviation from power law, spatial data of forest canopy gaps, and three lattice models. Ecological Modelling, 198(3–4):399–408, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002420.

Saldana:2007:PCS

[SI07] A. Saldaña and J. J. Ibáñez. Pedodiversity, connectance and spatial variability of soil properties, what is the relationship? Ecological Modelling, 208(2–4):342–352, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007003134.

Sato:2007:SDN

[SIK07] Hisashi Sato, Akihiko Itoh, and Takashi Kohyama. SEIB– DGVM: a new dynamic global vegetation model using a spatially explicit individual-based approach. Ecological Modelling, 200(3–4):279–307, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004145.

Silvert:2000:FIE

[Sil00] William Silvert. Fuzzy indices of environmental conditions. Ecological Modelling, 130(1–3):111–119, June 1, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380000002040.

Silverman:2004:GMM

[Sil04] Emily D. Silverman. A group movement model for waterfowl aggregation. Ecological Modelling, 175(4):411–424, August 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005362.

Silverman:2007:MEG

[Sil07] B. David Silverman. Modeling the effect of growth rate and survivability trade-offs on species coexistence and spatial topology at a traveling invasive wave-front. Ecological Modelling, 202(3–4):454–464, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005461.

Singkran:2007:AEM

[Sin07] Nuanchan Singkran. An abundance exchange model of fish assemblage response to changing habitat along embayment– stream gradients of Lake Ontario, New York. Ecological Modelling, 201(3–4):453–467, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004972.

Sato:2007:MIM

[SIS^{+}07] Tatsuaki Sato, Yugo Imazu, Takuya Sakawa, Takahiro Kazama, Takashi Wakabayashi, and Thomas Uhrenholdt. Modeling of integrated marine ecosystem including the generation-tracing type scallop growth model. Ecological Modelling, 208(2–4):263–285, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003092.

Store:2003:GBM

[SJ03] Ron Store and Jukka Jokimäki. A GIS-based multi-scale approach to habitat suitability modeling. Ecological Modelling, 169(1):1–15, November 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002035.

Schippers:2005:RTS

[SJ05] Peter Schippers and Eelke Jongejans. Release thresholds strongly determine the range of seed dispersal by wind. Ecological Modelling, 185(1):93–103, June 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005988.

Schneeberger:2006:EDR

[SJ06] Anne Schneeberger and Vincent A. A. Jansen. The estimation of dispersal rates using the covariance of local populations. Ecological Modelling, 196(3–4):434–446, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000950.

Svensson:2008:BCM

[SJG^{+}08] Magnus Svensson, Per-Erik Jansson, David Gustafsson, Dan Berggren Kleja, Ola Langvall, and Anders Lindroth. Bayesian calibration of a model describing carbon, water and heat fluxes for a Swedish boreal forest stand. Ecological Modelling, 213(3–4):331–344, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000069.

Strand:2005:MBR

[SJH05] Espen Strand, Christian Jørgensen, and Geir Huse. Modelling buoyancy regulation in fishes with swimbladders: bioenergetics and behaviour. Ecological Modelling, 185 (2–4):309–327, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000049.

Sun:2008:DCS

[SJLL08] Gui-Quan Sun, Zhen Jin, Quan-Xing Liu, and Li Li. Dynamical complexity of a spatial predator–prey model with migration. Ecological Modelling, 219(1–2):248–255, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008004079.

Sullivan:2003:MWS

[SJM03] Annett B. Sullivan, Henriette I. Jager, and Ralph Myers. Modeling white sturgeon movement in a reservoir: the effect of water quality and sturgeon density. Ecological Modelling, 167(1–2):97–114, September 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001698.

Snowling:2001:EMU

[SK01] S. D. Snowling and J. R. Kramer. Evaluating modelling uncertainty for model selection. Ecological Modelling, 138(1–3):17–30, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003902.

Shimatani:2004:SAC

[SK04] Kenichiro Shimatani and Yasuhiro Kubota. Spatial analysis for continuously changing point patterns along a gradient and its application to an Abies sachalinensis population. Ecological Modelling, 180(2–3):359–369, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000400287X.

Stegert:2007:PZP

[SKCM07] Christoph Stegert, Markus Kreus, Francois Carlotti, and Andreas Moll. Parameterisation of a zooplankton population model for Pseudocalanus elongatus using stage durations from laboratory experiments. Ecological Modelling, 206(3–4):213–230, August 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002360.

Skiba:2003:MDI

[Ski03] Yuri N. Skiba. On a method of detecting the industrial plants which violate prescribed emission rates. Ecological Modelling, 159(2–3):125–132, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200279X.

Skirvin:2004:VPM

[Ski04] David. J. Skirvin. Virtual plant models of predatory mite movement in complex plant canopies. Ecological Modelling, 171(3):301–313, January 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003594.

Smit:2006:TBR

[SKJvdHG06] Mathijs G. D. Smit, Belinda J. Kater, Robbert G. Jak, and Martine J. van den Heuvel-Greve. Translating bioassay results to field population responses using a Leslie-matrix model for the marine amphipod Corophium volutator. Ecological Modelling, 196(3–4):515–526, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000688.

Sugiura:2003:MMM

[SKK^{+}03] Katsura Sugiura, Yukishige Kawasaki, Minoru Kinoshita, Atsumi Murakami, Hiroaki Yoshida, and Yoshio Ishikawa. A mathematical model for microcosms: formation of the colonies and coupled oscillation in population densities of bacteria. Ecological Modelling, 168(1–2):173–201, October 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001960.

Sung:2006:SAM

[SKM^{+}06] Kijune Sung, Jaisoo Kim, C. L. Munster, M. Yavuz Corapcioglu, Soyoung Park, M. C. Drew, and Y. Y. Chang. A simple approach to modeling microbial biomass in the rhizosphere. Ecological Modelling, 190(3–4):277–286, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002620.

Schelhaas:2007:ITI

[SKP^{+}07] M. J. Schelhaas, K. Kramer, H. Peltola, D. C. van der Werf, and S. M. J. Wijdeven. Introducing tree interactions in wind damage simulation. Ecological Modelling, 207(2–4):197–209, October 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002578.

Shevtsov:2009:DEA

[SKP09] Jane Shevtsov, Caner Kazanci, and Bernard C. Patten. Dynamic environ analysis of compartmental systems: a computational approach. Ecological Modelling, 220(22):3219–3224, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000900492X.

Sirakoulis:2000:CAM

[SKT00] G. Ch Sirakoulis, I. Karafyllidis, and A. Thanailakis. A cellular automaton model for the effects of population movement and vaccination on epidemic propagation. Ecological Modelling, 133(3):209–223, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002945.

Stilma:2009:RAA

[SKvdW09] Eveline S. C. Stilma, Karel J. Keesman, and Wopke van der Werf. Recruitment and attrition of associated plants under a shading crop canopy: Model selection and calibration. Ecological Modelling, 220(8):1113–1125, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000313.

Seagle:2001:AFG

[SL01] Steven W. Seagle and Suh-Yuen Liang. Application of a forest gap model for prediction of browsing effects on riparian forest succession. Ecological Modelling, 144 (2–3):213–229, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003738.

Schulte:2002:MSV

[SL02] R. P. O. Schulte and E. A. Lantinga. Mechanistic simulation of the vertical structure of mixed swards. Ecological Modelling, 149(3):229–246, April 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004380.

Seghieri:2004:CVD

[SL04] Josiane Seghieri and Francis Laloë. Characterization of the variability of the daily course of leaf water potential in the dominant shrub species within Sahelian fallows in southwest Niger. Ecological Modelling, 173(2–3):271–281, April 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003004113.

Siracusa:2006:DCW

[SL06] G. Siracusa and A. D. La Rosa. Design of a constructed wetland for wastewater treatment in a Sicilian town and environmental evaluation using the emergy analysis. Ecological Modelling, 197(3–4):490–497, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001438.

Soethe:2007:RTB

[SLE07] N. Soethe, J. Lehmann, and C. Engels. Root tapering between branching points should be included in fractal root system analysis. Ecological Modelling, 207 (2–4):363–366, October 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007002918.

Simioni:2000:TPB

[SLGS00] G. Simioni, X. Le Roux, J. Gignoux, and H. Sinoquet. Treegrass: a 3D, process-based model for simulating plant interactions in tree–grass ecosystems. Ecological Modelling, 131(1):47–63, June 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000243X.

Shi:2006:LSM

[SLL^{+}06] Haijin Shi, Edward J. Laurent, Joseph LeBouton, Laila Racevskis, Kimberly R. Hall, Michael Donovan, Robert V. Doepker, Michael B. Walters, Frank Lupi, and Jianguo Liu. Local spatial modeling of white-tailed deer distribution. Ecological Modelling, 190(1–2):171–189, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000500219X.

Salvador:2005:DFO

[SLPP05] R. Salvador, F. Lloret, X. Pons, and J. Piñol. Does fire occurrence modify the probability of being burned again? A null hypothesis test from Mediterranean ecosystems in NE Spain. Ecological Modelling, 188(2–4):461–469, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000500092X.

Schulte:2003:APSa

[SLS03] R. P. O. Schulte, E. A. Lantinga, and P. C. Struik. Analysis of the production stability of mixed grasslands: I: a conceptual framework for the qualification of production stability in grassland ecosystems. Ecological Modelling, 159(1):43–69, January 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002612.

Strand:2009:SMS

[SLT^{+}09] Tara Strand, Brian Lamb, Harold Thistle, Eugene Allwine, and Holly Peterson. A simple model for simulation of insect pheromone dispersion within forest canopies. Ecological Modelling, 220(5):640–656, March 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005711.

Steinmann:2009:MPS

[SLZ09] K. Steinmann, H. P. Linder, and N. E. Zimmermann. Modelling plant species richness using functional groups. Ecological Modelling, 220(7):962–967, April 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000258.

Spieles:2003:MMT

[SM03] Douglas J. Spieles and William J. Mitsch. A model of macroinvertebrate trophic structure and oxygen demand in freshwater wetlands. Ecological Modelling, 161(3):183–194, March 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002003423.

Sara:2004:CCM

[SM04a] G. Sarà and A. Mazzola. The carrying capacity for Mediterranean bivalve suspension feeders: evidence from analysis of food availability and hydrodynamics and their integration into a local model. Ecological Modelling, 179 (3):281–296, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004001395.

Scheller:2004:FGB

[SM04b] Robert M. Scheller and David J. Mladenoff. A forest growth and biomass module for a landscape simulation model, LANDIS: design, validation, and application. Ecological Modelling, 180(1):211–229, December 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003837.

Spromberg:2006:RCT

[SM06] Julann A. Spromberg and James P. Meador. Relating chronic toxicity responses to population-level effects: a comparison of population-level parameters for three salmon species as a function of low-level toxicity. Ecological Modelling, 199(3):240–252, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002286.

Savina:2007:GBC

[SM07a] Marie Savina and Alain Ménesguen. The growth of the banded carpet shell (Paphia rhomboïdes) in a contrasted region, the English Channel (Western Europe): a modelling study. Ecological Modelling, 205(1–2):39–51, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007000300.

Schaefer:2007:GRS

[SM07b] James A. Schaefer and Stephen J. Mayor. Geostatistics reveal the scale of habitat selection. Ecological Modelling, 209(2–4):401–406, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003201.

Schimit:2009:BRN

[SM09] P. H. T. Schimit and L. H. A. Monteiro. On the basic reproduction number and the topological properties of the contact network: an epidemiological study in mainly locally connected cellular automata. Ecological Modelling, 220(7):1034–1042, April 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000386.

Scott:2006:EPS

[SMB^{+}06] Beth E. Scott, Gudrun Marteinsdottir, Gavin A. Begg, Peter J. Wright, and Olav Sigurd Kjesbu. Effects of population size/age structure, condition and temporal dynamics of spawning on reproductive output in Atlantic cod (Gadus morhua). Ecological Modelling, 191(3–4):383–415, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002772.

Stephenson:2006:MEP

[SMET06] C. M. Stephenson, M. L. MacKenzie, C. Edwards, and J. M. J. Travis. Modelling establishment probabilities of an exotic plant, Rhododendron ponticum, invading a heterogeneous, woodland landscape using logistic regression with spatial autocorrelation. Ecological Modelling, 193(3–4):747–758, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004783.

Scrinzi:2007:DNN

[SMG07] Gianfranco Scrinzi, Laura Marzullo, and David Galvagni. Development of a neural network model to update forest distribution data for managed alpine stands. Ecological Modelling, 206(3–4):331–346, August 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001950.

Senzia:2002:MNT

[SMM^{+}02] M. A. Senzia, A. W. Mayo, T. S. A. Mbwette, J. H. Y. Katima, and S. E. Jørgensen. Modelling nitrogen transformation and removal in primary facultative ponds. Ecological Modelling, 154(3):207–215, September 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000182.

Skalski:2008:EAM

[SMR08] John R. Skalski, Joshua J. Millspaugh, and Kristen E. Ryding. Effects of asymptotic and maximum age estimates on calculated rates of population change. Ecological Modelling, 212(3–4):528–535, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006047.

Seele:2000:CTI

[SMSR00] Jan Seele, Manuela Mayr, Frank Staab, and Uta Raeder. Combination of two indication systems in pre-alpine lakes — diatom index and macrophyte index. Ecological Modelling, 130(1–3):145–149, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002076.

Seppelt:2009:CSC

[SMSV09] Ralf Seppelt, Felix Müller, Boris Schröder, and Martin Volk. Challenges of simulating complex environmental systems at the landscape scale: a controversial dialogue between two cups of espresso. Ecological Modelling, 220(24):3481–3489, December 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009006176.

Sintes:2007:CGD

[SMTR07] Tomás Sintes, E. Moragues, A. Traveset, and J. Rita. Clonal growth dynamics of the invasive Carpobrotus affine acinaciformis in Mediterranean coastal systems: a non-linear model. Ecological Modelling, 206(1–2):110–118, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001615.

Shepherd:2002:SEW

[SMW02] A. Shepherd, S. M. McGinn, and G. C. L. Wyseure. Simulation of the effect of water shortage on the yields of winter wheat in North-East England. Ecological Modelling, 147(1):41–52, January 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004057.

Simas:2001:EGC

[SNF01] T. Simas, J. P. Nunes, and J. G. Ferreira. Effects of global climate change on coastal salt marshes. Ecological Modelling, 139(1):1–15, March 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002265.

Snover:2008:CUB

[Sno08] Melissa L. Snover. Comments on “Using Bayesian statespace modelling to assess the recovery and harvest potential of the Hawaiian green sea turtle stock”. Ecological Modelling, 212(3–4):545–549, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006023.

Smith:2006:LBB

[SNRW06] A. Smith, V. Nikora, A. Ross, and G. Wake. A lattice Boltzmann-based model of plankton–flow interaction around a mussel cluster. Ecological Modelling, 192 (3–4):645–657, February 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003273.

Sukhanov:2002:DFP

[SO02] V. V. Sukhanov and A. M. Omelko. Dynamics of feeding preferences by a predator. Ecological Modelling, 154 (3):203–206, September 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380002000078.

Sanchez:2004:EFC

[SO04] Francisco Sánchez and Ignacio Olaso. Effects of fisheries on the Cantabrian Sea shelf ecosystem. Ecological Modelling, 172(2–4):151–174, March 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003673.

Stahl:2006:RIA

[SO06] Justyn T. Stahl and Madan K. Oli. Relative importance of avian life-history variables to population growth rate. Ecological Modelling, 198(1–2):23–39, September 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006001669.

Seaquist:2003:RSB

[SOA03] J. W. Seaquist, L. Olsson, and J. Ardö. A remote sensingbased primary production model for grassland biomes. Ecological Modelling, 169(1):131–155, November 1, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002679.

Sasai:2007:CTC

[SOHY07] T. Sasai, K. Okamoto, T. Hiyama, and Y. Yamaguchi. Comparing terrestrial carbon fluxes from the scale of a flux tower to the global scale. Ecological Modelling, 208(2–4):135–144, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002979.

Shindo:2003:MBE

[SOK03] Junko Shindo, Katsuo Okamoto, and Hiroyuki Kawashima. A model-based estimation of nitrogen flow in the food production–supply system and its environmental effects in East Asia. Ecological Modelling, 169(1):197–212, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003002709.

Shindo:2006:PEE

[SOK06] Junko Shindo, Katsuo Okamoto, and Hiroyuki Kawashima. Prediction of the environmental effects of excess nitrogen caused by increasing food demand with rapid economic growth in eastern Asian countries, 1961–2020. Ecological Modelling, 193(3–4):703–720, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004801.

Schild:2001:CBP

[SP01] R. Schild and D. Prochnow. Coupling of biomass production and sedimentation of suspended sediments in eutrophic rivers. Ecological Modelling, 145(2–3):263–274, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003957.

Stockwell:2002:ESS

[SP02] David R. B. Stockwell and A. Townsend Peterson. Effects of sample size on accuracy of species distribution models. Ecological Modelling, 148(1):1–13, February 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100388X.

Sarkar:2006:ESM

[SPB^{+}06] R. R. Sarkar, S. V. Petrovskii, M. Biswas, A. Gupta, and J. Chattopadhyay. An ecological study of a marine plankton community based on the field data collected from Bay of Bengal. Ecological Modelling, 193(3–4):589–601, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004266.

Solidoro:2005:NPD

[SPC05] Cosimo Solidoro, Roberto Pastres, and Gianpiero Cossarini. Nitrogen and plankton dynamics in the lagoon of Venice. Ecological Modelling, 184(1):103–123, May 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005873.

Sagnard:2007:MSD

[SPD^{+}07] Fabrice Sagnard, Christian Pichot, Philippe Dreyfus, Pedro Jordano, and Bruno Fady. Modelling seed dispersal to predict seedling recruitment: Recolonization dynamics in a plantation forest. Ecological Modelling, 203(3–4):464–474, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000600634X.

Susani:2006:CBT

[SPJB06] Ludovico Susani, Federico M. Pulselli, Sven E. Jørgensen, and Simone Bastianoni. Comparison between technological and ecological exergy. Ecological Modelling, 193 (3–4):447–456, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004011.

Schramski:2009:RTT

[SPK^{+}09] J. R. Schramski, B. C. Patten, C. Kazanci, D. K. Gattie, and N. N. Kellam. The Reynolds transport theorem: Application to ecological compartment modeling and case study of ecosystem energetics. Ecological Modelling, 220(22):3225–3232, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005699.

Shan:2006:MLP

[SPM06] Y. Shan, D. Paull, and R. I. McKay. Machine learning of poorly predictable ecological data. Ecological Modelling, 195(1–2):129–138, May 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005843.

Sun:2008:SCE

[SPM^{+}08] Jianfeng Sun, Changhui Peng, Harry McCaughey, Xiaolu Zhou, Valerie Thomas, Frank Berninger, Benoît. St-Onge, and Dong Hua. Simulating carbon exchange of Canadian boreal forests: II. Comparing the carbon budgets of a boreal mixedwood stand to a black spruce forest stand. Ecological Modelling, 219(3–4):276–286, December 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003591.

Sage:2003:IMM

[SPS03] Richard W. Sage, Bernard C. Patten, and Paulette A. Salmon. Institutionalized model-making and ecosystembased management of exploited resource populations: a comparison with instrument flight. Ecological Modelling, 170(2–3):107–128, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002205.

Salvador:2001:GSA

[SPTP01] R. Salvador, J. Piñol, S. Tarantola, and E. Pla. Global sensitivity analysis and scale effects of a fire propagation model used over Mediterranean shrublands. Ecological Modelling, 136(2–3):175–189, January 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004191.

Scheibe:2002:FIB

[SR02] Timothy D. Scheibe and Marshall C. Richmond. Fish individual-based numerical simulator (FINS): a particlebased model of juvenile salmonid movement and dissolved gas exposure history in the Columbia River basin. Ecological Modelling, 147(3):233–252, January 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004240.

Sable:2008:CIB

[SR08] Shaye E. Sable and Kenneth A. Rose. A comparison of individual-based and matrix projection models for simulating yellow perch population dynamics in Oneida Lake, New York, USA. Ecological Modelling, 215(1–3):105–121, July 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000914.

Segarra:2005:GTS

[SRA05] José Segarra, José Raventós, and Miguel F. Acevedo. Growth of tropical savanna grass plants in competition: a shoot population model. Ecological Modelling, 189(3– 4):270–288, December 10, 2005. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002085.

Sexton:2006:HAQ

[SRB06] Joseph O. Sexton, R. Douglas Ramsey, and Dale L. Bartos. Habitone analysis of quaking aspen in the Utah book cliffs: Effects of site water demand and conifer cover. Ecological Modelling, 198(3–4):301–311, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002092.

Sprugel:2009:SEM

[SRG^{+}09] Douglas G. Sprugel, Katherine G. Rascher, Rolf Gersonde, Martin Dovčiak, James A. Lutz, and Charles B. Halpern. Spatially explicit modeling of overstory manipulations in young forests: Effects on stand structure and light. Ecological Modelling, 220(24):3565–3575, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005328.

Srinivasu:2004:RSE

[Sri04] P. D. N. Srinivasu. Regime shifts in eutrophied lakes: a mathematical study. Ecological Modelling, 179(1): 115–130, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002650.

Steele:2006:MAD

[SRK06] Brian M. Steele, Swarna K. Reddy, and Robert E. Keane. A methodology for assessing departure of current plant communities from historical conditions over large landscapes. Ecological Modelling, 199(1):53–63, November 1, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006002754.

Sekimura:2000:EPD

[SRL^{+}00] Toshio Sekimura, Tiina Roose, Bo Li, Philip K. Maini, Jun ichirou Suzuki, and Toshihiko Hara. The effect of population density on shoot morphology of herbs in relation to light capture by leaves. Ecological Modelling, 128(1):51–62, April 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380099002264.

Steele:2005:RSA

[SRN05] B. M. Steele, S. K. Reddy, and R. R. Nemani. A regression strategy for analyzing environmental data generated by spatio-temporal processes. Ecological Modelling, 181(2–3):93–108, January 20, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003539.

Sundermeyer:2006:AEC

[SRR06] Miles A. Sundermeyer, Brian J. Rothschild, and Allan R. Robinson. Assessment of environmental correlates with the distribution of fish stocks using a spatially explicit model. Ecological Modelling, 197(1–2):116–132, August 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006001347.

Singh:2004:RHS

[SRRS04] Brajendra K. Singh, J. Subba Rao, R. Ramaswamy, and Somdatta Sinha. The role of heterogeneity on the spatiotemporal dynamics of host–parasite metapopulation. Ecological Modelling, 180(2–3):435–443, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002947.

Shirley:2003:ISD

[SRS^{+}03] Mark D. F. Shirley, Steve P. Rushton, Graham C. Smith, Andrew B. South, and Peter W. W. Lurz. Investigating the spatial dynamics of bovine tuberculosis in badger populations: evaluating an individual-based simulation model. Ecological Modelling, 167(1–2):139–157, September 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001674.

Sahoo:2005:PPG

[SRW05] G. B. Sahoo, C. Ray, and H. F. Wade. Pesticide prediction in ground water in North Carolina domestic wells using artificial neural networks. Ecological Modelling, 183(1):29–46, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004004296.

Steinborn:2000:EIS

[SS00] Wolf Steinborn and Yuri Svirezhev. Entropy as an indicator of sustainability in agro-ecosystems: North Germany case study. Ecological Modelling, 133(3):247–257, September 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003239.

Svirezhev:2001:ESR

[SS01] Yuri M. Svirezhev and Wolf H. Steinborn. Exergy of solar radiation: information approach. Ecological Modelling, 145(2–3):101–110, November 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004094.

Schroder:2006:APP

[SS06a] B. Schröder and R. Seppelt. Analysis of pattern–process interactions based on landscape models — overview, general concepts, and methodological issues. Ecological Modelling, 199(4):505–516, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002936.

Seppelt:2006:PPD

[SS06b] Ralf Seppelt and Boris Schröder. Pattern and processes of dynamic mosaic landscapes. Ecological Modelling, 199(4):377–378, December 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002821.

Schizas:2007:WER

[SS07] Dimitrios Schizas and George Stamou. What ecosystems really are — physicochemical or biological entities? Ecological Modelling, 200(1–2):178–182, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003449.

Sellami:2008:MHM

[SS08a] Mohamed Habib Sellami and Mohamed Salah Sifaoui. Modelling of heat and mass transfer inside a traditional oasis: Experimental validation. Ecological Modelling, 210 (1–2):144–154, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007003699.

Sohl:2008:UFS

[SS08b] Terry Sohl and Kristi Sayler. Using the FORE-SCE model to project land-cover change in the southeastern United States. Ecological Modelling, 219(1–2):49–65, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003876.

Shaman:2006:HDM

[SSCS06] Jeffrey Shaman, Marc Spiegelman, Mark Cane, and Marc Stieglitz. A hydrologically driven model of swamp water mosquito population dynamics. Ecological Modelling, 194(4):395–404, April 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005005399.

Suarez-Seoane:2008:MEN

[SSdlMP^{+}08] Susana Suárez-Seoane, Eladio L. García de la Morena, Manuel B. Morales Prieto, Patrick E. Osborne, and Eduardo de Juana. Maximum entropy niche-based modelling of seasonal changes in little bustard (Tetrax tetrax) distribution. Ecological Modelling, 219(1–2):17–29, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003633.

Savenkoff:2007:EFP

[SSH^{+}07] C. Savenkoff, D. P. Swain, J. M. Hanson, M. Castonguay, M. O. Hammill, H. Bourdages, L. Morissette, and D. Chabot. Effects of fishing and predation in a heavily exploited ecosystem: Comparing periods before and after the collapse of groundfish in the southern Gulf of St. Lawrence (Canada). Ecological Modelling, 204(1–2):115–128, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006006685.

Svirezhev:2008:MLT

[SSH08] Yu. M. Svirezhev and A. Svirejeva-Hopkins. The model of long-term evolution of the carbon cycle. Ecological Modelling, 216(2):114–126, August 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001300.

Sales:2007:ISD

[SSK^{+}07] Marcio H. Sales, Carlos M. Souza, Phaedon C. Kyriakidis, Dar A. Roberts, and Edson Vidal. Improving spatial distribution estimation of forest biomass with geostatistics: a case study for Rondônia, Brazil. Ecological Modelling, 205(1–2):221–230, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000932.

Sohma:2008:BPC

[SSKN08] Akio Sohma, Yasuyuki Sekiguchi, Tomohiro Kuwae, and Yoshiyuki Nakamura. A benthic–pelagic coupled ecosystem model to estimate the hypoxic estuary including tidal flat — model description and validation of seasonal/daily dynamics. Ecological Modelling, 215(1–3):10–39, July 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008000963.

Sturtevant:2009:SDM

[SSM^{+}09] Brian R. Sturtevant, Robert M. Scheller, Brian R. Miranda, Douglas Shinneman, and Alexandra Syphard. Simulating dynamic and mixed-severity fire regimes: a processbased fire extension for LANDIS-II. Ecological Modelling, 220(23):3380–3393, December 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900533X.

Shvidenko:2007:SEM

[SSNB07] Anatoly Shvidenko, Dmitry Schepaschenko, Sten Nilsson, and Yuri Bouloui. Semi-empirical models for assessing biological productivity of Northern Eurasian forests. Ecological Modelling, 204(1–2):163–179, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700004X.

Svoray:2008:STE

[SSNHP08] Tal Svoray, Rakefet Shafran-Nathan, Zalmen Henkin, and Avi Perevolotsky. Spatially and temporally explicit modeling of conditions for primary production of annuals in dry environments. Ecological Modelling, 218(3– 4):339–353, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008003827.

Svirezhev:2003:ESR

[SSP03] Yuri M. Svirezhev, Wolf H. Steinborn, and Valeri L. Pomaz. Exergy of solar radiation: global scale. Ecological Modelling, 169(2–3):339–346, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002801.

Silva-Santos:2008:TSD

[SSPL^{+}08] Pedro Silva-Santos, Miguel Ângelo Pardal, Ricardo Jorge Lopes, Tiago Múrias, and João Alexandre Cabral. Testing the Stochastic Dynamic Methodology (StDM) as a management tool in a shallow temperate estuary of south Europe (Mondego, Portugal). Ecological Modelling, 210 (4):377–402, February 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004565.

Soboleva:2003:QTS

[SSPR03] T. K. Soboleva, P. R. Shorten, A. B. Pleasants, and A. L. Rae. Qualitative theory of the spread of a new gene into a resident population. Ecological Modelling, 163(1–2):33–44, May 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002003575.

Sabbadin:2007:DRS

[SSR07] Régis Sabbadin, Danny Spring, and Charles-Elie Rabier. Dynamic reserve site selection under contagion risk of deforestation. Ecological Modelling, 201(1):75–81, February 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003577.

Suzuki:2000:MSD

[SSS00] Motoyuki Suzuki, Masaki Sagehashi, and Akiyoshi Sakoda. Modelling the structural dynamics of a shallow and eutrophic water ecosystem based on mesocosm observations. Ecological Modelling, 128(2–3):221–243, April 20, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099002318.

Sadykova:2009:MCP

[SSS^{+}09] Dinara Sadykova, Jostein Skurdal, Alexander Sadykov, Trond Taugbol, and Dag O. Hessen. Modelling crayfish population dynamics using catch data: a size-structured model. Ecological Modelling, 220(20):2727–2733, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009004839.

Stange:2007:NAC

[Sta07] Claus Florian Stange. A novel approach to combine response functions in ecological process modelling. Ecological Modelling, 204(3–4):547–552, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000233.

Sulkava:2006:SRA

[STH06] Mika Sulkava, Jarkko Tikka, and Jaakko Hollmén. Sparse regression for analyzing the development of foliar nutrient concentrations in coniferous trees. Ecological Modelling, 191(1):118–130, January 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003807.

Stirn:2006:IFA

[Sti06] Lidija Zadnik Stirn. Integrating the fuzzy analytic hierarchy process with dynamic programming approach for determining the optimal forest management decisions. Ecological Modelling, 194(1–3):296–305, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005005272.

Stillman:2008:MIB

[Sti08] Richard A. Stillman. MORPH — an individual-based model to predict the effect of environmental change on foraging animal populations. Ecological Modelling, 216 (3–4):265–276, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008002020.

Shiyomi:2000:RPB

[STK00] Masae Shiyomi, Shigeo Takahashi, and Hiromitsu Kirita. Roles of plant biomass and vegetational heterogeneity, and energy-matter cycling in grassland sustainability. Ecological Modelling, 132(1–2):135–149, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003112.

Stockwell:2006:IEN

[Sto06] David R. B. Stockwell. Improving ecological niche models by data mining large environmental datasets for surrogate models. Ecological Modelling, 192(1–2):188–196, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002905.

Straskraba:2001:ENC

[Str01a] Milan Straškraba. Erratum to “Natural control mechanisms in models of aquatic ecosystems”: [Ecol. Model. 140/ 3 (2001) 195–206]. Ecological Modelling, 142(1–2):187, August 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003696.

Straskraba:2001:NCM

[Str01b] Milan Straškraba. Natural control mechanisms in models of aquatic ecosystems. Ecological Modelling, 140(3):195–205, June 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003386.

Strigul:2009:CIE

[Str09] Nikolay Strigul. Can imitation explain dialect origins? Ecological Modelling, 220(20):2624–2639, October 24, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009004864.

Schultz:2008:EBD

[SU08] Patrick Schultz and Noel R. Urban. Effects of bacterial dynamics on organic matter decomposition and nutrient release from sediments: a modeling study. Ecological Modelling, 210(1–2):1–14, January 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003523.

Suh:2005:TME

[Suh05] Sangwon Suh. Theory of materials and energy flow analysis in ecology and economics. Ecological Modelling, 189(3–4):251–269, December 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002139.

Suh:2006:ETM

[Suh06] Sangwon Suh. Erratum to “Theory of materials and energy flow analysis in ecology and economics” [Ecol. Modell. 189 (3–4) (2005) 251–269]. Ecological Modelling, 195(3–4):410, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006000408.

Seppelt:2002:OML

[SV02] Ralf Seppelt and Alexey Voinov. Optimization methodology for land use patterns using spatially explicit landscape models. Ecological Modelling, 151(2–3):125–142, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004550.

Seppelt:2003:OML

[SV03a] Ralf Seppelt and Alexey Voinov. Optimization methodology for land use patterns — evaluation based on multiscale habitat pattern comparison. Ecological Modelling, 168(3):217–231, October 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003001388.

Steinhardt:2003:MSL

[SV03b] Uta Steinhardt and Martin Volk. Meso-scale landscape analysis based on landscape balance investigations: problems and hierarchical approaches for their resolution. Ecological Modelling, 168(3):251–265, October 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001406.

Strand:2009:SEM

[SVB09] Eva K. Strand, Lee A. Vierling, and Stephen C. Bunting. A spatially explicit model to predict future landscape composition of aspen woodlands under various management scenarios. Ecological Modelling, 220(2):175–191, January 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008004419.

Sun:2006:MBC

[SvdWB^{+}06] Xiulian Sun, Wopke van der Werf, Felix J. J. A. Bianchi, Zhihong Hu, and Just M. Vlak. Modelling biological control with wild-type and genetically modified baculoviruses in the Helicoverpa armigera–cotton system. Ecological Modelling, 198(3–4):387–398, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002407.

Svirezhev:2000:LVM

[Svi00a] Yuri Svirezhev. Lotka–Volterra models and the global vegetation pattern. Ecological Modelling, 135(2–3):135– 146, December 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380000003550.

Svirezhev:2000:TE

[Svi00b] Yuri M. Svirezhev. Thermodynamics and ecology. Ecological Modelling, 132(1–2):11–22, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000301X.

Svirezhev:2002:SSD

[Svi02] Yuri M. Svirezhev. Simple spatially distributed model of the global carbon cycle and its dynamic properties. Ecological Modelling, 155(1):53–69, September 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000662.

Svirezhev:2004:ETI

[Svi04] Yuri M. Svirezhev. Ecosystem as a text: information analysis of the global vegetation pattern. Ecological Modelling, 174(1–2):19–24, May 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005659.

Svirezhev:2008:NME

[Svi08] Yuri M. Svirezhev. Nonlinearities in mathematical ecology: Phenomena and models: Would we live in Volterra’s world?

Ecological Modelling, 216(2):89–101, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001270.

Stigter:2004:OHT

[SvL04] J. D. Stigter and F. van Langevelde. Optimal harvesting in a two-species model under critical depensation: the case of optimal harvesting in semi-arid grazing systems. Ecological Modelling, 179(2):153–161, November 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002534.

Semovski:2004:SEN

[SVS04] S. V. Semovski, E. Verheyen, and D. Y. Sherbakov. Simulating the evolution of neutrally evolving sequences in a population under environmental changes. Ecological Modelling, 176(1–2):99–107, August 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005593.

Stoyan:2001:EFD

[SW01] Dietrich Stoyan and Sven Wagner. Estimating the fruit dispersion of anemochorous forest trees. Ecological Modelling, 145(1):35–47, November 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003854.

Song:2003:RFE

[SW03] Conghe Song and Curtis E. Woodcock. A regional forest ecosystem carbon budget model: impacts of forest age structure and landuse history. Ecological Modelling, 164(1):33–47, June 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003000139.

Swanson:2006:CCC

[Swa06] Richard E. Swanson. Comments on “Climate change: detection and attribution of trends from long-term geologic data” by C. Loehle [Ecological Modelling 171 (4) (2004)

433–450]. Ecological Modelling, 192(1–2):314–316, February 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002899.

Sherratt:2006:WFR

[SWB06] Thomas N. Sherratt, David M. Wilkinson, and Roderick S. Bain. Why fruits rot, seeds mold and meat spoils: a reappraisal. Ecological Modelling, 192(3–4):618–626, February 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003753.

Swain:2008:UHF

[SWBH08] D. L. Swain, T. Wark, and G. J. Bishop-Hurley. Using high fix rate GPS data to determine the relationships between fix rate, prediction errors and patch selection. Ecological Modelling, 212(3–4):273–279, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005583.

Song:2001:SOC

[SWBJ01] Weiguo Song, Fan Weicheng, Wang Binghong, and Zhou Jianjun. Self-organized criticality of forest fire in China. Ecological Modelling, 145(1):61–68, November 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003830.

Smith:2007:SMR

[SWCO07] J. M. D. Smith, J. P. Ward, L. E. Child, and M. R. Owen. A simulation model of rhizome networks for Fallopia japonica (Japanese knotweed) in the United Kingdom. Ecological Modelling, 200(3–4):421–432, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003863.

Shen:2005:SDP

[SWK^{+}05] Weijun Shen, Jianguo Wu, Paul R. Kemp, James F. Reynolds, and Nancy B. Grimm. Simulating the dynamics of primary productivity of a Sonoran ecosystem:

Model parameterization and validation. Ecological Modelling, 189(1–2):1–24, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001936.

Song:2006:TTP

[SWSF06] Weiguo Song, Jian Wang, Kohyu Satoh, and Weicheng Fan. Three types of power-law distribution of forest fires in Japan. Ecological Modelling, 196(3–4):527–532, July 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006000962.

Smith:2007:QCP

[SY07] S. Lan Smith and Yasuhiro Yamanaka. Quantitative comparison of photoacclimation models for marine phytoplankton. Ecological Modelling, 201(3–4):547–552, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004431.

Stapleton:2004:MME

[SYC04] Lee M. Stapleton, Scott D. Young, and Neil M. J. Crout. Have missing markets for ecological goods and services affected modelling of terrestrial C and N fluxes? Ecological Modelling, 179(4):569–574, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002819.

Su:2009:UEH

[SYCU09] M. R. Su, Z. F. Yang, B. Chen, and S. Ulgiati. Urban ecosystem health assessment based on emergy and set pair analysis — a comparative study of typical Chinese cities. Ecological Modelling, 220(18):2341–2348, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003974.

Svirezhev:2003:FGG

[SZ03] Yuri Svirezhev and Nicolai Zavalishin. “Forest–grass” global vegetation model with forest age structure. Ecological Modelling, 160(1–2):1–12, February 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002120.

Szilagyi:2000:CVI

[Szi00] Jozsef Szilagyi. Can a vegetation index derived from remote sensing be indicative of areal transpiration? Ecological Modelling, 127(1):65–79, March 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002008.

Su:2004:DMA

[SZL^{+}04] Fenzhen Su, Chenghu Zhou, Vincent Lyne, Yunyan Du, and Wenzhong Shi. A data-mining approach to determine the spatio-temporal relationship between environmental factors and fish distribution. Ecological Modelling, 174(4):421–431, June 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003004666.

Tucker:2005:USV

[TA05] Brian C. Tucker and Madhur Anand. On the use of stationary versus hidden Markov models to detect simple versus complex ecological dynamics. Ecological Modelling, 185(2–4):177–193, July 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004006027.

Telesca:2005:TSP

[TAL^{+}05] Luciano Telesca, Giuseppe Amatulli, Rosa Lasaponara, Michele Lovallo, and Adriano Santulli. Time-scaling properties in forest-fire sequences observed in Gargano area (southern Italy). Ecological Modelling, 185(2–4):531–544, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000177.

Tanaka:2002:MLM

[Tan02] Katsunori Tanaka. Multi-layer model of CO_{2} exchange in a plant community coupled with the water budget of leaf surfaces. Ecological Modelling, 147(1):85–104, January 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004136.

Tomenko:2007:MCW

[TAP07] Volodymyr Tomenko, Sirajuddin Ahmed, and Viktor Popov. Modelling constructed wetland treatment system performance. Ecological Modelling, 205(3–4):355–364, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001172.

Tarkhanov:2008:LSA

[Tar08] N. Tarkhanov. Lyapunov stability for an age-structured population model. Ecological Modelling, 216(2):232–239, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001464.

Tan:2005:CLS

[TB05] Can Ozan Tan and Meryem Beklioglu. Catastrophic-like shifts in shallow Turkish lakes: a modeling approach. Ecological Modelling, 183(4):425–434, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005010.

Tan:2006:MCN

[TB06a] Can Ozan Tan and Meryem Beklioglu. Modeling complex nonlinear responses of shallow lakes to fish and hydrology using artificial neural networks. Ecological Modelling, 196(1–2):183–194, July 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000561.

Tilley:2006:DEA

[TB06b] David Rogers Tilley and Mark T. Brown. Dynamic emergy accounting for assessing the environmental benefits of subtropical wetland stormwater management systems. Ecological Modelling, 192(3–4):327–361, February 25, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005003728.

Teriokhin:2008:IDD

[TB08] Anatoly T. Teriokhin and Elena V. Budilova. The impact of different density stresses on the dynamics of two competitive populations. Ecological Modelling, 212(1–2):5–9, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700511X.

Thiffault:2007:ISA

[TBP^{+}07] Evelyne Thiffault, Nicolas Bélanger, David Paré, William H. Hendershot, and Alison Munson. Investigating the soil acid–base status in managed boreal forests using the SAFE model. Ecological Modelling, 206(3–4):301–321, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001913.

Torres:2008:CFS

[TBPF08] I. López Torres, C. Fullana Belda, S. F. Ortuño Pérez, and A. J. Martín Fernández. Choosing Fagus sylvatica L. matrix model dimension by sensitivity analysis of the population growth rate with respect to the width of the diameter classes. Ecological Modelling, 218(3–4):307–314, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003487.

Tatarinov:2009:LTS

[TC09] Fyodor A. Tatarinov and Emil Cienciala. Long-term simulation of the effect of climate changes on the growth of main Central-European forest tree species. Ecological Modelling, 220(21):3081–3088, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000805.

Turner:2002:OMS

[TCD02] B. J. Turner, O. Chikumbo, and S. M. Davey. Optimisation modelling of sustainable forest management at the regional level: an Australian example. Ecological Modelling, 153(1–2):157–179, July 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005087.

Teel:2003:SBH

[TCGL03] P. D. Teel, M. S. Corson, W. E. Grant, and M. T. Longnecker. Simulating biophysical and human factors that affect detection probability of cattle-fever ticks (Boophilus spp.) in semi-arid thornshrublands of South Texas. Ecological Modelling, 170(1):29–43, December 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002965.

Todorovski:2006:IKD

[TD06] Ljupčo Todorovski and Sašo Džeroski. Integrating knowledgedriven and data-driven approaches to modeling. Ecological Modelling, 194(1–3):3–13, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500493X.

Teh:2008:SMP

[TDdSLS^{+}08] Su Yean Teh, Donald L. DeAngelis, Leonel da Silveira Lobo Sternberg, Fernando R. Miralles-Wilhelm, Thomas J. Smith, and Hock-Lye Koh. A simulation model for projecting changes in salinity concentrations and species dominance in the coastal margin habitats of the Everglades. Ecological Modelling, 213(2):245–256, May 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006333.

Tomlinson:2007:FSM

[TDHO07] Kyle W. Tomlinson, James G. Dominy, John W. Hearne, and Timothy G. O’Connor. A functional-structural model for growth of clonal bunchgrasses. Ecological Modelling, 202(3–4):243–264, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006005370.

Tichit:2007:CVM

[TDL^{+}07] M. Tichit, L. Doyen, J. Y. Lemel, O. Renault, and D. Durant. A co-viability model of grazing and bird community management in farmland. Ecological Modelling, 206(3–4):277–293, August 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700186X.

Todorovski:2003:UED

[TDLP03] Ljupčo Todorovski, Sašo Džeroski, Pat Langley, and Christopher Potter. Using equation discovery to revise an Earth ecosystem model of the carbon net production. Ecological Modelling, 170(2–3):141–154, December 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002229.

Tews:2006:LPM

[TEMJ06] Jörg Tews, Alexandra Esther, Sue J. Milton, and Florian Jeltsch. Linking a population model with an ecosystem model: Assessing the impact of land use and climate change on savanna shrub cover dynamics. Ecological Modelling, 195(3–4):219–228, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006113.

Thornley:2005:OEL

[TF05] John H. M. Thornley and James France. An openended logistic-based growth function. Ecological Modelling, 184(2–4):257–261, June 1, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005514.

Turley:2009:DCU

[TF09] Marianne C. Turley and E. David Ford. Definition and calculation of uncertainty in ecological process models. Ecological Modelling, 220(17):1968–1983, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003159.

Tews:2007:PNE

[TFF07] Joerg Tews, Michael A. D. Ferguson, and Lenore Fahrig. Potential net effects of climate change on High Arctic Peary caribou: Lessons from a spatially explicit simulation model. Ecological Modelling, 207(2–4):85–98, October 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007002359.

Touratier:2001:SMR

[TFM01] Franck Touratier, John G. Field, and Coleen L. Moloney. A stoichiometric model relating growth substrate quality (C:N:P ratios) to N:P ratios in the products of heterotrophic release and excretion. Ecological Modelling, 139(2–3):265–291, April 30, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100237X.

Tsubo:2007:WBM

[TFTO07] M. Tsubo, S. Fukai, T. P. Tuong, and M. Ouk. A water balance model for rainfed lowland rice fields emphasising lateral water movement within a toposequence. Ecological Modelling, 204(3–4):503–515, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000543.

Tanaka:2006:PCE

[TH06a] Katsunori Tanaka and Shoji Hashimoto. Plant canopy effects on soil thermal and hydrological properties and soil respiration. Ecological Modelling, 196(1–2):32–44, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000536.

Tietjen:2006:MDM

[TH06b] Britta Tietjen and Andreas Huth. Modelling dynamics of managed tropical rainforests — an aggregated approach. Ecological Modelling, 199(4):421–432, December 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006002869.

Tang:2008:BIS

[TH08] Sanyi Tang and Elizabeth A. Heron. Bayesian inference for a stochastic logistic model with switching points. Ecological Modelling, 219(1–2):153–169, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004092.

Tomlinson:2002:AEE

[THA02] K. W. Tomlinson, J. W. Hearne, and R. R. Alexander. An approach to evaluate the effect of property size on land-use options in semi-arid rangelands. Ecological Modelling, 149(1–2):85–95, March 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005166.

Topping:2003:AAB

[THJ^{+}03] Chris J. Topping, Tine S. Hansen, Thomas S. Jensen, Jane U. Jepsen, Frank Nikolajsen, and Peter Odderskær. ALMaSS, an agent-based model for animals in temperate European landscapes. Ecological Modelling, 167 (1–2):65–82, September 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000300173X.

Tian:2006:ETS

[Tia06a] Rucheng C. Tian. Erratum to “Toward standard parameterizations in marine biological modeling” [Ecol. Modell. 193 (2006) 363–386]. Ecological Modelling, 199(1):128, November 1, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006003619.

Tian:2006:TSP

[Tia06b] Rucheng C. Tian. Toward standard parameterizations in marine biological modeling. Ecological Modelling, 193 (3–4):363–386, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004473.

Tiezzi:2005:TTT

[Tie05] E. Tiezzi. Towards a thermodynamic theory for ecological systems. Ecological Modelling, 188(2–4):492–493, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000500428X.

Todd:2001:SUS

[TIJ^{+}01] Charles R. Todd, Pablo Inchausti, Simone Jenkins, Mark A. Burgman, and Meei Pyng Ng. Structural uncertainty in stochastic population models: delayed development in the eastern barred bandicoot, Perameles gunnii. Ecological Modelling, 136(2–3):237–254, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004270.

Tilley:2004:HOC

[Til04] David Rogers Tilley. Howard T. Odum’s contribution to the laws of energy. Ecological Modelling, 178(1–2):121–125, October 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003005519.

Tisdell:2006:KAS

[Tis06] Clem Tisdell. Knowledge about a species’ conservation status and funding for its preservation: Analysis. Ecological Modelling, 198(3–4):515–519, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002080.

Tinker:2001:TSD

[TK01] Daniel B. Tinker and Dennis H. Knight. Temporal and spatial dynamics of coarse woody debris in harvested and unharvested lodgepole pine forests. Ecological Modelling, 141(1–3):125–149, July 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002691.

Tollner:2007:DET

[TK07] Ernest W. Tollner and Caner Kazanci. Defining an ecological thermodynamics using discrete simulation approaches. Ecological Modelling, 208(1):68–79, October 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002530.

Terui:2008:PDM

[TK08] Takeshi Terui and Michio J. Kishi. Population dynamics model of Copepoda (Neocalanus cristatus) in the northwestern subarctic Pacific. Ecological Modelling, 215(1–3):77–88, July 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008000902.

Tobor-Kaplon:2007:EII

[TKHS^{+}07] Maria A. Tobor-Kapłon, Remko Holtkamp, Ursula M. Scharler, Agnieszka Doroszuk, Frans J. A. Kuenen, Jaap Bloem, and Peter C. de Ruiter. Evaluation of information indices as indicators of environmental stress in terrestrial soils. Ecological Modelling, 208(1):80–90, October 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002542.

Troost:2007:BAE

[TKK07] T. A. Troost, B. W. Kooi, and S. A. L. M. Kooijman. Bifurcation analysis of ecological and evolutionary processes in ecosystems. Ecological Modelling, 204(1–2):253–268, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700018X.

Tollner:2009:CSA

[TKSP09] E. W. Tollner, C. Kazanci, J. R. Schramski, and B. C. Patten. Control system approaches to ecological systems analysis: Invariants and frequency response. Ecological Modelling, 220(22):3233–3240, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005304.

Tyre:2007:IMM

[TKTB07] Andrew Tyre, Gregory D. Kerr, Brigitte Tenhumberg, and C. Michael Bull. Identifying mechanistic models of spatial behaviour using pattern-based modelling: an example from lizard home ranges. Ecological Modelling, 208(2–4):307–316, November 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003109.

Tuchkovenko:2003:MMO

[TL03] Yuri S. Tuchkovenko and Serguei A. Lonin. Mathematical model of the oxygen regime of Cartagena Bay. Ecological Modelling, 165(1):91–106, July 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000644.

Tang:2001:SEE

[TLW01] Shouzheng Tang, Yong Li, and Yonghe Wang. Simultaneous equations, error-in-variable models, and model integration in systems ecology. Ecological Modelling, 142(3):285–294, August 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000100326X.

Turechek:2001:ESP

[TM01] William W. Turechek and Laurence V. Madden. Effect of scale on plant disease incidence and heterogeneity in a spatial hierarchy. Ecological Modelling, 144(1):77–95, October 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003507.

Tanaka:2005:EBS

[TM05a] Katsumasa Tanaka and Fred T. Mackenzie. Ecosystem behavior of southern Kaneohe Bay, Hawaii: a statistical and modelling approach. Ecological Modelling, 188(2–4):296–326, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500164X.

Taulis:2005:EWW

[TM05b] M. E. Taulis and M. W. Milke. Estimation of WGEN weather generation parameters in arid climates. Ecological Modelling, 184(2–4):177–191, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005459.

Tixier:2004:SPC

[TMD04] P. Tixier, E. Malezieux, and M. Dorel. SIMBA-POP: a cohort population model for long-term simulation of banana crop harvest. Ecological Modelling, 180(2–3): 407–417, December 25, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002911.

Tracol:2006:TSG

[TMHJ06] Y. Tracol, E. Mougin, P. Hiernaux, and L. Jarlan. Testing a Sahelian grassland functioning model against herbage mass measurements. Ecological Modelling, 193 (3–4):437–446, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004023.

Tews:2004:MSD

[TMJ04] Jörg Tews, Kirk Moloney, and Florian Jeltsch. Modeling seed dispersal in a variable environment: a case study of the fleshy-fruited savanna shrub Grewia flava. Ecological Modelling, 175(1):65–76, June 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300468X.

Thapanand:2007:TPF

[TMJ^{+}07] Thanitha Thapanand, Jacques Moreau, Tuantong Jutagate, Prachit Wongrat, Teera Lekchonlayut, Charumas Meksumpun, Sommai Janekitkarn, Arunee Rodloi, Varunthat Dulyapruk, and Ladda Wongrat. Towards possible fishery management strategies in a newly impounded man-made lake in Thailand. Ecological Modelling, 204(1–2):143–155, May 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007000026.

Tiemeyer:2007:MDS

[TMLV07] Bärbel Tiemeyer, Roger Moussa, Bernd Lennartz, and Marc Voltz. MHYDAS-DRAIN: a spatially distributed model for small, artificially drained lowland catchments. Ecological Modelling, 209(1):2–20, November 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003559.

Tremblay:2005:ADT

[TMM05] M. Tremblay, C. Messier, and D. J. Marceau. Analysis of deciduous tree species dynamics after a severe ice storm using SORTIE model simulations. Ecological Modelling, 187(2–3):297–313, September 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000657.

Termansen:2006:UGA

[TMP06] Mette Termansen, Colin J. McClean, and Christopher D. Preston. The use of genetic algorithms and Bayesian classification to model species distributions. Ecological Modelling, 192(3–4):410–424, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500356X.

Trnka:2007:ECB

[TMS^{+}07] M. Trnka, F. Muška, D. Semerádová, M. Dubrovský, E. Kocmánková, and Z. Žalud. European corn borer life stage model: Regional estimates of pest development and spatial distribution under present and future climate. Ecological Modelling, 207(2–4):61–84, October 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007002414.

Todd:2001:GUC

[TN01] Charles R. Todd and Meei Pyng Ng. Generating unbiased correlated random survival rates for stochastic population models. Ecological Modelling, 144(1):1–11, October 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000100343X.

Tullos:2006:QMA

[TN06] Desiree D. Tullos and Michael Neumann. A qualitative model for analyzing the effects of anthropogenic activities in the watershed on benthic macroinvertebrate communities. Ecological Modelling, 196(1–2):209–220, July 10, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006000597.

Tsehaye:2008:EOF

[TN08] Iyob Tsehaye and Leopold A. J. Nagelkerke. Exploring optimal fishing scenarios for the multispecies artisanal fisheries of Eritrea using a trophic model. Ecological Modelling, 212(3–4):319–333, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005868.

Taguchi:2009:EBW

[TN09] Koichi Taguchi and Kisaburo Nakata. Evaluation of biological water purification functions of inland lakes using an aquatic ecosystem model. Ecological Modelling, 220(18):2255–2271, September 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003305.

Todd:2004:DDU

[TNK04] Charles R. Todd, Simon J. Nicol, and John D. Koehn. Density-dependence uncertainty in population models for the conservation management of trout cod, Maccullochella macquariensis. Ecological Modelling, 171(4):359–380, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003003545.

Tonnang:2009:AIB

[TNO^{+}09a] Henri E. Z. Tonnang, Lev V. Nedorezov, Horace Ochanda, John Owino, and Bernhard Löhr. Assessing the impact of biological control of Plutella xylostella through the application of Lotka–Volterra model. Ecological Modelling, 220(1):60–70, January 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004298.

Tonnang:2009:EDH

[TNO^{+}09b] Henri E. Z. Tonnang, Lev V. Nedorezov, John O. Owino, Horace Ochanda, and Bernhard Löhr. Evaluation of discrete host–parasitoid models for diamondback moth and Diadegma semiclausum field time population density series. Ecological Modelling, 220(15):1735–1744, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002646.

Tole:2006:CRS

[Tol06] Lise Tole. Choosing reserve sites probabilistically: a Colombian Amazon case study. Ecological Modelling, 194(4):344–356, April 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004904.

Tran:2009:DMI

[TOS09] Liem T. Tran, Robert V. O’Neill, and Elizabeth R. Smith. Determine the most influencing stressors and the most susceptible resources for environmental integrated assessment. Ecological Modelling, 220(18):2335–2340, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009003937.

Tonitto:2006:DST

[TP06] Christina Tonitto and Thomas M. Powell. Development of a spatial terrestrial nitrogen model for application to Douglas-fir forest ecosystems. Ecological Modelling, 193(3–4):340–362, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004138.

Tiezzi:2008:EAL

[TP08] Enzo Tiezzi and Riccardo Maria Pulselli. An entropic approach to living systems. Ecological Modelling, 216(2):229–231, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008001452.

Tison:2007:PDR

[TPC^{+}07] J. Tison, Y.-S. Park, M. Coste, J. G. Wasson, F. Rimet, L. Ector, and F. Delmas. Predicting diatom reference communities at the French hydrosystem scale: a first step towards the definition of the good ecological status. Ecological Modelling, 203(1–2):99–108, April 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005618.

Tournoud:2006:OQN

[TPCS06] Marie-George Tournoud, Sylvain Payraudeau, Flavie Cernesson, and Christian Salles. Origins and quantification of nitrogen inputs into a coastal lagoon: Application to the Thau lagoon (France). Ecological Modelling, 193(1–2):19–33, March 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005004540.

Texeira:2008:HDF

[TPJ08] Marcos Texeira, José M. Paruelo, and Esteban Jobbagy. How do forage availability and climate control sheep reproductive performance?: an analysis based on artificial neural networks and remotely sensed data. Ecological Modelling, 217(1–2):197–206, September 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003086.

Thebault:2003:CBT

[TR03] J. M. Thébault and S. Rabouille. Comparison between two mathematical formulations of the phytoplankton specific growth rate as a function of light and temperature, in two simulation models (Aster & Yoyo). Ecological Modelling, 163(1–2):145–151, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004040.

Thornton:2005:EMS

[TR05] Peter E. Thornton and Nan A. Rosenbloom. Ecosystem model spin-up: Estimating steady state conditions in a coupled terrestrial carbon and nitrogen cycle model. Ecological Modelling, 189(1–2):25–48, November 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001948.

Tixier:2006:MPD

[TRDM06] P. Tixier, J.-M. Risède, M. Dorel, and E. Malézieux. Modelling population dynamics of banana plant-parasitic nematodes: a contribution to the design of sustainable cropping systems. Ecological Modelling, 198(3–4):321–331, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002134.

Tan:2003:GEC

[TS03] Guoxin Tan and Ryosuke Shibasaki. Global estimation of crop productivity and the impacts of global warming by GIS and EPIC integration. Ecological Modelling, 168(3):357–370, October 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003001467.

Touloumis:2009:MAD

[TS09] K. Touloumis and G. P. Stamou. A metapopulation approach of the dynamics of arthropods from Mediterraneantype ecosystems. Ecological Modelling, 220(8):1105–1112, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009000817.

Tang:2009:EEP

[TSBH09] Guoping Tang, Sarah L. Shafer, Patrick J. Bartlein, and Justin O. Holman. Effects of experimental protocol on global vegetation model accuracy: a comparison of simulated and observed vegetation patterns for Asia. Ecological Modelling, 220(12):1481–1491, June 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001999.

Tschirhart:2002:RCA

[Tsc02] John Tschirhart. Resource competition among plants: from maximizing individuals to community structure. Ecological Modelling, 148(2):191–212, February 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004392.

Tschirhart:2004:NAS

[Tsc04] John Tschirhart. A new adaptive system approach to predator–prey modeling. Ecological Modelling, 176(3– 4):255–276, September 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000481.

Trancoso:2005:MMU

[TSF^{+}05] A. R. Trancoso, S. Saraiva, L. Fernandes, P. Pina, P. Leitão, and R. Neves. Modelling macroalgae using a 3D hydrodynamic-ecological model in a shallow, temperate estuary. Ecological Modelling, 187(2–3):232–246, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005000918.

Thornley:2007:OEL

[TSF07] John H. M. Thornley, John J. Shepherd, and J. France. An open-ended logistic-based growth function: Analytical solutions and the power-law logistic model. Ecological Modelling, 204(3–4):531–534, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006648.

Trolle:2008:WFD

[TSJ08] Dennis Trolle, Henrik Skovgaard, and Erik Jeppesen. The water framework directive: Setting the phosphorus loading target for a deep lake in Denmark using the 1D lake ecosystem model DYRESM–CAEDYM. Ecological Modelling, 219(1–2):138–152, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004080.

Travers:2009:TWC

[TSJ^{+}09] M. Travers, Y.-J. Shin, S. Jennings, E. Machu, J. A. Huggett, J. G. Field, and P. M. Cury. Two-way coupling versus one-way forcing of plankton and fish models to predict ecosystem changes in the Benguela. Ecological Modelling, 220(21):3089–3099, November 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009005766.

Tyutyunov:2002:RAH

[TSJA02] Yuri Tyutyunov, Inna Senina, Christian Jost, and Roger Arditi. Risk assessment of the harvested pike-perch population of the Azov Sea. Ecological Modelling, 149(3):297–311, April 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001004781.

Tollner:2009:INP

[TSKP09] E. W. Tollner, J. R. Schramski, C. Kazanci, and B. C. Patten. Implications of network particle tracking (NPT) for ecological model interpretation. Ecological Modelling, 220(16):1904–1912, August 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009002774.

Torres-Sorando:2003:PDO

[TSZdRR03] Lourdes J. Torres-Sorando, Donis Zacarias, Evelyn Zoppi de Roa, and Diego J. Rodríguez. Population dynamics of Oithona hebes (Copepoda: Cyclopoida) in a coastal estuarine lagoon of Venezuela: a stage-dependent matrix growth model. Ecological Modelling, 161(3):159–168, March 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003551.

Tasser:2005:NMP

[TT05] E. Tasser and Ulrike Tappeiner. New model to predict rooting in diverse plant community compositions. Ecological Modelling, 185(2–4):195–211, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004006039.

Tappeiner:2001:GBM

[TTA^{+}01] Ulrike Tappeiner, Gottfried Tappeiner, Janette Aschenwald, Erich Tasser, and Bertram Ostendorf. GIS-based modelling of spatial pattern of snow cover duration in an alpine area. Ecological Modelling, 138(1–3):265–275, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000004075.

Thompson:2003:MBA

[TTA^{+}03] C. J. Thompson, B. J. P. Thompson, P. K. Ades, R. Cousens, P. Garnier-Gere, K. Landman, E. Newbigin, and M. A. Burgman. Model-based analysis of the likelihood of gene introgression from genetically modified crops into wild relatives. Ecological Modelling, 162(3):199–209, April 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003472.

Tyre:2006:ILS

[TTB06] Andrew J. Tyre, Brigitte Tenhumberg, and C. Michael Bull. Identifying landscape scale patterns from individual scale processes. Ecological Modelling, 199(4):442–450, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002882.

Thayer:2007:WEU

[TTHH07] Sarah A. Thayer, William W. Taylor, Daniel B. Hayes, and Robert C. Haas. Weight of evidence for underlying dynamics of yellow perch in Saginaw Bay, Lake Huron. Ecological Modelling, 206(1–2):31–40, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001457.

Tyson:2007:MDS

[TTJ07] Rebecca Tyson, Howard Thistlewood, and Gary J. R. Judd. Modelling dispersal of sterile male codling moths, Cydia pomonella, across orchard boundaries. Ecological Modelling, 205(1–2):1–12, July 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006673.

Tuomi:2009:LLD

[TTJ^{+}09] M. Tuomi, T. Thum, H. Järvinen, S. Fronzek, B. Berg, M. Harmon, J. A. Trofymow, S. Sevanto, and J. Liski. Leaf litter decomposition — estimates of global variability based on Yasso07 model. Ecological Modelling, 220(23):3362–3371, December 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000900386X.

Topalova:2009:MRM

[TTPS09] Yana Topalova, Yovana Todorova, Antoaneta Panova, and Irina Schneider. Modelling of the relationship moisture content to nutrient transformation rate in river sediments. Ecological Modelling, 220(23):3325–3330, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009005717.

Tudor:2001:ICS

[Tud01] Mladen Tudor. Inter-comparison of some equations for evaluating fish daily ration by numerical experiment with an impulse-input feeding model. Ecological Modelling, 136 (2–3):167–174, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000000418X.

Tuomi:2008:HSR

[TVK^{+}08] Mikko Tuomi, Pekka Vanhala, Kristiina Karhu, Hannu Fritze, and Jari Liski. Heterotrophic soil respiration — comparison of different models describing its temperature dependence. Ecological Modelling, 211(1–2):182–190, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700453X.

Thompson:2000:PMP

[TVS00] R. E. Thompson, E. O. Voit, and G. I. Scott. A probabilistic model for predicting distributions of PAH ratios between oysters and marine sediments. Ecological Modelling, 135(2–3):231–242, December 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003677.

Tuljapurkar:2000:ETS

[TW00] Shripad Tuljapurkar and Pamela Wiener. Escape in time: stay young or age gracefully? Ecological Modelling, 133(1–2):143–159, August 25, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000288X.

Tang:2002:PEP

[TW02] Shouzheng Tang and Yonghe Wang. A parameter estimation program for the error-in-variable model. Ecological Modelling, 156(2–3):225–236, November 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001734.

Timoshevskii:2003:NME

[TYK03] Andrei Timoshevskii, Vladimir Yeremin, and Sergey Kalkuta. New method for ecological monitoring based on the method of self-organising mathematical models. Ecological Modelling, 162(1–2):1–13, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004039.

Tung:2007:MST

[TYLL07] Ching-Pin Tung, Yi-Chen E. Yang, Tsung-Yu Lee, and Ming-Hsu Li. Modification of a stream temperature model with Beer’s law and application to GaoShan Creek in Taiwan. Ecological Modelling, 200(1–2):217–224, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006003644.

Toda:2009:FPI

[TYS^{+}09] Motomu Toda, Masayuki Yokozawa, Akihiro Sumida, Tsutomu Watanabe, and Toshihiko Hara. Foliage profiles of individual trees determine competition, self-thinning, biomass and NPP of a Cryptomeria japonica forest stand: a simulation study based on a stand-scale process-based forest model. Ecological Modelling, 220(18):2272–2280, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009003342.

Tanaka:2009:WBC

[TYT^{+}09] Katsunori Tanaka, Natsuko Yoshifuji, Nobuaki Tanaka, Katsushige Shiraki, Chatchai Tantasirin, and Masakazu Suzuki. Water budget and the consequent duration of canopy carbon gain in a teak plantation in a dry tropical region: Analysis using a soil–plant–air continuum multilayer model. Ecological Modelling, 220(12):1534–1543, June 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001975.

Tao:2005:RSC

[TYZ^{+}05] Fulu Tao, Masayuki Yokozawa, Zhao Zhang, Yinlong Xu, and Yousay Hayashi. Remote sensing of crop production in China by production efficiency models: models comparisons, estimates and uncertainties. Ecological Modelling, 183(4):385–396, May 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004983.

Tian:2005:MPD

[TYZC05] Yongzhong Tian, Tianxiang Yue, Lifen Zhu, and Nicholas Clinton. Modeling population density using land cover data. Ecological Modelling, 189(1–2):72–88, November 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001900.

Uchmanski:2000:RPA

[Uch00] Janusz Uchmański. Resource partitioning among competing individuals and population persistence: an individualbased model. Ecological Modelling, 131(1):21–32, June 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000002428.

Uchida:2007:SFW

[UDB07] Satoshi Uchida, Barbara Drossel, and Ulrich Brose. The structure of food webs with adaptive behaviour. Ecological Modelling, 206(3–4):263–276, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001822.

Ulanowicz:2006:EIA

[UJF06] Robert E. Ulanowicz, Sven Erik Jørgensen, and Brian D. Fath. Exergy, information and aggradation: an ecosystems reconciliation. Ecological Modelling, 198(3–4):520–524, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006002626.

Ulanowicz:2005:JJK

[Ula05] Robert E. Ulanowicz. James j. Kay (1955–2004). Ecological Modelling, 181(2–3):285–286, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003758.

Ulanowicz:2009:DNE

[Ula09] Robert E. Ulanowicz. The dual nature of ecosystem dynamics. Ecological Modelling, 220(16):1886–1892, August 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009002695.

Ulgiati:2004:HOC

[Ulg04] Sergio Ulgiati. H. T. Odum and E. C. Odum, the prosperous way down. Ecological Modelling, 178(1–2):247–250, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003005520.

Ulgiati:2006:OIS

[URB06] Sergio Ulgiati, Marco Raugei, and Silvia Bargigli. Overcoming the inadequacy of single-criterion approaches to life cycle assessment. Ecological Modelling, 190(3–4):432–442, January 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002589.

Ursino:2009:ABG

[Urs09] Nadia Ursino. Above and below ground biomass patterns in arid lands. Ecological Modelling, 220(11):1411–1418, June 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001549.

Ulanowicz:2008:LIM

[US08] Robert E. Ulanowicz and Ursula M. Scharler. Leastinference methods for constructing networks of trophic flows. Ecological Modelling, 210(3):278–286, January 24, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007004073.

Unnasch:2006:DTM

[USK^{+}06] Robert S. Unnasch, Tonya Sprenger, Charles R. Katholi, Eddie W. Cupp, Geoffrey E. Hill, and Thomas R. Unnasch. A dynamic transmission model of eastern equine encephalitis virus. Ecological Modelling, 192(3–4):425–440, February 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003583.

Uusitalo:2007:ACB

[Uus07] Laura Uusitalo. Advantages and challenges of Bayesian networks in environmental modelling. Ecological Modelling, 203(3–4):312–318, May 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006089.

Umgiesser:2001:HSM

[UZ01] G. Umgiesser and L. Zampato. Hydrodynamic and salinity modeling of the Venice channel network with coupled 1-D–2-D mathematical models. Ecological Modelling, 138(1–3):75–85, March 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000394X.

Uttieri:2005:FCT

[UZSM05] Marco Uttieri, Enrico Zambianchi, J. Rudi Strickler, and Maria Grazia Mazzocchi. Fractal characterization of threedimensional zooplankton swimming trajectories. Ecological Modelling, 185(1):51–63, June 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005952.

Verhoef:2000:SSD

[VA00] A. Verhoef and S. J. Allen. A SVAT scheme describing energy and CO_{2} fluxes for multi-component vegetation: calibration and test for a Sahelian savannah. Ecological Modelling, 127(2–3):245–267, March 30, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002136.

VanBroekhoven:2006:FRB

[VADV06] Ester Van Broekhoven, Veronique Adriaenssens, Bernard De Baets, and Piet F. M. Verdonschot. Fuzzy rule-based macroinvertebrate habitat suitability models for running waters. Ecological Modelling, 198(1–2):71–84, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000600175X.

Vandermeer:2004:WBQ

[Van04] John Vandermeer. Wada basins and qualitative unpredictability in ecological models: a graphical interpretation. Ecological Modelling, 176(1–2):65–74, August 15, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003005568.

Vandermeer:2008:NCP

[Van08a] John Vandermeer. The niche construction paradigm in ecological time. Ecological Modelling, 214(2–4):385–390, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001506.

Vanoverbeke:2008:MIP

[Van08b] Joost Vanoverbeke. Modeling individual and population dynamics in a consumer–resource system: Behavior under food limitation and crowding and the effect on population cycling in Daphnia. Ecological Modelling, 216(3–4):385–401, September 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002342.

Vayenas:2005:DTP

[VATP05] D. V. Vayenas, G. Aggelis, V. Tsagou, and S. Pavlou. Dynamics of a two-prey–one-predator system with predator switching regulated by a catabolic repression control-like mode. Ecological Modelling, 186(3):345–357, August 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000578.

Varley:2006:AMR

[VB06] Nathan Varley and Mark S. Boyce. Adaptive management for reintroductions: Updating a wolf recovery model for Yellowstone National Park. Ecological Modelling, 193(3–4):315–339, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004412.

vonBloh:2008:EBE

[vBBE^{+}08] Werner von Bloh, Christine Bounama, Klaus Eisenack, Brigitte Knopf, and Oliver Walkenhorst. Estimating the biogenic enhancement factor of weathering using an inverse viability method. Ecological Modelling, 216(2):245–251, August 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008001488.

VandenBrink:2006:UEM

[VBD06] Paul J. Van den Brink, Colin D. Brown, and Igor G. Dubus. Using the expert model PERPEST to translate measured and predicted pesticide exposure data into ecological risks. Ecological Modelling, 191(1):106–117, January 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005003790.

Vanhoutte-Brunier:2008:MKM

[VBFM^{+}08] Alice Vanhoutte-Brunier, Liam Fernand, Alain Ménesguen, Sandra Lyons, Francis Gohin, and Philippe Cugier. Modelling the Karenia mikimotoi bloom that occurred in the western English Channel during summer 2003. Ecological Modelling, 210(4):351–376, February 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004280.

Vospernik:2007:EVS

[VBRS07] Sonja Vospernik, Mike Bokalo, Friedrich Reimoser, and Hubert Sterba. Evaluation of a vegetation simulator for roe deer habitat predictions. Ecological Modelling, 202(3–4):265–280, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006005333.

Vega-Cendejas:2001:EFM

[VCAS01] M. E. Vega-Cendejas and F. Arreguín-Sánchez. Energy fluxes in a mangrove ecosystem from a coastal lagoon in Yucatan Peninsula, Mexico. Ecological Modelling, 137(2–3):119–133, February 20, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000421X.

Valandro:2003:WHB

[VCC03] L. Valandro, R. Caimmi, and L. Colombo. What is hidden behind the concept of ecosystem efficiency in energy transformation? Ecological Modelling, 170(2–3):185–191, December 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380003002254.

Vignati:2005:ETD

[VCD05] D. A. L. Vignati, M. Camusso, and J. Dominik. Estimation of the truly dissolved concentrations of Cd, Cu, Ni, and Zn in contrasting aquatic environments with a simple empirical model. Ecological Modelling, 184(1):125–139, May 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004005885.

Vitale:2007:RAP

[VCM07] Marcello Vitale, Francesca Capogna, and Fausto Manes. Resilience assessment on Phillyrea angustifolia L. maquis undergone to experimental fire through a big-leaf modelling approach. Ecological Modelling, 203(3–4):387–394, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006006272.

Vilenkin:2009:PST

[VCP09] Boris Vilenkin, Vladimir I. Chikatunov, and Tomáš Pavlíček. The pattern of species turnover resulting from stochastic population dynamics: the model and field data. Ecological Modelling, 220(5):657–661, March 10, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008005851.

Vincenzi:2006:GBH

[VCRD06] Simone Vincenzi, Graziano Caramori, Remigio Rossi, and Giulio A. De Leo. A GIS-based habitat suitability model for commercial yield estimation of Tapes philippinarum in a Mediterranean coastal lagoon (Sacca di Goro, Italy). Ecological Modelling, 193(1–2):90–104, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004588.

vandenBerg:2002:WUC

[vdBDR02] M. van den Berg, P. M. Driessen, and R. Rabbinge. Water uptake in crop growth models for land use systems analysis: II. Comparison of three simple approaches. Ecological Modelling, 148(3):233–250, March 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004367.

vanderHoff:2009:DSB

[vdHGF09] Quay van der Hoff, Johanna C. Greeff, and Temple H. Fay. Defining a stability boundary for three species competition models. Ecological Modelling, 220(20):2640–2645, October 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009005274.

vanderPerk:2000:SMT

[vdPLG^{+}00] M. van der Perk, T. Lev, A. G. Gillett, J. P. Absalom, P. A. Burrough, N. M. J. Crout, E. K. Garger, N. Semiochkina, Y. V. Stephanishin, and G. Voigt. Spatial modelling of transfer of long-lived radionuclides from soil to agricultural products in the chernigov region, Ukraine. Ecological Modelling, 128(1):35–50, April 3, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002252.

vanderPeijl:2000:SEN

[vdPvOV00] M. J. van der Peijl, M. M. P. van Oorschot, and J. T. A. Verhoeven. Simulation of the effects of nutrient enrichment on nutrient and carbon dynamics in a river marginal wetland. Ecological Modelling, 134(2–3):169–184, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003598.

Vuilleumier:2007:LSA

[VF07] Séverine Vuilleumier and Pierre Fontanillas. Landscape structure affects dispersal in the greater white-toothed shrew: Inference between genetic and simulated ecological distances. Ecological Modelling, 201(3–4):369–376, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006004819.

Verdesca:2006:EEA

[VFTB06] D. Verdesca, M. Federici, L. Torsello, and R. Basosi. Exergyeconomic accounting for sea-coastal systems: a novel approach. Ecological Modelling, 193(1–2):132–139, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004618.

Voinov:2008:LSP

[VG08] Alexey Voinov and Erica J. Brown Gaddis. Lessons for successful participatory watershed modeling: a perspective from modeling practitioners. Ecological Modelling, 216(2):197–207, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008001403.

Valavanis:2004:GEM

[VGK^{+}04] Vasilis D. Valavanis, Stratis Georgakarakos, Argyris Kapantagakis, Andreas Palialexis, and Isidora Katara. A GIS environmental modelling approach to essential fish habitat designation. Ecological Modelling, 178(3–4):417–427, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004001747.

VanKirk:2007:DMP

[VH07] Robert W. Van Kirk and Sheryl L. Hill. Demographic model predicts trout population response to selenium based on individual-level toxicity. Ecological Modelling, 206 (3–4):407–420, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007002001.

Vergara:2009:LEE

[VH09] Pablo M. Vergara and Ingo Hahn. Linking edge effects and patch size effects: Importance of matrix nest predators. Ecological Modelling, 220(9–10):1189–1196, May 17, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009001215.

Vallino:2005:EEG

[VHG05] J. J. Vallino, C. S. Hopkinson, and R. H. Garritt. Estimating estuarine gross production, community respiration and net ecosystem production: a nonlinear inverse technique. Ecological Modelling, 187(2–3):281–296, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005000633.

vanHorssen:2002:USA

[vHPS02] P. W. van Horssen, E. J. Pebesma, and P. P. Schot. Uncertainties in spatially aggregated predictions from a logistic regression model. Ecological Modelling, 154(1–2):93–101, August 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002000601.

Villanueva:2008:MTI

[VIK^{+}08] Maria Concepcion S. Villanueva, Mwapu Isumbisho, Boniface Kaningini, Jacques Moreau, and Jean-Claude Micha. Modeling trophic interactions in Lake Kivu: What roles do exotics play? Ecological Modelling, 212(3–4):422–438, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007005649.

Villa:2001:IMA

[Vil01] Ferdinando Villa. Integrating modelling architecture: a declarative framework for multi-paradigm, multi-scale ecological modelling. Ecological Modelling, 137(1):23–42, February 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000004221.

Vilizzi:2005:LMD

[Vil05] Lorenzo Vilizzi. The linear model diagram: a graphical method for the display of factor relationships in experimental design. Ecological Modelling, 184(2–4):263–275, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004005526.

Vilenkin:2006:ERR

[Vil06] B. Vilenkin. Ecological reading of random numbers. Ecological Modelling, 195(3–4):385–392, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006241.

Verburg:2001:SBB

[VJ01] P. S. J. Verburg and D. W. Johnson. A spreadsheetbased biogeochemical model to simulate nutrient cycling processes in forest ecosystems. Ecological Modelling, 141(1–3):185–200, July 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001002733.

Vasas:2006:TKS

[VJ06] Vera Vasas and Ferenc Jordán. Topological keystone species in ecological interaction networks: Considering link quality and non-trophic effects. Ecological Modelling, 196(3–4):365–378, July 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006000755.

Vladusic:2006:MLG

[VKB06] Daniel Vladušič, Boris Kompare, and Ivan Bratko. Modelling Lake Glumsø with Q2 learning. Ecological Modelling, 191(1):33–46, January 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005003479.

Vallina:2008:PUN

[VL08] S. M. Vallina and C. Le Quéré. Preferential uptake of NH4+ over NO_{3}^{−} in marine ecosystem models: a simple and more consistent parameterization. Ecological Modelling, 218(3–4):393–397, November 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003268.

Villanueva:2006:CAT

[VLA^{+}06] M. C. Villanueva, P. Lalèyè, J.-J. Albaret, R. Laë, L. Tito de Morais, and J. Moreau. Comparative analysis of trophic structure and interactions of two tropical lagoons. Ecological Modelling, 197(3–4):461–477, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001402.

Viotti:2002:AUP

[VLD02] P. Viotti, G. Liuti, and P. Di Genova. Atmospheric urban pollution: applications of an artificial neural network (ANN) to the city of Perugia. Ecological Modelling, 148(1):27–46, February 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004343.

vanLangevelde:2008:IIR

[vLDHP08] Frank van Langevelde, Michael Drescher, Ignas M. A. Heitkönig, and Herbert H. T. Prins. Instantaneous intake rate of herbivores as function of forage quality and mass: Effects on facilitative and competitive interactions. Ecological Modelling, 213(3–4):273–284, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006485.

vanLeeuwen:2007:RTC

[vLLv^{+}07] Edwin van Leeuwen, Gissell Lacerot, Egbert H. van Nes, Lia Hemerik, and Marten Scheffer. Reduced top–down control of phytoplankton in warmer climates can be explained by continuous fish reproduction. Ecological Modelling, 206(1–2):205–212, August 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001937.

Venturino:2001:RPB

[VM01] Ezio Venturino and Alexander B. Medvinsky. The role of periodic boundary forcing in plankton pattern formation. Ecological Modelling, 140(3):255–270, June 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003180.

Vuilleumier:2006:ADM

[VM06] Séverine Vuilleumier and Richard Metzger. Animal dispersal modelling: Handling landscape features and related animal choices. Ecological Modelling, 190(1–2):159–170, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000500222X.

Valleriani:2007:MSG

[VM07] Angelo Valleriani and Torsten Meene. Multilevel selection in a gradient. Ecological Modelling, 208(2–4):159–164, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007002992.

Vaclavik:2009:ISD

[VM09] Tomáš Václavík and Ross K. Meentemeyer. Invasive species distribution modeling (iSDM): Are absence data and dispersal constraints needed to predict actual distributions? Ecological Modelling, 220(23):3248–3258, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009005742.

Vazquez:2005:CAH

[VMG05] E. Vidal Vázquez, J. G. Vivas Miranda, and A. Paz González. Characterizing anisotropy and heterogeneity of soil surface microtopography using fractal models. Ecological Modelling, 182(3–4):337–353, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001681.

Vincent:2006:CES

[VML^{+}06] Céline Vincent, David Mouillot, Michel Lauret, Thang Do Chi, Marc Troussellier, and Catherine Aliaume. Contribution of exotic species, environmental factors and spatial components to the macrophyte assemblages in a Mediterranean lagoon (Thau lagoon, Southern France). Ecological Modelling, 193(1–2):119–131, March 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005004606.

Vorpahl:2009:MST

[VMR09] Peter Vorpahl, Sylvia Moenickes, and Otto Richter. Modelling of spatio-temporal population dynamics of earthworms under wetland conditions — an integrated approach. Ecological Modelling, 220(24):3647–3657, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001860.

vanNoordwijk:2002:STO

[vN02] Meine van Noordwijk. Scaling trade-offs between crop productivity, carbon stocks and biodiversity in shifting cultivation landscape mosaics: the FALLOW model. Ecological Modelling, 149(1–2):113–126, March 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100518X.

vanNes:2008:MED

[vNL^{+}08] Egbert H. van Nes, Ruurd Noordhuis, Eddy H. H. R. Lammens, Rob Portielje, Bart Reeze, and Edwin T. H. M. Peeters. Modelling the effects of diving ducks on zebra mussels Dreissena polymorpha in lakes. Ecological Modelling, 211(3–4):481–490, March 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007005273.

vanNes:2002:PIB

[vNLS02] Egbert H. van Nes, Eddy H. R. R. Lammens, and Marten Scheffer. PISCATOR, an individual-based model to analyze the dynamics of lake fish communities. Ecological Modelling, 152(2–3):261–278, July 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000054.

vanNoordwijk:2002:FBA

[vNM02] Meine van Noordwijk and Rachmat Mulia. Functional branch analysis as tool for fractal scaling above- and belowground trees for their additive and non-additive properties. Ecological Modelling, 149(1–2):41–51, March 15, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001005130.

vanNes:2003:AAM

[vNS03] Egbert H. van Nes and Marten Scheffer. Alternative attractors may boost uncertainty and sensitivity in ecological models. Ecological Modelling, 159(2–3):117–124, January 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002002788.

vanNes:2003:CSE

[vNSvdBC03] Egbert H. van Nes, Marten Scheffer, Marcel S. van den Berg, and Hugo Coops. Charisma: a spatial explicit simulation model of submerged macrophytes. Ecological Modelling, 159(2–3):103–116, January 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002752.

vanOijen:2004:SED

[vODFS04] M. van Oijen, M. F. Dreccer, K.-H. Firsching, and B. J. Schnieders. Simple equations for dynamic models of the effects of CO_{2} and O_{3} on light-use efficiency and growth of crops. Ecological Modelling, 179(1):39–60, November 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000400256X.

Villanueva:2006:TRR

[VOM06] Maria Concepcion Villanueva, Maxime Ouedraogo, and Jacques Moreau. Trophic relationships in the recently impounded Bagré reservoir in Burkina Faso. Ecological Modelling, 191(2):243–259, January 31, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002735.

Vayenas:2001:CDM

[VP01] Dimitris V. Vayenas and Stavros Pavlou. Chaotic dynamics of a microbial system of coupled food chains. Ecological Modelling, 136(2–3):285–295, January 20, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000004373.

Villavicencio-Pulido:2008:LQV

[VPB08] Geiser Villavicencio-Pulido and Ignacio Barradas. Latency and quarantine vs. backward bifurcation. Ecological Modelling, 214(1):59–64, June 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000264.

Viscido:2005:EPS

[VPG05] Steven V. Viscido, Julia K. Parrish, and Daniel Grünbaum. The effect of population size and number of influential neighbors on the emergent properties of fish schools. Ecological Modelling, 183(2–3):347–363, April 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004958.

Viscido:2007:FIS

[VPG07] Steven V. Viscido, Julia K. Parrish, and Daniel Grünbaum. Factors influencing the structure and maintenance of fish schools. Ecological Modelling, 206(1–2):153–165, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001755.

vanPuijenbroek:2004:IMN

[vPJK04] P. J. T. M. van Puijenbroek, J. H. Janse, and J. M. Knoop. Integrated modelling for nutrient loading and ecology of lakes in the Netherlands. Ecological Modelling, 174(1–2):127–141, May 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004000031.

Vaniscotte:2009:MSD

[VPR^{+}09] Amélie Vaniscotte, David R. J. Pleydell, Francis Raoul, Jean Pierre Quéré, Qiu Jiamin, Qian Wang, Li Tiaoying, Nadine Bernard, Michael Coeurdassier, Pierre Delattre, Kenichi Takahashi, Jean-Christophe Weidmann, and Patrick Giraudoux. Modelling and spatial discrimination of small mammal assemblages: an example from western Sichuan (China). Ecological Modelling, 220(9–10):1218–1231, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001483.

Viotti:2005:CTU

[VPSG05] P. Viotti, M. Petrangeli Papini, N. Stracqualursi, and C. Gamba. Contaminant transport in an unsaturated soil: laboratory tests and numerical simulation model as procedure for parameters evaluation. Ecological Modelling, 182(2):131–148, March 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004004090.

Vasiliadou:2009:DCT

[VPV09] I. A. Vasiliadou, S. Pavlou, and D. V. Vayenas. Dynamics of a chemostat with three competitive hydrogen oxidizing denitrifying microbial populations and their efficiency for denitrification. Ecological Modelling, 220(8):1169–1180, April 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009000854.

Vaca-Rodriguez:2006:AEP

[VREA06] Juan Guillermo Vaca-Rodríguez and Roberto Ramón Enríquez-Andrade. Analysis of the eastern Pacific yellowfin tuna fishery based on multiple management objectives. Ecological Modelling, 191(2):275–290, January 31, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005002711.

VanNes:2005:SIC

[VS05] Egbert H. Van Nes and Marten Scheffer. A strategy to improve the contribution of complex simulation models to ecological theory. Ecological Modelling, 185(2–4):153–164, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000400612X.

Vabo:2008:ESS

[VS08a] Rune Vabø and Georg Skaret. Emerging school structures and collective dynamics in spawning herring: a simulation study. Ecological Modelling, 214(2–4):125–140, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000525.

VanGossum:2008:DMS

[VS08b] Hans Van Gossum and Thomas N. Sherratt. A dynamical model of sexual harassment in damselflies and its implications for female-limited polymorphism. Ecological Modelling, 210(1–2):212–220, January 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004012.

Vitale:2003:MLG

[VSFM03] Marcello Vitale, Mauro Scimone, Enrico Feoli, and Fausto Manes. Modelling leaf gas exchanges to predict functional trends in Mediterranean Quercus ilex forest under climatic changes in temperature. Ecological Modelling, 166(1–2):123–134, August 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001297.

Verbeeck:2008:MYM

[VSG^{+}08] Hans Verbeeck, Roeland Samson, André Granier, Pierre Montpied, and Raoul Lemeur. Multi-year model analysis of GPP in a temperate beech forest in France. Ecological Modelling, 210(1–2):85–103, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003675.

VanDerWal:2009:SPA

[VSGW09] Jeremy VanDerWal, Luke P. Shoo, Catherine Graham, and Stephen E. Williams. Selecting pseudo-absence data for presence-only distribution modeling: How far should you stray from what you know? Ecological Modelling, 220(4):589–594, February 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005486.

Vernier:2008:GSH

[VSHC08] Pierre R. Vernier, Fiona K. A. Schmiegelow, Susan Hannon, and Steve G. Cumming. Generalizability of songbird habitat models in boreal mixedwood forests of Alberta. Ecological Modelling, 211(1–2):191–201, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004541.

Valenti:2008:EMP

[VTB^{+}08] D. Valenti, L. Tranchina, M. Brai, A. Caruso, C. Cosentino, and B. Spagnolo. Environmental metal pollution considered as noise: Effects on the spatial distribution of benthic foraminifera in two coastal marine areas of Sicily (Southern Italy). Ecological Modelling, 213(3–4):449–462, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000549.

Vinatier:2009:CSE

[VTL^{+}09] Fabrice Vinatier, Philippe Tixier, Christophe Le Page, Pierre-François Duyck, and Françoise Lescourret. COS-MOS, a spatially explicit model to simulate the epidemiology of Cosmopolites sordidus in banana fields. Ecological Modelling, 220(18):2244–2254, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004062.

Verstraeten:2006:TWL

[VVF06] Willem W. Verstraeten, Frank Veroustraete, and Jan Feyen. On temperature and water limitation of net ecosystem productivity: Implementation in the C-Fix model. Ecological Modelling, 199(1):4–22, November 1, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002778.

vanWijk:2007:PEF

[vW07] M. T. van Wijk. Predicting ecosystem functioning from plant traits: Results from a multi-scale ecophysiological modeling approach. Ecological Modelling, 203(3–4):453–463, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006006326.

vanWijk:2002:CDM

[vWBV02] M. T. van Wijk, W. Bouten, and J. M. Verstraten. Comparison of different modelling strategies for simulating gas exchange of a Douglas-fir forest. Ecological Modelling, 158(1–2):63–81, December 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001746.

vonWaldow:2008:MEE

[vWSH08] Harald von Waldow, Martin Scheringer, and Konrad Hungerbühler. Modelled environmental exposure to persistent organic chemicals is independent of the time course of emissions: Proof and significance for chemical exposure assessments. Ecological Modelling, 219(1–2):256–259, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000800416X.

Verbunt:2005:HIL

[VZG05] M. Verbunt, M. Groot Zwaaftink, and J. Gurtz. The hydrologic impact of land cover changes and hydropower stations in the alpine Rhine basin. Ecological Modelling, 187(1):71–84, September 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005000414.

Wood:2002:GIM

[WA02] Simon N. Wood and Nicole H. Augustin. GAMs with integrated model selection using penalized regression splines and applications to environmental modelling. Ecological Modelling, 157(2–3):157–177, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000200193X.

Webby:2007:MAV

[WAB^{+}07] R. B. Webby, P. T. Adamson, J. Boland, P. G. Howlett, A. V. Metcalfe, and J. Piantadosi. The Mekong — applications of value at risk (VaR) and conditional value at risk (CVaR) simulation to the benefits, costs and consequences of water resources development in a large river basin. Ecological Modelling, 201(1):89–96, February 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003590.

Walters:2001:LPP

[Wal01] Steven Walters. Landscape pattern and productivity effects on source–sink dynamics of deer populations. Ecological Modelling, 143(1–2):17–32, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003520.

Wallace:2004:PIH

[Wal04] Robert G. Wallace. Projecting the impact of HAART on the evolution of HIV’s life history. Ecological Modelling, 176(3–4):227–253, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400047X.

Wang:2005:DSA

[Wan05] Shusen Wang. Dynamics of surface albedo of a boreal forest and its simulation. Ecological Modelling, 183(4):477–494, May 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004005289.

Wang:2007:LSE

[Wan07] Guiming Wang. On the latent state estimation of nonlinear population dynamics using Bayesian and non-Bayesian state-space models. Ecological Modelling, 200 (3–4):521–528, January 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004133.

Ward:2008:RCF

[War08] Eric J. Ward. A review and comparison of four commonly used Bayesian and maximum likelihood model selection tools. Ecological Modelling, 211(1–2):1–10, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005558.

Wilhelm:2000:GFD

[WB00] Thomas Wilhelm and Rainer Brüggemann. Goal functions for the development of natural systems. Ecological Modelling, 132(3):231–246, August 5, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002465.

Weisberg:2005:MIE

[WBB05] Peter J. Weisberg, Francesco Bonavia, and Harald Bugmann. Modeling the interacting effects of browsing and shading on mountain forest tree regeneration (Picea abies). Ecological Modelling, 185(2–4):213–230, July 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004006155.

Wehde:2001:IOC

[WBH01] Henning Wehde, Jan O. Backhaus, and Else Nøst Hegseth. The influence of oceanic convection in primary production. Ecological Modelling, 138(1–3):115–126, March 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003975.

Wohlfahrt:2003:CSV

[WBN^{+}03] Georg Wohlfahrt, Michael Bahn, Christian Newesely, Sigrid Sapinsky, Ulrike Tappeiner, and Alexander Cernusca. Canopy structure versus physiology effects on net photosynthesis of mountain grasslands differing in land use. Ecological Modelling, 170(2–3):407–426, December 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002424.

Whipple:2007:IED

[WBP^{+}07] Stuart J. Whipple, Stuart R. Borrett, Bernard C. Patten, David K. Gattie, John R. Schramski, and Seth A. Bata. Indirect effects and distributed control in ecosystems: Comparative network environ analysis of a seven-compartment model of nitrogen flow in the Neuse River estuary, USA — time series analysis. Ecological Modelling, 206(1–2):1–17, August 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007001202.

Wunder:2006:OTM

[WBR^{+}06] Jan Wunder, Christof Bigler, Björn Reineking, Lorenz Fahse, and Harald Bugmann. Optimisation of tree mortality models based on growth patterns. Ecological Modelling, 197(1–2):196–206, August 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001025.

Wyse:2007:SMM

[WBR07] Ana Paula P. Wyse, Luiz Bevilacqua, and Marat Rafikov. Simulating malaria model for different treatment intensities in a variable environment. Ecological Modelling, 206(3–4):322–330, August 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007001871.

Wang:2008:PGD

[WBR08] Ying Wang, William L. Bauerle, and Robert F. Reynolds. Predicting the growth of deciduous tree species in response to water stress: FVS-BGC model parameterization, application, and evaluation. Ecological Modelling, 217(1–2):139–147, September 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002950.

Woodbury:2002:UPE

[WBS^{+}02] Peter B. Woodbury, Ronald M. Beloin, Dennis P. Swaney, Brian E. Gollands, and David A. Weinstein. Using the ECLPSS software environment to build a spatially explicit component-based model of ozone effects on forest ecosystems. Ecological Modelling, 150(3):211–238, May 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004653.

Wohlfahrt:2000:MWP

[WBTC00] G. Wohlfahrt, M. Bahn, U. Tappeiner, and A. Cernusca. A model of whole plant gas exchange for herbaceous species from mountain grassland sites differing in land use. Ecological Modelling, 125(2–3):173–201, January 15, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099001805.

Wickramasuriya:2009:DSC

[WBvDHZ09] Rohan Chandralal Wickramasuriya, Arnold K. Bregt, Hedwig van Delden, and Alex Hagen-Zanker. The dynamics of shifting cultivation captured in an extended constrained cellular automata land use model. Ecological Modelling, 220(18):2302–2309, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003883.

Walker:2003:SME

[WC03] Sean E. Walker and William H. Cade. A simulation model of the effects of frequency dependence, density dependence and parasitoid flies on the fitness of male field crickets. Ecological Modelling, 169(1):119–130, November 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003002667.

Walters:2007:ARF

[WC07] Carl Walters and Villy Christensen. Adding realism to foraging arena predictions of trophic flow rates in Ecosim ecosystem models: Shared foraging arenas and bout feeding. Ecological Modelling, 209(2–4):342–350, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003493.

Wang:2002:NCM

[WCC02] Shusen Wang, Wenjun Chen, and Josef Cihlar. New calculation methods of diurnal distribution of solar radiation and its interception by canopy over complex terrain. Ecological Modelling, 155(2–3):191–204, October 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001229.

Wang:2008:MBS

[WCH08] Jihuai Wang, Svend Christensen, and Preben Klarskov Hansen. A method for building spatial model of annual weed seed dispersal from experimental data and its application to simulating Bromus sterilis population dispersal. Ecological Modelling, 210(4):446–452, February 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004176.

Wang:2002:SSD

[WCHM02] Hongqing Wang, Joseph D. Cornell, Charles A. S. Hall, and David P. Marley. Spatial and seasonal dynamics of surface soil carbon in the Luquillo Experimental Forest, Puerto Rico. Ecological Modelling, 147(2):105–122, January 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004148.

White:2000:ENZ

[WCS00] J. D. White, N. C. Coops, and N. A. Scott. Estimates of New Zealand forest and scrub biomass from the 3-PG model. Ecological Modelling, 131(2–3):175–190, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002519.

Wu:2002:SEH

[WD02] Jianguo Wu and John L. David. A spatially explicit hierarchical approach to modeling complex ecological systems: theory and applications. Ecological Modelling, 153(1–2):7–26, July 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004999.

Wiens:2008:TWF

[WDBK08] Trevor S. Wiens, Brenda C. Dale, Mark S. Boyce, and G. Peter Kershaw. Three way k-fold cross-validation of resource selection functions. Ecological Modelling, 212(3–4):244–255, April 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007005339.

Whigham:2006:ESP

[WDR06] P. A. Whigham, G. Dick, and F. Recknagel. Exploring seasonal patterns using process modelling and evolutionary computation. Ecological Modelling, 195(1–2):146–152, May 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005005867.

Wefering:2000:UAA

[WDW00] Frank M. Wefering, Leon E. Danielson, and Nancy M. White. Using the AMOEBA approach to measure progress toward ecosystem sustainability within a shellfish restoration project in North Carolina. Ecological Modelling, 130(1–3):157–166, June 1, 2000. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380000002052.

Wegehenkel:2000:TMS

[Weg00] Martin Wegehenkel. Test of a modelling system for simulating water balances and plant growth using various different complex approaches. Ecological Modelling, 129(1):39–64, May 3, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000002210.

Welk:2004:CRP

[Wel04] Erik Welk. Constraints in range predictions of invasive plant species due to non-equilibrium distribution patterns: Purple loosestrife (Lythrum salicaria) in North America. Ecological Modelling, 179(4):551–567, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002807.

Wang:2001:DUE

[WEW01] Guiming Wang, W. Daniel Edge, and Jerry O. Wolff. Demographic uncertainty in ecological risk assessments. Ecological Modelling, 136(1):95–102, January 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003781.

Walder:2008:ALH

[WFB^{+}08] Konrad Wälder, Nico Frischbier, Michael Bredemeier, Wolfgang Näther, and Sven Wagner. Analysis of O_{F} -layer humus mass variation in a mixed stand of European beech and Norway spruce: an application of structural equation modelling. Ecological Modelling, 213(3–4):319–330, May 24, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008000045.

Wainwright:2007:CTL

[WFHP07] Thomas C. Wainwright, Leah R. Feinberg, Rian C. Hooff, and William T. Peterson. A comparison of two lower trophic models for the California Current system. Ecological Modelling, 202(1–2):120–131, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004613.

Weber:2001:LTL

[WFM01] Andrea Weber, Nicola Fohrer, and Detlev Möller. Longterm land use changes in a mesoscale watershed due to socio-economic factors — effects on landscape structures and functions. Ecological Modelling, 140(1–2):125–140, May 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002617.

Wu:2005:PEE

[WFT05] Jun Wu, Makoto Fukuhara, and Tatsuoki Takeda. Parameter estimation of an ecological system by a neural network with residual minimization training. Ecological Modelling, 189(3–4):289–304, December 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002176.

Weinstoerffer:2000:ACL

[WG00] Jean Weinstoerffer and Philippe Girardin. Assessment of the contribution of land use pattern and intensity to landscape quality: use of a landscape indicator. Ecological Modelling, 130(1–3):95–109, June 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000000209X.

Wilmers:2004:SEW

[WG04] Christopher C. Wilmers and Wayne M. Getz. Simulating the effects of wolf-elk population dynamics on resource flow to scavengers. Ecological Modelling, 177(1–2):193–208, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004001164.

Wang:2007:SSM

[WG07a] Guiming Wang and Lowell L. Getz. State-space models for stochastic and seasonal fluctuations of vole and shrew populations in east-central Illinois. Ecological Modelling, 207(2–4):189–196, October 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002608.

Wang:2007:HRD

[WG07b] Magnus Wang and Volker Grimm. Home range dynamics and population regulation: an individual-based model of the common shrew Sorex araneus. Ecological Modelling, 205(3–4):397–409, July 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001226.

White:2008:MMV

[WGB^{+}08] Joseph D. White, Kevin J. Gutzwiller, Wylie C. Barrow, Lori Johnson Randall, and Pamela Swint. Modeling mechanisms of vegetation change due to fire in a semi-arid ecosystem. Ecological Modelling, 214(2–4):181–200, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000756.

Wang:2002:STP

[WGS^{+}02] Guangxing Wang, George Gertner, Vivek Singh, Svetlana Shinkareva, Pablo Parysow, and Alan Anderson. Spatial and temporal prediction and uncertainty of soil loss using the revised universal soil loss equation: a case study of the rainfall–runoff erosivity R factor. Ecological Modelling, 153(1–2):143–155, July 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005075.

Woodall:2005:ASA

[WGT05] C. W. Woodall, P. L. Grambsch, and W. Thomas. Applying survival analysis to a large-scale forest inventory for assessment of tree mortality in Minnesota. Ecological Modelling, 189(1–2):199–208, November 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500195X.

Willems:2008:WWP

[WGV^{+}08] Wouter Willems, Peter Goethals, Dries Van den Eynde, Gert Van Hoey, Vera Van Lancker, Els Verfaillie, Magda Vincx, and Steven Degraer. Where is the worm? Predictive modelling of the habitat preferences of the tubebuilding polychaete Lanice conchilega. Ecological Modelling, 212(1–2):74–79, March 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007005182.

Wang:2001:MPC

[WGVB01] S. Wang, R. F. Grant, D. L. Verseghy, and T. A. Black. Modelling plant carbon and nitrogen dynamics of a boreal aspen forest in CLASS — the Canadian land surface scheme. Ecological Modelling, 142(1–2):135–154, August 1, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001002848.

Weclaw:2004:SCM

[WH04] Piotr Weclaw and Robert J. Hudson. Simulation of conservation and management of woodland caribou. Ecological Modelling, 177(1–2):75–94, September 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001085.

Winkler:2007:SAD

[WH07] Eckart Winkler and Thilo Heinken. Spread of an antdispersed annual herb: an individual-based simulation study on population development of Melampyrum pratense L. Ecological Modelling, 203(3–4):424–438, May 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006314.

Wu:2009:DAO

[WH09] Xingjie Wu and Wentao Huang. Dynamic analysis of a one-prey multi-predator impulsive system with ivlev-type functional. Ecological Modelling, 220(6):774–783, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009000325.

Wang:2008:MOG

[WHH^{+}08] Hongqing Wang, Wenrui Huang, Mark A. Harwell, Lee Edmiston, Elijah Johnson, Ping Hsieh, Katherine Milla, John Christensen, Jessica Stewart, and Xiaohai Liu. Modeling oyster growth rate by coupling oyster population and hydrodynamic models for Apalachicola Bay, Florida, USA. Ecological Modelling, 211(1–2):77–89, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004267.

Wang:2007:MSS

[WHHH07] Hongqing Wang, Y. Ping Hsieh, Mark A. Harwell, and Wenrui Huang. Modeling soil salinity distribution along topographic gradients in tidal salt marshes in Atlantic and Gulf coastal regions. Ecological Modelling, 201 (3–4):429–439, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004959.

Whigham:2000:IMD

[Whi00] P. A. Whigham. Induction of a marsupial density model using genetic programming and spatial relationships. Ecological Modelling, 131(2–3):299–317, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002489.

White:2005:EAE

[Whi05] Ben White. An economic analysis of ecological monitoring. Ecological Modelling, 189(3–4):241–250, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005002097.

Winslow:2001:GAM

[WHP01] Jerome C. Winslow, E. Raymond Hunt, and Stephen C. Piper. A globally applicable model of daily solar irradiance estimated from air temperature and precipitation data ^{1}. Ecological Modelling, 143(3):227–243, September 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003416.

Winslow:2003:ISW

[WHP03] Jerome C. Winslow, E. Raymond Hunt, and Stephen C. Piper. The influence of seasonal water availability on global C_{3} versus C_{4} grassland biomass and its implications for climate change research. Ecological Modelling, 163(1–2):153–173, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004155.

Wattenbach:2005:SAI

[WHW^{+}05] Martin Wattenbach, Fred Hattermann, Reinhard Weng, Frank Wechsung, Valentina Krysanova, and Franz Badeck. A simplified approach to implement forest eco-hydrological properties in regional hydrological modelling. Ecological Modelling, 187(1):40–59, September 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000402.

Wang:2003:DOD

[WHX^{+}03] H. Wang, M. Hondzo, C. Xu, V. Poole, and A. Spacie. Dissolved oxygen dynamics of streams draining an urbanized and an agricultural catchment. Ecological Modelling, 160(1–2):145–161, February 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003241.

Wu:2006:PTS

[WHZ06] Wei Wu, Charles Hall, and Lianjun Zhang. Predicting the temporal and spatial probability of orographic cloud cover in the Luquillo Experimental Forest in Puerto Rico using generalized linear (mixed) models. Ecological Modelling, 192(3–4):473–498, February 25, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003777.

Wang:2009:HAT

[WIH^{+}09] Weile Wang, Kazuhito Ichii, Hirofumi Hashimoto, Andrew R. Michaelis, Peter E. Thornton, Beverly E. Law, and Ramakrishna R. Nemani. A hierarchical analysis of terrestrial ecosystem model Biome-BGC: Equilibrium analysis and model calibration. Ecological Modelling, 220(17):2009–2023, September 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003111.

Wilder:2001:EIC

[Wil01] J. W. Wilder. Effect of initial condition sensitivity and chaotic transients on predicting future outbreaks of gypsy moths. Ecological Modelling, 136(1):49–66, January 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003859.

Wilhelm:2003:EDM

[Wil03] Thomas Wilhelm. An elementary dynamic model for non-binary food webs. Ecological Modelling, 168(1– 2):145–152, October 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003002072.

Wilhelm:2005:SEN

[Wil05] Thomas Wilhelm. A study of equal node models for food webs. Ecological Modelling, 181(2–3):215–227, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003680.

Williams:2007:OMN

[Wil07] Byron K. Williams. Optimal management of non-Markovian biological populations. Ecological Modelling, 200(1–2):234–242, January 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006003668.

Willis:2008:SMU

[Wil08] Jay Willis. Simulation model of universal law of school size distribution applied to southern bluefin tuna (Thunnus maccoyii) in the Great Australian Bight. Ecological Modelling, 213(1):33–44, April 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000367.

Williams:2009:MDP

[Wil09] Byron K. Williams. Markov decision processes in natural resources management: Observability and uncertainty. Ecological Modelling, 220(6):830–840, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000246.

Wimberly:2004:FFL

[Wim04] Michael C. Wimberly. Fire and forest landscapes in the Georgia Piedmont: an assessment of spatial modeling assumptions. Ecological Modelling, 180(1):41–56, December 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003485.

Werner:2007:SNM

[WIMK07] Francisco E. Werner, Shin-Ichi Ito, Bernard A. Megrey, and Michio J. Kishi. Synthesis of the NEMURO model studies and future directions of marine ecosystem modeling. Ecological Modelling, 202(1–2):211–223, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004844.

Wirtz:2000:SOS

[Wir00] Kai W. Wirtz. Second order up-scaling: theory and an exercise with a complex photosynthesis model. Ecological Modelling, 126(1):59–71, February 1, 2000. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438009900174X.

Witting:2002:EDE

[Wit02] Lars Witting. Evolutionary dynamics of exploited populations selected by density dependent competitive interactions. Ecological Modelling, 157(1):51–68, November 15, 2002. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380002001722.

Wolesensky:2005:MDM

[WJL05] William Wolesensky, Anthony Joern, and J. David Logan. A model of digestion modulation in grasshoppers. Ecological Modelling, 188(2–4):358–373, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001663.

Wichmann:2003:ERP

[WJM^{+}03] Matthias C. Wichmann, Karin Johst, Kirk A. Moloney, Christian Wissel, and Florian Jeltsch. Extinction risk in periodically fluctuating environments. Ecological Modelling, 167(3):221–231, September 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001364.

With:2004:ELS

[WK04] Kimberly A. With and Anthony W. King. The effect of landscape structure on community self-organization and critical biodiversity. Ecological Modelling, 179(3):349–366, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004002546.

Wang:2003:UCM

[WKL^{+}03] Jihuai Wang, M. J. Kropff, B. Lammert, S. Christensen, and P. K. Hansen. Using CA model to obtain insight into mechanism of plant population spread in a controllable system: annual weeds as an example. Ecological Modelling, 166(3):277–286, August 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001649.

Wei:2003:DSL

[WKZ03] X. Wei, J. P. Kimmins, and G. Zhou. Disturbances and the sustainability of long-term site productivity in lodgepole pine forests in the central interior of British Columbia — an ecosystem modeling approach. Ecological Modelling, 164(2–3):239–256, June 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000620.

Wang:2004:SVE

[WKZP04] Kai-Yun Wang, Seppo Kellomäki, Tianshan Zha, and Heli Peltola. Seasonal variation in energy and water fluxes in a pine forest: an analysis based on eddy covariance and an integrated model. Ecological Modelling, 179(3):259–279, November 30, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004001036.

Wildhaber:2004:IHC

[WL04] Mark L. Wildhaber and Peter J. Lamberson. Importance of the habitat choice behavior assumed when modeling the effects of food and temperature on fish populations. Ecological Modelling, 175(4):395–409, August 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005349.

Wang:2006:GLO

[WL06] Dong Wang and Wei-Zhen Lu. Ground-level ozone prediction using multilayer perceptron trained with an innovative hybrid approach. Ecological Modelling, 198 (3–4):332–340, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006002389.

Wackernagel:2004:GRM

[WLB^{+}04] H. Wackernagel, C. Lajaunie, N. Blond, C. Roth, and R. Vautard. Geostatistical risk mapping with chemical transport model output and ozone station data. Ecological Modelling, 179(2):177–185, November 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002571.

Watanabe:2005:BEK

[WLB^{+}05] Karen H. Watanabe, Hsin-I. Lin, Henry L. Bart, Peter Martinat, Jay C. Means, Myra L. Kunas, and Deborah A. Grimm. Bayesian estimation of kinetic rate constants in a food-web model of polycyclic aromatic hydrocarbon bioaccumulation. Ecological Modelling, 181(2–3):229–246, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004003692.

Walsh:2007:DML

[WLG07] Andrew S. Walsh, Thomas A. Louis, and Gregory E. Glass. Detecting multiple levels of effect during survey sampling using a Bayesian approach: Point prevalence estimates of a hantavirus in hispid cotton rats (Sigmodon hispidus). Ecological Modelling, 205(1–2):29–38, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000440.

Wu:2000:ELA

[WLJ00] Jianguo Wu, Yuanbo Liu, and Dennis E. Jelinski. Effects of leaf area profiles and canopy stratification on simulated energy fluxes: the problem of vertical spatial scale. Ecological Modelling, 134(2–3):283–297, October 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000003537.

White:2008:PAA

[WLL^{+}08] Thomas White, Nancy Luckai, Guy R. Larocque, Werner A. Kurz, and Carolyn Smyth. A practical approach for assessing the sensitivity of the carbon budget model of the Canadian forest sector (CBM-CFS3). Ecological Modelling, 219(3–4):373–382, December 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003608.

Wang:2004:FER

[WLLY04] Wan-Xiong Wang, Zi-Zhen Li, Yi-Min Li, and Fengxiang Yang. Fuzzy expression related to biotic community. Ecological Modelling, 179(3):367–372, November 30, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002662.

Westerberg:2008:EDC

[WLNW08] Lars Westerberg, Tom Lindström, Elna Nilsson, and Uno Wennergren. The effect on dispersal from complex correlations in small-scale movement. Ecological Modelling, 213(2):263–272, May 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006497.

Williamson:2006:ENT

[WLSE06] M. S. Williamson, T. M. Lenton, J. G. Shepherd, and N. R. Edwards. An efficient numerical terrestrial scheme (ENTS) for Earth system modelling. Ecological Modelling, 198(3–4):362–374, October 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002432.

Wu:2009:ISA

[WLZ^{+}09] Daqian Wu, Jian Liu, Gaosheng Zhang, Wenjuan Ding, Wei Wang, and Renqing Wang. Incorporating spatial autocorrelation into cellular automata model: an application to the dynamics of Chinese tamarisk (Tamarix chinensis lour.). Ecological Modelling, 220(24):3490–3498, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009001902.

Wang:2000:DEM

[WM00] Naiming Wang and William J. Mitsch. A detailed ecosystem model of phosphorus dynamics in created riparian wetlands. Ecological Modelling, 126(2–3):101–130, February 28, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000000260X.

Wu:2002:MCE

[WM02] Jianguo Wu and Danielle Marceau. Modeling complex ecological systems: an introduction. Ecological Modelling, 153(1–2):1–6, July 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004987.

Wegehenkel:2006:CGS

[WM06] Martin Wegehenkel and Wilfried Mirschel. Crop growth, soil water and nitrogen balance simulation on three experimental field plots using the O pus model — a case study. Ecological Modelling, 190(1–2):116–132, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002292.

Wang:2008:SHP

[WM08] Qian Wang and George P. Malanson. Spatial hyperdynamism in a post-disturbance simulated forest. Ecological Modelling, 215(4):337–344, July 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001804.

Weir:2008:LFB

[WMH08] Jackie N. Weir, Shawn F. Morrison, and David S. Hik. Linking foraging behavior to population density: an assessment of GMM models for Dall sheep. Ecological Modelling, 211(3–4):396–402, March 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004814.

Wu:2007:SIR

[WMM^{+}07] L. Wu, M. B. McGechan, N. McRoberts, J. A. Baddeley, and C. A. Watson. SPACSYS: Integration of a 3D root architecture component to carbon, nitrogen and water cycling — model description. Ecological Modelling, 200(3–4):343–359, January 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003802.

Wang:2009:MIB

[WMSW09] Hao Wang, Wendy Morrison, Abhinav Singh, and Howard (Howie) Weiss. Modeling inverted biomass pyramids and refuges in ecosystems. Ecological Modelling, 220(11):1376–1382, June 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009001823.

Walder:2009:IIM

[WNW09] Konrad Wälder, Wolfgang Näther, and Sven Wagner. Improving inverse model fitting in trees — anisotropy, multiplicative effects, and Bayes estimation. Ecological Modelling, 220(8):1044–1053, April 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000970.

Walley:2001:UPR

[WO01a] William J. Walley and Mark A. O’Connor. Unsupervised pattern recognition for the interpretation of ecological data. Ecological Modelling, 146(1–3):219–230, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380001003088.

Werner:2001:NNN

[WO01b] Heinrich Werner and Michael Obach. New neural network types estimating the accuracy of response for ecological modelling. Ecological Modelling, 146(1–3):289–298, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003143.

Westerberg:2005:MEE

[WÖW05] Lars Westerberg, Örjan Östman, and Uno Wennergren. Movement effects on equilibrium distributions of habitat generalists in heterogeneous landscapes. Ecological Modelling, 188(2–4):432–447, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001213.

Walker:2006:SME

[WPBM06] David M. Walker, F. Javier Pérez-Barbería, and Glenn Marion. Stochastic modelling of ecological processes using hybrid Gibbs samplers. Ecological Modelling, 198 (1–2):40–52, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006001682.

Wafa:2004:MGF

[WPD04] Aboul Hosn Wafa, Dutilleul Pierre, and Boisclair Daniel. Modelling growth and food intake rhythms of brook trout Salvelinus fontinalis under the effects of density and ration. Ecological Modelling, 175(4):385–394, August 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003005106.

Wauters:2007:RTS

[WPMT07] Luc A. Wauters, Damiano G. Preatoni, Ambrogio Molinari, and Guido Tosi. Radio-tracking squirrels: Performance of home range density and linkage estimators with small range and sample size. Ecological Modelling, 202(3–4):333–344, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006005369.

Whigham:2001:IAE

[WR01a] P. A. Whigham and Friedrich Recknagel. An inductive approach to ecological time series modelling by evolutionary computation. Ecological Modelling, 146(1– 3):275–287, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001003131.

Whigham:2001:PCF

[WR01b] P. A. Whigham and Friedrich Recknagel. Predicting chlorophyll- a in freshwater lakes by hybridising processbased models and genetic algorithms. Ecological Modelling, 146(1–3):243–251, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003106.

Wilson:2001:TGA

[WR01c] Hugh Wilson and Friedrich Recknagel. Towards a generic artificial neural network model for dynamic predictions of algal abundance in freshwater lakes. Ecological Modelling, 146(1–3):69–84, December 1, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001002976.

Witten:2003:CTA

[WR03] G. Q. Witten and F. D. Richardson. Competition of three aggregated microbial species for four substrates in the rumen. Ecological Modelling, 164(2–3):121–135, June 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003836.

Wagenschein:2008:MIR

[WR08] Dierk Wagenschein and Michael Rode. Modelling the impact of river morphology on nitrogen retention — a case study of the Weisse Elster River (Germany). Ecological Modelling, 211(1–2):224–232, February 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004723.

Weber:2008:SSD

[WRB08] Pascale Weber, Andreas Rigling, and Harald Bugmann. Sensitivity of stand dynamics to grazing in mixed Pinus sylvestris and Quercus pubescens forests: a modelling study.

Ecological Modelling, 210(3):301–311, January 24, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007004097.

Wernsdorfer:2008:IUT

[WRC^{+}08] H. Wernsdörfer, V. Rossi, G. Cornu, S. Oddou-Muratorio, and S. Gourlet-Fleury. Impact of uncertainty in tree mortality on the predictions of a tropical forest dynamics model. Ecological Modelling, 218(3–4):290–306, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003645.

Walter:2001:PEE

[WRCB01] Mark Walter, Friedrich Recknagel, Craig Carpenter, and Myriam Bormans. Predicting eutrophication effects in the Burrinjuck Reservoir (Australia) by means of the deterministic model SALMO and the recurrent neural network model ANNA. Ecological Modelling, 146(1–3):97–113, December 1, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000100299X.

Wagner:2007:CMI

[WRP07] L. D. Wagner, J. V. Ross, and H. P. Possingham. Catastrophe management and inter-reserve distance for marine reserve networks. Ecological Modelling, 201(1):82–88, February 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006003589.

Wilkinson:2002:MMP

[WS02] D. Wilkinson and G. C. Smith. A model of the mite parasite, Varroa destructor, on honeybees (Apis mellifera) to investigate parameters important to mite population growth. Ecological Modelling, 148(3):263–275, March 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001004409.

Waller:2003:MCA

[WSCR03] Lance A. Waller, David Smith, James E. Childs, and Leslie A. Real. Monte Carlo assessments of goodnessof-fit for ecological simulation models. Ecological Modelling, 164(1):49–63, June 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003000115.

Willocquet:2000:DEM

[WSF^{+}00] Laetitia Willocquet, Serge Savary, Luzviminda Fernandez, Francisco Elazegui, and Paul Teng. Development and evaluation of a multiple-pest, production situation specific model to simulate yield losses of rice in tropical Asia. Ecological Modelling, 131(2–3):133–159, July 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380000002714.

Willocquet:2002:SVR

[WSF^{+}02] L. Willocquet, S. Savary, L. Fernandez, F. A. Elazegui, N. Castilla, D. Zhu, Q. Tang, S. Huang, X. Lin, H. M. Singh, and R. K. Srivastava. Structure and validation of RI-CEPEST, a production situation-driven, crop growth model simulating rice yield response to multiple pest injuries for tropical Asia. Ecological Modelling, 153(3):247–268, August 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002000145.

Wilson:2008:MAP

[WSP08] Duncan S. Wilson, Margo A. Stoddard, and Klaus J. Puettmann. Monitoring amphibian populations with incomplete survey information using a Bayesian probabilistic model. Ecological Modelling, 214(2–4):210–218, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000768.

Weiss:2006:SWT

[WSS^{+}06] Robert Weiss, Narasinha J. Shurpali, Tapani Sallantaus, Raija Laiho, Jukka Laine, and Jukka Alm. Simulation of water table level and peat temperatures in boreal peatlands. Ecological Modelling, 192(3–4):441–456, February 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003534.

Wiegand:2008:RSI

[WSWL08] Kerstin Wiegand, David Saltz, David Ward, and Simon A. Levin. The role of size inequality in self-thinning: a patternoriented simulation model for arid savannas. Ecological Modelling, 210(4):431–445, February 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007004152.

Wu:2007:GSM

[WSY^{+}07] Wenbin Wu, Ryosuke Shibasaki, Peng Yang, Guoxin Tan, Kan ichiro Matsumura, and Kenji Sugimoto. Globalscale modelling of future changes in sown areas of major crops. Ecological Modelling, 208(2–4):378–390, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003171.

Wramneby:2008:PUM

[WSZS08] Anna Wramneby, Benjamin Smith, Sönke Zaehle, and Martin T. Sykes. Parameter uncertainties in the modelling of vegetation dynamics — effects on tree community structure and ecosystem functioning in European forest biomes. Ecological Modelling, 216(3–4):277–290, September 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008002032.

Warren:2004:TSM

[WT04a] John Warren and Chris Topping. A trait specific model of competition in a spatially structured plant community. Ecological Modelling, 180(4):477–485, December 31, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004002960.

Wu:2004:SOM

[WT04b] Fu-Chun Wu and Yin-Phan Tsang. Second-order Monte Carlo uncertainty/variability analysis using correlated model parameters: application to salmonid embryo survival risk assessment. Ecological Modelling, 177(3–4):393–414, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004001474.

Witting:2000:MOC

[WTL00] Lars Witting, Jürgen Tomiuk, and Volker Loeschcke. Modelling the optimal conservation of interacting species. Ecological Modelling, 125(2–3):123–144, January 15, 2000. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380099001775.

Wanko:2009:ADI

[WTMG09] Adrien Wanko, Gabriela Tapia, Robert Mosé, and Caroline Gregoire. Adsorption distribution impact on preferential transport within horizontal flow constructed wetland (HFCW). Ecological Modelling, 220(23):3342–3352, December 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009005730.

Wolff:2007:CFM

[WTMY07] Matthias Wolff, Marc Taylor, Jaime Mendo, and Carmen Yamashiro. A catch forecast model for the Peruvian scallop (Argopecten purpuratus) based on estimators of spawning stock and settlement rate. Ecological Modelling, 209(2–4):333–341, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700347X.

Wang:2006:ETD

[WTS^{+}06] Quan Wang, John Tenhunen, Markus Schmidt, Olimpia Kolcun, Matthias Droesler, and Markus Reichstein. Estimation of total, direct and diffuse PAR under clear skies in complex alpine terrain of the National Park Berchtesgaden, Germany. Ecological Modelling, 196(1–2):149–162, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000639.

Wallentin:2008:UAT

[WTST08] Gudrun Wallentin, Ulrike Tappeiner, Josef Strobl, and Erich Tasser. Understanding alpine tree line dynamics: an individual-based model. Ecological Modelling, 218(3–4):235–246, November 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003451.

Wan:2005:IEM

[WV05] Zhenwen Wan and Joe Vallino. An inverse ecosystem model of year-to-year variations with first order approximation to the annual mean fluxes. Ecological Modelling, 187(4):369–388, October 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001183.

Wan:2008:SIA

[WVP08] Zhenwen Wan, Joseph J. Vallino, and Bruce J. Peterson. Study of the inter-annual food web dynamics in the Kuparuk River with a first-order approximation inverse model. Ecological Modelling, 211(1–2):97–112, February 24, 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007004310.

Wolfer:2006:MCG

[WvS06] Susanne R. Wolfer, Egbert H. van Nes, and Dietmar Straile. Modelling the clonal growth of the rhizomatous macrophyte Potamogeton perfoliatus. Ecological Modelling, 192(1–2):67–82, February 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003285.

Westerberg:2003:PSD

[WW03] Lars Westerberg and Uno Wennergren. Predicting the spatial distribution of a population in a heterogeneous landscape. Ecological Modelling, 166(1–2):53–65, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001182.

Wild:2008:KGC

[WW08] Jan Wild and Eckart Winkler. Krummholz and grassland coexistence above the forest-line in the Krkonoše Mountains: Grid-based model of shrub dynamics. Ecological Modelling, 213(3–4):293–307, May 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000021.

Watts:2009:ERI

[WW09] Michael J. Watts and S. P. Worner. Estimating the risk of insect species invasion: Kohonen self-organising maps versus k-means clustering. Ecological Modelling, 220(6):821–829, March 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008005966.

Wulder:2007:ULS

[WWC^{+}07] Michael A. Wulder, Joanne C. White, Nicholas C. Coops, Trisalyn Nelson, and Barry Boots. Using local spatial autocorrelation to compare outputs from a forest growth model. Ecological Modelling, 209(2–4):264–276, December 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007003419.

Wang:2005:EMS

[WWL^{+}05] Hong-Zhu Wang, Hai-Jun Wang, Xiao-Min Liang, Le-Yi Ni, Xue-Qin Liu, and Yong-De Cui. Empirical modelling of submersed macrophytes in Yangtze lakes. Ecological Modelling, 188(2–4):483–491, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001237.

Wagner:2004:DFD

[WWRZ04] Sven Wagner, Konrad Wälder, Eric Ribbens, and André Zeibig. Directionality in fruit dispersal models for anemochorous forest trees. Ecological Modelling, 179 (4):487–498, December 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004002741.

Wang:2009:ASG

[WX09] Yuankun Wang and Ziqiang Xia. Assessing spawning ground hydraulic suitability for Chinese sturgeon (Acipenser sinensis) from horizontal mean vorticity in Yangtze River. Ecological Modelling, 220(11):1443–1448, June 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001811.

Wan-Xiong:2003:FDS

[WXYMZZF03] Wang Wan-Xiong, Li Yi-Min, Li Zi-Zhen, and Yang Fengxiang. A fuzzy description on some ecological concept. Ecological Modelling, 169(2–3):361–366, November 15, 2003. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002795.

West:2007:PSQ

[WYMS07] Andrew D. West, Michael G. Yates, Selwyn McGrorty, and Richard A. Stillman. Predicting site quality for shorebird communities: a case study on the wash embayment, UK. Ecological Modelling, 202(3–4):527–539, April 10, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006005709.

Watanabe:2005:SMR

[WYO05] Kozo Watanabe, Chihiro Yoshimura, and Tatsuo Omura. Stochastic model for recovery prediction of macroinvertebrates following a pulse-disturbance in river. Ecological Modelling, 189(3–4):396–412, December 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002188.

Wang:2009:ACS

[WYT09] Shusen Wang, Yan Yang, and Alexander P. Trishchenko. Assessment of canopy stomatal conductance models using flux measurements. Ecological Modelling, 220(17): 2115–2118, September 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003172.

Wang:2001:DMA

[WZ01] Yeqiao Wang and Xinsheng Zhang. A dynamic modeling approach to simulating socioeconomic effects on landscape changes. Ecological Modelling, 140(1–2):141–162, May 30, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001002629.

Wintrebert:2005:JMB

[WZC^{+}05] Claire M. A. Wintrebert, A. H. Zwinderman, E. Cam, R. Pradel, and J. C. van Houwelingen. Joint modelling of breeding and survival in the kittiwake using frailty models. Ecological Modelling, 181(2–3):203–213, January 20, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003679.

Wang:2008:MSN

[WZJ08] Yuhui Wang, Guangsheng Zhou, and Bingrui Jia. Modeling SOC and NPP responses of meadow steppe to different grazing intensities in Northeast China. Ecological Modelling, 217(1–2):72–78, September 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002494.

Wan:2009:MAT

[WZKL09] Liqiang Wan, Baisen Zhang, Peter Kemp, and Xianglin Li. Modelling the abundance of three key plant species in New Zealand hill-pasture using a decision tree approach. Ecological Modelling, 220(15):1819–1825, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009002877.

Wang:2005:EEE

[WZN05] Lingmei Wang, Jintun Zhang, and Weidou Ni. Emergy evaluation of eco-industrial park with power plant. Ecological Modelling, 189(1–2):233–240, November 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005001225.

Wang:2005:DSS

[WZW05] Zong-Ling Wang, Da-Yong Zhang, and Gang Wang. Does spatial structure facilitate coexistence of identical competitors? Ecological Modelling, 181(1):17–23, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003552.

Xing:2007:SNE

[XBM^{+}07] Zisheng Xing, Charles P.-A. Bourque, Fanrui Meng, Tianshan Zha, Roger M. Cox, and D. Edwin Swift. A simple net ecosystem productivity model for gap filling of towerbased fluxes: an extension of Landsberg’s equation with modifications to the light interception term. Ecological Modelling, 206(3–4):250–262, August 24, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007001792.

Xing:2008:MEM

[XBM^{+}08a] Zisheng Xing, Charles P.-A. Bourque, Fan-Rui Meng, Roger M. Cox, D. Edwin Swift, Tianshan Zha, and Lien Chow. Modification of an ecosystem model for filling medium-sized gaps in tower-based estimates of net ecosystem productivity. Ecological Modelling, 213(1):86–97, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007006278.

Xing:2008:PBM

[XBM^{+}08b] Zisheng Xing, Charles P.-A. Bourque, Fan-Rui Meng, Roger M. Cox, D. Edwin Swift, Tianshan Zha, and Lien Chow. A process-based model designed for filling of large data gaps in tower-based measurements of net ecosystem productivity. Ecological Modelling, 213(2):165–179, May 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007006266.

Xiao:2007:DEH

[XCW07] Sa Xiao, Shu-Yan Chen, and Gang Wang. Does the ESS height of plant population still exist with the inclusion of spatial structure? — an individual-based model research. Ecological Modelling, 204(1–2):213–218, May 24, 2007. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380007000063.

Xu:2009:UAT

[XG09] Chonggang Xu and George Z. Gertner. Uncertainty analysis of transient population dynamics. Ecological Modelling, 220(3):283–293, February 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800481X.

Xu:2009:CLU

[XGL^{+}09] Xia Xu, Qiong Gao, Ying-Hui Liu, Jing-Ai Wang, and Yong Zhang. Coupling a land use model and an ecosystem model for a crop-pasture zone. Ecological Modelling, 220(19):2503–2511, October 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003184.

Xiong:2008:CMC

[XHC^{+}08] Wei Xiong, Ian Holman, Declan Conway, Erda Lin, and Yue Li. A crop model cross calibration for use in regional climate impacts studies. Ecological Modelling, 213(3–4):365–380, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008000094.

Xu:2004:AEC

[XHH^{+}04] Chonggang Xu, Hong S. He, Yuanman Hu, Yu Chang, David R. Larsen, Xiuzhen Li, and Rencang Bu. Assessing the effect of cell-level uncertainty on a forest landscape model simulation in northeastern China. Ecological Modelling, 180(1):57–72, December 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003497.

Xu:2005:LHS

[XHH^{+}05] Chonggang Xu, Hong S. He, Yuanman Hu, Yu Chang, Xiuzhen Li, and Rencang Bu. Latin hypercube sampling and geostatistical modeling of spatial uncertainty in a spatially explicit forest landscape model simulation. Ecological Modelling, 185(2–4):255–269, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004006209.

Xiao:2000:GTF

[Xia00a] Yongshun Xiao. A general theory of fish stock assessment models. Ecological Modelling, 128(2–3):165–180, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000000199X.

Xiao:2000:IBA

[Xia00b] Yongshun Xiao. An individual-based approach to evaluating experimental designs for estimating rates of fish movement from tag recoveries. Ecological Modelling, 128(2–3):149–163, April 20, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438009900229X.

Xiao:2000:MTD

[Xia00c] Yongshun Xiao. Modelling temperature-dependency in biology by generalizing temperature coefficient Q_{10}. Ecological Modelling, 127(2–3):283–290, March 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380099002185.

Xiao:2002:RAM

[Xia02] Yongshun Xiao. Relationship among models for yield per recruit analysis, models for demographic analysis, and ageand time-dependent stock assessment models. Ecological Modelling, 155(2–3):95–125, October 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002000285.

Xiao:2003:DEL

[Xia03] Yongshun Xiao. Doppler effect in larval biology: theory and applications. Ecological Modelling, 165(1):1–22, July 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003563.

Xiao:2004:MLB

[Xia04] Yongshun Xiao. Modelling the learning behaviour of fishers: learning more from their successes than from their failures. Ecological Modelling, 171(1–2):3–20, January 1, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003002722.

Xiao:2005:CER

[Xia05a] Yongshun Xiao. Catch equations: restoring the missing terms in the nominally generalized baranov catch equation. Ecological Modelling, 181(4):535–556, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004003965.

Xiao:2005:EBW

[Xia05b] Yongshun Xiao. An equation of biological windows and its applications. Ecological Modelling, 181(4):521–533, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380004003953.

Xiao:2006:CEC

[Xia06] Yongshun Xiao. Catch equations: calculating the instantaneous rate of fishing mortality from catch and back. Ecological Modelling, 193(3–4):225–252, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000500400X.

Xiao:2007:FEM

[Xia07] Yongshun Xiao. The fundamental equations of multi-species virtual population analysis and its variants. Ecological Modelling, 201(3–4):477–494, March 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006005114.

Xue:2008:CLP

[XIX^{+}08] Huijie Xue, Lewis Incze, Danya Xu, Nicholas Wolff, and Neal Pettigrew. Connectivity of lobster populations in the coastal Gulf of Maine: Part i: Circulation and larval transport potential. Ecological Modelling, 210(1–2):193–211, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007004000.

Xu:2007:PBM

[XJM07] Shangping Xu, Peter R. Jaffé, and Denise L. Mauzerall. A process-based model for methane emission from flooded rice paddy systems. Ecological Modelling, 205(3–4):475–491, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007001469.

Xie:2001:DKB

[XLD01] Jialong Xie, Jianguo Liu, and Robert Doepker. DeerKBS: a knowledge-based system for white-tailed deer management. Ecological Modelling, 140(1–2):177–192, May 30, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380001002642.

Xu:2004:MCE

[XLZ^{+}04] F. L. Xu, K. C. Lam, Z. Y. Zhao, W. Zhan, Y. David Chen, and S. Tao. Marine coastal ecosystem health assessment: a case study of the Tolo Harbour, Hong Kong, China. Ecological Modelling, 173(4):355–370, April 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004241.

Xenakis:2008:SUA

[XRM08] Georgios Xenakis, Duncan Ray, and Maurizio Mencuccini. Sensitivity and uncertainty analysis from a coupled 3-PG and soil organic matter decomposition model. Ecological Modelling, 219(1–2):1–16, November 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003694.

Xia:2008:SWC

[XS08] Y. Q. Xia and M. A. Shao. Soil water carrying capacity for vegetation: a hydrologic and biogeochemical process model solution. Ecological Modelling, 214(2–4):112–124, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000537.

Xie:2009:CTM

[XSY^{+}09] Yichun Xie, Zongyao Sha, Mei Yu, Yongfei Bai, and Lei Zhang. A comparison of two models with Landsat data for estimating above ground grassland biomass in Inner Mongolia, China. Ecological Modelling, 220(15):1810–1818, August 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009002786.

Xu:2001:GBM

[XTDL01] Fu-Liu Xu, Shu Tao, R. W. Dawson, and Beng-Gang Li. A GIS-based method of lake eutrophication assessment. Ecological Modelling, 144(2–3):231–244, October 15, 2001. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000100374X.

Xu:2003:FEM

[Xu03] Fu-Liu Xu. Fundamentals of Ecological Modelling, 3rd Edition: S. E. Jørgensen, G. Bendoricchio (Eds.), Elsevier, 2001, 530 pages; ISBN: 0-08-044028-2 (Paperback edition); ISBN: 0-08-044015-0 (Hardbound edition). Ecological Modelling, 166(3):295–296, August 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001327.

Xiao:2003:DEE

[XV03] Yanni Xiao and Frank Van Den Bosch. The dynamics of an eco-epidemic model with biological control. Ecological Modelling, 168(1–2):203–214, October 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003001972.

Xue:2004:PLA

[XWB04] Qingwu Xue, Albert Weiss, and P. Stephen Baenziger. Predicting leaf appearance in field-grown winter wheat: evaluating linear and non-linear models. Ecological Modelling, 175(3):261–270, July 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004800.

Xu:2006:TME

[XZD^{+}06] Fu-Liu Xu, Shan-Shan Zhao, Richard W. Dawson, Jun-Yi Hao, Ying Zhang, and Shu Tao. A triangle model for evaluating the sustainability status and trends of economic development. Ecological Modelling, 195(3–4):327–337, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005006095.

Xu:2007:MMA

[XZS^{+}07] Lingling Xu, Xianzhou Zhang, Peili Shi, Wenhua Li, and Yongtao He. Modeling the maximum apparent quantum use efficiency of alpine meadow ecosystem on Tibetan Plateau. Ecological Modelling, 208(2–4):129–134, November 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007002864.

Xu:2005:EHI

[XZZ^{+}05] Fu-Liu Xu, Zhen-Yan Zhao, Wei Zhan, Shan-Shan Zhao, R. W. Dawson, and Shu Tao. An ecosystem health index methodology (EHIM) for lake ecosystem health assessment. Ecological Modelling, 188(2–4):327–339, November 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005001419.

Yuan:2007:EWE

[YABM07] Fengming Yuan, M. Altaf Arain, T. Andrew Black, and Kai Morgenstern. Energy and water exchanges modulated by soil–plant nitrogen cycling in a temperate Pacific Northwest conifer forest. Ecological Modelling, 201 (3–4):331–347, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000600487X.

Yard:2005:ITC

[YBM^{+}05] Michael D. Yard, Glenn E. Bennett, Steve N. Mietz, Lewis G. Coggins, Lawrence E. Stevens, Susan Hueftle, and Dean W. Blinn. Influence of topographic complexity on solar insolation estimates for the Colorado River, Grand Canyon, AZ. Ecological Modelling, 183(2–3):157–172, April 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004004375.

Yu:2001:MLTb

[YCVA01] Zicheng Yu, Ian D. Campbell, Dale H. Vitt, and Michael J. Apps. Modelling long-term peatland dynamics. I. Concepts, review, and proposed design. Ecological Modelling, 145(2–3):197–210, November 15, 2001. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100391X.

Yassemi:2008:DII

[YDS08] S. Yassemi, S. Dragićević, and M. Schmidt. Design and implementation of an integrated GIS-based cellular automata model to characterize forest fire behaviour. Ecological Modelling, 210(1–2):71–84, January 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003626.

Yearsley:2004:TPD

[Yea04] Jonathan M. Yearsley. Transient population dynamics and short-term sensitivity analysis of matrix population models. Ecological Modelling, 177(3–4):245–258, October 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004001188.

Yamada:2003:EIE

[YEMZ03] Kuniko Yamada, Jane Elith, Michael McCarthy, and André Zerger. Eliciting and integrating expert knowledge for wildlife habitat modelling. Ecological Modelling, 165(2–3):251–264, July 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000772.

Yearsley:2003:SAE

[YFH03] Jon M. Yearsley, David Fletcher, and Christine Hunter. Sensitivity analysis of equilibrium population size in a densitydependent model for short-tailed shearwaters. Ecological Modelling, 163(1–2):119–129, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004167.

Yue:2006:SMT

[YFLW06] Tian Xiang Yue, Ze Meng Fan, Ji Yuan Liu, and Bo Xin Wei. Scenarios of major terrestrial ecosystems in China. Ecological Modelling, 199(3):363–376, December 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006002365.

Yamamoto:2004:PNS

[YH04] Tamiji Yamamoto and Gen Hatta. Pulsed nutrient supply as a factor inducing phytoplankton diversity. Ecological Modelling, 171(3):247–270, January 15, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003363.

Yen:2004:LSM

[YHC04] P. P. W. Yen, F. Huettmann, and F. Cooke. A large-scale model for the at-sea distribution and abundance of Marbled Murrelets (Brachyramphus marmoratus) during the breeding season in coastal British Columbia, Canada. Ecological Modelling, 171(4):395–413, February 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003570.

Yang:2004:HFF

[YHG04] Jian Yang, Hong S. He, and Eric J. Gustafson. A hierarchical fire frequency model to simulate temporal patterns of fire regimes in LANDIS. Ecological Modelling, 180(1):119–133, December 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003783.

Yamamoto:2008:ENE

[YHTH08] Tamiji Yamamoto, Naoyuki Hiraga, Kuniaki Takeshita, and Toshiya Hashimoto. An estimation of net ecosystem metabolism and net denitrification of the Seto Inland Sea, Japan. Ecological Modelling, 215(1–3):55–68, July 2008. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000800094X.

Yule:2009:PNA

[YJJG09] Jeffrey V. Yule, Christopher X. J. Jensen, Aby Joseph, and Jimmie Goode. The puzzle of North America’s Late Pleistocene megafaunal extinction patterns: Test of new explanation yields unexpected results. Ecological Modelling, 220(4):533–544, February 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005358.

Yamazaki:2000:DAB

[YK00] Atsuko K. Yamazaki and Daniel Kamykowski. A dinoflagellate adaptive behavior model: response to internal biochemical cues. Ecological Modelling, 134(1):59–72, September 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380000003367.

Yurekli:2005:ALS

[YKO05] Kadri Yurekli, Ahmet Kurunc, and Fazli Ozturk. Application of linear stochastic models to monthly flow data of Kelkit Stream. Ecological Modelling, 183(1):67–75, April 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380004004314.

Yurova:2007:CSO

[YL07] Alla Yu. Yurova and Harry Lankreijer. Carbon storage in the organic layers of boreal forest soils under various moisture conditions: a model study for Northern Sweden sites. Ecological Modelling, 204(3–4):475–484, June 16, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000567.

Yue:2001:CHL

[YLJ^{+}01] Tianxiang Yue, Jiyuan Liu, Sven Erik Jørgensen, Zhiqiang Gao, Shihuang Zhang, and Xiangzheng Deng. Changes of holdridge life zone diversity in all of China over half a century. Ecological Modelling, 144(2–3):153–162, October 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001003702.

Yue:2003:LCD

[YLJY03] Tian Xiang Yue, Ji Yuan Liu, Sven Erik Jørgensen, and Qin Hua Ye. Landscape change detection of the newly created wetland in Yellow River Delta. Ecological Modelling, 164(1):21–31, June 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002003915.

Yue:2005:CED

[YLL^{+}05] Tian Xiang Yue, Ji Yuan Liu, Zheng Qing Li, Si Qing Chen, Sheng Nan Ma, Yong Zhong Tian, and Feng Ge. Considerable effects of diversity indices and spatial scales on conclusions relating to ecological diversity. Ecological Modelling, 188(2–4):418–431, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000906.

Yu:2002:SLP

[YLLW02] Qiang Yu, Yunfen Liu, Jiandong Liu, and Tianduo Wang. Simulation of leaf photosynthesis of winter wheat on Tibetan Plateau and in North China Plain. Ecological Modelling, 155(2–3):205–216, October 1, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001230.

Yang:2003:EWT

[YLSH03] Hui Yang, Yourun Li, Jingzhu Shen, and Shanying Hu. Evaluating waste treatment, recycle and reuse in industrial system: an application of the eMergy approach. Ecological Modelling, 160(1–2):13–21, February 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002880.

Yee:2002:VGA

[YM02] Thomas W. Yee and Monique Mackenzie. Vector generalized additive models in plant ecology. Ecological Modelling, 157(2–3):141–156, November 30, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002001928.

Yang:2005:MPR

[YM05] Zongjian Yang and David J. Midmore. Modelling plant resource allocation and growth partitioning in response to environmental heterogeneity. Ecological Modelling, 181(1):59–77, January 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380004003588.

Yniguez:2008:AMG

[YMD08] Aletta T. Yñiguez, John W. McManus, and Donald L. DeAngelis. Allowing macroalgae growth forms to emerge: Use of an agent-based model to understand the growth and spread of macroalgae in Florida coral reefs, with emphasis on Halimeda tuna. Ecological Modelling, 216(1):60–74, August 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000800207X.

Yamamura:2003:DRW

[YMT03] Kohji Yamamura, Seiichi Moriya, and Koichi Tanaka. Discrete random walk model to interpret the dispersal parameters of organisms. Ecological Modelling, 161(3):151–157, March 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002003459.

Yoshida:2003:EDS

[Yos03] Katsuhiko Yoshida. Evolutionary dynamics of species diversity in an interaction web system. Ecological Modelling, 163(1–2):131–143, May 1, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002004179.

Yoshida:2006:EIS

[Yos06a] Katsuhiko Yoshida. Effect of the intensity of stochastic disturbance on temporal diversity patterns — a simulation study in evolutionary time scale. Ecological Modelling, 196(1–2):103–115, July 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006228.

Yoshida:2006:ICP

[Yos06b] Katsuhiko Yoshida. Intra-clade predation facilitates the evolution of larger body size. Ecological Modelling, 196(3–4):533–539, July 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006000949.

Yoshida:2008:ECV

[Yos08] Katsuhiko Yoshida. Evolutionary cause of the vulnerability of insular communities. Ecological Modelling, 210 (4):403–413, February 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007004127.

Yemshanov:2002:SES

[YP02] Dennis Yemshanov and Ajith H. Perera. A spatially explicit stochastic model to simulate boreal forest cover transitions: general structure and properties. Ecological Modelling, 150(1–2):189–209, April 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000100480X.

Yin:2005:LLS

[YPR^{+}05] Zuo-Yun Yin, Shao-Lin Peng, Hai Ren, Qinfeng Guo, and Zhang-He Chen. LogCauchy, log-sech and lognormal distributions of species abundances in forest communities. Ecological Modelling, 184(2–4):329–340, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005587.

Yee:2008:CEI

[YSB08] Susan Harrell Yee, Deborah L. Santavy, and Mace G. Barron. Comparing environmental influences on coral bleaching across and within species using clustered binomial regression. Ecological Modelling, 218(1–2):162–174, October 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008003141.

Ye:2006:ECY

[YSG^{+}06] Xujun Ye, Kenshi Sakai, Leroy Ortega Garciano, Shin-Ichi Asada, and Akira Sasao. Estimation of citrus yield from airborne hyperspectral images using a neural network model. Ecological Modelling, 198(3–4):426–432, October 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006002547.

Yang:2000:EAB

[YSM00] M. D. Yang, R. M. Sykes, and C. J. Merry. Estimation of algal biological parameters using water quality modeling and SPOT satellite data. Ecological Modelling, 125(1):1–13, January 1, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380099000654.

Yang:2006:MNW

[YSM^{+}06] Xuefei Yang, Andrew K. Skidmore, David R. Melick, Zhekun Zhou, and Jianchu Xu. Mapping non-wood forest product (matsutake mushrooms) using logistic regression and a GIS expert system. Ecological Modelling, 198(1–2):208–218, September 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001839.

Yu:2001:MLTa

[YTCV01] Zicheng Yu, M. R. Turetsky, I. D. Campbell, and D. H. Vitt. Modelling long-term peatland dynamics. II. Processes and rates as inferred from litter and peat-core data. Ecological Modelling, 145(2–3):159–173, November 15, 2001. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001003878.

Yang:2003:MIT

[YTH03] Yuqing Yang, Stephen J. Titus, and Shongming Huang. Modeling individual tree mortality for white spruce in Alberta. Ecological Modelling, 163(3):209–222, May 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000085.

Yue:2008:SMH

[YTLF08] Tian-Xiang Yue, Yong-Zhong Tian, Ji-Yuan Liu, and Ze-Meng Fan. Surface modeling of human carrying capacity of terrestrial ecosystems in China. Ecological Modelling, 214(2–4):168–180, June 24, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008000586.

Yamanaka:2003:IBM

[YTS03] Takehiko Yamanaka, Sadahiro Tatsuki, and Masakazu Shimada. An individual-based model for sex-pheromone-oriented flight patterns of male moths in a local area. Ecological Modelling, 161(1–2):35–51, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380002002910.

Yoshimura:2009:MEG

[YTT^{+}09] Jin Yoshimura, Yumi Tanaka, Tatsuya Togashi, Shigehide Iwata, and Kei ichi Tainaka. Mathematical equivalence of geometric mean fitness with probabilistic optimization under environmental uncertainty. Ecological Modelling, 220(20):2611–2617, October 24, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009004517.

Yunusi:2008:ISN

[Yun08] Mahmadyusuf K. Yunusi. Investigation of some nonlinear singular model ecosystems and new concerned mathematical problems. Ecological Modelling, 216(2):172–177, August 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008001373.

Yang:2002:OCM

[YWHW02] Xin Yang, Mingxing Wang, Yao Huang, and Yuesi Wang. A one-compartment model to study soil carbon decomposition rate at equilibrium situation. Ecological Modelling, 151(1):63–73, May 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380001004859.

Yue:2005:SMH

[YWL^{+}05] Tian Xiang Yue, Ying An Wang, Ji Yuan Liu, Shu Peng Chen, Dong Sheng Qiu, Xiang Zheng Deng, Ming Liang Liu, Yong Zhong Tian, and Bian Ping Su. Surface modelling of human population distribution in China. Ecological Modelling, 181(4):461–478, February 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004003722.

Yi:2006:SCI

[YY06] Xianfeng Yi and Yueqin Yang. A stable carbon isotopic approach for understanding the CO_{2} flux at the haibei alpine meadow ecosystem — a simple model. Ecological Modelling, 193(3–4):796–800, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003613.

Yoshie:2007:PSS

[YYR^{+}07] Naoki Yoshie, Yasuhiro Yamanaka, Kenneth A. Rose, David L. Eslinger, Daniel M. Ware, and Michio J. Kishi. Parameter sensitivity study of the NEMURO lower trophic level marine ecosystem model. Ecological Modelling, 202(1–2):26–37, March 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004546.

Yi:2006:MTP

[YYZ06] Xianfeng Yi, Yueqin Yang, and Xiaoai Zhang. Modeling trophic positions of the alpine meadow ecosystem combining stable carbon and nitrogen isotope ratios. Ecological Modelling, 193(3–4):801–808, March 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003315.

Ying:2007:CAG

[YZC^{+}07] Xiong Ying, Guang-Ming Zeng, Gui-Qiu Chen, Lin Tang, Ke-Lin Wang, and Dao-You Huang. Combining AHP with GIS in synthetic evaluation of eco-environment quality — a case study of Hunan Province, China. Ecological Modelling, 209(2–4):97–109, December 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007003183.

Yonow:2004:MPD

[YZS^{+}04] T. Yonow, M. P. Zalucki, R. W. Sutherst, B. C. Dominiak, G. F. Maywald, D. A. Maelzer, and D. J. Kriticos. Modelling the population dynamics of the Queensland fruit fly, Bactrocera (Dacus) tryoni: a cohort-based approach incorporating the effects of weather. Ecological Modelling, 173(1):9–30, March 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003003065.

Zhang:2009:BHF

[ZA09] Weitao Zhang and George B. Arhonditsis. A Bayesian hierarchical framework for calibrating aquatic biogeochemical models. Ecological Modelling, 220(18):2142–2161, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009003925.

Zhu:2003:FNCa

[ZAM^{+}03a] Zhanxue Zhu, Paul A. Arp, Fanrui Meng, Charles P.-A. Bourque, and Neil W. Foster. A forest nutrient cycling and biomass model (ForNBM) based on year-round, monthly weather conditions, Part I: assumption, structure and processing. Ecological Modelling, 169(2–3):347–360, November 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002813.

Zhu:2003:FNCb

[ZAM^{+}03b] Zhanxue Zhu, Paul A. Arp, Fanrui Meng, Charles P.-A. Bourque, and Neil W. Foster. A forest nutrient cycling and biomass model (ForNBM) based on year-round, monthly weather conditions: Part II: Calibration, verification, and application. Ecological Modelling, 170(1):13–27, December 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003002849.

Zhu:2005:MSW

[ZAM^{+}05] Zhanxue Zhu, Paul A. Arp, Asit Mazumder, Fanrui Meng, Charles P.-A. Bourque, and Neil W. Foster. Modeling stream water nutrient concentrations and loadings in response to weather condition and forest harvesting. Ecological Modelling, 185(2–4):231–243, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004006179.

Zhu:2007:MRS

[ZAM^{+}07] Zhanxue Zhu, Paul A. Arp, Fanrui Meng, Charles P.-A. Bourque, and Asit Mazumder. Modeling response of soil productivity to biogeochemical cycling and atmospheric acid deposition in the Hayden Brook watershed (Canada) using the ForNBM model. Ecological Modelling, 205(3–4):410–422, July 24, 2007. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380007001196.

Zavala:2004:IDT

[Zav04] Miguel A. Zavala. Integration of drought tolerance mechanisms in Mediterranean sclerophylls: a functional interpretation of leaf gas exchange simulators. Ecological Modelling, 176(3–4):211–226, September 1, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004000468.

Zavalishin:2008:DCA

[Zav08] Nikolay N. Zavalishin. Dynamic compartment approach for modelling regimes of carbon cycle functioning in bog ecosystems. Ecological Modelling, 213(1):16–32, April 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380007006370.

Zhou:2004:NNI

[ZB04] Mo Zhou and Joseph Buongiorno. Nonlinearity and noise interaction in a model of forest growth. Ecological Modelling, 180(2–3):291–304, December 25, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004002832.

Zaldivar:2009:MAR

[ZBD^{+}09] J. M. Zaldívar, F. S. Bacelar, S. Dueri, D. Marinov, P. Viaroli, and E. Hernández-García. Modeling approach to regime shifts of primary production in shallow coastal ecosystems. Ecological Modelling, 220 (21):3100–3110, November 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009000738.

Zuo:2003:STM

[ZBL03] Heping Zuo, Adrian L. Bugg, and Greg Laughlin. Spatial and temporal modelling of pasture and plantation production for integrated catchment planning in south-eastern Australia. Ecological Modelling, 166(1–2):75–86, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003001030.

Zhang:2007:NNM

[ZBL07] WenJun Zhang, ChangJun Bai, and GuoDao Liu. Neural network modeling of ecosystems: a case study on cabbage growth system. Ecological Modelling, 201 (3–4):317–325, March 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380006004480.

Zhou:2007:DAG

[ZBSA07] Xinhua Zhou, James R. Brandle, Michele M. Schoeneberger, and Tala Awada. Developing above-ground woody biomass equations for open-grown, multiple-stemmed tree species: Shelterbelt-grown Russian-olive. Ecological Modelling, 202(3–4):311–323, April 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000600531X.

Zhao:2005:DDM

[ZBW05] Dehai Zhao, Bruce Borders, and Machelle Wilson. A density-dependent matrix model for bottomland hardwood stands in the Lower Mississippi Alluvial Valley. Ecological Modelling, 184(2–4):381–395, June 1, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005770.

Zhao:2006:MNE

[ZBWR06] Dehai Zhao, Bruce Borders, Machelle Wilson, and Stephen L. Rathbun. Modeling neighborhood effects on the growth and survival of individual trees in a natural temperate speciesrich forest. Ecological Modelling, 196(1–2):90–102, July 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380006000585.

Zhang:2009:SDM

[ZBZ^{+}09] Guilian Zhang, Carolina Boom, Guigang Zhang, Xinwei Liu, Qin Du, and Shaolin Peng. Simulating dynamics of managed monsoon evergreen broad-leaved forest in south China. Ecological Modelling, 220(18):2218–2230, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009003329.

Zheng:2000:EEF

[ZC00] Daolan Zheng and Jiquan Chen. Edge effects in fragmented landscapes: a generic model for delineating area of edge influences (D-AEI). Ecological Modelling, 132(3):175–190, August 5, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000002544.

Zhang:2007:MEE

[ZC07] Yuying Zhang and Yong Chen. Modeling and evaluating ecosystem in 1980s and 1990s for American lobster (Homarus americanus) in the Gulf of Maine. Ecological Modelling, 203(3–4):475–489, May 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006006363.

Zhang:2008:TDE

[ZCB08] Hongyan Zhang, David A. Culver, and Leon Boegman. A two-dimensional ecological model of Lake Erie: Application to estimate dreissenid impacts on large lake plankton populations. Ecological Modelling, 214(2–4):219–241, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000781.

Zhang:2008:TMT

[ZCG^{+}08] Jun Zhang, Ziyan Chu, Ying Ge, Xiaolu Zhou, Hong Jiang, Jie Chang, Changhui Peng, Jiawen Zheng, Bo Jiang, Jinru Zhu, and Shuquan Yu. TRIPLEX model testing and application for predicting forest growth and biomass production in the subtropical forest zone of China’s Zhejiang Province. Ecological Modelling, 219(3–4):264–275, December 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008003529.

Zubillaga:2007:MFS

[ZCKR07] M. Mercedes Zubillaga, Miguel L. Cabrera, David E. Kissel, and John A. Rema. Modeling field-scale N mineralization in coastal plain soils (USA). Ecological Modelling, 207(2–4):243–250, October 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007002761.

Zhu:2009:OEM

[ZCQ^{+}09] Lin Zhu, Jing M. Chen, Qiming Qin, Jianping Li, and Lianxi Wang. Optimization of ecosystem model parameters using spatio-temporal soil moisture information. Ecological Modelling, 220(18):2121–2136, September 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009003196.

Zhao:2009:NWF

[ZCY09] X. Zhao, B. Chen, and Z. F. Yang. National water footprint in an input–output framework — a case study of China 2002. Ecological Modelling, 220(2):245–253, January 24, 2009. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008004584.

Zheng:2004:DWQ

[ZCZ04] Lianyuan Zheng, Changsheng Chen, and Frank Y. Zhang. Development of water quality model in the Satilla River Estuary, Georgia. Ecological Modelling, 178(3–4):457–482, November 1, 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380004002042.

Zavala:2005:MMT

[ZdlP05] M. A. Zavala and R. Bravo de la Parra. A mechanistic model of tree competition and facilitation for Mediterranean forests: Scaling from leaf physiology to stand dynamics. Ecological Modelling, 188(1):76–92, October 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005002437.

Zemlys:2003:RPI

[ZDR03] Petras Zemlys, Darius Daunys, and Arturas Razinkovas. Revision pre-ingestive selection efficiency definition for suspension feeding bivalves: facilitating the material fluxes modelling. Ecological Modelling, 166(1–2):67–74, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000991.

Zlatev:2008:LSC

[ZEGS08] Zahari Zlatev, Adolf Ebel, Krassimir Georgiev, and Roberto San José. Large scale computations in environmental modelling: Editorial introduction. Ecological Modelling, 217(3–4):207–208, October 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002573.

Zeide:2004:IUG

[Zei04] Boris Zeide. Intrinsic units in growth modeling. Ecological Modelling, 175(3):249–259, July 15, 2004. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003004794.

Zhang:2005:MCE

[ZGF^{+}05] Y. Zhang, R. F. Grant, L. B. Flanagan, S. Wang, and D. L. Verseghy. Modelling CO_{2} and energy exchanges in a northern semiarid grassland using the carbon- and nitrogencoupled Canadian Land Surface Scheme (C-CLASS). Ecological Modelling, 181(4):591–614, February 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380004004028.

Zhang:2005:SRA

[ZGH05] Lianjun Zhang, Jeffrey H. Gove, and Linda S. Heath. Spatial residual analysis of six modeling techniques. Ecological Modelling, 186(2):154–177, August 15, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000153.

Zhou:2008:IEA

[ZGL08] Demin Zhou, Huili Gong, and Zhaoli Liu. Integrated ecological assessment of biophysical wetland habitat in water catchments: Linking hydro-ecological modelling with geo-information techniques. Ecological Modelling, 214(2–4):411–420, June 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008001555.

Gong:2009:UES

[zGsLcXwZ09] Jian zhou Gong, Yan sui Liu, Bei cheng Xia, and Guan wei Zhao. Urban ecological security assessment and forecasting, based on a cellular automata model: a case study of Guangzhou, China. Ecological Modelling, 220 (24):3612–3620, December 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009007042.

Zang:2006:AMR

[ZH06] Shuying Zang and Xi Huang. An aggregated multivariate regression land-use model and its application to land-use change processes in the Daqing region (northeast China). Ecological Modelling, 193(3–4):503–516, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004102.

Zhang:2000:MMW

[Zha00] Zhibin Zhang. Mathematical models of wildlife management by contraception. Ecological Modelling, 132(1–2):105–113, July 30, 2000. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380000003082.

Zhang:2003:CTD

[Zha03a] Xiaoxian Zhang. Comment on “A three-dimensional spatial model for plant competition in an heterogeneous soil environment” by Biondini. Ecological Modelling, 166 (1–2):185–186, August 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380003001157.

Zhang:2003:MCA

[Zha03b] Zhibin Zhang. Mutualism or cooperation among competitors promotes coexistence and competitive ability. Ecological Modelling, 164(2–3):271–282, June 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000693.

Zhang:2006:BRE

[Zha06] J. Zhang. Book review: Eco-exergy as Sustainability, S. E. Jørgensen. WIT Press (2006), 208 pp., Price: £73 (Hardbound), ISBN: 1-84564-059-4. Ecological Modelling, 199(1):127, November 1, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003607.

Zhang:2007:OPS

[Zha07] Zhibin Zhang. The outbreak pattern of SARS cases in China as revealed by a mathematical model. Ecological Modelling, 204(3–4):420–426, June 16, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000700049X.

Zhang:2007:OEU

[ZHL07] Yanyu Zhang, Xiaozhuo Han, and Zizhen Li. Overcrowding effect, unstable or stable. Ecological Modelling, 200(1–2):269–270, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004121.

Zhou:2002:SES

[ZHMN02] X. Zhou, C. W. Hoy, S. A. Miller, and L. R. Nault. Spatially explicit simulation of aster yellows epidemics and control on lettuce. Ecological Modelling, 151(2–3):293–307, June 1, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380001004926.

Zhu:2005:COA

[ZHW05] Mu Zhu, Trevor J. Hastie, and Guenther Walther. Constrained ordination analysis with flexible response functions. Ecological Modelling, 187(4):524–536, October 10, 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005000876.

Zhang:2003:HVS

[ZJBI03] Jingjie Zhang, Sven Erik Jørgensen, Meryem Beklioglu, and Ozlem Ince. Hysteresis in vegetation shift — Lake Mogan prognoses. Ecological Modelling, 164(2–3):227–238, June 15, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380003000504.

Zhang:2007:GRW

[ZJC^{+}07] Xiaoxian Zhang, Scott N. Johnson, John W. Crawford, Peter J. Gregory, and Iain M. Young. A general random walk model for the leptokurtic distribution of organism movement: Theory and application. Ecological Modelling, 200(1–2):79–88, January 10, 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006003437.

Znidarsic:2006:ACC

[ŽJDK06] Martin Žnidaršič, Aleks Jakulin, Sašo Džeroski, and Christian Kampichler. Automatic construction of concept hierarchies: the case of foliage-dwelling spiders. Ecological Modelling, 191(1):144–158, January 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005003820.

Zhang:2006:MMS

[ZJG^{+}06] Xiaoxian Zhang, Scott N. Johnson, Peter J. Gregory, John W. Crawford, Iain M. Young, Philip J. Murray, and Steve C. Jarvis. Modelling the movement and survival of the root-feeding clover weevil, Sitona lepidus, in the root-zone of white clover. Ecological Modelling, 190(1–2):133–146, January 10, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380005002310.

Zhang:2004:ESD

[ZJM04] Jingjie Zhang, Sven Erik Jørgensen, and Henrik Mahler. Examination of structurally dynamic eutrophication model. Ecological Modelling, 173(4):313–333, April 15, 2004. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380003004137.

Zhang:2006:SDC

[ZJN^{+}06] Guilian Zhang, Hong Jiang, Guodong Niu, Xinwei Liu, and Shaolin Peng. Simulating the dynamics of carbon and nitrogen in litter-removed pine forest. Ecological Modelling, 195(3–4):363–376, June 15, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005006253.

Zhang:2003:SDM

[ZJTB03] Jingjie Zhang, Sven Erik Jørgensen, Can Ozan Tan, and Meryem Beklioglu. A structurally dynamic modelling — Lake Mogan, Turkey as a case study. Ecological Modelling, 164(2–3):103–120, June 15, 2003. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003000516.

Zhang:2008:ITM

[ZK08] Xin Zhang and Russell G. Kreis. Importance of temperature in modeling food web bioaccumulation in large aquatic systems. Ecological Modelling, 218(3–4):315–322, November 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008003621.

Zell:2009:PCD

[ZKH09] Jürgen Zell, Gerald Kändler, and Marc Hanewinkel. Predicting constant decay rates of coarse woody debris — a meta-analysis approach with a mixed model. Ecological Modelling, 220(7):904–912, April 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009000672.

Li:2005:IPP

[zLGH^{+}05] Zi zhen Li, Meng Gao, Cang Hui, Xiao zhuo Han, and Honghua Shi. Impact of predator pursuit and prey evasion on synchrony and spatial patterns in metapopulation. Ecological Modelling, 185(2–4):245–254, July 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004006192.

Zhang:2006:SDD

[ZLH06] Feng Zhang, Zizhen Li, and Cang Hui. Spatiotemporal dynamics and distribution patterns of cyclic competition in metapopulation. Ecological Modelling, 193 (3–4):721–735, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005004795.

Zhao:2005:MME

[ZLL05] Sheng Zhao, Zizhen Li, and Wenlong Li. A modified method of ecological footprint calculation and its application. Ecological Modelling, 185(1):65–75, June 10, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005964.

Zaniewski:2002:PSS

[ZLO02] A. Elizabeth Zaniewski, Anthony Lehmann, and Jacob McC. Overton. Predicting species spatial distributions using presence-only data: a case study of native New Zealand ferns. Ecological Modelling, 157(2–3):261–280, November 30, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002001990.

Zou:2009:MPP

[ZLW09] Rui Zou, Wu-Seng Lung, and Jing Wu. Multiplepattern parameter identification and uncertainty analysis approach for water quality modeling. Ecological Modelling, 220(5):621–629, March 10, 2009. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008005723.

Zhang:2009:EMV

[ZLX09] Yong Zhang, Laifu Liu, and Rumei Xu. The effect of migration on the viability, dynamics and structure of two coexisting metapopulations. Ecological Modelling, 220 (3):272–282, February 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380008004778.

Zhang:2002:SML

[ZLZM02] Yu Zhang, Changsheng Li, Xiuji Zhou, and Berrien Moore. A simulation model linking crop growth and soil biogeochemistry for sustainable agriculture. Ecological Modelling, 151(1):75–108, May 15, 2002. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005270.

Zlatev:2008:ICC

[ZM08] Zahari Zlatev and Lars Moseholm. Impact of climate changes on pollution levels in Denmark. Ecological Modelling, 217(3–4):305–319, October 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008002639.

Zhang:2008:MFL

[ZMB^{+}08] C. F. Zhang, F.-R. Meng, J. S. Bhatti, J. A. Trofymow, and Paul A. Arp. Modeling forest leaf-litter decomposition and N mineralization in litterbags, placed across Canada: a 5-model comparison. Ecological Modelling, 219(3–4):342–360, December 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008003517.

Zhao:2006:GAM

[ZNC^{+}06] Chuanyan Zhao, Zhongren Nan, Guodong Cheng, Junhua Zhang, and Zhaodong Feng. GIS-assisted modelling of the spatial distribution of Qinghai spruce (Picea crassifolia) in the qilian mountains, northwestern China based on biophysical parameters. Ecological Modelling, 191 (3–4):487–500, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002863.

Zhou:2006:FPA

[ZPD06] Xiaolu Zhou, Changhui Peng, and Qing-Lai Dang. Formulating and parameterizing the allocation of net primary productivity for modeling overmature stands in boreal forest ecosystems. Ecological Modelling, 195(3–4):264–272, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005006058.

Zhou:2008:SCE

[ZPD^{+}08] Xiaolu Zhou, Changhui Peng, Qing-Lai Dang, Jianfeng Sun, Haibin Wu, and Dong Hua. Simulating carbon exchange in Canadian boreal forests: I. Model structure, validation, and sensitivity analysis. Ecological Modelling, 219(3–4):287–299, December 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000800361X.

Zadereev:2003:IBM

[ZPGG03] Egor S. Zadereev, Igor G. Prokopkin, Vladimir G. Gubanov, and Michail V. Gubanov. Individual-based model of the reproduction cycle of Moina macrocopa (Crustacea: Cladocera). Ecological Modelling, 162(1–2):15–31, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003484.

Zhao:2007:SVS

[ZPK^{+}07] Ying Zhao, Stephan Peth, Julia Krümmelbein, Rainer Horn, Zhongyan Wang, Markus Steffens, Carsten Hoffmann, and Xinhua Peng. Spatial variability of soil properties affected by grazing intensity in inner Mongolia grassland. Ecological Modelling, 205(1–2):241–254, July 10, 2007. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007000956.

Zheng:2009:MDD

[ZPO09] Chaozhi Zheng, Juho Pennanen, and Otso Ovaskainen. Modelling dispersal with diffusion and habitat selection: Analytical results for highly fragmented landscapes. Ecological Modelling, 220(12):1495–1505, June 24, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380009001203.

Zador:2006:BPE

[ZPP06] Stephani G. Zador, John F. Piatt, and André E. Punt. Balancing predation and egg harvest in a colonial seabird: a simulation model. Ecological Modelling, 195(3–4):318–326, June 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000500565X.

Zhou:2005:PSP

[ZPW05] X. Zhou, N. Persaud, and H. Wang. Periodicities and scaling parameters of daily rainfall over semi-arid Botswana. Ecological Modelling, 182(3–4):371–378, March 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000400170X.

Zeller:2004:MSC

[ZR04] Dirk Zeller and Jákup Reinert. Modelling spatial closures and fishing effort restrictions in the Faroe Islands marine ecosystem. Ecological Modelling, 172(2–4):403–420, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000300382X.

Zetina-Rejon:2004:EHS

[ZRASC04] Manuel J. Zetina-Rejón, Francisco Arreguín-Sánchez, and Ernesto A. Chávez. Exploration of harvesting strategies for the management of a Mexican coastal lagoon fishery. Ecological Modelling, 172(2–4):361–372, March 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S030438000300379X.

Zhao:2008:PCP

[ZRCA08] Jingyang Zhao, Maryam Ramin, Vincent Cheng, and George B. Arhonditsis. Plankton community patterns across a trophic gradient: the role of zooplankton functional groups. Ecological Modelling, 213(3–4):417–436, May 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380008000379.

Zhang:2008:MCC

[ZRR^{+}08] Xiaomi Zhang, Kenneth R. Rygwelski, Ronald Rossmann, James J. Pauer, and Russell G. Kreis. Model construct and calibration of an integrated water quality model (LM2-toxic) for the Lake Michigan Mass Balance Project. Ecological Modelling, 219(1–2):92–106, November 24, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008004031.

Zhang:2009:INE

[ZS09] Wei Zhang and Scott M. Swinton. Incorporating natural enemies in an economic threshold for dynamically optimal pest management. Ecological Modelling, 220(9–10):1315–1324, May 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009000830.

Zaldivar:2008:CRS

[ZSD^{+}08] José-Manuel Zaldívar, Fernanda Strozzi, Sibylle Dueri, Dimitar Marinov, and Joseph P. Zbilut. Characterization of regime shifts in environmental time series with recurrence quantification analysis. Ecological Modelling, 210(1–2):58–70, January 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380007003638.

Zammit:2005:MEL

[ZSKV05] C. Zammit, M. Sivapalan, P. Kelsey, and N. R. Viney. Modelling the effects of land-use modifications to control nutrient loads from an agricultural catchment in Western Australia. Ecological Modelling, 187(1):60–70, September 2005. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005000384.

Zharova:2008:AAF

[ZSPV08] Nadezhda Zharova, Adriano Sfriso, Bruno Pavoni, and Alexey Voinov. Analysis of annual fluctuations of C. nodosa in the Venice lagoon: Modeling approach. Ecological Modelling, 216(2):134–144, August 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380008001324.

Zaldivar:2006:P

[ZSTM06] J. M. Zaldívar, C. Solidoro, G. Tsirtsis, and D. Mouillot. Preface. Ecological Modelling, 193(1–2):1–3, March 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004527.

Zhou:2006:MVC

[ZSZ06] Z. C. Zhou, Z. P. Shangguan, and D. Zhao. Modeling vegetation coverage and soil erosion in the Loess Plateau Area of China. Ecological Modelling, 198(1–2):263–268, September 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380006002067.

Zhuravlev:2008:MEE

[ZTF08] Yu. N. Zhuravlev, A. V. Tuzinkevich, and E. Ya. Frisman. Modelling the early events of primordial life. Ecological Modelling, 212(3–4):536–544, April 10, 2008. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380007006035.

Zhou:2003:PMM

[ZTXD03] Ziyong Zhou, Shu Tao, Fuliu Xu, and R. Dawson. A physical–mathematical model for the transport of heavy metals and toxic matter from point sources by geogas microbubbles. Ecological Modelling, 161(1–2):139–149, March 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380002003228.

Zucchetto:2004:RHO

[Zuc04] James J. Zucchetto. Reflections on Howard T. Odum’s paper: Energy, Ecology and Economics, Ambio, 1973. Ecological Modelling, 178(1–2):195–198, October 2004. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380003005532.

Zuenko:2007:ALT

[Zue07] Yury I. Zuenko. Application of a lower trophic level model to a coastal sea ecosystem. Ecological Modelling, 202(1–2):132–143, March 2007. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006004625.

Zelenev:2005:MWL

[ZvBS05] V. V. Zelenev, A. H. C. van Bruggen, and A. M. Semenov. Modeling wave-like dynamics of oligotrophic and copiotrophic bacteria along wheat roots in response to nutrient input from a growing root tip. Ecological Modelling, 188(2–4):404–417, November 10, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000864.

Zhang:2005:MPN

[ZVK05] Baisen Zhang, Ian Valentine, and Peter Kemp. Modelling the productivity of naturalised pasture in the North Island, New Zealand: a decision tree approach. Ecological Modelling, 186(3):299–311, August 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000542.

Zabadal:2005:SUP

[ZVLP05] J. R. S. Zabadal, M. T. M. B. Vilhena, S. Q. Bogado Leite, and C. A. Poffal. Solving unsteady problems in water pollution using Lie symmetries. Ecological Modelling, 186(3):271–279, August 25, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380004005563.

Zhang:2002:STM

[ZW02] Huayong Zhang and Jianguo Wu. A statistical thermodynamic model of the organizational order of vegetation. Ecological Modelling, 153(1–2):69–80, July 15, 2002. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380001005026.

Zhou:2006:OLS

[ZW06] Shu-Rong Zhou and Gang Wang. One large, several medium, or many small? Ecological Modelling, 191 (3–4):513–520, February 5, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380005002875.

Zhao:2005:SEG

[ZWCL05] Chengyi Zhao, Yuchao Wang, Xi Chen, and Baoguo Li. Simulation of the effects of groundwater level on vegetation change by combining FEFLOW software. Ecological Modelling, 187(2–3):341–351, September 25, 2005. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000803.

Zhao:2005:ADA

[ZWXF05] Liang Zhao, Hao Wei, Yongfu Xu, and Shizuo Feng. An adjoint data assimilation approach for estimating parameters in a three-dimensional ecosystem model. Ecological Modelling, 186(2):235–250, August 15, 2005. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380005000293.

Zhan:2003:AAE

[ZXC03] Xiwu Zhan, Yongkang Xue, and G. James Collatz. An analytical approach for estimating CO_{2} and heat fluxes over the Amazonian region. Ecological Modelling, 162(1–2):97–117, April 1, 2003. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380002004052.

Zhang:2006:AUE

[ZYL06] Yan Zhang, Zhifeng Yang, and Wei Li. Analyses of urban ecosystem based on information entropy. Ecological Modelling, 197(1–2):1–12, August 10, 2006. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304380006001013.

Zhang:2006:MEI

[ZYY06] Yan Zhang, Zhifeng Yang, and Xiangyi Yu. Measurement and evaluation of interactions in complex urban ecosystem. Ecological Modelling, 196(1–2):77–89, July 10, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380006000573.

Zhang:2009:ENE

[ZYY09a] Yan Zhang, Zhifeng Yang, and Xiangyi Yu. Ecological network and emergy analysis of urban metabolic systems: Model development, and a case study of four Chinese cities. Ecological Modelling, 220(11):1431–1442, June 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009000842.

Zhang:2009:EUM

[ZYY09b] Yan Zhang, Zhifeng Yang, and Xiangyi Yu. Evaluation of urban metabolism based on emergy synthesis: a case study for Beijing (China). Ecological Modelling, 220 (13–14):1690–1696, July 17, 2009. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009002294.

Zhang:2008:NNM

[ZZ08] WenJun Zhang and XiYan Zhang. Neural network modeling of survival dynamics of holometabolous insects: a case study. Ecological Modelling, 211(3–4):433–443, March 10, 2008. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438000700498X.

Zhang:2009:EDS

[ZZL^{+}09] Hui Zhang, Feng Zhang, Zizhen Li, Meng Gao, and Wenlong Li. Evolutionary diversity and spatiotemporal dynamics of a spatial game. Ecological Modelling, 220(19):2353–2364, October 10, 2009. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009004037.

Zhang:2006:VAT

[ZZLC06] Shuqing Zhang, Junyan Zhang, Fang Li, and Roger Cropp. Vector analysis theory on landscape pattern (VATLP). Ecological Modelling, 193(3–4):492–502, March 15, 2006. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380005004035.

Zabkar:2006:QPO

[ŽŽV^{+}06] Jure Žabkar, Rahela Žabkar, Daniel Vladušič, Danijel Čemas, Dorian Šuc, and Ivan Bratko. Q2 prediction of ozone concentrations. Ecological Modelling, 191(1):68–82, January 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005003509.

Zhang:2006:SAT

[ZZZ06] Chunrui Zhang, Yuangang Zu, and Baodong Zheng. Stability analysis in a two-dimensional life energy system model with delay. Ecological Modelling, 193(3–4):691–702, March 15, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380005004813.